A Complete Bibliography of Publications in the Journal of Cell Biology: 2015–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

02 May 2023 Version 1.07

Title word cross-reference

1 [ARV^{+}18, FFG^{+}18, SPE^{+}17a, VMR^{+}19]. 2 [MF16b]. 3 [CZW^{+}18, GTMZ^{+}15, KBT^{+}19, LJS^{+}16a, LJS^{+}16b, MF16b, NKH^{+}19, NPC17, PHKY17, PSL^{+}17, SPJ^{+}15, TYD^{+}15, VZB19]. 4 [HHM15, JJB^{+}19, KT15a, KT15b, NMN^{+}15]. 5 [DZB^{+}18, GWZ^{+}19a]. 6 [Les15o]. ^{+} [BKH^{+}15, CWL^{+}17, CNC^{+}18, LRD19, MPMP16, OBY^{+}15, WZG^{+}17, WRGB^{+}15]. ^{2+} [CCQ^{+}18, CBM^{+}16, LE16, MPMP16, MWSM18, MWSM19, MPW^{+}19, RGOS^{+}16, SZL^{+}16, Sør17, SBP^{+}16, WZG^{+}17, WWT18]. ^{Cdc55} [JRH^{+}16]. Dpb11 [LCD^{+}17]. ^{Fizzy} [DLM^{+}15]. ^{G12D} [XWZ^{+}15]. ^{PCH−2} [NHCB15]. ^{Rad9} [LCD^{+}17]. ^{ZYG11} [BHS^{+}16, Bra16]. _{1} [CNC^{+}18]. _{2} [CCLL17, CWZ^{+}15, Dic17, GPD^{+}19, HQW15, JJW17, KML^{+}15, LLW^{+}15, MBC^{+}19, DR19, SKZ^{+}18a, Yud19]. _{2A} [XJG^{+}17]. _{3} [RSC^{+}19]. α [BNS^{+}17, BKG^{+}15, BAGM17, CST^{+}16, CHI^{+}15, CKS^{+}15, DR16, FBBRCA^{+}18, FRP^{+}17, GSP^{+}18, HDA^{+}17, JCK^{+}19, KT15a, KT15b, LLC^{+}17, LBV^{+}17, MB17a, MSS^{+}17, NNK^{+}15, PAC^{+}15, Qi17, RKK^{+}18, STR^{+}18, SSV^{+}18, SFZ^{+}17, TTU^{+}17, WWZ^{+}18, WIS^{+}17, YTGA16, ZT15]. β [AGB^{+}19, ACG^{+}17, ARV^{+}18, BAGM17, CHS^{+}17, CSG^{+}15, CIK^{+}17, DVS^{+}17, DKA^{+}16, FWL^{+}17, FVF^{+}16, GBD^{+}18, HAK^{+}15, JLB^{+}18, JJB^{+}19, Les15u, LJ17b, LDR^{+}19, LSS^{+}15, LLC^{+}17, MA17, MSS^{+}17, PhHS^{+}16, PTK16, PLH18, PW19, PAC^{+}15, RRM^{+}17, Sho15s, SPJ^{+}15, SHVO^{+}18, SLG^{+}18, TSK^{+}18, TSK^{+}19, VXF^{+}15, WWZ^{+}18, WEQ^{+}15, WGHE^{+}18, XWZ^{+}15, XMJ^{+}19, XTT^{+}18, YYZ^{+}15, ZQZ19, ZT15]. β_{1−42} [QYC^{+}17]. δ [DVS^{+}17, GSD^{+}15, RKK^{+}18, WDM^{+}15]. γ [CKS^{+}15, IZZ^{+}18, cLNF^{+}16, LDR^{+}19, LBV^{+}17, MSS^{+}17, MSL16, SVD^{+}15, Sho15-68, SKZ^{+}18b]. κ [Hu15, LAMACE^{+}17, MCS^{+}15, YGMR^{+}17, ZLG^{+}15, dVGO^{+}16]. μ [NEW^{+}17].

-1 [RKK^{+}18]. -actin [MSS^{+}17, SVD^{+}15, Sho15-68, SPJ^{+}15]. -actinin- [KT15a, KT15b]. -appendage [FRP^{+}17]. -arrestin [HDA^{+}17, PhHS^{+}16]. -barrel [JLB^{+}18, WEQ^{+}15]. -binding [RSC^{+}19]. -Catenin [CHI^{+}15, WGHE^{+}18, WIS^{+}17, GBD^{+}18, LJ17b, RRM^{+}17, MB17a]. -dependent [KT15a, KT15b, CBM^{+}16]. -directed [GSD^{+}15]. -Glycan-dependent [LGH^{+}18]. -induced [VXF^{+}15, XWZ^{+}15]. -Integrins [JCK^{+}19]. -kinase [GWZ^{+}19a, JJB^{+}19]. -like [DVS^{+}17, ZQZ19]. -phosphatase [DZB^{+}18, NMN^{+}15]. -phosphate [HHM15, Les15o, LJS^{+}16a, LJS^{+}16b]. -PIX [LSS^{+}15]. -QC [MPA^{+}16]. -Secretase [IZZ^{+}18, LDR^{+}19]. -secretases [CKS^{+}15]. -stimulated [MWSM18, MWSM19]. -synuclein [CST^{+}16, DR16]. -TAT1 [FBBRCA^{+}18]. -tubulin [BNS^{+}17, cLNF^{+}16, MSL16, SKZ^{+}18b]. -TuSC [cLNF^{+}16].

/H [MPMP16].

1 [AZS^{+}15, BKH^{+}15, CWL^{+}17, CRPSC^{+}19, CPBG19, CST^{+}16, CYL^{+}18, CLBB15, CBB15, DDAR^{+}16, DR16, DKA^{+}16, DLBMA^{+}15, FSB^{+}15, GCJ^{+}15, GWZ^{+}19b, HBS^{+}15, HAR^{+}15, JH19, JNW15, KCB^{+}16, LAMACE^{+}17, MSK^{+}18, MFP17, MDC^{+}16, NEW^{+}17, NNK^{+}15, OOT^{+}18, QCC^{+}19, SCNTC^{+}18, SRF19, Sed15x, Sho15-48, SENL^{+}15, SLH17, TNP^{+}15, TCD^{+}15, XMJ^{+}19, YHG^{+}17, YKKB17]. 1/APPL1 [LRM^{+}19]. 1/Bora [TNP^{+}15]. 1/BUB [KMLG^{+}15, KMLG^{+}16]. 1/p38 [DKA^{+}16]. 10 [LWH^{+}18]. 103/ [PPK^{+}16]. 11 [CNN^{+}17, Les17]. 14-3-3 [BBS^{+}17, Das17, iNLM^{+}19]. 16 [WZR19]. 17 [Juh16, MLMF16]. 170 [JNW15]. 1A [BRY^{+}19, KCB^{+}16]. 1b [IYP^{+}18, PLD^{+}15, HBDW^{+}15]. 1E [THA^{+}16].

2 [BBSA^{+}16, BVR^{+}17, CLV17, CLO^{+}19, DMS^{+}15, DKA^{+}16, EMRS^{+}18, EPF16, FVF^{+}16, GCC^{+}18, KMK^{+}17a, KMK^{+}17b, NWD^{+}19, SAF^{+}19, WWZ^{+}17]. 2/SETDB1 [DMG^{+}19]. 200 [HBWY18]. 203 [LCM^{+}16].

3 [BVR^{+}17, DDAR^{+}16, HBDW^{+}15, KSM^{+}18, LRBB15, Les15t, LCZ^{+}16, MNL^{+}16, Sho16b, WKM^{+}15, Zha19, KMLG^{+}15, KMLG^{+}16]. 3-dependent [ACG^{+}17]. 3-independent [PAC^{+}15]. 34 [LLL^{+}18]. 36.1 [GWZ^{+}19a]. 3a [KCB^{+}16]. 3p [HZH^{+}15].

4 [LTS17, MWSM18, MWSM19, PUTM15, TCP^{+}15]. 4/ [SSM^{+}18]. 40S [KPA^{+}20, GSD^{+}15, KPA^{+}16]. 413/ [LWH^{+}18]. 43 [FSF^{+}15, Les15-31]. 45a [LFK^{+}17a, Sho17k]. 4E [KVK^{+}17]. 4E-BP [KVK^{+}17].

5 [RHC^{+}16, LWH^{+}18, SNGO16]. 5-bisphosphate [GCJ^{+}15]. 5’-Inositol [RHC^{+}16]. 51 [OMKM16]. 53BP1 [LDU^{+}16, BCMM^{+}19, Can19, LCD^{+}17, MAK^{+}16]. 5P [JJW17].

6 [ABPS17, CIS^{+}17, CYL^{+}18, dVGO^{+}16]. 60S [BMW^{+}18]. 65 [DMG^{+}19]. 6a [CYT^{+}18]. 6B [SXT16].

7 [CSC^{+}15, Sho15-34, ZQZ19]. 7/ [ZLG^{+}15]. 70/110 [PXN18].

8 [GLSS^{+}15b, GLSS^{+}15a, KKP^{+}17, SKZ^{+}18a]. 84 [SJ16].

A-type [HLW^{+}15]. A/C [CMM^{+}15]. A2 [GDD^{+}15, LFT^{+}16, ZYA^{+}17]. A2-dependent [GDD^{+}15]. A3 [Sho15o, SCP^{+}15]. AAA [SMA^{+}19]. aberrant [MSŁK^{+}18, ZLZD16]. Abl [HLLK19]. ablation [CKM^{+}16]. abscission [CWL^{+}16, JPF^{+}16, KFAMR17]. Absence [ZT15, DRMW17, MSCS19]. abundance [RDH^{+}19]. ACBD5 [CCH^{+}17, HCC^{+}17]. accent [Sho15-46]. access [Les16a]. accessibility [SPJ^{+}15]. Accessorizing [CWG15]. accounts [LDM17]. accumulation [DMC^{+}16, DRL^{+}19, KZW^{+}18]. accurate [EVR^{+}19, MDOS19]. acentric [KEV^{+}17]. acentriolar [BCS^{+}17]. acetylation [FBBRCA^{+}18]. achieve [CKJ^{+}15]. acid [CNRR^{+}17, CWI^{+}19, GWZ^{+}19a, HGF^{+}18, KNQ^{+}19, MPN^{+}18, PCK^{+}17, PKKB17, SLH17, VKT^{+}15, YPY^{+}15, vLvdKR18]. acidic [MPMP16, NEW^{+}17]. acidification [TCZ^{+}16]. acids [MF18]. acridine [PCK^{+}17]. across [CSF^{+}17, CSF^{+}18, DW17, FSF^{+}15, KST^{+}19, SD17, VMR^{+}19]. act [EBMW^{+}18, EWL16, KMBO^{+}15, RGMM18, TF19]. Actin [BCM^{+}18, CHP^{+}17, DLT^{+}18, LTG^{+}18, Les15a, NC18, Sho15a, Sho15b, VQ17, WB18, AHS^{+}18, BSL^{+}15, BBSA^{+}16, BHDK17, CJS^{+}18, CDT^{+}19, CHI^{+}15, CSA19, CLO^{+}19, CBB15, DPS^{+}18, DWH^{+}17a, DQB^{+}16, DMH^{+}15, DN16, DBG^{+}15, ES18, FML^{+}17, FLLM17, GDD^{+}15, GSP^{+}18, GTW^{+}15, GSKL^{+}18, GDB^{+}15, GFWG15, HH16, HAK^{+}15, HQW15, HM19, IYP^{+}18, ISL^{+}18, JH19, JKA^{+}15, JBE^{+}17, JIB^{+}19, KKD^{+}16, KQM^{+}19, KST^{+}19, LMR^{+}17, Les15w, LXR^{+}15, LZD^{+}16, LSS^{+}15, MBS^{+}18, MCGM15a, MCGM15b, MBT16, MSS^{+}17, NKP^{+}15, OLT^{+}19, PM18, PKH^{+}19, PMRM17, PD19, PPR^{+}19, RCS^{+}19, RHPH^{+}18, RHH^{+}18, RSCR15, Roy16, SHW^{+}17, Sch17a, SVD^{+}15, SSH^{+}15, SZSS18, Sho15d, Sho15s, Sho15-28, Sho15-68, SHH^{+}16, SPJ^{+}15, SDP^{+}15a, SDP^{+}15b, TBK^{+}16, TLMG^{+}15, Ver18, YEM^{+}19, YHS^{+}15, YSM^{+}17, YKKB17, ZAAN17, vGWC^{+}18, HR16, LW16a, Sch17a, SK18b, FBBRCA^{+}18]. actin-based [HH16]. actin-binding [GDB^{+}15, OLT^{+}19]. actin-cytoskeletal [RSCR15]. actin-dependent [DQB^{+}16, YEM^{+}19]. actin-filled [FLLM17]. actin-mediated [MBS^{+}18, vGWC^{+}18]. actinin [KT15a, KT15b]. action [KHS^{+}16, MSvO17, Sho16-37, Sho18e]. actions [MSE^{+}17, vHGD^{+}15]. activate [LMPG^{+}15, PYO^{+}18]. activated [CCS^{+}19, GSP^{+}18, GSM^{+}15, NWFY15, WHS^{+}19, WIS^{+}17, YWW17, YNN18]. activates [BNB^{+}15, CZZ^{+}15, IBG^{+}15, KOR^{+}19, LM15, MSK^{+}18, MCL^{+}15, TJF18, WFOA15]. Activating [PHKY17, SENL^{+}15, YHG^{+}17]. activation [ATH^{+}19, AGGSF^{+}16, APS^{+}17, AIS^{+}18, ANM^{+}19, BLG^{+}15, Bob17, BLZ^{+}15, CWL^{+}17, CPCtR^{+}15, CLBB15, DPS^{+}18, DCO^{+}12, DCO^{+}16, FTAB^{+}15, GBD^{+}18, GLC^{+}19, GWL^{+}19, GAS^{+}15, HHCK19, HMM^{+}19, HHH^{+}19, HB16, HLEM^{+}18, IKK^{+}18, KSL^{+}17, LH15, LS18, LWH^{+}18, LBV^{+}17, MMW^{+}19, MAK^{+}16, MCS^{+}15, MCGC^{+}15, MJSB16, NKP^{+}15, PLS^{+}15, PTMP^{+}15, QZX19, SV16, SSRG18, SK16b, SQC^{+}16, SS18, TNP^{+}15, TF16, TGQ^{+}17, UFT^{+}15, VXF^{+}15, WG16, WXC^{+}18, WWZ^{+}18, WWZ^{+}17, Woo18, WKW^{+}15, YTL15, dlRHM^{+}18]. activator [BC19, GSKL^{+}18, GTD^{+}18, ODH19, RYS^{+}15]. Active [HLST19, SHVO^{+}18, CL19, GBB^{+}19, KMK^{+}17a, KMK^{+}17b, cLNF^{+}16, MWT^{+}16, NHA^{+}19, SES^{+}19, THM^{+}19]. Activity [GSS^{+}17, AFT^{+}19, BMS^{+}17, CIK^{+}17, CYL^{+}18, DKM^{+}15, ESS^{+}17, FKL^{+}18a, FKL^{+}18b, GHD^{+}17, GBD^{+}18, GLJ^{+}17, GLS^{+}15, HZH^{+}15, HCN^{+}15, IKK^{+}18, KBB^{+}15, KBB^{+}16, LJ17a, LRBB15, Log17, LWF^{+}15, MKD^{+}18, MGT^{+}19, MpDN^{+}17, MCL^{+}15, RGM^{+}16, SSH^{+}15, TBJ^{+}17, TF19, VMR^{+}19, WDM^{+}15, ZAT^{+}19, ZAT^{+}17]. Activity-dependent [GSS^{+}17, HZH^{+}15]. actomyosin [CMMB^{+}15, HLHFG15, MHY^{+}16, NLBA^{+}15, OKN^{+}16, XS16, Mar16b]. acts [CSC^{+}15, FG15, HB18, LWH^{+}18, SBC^{+}16a, SBC^{+}16b, ZQZ19]. acute [CHZ^{+}17, GD16, NS15, NNK^{+}15]. ADAM10 [DCO^{+}12, DCO^{+}16]. ADAM10/Kuzbanian [DCO^{+}12, DCO^{+}16]. adapt [PXN18]. adaptability [Sho17j]. adapter [QZX19, WV18b]. adapters [BhHS^{+}17]. adaptor [DKR^{+}19a, DKR^{+}19b, MYT^{+}16, NEW^{+}17, SV16, SD19, VMR^{+}19, WHS^{+}19, ZY16]. adaptors [BDK^{+}18]. adapts [OI18b]. added [Sho15-30]. Adding [Sho15c]. addition [CG17, MOM^{+}18]. adducts [ABGG16]. Adenomatous [JBE^{+}17]. adenosine [XJG^{+}17]. adenylyl [CS16a]. Adherens [SOII18, BPH^{+}18, CAP^{+}16, ES18, GPAA^{+}18, KLS^{+}19, TE15, TCD^{+}15, WW16]. Adhesion [Bea16, KG15, AMS^{+}17, BP19c, CTI^{+}19, CLBB15, DBC^{+}15, DCM^{+}17, ES18, EVR^{+}19, HBWY18, HHS^{+}16, JKA^{+}15, JAHH18, JBE^{+}17, JIB^{+}19, KSG^{+}16, KS17, KOV^{+}16a, KOV^{+}16b, LLK^{+}17, LDM17, LCM^{+}16, LM19, LBJ^{+}19, MCD^{+}19, POE^{+}16, PMG^{+}17, RBZ18, SSPD15, WCL^{+}18, WIS^{+}17, ZB18]. adhesion-based [JKA^{+}15]. adhesions [FBPN^{+}18, FKG^{+}19, GGF^{+}19, PPR^{+}19, Sho15i, Sho16s, SHVO^{+}18, SZR^{+}15, TLMG^{+}15]. adhesive [HVH^{+}19]. adipocyte [SQB^{+}15]. adipocytes [BBC^{+}16]. adipogenesis [EW17]. adipogenic [OBS^{+}17]. adjacent [NF19]. Adult [GI19, LCZ^{+}16, UGHB^{+}16, WRGB^{+}15]. Advances [RS19]. Advisory [Mar19]. aerobic [ALY^{+}17]. afadin [CAP^{+}16]. affect [LRS^{+}17]. affected [LSMZ^{+}18]. affecting [NDL17]. affinity [CBB15]. after [GCZ^{+}19, HSN^{+}16, LWZ^{+}18, LDU^{+}16, MBG^{+}18b, MAK^{+}16, MpDN^{+}17, RZS^{+}15, SG17, Sho16d, TCP^{+}18, XPZ^{+}19]. again [FD18]. against [AMT^{+}15, BGJ^{+}16, ES18, LAMACE^{+}17, LUC^{+}15, PVP^{+}19]. age [Sed15a, Sho17h, TALR^{+}19, WS18]. agent [FKW^{+}17]. aggregate [OCS15]. aggregates [BCH^{+}17, BPW^{+}17]. aggregation [CST^{+}16, CN15, GUM^{+}18, HKG17, MTM^{+}17, Sho16r, Sho16v, ZLZD16]. aggrephagy [LLW^{+}17]. aging [HTLG18, KVK^{+}17, KJH18, KPEJ17, MG18, NWW17, O’D18b, SM18, KF18]. aging-induced [NWW17]. ahead [Les15s]. aids [JHF^{+}15]. Aip1 [BRY^{+}19]. Aip1/Wdr1 [BRY^{+}19]. Aip1/Wdr1-deficient [BRY^{+}19]. airway [SCK^{+}19, SCK^{+}23]. Ajuba [RBZ18]. AKT [TF19, TGQ^{+}17, PLS^{+}15]. Alan [Mar15]. alarm [Sho17i]. Alberto [Cas17a]. align [HTK^{+}16]. aligning [EAW^{+}17]. alignment [BRH^{+}16, FMS^{+}19, iNLM^{+}19, OM19, ZGZ^{+}15]. alive [FV17]. ALIX [CWL^{+}16]. All-access [Les16a]. alleviate [LCTP17]. allosteric [KSL^{+}17]. allow [BSP16, DSC^{+}18]. allows [BGH18, ITN^{+}17]. Alm1 [SPGB^{+}17]. along [BJL^{+}18, GTMZ^{+}15, Sho15a]. ALS-associated [MCH^{+}18]. ALS-linked [CGBD^{+}17]. Altan [Pow15h]. alter [TVG^{+}19]. alterations [CYH^{+}16]. Altered [YBZ^{+}18, YKO^{+}16]. Altering [LTRW15, GM16, WCL^{+}18]. alternate [Sho15-47]. Alternative [RYS^{+}15, Cas17a, GDL^{+}15, VLZ15]. alters [MGJ^{+}16, OBS^{+}17]. Alushin [O’D19d]. Alzheimer [HHS18]. amazing [Sed16d]. AMD [JERL^{+}15]. AMD-like [JERL^{+}15]. ameliorates [CKM^{+}16]. AMIGO2 [PLS^{+}15]. amino [CNRR^{+}17, GWZ^{+}19a, KNQ^{+}19, MF18, PKKB17, YPY^{+}15, vLvdKR18]. amino-acid [vLvdKR18]. amoeboid [TG17]. among [LS18]. AMPA [BNB^{+}15, FRP^{+}17, HZH^{+}15]. amphipathic [CWCG19, HGF^{+}18]. AMPK [GLJ^{+}17]. Amplification [DN16, DRL^{+}19, DSH^{+}18, LMC^{+}18, RMS^{+}18]. amplifies [LRM^{+}19, Sho15e]. amyloid [GWF17]. Ana [Sil16a]. Ana-Maria [Sil16a]. Ana2 [MBG^{+}18a]. analogs [KSM^{+}17]. Analysis [NP15, SKG^{+}16, AATP17, CTS^{+}18, DSC^{+}18, GSC^{+}16, HKK^{+}19, JSB^{+}18, KBB^{+}17, NDC^{+}19, QPZ^{+}17, RLS18a, RLS18b, SSdLA^{+}15, UBR^{+}17, WMK^{+}16]. analyzing [BMP^{+}18]. anaphase [CKKG17, FMS^{+}19, JHF^{+}15, KBKW19, KMLG^{+}15, KMLG^{+}16, KJTY19, LWZ^{+}19, WV18b]. anaphase-promoting [KJTY19]. anaphylaxis [MDC^{+}16]. anastasis [SGB^{+}17]. Ancestral [GGWL^{+}19]. anchor [CMA19, KY15, KL17, LM15, Les15e, PLS^{+}15, PD19, Sho16d, LGH^{+}18]. anchorage [LLS^{+}18]. anchored [LKE15, SLAR^{+}16]. anchoring [CWG15, SWC^{+}17, SDP^{+}15a, SDP^{+}15b]. anchors [NDRJ15, PKC^{+}16, SER^{+}15, YIT15]. ancient [vGWC^{+}18]. Andrea [Cas17b]. Andrew [Mar17]. anemia [MCOGD^{+}17]. Aneuploidy [RMB^{+}18]. angiogenesis [LLC^{+}17, PLS^{+}15, Sho15-71, TCD^{+}15]. Angiomotin [WQD^{+}18]. angulin [SLM^{+}15]. angulin-1 [SLM^{+}15]. anillin [PUTM15, AGL^{+}15]. animal [VM19]. animals [MSE^{+}17]. animated [Nel17]. anisotropic [SOW^{+}17]. Ann [O’D17a]. Anne [Sed15a]. Annexin [GDD^{+}15, DQB^{+}16]. anomalies [MSŁK^{+}18]. antagonism [MOJ16]. antagonistic [ED17, KD17a]. antagonistically [LRS^{+}17]. Anti [TG19, OBS^{+}17, PBG18, PSCS16, SG17]. anti-adipogenic [OBS^{+}17]. anti-fission [PSCS16]. Anti-mitotic [TG19]. anti-mouse [PBG18]. anti-rabbit [PBG18]. anti-resection [SG17]. antiaging [Pow15d]. antibody [TCP^{+}18]. antigen [CPCtR^{+}15, SDI^{+}19, ST17]. antigen-presenting [ST17]. antigens [PLD17]. antimicrobial [CYT^{+}18]. antioxidative [HGG^{+}17]. antisense [PST18]. Antonina [Spe17b]. AP [CPBG19, NEW^{+}17]. AP-1 [NEW^{+}17]. AP-1-dependent [CPBG19]. AP2 [FRP^{+}17, KSL^{+}17]. apart [TH18]. APC [JIB^{+}19, LJ17a, PLG^{+}15, YTL15]. APC-mediated [JIB^{+}19]. APC/C [YTL15]. Apical [RMOG17, SWPS^{+}19, CTI^{+}19, EKP^{+}19, FLG^{+}18, GSP^{+}18, HTK^{+}16, MMW^{+}19, MVJ^{+}19, NiYT^{+}16, PVP18, SBS^{+}18, Sho15-57, SOW^{+}17, SCK^{+}19, SCK^{+}23, SHO^{+}15-74, TNK18, VKJ^{+}15, Jan18]. apical-directed [NiYT^{+}16]. apically [ZDSM^{+}18]. apicosome [RMOG17, TST^{+}17]. apoptosis [PCK^{+}17, PCP17, VZFG^{+}18]. apoptotic [CWZ^{+}15, DCB^{+}15, OMKM16, Sho15-70, SGB^{+}17, WV18a, YHG^{+}17]. apoptotic/senescent [DCB^{+}15]. App [Sho17l]. apparatus [FZD^{+}19, KOK^{+}19, ZHP^{+}19]. appendage [FRP^{+}17]. APPL [KMBO^{+}15]. APPL1 [LRM^{+}19]. application [LW17]. applications [ISK^{+}15]. applied [KST^{+}19]. apposition [DSS^{+}15]. approach [Pow15i, Sil16b]. approaches [GLS^{+}17]. Approximated [MBF17]. Arabidopsis [VML^{+}17]. architectural [NGG^{+}16]. architecture [CMM^{+}15, CO19, DWH^{+}17b, EW17, KP18, MPA^{+}16, NKH^{+}19, RHPH^{+}18, SEMP15, VAB^{+}18, WHP^{+}18]. arcs [MHY^{+}16]. ARF [CYL^{+}18]. ARF-6 [CYL^{+}18]. ARF1 [Gen17, RLJ^{+}17]. ARF4 [EPF16]. ARF6 [MCCL^{+}15, ZDSM^{+}18, HOH^{+}16, RSC^{+}19, Sho16a, ZRDP19]. ARF6-JIP3 [MCCL^{+}15]. ARF6-JIP3/ [MCCL^{+}15]. Arfaptin [JJB^{+}19]. ARFRP1 [IB19a, IB19b]. ARHGEF17 [IWM^{+}16, MF16a]. arises [LMC^{+}18]. ARL1 [IB19a, IB19b]. Arl2 [CKX^{+}16]. Arl2- [CKX^{+}16]. ARL5 [IB19a, IB19b]. Arl8b [FdAV^{+}17, MAJ^{+}17]. Arl8b- [FdAV^{+}17]. arm [RSG^{+}15]. arms [TWD^{+}17]. ARNO [RLJ^{+}17]. ARP [Too18]. Arp2 [BVR^{+}17, HBDW^{+}15, LRBB15, Les15t, PAC^{+}15]. Arp2/3 [BVR^{+}17, HBDW^{+}15, LRBB15, Les15t, PAC^{+}15]. Arp2/3-independent [PAC^{+}15]. ARPP19 [HGC^{+}19]. array [LCP^{+}15]. arrays [NLS^{+}18]. arrest [AGGSF^{+}16, CO19, MHA^{+}16, MAK^{+}16]. arrestin [HDA^{+}17, PhHS^{+}16]. arrests [LDU^{+}16]. art [Inf18a, Sed15r]. Art1 [LHT^{+}19].

ARTD1-mediated [HGA^{+}17]. arteriogenic [VCD^{+}15]. ASAR [PST18]. ASB11 [CHL^{+}19]. ASB7 [UOT^{+}16]. ASC [BS17a, KST^{+}17a, KST^{+}17b]. Asp [IG15, SZF^{+}15]. aspects [Sch15]. Aspergillus [SMOO17]. assay [BDLB15, GRU18]. assemble [GSP^{+}18, SRT^{+}18, SJL^{+}19, WTB^{+}19]. assembled [LHA^{+}15]. assembles [CBH^{+}15, KLS^{+}19, RBP^{+}17, THM^{+}19, WDW^{+}17]. assemblies [LOG15, VLP^{+}15]. Assembling [SG19]. Assembly [GFH^{+}16, MCM^{+}17, SV16, ATS19, AMS^{+}17, ACG^{+}17, BPH^{+}15, BTV16, BYUJ17, BMW^{+}18, BNKB15, BHS^{+}19, BCS^{+}17, CHS^{+}17, CM16, CYMS^{+}19, CGD^{+}18, CSC^{+}15, CED^{+}15, CPB^{+}16, DN17, FML^{+}17, FCLoS19, GFvA^{+}15, GBK^{+}17, GSKL^{+}18, GCJ^{+}15, GSD^{+}15, GHS16a, GHS16b, GFWG15, HK15, HBS^{+}15, HM19, IWM^{+}16, JBE^{+}17, JIB^{+}19, KCB^{+}16, KY15, KWB^{+}15, KL17, KD17b, LFK^{+}17a, LXR^{+}15, MBG^{+}18a, MCL^{+}15, OSR^{+}15, PKS^{+}19, PD19, PCP17, PMP^{+}17, RLJ^{+}17, RGR^{+}18, RO18, SSdLA^{+}15, SHW^{+}17, SRI^{+}19, SSPD15, SSR^{+}17, Sho15r, Sho15-58, Sho16-31, SWPS^{+}19, TT19, VPD^{+}16, VQ17, WZR19, WFS15, Woo18, YLW^{+}15, YAHH15, ZNR^{+}18]. assess [MTC17]. Assessing [Sho15d, BG18]. assists [SPWM15]. associate [KPA^{+}16, KPA^{+}20, NKH^{+}19]. associated [AIK^{+}16, ACG^{+}17, BPH^{+}18, BBMM^{+}16, CSF^{+}17, CSF^{+}18, DLH^{+}19, EAW^{+}17, GLS^{+}15, GG16, HKT^{+}17, HV17, JNW15, KWB^{+}15, LEM17, LSJY15, LZC^{+}15, Lov18, MGJ^{+}16, SPE^{+}17a, TVG^{+}19, WLJ16, CNA^{+}17, MCH^{+}18]. associates [AGB^{+}19]. association [ARB^{+}19, BGH18, FdSR^{+}17, FTDC17, GWL^{+}19, NHG^{+}18, zLSSS^{+}18]. associations [CCH^{+}17]. aster [DBG^{+}15]. Asterless [GJFR16, Sho15e, KGN^{+}15]. asters [Sho16j, TKM16]. astral [KNPC16]. astrocytes [SQC^{+}16]. Astrocytic [HS16]. Asymmetric [PSC^{+}15, SXE^{+}19, AZ19, BCH^{+}17, CWL^{+}17, CKX^{+}16, JDZ^{+}16, KZW^{+}18, OWW^{+}19, PUTM15, VY18]. Asymmetrically [BGJ^{+}16]. asymmetry [RSvW^{+}15]. ataxia [DLH^{+}19, MNL^{+}16]. Ataxin [MNL^{+}16, Sho16b]. Ataxin-3 [MNL^{+}16, Sho16b]. ATF4 [HGM^{+}19, KVK^{+}17, QPZ^{+}17, Sho17a]. ATF4-dependent [HGM^{+}19]. ATG [MLJ^{+}16, SD16b]. ATG2 [Kti19, VYB^{+}19, GSRG^{+}18]. Atg2-dependent [GSRG^{+}18]. Atg5 [NWFY15]. Atg8 [KJF^{+}18, NPU^{+}16]. ATG8s [Mar16a]. ATG9 [SE19, GSRG^{+}18]. ATG9A [JJB^{+}19]. atherosclerosis [TGCO15]. atlastin [LKM^{+}15a, Sho15-54, WHL17, WMH^{+}18]. atlastin-mediated [LKM^{+}15a]. Atlastins [LLAC18a, LLAC18b]. ATM [CCS^{+}19, Sch19]. Atoh1 [WRGB^{+}15]. ATP [CCS^{+}19, XJG^{+}17]. ATP13A2 [WTC^{+}19]. ATP8B1 [BDZ^{+}15]. ATP8B1-mediated [BDZ^{+}15]. ATPase [SMA^{+}19, UFT^{+}15]. ATPase/ [WHB^{+}18]. ATPase/ubiquitin [SMA^{+}19]. ATR [ATH^{+}19, BC19]. atrophy [ARV^{+}18, PPB^{+}15]. attached [KD19]. attachment [GCL^{+}15, RVS^{+}19, YAHH15, ZYA^{+}17]. attachments [ASZ^{+}18, DRMW17, FD18, KD17b, LM19]. attenuate [TAQ^{+}19]. attenuating [SBP^{+}16]. attraction [Sho16-28]. Atypical [MGJ^{+}16, GAS^{+}15]. auditory [LMdM^{+}16, PCM16]. augments [DGS^{+}18].

Augmin [DRL^{+}19, SKZ^{+}18b]. Aurora [ASZ^{+}18, AFT^{+}19, BHS^{+}19, BCS^{+}17, BCMM^{+}19, DMB^{+}18, EJK^{+}16, HLEM^{+}18, IGK^{+}16, MSK^{+}18, PTMP^{+}15, RGM^{+}16, RSG^{+}15, Sed15f, TWD^{+}17, Woo18]. autoimmunity [dVGO^{+}16]. Autoinhibition [KYN^{+}18, QZX19]. autointegration [HSK^{+}19]. Automated [BFPD19]. autonomous [CHH^{+}15]. autonomously [WYHG17]. Autophagic [TGK^{+}19, LTB^{+}17]. Autophagosomal [MJN^{+}18, BPL^{+}18, SKN19]. Autophagosome [ZZ19, BPL^{+}18, GRU18, HM19, JJB^{+}19, Kti19, MSV^{+}19, MHI^{+}18, NPU^{+}16, VYB^{+}19, WTC^{+}19, iYJF^{+}16, ZWZ^{+}19]. autophagosomes [CZL^{+}15, KJF^{+}18, NNH17, Sho15f, Sho16s]. Autophagy [GLL^{+}18a, MTGG18, VV17b, WYHG17, CD18, CYH^{+}16, GSCIL^{+}15, GTD^{+}18, HSZ^{+}18, KSG^{+}16, KJC^{+}15, KH19, LLAC18a, LLAC18b, Mar16a, MLJ^{+}16, MOS^{+}18, NCV^{+}16, PPK^{+}16, SSRG18, SE19, Sho15x, SD16b, TCP^{+}15, TCZ^{+}16, VTG^{+}16, VMP16, WTSA17, WCY^{+}16a, WCY^{+}16b, gXNG^{+}15, gXNG^{+}16, ZQZ19, ZZ16, ZZW^{+}19]. autophagy-deficient [WCY^{+}16a, WCY^{+}16b]. autophagy-dependent [KSG^{+}16]. autophagy-related [CD18]. Auwerx [Inf18c]. availability [AB18, CCBC19, PKKB17]. available [Ava18]. avenues [MG18]. avidity [GPD^{+}19]. avoid [ML15b, RM16, Sho16z]. avoiding [JW19]. award [Sho16a]. away [Bra16, Rab17, VHB18]. axis [EPF16, GWZ^{+}19b, iHMM^{+}17, LDU^{+}16, MBS^{+}18, PTR^{+}19]. Axon [BBW16, AWS^{+}16, AGB^{+}19, DKM^{+}15, FFG^{+}18, FSF^{+}15, GKGK16, HR16, IYP^{+}18, KKC^{+}19, KBT^{+}15, LZH^{+}18, MvVV^{+}16, MSS^{+}17, PC17, SEMP15, XTT^{+}18, ZAT^{+}19, ZYL^{+}16, vBMG^{+}15]. Axonal [CZL^{+}15, Sho15f, BBW16, CDT^{+}19, GHD^{+}17, GWF17, NNH17, NWD^{+}19, OFP^{+}19, Roy16, Sed15r, VXF^{+}15, WFOA15, WTSA17, WRH^{+}16, YKO^{+}16, ZZW^{+}19]. axoneme [SSPD15, VPD^{+}16, ZHP^{+}19]. axons [GTW^{+}15, Sho15a].

B [FKO^{+}18, Sho15g, ST17, ASZ^{+}18, AFT^{+}19, BHS^{+}19, EJK^{+}16, HHCK19, HHH^{+}19, Hu15, JGCAC^{+}15, LAMACE^{+}17, Les15o, MCS^{+}15, MCGC^{+}15, OSK^{+}15, PLD17, RGM^{+}16, RSG^{+}15, SDI^{+}19, Sed15f, TCP^{+}18, TWD^{+}17, WXC^{+}18, ZLG^{+}15, dVGO^{+}16]. B-dependent [LAMACE^{+}17]. B1 [BHS^{+}16, LTC^{+}18, SHC^{+}18, dVGO^{+}16]. B2 [LTC^{+}18]. B3 [KBKW19, LWZ^{+}19]. B55 [MBG^{+}18b, CHB^{+}16]. B56 [HBM^{+}19]. BACE1 [LDR^{+}19]. Back [TL17, XS16, CV19, Jor16g, JGCAC^{+}15, Sho15x]. Back-to-back [XS16]. bacteria [BLL15, CSM17]. bacterial [BHDK17, ISK^{+}15, TLMG^{+}15, VQ17, ZWB^{+}19]. Badovinac [Pow15j]. BAF [KL19]. BAG3 [ALY^{+}17]. BAIAP3 [Sør17, ZJM^{+}17]. BAK [CLV17, iHMM^{+}17]. balance [DCB^{+}15, EMRS^{+}18, Mes16, Sed15t, Sho16l, SAO^{+}17]. Balanced [Les15b]. balances [MTGG18]. Balancing [Wil15, MVJ^{+}19, TF19]. Balla [O’D18e]. ballet [Hen19]. bang [FLG^{+}18, TNK18, Jan18]. Bao [O’D18a]. Bao-Liang [O’D18a]. BAR [KJON^{+}17, SE19, SZK^{+}19, UMC^{+}15, UMC^{+}17, WYV^{+}19, WMB^{+}15]. barcoding [BCG^{+}19]. Bard [Jor16e]. barrel [JLB^{+}18, WEQ^{+}15]. Barres [AD18]. barrier [HSK^{+}19, JKD^{+}19, KHRL17, KSM^{+}18, KBB^{+}15, KBB^{+}16, MRGWB^{+}16, NIN^{+}19, Sho15-40, SSE18, SCK^{+}19, SCK^{+}23, SLM^{+}15, TCD^{+}15]. barrier-to-autointegration [HSK^{+}19]. Barriers [TG15, HNF^{+}18, Les16d, TE15]. BARs [SDHC17]. Basal [LSMZ^{+}18, MGT^{+}19, BGJ^{+}16, GBRH15, HTK^{+}16, PVP18, RDH^{+}19]. based [DATI18, FLLM17, HH16, ISK^{+}15, JKA^{+}15, KDR^{+}19, KSM^{+}17, RGOS^{+}16, TYK19, THG19]. Basement [CC19, CPB^{+}16, JCK^{+}19]. basic [Sho16c]. basis [ATRG19, GFvA^{+}15, MKA^{+}17]. basket [Les16h, NGG^{+}16, SBR^{+}15]. baton [O’D18d]. Bayonets [ZB19]. BBF2H7 [ITN^{+}17]. BBSome [YNN18]. Bcl [CLV17]. Bcl-2 [CLV17]. BDNF [BLZ^{+}15, FTAB^{+}15, ODH19]. BDNF-signaling [ODH19]. be [Jor16h, LPWK15]. BEACH [LLW^{+}17]. BEACH-containing [LLW^{+}17]. beacon [PH16]. Bear [Cas16b]. Beata [Inf18a]. beating [BGJ^{+}16]. becomes [Sør17]. before [LMC^{+}18, SSRG18]. Beginning [Ger15]. behave [Les15b]. behavior [CRZ^{+}16, IBG^{+}15, Wil15]. behind [OM19]. Ben [AD18]. bending [TBJ^{+}17]. bent [MOM^{+}18]. best [NA16]. bet [DAG^{+}15, vHGD^{+}15]. Bet1 [MHA^{+}19]. BethAnn [IO18]. better [Les15b, NA16]. Between [Lov18, ABF^{+}16, BFPD19, BDLB15, CANG^{+}17, CKS^{+}15, CCH^{+}17, GSRG^{+}18, HGL^{+}17, Inf18a, KTK^{+}18, MGSO^{+}18, PMRM17, PUY^{+}19, SLW^{+}18, SCG17, SZR^{+}15, TE15, UDH^{+}16, VGB^{+}17]. Beware [NF19]. Beyond [CD18, CC19, DR16, Gar15a]. Bhalla [Sed16d]. bias [JW19, Les15d]. biased [CDT^{+}19]. BICD1 [AGB^{+}19]. BICD2 [HV17]. Bidirectional [BMF^{+}18, RFG19]. bifurcated [CKKG17]. big [FB15, FLG^{+}18, FA16, MC15, Jan18, TNK18]. BIK [CHL^{+}19]. Bin1 [NiYT^{+}16]. binders [HCML15]. Binding [MCL^{+}15, BHB^{+}18, BDLB15, BPW15, Bob17, BS17b, CCQ^{+}18, CBH^{+}15, FCB^{+}09, FCB^{+}19, GFH^{+}16, GDB^{+}15, GLC^{+}19, HKM^{+}15, HLLK19, KGN^{+}15, KDV^{+}15, MDOS19, OLT^{+}19, PKC^{+}16, QCC^{+}19, QZY^{+}19, RSC^{+}19, SER^{+}15, SMK^{+}18, SG17, SiYM^{+}18, WHiO^{+}19, WV18b, HSN^{+}16, LHY^{+}19, NIdG^{+}18, NPÖ^{+}17, YLND^{+}16]. binds [BBSA^{+}16, CPEE^{+}15, FLG^{+}18, GDB^{+}15, HBDW^{+}15, MAJ^{+}17, YVM18, ZWS^{+}16, vBMG^{+}15]. bioavailability [SAF^{+}19]. biochemical [ECAB^{+}16]. bioenergetics [BBW16, QJP^{+}17]. biogenesis [BJB^{+}18, Boh18, CGD^{+}18, FWL^{+}17, HSB^{+}19, HAR^{+}15, JLB^{+}18, JHC^{+}16, MGE^{+}15, NP15, Sho16-29, TTC^{+}16, TF16, VYB^{+}19, VKT^{+}15, WEQ^{+}15, ZNR^{+}18, ZWB^{+}19]. biologist [Mar15, O’D17c, She15]. biology [BH15, Cas16a, CZP16, DD18, Fuc15, Gar15b, GGR15, GD16, HCML15, ISK^{+}15, JDG16, MXS17, May15, MHW19, O’D17d, O’D18a, OI18a, O’D19h, RSS15, SSC^{+}19, Sch15, SQ15, SK16b, SKG17, TGCO15, Tar15, TMK18, Tra18, YH15, vS15, Hal15]. Biophysical [HSK^{+}16, ECAB^{+}16]. biorientation [FTDC17, LJ17a, RGM^{+}16]. biosensor [GPD^{+}19, OSL^{+}19]. biosensors [OSL^{+}19]. biosynthesis [Mes16]. biosynthetic [SJL^{+}19]. bipartite [SCL^{+}19]. Biphasic [FLN^{+}10, NIdG^{+}18, FLN^{+}16]. bipolarity [ZLZD16]. biportin [ATRG19]. Bipotent [TGJ^{+}17]. bisphosphate [GCJ^{+}15]. bistability [DSSF^{+}15]. bite [Sho16s]. black [THG19]. blastocyst [BMC15]. blaze [Sho15a]. Bleach [Les15c]. bleb [DATI18]. bleb-based [DATI18]. blebbing [HHBG17, MWB^{+}19]. blebs [CSA19]. BLM [CNA^{+}17, DKS15, PMHB17]. Blobel [Tra18]. blobs [NC18]. BLOC [FC16, DMS^{+}15, DDAR^{+}16, MFP17]. BLOC-1 [DDAR^{+}16, MFP17]. BLOC-2 [DMS^{+}15]. BLOC-3 [DDAR^{+}16]. block [XMJ^{+}19]. blocking [VXF^{+}15]. blocks [KKC^{+}19, VLP^{+}15]. Blood [FG16, NIN^{+}19, SLM^{+}15]. Blos1 [BMM^{+}19]. blue [BP19a, BP19b]. blueprint [KWB^{+}15]. Blume [Sed15p]. BMP [FVF^{+}16, FG16, VAKB^{+}18]. BMP-2 [FVF^{+}16]. BMPR2 [GWZ^{+}19b]. bMunc13 [KMK^{+}17a, KMK^{+}17b]. bMunc13-2 [KMK^{+}17a, KMK^{+}17b]. Bnip3 [GDL^{+}15]. Board [Mar19]. bodies [BGJ^{+}16, BBK16, HTK^{+}16, MPW^{+}19, SD16a, YIT15]. body [DS16a, HAR^{+}15, MYT^{+}16, O’D19i, RND^{+}17, Sho15h, SEMP15, SHO^{+}18g, TTC^{+}16, WMK^{+}16]. bombs [ZB19]. bond [Mok16, RPMC^{+}16]. bonding [FC15]. bones [CIS^{+}17]. Bonnet [Pow15h]. boost [SD16a]. boosts [FG16, SAK^{+}18]. Bora [TNP^{+}15]. BORC [FdAV^{+}17, PKKB17]. BORC-dependent [FdAV^{+}17]. border [MRGWB^{+}16]. Borealin [ARB^{+}19]. Borrelia [KSGL19]. Both [LBB^{+}15, Bro16, LDM15, cLNF^{+}16, Sho15b, Sho15-59, WRH^{+}16, ZPT^{+}15]. bound [FVF^{+}16, RVS^{+}19]. boundaries [Wil15]. box [THG19]. BP [KVK^{+}17]. Brain [TE15, NIN^{+}19, PCP17, Sch19, Sho15-29, SLM^{+}15]. Brajendra [O’D19a]. brake [LL19]. branch [MSS^{+}17]. Branched [ES18, HAK^{+}15, HVH^{+}19, HQW15, RHH^{+}18, SHH^{+}16]. branches [Les15w, Sho15s, Sho15-28]. branching [AATP17, ISL^{+}18, NC18, RSCR15, VM19, WSDY17]. Brangwynne [Jor16b]. BRCA1 [ABGG16]. BRCA2 [RZS^{+}15]. break [BLL15, KHA^{+}18, Les15y, PMHB17, SJ16]. breakage [DPS^{+}18]. Breaking [NL16, Inf19b, LOG15]. breaks [AWL18, CG17, DLM^{+}15, Pri17]. breakup [FA16]. breast [CBF^{+}18, DCM^{+}17, GLL^{+}18b, JPC^{+}17, Lin15, STR^{+}18]. bridge [AGL^{+}15, DPS^{+}18, SER^{+}15, Sho15h]. bridges [FA16]. Bridging [Inf18a]. Bringing [Pow16d]. brink [SGB^{+}17]. BRISC [Sho15-58, YLW^{+}15]. Brl1 [ZNR^{+}18]. broad [APK^{+}18, ISK^{+}15]. broad-spectrum [APK^{+}18]. broken [DLM^{+}15]. Brr6 [ZNR^{+}18]. Bruchpilot [SES^{+}19]. Brunet [Sed15a]. BRWD1 [Les15d, PBG^{+}15]. Bsg25D [RAS^{+}19]. BUB [KMLG^{+}15, KMLG^{+}16]. BUB-1 [KMLG^{+}15, KMLG^{+}16]. BUB-1/BUB-3 [KMLG^{+}15, KMLG^{+}16]. BUB-3 [KMLG^{+}15, KMLG^{+}16]. Bub3 [DLM^{+}15, YTL15]. BubR1 [DLM^{+}15]. BubR1-dependent [DLM^{+}15]. Buckle [KS17]. bud [NDRJ15]. budding [AFXS16, AUTM16, COGP15, DTW^{+}16, LK17, LSJY15, SBR^{+}15, SHO^{+}18g, SLD^{+}15]. buds [Sho15p]. buffer [LRD19]. buffering [MDOS19]. bug [O’D17b]. BuGZ [HLEM^{+}18, Woo18]. build [NL16, Sho16e, YKKB17]. Building [LBD18, O’D19i, Sho15h, HKM^{+}15, PVP18, VLP^{+}15, Inf19b]. builds [Sho15-48]. Built [SD16a]. bulk [EKP^{+}19]. bulky [SNOBM16]. bundle [CLO^{+}19]. bundles [JKA^{+}15, LMdM^{+}16]. bundling [GDD^{+}15, RBC^{+}17]. burden [Sho15-54]. Burning [FA16]. bypass [PHA^{+}17]. bypassed [LPWK15]. bystander [HMC^{+}16].

C [KWB^{+}15, CMM^{+}15, CAI^{+}15, Col19, HESKK15a, HESKK15b, MRWM18, NL16, RGR^{+}18, SER^{+}15, WHiO^{+}19, YTL15, FZD^{+}19, ZCL^{+}15]. C-terminal [NL16, SER^{+}15]. C-tubule [RGR^{+}18]. C. [AGL^{+}15, BNKB15, BCMG19, CSC^{+}15, GGWL^{+}19, KMLG^{+}15, KMLG^{+}16, KH19, LPGB16, Les16d, MRMM18, PMRM17, SFG^{+}17, SSPD15, SSR^{+}17, TNP^{+}15, YHG^{+}17, ZQZ19]. C1 [TVG^{+}19]. C1a [FZD^{+}19]. C1a-e-c [FZD^{+}19]. C2 [ZJM^{+}17]. Ca [CCQ^{+}18, CBM^{+}16, LE16, MPMP16, MWSM18, MWSM19, MPW^{+}19, RGOS^{+}16, SZL^{+}16, Sør17, SBP^{+}16, WZG^{+}17, WWT18]. CA3 [BLZ^{+}15]. Cab45 [CBM^{+}16]. cable [PKH^{+}19]. cables [SHW^{+}17]. Cadherin [KNL^{+}17, Sho15i, Blu15a, CHI^{+}15, KLS^{+}19, PBL^{+}16, SDP^{+}15a, SDP^{+}15b, SXT16, BMC15, CBH^{+}15, DSvNA^{+}15a, DSvNA^{+}15b, GBD^{+}18, GDB^{+}15, HLHFG15, JKD^{+}19, RMS^{+}18, Sho15v, VHB18]. Cadherin-6B [SXT16]. cadherin/ [BKG^{+}15]. cadherins [KHS^{+}16, SPE^{+}17a]. Caenorhabditis [DRMW17, KFAMR17, LYO15, ZAAN17]. Calcium [CZZ^{+}15, VMP16, BZG^{+}17, CJS^{+}18, Col19, GSM^{+}15, KBJ16, MJSB16, RYS^{+}15, SD19, SK18b, TVG^{+}19, WHS^{+}19]. calcium-activated [WHS^{+}19]. calmodulin [CZZ^{+}15]. calpain [ARV^{+}18]. calpain- [ARV^{+}18]. Calreticulin [SQB^{+}15]. CaM [SZF^{+}15]. cAMP [CS16a, GCVAGS^{+}18, IdSCB^{+}16, IKK^{+}18]. Can [Ava18, Bro16, LPWK15, LTC^{+}18, MG16, PCK^{+}17, VHB18, Ver16]. canalizes [DKMV15]. Cancer [ACG^{+}17, BBMM^{+}16, EAW^{+}17, RMTR17, ALY^{+}17, BS18, CC19, CBF^{+}18, DMC^{+}16, DCM^{+}17, GLL^{+}18b, GN18, JPC^{+}17, KKP^{+}17, Les15-30, Lin15, MB15, MCCL^{+}15, MSV16, MWSM18, MWSM19, MTC17, NF19, NKW^{+}19, O’D18g, PAC^{+}15, QSZ^{+}17a, QSZ^{+}17b, RHC^{+}16, RGOS^{+}16, RMS^{+}18, RRM^{+}17, STR^{+}18, Sch17b, Sed15b, Sed15l, Sho16i, TG19, TF19, TMFR^{+}19, VWM^{+}18, ZRDP19, vV17a]. Cancer-associated [ACG^{+}17, BBMM^{+}16, EAW^{+}17]. candle [O’D19c]. candles [O’D19c]. cannibalism [Pas19]. canonical [DRMW17, DGS^{+}18, HB18]. can’t [Kaw17]. capacity [LRD19, MpDN^{+}17, PBL^{+}19]. capillary [KJZ^{+}19]. capillary-like [KJZ^{+}19]. capping [AKD^{+}17]. Caprin1 [KPA^{+}16, KPA^{+}20]. caps [JH19]. capsid [IZZ^{+}18]. capture [BCM^{+}18, HK15, JIB^{+}19, Ver18]. carbon [Sho16-37]. carboxyl [CAA^{+}17]. carcinogenesis [ZLG^{+}15]. carcinoma [LAMACE^{+}17]. cardiac [ASPY^{+}16, BFS^{+}19, CMTH^{+}15, MT19]. Cardiolipin [VGB^{+}17, RXEB^{+}19]. cardiomyocyte [AGGSF^{+}16, DV16]. cardiomyocytes [ASPY^{+}16, DKA^{+}16]. cardiomyopathies [MHW19]. cardiomyopathy [CRC^{+}15]. cardiovascular [CLL^{+}16]. Career [Mar19, O’D17d, Sil16a]. Cargo [KJON^{+}17, MFVS18, VKJ^{+}15, CPBG19, CBM^{+}16, CCY^{+}19, DMS^{+}15, GM16, ITN^{+}17, KMBO^{+}15, KOK^{+}19, LHT^{+}19, MAJ^{+}17, MGJ^{+}16, MPW^{+}19, QZX19, SSM^{+}18, SV16, SDHC17, WV18b]. cargo- [ITN^{+}17]. cargo-adapter [QZX19]. Cargo-selective [KJON^{+}17, VKJ^{+}15]. cargo-sorting [KMBO^{+}15]. cargo-specific [LHT^{+}19]. cargos [YDM^{+}18]. carrier [DWB^{+}17]. carriers [CGPB17, CCY^{+}19, DDAR^{+}16, GYK^{+}17, MSCS19, RHH^{+}18]. cartilage [HPE^{+}19]. cartilage-mediated [HPE^{+}19]. cartography [Tar15]. Cas9 [LYO15, MTN^{+}16]. cascade [CKKG17]. case [Les16h]. caspase [APS^{+}17, GSP^{+}18, OR17, KKP^{+}17]. caspase-2 [APS^{+}17, OR17]. Caspase-8 [KKP^{+}17]. CAST [HKG^{+}18]. casts [Sho15-33]. catabolism [SWS^{+}19]. catabolite [ZWW^{+}19]. catalyzed [CR18]. catastrophe [GCL^{+}15, gXNG^{+}15, gXNG^{+}16]. Catch [Das17]. Catching [SS16, O’D17b]. Catenin [CHI^{+}15, WGHE^{+}18, WIS^{+}17, CSG^{+}15, GBD^{+}18, LJ17b, RRM^{+}17, MB17a, BKG^{+}15]. cation [LgYL^{+}18]. CatSper [EMB^{+}15]. caught [Sed15c]. cause [BHS18]. causes [DSH^{+}18, GBD^{+}18, HKG17, HGM^{+}19, YBZ^{+}18, ZT15, MG18]. causing [OBS^{+}17, Van19]. Caveolae [CMTH^{+}15, JSB^{+}18, Sho18a, TSB^{+}18]. caveolin [LNH^{+}15]. cavin [LNH^{+}15]. CAX [LE16]. Cayetano [Sed15b]. caz [MCH^{+}18]. Cbl [SSV^{+}18]. Cbx4 [CE16, MLR^{+}16]. CCM2 [DLZ^{+}15]. CCM3 [DLZ^{+}15]. CCNB1 [APHH^{+}19, HAPC^{+}19]. CCP1 [GHKW^{+}19]. CD16 [SAK^{+}18]. CD8 [CWL^{+}17, OBY^{+}15]. Cdc12 [WMB^{+}15]. Cdc15 [UMC^{+}15, UMC^{+}17, WMB^{+}15]. Cdc20 [DLM^{+}15, YTL15]. Cdc20-dependent [YTL15]. Cdc25 [HHCK19]. Cdc31 [MP17b]. Cdc31/centrin [MP17b]. Cdc31p [DOA^{+}17]. Cdc42 [BNB^{+}15, BDZ^{+}15, CM18, PBL^{+}16, SSH^{+}15, Sho15p, WKW^{+}15, ZAAN17]. Cdc42-dependent [SSH^{+}15]. Cdc42-GTP [CM18]. Cdc42-mediated [PBL^{+}16]. Cdc42p [SHR17]. Cdc48 [ZY16]. Cdc55 [Les16b]. CDK [JJW17]. CDK-dependent [JJW17]. CDK1 [APHH^{+}19, HAPC^{+}19, JAHH18, SKW^{+}19, WHiO^{+}19, HHCK19, HHH^{+}19, TNP^{+}15, WV18b]. CDK1-CCNB1 [APHH^{+}19, HAPC^{+}19]. CDK1-mediated [SKW^{+}19, WHiO^{+}19]. CDK2 [PTR^{+}19]. Cdr2 [AOL^{+}18]. Cdt1 [ASZ^{+}18]. Celebrating [Hal15]. Cell [AOL^{+}18, BCMG19, Fuc15, GPAA^{+}18, JAHH18, JSB^{+}18, LDM17, LSPC16, LM19, MTC^{+}19, MVJ^{+}19, MHW19, MKA^{+}19, NTT^{+}15, OLL^{+}17, Tar15, YEM^{+}19, ZB18, AGGSF^{+}16, AZ19, ASPY^{+}16, BDAW15, BSK^{+}19, BBMM^{+}16, BH15, BVR^{+}17, BDW19, Bro16, BJL^{+}18, CSO^{+}19, CWL^{+}17, CPCtR^{+}15, CNRR^{+}17, Cas16b, CSS^{+}18, CZP16, CIK^{+}17, CHL^{+}19, CWZ^{+}15, CEM^{+}15, CLBB15, CSM17, DBC^{+}15, DPGS^{+}18, DD18, DSS^{+}15, DMC^{+}17, DAG^{+}15, DGS^{+}18, DVS^{+}17, DBG^{+}15, DK16, ES18, EAW^{+}17, FG15, FB15, FAH^{+}17, FK17, FC19, GSP^{+}18, GKK16a, GKK16b, GCZ^{+}19, GPPJ^{+}18, GDL^{+}15, Gar15b, GGC^{+}17, GM18, GSD^{+}15, GP17, GHKW^{+}19, GSCIL^{+}15, GKGK16, GGL^{+}19, GGR15, GAS^{+}18, GD16, GTMZ^{+}15, GCC^{+}18, GWZ^{+}19b, HGC^{+}19, HKH16, Har16, HCML15, HBS^{+}15, HTK^{+}16, HBWY18, HH18, HB16, IKRMN16, IM16, ISK^{+}15, Jan18, JDG16, JOJG16, JDZ^{+}16, JBE^{+}17]. cell [KZW^{+}18, KPEJ17, KNL^{+}17, LAMACE^{+}17, LDM15, LLK^{+}17, LR18, LCM^{+}16, LL17, LK17, LL19, LJ17b, LLZ^{+}19, LBD18, LDMW^{+}15, LWF^{+}15, MBS^{+}18, MRGWB^{+}16, MLR^{+}16, MMW^{+}19, Mar15, MOJ16, MCD^{+}19, MXS17, MHG^{+}19, May15, MPMP16, MGA19, MKD^{+}18, MJSB16, MF16b, MDC^{+}16, MHY^{+}16, NKP^{+}15, NAFM^{+}17, NLBA^{+}15, O’D17c, O’D19h, OKN^{+}16, OBY^{+}15, OSK^{+}15, PLG^{+}15, PLS^{+}15, PAC^{+}15, PA19, PGRY^{+}19, PSC^{+}15, PCK^{+}17, PBL^{+}16, PLD^{+}15, PPR^{+}19, QCC^{+}19, RHC^{+}16, RHPH^{+}18, RGOS^{+}16, RBZ18, RLM^{+}15, RM16, RSS15, SPH^{+}19, SLW^{+}18, SM16, SBS^{+}18, SXT16, Sch19, Sch17b, SSPD15, SSC^{+}19, SQ15, Sed15e, Sed15t, SK16b, SS19, She15, Sho15g, Sho15-62, Sho16c, Sho18a, SAT^{+}17, SKG17, SR17a, SK18a, SHR17, ST17, SBC^{+}16a, SBC^{+}16b, SAK^{+}18, SNB^{+}18, SDP^{+}15a, SDP^{+}15b, SGB^{+}17, SYK^{+}17, TH18, THG19, THA^{+}16, TB16]. cell [TG17, TMK18, TCD^{+}15, Tra18, UMC^{+}15, UMC^{+}17, UGHB^{+}16, VRK^{+}17, VTG^{+}16, VY18, VLZ15, VZFG^{+}18, WG16, WZC^{+}15, WXC^{+}18, WYHG17, WSDY17, WCL^{+}18, WPA^{+}18, WFS15, Wil15, WV18a, WRGB^{+}15, gXNG^{+}15, gXNG^{+}16, XPZ^{+}19, YYZ^{+}15, YH15, YGW^{+}17, YHG^{+}17, YLND^{+}16, ZRDP19, ZTR^{+}17, ZAAN17, ZDSM^{+}18, ZCL^{+}15, dlFEvW^{+}15, vHGD^{+}15, vS15, Hal15]. Cell-cycle [MKA^{+}19, AGGSF^{+}16]. cell-derived [ASPY^{+}16]. Cell-free [JSB^{+}18, Sho18a, WFS15]. cell-to-cell [MOJ16]. Cells [Sho16d, ALY^{+}17, BMM^{+}19, BPH^{+}19, BRACA^{+}16, Blo19, BPS^{+}15, BUPC19, CAKL16, CPCtR^{+}15, CNC^{+}18, CSG^{+}15, CV19, CMTH^{+}15, CEM^{+}15, Col18, CDF^{+}18, DMC^{+}16, DSC^{+}18, DMH^{+}15, DVS^{+}17, ESS^{+}17, FWL^{+}17, FC19, FKL^{+}18a, FKL^{+}18b, FJ17, GBRH15, GLL^{+}18a, GLL^{+}18b, GCVAGS^{+}18, GN18, GAS^{+}18, Haw18, HHT^{+}16, HKK^{+}19, HKT^{+}17, HMC^{+}16, IGK^{+}16, Inf18b, IBG^{+}15, JNW15, KF18, KKP^{+}17, KdBKvdK15, KOIT^{+}16, Les15p, Les15o, Les15r, Les15y, LT18, MTN^{+}16, MB17a, MA17, MHA^{+}16, MWSM18, MWSM19, MpDN^{+}17, MPN^{+}18, MT19, NiYT^{+}16, NF19, Nie19, OSW^{+}17, O’D17a, OCS15, OPP^{+}18, OFP^{+}19, Ott16, PW19, PVP^{+}19, PKN^{+}15, PHKY17, PMW18, PBS^{+}16, PCM16, PMG^{+}17, Pow16b, QSZ^{+}17a, QSZ^{+}17b, RMB^{+}18, RZS^{+}15, RMS^{+}18, RMTR17, SDI^{+}19, STR^{+}18, SS16, SZF^{+}15, SSH^{+}15, Sed15l, Sed15n, Sho15-29, Sho15-70, Sho16i, Sho16q, Sho16v, Sho16-27, Sho17g, Sho17k]. cells [Sho18d, SRT^{+}18, SSE18, SKG^{+}16, SKO^{+}15, ST17, SAK^{+}18, SCP^{+}17, SHO^{+}15-74, SMN^{+}16, TGJ^{+}17, TCP^{+}18, TST^{+}17, TBL^{+}15, TCWM18, TS15b, TSJ^{+}15, TMFR^{+}19, TZC^{+}15, TALR^{+}19, TSK^{+}18, TSK^{+}19, UDH^{+}16, VM19, VBJ^{+}18a, VBJ^{+}18b, VPD^{+}16, VAKB^{+}18, VKJ^{+}15, VZFG^{+}18, WCY^{+}16a, WCY^{+}16b, WTB^{+}19, WKM^{+}15, WHB^{+}18, XMJ^{+}19, XTS^{+}15, YYM^{+}18, YTTH^{+}17, YLW^{+}15, YHS^{+}15, ZDM^{+}15, ZJM^{+}17, ZGDS^{+}16, ZZMC^{+}15, ZCH^{+}18, dVGO^{+}16, vV17a, SW18]. Cellular [BMS^{+}17, Bea16, Blu15b, BOL17, CHL^{+}19, FCB^{+}09, FCB^{+}19, FA16, GY18, GF16, HF15, KHS^{+}16, KJH18, KBJ16, LZ16, LDR^{+}19, Mar17, MRM18, MSvO17, Pas19, Pow15g, RC15, SBM17, TGCO15, Tar15]. CENP [BGH18, KWB^{+}15, LRS^{+}17, LM16, LBB^{+}15, MHSD^{+}15, NAFM^{+}17, WHiO^{+}19, WFS15]. CENP-A [LRS^{+}17, WHiO^{+}19]. CENP-C [WHiO^{+}19, KWB^{+}15]. censuses [Les15v]. center [AUTM16, LZ16, Sho16j, CM16]. centering [SXE^{+}19, TKM16, ZCH^{+}18, AZ19]. centers [JhZbYmP15]. central [FZD^{+}19, GJW^{+}17, KO19, LPHH16, WPA^{+}18, ZHP^{+}19, vBMG^{+}15]. centralspindlin [ABP^{+}19]. centrin [MP17b]. Centrin2 [PM15]. centriole [AWS^{+}18, GJFR16, KMC^{+}19, LUC^{+}15, LBD18, LTS17, MBG^{+}18a, MCL^{+}15, TYK19]. Centrioles [SSR^{+}17, BPSK^{+}16, Ver16]. Centrobin [OTG^{+}18, RGR^{+}18]. Centromere [BGH18, VGA^{+}15, AFT^{+}19, HKT^{+}17, KWB^{+}15, LBB^{+}15, WFS15]. centromere-associated [HKT^{+}17]. centromeres [EJK^{+}16, FFÁTC15, LM16, Sho15-73, Sho16-36]. centromeric [NAFM^{+}17, Sho15-65]. centrosomal [CANG^{+}17]. Centrosome [DSH^{+}18, LMC^{+}18, PCP17, Sho17b, BYUJ17, BCMG19, Cas16a, CGY^{+}19, LDU^{+}16, LJP^{+}15, LSJY15, LTS17, MAK^{+}16, PSL^{+}17, PSP^{+}15, RFO^{+}16, RMS^{+}18, SZF^{+}15, YYM^{+}18]. centrosome-unattached [RFO^{+}16]. centrosomes [BKR^{+}19, Les15v, O’D17f, PTMP^{+}15, Sed16a, VHB18]. CEP83 [LLY^{+}19]. ceramide [LCTP17]. Cerebellar [DLH^{+}19]. cerevisiae [LKM^{+}15a, LTRW15, YTL15]. chain [FML^{+}17, HPE^{+}19, LDMW^{+}15, MFVS18, OKK^{+}15, Sho15-37, Sho15-49]. chains [GDV19, LYO15]. chairs [SG17]. challenge [AR15]. challenges [LW17]. Challenging [MT19]. Chang [Jor16d]. change [UGG18]. changes [BMP^{+}18, HSK^{+}16, IZBH^{+}17, RZS^{+}15, SOII18, VCD^{+}15, VMP16]. channel [GGC^{+}17, GLS^{+}15, LgYL^{+}18, VMP16, WZG^{+}17, Zhu17]. channels [BZG^{+}17, GSM^{+}15, KBB^{+}17, Kti19, RYS^{+}15, Sed15g]. chaperone [CST^{+}16, DKM^{+}15, GHD^{+}17, GTD^{+}18, GUM^{+}18, LJ17b, QJP^{+}17, gXNG^{+}15, gXNG^{+}16, zLSSS^{+}18]. chaperone-facilitated [GUM^{+}18]. chaperone-mediated [GTD^{+}18, gXNG^{+}15, gXNG^{+}16]. Chaperones [Sed15c, CGD^{+}18, DR16, JLB^{+}18, MCM^{+}17, PXN18]. characteristics [LgYL^{+}18, MTC17]. Characterization [CXZ^{+}18, RND^{+}17]. characterizing [BKG^{+}15]. Charting [Pow16c]. Chasing [Pow16b]. CHCHD2 [ZGDS^{+}16]. check [GI19, Hu15, Sho17g]. Checkpoint [HBM^{+}19, APHH^{+}19, Blo19, BNKB15, BDW19, HAPC^{+}19, Hui19, IWM^{+}16, KY15, KD17b, MHG^{+}19, MGT^{+}19, MWF^{+}15, NHCB15, PDZ18, PCP17, RVS^{+}19]. Chemical [DD18]. chemokine [BKH^{+}15, CAI^{+}15]. chemomechanical [YBZ^{+}18]. chemotactic [KMJ^{+}18]. chemotaxis [BLG^{+}15]. Chemotherapy [TMFR^{+}19]. Chemotherapy-induced [TMFR^{+}19]. Chen [O’D17d, O’D18b]. Cheney [Jor16i]. Chiara [Sed15d]. chill [Les15-30]. chimeras [BKG^{+}15]. chip [DV16]. Chipuk [OI18a]. CHK1 [MGT^{+}19, DPS^{+}18]. Chlamydia [CRS^{+}17]. Chloride [Sho16e, CPB^{+}16]. chloroplast [ZWB^{+}19]. Chmp4c [PDZ18]. cholesterol [BNB^{+}15, GLL^{+}18a, O’D18a, SiYM^{+}18, dlRHM^{+}18]. Choosing [Sho17c]. choreography [Sed15v]. ChREBP [CIK^{+}17]. chromatid [SNB^{+}18]. chromatids [TH18]. Chromatin [BFS^{+}19, CM16, KPGG^{+}19, AIK^{+}16, BGH18, DPS^{+}18, EW17, GCA^{+}17, HLW^{+}15, HLST19, LT19a, MGA19, MWW^{+}16, MT19, NHA^{+}19, NAFM^{+}17, PLH18, STF18, Sed15v, Sed16b, Sed16e, Sho16-37, WFS15, WWTF17]. chromatin-modifying [Sed16e]. chromokinesin [KEV^{+}17, TWD^{+}17]. chromosomal [ARB^{+}19, ABPS17, FTDC17, IBFDB18, MTC^{+}19]. Chromosome [LJ17a, MSŁK^{+}18, Sho15j, ARB^{+}19, BRH^{+}16, BTV16, BG19, CM16, CG17, DW17, DTW^{+}16, EGY^{+}19, FTDC17, FMS^{+}19, KEV^{+}17, LDM15, LRS^{+}17, MDOS19, MGSO^{+}18, MHSD^{+}15, MH15, OM19, PCF^{+}19, PBG^{+}15, PST18, QZY^{+}19, RGM^{+}16, RSG^{+}15, SCNTC^{+}18, SPK^{+}18, SKW^{+}19, SJ16, TWD^{+}17, Ver18, WHP^{+}18, ZGZ^{+}15]. chromosome-wide [PST18]. chromosomes [BCM^{+}18, DLM^{+}15, Les15k, MBR19, MGSO^{+}18, PUY^{+}19, Sed16d]. Chronos [NWW17]. chylomicrons/VLDLs [SNOBM16]. CI [SDHC17]. Cilia [Les15e, ALLA18, DLT^{+}18, DCF^{+}17, GDB^{+}17, LTG^{+}18, LYO15, LLY^{+}19, LVG^{+}18, MFP17, Ott16, SSR^{+}17, Sho15-67, Sho16k, Sho16t, VAKB^{+}18, WHC^{+}19, YNN18, YSM^{+}17]. cilia-mediated [DLT^{+}18]. Ciliary [zLSSS^{+}18, BhHS^{+}17, BGJ^{+}16, CKJ^{+}15, CHH^{+}15, FZD^{+}19, MG17, RDO^{+}15, SPD^{+}17, SSV^{+}18, SSPD15, ZHP^{+}19]. ciliated [KZW^{+}18]. ciliogenesis [BRACA^{+}16, FKO^{+}18, IGK^{+}16, LHY^{+}19, LTS17, OTG^{+}18, PSL^{+}17]. cilium [CHH^{+}15, DER^{+}18, PhHS^{+}16, PM15]. cilium-autonomous [CHH^{+}15]. CIN [Sho16i]. CIN85 [YYZ^{+}15]. circadian [CZP16]. circuit [FBBRCA^{+}18, ZLG^{+}15]. circuitry [TZC^{+}15, vBMG^{+}15]. citron [JPF^{+}16]. CK1 [GSD^{+}15, WDM^{+}15]. CLAMP [KZW^{+}18]. CLAMP/ [KZW^{+}18]. CLASP [LNS^{+}19]. CLASPing [BP19a, BP19b]. Class [GCVAGS^{+}18, LMdM^{+}16, GPD^{+}19]. classes [CVL^{+}19]. classical [Pug15]. Clathrin [LBJ^{+}19, CYMS^{+}19, DSC^{+}18, FML^{+}17, FWH^{+}16, KSL^{+}17, MFVS18, NEW^{+}17, OMK^{+}17, PD19, WLC^{+}17]. clathrin-coated [MFVS18]. Clathrin-containing [LBJ^{+}19]. clathrin-dependent [OMK^{+}17]. clathrin-mediated [DSC^{+}18, FML^{+}17, FWH^{+}16, KSL^{+}17, PD19, WLC^{+}17]. Claudine [Jor16a]. Claudins [ONT^{+}19]. clean [Kon17]. clear [MSK^{+}18]. clearance [CWZ^{+}15, GSCIL^{+}15]. Clearing [Sed15e]. clears [SZE19]. cleavage [FLN^{+}10, FLN^{+}16, GSP^{+}18, JH19, LW16b, SDW^{+}19, XS16]. clients [PXN18]. Cliff [Jor16b]. CLIP [JNW15, MRK^{+}18]. CLIP-170 [JNW15]. cloaked [Hyr15]. clock [AWS^{+}18]. close [Inf18b, MB17a, SZ17a]. closer [Jor16f, SA19]. closure [DKMV15, Mar16b, TLH^{+}19, XS16, ZWZ^{+}19]. clues [KJ16]. CLUH [SPMM^{+}17]. cluster [NEW^{+}17]. clustering [CHI^{+}15, CPEE^{+}15, RHCS^{+}16]. clusters [CSF^{+}17, CSF^{+}18, RZS^{+}15, Sho15i]. Cnn [LJP^{+}15]. CNS [NLH^{+}19]. co [DMG^{+}19, zLSSS^{+}18]. co-chaperone [zLSSS^{+}18]. co-factor [DMG^{+}19]. coat [Gli17]. coated [GYK^{+}17, MFVS18]. coats [RBP^{+}17]. Cobl [ISL^{+}18]. Cobl-like [ISL^{+}18]. Cocaine [NLH^{+}19]. Cocaine-induced [NLH^{+}19]. cochaperone [ABPS17]. code [DK16, MSC19, O’D19e]. coding [NPÖ^{+}17]. Coenzyme [SJL^{+}19, MMB^{+}15]. coexpansion [WF15]. cofactor [PNE^{+}19]. cofilin [ZAT^{+}19, HBDW^{+}15]. Cofilin-dependent [HBDW^{+}15]. cofilin-mediated [LZD^{+}16]. cognate [SKL^{+}18]. cohesin [CTS^{+}18, QZY^{+}19]. cohesion [PSP^{+}15, SZF^{+}15]. cohibin [MRK^{+}18]. cohort [Sho15k]. coincidence [DWH^{+}17a]. Cold [XSJ18, Sed15g]. Cold-induced [XSJ18]. Cole [Pow15a]. coli [JBE^{+}17, DBS18]. Collagen [SCL^{+}16, ASM^{+}15, CPB^{+}16, ITN^{+}17, JCK^{+}19, Sho15m, Sho16e, Sho16g]. Collagen-derived [SCL^{+}16]. collateral [MSS^{+}17, RM16]. collectins [JNS^{+}19]. Collective [SM16, DPGS^{+}18, HKH16, MBS^{+}17, PBL^{+}16, SBC^{+}16a, SBC^{+}16b, SMN^{+}16, WCL^{+}18, ZTR^{+}17]. collectively [LM19]. collude [DR19]. colonic [AMT^{+}15]. columnar [LDM17]. columns [CED^{+}15]. come [Jor16d, O’D19d, Sho16-33]. comes [KBJ16, WS18]. Comestibles [MA17]. Coming [FC16]. command [LZ16, LS16]. command-and-control [LZ16]. commandeers [Sho15-41]. COMMD9 [LKM^{+}15b, Sho15-41]. COMMD9-dependent [LKM^{+}15b]. commitment [CANG^{+}17, Col18, DSSF^{+}15, Sho15-69, SQB^{+}15]. communicate [ML15b]. compaction [EGY^{+}19, FMS^{+}19, KP18]. compartment [BFS^{+}19, VV17b]. Compartmentalization [LPGB16, AWS^{+}16, MSC19]. Compartmentalizing [JBMM16]. compartments [CZW^{+}18, CXZ^{+}18, KMBO^{+}15, KJ16]. compensate [LTC^{+}18]. compensation [Góm17]. compete [DATI18, SG17, Sho16y]. competence [WWW^{+}18, YVlMS18]. Competitive [BDK^{+}18]. complete [Bob17]. completion [SOP^{+}16]. complex [ARB^{+}19, AHA^{+}19, BSK^{+}19, BPW15, CWG15, CGPB17, CGY^{+}19, CTI^{+}19, CRA^{+}19, CBF^{+}18, Con16, CSC^{+}15, CBH^{+}15, DOA^{+}17, DQB^{+}16, DWH^{+}17b, EEE^{+}16, FTDC17, GFvA^{+}15, GPS^{+}17, GBM^{+}15, HK15, HHS^{+}16, IBFDB18, JRH^{+}16, KHRL17, KCB^{+}16, KMLG^{+}15, KMLG^{+}16, KJTY19, KSM^{+}18, LPRW17, LRBB15, Les15g, LHA^{+}15, LKE15, cLNF^{+}16, LDR^{+}19, LTRW15, Mes16, MKA^{+}17, MDC^{+}16, NNH17, NDRJ15, NGG^{+}16, NIS^{+}16, RPMC^{+}16, RGM^{+}16, SCNTC^{+}18, SZF^{+}15, SMC^{+}15, Sho15c, Sho15-31, Sho15-43, SHO^{+}18g, SKZ^{+}18b, SBC^{+}16a, SBC^{+}16b, SMOO17, SCL^{+}19, TJF18, TE15, TRM^{+}16, VGB^{+}17, WSP^{+}18, YLW^{+}15, YIT15, ZNR^{+}18]. complex-dependent [NIS^{+}16]. complexes [ACRM17, DMD19, KPA^{+}16, KPA^{+}20, KNL^{+}17, LPRW17, LTC^{+}16, LR18, LBJ^{+}19, MSL16, RGMM18, RND^{+}17, Sho15-51]. Complexin [SES^{+}19]. complexity [Sho15c, Spe17b]. component [CTN^{+}19, MCM^{+}17, MST^{+}15, TBJ^{+}17]. components [AKTR18, CGY^{+}19, MSK^{+}18, SPGB^{+}17, WF15]. composed [MYT^{+}16]. composition [HHS^{+}16, IZBH^{+}17, JH19, KBB^{+}15, KBB^{+}16, SOII18]. Comprehensive [HKK^{+}19]. compression [KS17]. compromises [XIZ^{+}18]. Computer [FLS^{+}16, GLS^{+}17]. concatenation [BLPV^{+}17]. concentrates [DBG^{+}15]. Concentrating [ML15a, TTC^{+}16]. concept [MBT16]. Concerted [MSE^{+}17, vHGD^{+}15]. concurrently [iYJF^{+}16]. condensation [KPA^{+}16, KPA^{+}20, MH15, SPK^{+}18]. condensin [Ger18]. Condensins [MHH18, WHP^{+}18]. conditions [DTW^{+}16, KP18]. conducted [VKJ^{+}15]. cone [BFPD19, CG16, IYP^{+}18, WRH^{+}16]. confer [YGMR^{+}17]. confers [PTK16]. Confinement [MWB^{+}19, GBD^{+}18, HH16, SMN^{+}16]. confinement-dependent [GBD^{+}18]. conflicts [UDH^{+}16]. conformation [DTW^{+}16, OBS^{+}17, SMK^{+}18]. Conformational [FSB^{+}15, WHL17]. conformer [WWZ^{+}18]. conformer-specific [WWZ^{+}18]. Congressing [ACRM17]. congression [MHSD^{+}15, SMF^{+}15]. conjugates [KJ16, PAM^{+}16]. connecting [DER^{+}18, TSB^{+}18, FKO^{+}18]. connection [MW17]. connectivity [DOH^{+}17]. connects [SD19, TLMG^{+}15]. connexin [KDM^{+}18]. connexin-43 [KDM^{+}18]. consequences [MG18]. conservation [DW17, SD17]. conserved [BHS^{+}16, COGP15, CRZ^{+}16, DBS18, DZL^{+}15, FLLM17, LCD^{+}17, LNH^{+}15, TBJ^{+}17]. conspiracy [BK19]. constituent [RDN^{+}19]. Constitutive [HKT^{+}17, KWB^{+}15]. constrain [Bro16, LWF^{+}15, MSV^{+}19]. constraining [CLBB15]. constricted [XPZ^{+}19]. constriction [CJS^{+}18, Jan18, MMW^{+}19, MVJ^{+}19, RHH^{+}18]. construction [Gen17]. consult [Sho17b]. contact [BDK^{+}18, CMMB^{+}15, FKL^{+}18a, FKL^{+}18b, GSRG^{+}18, GKGK16, GBM^{+}15, Hen19, KBJ16, KLHC^{+}18, MKD^{+}18, MYN^{+}17, PHA^{+}17, SA19, SZ17a, SKZ^{+}18a, SDP^{+}15a, SDP^{+}15b, SLPW19, VMR^{+}19]. contact-dependent [CMMB^{+}15]. contacts [AEP^{+}17, DLH^{+}19, DPGS^{+}18, DSS^{+}15, GY18, MS19a, MS19b, MST^{+}15, SBS^{+}18, Sho15-30, SK18b, SJL^{+}19, VRM^{+}19]. contain [CST^{+}17, KdBKvdK15]. containing [LLW^{+}17, LBJ^{+}19, NDRJ15, TCP^{+}15, ZJM^{+}17]. content [HAK^{+}15, SZSS18]. contractile [KTM19, MSK^{+}18, SOP^{+}16, WMB^{+}15]. contractility [AHA^{+}19, KT15a, KT15b, NWP^{+}16, Wu17]. contraction [CHP^{+}17, FTAB^{+}15, GKC^{+}17, JhZbYmP15, MXV^{+}16, TY16]. contractions [MRMM18]. contribute [GCZ^{+}19, HNF^{+}18, Mar16a, SFG^{+}17, SKN19]. contributes [LXR^{+}15, SMK^{+}18]. Contribution [NEW^{+}17, RVS^{+}19, VPD^{+}16]. contributions [ECAB^{+}16]. Control [AHS^{+}18, DWH^{+}17a, LLW^{+}15, SPK^{+}18, AZS^{+}15, BCH^{+}17, BSP16, CD18, CBAP^{+}17, Can17, CS16b, CYMS^{+}19, CE16, CED^{+}15, DPGS^{+}18, DSSF^{+}15, DZB^{+}18, FG15, GJFR16, GWZ^{+}19a, GSKL^{+}18, GN18, GSM^{+}15, GCC^{+}18, HGC^{+}19, HHM15, HCN^{+}15, HCS^{+}18, HB18, IM16, LL17, LOG15, Les15x, LM19, LZ16, LFK^{+}17b, LVG^{+}18, MKA^{+}19, iNLM^{+}19, NGG^{+}16, PXN18, PLD^{+}15, RLM^{+}15, SSM^{+}18, SZE19, SAF^{+}19, SG18a, SG18b, SLAR^{+}16, SK18b, SB19, TWD^{+}17, UBBSM15, WV18a, YHS^{+}15, YLND^{+}16, ZRDP19, vdVFM^{+}17]. controlled [ABF^{+}16, ANM^{+}19, MCM^{+}17, RLJ^{+}17, TJMM^{+}18, WF15, vHGD^{+}15]. controlling [CST^{+}16, DCM^{+}17, DLBMA^{+}15, SPMM^{+}17, SHH^{+}16, WBNH18, WWZ^{+}17]. controls [ALLA18, BHS^{+}19, CW17, CIK^{+}17, CCLL17, CLL^{+}16, CSYB^{+}17, CRA^{+}19, CKKG17, CHB^{+}16, CCBC19, Das17, DCB^{+}15, DLT^{+}18, EVR^{+}19, FBBRCA^{+}18, FVF^{+}16, FC19, GDB^{+}17, GPPJ^{+}18, GCJ^{+}15, GCH15, GGL^{+}19, HKH16, HAK^{+}15, HQW15, HPW^{+}17, HKT^{+}17, HDA^{+}17, HAR^{+}15, JH19, JPF^{+}16, KCB^{+}16, KKP^{+}17, KSM^{+}18, KQM^{+}19, LSPC16, LLC^{+}17, LTS17, LDG^{+}15, MRGWB^{+}16, MpDN^{+}17, MDC^{+}16, NHG^{+}18, NGX^{+}19, OTG^{+}18, OSK^{+}15, PLS^{+}15, PSC^{+}15, PAM^{+}16, PST18, PMG^{+}17, PKKB17, QJP^{+}17, RHPH^{+}18, RBM^{+}19, RSvW^{+}15, SDI^{+}19, SHW^{+}17, SVD^{+}15, SSH^{+}15, SEMP15, SCK^{+}19, SCK^{+}23, SiYM^{+}18, SKZ^{+}18a, SDP^{+}15a, SDP^{+}15b, SYK^{+}17, TCD^{+}15, WYHG17, WZR19, WQD^{+}18, WV18b, ZJM^{+}17, ZDSM^{+}18]. convene [Kon17]. convention [Sle16]. converge [HMC^{+}16, RM16]. conversion [CW17, LJS^{+}16a, LJS^{+}16b]. converts [RYS^{+}15, WWY^{+}18]. cooperate [BCS^{+}17, DAG^{+}15, SNOBM16, Sho16y]. cooperates [GWZ^{+}19a, LLS^{+}18, RAS^{+}19, SES^{+}19]. cooperating [WRV15]. cooperation [Mar17]. cooperatively [TAQ^{+}19]. coordinate [CAKL16, CWZ^{+}15, IB19a, IB19b, PUTM15]. Coordinated [EVR^{+}19, LZC^{+}15, HTK^{+}16, LCD^{+}17]. coordinately [HBWY18, MKS17, ONT^{+}19, PPK^{+}16]. coordinates [BCM^{+}18, GBRH15, LLW^{+}17, MF18, MKD^{+}18, NiYT^{+}16, RMMS^{+}17, SJJ^{+}19, YSM^{+}17]. Coordinating [Jor16c]. coordination [CZP16, MCGM15a, MCGM15b]. cop [Sho16-27]. COPII [Far16, Gli17, GYK^{+}17, RBP^{+}17, SKN19]. COPII-coated [GYK^{+}17]. COPII-dependent [Far16]. cord [CBAP^{+}17]. core [NNH17, PTK16, ZJM^{+}17]. Coronin [BRY^{+}19, HBDW^{+}15]. corrals [LTG^{+}18]. correct [DLM^{+}15, IKRMN16]. Correction [BP19a, CSF^{+}18, DMH^{+}15, DCO^{+}16, DKR^{+}19a, FLN^{+}16, FLG^{+}19, FCB^{+}19, FKL^{+}18a, GKK16a, GHS16a, IB19b, KT15a, KMK^{+}17a, KM18a, KPA^{+}20, KMLG^{+}16, KBB^{+}16, KOV^{+}16a, KST^{+}17a, LLAC18b, LJS^{+}16a, MS19a, MSW^{+}17, MWSM19, QSZ^{+}17a, RLS18a, SG18a, SCK^{+}23, SBC^{+}16a, TSK^{+}19, VBJ^{+}18a, WCY^{+}16a, gXNG^{+}16, XRH^{+}18a, HBM^{+}19]. Correlative [XRH^{+}18b, XRH^{+}18a]. Cortactin [GM16, SHH^{+}16, HQW15, KBT^{+}15]. cortex [MSK^{+}18, YKKB17]. Cortical [CMA19, JDZ^{+}16, AOL^{+}18, CMMB^{+}15, CSA19, DOH^{+}17, GM18, JH19, KL17, LM15, LSMG18, NLS^{+}18, NDRJ15, OKN^{+}16, Sho15-33, YVlMS18]. cotranslational [WYoS17]. count [KD19, PCM16]. counteracting [BCMM^{+}19, PMHB17]. couple [VGA^{+}15]. coupled [BSP16, FdSR^{+}17, IdSCB^{+}16, PhHS^{+}16]. couples [ATH^{+}19, CHI^{+}15, KT15a, KT15b, TZC^{+}15]. Coupling [ASPY^{+}16, AB18, DV16, FSB^{+}15, LHB^{+}18, POE^{+}16, SBP^{+}16, YKO^{+}16, YBZ^{+}18, Cas17a]. course [Les15o]. coverage [Les16a]. covered [KL19]. CP110 [PM15]. CRACker [SD19]. CRACR2a [WHS^{+}19]. cranial [HB18]. created [LNS^{+}19]. creates [HAPC^{+}19, OKN^{+}16]. creation [LNS^{+}19]. crest [HB18, MBS^{+}17, SXT16, SMN^{+}16]. CRHR1 [IdSCB^{+}16]. crime [CB16]. crinophagy [CLH^{+}18, Sho18b]. CRIPTO [LFT^{+}16]. crisis [BHS18]. CRISP1 [EMB^{+}15]. CRISPR [LYO15, MTN^{+}16, MHI^{+}18, Sho16-35]. CRISPR-Cas9 [MTN^{+}16]. CRISPRi [BPH^{+}19]. cristae [OMKM16, Sho16p, TBJ^{+}17]. critical [CWL^{+}17, DB15a, LLW^{+}15, LNH^{+}15, OKY^{+}16, PBG^{+}15, RBC^{+}17, TYD^{+}15, UFT^{+}15, Zhu17]. critically [IdSCB^{+}16]. CRL2 [BHS^{+}16, Bra16]. CRMP [YKKB17]. CRMP-1 [YKKB17]. Cross [KQM^{+}19, AHS^{+}18, MBT16, SZR^{+}15, TF16, WB18, ZAT^{+}17]. Cross-linker [KQM^{+}19, AHS^{+}18]. cross-linking [ZAT^{+}17]. Crossed [NPC17]. crosses [Les15-31]. crossover [WHL17, WMH^{+}18]. crossroads [Gen17, HS16, SD16b]. Crosstalk [TE15]. crowd [GN18]. crowding [SPWM15, SZK^{+}19]. crucial [FKO^{+}18, NPU^{+}16, PKH^{+}19, RPMC^{+}16, Sho16p, WEQ^{+}15]. Crumbs [DK17, PMRMS17, Sho15-61, VLZ15]. Cryo [HVH^{+}19, SAB^{+}18, HGL^{+}17]. Cryo-EM [HVH^{+}19]. cryotomography [NDC^{+}19]. Cs [O’D16a]. cTAGE5 [FWL^{+}17, SNOBM16]. CTD [EJK^{+}16]. CtIP [ABGG16]. cues [Bro16, BJL^{+}18]. Cul4 [PNE^{+}19]. Cul5 [CHL^{+}19]. Cul5-ASB11 [CHL^{+}19]. Cullin [KSM^{+}18]. Cullin-3 [KSM^{+}18]. cultures [MF16b]. curbs [HLST19]. curvature [BJO^{+}16, CWCG19, DWH^{+}17a, JDG16, LMM16, McM19, SHR17, XIZ^{+}18]. curved [MOM^{+}18]. cut [Rab17]. Cuticle [KH19]. Cutting [BP19a, BP19b, CGT16]. CXCL10 [NLH^{+}19]. CXCR4 [BKH^{+}15]. cyclase [CS16a]. cycle [AGGSF^{+}16, CEM^{+}15, GP17, HGC^{+}19, JAHH18, LL19, LSPC16, MTC^{+}19, MHG^{+}19, MGA19, MKA^{+}19, NAFM^{+}17, OLL^{+}17, SLW^{+}18, SNB^{+}18, TH18, WG16, XPZ^{+}19, YYM^{+}18, ZB18]. cycle-dependent [LSPC16, OLL^{+}17]. Cyclin [KBKW19, LTC^{+}18, LWZ^{+}19, SHC^{+}18, ZYA^{+}17, BHS^{+}16, GMTL18, HHCK19, HHH^{+}19, LTC^{+}18]. cyclin-dependent [GMTL18]. Cycling [TY16, ABF^{+}16, DDAR^{+}16, ZZ16]. cystogenesis [DSH^{+}18]. Cysts [SWPS^{+}19]. cytocortex [TNK18]. cytogenetic [SWD^{+}19]. cytogenetic-scale [SWD^{+}19]. cytohesin [RSC^{+}19]. cytohesin-1 [RSC^{+}19]. cytokinesis [ABP^{+}19, DPS^{+}18, DOH^{+}17, DKR^{+}19b, JDZ^{+}16, LSPC16, MSK^{+}18, Pol17, RBC^{+}17, SOP^{+}16, SWC^{+}17, DKR^{+}19a]. Cytokinetic [Sho16f, BDW19, CHP^{+}17, CWL^{+}16, WG16]. cytological [CZW^{+}18]. Cytomatrix [HKG^{+}18, SES^{+}19]. cytoplasm [ABF^{+}16, PH16]. cytoplasmic [BYMS^{+}19, BBK16, CTN^{+}19, KDR^{+}19, KJC^{+}15, SFG^{+}17, SMA^{+}19]. Cytoskeletal [BGKL15, HNF^{+}18, KJZ^{+}19, MCD^{+}19, RSCR15, SZR^{+}15, SAO^{+}17]. cytoskeleton [AZS^{+}15, BJO^{+}16, CSA19, CLBB15, HTK^{+}16, MBT16, NKP^{+}15, O’D17b, Sho15-33, SAT^{+}17, TLMG^{+}15]. cytosol [ZWS^{+}16]. Cytosolic [JLB^{+}18, GHD^{+}17, GSM^{+}15, HBS^{+}15, PXN18, ZLZD16]. Cytotoxic [CMB^{+}18, DAG^{+}15]. cytotoxicity [HMC^{+}16, MPH^{+}15].

D [BSL^{+}15, CZW^{+}18, GTMZ^{+}15, MF16b, NKH^{+}19, NPC17, PHKY17, PSL^{+}17, SPJ^{+}15, TYD^{+}15, VZB19, dlFEvW^{+}15]. DAAM1 [NIS^{+}16, YHS^{+}15]. DAF [ZQZ19]. DAF-7 [ZQZ19]. Dam1C [NDC^{+}19]. Dam1C/DASH [NDC^{+}19]. damage [BSP^{+}17, CR18, Gek17, KH19, OR17, OLL^{+}17, PVP^{+}19, PKN^{+}15, RZS^{+}15, RM16, SG17, WZC^{+}15, WBNH18, XPZ^{+}19, XTS^{+}15]. damage-induced [Gek17]. Damaged [Sho15l, BJB^{+}18, DBS18, PSCS16]. DAN [MBS^{+}17, Inf18b]. Dangerous [CG17]. Danica [O’D18b]. Dap12 [ZT15]. dark [BBHBFSF18, Sho18f]. DASH [NDC^{+}19]. daughter [AWS^{+}18, BPSK^{+}16, LK17, LTS17, PKN^{+}15]. daughters [Sed15b]. David [Jor16c]. Davis [Inf18b]. DC2 [SCG17]. DCAF12 [PNE^{+}19]. DDA3 [UOT^{+}16]. deacetylases [GCVAGS^{+}18]. Deacetylation [KKC^{+}19]. death [CF15, DMC^{+}17, DWB^{+}17, DGS^{+}18, GPAA^{+}18, GDL^{+}15, O’D17g, PCK^{+}17, SGB^{+}17, gXNG^{+}15, gXNG^{+}16]. decay [CTS^{+}18, SPMM^{+}17]. Deciphering [FWH^{+}16, O’D16b]. decision [AS17, BOL17, Sho17b]. decisions [HH18]. Decoding [Spe17b]. decondensation [KPGG^{+}19]. deconstructing [PVP18]. decreased [WGHE^{+}18]. decreases [MNL^{+}16]. Decrypting [Sho16g, Cas17b]. deep [Fuc15, GTW^{+}15, SK16a]. Defective [BLO^{+}16, CS16b, ZWB^{+}19]. Defects [AEP^{+}17, Blo19, CNRR^{+}17, MNL^{+}16, OSW^{+}17, RSG^{+}15]. defense [Sed15j]. deficiency [MHG^{+}19, VGB^{+}17]. deficient [BRY^{+}19, CRC^{+}15, WCY^{+}16a, WCY^{+}16b, YKO^{+}16]. deficits [VXF^{+}15, ZYL^{+}16]. define [Sed16c]. defines [BFPD19, MOJ16, RFO^{+}16, WFS15]. degeneration [BBW16, Qi17, WFOA15, WTSA17]. Degradation [BMM^{+}19, gXNG^{+}15, YDM^{+}18, BHS^{+}16, DCM^{+}17, KDM^{+}18, KJTY19, LH19, LTB^{+}17, MTGG18, MOS^{+}18, MRK^{+}18, Nie19, PA19, POE^{+}16, PMP^{+}17, SPGB^{+}17, STR^{+}18, SIO^{+}16, TGK^{+}19, UOT^{+}16, WWZ^{+}17, WLJ16, YHG^{+}17, gXNG^{+}16]. degrades [WLJ18]. degrading [MB15]. degranulation [MDC^{+}16]. degron [HESKK15a, HESKK15b]. Dejana [Pow16a]. Dekker [O’D16a]. delamination [SR17a]. delays [CSG^{+}15, GCL^{+}15, KMRD^{+}16, Ver18]. deletion [FWL^{+}17]. Delineating [RVS^{+}19, Cas16b]. deliver [Sho17g]. delivers [MLMF16, Sho17e]. delivery [BNB^{+}15, DMS^{+}15, FC16, ISK^{+}15, Lin15, PM18, Sho15-57, SHO^{+}15-74]. Delta [Sho16o]. demand [HSK^{+}16]. dementia [WLM^{+}15]. demethylase [GCA^{+}17, Pri17, UBBSM15]. demonstrates [SLD^{+}15]. Demystifying [Sed15f]. dendrite [KYN^{+}18, NC18]. dendrite-specific [KYN^{+}18]. dendrites [Bro19, GSS^{+}17]. Dendritic [Nie19, PM18, BSL^{+}15, BJL^{+}18, CPCtR^{+}15, CLBB15, FTS^{+}19, GSS^{+}17, ISL^{+}18, LMR^{+}17, LLL^{+}15, LSS^{+}15, OPP^{+}18, Qi17, Sch17a, SSH^{+}15, TTU^{+}17, VRK^{+}17, WQD^{+}18, YDM^{+}18]. DENND2B [IBG^{+}15]. Dense [ASM^{+}15, Sho15m, NNH17, ZJM^{+}17]. dense-core [ZJM^{+}17]. density [JPC^{+}17]. Dent [Sed16e]. dependencies [SSdLA^{+}15]. dependent [ASZ^{+}18, AZS^{+}15, APHH^{+}19, AOL^{+}18, AWL18, APS^{+}17, ACG^{+}17, ACRM17, AIS^{+}18, BCMG19, CPBG19, CKX^{+}16, CVL^{+}19, CMMB^{+}15, CRS^{+}17, CBM^{+}16, DQB^{+}16, DLM^{+}15, DMH^{+}15, DKA^{+}16, DLBMA^{+}15, DCF^{+}17, EEE^{+}16, Far16, FBPN^{+}18, FdAV^{+}17, GDD^{+}15, GTW^{+}15, GBD^{+}18, GLJ^{+}17, GMTL18, GLC^{+}19, GSRG^{+}18, GSS^{+}17, GWF17, HHBG17, HBDW^{+}15, HZH^{+}15, HGM^{+}19, JJW17, KT15a, KT15b, KSG^{+}16, LAMACE^{+}17, LRH^{+}15, LOG15, LMR^{+}17, LSPC16, LKM^{+}15b, LGH^{+}18, LDP^{+}15, MTC^{+}19, MPH^{+}15, MCGM15a, MCGM15b, MLMF16, MF18, MCGC^{+}15, MDC^{+}16, NIS^{+}16, OMK^{+}17, OMKM16, OLL^{+}17, QYC^{+}17, SPD^{+}17, SRF19, SSL^{+}17, SSH^{+}15, SDHC17, Sør17, SJL^{+}19, TBK^{+}16, WFOA15, WZG^{+}17, WW16, WF15, YEM^{+}19, YTL15, YDM^{+}18, YSR^{+}18, ZAT^{+}19, ZWZ^{+}19, ZCL^{+}15]. depends [CMB^{+}18, JNS^{+}19, RSC^{+}19]. DepHining [Ham18]. dephosphorylation [CHB^{+}16, LHT^{+}19, PS16]. depletion [BRY^{+}19, CCS^{+}19, GLSS^{+}15b, GLSS^{+}15a, HDA^{+}17, MLJ^{+}16, PTMP^{+}15]. depolymerase [BRH^{+}16]. depolymerization [ARV^{+}18]. depolymerizing [VGA^{+}15]. deposition [KS17, LSPC16]. deprived [HSK^{+}16]. derived [AKTR18, ASPY^{+}16, DMC^{+}16, Juh16, SCL^{+}16]. Designing [JW19]. desmin [ARV^{+}18]. Desmoplakin [KDM^{+}18, AZS^{+}15, Les15f]. desmosomes [Sho16h]. destined [MRK^{+}18]. destruction [NOS^{+}15]. Detaching [Lac19]. detachment [ACRM17]. detect [DRMW17]. detected [XTT^{+}18]. determinants [EBMW^{+}18, UKHK15, VRK^{+}17, VRM^{+}19]. determination [LLS^{+}18]. determine [YEM^{+}19]. determined [BYUJ17, OCS15]. determines [CHL^{+}19, CGBD^{+}17, HPB19, JOJG16, MWW^{+}16, PhHS^{+}16, RGOS^{+}16, SRI^{+}19, WXC^{+}18, WCY^{+}16a, WCY^{+}16b]. Determining [PS16]. Detyrosinated [MSV^{+}19]. detyrosination [YCSJ^{+}17]. Deubiquitinating [RDH^{+}19, YLW^{+}15]. Deursen [Pow15d]. developing [CIS^{+}17, DGS^{+}18, MRO^{+}15, MLR^{+}16, SCL^{+}16]. development [DSH^{+}18, GJFR16, GFH^{+}16, GGL^{+}19, HGC^{+}19, HKG^{+}18, HCN^{+}15, KVK^{+}17, LLS^{+}16, LMR^{+}17, LXJ^{+}17, OWW^{+}19, PA19, PSC^{+}15, RGR^{+}18, SM16, Sho15-50, SCK^{+}19, SCK^{+}23, SCP^{+}15, SLG^{+}18, THG19, TS15a, ZWW^{+}19, dVGO^{+}16, vS15]. development-specific [GFH^{+}16]. Developmental [SJJ^{+}19, DD18, DSSF^{+}15, ITN^{+}17, She15]. Developmentally [LHY^{+}19, CLH^{+}18]. devices [EWL16]. DGCR8 [CSYB^{+}17]. Dia1 [FBPN^{+}18]. Dia1-dependent [FBPN^{+}18]. diabetes [PW19]. diabetic [CIK^{+}17, ZPT^{+}15]. diacidic [CGBD^{+}17]. Dialogue [CANG^{+}17]. Diaphanous [LM16]. diaphragms [CRPSC^{+}19]. DICER [CR18, BSP^{+}17]. DICER- [CR18]. Dickkopf [DMC^{+}16]. Dickkopf-related [DMC^{+}16]. dictate [JCK^{+}19]. dictated [SFA^{+}19]. dictates [GB18, Sho15n]. differ [RCS^{+}19]. differences [HGL^{+}17, LS18]. Different [IdSCB^{+}16, DTW^{+}16, HCS^{+}18, LDM17, PKC^{+}16, RGMM18]. Differential [CM18, MSS^{+}17, AB18, BLG^{+}15, GAS^{+}18, OKN^{+}16]. differentially [GLL^{+}18b, HGC^{+}19, KLHC^{+}18, NLBA^{+}15, ZPT^{+}15]. differentiation [BSK^{+}19, BMP^{+}18, BMS^{+}17, CWL^{+}17, CSG^{+}15, CEM^{+}15, CRK^{+}17, DSC^{+}18, DAG^{+}15, EPF16, GCH15, GCC^{+}18, GWZ^{+}19b, MSL16, OBY^{+}15, PAM^{+}16, RSCR15, Sed15e, Sho16c, SQB^{+}15, UGHB^{+}16, WYHG17, YGW^{+}17, ZGDS^{+}16, dlFEvW^{+}15]. diffuses [RZS^{+}15]. Diffusion [UKHK15, TRM^{+}16, TG15]. digested [Les16c]. Digging [TG17]. dimensions [SB17]. dimer [WMH^{+}18]. Dimerization [Sho15n, WBL^{+}15]. dimers [MB17a]. dine [TS15b]. DIP [NWD^{+}19]. DIP-2 [NWD^{+}19]. diploid [YYM^{+}18]. Direct [VM19, CYT^{+}18, FRP^{+}17, HLLK19, JKA^{+}15, KTK^{+}18, KMJ^{+}18, PBL^{+}19, WMB^{+}15]. Directed [HLW^{+}15, SW18, CEM^{+}15, DRL^{+}19, FLN^{+}10, FLN^{+}16, GSD^{+}15, NiYT^{+}16, PPR^{+}19, YTTH^{+}17, YVM18]. directing [YGW^{+}17]. directional [EAW^{+}17]. directionality [CLL^{+}16]. directly [ALY^{+}17, BBSA^{+}16, DOA^{+}17, SKZ^{+}18b, vBMG^{+}15]. directs [APS^{+}17, CWI^{+}19, HLHFG15, KDA^{+}18, MVJ^{+}19, SWS^{+}19, SHR17, TLH^{+}19]. disables [PSCS16]. disaggregase [OCS15]. disassembles [SAK^{+}18]. disassembly [IBFDB18, LDG^{+}15, WHC^{+}19, WW16, WMH^{+}18]. disc [MG17, Pug15]. Discrete [CAKL16, MTC^{+}19]. Discs [DSA15, FLG^{+}18, SPD^{+}17]. Disease [HV17, TVG^{+}19, CS16b, DLH^{+}19, HHS18, KJH18, NPC17, SS19, TS15a, VV17b]. Disease-associated [HV17, TVG^{+}19]. diseases [HPE^{+}19, KM17, KM18a, MB17b, VZ17]. Dishevelled [LHY^{+}19]. disjunction [RSG^{+}15]. dismutases [WBNH18]. disordered [BA18, SZK^{+}19]. dispensable [SSPD15, SSR^{+}17]. dispersal [SDW^{+}19]. dispersion [RSG^{+}15]. displays [TBJ^{+}17]. disrupt [YCSJ^{+}17]. Disrupted [Sho16h]. disrupting [ZWW^{+}19]. disrupts [DSH^{+}18]. dissection [MP17a]. dissemination [SSE18]. distal [DER^{+}18]. distance [MTC^{+}19, MS19a, MS19b]. distance-dependent [MTC^{+}19]. distant [BLL15]. Distinct [AFXS16, BPSK^{+}16, BSP16, GDV19, CGPB17, CKKG17, DDAR^{+}16, HKH16, IB19a, JCK^{+}19, KGN^{+}15, LPRW17, LTG^{+}18, MSE^{+}17, MSL16, RGMM18, SSL^{+}17, SPWM15, SHVO^{+}18, VLP^{+}15, WWZ^{+}18]. distinction [LDP^{+}15]. distinguish [PCK^{+}17]. distorting [TSFP^{+}15]. distribution [KST^{+}19, SiYM^{+}18, WZR19]. disulfide [FC15, Mok16, RPMC^{+}16]. dive [SK16a]. Divergent [MSL16]. Diverse [ZTR^{+}17, NDL17, RNP^{+}17]. diversify [Sho17d]. diverts [MG17]. dives [Sed16a]. divide [Inf18a, Les15-31]. dividing [Gra16]. Division [CG16, AZ19, CJS^{+}18, CKX^{+}16, FK17, JDZ^{+}16, Les15a, PUTM15, PKH^{+}19, PSC^{+}15, SXE^{+}19, UMC^{+}15, UMC^{+}17, VY18, iYJF^{+}16, ZB18]. divisions [LDM17]. Dkk1 [DMC^{+}16]. DLC1 [TGQ^{+}17, TAQ^{+}19]. Dlg1 [AHA^{+}19]. Dlx3 [UBBSM15]. DNA [ATH^{+}19, AWL18, ABGG16, Bob17, BG18, BHS18, BSP^{+}17, BCMM^{+}19, Can19, CR18, Col18, DLM^{+}15, EMRS^{+}18, Gek17, Ger15, GCA^{+}17, GRB19, GCW^{+}16, LVF^{+}15, LL19, LS16, LCD^{+}17, Lov18, MTGG18, MN17, OR17, OLL^{+}17, PMRM17, PMHB17, PVP^{+}19, PKN^{+}15, PUY^{+}19, Pri17, RZS^{+}15, RS19, RLS18a, RLS18b, SG19, Sed15q, SG17, TSFP^{+}15, WZC^{+}15, WSP^{+}18, XIZ^{+}18, XPZ^{+}19, XTS^{+}15, YGMR^{+}17, YTGA16, vV17a]. DNA2 [TBL^{+}15]. Dnm1-independent [Gra16]. dNTP [QSZ^{+}17a, QSZ^{+}17b]. Do [Haw18, Mar16b, NA16, Sed15u, Sed16d]. Doa1 [WLJ16, ZY16]. Doa10 [HESKK15a, HESKK15b]. DOCK7 [NYW^{+}17]. docking [GDD^{+}15, NGX^{+}19]. does [PKS^{+}19, SD17, TT19]. doesn’t [Les15n]. Doing [NA16]. Domain [LL17, BPH^{+}18, BA18, CE16, DZL^{+}15, GLC^{+}19, GUM^{+}18, ISL^{+}18, KCB^{+}16, LBB^{+}15, LDG^{+}15, NL16, RC15, SOW^{+}17, SLG^{+}18, TCP^{+}15, WYV^{+}19, ZJM^{+}17]. domain-containing [TCP^{+}15].

Domain-specific [LL17]. domains [CST^{+}17, HKM^{+}15, KGN^{+}15, KNQ^{+}19, PKC^{+}16, SWD^{+}19, SG19, SZK^{+}19, SJL^{+}19, XRH^{+}18a, XRH^{+}18b]. dominant [LPWK15]. donor [DV16]. dorsal [DKMV15, SHW^{+}17, SZL^{+}16]. dosage [Góm17]. double [BLL15, BSP^{+}17, PMHB17, SJ16]. double-strand [BLL15, PMHB17]. double-stranded [BSP^{+}17]. doublets [BMF^{+}18]. down [Inf19b, NL16, PMP^{+}17, vV17a]. down-regulation [PMP^{+}17]. downs [ZZ16]. downstream [BSP16, GFWG15, NWFY15, PSC^{+}15, SYK^{+}17]. DPP [DKMV15, LWF^{+}15]. DPP-mediated [DKMV15]. Draper [WV18a]. Draxin [HB18]. drift [HKT^{+}17]. Drive [BK19, GGF^{+}19, JBE^{+}17, MRMM18, MSE^{+}17, NLBA^{+}15, OKN^{+}16, SMF^{+}15, SSL^{+}17, Sho16h, SZK^{+}19, SBC^{+}16a, SBC^{+}16b, TNP^{+}15, XS16, YTTH^{+}17]. driven [CPBG19, FFG^{+}18, HOH^{+}16, KL17, MB17b, SNGO16, LHA^{+}15, PGMM^{+}19]. driver [NKW^{+}19]. drives [BNB^{+}15, CBB15, DRL^{+}19, GUM^{+}18, KS17, KJZ^{+}19, KSGL19, LRH^{+}15, LCP^{+}15, LEM17, Lin15, LE16, PSL^{+}17, QYC^{+}17, TBL^{+}15]. driving [WHL17]. drop [Les15e]. droplet [Boh18, COGP15, DLH^{+}19, EBMW^{+}18, GBM^{+}15, HSB^{+}19, NO19, OKY^{+}16, SWS^{+}19, SAB^{+}18, VTG^{+}16, XLW^{+}18, DLH^{+}19]. droplets [CWI^{+}19, FW16, GBK^{+}17, GSB^{+}15, KOR^{+}19, NO19, SMA^{+}19, TJMM^{+}18]. Drosophila [EG19, SDW^{+}19, TCWM18, TNK18, CKJ^{+}15, CLH^{+}18, DSS^{+}15, DKMV15, FAH^{+}17, FLG^{+}18, JH19, KO19, KVK^{+}17, KDA^{+}18, KPEJ17, LSMZ^{+}18, LZC^{+}15, LPHH16, LLS^{+}18, LWF^{+}15, MBG^{+}18b, OSW^{+}17, POTZ15, RGR^{+}18, RMB^{+}18, RAS^{+}19, SOW^{+}17, TSJ^{+}15, VPD^{+}16, WLM^{+}15, XS16]. Drp1 [Gra16, JCF^{+}17, OMKM16, iYJF^{+}16]. Drp1-dependent [OMKM16]. Drp1/Dnm1 [Gra16]. Drp1/Dnm1-independent [Gra16]. drug [HOH^{+}16, LLZ^{+}19]. druggable [NKW^{+}19]. DSCR1 [CG16, WRH^{+}16]. dTBC1D7 [RHJW18]. Dual [MSK^{+}19, NYW^{+}17, WWTF17, ABP^{+}19, WBNH18]. Duchenne [NWP^{+}16]. due [MSŁK^{+}18]. Duménil [Sil16a]. Dumont [Pow16e]. duplication [KMC^{+}19, LUC^{+}15, RND^{+}17, TYK19]. duration [LK17, MAK^{+}16]. during [ABP^{+}19, AGB^{+}19, AIS^{+}18, BPH^{+}15, BMP^{+}18, BMC15, BVR^{+}17, BDZ^{+}15, BCMM^{+}19, CPP^{+}18, CM18, CHL^{+}19, CMTH^{+}15, CHP^{+}17, CO19, CHH^{+}15, CPB^{+}16, CHB^{+}16, DSC^{+}18, DSSF^{+}15, DB15b, DKS15, DKMV15, EJK^{+}16, EMB^{+}15, EPF16, FML^{+}17, FDR^{+}16, FMS^{+}19, FCLoS19, FC16, GMTL18, GKGK16, GGL^{+}19, GCW^{+}16, HM19, HLHFG15, IZZ^{+}18, IZBH^{+}17, JPC^{+}17, JH19, JAHH18, JDZ^{+}16, KVK^{+}17, KCB^{+}16, KTM19, KPEJ17, KFAMR17, Lac19, LLS^{+}16, LSPC16, Les15-32, LWH^{+}18, LDP^{+}15, MSK^{+}18, MOJ16, MGSO^{+}18, MXV^{+}16, MHG^{+}19, MvVV^{+}16, MTGG18, MSW^{+}07, MSW^{+}17, MKD^{+}18, MSK^{+}19, MSL16, NVP17, NWFY15, NPU^{+}16, OBY^{+}15, OWW^{+}19, PUTM15, PTR^{+}19, PVP^{+}19, PD19, PSC^{+}15, PPR^{+}19, RBZ18, SvZS^{+}16, SHW^{+}17, SRF19, SRT^{+}18, SOW^{+}17, SK18a, SCK^{+}19, SCK^{+}23, SKO^{+}15, SBC^{+}16a, SBC^{+}16b, SNB^{+}18, SCP^{+}15, TYD^{+}15, TF16, UGG18, VRK^{+}17, VTG^{+}16, WMH^{+}18, XS16]. during [iYJF^{+}16, ZZMC^{+}15]. duties [Les15q]. Dyche [Pow15b]. dye [PCK^{+}17]. dynactin [HV17, SV16]. Dynamic [CYMS^{+}19, DJV^{+}16, GM18, GP17, IZBH^{+}17, SBM^{+}19, BMP^{+}18, BP19a, BP19b, CSA19, EGY^{+}19, GTW^{+}15, GKC^{+}17, LJ16, LLL^{+}15, NIdG^{+}18, PPR^{+}19, QYC^{+}17, Sil16a, SSE18, XRH^{+}18a, XRH^{+}18b, CJ16]. dynamically [THM^{+}19]. Dynamics [AWL18, KST^{+}17b, MBT16, SNB^{+}18, TH18, AHS^{+}18, AZS^{+}15, ANM^{+}19, BFS^{+}19, BGKL15, BS17a, BS17b, CHI^{+}15, DMB^{+}18, FFG^{+}18, FC15, FWH^{+}16, FLS^{+}16, HPB19, HNF^{+}18, HQW15, HLLK19, HCN^{+}15, IYP^{+}18, JhZbYmP15, KHA^{+}18, Les16i, LT18, LSS^{+}15, MTN^{+}16, MTC^{+}19, MNLB16, MGA19, MRMM18, MH15, MC16, MCOGD^{+}17, MG16, NVP17, NIN^{+}19, OG16, PIA16, PBS^{+}16, RSCR15, SLW^{+}18, SRI^{+}19, Sch17a, SSH^{+}15, Sho15-55, SHH^{+}16, Sle16, SK18b, SAO^{+}17, TWD^{+}17, TSB^{+}18, UOT^{+}16, WSDY17, WKM^{+}15, WBL^{+}15, YSM^{+}17, vdVFM^{+}17, KST^{+}17a]. dynamin [FRP^{+}17, MGE^{+}15, VAB^{+}18, YSW^{+}15, CLO^{+}19, LRM^{+}19]. dynamin-1 [LRM^{+}19]. dynamin-1/APPL1 [LRM^{+}19]. Dynamin-2 [CLO^{+}19]. dynamin-like [YSW^{+}15]. dynamin-related [MGE^{+}15, VAB^{+}18]. Dynein [EFM17, MW17, ODH19, AGB^{+}19, CZL^{+}15, DKR^{+}19a, DKR^{+}19b, FML^{+}17, GPS^{+}17, GDV19, HV17, KDR^{+}19, KL17, LM15, LYO15, QZX19, SMF^{+}15, SFG^{+}17, SV16, SD19, Sho15-42, UFT^{+}15, WHS^{+}19, WV18b, zLSSS^{+}18]. dynein-1 [KDR^{+}19]. dynein-mediated [SD19]. DYRK [UMC^{+}15, UMC^{+}17]. DYRK-family [UMC^{+}15, UMC^{+}17]. dysfunction [AGGSF^{+}16, HGM^{+}19, KM17, KM18a, Pow15a, Qi17, TTU^{+}17, YKO^{+}16]. dysplasia [RMB^{+}18]. dysregulation [PC17]. dystroglycan [CPEE^{+}15]. dystrophy [CKM^{+}16, NWP^{+}16].

E-cadherin [BMC15, BKG^{+}15, GBD^{+}18, HLHFG15, RMS^{+}18, VHB18]. E-cadherin/ [BKG^{+}15]. E-catenin [BKG^{+}15]. E2 [CBAP^{+}17]. E2F1 [ZCL^{+}15]. E3 [CHL^{+}19, GCW^{+}16, SvZS^{+}16, SSV^{+}18, WXFS17, XWZ^{+}15]. each [ES18, LBD18]. Early [AUTM16, JJW17, AIS^{+}18, Góm17, GSB^{+}15, HCN^{+}15, LPGB16, LJS^{+}16a, LJS^{+}16b, MP17a, MRMM18, O’D18b, SERP16, SRT^{+}18, SKL^{+}18, Mar19]. early-to-late [LJS^{+}16a, LJS^{+}16b]. earns [BS17a]. ease [Les15-28]. easily [Les16c]. eat [FV17, Les15-27]. eating [CD18]. Eaton [KS19]. EB1 [CCQ^{+}18, WRV15, WWT18, WKM^{+}15, YWdH^{+}17, YVM18]. EB1/3 [WKM^{+}15]. EB3 [YWdH^{+}17]. ECM [ECAB^{+}16, Les16c, POE^{+}16]. Ectopic [DVS^{+}17, NWD^{+}19]. ectosome [MG17, SPD^{+}17]. edge [CGT16, IBG^{+}15, PMG^{+}17]. edited [DSC^{+}18]. editing [LJ16, MSŁK^{+}18]. editorial [Les15q]. Edward [OI18a]. EEA [BKH^{+}15]. EEA-1 [BKH^{+}15]. EFA6B [ZDSM^{+}18]. effect [KJ16]. effector [BBC^{+}16, CW17, DCB^{+}15, LXJ^{+}17, LWH^{+}18, MAJ^{+}17, OBY^{+}15, PLD^{+}15]. efficiency [BHB^{+}18]. efficient [MKS17, MHA^{+}19, Sed15e, SDW^{+}19, TTC^{+}16]. efficiently [VGA^{+}15]. Efraín [O’D19b]. EGFR [CMMB^{+}15, Far16, SIO^{+}16]. egg [RBR19]. egress [TGK^{+}19]. EHBP1L1 [NiYT^{+}16]. EHD2 [TSB^{+}18]. Ehmt1 [PTR^{+}19]. eIF2 [Qi17, TTU^{+}17]. elastic [WRV15]. Electron [NDC^{+}19, BCG^{+}19, SAB^{+}18, VGB^{+}17]. elegans [AGL^{+}15, BNKB15, BCMG19, CSC^{+}15, DRMW17, GGWL^{+}19, KMLG^{+}15, KMLG^{+}16, KFAMR17, KH19, LPGB16, Les16d, LYO15, MRMM18, PMRM17, SFG^{+}17, SSPD15, SSR^{+}17, TNP^{+}15, YHG^{+}17, ZQZ19, ZAAN17]. element [PLH18]. elements [BPH^{+}15, VZB19]. Elena [Sed15g]. Elevated [WCY^{+}16b, WCY^{+}16a]. Elias [Sed15h]. elicits [KH19]. eliminate [BPSK^{+}16]. Elisabetta [Pow16a]. ELKS [HKG^{+}18]. ELKS1 [KMK^{+}17a, KMK^{+}17b, Sho17d]. elongation [ST16b, YKKB17]. Elp3 [LRH^{+}15]. embryo [BCMG19, JH19, MRMM18, XS16]. Embryonic [Mar16b, BSK^{+}19, CSG^{+}15, HLHFG15, KFAMR17, TGJ^{+}17, UGHB^{+}16, WRGB^{+}15, XTS^{+}15]. embryonically [LYO15]. embryos [LT19a, LPGB16, LFT^{+}16, SFG^{+}17, SHC^{+}18, TNP^{+}15]. emerges [UGG18]. Emerging [VY18, MBT16]. EMT [HB18]. enable [JLB^{+}18, KNQ^{+}19, MCD^{+}19]. enables [BLL15, CWCG19, CMA19, KSG^{+}16, SLG^{+}18]. Enabling [NW19, HAPC^{+}19, YNN18]. encoding [ADBST^{+}15]. encounter [VGY^{+}17]. end [AHS^{+}18, Bro19, BHDK17, DRMW17, FFG^{+}18, FTS^{+}19, FR16, Ger15, JNW15, KD17b, LTC^{+}16, MGW18, Ric18, SMF^{+}15, SFA^{+}19, VQ17, YTTH^{+}17, YWdH^{+}17, YVM18]. end-on [DRMW17, KD17b]. endocrine [FLG^{+}15, FLG^{+}19, KOIT^{+}16]. endocytic [CYL^{+}18, CXZ^{+}18, DSC^{+}18, HHT^{+}16, HHM15, KMBO^{+}15, LWH^{+}18, LTB^{+}17, NMN^{+}15, Sho15-45, WHS^{+}19]. Endocytosis [MCGM15a, MCGM15b, CMB^{+}18, CYMS^{+}19, DSC^{+}18, EKP^{+}19, FML^{+}17, FWH^{+}16, FRP^{+}17, FJ17, GKK16a, GKK16b, HDA^{+}17, HLHFG15, KSL^{+}17, LHT^{+}19, PD19, PMRMS17, RBM^{+}19, Sch17b, Sho17i, WLC^{+}17]. Endocytosis-dependent [MCGM15a, MCGM15b]. endogenous [GTD^{+}18, PBS^{+}16]. endolysosomal [CZZ^{+}15, CGPB17, MLMF16, HZB^{+}15]. endolysosomes [Juh16]. endophagosomes [NHG^{+}18]. endoplasmic [GSRG^{+}18, GSB^{+}15, HSB^{+}19, JCF^{+}17, KML^{+}15, LPGB16, LLAC18a, LLAC18b, LGH^{+}18, MHS^{+}18, NDRJ15, PYO^{+}18, Pow15f, SNOBM16]. Endosomal [HCS^{+}18, LKM^{+}15b, BDLB15, CR17, DMS^{+}15, FdAV^{+}17, HQW15, KNQ^{+}19, LPWK15, MP17a, MBS^{+}18, MCCL^{+}15, MGJ^{+}16, MOS^{+}18, MBC^{+}19, SIO^{+}16, Sho15-41, DR19, Lin15]. Endosome [DNMB16, CW17, CCY^{+}19, DOA^{+}17, LRM^{+}19, LJS^{+}16a, LJS^{+}16b, Sho15-56, AEP^{+}17]. endosomes [BKH^{+}15, BLPV^{+}17, CZL^{+}15, GAS^{+}15, GSB^{+}15, HHT^{+}16, KMBO^{+}15, KM18b, Les16f, Les16j, MPH^{+}15, MFP17, ODH19, SERP16, SKL^{+}18, VBL^{+}18, WZR19]. Endothelial [VCD^{+}15, CMTH^{+}15, CBH^{+}15, DLZ^{+}15, GGC^{+}17, HKH16, HNF^{+}18, JKD^{+}19, KSM^{+}18, KLS^{+}19, MRGWB^{+}16, Sho15-29, TCD^{+}15, VAKB^{+}18]. endothelium [YGW^{+}17]. endotube [KJZ^{+}19]. ends [IG15, KNPC16, LNS^{+}19, LVG^{+}18, NLS^{+}18, Sho15-59, Wor19, YIT15]. enemy [DR19]. energetic [HSK^{+}16]. energetics [Sch19]. Energy [LS18, CRK^{+}17, DN17, WHL17, ZYL^{+}16]. enforces [BDW19]. engage [Pas19, Sho15-52]. engagement [KCB^{+}16, SAK^{+}18]. Engineered [RFG19, FTDC17]. engulf [TMFR^{+}19]. engulfment [Log17, WV18a]. enhance [CLO^{+}19, TMFR^{+}19]. enhances [BHB^{+}18, GM16, HGG^{+}17, TCP^{+}15, YKKB17]. enhancing [ZYL^{+}16]. enriched [BKR^{+}19]. Enrichment [BHS^{+}19, LMR^{+}17]. ENSA [HGC^{+}19]. Ensconsin [RAS^{+}19]. ensure [IKRMN16, RGM^{+}16]. ensures [ABPS17, BG19, CYL^{+}18, FMS^{+}19, MDOS19, MGA19, SOP^{+}16, SCL^{+}19, ZAT^{+}17]. entangled [MGSO^{+}18]. enter [HHT^{+}16, SS16]. enterocyte [BDZ^{+}15]. enterocytes [EKP^{+}19]. enters [Sho15-63]. entorhinal [AMS^{+}17]. entorhinal-hippocampal [AMS^{+}17]. entotic [HHBG17]. entrance [CANG^{+}17]. entry [CCQ^{+}18, JHF^{+}15, LL19, MHG^{+}19, SBP^{+}16, TNP^{+}15, WWT18]. Enucleated [GAS^{+}18]. envelope [CGY^{+}19, DWH^{+}17b, HH16, KL19, LW16a, MBG^{+}18b, SPWM15, SER^{+}15, SKG^{+}16, ZWB^{+}19, GCH15]. envelope/ER [GCH15]. environment [Sho16-27]. enzyme [Les15-27, RDH^{+}19, WDW^{+}17, YLW^{+}15]. enzymes [MB15, MRM18, OSK^{+}15]. Eph [JPF^{+}16, FC19, GKGK16, OKN^{+}16, Pas16, Pas19, GGL^{+}19]. Eph-mediated [JPF^{+}16]. Eph/ [GGL^{+}19]. Eph/ephrin [OKN^{+}16]. EphA2 [CB16, NBG^{+}16]. EphB [PLD^{+}15]. EphB2 [GKK16a, GKK16b]. ephrin [GKK16a, GKK16b, Sho15o, SCP^{+}15, GKGK16, GGL^{+}19, OKN^{+}16]. Ephrin-A3 [Sho15o, SCP^{+}15]. ephrins [Pas16, Pas19]. epic [ZZ19]. epidermal [CE16, CRA^{+}19, EPF16, KBB^{+}15, KBB^{+}16, MOJ16, NTT^{+}15]. epidermis [LLC^{+}17, ZAAN17]. epigenetic [IZBH^{+}17, OBS^{+}17, UBBSM15, VWM^{+}18]. Epigenetics [YVlMS18]. Epithelia [Les16c, KZW^{+}18, SLW^{+}18, SBS^{+}18, Sho17e]. Epithelial [AMT^{+}15, AUTM16, AHA^{+}19, BPH^{+}18, BRACA^{+}16, CYL^{+}18, CTI^{+}19, ECAB^{+}16, FBPN^{+}18, FKL^{+}18a, FKL^{+}18b, GSP^{+}18, GPAA^{+}18, GPPJ^{+}18, HKK^{+}19, IM16, KT15a, KT15b, KNL^{+}17, MLR^{+}16, MXV^{+}16, MF16b, NiYT^{+}16, NIS^{+}16, ONT^{+}19, Ott16, RBZ18, RMS^{+}18, SXT16, Sho18d, SSE18, SHO^{+}15-74, TZC^{+}15, UGHB^{+}16, VWM^{+}18, VKJ^{+}15, YGW^{+}17]. epithelial-to-mesenchymal [SXT16]. epithelial/mesenchymal [VWM^{+}18]. epithelium [GBRH15, KHS^{+}16, MLR^{+}16, PMRMS17, RSCR15]. epitranscriptomics [YVlMS18]. Epo1p [NDRJ15]. EPS8 [GDB^{+}15]. equatorial [ZCH^{+}18]. ER-localized [MST^{+}15]. ER-mitochondria [SK18b]. ER-mitochondrial [LPWK15]. ER-resident [CRN^{+}19]. ER-to-Golgi [MSCS19]. ERAD [HESKK15a, HESKK15b, NOS^{+}15]. Erasing [Pri17]. Erbin [LWH^{+}18, XWZ^{+}15, CTI^{+}19]. erects [Sho15-40]. Erika [O’D18c]. Erin [O’D17b]. ERK [THA^{+}16, TAQ^{+}19]. ERMES [KTK^{+}18]. ERO1 [KML^{+}15]. ERO1-independent [KML^{+}15]. error [HBM^{+}19]. errors [DRMW17, TSFP^{+}15]. Erv41 [Les15g, SMC^{+}15]. Erv46 [Les15g, SMC^{+}15]. escape [Góm17]. escaped [SMC^{+}15]. Escherichia [DBS18]. ESCRT [BLPV^{+}17, CWI^{+}19, CWL^{+}16, MHS^{+}18, OMK^{+}17, PDZ18, TLH^{+}19, ZWZ^{+}19]. ESCRT- [OMK^{+}17]. ESCRT-I [CWL^{+}16]. ESCRT-I/II [CWL^{+}16]. ESCRT-III [CWL^{+}16, MHS^{+}18]. ESCRT-mediated [BLPV^{+}17]. essential [CST^{+}17, DKR^{+}19a, DKR^{+}19b, EEE^{+}16, GGF^{+}19, HKM^{+}15, IWM^{+}16, KYN^{+}18, LYO15, MRO^{+}15, MPN^{+}18, NW19, OMKM16, RGR^{+}18, SLD^{+}15, SHC^{+}18, SFZ^{+}17, dVGO^{+}16, vHGD^{+}15]. establish [LS18]. establishes [GSRG^{+}18, LLK^{+}17]. establishing [SEMP15]. establishment [LBB^{+}15, WKW^{+}15]. estrogen [STR^{+}18]. ETAA1 [ATH^{+}19, BC19, BG19]. ETAA1-mediated [ATH^{+}19]. Eugenia [Pow16b]. eukaryotes [DW17, SD17]. Eva [Sed15i]. evagination [DSA15, Pug15]. Even [Les15h]. event [DCP^{+}19]. events [BDW19]. every [Ava18]. eviction [AIS^{+}18, HGA^{+}17]. Evidence [OMK^{+}17]. EVL [YKKB17]. EVL-mediated [YKKB17]. evolutionarily [FLLM17]. evolutionary [KD17a, O’D19h, TG17]. evolving [Sch17b, CR17]. Ewald [Mar17]. excessive [VAKB^{+}18]. exchange [MPMP16, RLJ^{+}17, RSC^{+}19, ZTR^{+}17]. Excision [TSFP^{+}15, CR18, GCW^{+}16, WHC^{+}19]. Excitable [MRMM18, GKC^{+}17]. excitation [FGR^{+}18]. excitatory [SQC^{+}16]. Excitement [O’D19f]. excitotoxic [DWB^{+}17]. exclusion [SPWM15]. exerted [MSŁK^{+}18]. exerts [vdVFM^{+}17]. exhibits [OSW^{+}17]. exit [CANG^{+}17, CSYB^{+}17, CHB^{+}16, LCP^{+}15, LFK^{+}17b, MKS17, NP15, PS16, RBP^{+}17, SSM^{+}18, Sho18c]. exiting [Blo19]. exits [Sho16k]. exocrine [KOIT^{+}16]. exocyst [AKTR18, SDI^{+}19]. exocytosis [DB15b, GDD^{+}15, MCCL^{+}15, Sho15-66, UGG18, VKJ^{+}15, vGWC^{+}18]. exon [BPW15]. exosome [BSK^{+}19, GM16, HAR^{+}15, MWSM18, MWSM19, SHH^{+}16, VBJ^{+}18a, VBJ^{+}18b]. Exosomes [GKGK16, HHT^{+}16, Pas16, BJL^{+}18, KKP^{+}17, Rab17, SS16]. expand [Pas16, Sho15-27]. expanding [SF15]. expansion [Jan18, MWB^{+}19, RSS15]. Expecto [Bro19]. experiment [JW19]. Exploring [Cas16a, O’D17g, Sed16b]. export [ATRG19, BYMS^{+}19, BMS^{+}17, CGPB17, DMV^{+}19, EMRS^{+}18, ITN^{+}17, LFK^{+}17b, SSM^{+}18, SNOBM16, SLD^{+}15, SHC^{+}18, WWW^{+}18, vdVFM^{+}17]. exportin [APK^{+}18]. express [Góm17]. expressing [DSC^{+}18]. expression [AIS^{+}18, CAI^{+}15, CCBC19, DLZ^{+}15, DVS^{+}17, DKA^{+}16, FBBRCA^{+}18, HPB19, HZH^{+}15, Jor16c, JBMM16, LWF^{+}15, MN17, STF18, UBBSM15]. extended [MAK^{+}16]. extends [GRB19]. extension [MvVV^{+}16, WRH^{+}16]. extent [MLJ^{+}16]. Extracellular [CPB^{+}16, KMJ^{+}18, SCP^{+}17, DN17, SAO^{+}17, SW18]. extrachromosomally [CSF^{+}17, CSF^{+}18]. extract [GSC^{+}16, PLD17]. extraction [SDI^{+}19, ST17]. extracts [RBR19]. Extramitochondrial [RXEB^{+}19]. extravasate [TCWM18]. extrusion [GSP^{+}18, SBS^{+}18]. eye [O’D19a]. Ezrin [HHBG17]. Ezrin-dependent [HHBG17].

F [BCM^{+}18, AHS^{+}18, BSL^{+}15, CBB15, GTW^{+}15, JKA^{+}15, LXR^{+}15, PMRM17, RCS^{+}19, UMC^{+}15, UMC^{+}17, VQ17, VGA^{+}15, WMB^{+}15]. F-Actin [BCM^{+}18, AHS^{+}18, BSL^{+}15, CBB15, GTW^{+}15, JKA^{+}15, LXR^{+}15, PMRM17, RCS^{+}19]. F1 [GLC^{+}19]. FA [GPPJ^{+}18]. FABP4 [VBL^{+}18]. faced [DK17]. facilitate [CPCtR^{+}15, DOH^{+}17, HGA^{+}17, KEV^{+}17, SZR^{+}15]. facilitated [GUM^{+}18]. facilitates [BHS^{+}16, BPW^{+}17, COGP15, DLM^{+}15, DKS15, LCTP17, MPH^{+}15, TTC^{+}16, TGK^{+}19, WTC^{+}19, ZSdO^{+}15]. facilitating [DS16b, GLL^{+}18a, MSV^{+}19, VGY^{+}17]. Facilitation [ZYL^{+}16]. factor [Can17, CNRR^{+}17, CHS^{+}17, CR18, CRK^{+}17, CGD^{+}18, DMG^{+}19, DMV^{+}19, Ger18, HSK^{+}19, HSK^{+}16, IWM^{+}16, LPRW17, LT18, MTM^{+}17, NBG^{+}16, PLH18, PBL^{+}19, RLJ^{+}17, RHCS^{+}16, RSC^{+}19, SPK^{+}18, XTS^{+}15, YGMR^{+}17, YPY^{+}15, vBMG^{+}15]. factors [BPH^{+}19, BPH^{+}15, BMW^{+}18, BHDK17, CDF^{+}18, DUL^{+}19, DAG^{+}15, EBMW^{+}18, GLS^{+}15, IB19a, IB19b, MCOGD^{+}17, MG16, NDL17, NWFY15, NP15, SG17, TTC^{+}16, VQ17, XIZ^{+}18, ZTR^{+}17]. failure [LUC^{+}15]. FAK [GBD^{+}18, GLL^{+}18b, HHS^{+}16, KG15, LCM^{+}16]. Fam20C [GGWL^{+}19]. FAM92A1 [WYV^{+}19]. familiar [Les15x]. family [BhHS^{+}17, Blu15a, CLV17, COGP15, DATI18, GGWL^{+}19, GLS^{+}15, HBWY18, HKK^{+}19, HMM^{+}19, HLLK19, JHC^{+}16, NPU^{+}16, POE^{+}16, SID^{+}18, YTTH^{+}17, UMC^{+}15, UMC^{+}17, VLZ15]. FANCD2 [MCOGD^{+}17]. FANCI [MCOGD^{+}17]. Fanconi [MCOGD^{+}17]. Fanni [Cas16a]. far [MB17a]. farnesylated [ZWS^{+}16]. farnesylation [MWF^{+}15]. Fas [GPAA^{+}18]. Fast [GB18, SS16]. fat [RSS15, MBF17]. Fat2 [SBC^{+}16a, SBC^{+}16b, ST16b]. fate [CHL^{+}19, HH18, KDA^{+}18, PSC^{+}15, Sho17e, WCY^{+}16a, WCY^{+}16b, ZJM^{+}17]. fats [Boh18]. fattest [Sho16-30]. fatty [CWI^{+}19, MPN^{+}18]. Fc [LBV^{+}17]. feast [NF19]. feature [FCLoS19]. features [CSS^{+}18, GSC^{+}16, KOV^{+}16a, KOV^{+}16b, RFO^{+}16]. feed [DKMV15]. feed-forward [DKMV15]. Feedback [SZSS18, LRM^{+}19, LHA^{+}15, ZLG^{+}15]. feedforward [Hu15]. feely [Les15p]. Feldman [Sed16a]. female [PBG^{+}15]. fence [HR16]. Feng [Sed15j]. FERM [GPPJ^{+}18]. ferroptosis [BAGM17, DMC^{+}17]. ferroptosis-like [DMC^{+}17]. ferryman [Kti19]. fertilization [EMB^{+}15]. FGF [SAF^{+}19, SDW^{+}19]. FGF-2 [SAF^{+}19]. FGF2 [DCP^{+}19]. FGFR1 [FCB^{+}09, FCB^{+}19]. FHL1 [Sho18f, WWY^{+}18]. FHOD3 [PAC^{+}15]. fiber [LFK^{+}17a, LZH^{+}18, Sho15w, SCP^{+}15, OSR^{+}15]. fibers [LW16a, Sho15o]. fibrillar [ASM^{+}15]. fibroblast [FSB^{+}15]. fibroblasts [ACG^{+}17, EAW^{+}17, HAK^{+}15, JNS^{+}19]. Fibronectin [STR^{+}18, ACG^{+}17, EAW^{+}17, KG15]. fibrosis [Sho16h]. fibrotic [DKA^{+}16]. fidelity [ATH^{+}19, CTS^{+}18, FMS^{+}19, OM19]. Fidgetin [FFG^{+}18]. Fidgetin-like [FFG^{+}18]. field [BLZ^{+}15]. Fife [BZG^{+}17]. fight [DR19]. FIGNL1 [AGB^{+}19]. filament [ARV^{+}18, CHP^{+}17, FC15, HM19, ISL^{+}18, KKD^{+}16, LEM17, RBC^{+}17, VQ17]. filaments [DPGS^{+}18, LTG^{+}18, Pow15b, RHH^{+}18, WDW^{+}17]. Filamin [KST^{+}19]. fill [Kaw17, Sho15-73]. filled [FLLM17]. filopodia [HHT^{+}16, JPC^{+}17, MvVV^{+}16, SS16, Sho16-28, UBR^{+}17]. Filopodyan [UBR^{+}17]. FiloQuant [JPC^{+}17]. filter [BK19]. filtration [SQ15]. final [ALLA18]. find [Sho16j]. Finding [Pow15b, Pow15e]. Fine [NCV^{+}16, DZB^{+}18, TF19]. Fine-tuning [NCV^{+}16]. fingerprints [KF18]. FIP200 [WYHG17]. firm [Les16c]. first [Hal15, Pow15g, Sho15p]. fission [BYUJ17, BCH^{+}17, BPW^{+}17, CRS^{+}17, DBG^{+}15, LCP^{+}15, LXR^{+}15, OMKM16, PSCS16, RHH^{+}18, SPGB^{+}17, SPK^{+}18, Sed15w, Sho15q, Sho16p, SZK^{+}19, SK18b, TBK^{+}16, YAHH15, YIT15]. fit [Les15k, Sho16n]. fitness [RLM^{+}15]. fits [BA18]. Fitting [O’D18c]. Five [MG16]. FKB [ABPS17]. FKB-6 [ABPS17]. flagellum [BMF^{+}18, RGR^{+}18]. FLCN [MF18]. Fld1 [GBM^{+}15]. Fld1/Ldb16 [GBM^{+}15]. flies [AWS^{+}18, DCO^{+}12, DCO^{+}16]. flight [DSS^{+}15]. FLIM [VRM^{+}19]. flippase [RSvW^{+}15]. flipped [UFT^{+}15]. flips [Les15t]. flow [CDT^{+}19, CBB15, FG16, LLS^{+}18, SHW^{+}17, Sho15v, Sed15o]. Flower [CMB^{+}18]. fluid [BLO^{+}16, LTG^{+}18, Nie16]. fluidity [SDP^{+}15a, SDP^{+}15b]. fluorescence [BDAW15]. fluorescent [BCG^{+}19, KSM^{+}17, PABM16]. fluorescently [DSC^{+}18]. flux [FBX^{+}15, KBJ16, RGOS^{+}16]. Flying [O’D18f]. FMRP [Log17]. FNIP [MF18]. Focal [GGF^{+}19, FKG^{+}19, JBE^{+}17, JIB^{+}19, KSG^{+}16, KOV^{+}16a, KOV^{+}16b, MCD^{+}19, PPR^{+}19, Sho16s, SHVO^{+}18, TLMG^{+}15]. foci [DMG^{+}19, SHR17]. focus [ML15b, Sho15u]. focused [Les16b]. focuses [IG15]. folding [KTM19, KML^{+}15, LFK^{+}17a, LGH^{+}18, TSK^{+}18, TSK^{+}19]. follicle [PGMM^{+}19]. follicles [GI19]. follicular [dVGO^{+}16]. following [LUC^{+}15, SWD^{+}19]. Folsch [Jor16g]. force [ACRM17, BBHBFSF18, FTAB^{+}15, HB16, Jor16h, KTM19, KBT^{+}15, KOV^{+}16a, KOV^{+}16b, KST^{+}19, MVJ^{+}19, PLD17, RBZ18, WXC^{+}18]. force-dependent [ACRM17]. force-sensitive [RBZ18]. forces [AZ19, BGJ^{+}16, DPGS^{+}18, DN16, JhZbYmP15, Les15s, MSŁK^{+}18, Nel17, PBL^{+}16, Pow15e, SXE^{+}19, SWC^{+}17]. Forcible [NOS^{+}15]. forebrain [NYW^{+}17]. fork [BG18, Can19, CQB^{+}19, ZDM^{+}15]. forks [BCMM^{+}19, Les15i, RS19, RLS18a, RLS18b, TBL^{+}15]. form [BLZ^{+}15, CPCtR^{+}15, CGY^{+}19, DSA15, LPRW17, SPD^{+}17, Sho15-56]. Forman [CJ16]. formation [BSL^{+}15, BMC15, BS17b, CRPSC^{+}19, CHC^{+}18, CPB^{+}16, DPS^{+}18, DS16a, DJV^{+}16, DLZ^{+}15, FFÁTC15, GPAA^{+}18, GCZ^{+}19, GLL^{+}18b, HKG17, HM19, IYP^{+}18, ISL^{+}18, JSB^{+}18, KHA^{+}18, KBB^{+}15, KBB^{+}16, KST^{+}17a, KST^{+}17b, MSS^{+}17, MHI^{+}18, NWP^{+}16, NPU^{+}16, OKY^{+}16, ONT^{+}19, Ott16, PM15, RMMS^{+}17, SOII18, Sho15m, SLM^{+}15, SENL^{+}15, SAO^{+}17, SDP^{+}15a, SDP^{+}15b, SCL^{+}16, SAB^{+}18, TYK19, TCD^{+}15, VPD^{+}16, WPA^{+}18, WEQ^{+}15, WMB^{+}15, iYJF^{+}16]. Formin [MHY^{+}16, DBG^{+}15, GTW^{+}15, GSKL^{+}18, LM16, PAC^{+}15, WMB^{+}15, GFWG15, vGWC^{+}18]. formin-dependent [GTW^{+}15]. Formin-generated [MHY^{+}16]. Formin-like [GFWG15]. formin-mediated [GSKL^{+}18]. formin-nucleated [DBG^{+}15]. forming [JJB^{+}19, KSGL19, Kti19]. formins [DATI18]. forms [AHA^{+}19]. forward [DKMV15, Sed15s]. four [VMR^{+}19]. four-phosphate-adaptor-protein- [VMR^{+}19]. FoxO [MNLB16, OG16]. FOXO1 [Les15-29, ZPT^{+}15]. Fragile [Log17, OSW^{+}17]. framework [LS18]. Francesca [Pow15c]. Fred [Jor16d]. Fredberg [Sed15o]. Frederic [Jor16e]. free [JSB^{+}18, Sho18a, TG15, WFS15]. fresh [Les15i]. FRET [OSL^{+}19, VRM^{+}19]. fringes [Pow15b]. Fritz [O’D19h]. front [JGCAC^{+}15]. frontal [Les15r]. frontotemporal [WLM^{+}15]. Fuchs [O’D17g]. Fueling [Yel18]. fuels [KML^{+}15]. full [GKG^{+}18, Les15r, LDR^{+}19]. full-length [GKG^{+}18, LDR^{+}19]. function [AEP^{+}17, BKH^{+}15, BC19, BMC15, BCMM^{+}19, CKJ^{+}15, CQB^{+}19, CCS^{+}19, CWL^{+}16, CSYB^{+}17, CCY^{+}19, DMC^{+}16, DKM^{+}15, DB15a, DCF^{+}17, EJK^{+}16, FML^{+}17, GFvA^{+}15, GLL^{+}18b, Gen17, GGC^{+}17, GYK^{+}17, HGD^{+}15, KHRL17, KTK^{+}18, KSM^{+}18, LJP^{+}15, Les15z, LT19b, LTRW15, MCS^{+}15, RXEB^{+}19, RSvW^{+}15, Sed16a, Sho15-56, Sho16-36, TMK18, WZC^{+}15, WYV^{+}19, WF15]. functional [CKS^{+}15, CN15, NBG^{+}16, NGG^{+}16]. functionally [MSL16]. functions [ATRG19, BKG^{+}15, CD18, DSvNA^{+}15a, DSvNA^{+}15b, DMV^{+}19, FLG^{+}15, FLG^{+}19, IB19a, IB19b, MLJ^{+}16, MRM18, MGT^{+}19, NMN^{+}15, OSK^{+}15, PLD^{+}15, SSPD15, TAQ^{+}19]. Fundamental [Sch15]. fungal [VAB^{+}18]. Furin [BMC15]. furious [GB18]. furrow [FLN^{+}10, FLN^{+}16, LW16b, PUTM15]. furrows [WG16]. FUS [MCH^{+}18, YCSJ^{+}17]. fuse [BUPC19]. Fusion [BCH^{+}17, BPL^{+}18, CZZ^{+}15, CRC^{+}15, CZL^{+}15, CLO^{+}19, DOA^{+}17, DSS^{+}15, DS16b, DBG^{+}15, FR16, GRU18, GHKW^{+}19, LKM^{+}15a, MPH^{+}15, MJN^{+}18, MSW^{+}07, MSW^{+}17, MKD^{+}18, MSV^{+}19, NPU^{+}16, RHPH^{+}18, Sed15m, Sho15-44, SKL^{+}18, SHR17, VML^{+}17, WTC^{+}19, WHL17, WMH^{+}18, YSW^{+}15, vGWC^{+}18]. fusogenic [VKT^{+}15]. fusogens [VML^{+}17]. futile [AZ19]. future [SKG17, Tar15]. fuzzy [CBF^{+}18]. Fyn [FSB^{+}15]. FYVE [TCP^{+}15].

G [BSP16, CNC^{+}18, FdSR^{+}17, IdSCB^{+}16, LMPG^{+}15, MMW^{+}19, PhHS^{+}16, Sch17a, TLMG^{+}15]. G-protein [TLMG^{+}15]. G0 [Blo19]. G1 [MAK^{+}16, PKN^{+}15]. G2 [HHCK19, WV18b]. G2/M [HHCK19]. G3BP [KPA^{+}16, KPA^{+}20]. G3BP1 [TT19, ACG^{+}15, PKS^{+}19, SENL^{+}15]. G3BP1-S149 [TT19, PKS^{+}19]. GABAergic [CBAP^{+}17]. GABARAP [Mar16a, NPU^{+}16]. GABARAPL1 [SSRG18]. Gabriel [O’D19c]. Gag [HBS^{+}15, Sho15r]. Gage [Sil16b]. gain [WWW^{+}18]. galectin [KSGL19]. galectin-9 [KSGL19]. game [SG17]. gamete [VML^{+}17]. ganglion [IKRMN16, SZL^{+}16]. gap [KDM^{+}18, Sho16f, SOP^{+}16, CM18, TAQ^{+}19]. gastric [ZLG^{+}15]. gastrulation [Pow16c]. Gatekeepers [PW19]. Gatekeeping [Col18]. gating [RPMC^{+}16]. GATOR1 [MF18]. GATOR1-dependent [MF18]. Gauging [Sho15r]. GCN2 [KVK^{+}17]. GCS1 [VML^{+}17]. GEF [ANM^{+}19, DKM^{+}15, FDR^{+}16, Nie16]. GEF-H1 [FDR^{+}16]. gender [Les15d]. gene [AIK^{+}16, AIS^{+}18, CAI^{+}15, DKA^{+}16, FBBRCA^{+}18, Jor16c, JBMM16, LLS^{+}16, MSŁK^{+}18, MHI^{+}18, MN17, RHCS^{+}16, STF18, Sho16-34, TSB^{+}18, VZB19, WYHG17]. gene-regulatory [VZB19]. generate [LVF^{+}15]. generated [MHY^{+}16]. generates [GKC^{+}17, KHS^{+}16, LDR^{+}19]. generation [HB16]. generations [CSF^{+}17, CSF^{+}18]. genes [CMM^{+}15]. Genetic [MP17a, SSdLA^{+}15]. genetically [MCH^{+}18]. genetics [Pow16d]. Genome [DSC^{+}18, MHI^{+}18, SIBM17, ATH^{+}19, BPH^{+}19, CNA^{+}17, CZW^{+}18, HSN^{+}16, LUC^{+}15, NHA^{+}19, NPC17, UOT^{+}16, VZB19]. Genome-edited [DSC^{+}18]. Genome-wide [MHI^{+}18, SIBM17, BPH^{+}19]. genomes [CSF^{+}17, CSF^{+}18, SZ17b]. Genomic [LT19a, MTC^{+}19, MBR19, NKH^{+}19]. genomics [SKG17]. genotoxic [MTM^{+}17, ZDM^{+}15]. genuine [OKK^{+}15]. geography [May15]. geometry [LVG^{+}18]. Gergely [Cas16a]. germ [CAKL16, LL17, VPD^{+}16]. germline [AGL^{+}15, LLK^{+}17, LZC^{+}15, VZFG^{+}18]. get [Les15p, MB17a, RMTR17, Sho15l, Sho16q]. gets [Les15f, Les15j, SH17, Sho15x, Sho15-30, Sho15-45, Sho15-60, VR18]. Getting [KH19, NO19, O’D19g, Sed15j]. GGCX [FLG^{+}15, FLG^{+}19]. ghosts [KdBKvdK15]. Gihana [O’D19c]. Gin4 [RSvW^{+}15]. Girdin [LMPG^{+}15]. GIV [LMPG^{+}15]. GIV/Girdin [LMPG^{+}15]. give [Bea16, Les15y]. gives [KJ16, Sho15w]. gland [Sho15-50]. GlcNAc [BH15]. Glia [Sed15k, MRO^{+}15, PC17]. Glial [KO19, LRD19, RMMS^{+}17]. gliomedin [CPEE^{+}15]. Global [DTW^{+}16, GBB^{+}19, Sho16-37, ZDM^{+}15]. GLP [XMJ^{+}19]. GLP-1 [XMJ^{+}19]. GluA1 [HZH^{+}15]. glucose [HDA^{+}17]. Gluing [Sho18b]. GLUT4 [BBC^{+}16]. glutamate [FV17]. Glutamylation [MH15]. Glutathione [BS18]. glutathionylation [SAO^{+}17]. Glycan [LGH^{+}18]. glycinergic [CBAP^{+}17]. glycolysis [ALY^{+}17, Yel18]. Glycolytic [Sed15l, WDW^{+}17]. glycoprotein [NOS^{+}15, vBMG^{+}15]. glycoproteins [NOS^{+}15]. glycoproteomics [CVL^{+}19]. glycosylation [CVL^{+}19]. glycylation [GDB^{+}17]. Glypican [CIS^{+}17]. Glypican-6 [CIS^{+}17]. GMF [HAK^{+}15, Sho15s]. go [FW16, RMTR17, Sed15d]. goes [BH15, Sho15v, Sho15-39]. Going [MS19b, FC16, Les15r, MS19a]. Goley [O’D17b]. Golgi [CPBG19, CGPB17, CBM^{+}16, GNM16, IB19a, IB19b, KYN^{+}18, KOK^{+}19, LLL^{+}15, LCTP17, MSCS19, MHA^{+}19, SA19, Sed15p, VRM^{+}19, WDM^{+}15, YWdH^{+}17]. golgin [LLL^{+}15]. González [Sed15b]. good [Les15a]. GOP [SLH17, YHG^{+}17]. GOP-1 [SLH17, YHG^{+}17]. governing [CRK^{+}17]. governs [AUTM16, CKX^{+}16, MBS^{+}18, NVP17]. gp135 [SHO^{+}15-74]. gp210 [GCH15]. gp210/Nup210 [GCH15]. GPCR [AMS^{+}17, GAS^{+}15, LL17, MMW^{+}19, TCWM18, VBJ^{+}18a, VBJ^{+}18b]. GPCR-independent [MMW^{+}19]. GPCR-induced [GAS^{+}15]. GPCRs [Les16j, YNN18]. GPER1 [BLZ^{+}15]. GPI [LFT^{+}16, LGH^{+}18, Sed15d, SLAR^{+}16]. GPI-anchor [LGH^{+}18]. GPI-anchored [SLAR^{+}16]. Gpr161 [PhHS^{+}16, Sho16k]. GPR31 [FdSR^{+}17]. Gracefully [Sed15a]. Gracheva [Sed15g]. gradient [WTB^{+}19]. granular [MDC^{+}16]. granule [ACG^{+}15, CMB^{+}18, GDD^{+}15, KPA^{+}16, KPA^{+}20, NGX^{+}19, PIA16, PKS^{+}19, Sho16-31, SENL^{+}15, TT19]. granules [ATS19, ADBST^{+}15, HMC^{+}16, HCN^{+}15, JBMM16, Les15-28, NIdG^{+}18, PBL^{+}19, RM16]. great [Mar15, Sho15-39]. Greg [O’D19d]. GRHL2 [GBRH15]. grow [MOM^{+}18, Ver16]. Growing [Sed15m, Ric18]. Growth [CRK^{+}17, YPY^{+}15, BFPD19, Bro19, CIS^{+}17, CNRR^{+}17, CG16, CKX^{+}16, CHH^{+}15, DLH^{+}19, DRL^{+}19, DKM^{+}15, FTS^{+}19, FLG^{+}18, GI19, GSD^{+}15, GMTL18, HSK^{+}16, HPE^{+}19, IYP^{+}18, Jan18, JH19, KKC^{+}19, LDU^{+}16, LMdM^{+}16, Les16b, LLZ^{+}19, MBF17, MSS^{+}17, NBG^{+}16, PGMM^{+}19, PBL^{+}19, RHJW18, SAF^{+}19, Sho15d, Sho15-46, Sho17l, TNK18, WRH^{+}16, WLM^{+}15, WB18, XTT^{+}18, XLW^{+}18, ZAT^{+}19, vBMG^{+}15]. GSK3 [AZS^{+}15, ARV^{+}18, VXF^{+}15]. GSK3- [ARV^{+}18, AZS^{+}15]. GSK3B [WTSA17]. GSK3B-mediated [WTSA17]. GTP [CM18, LHY^{+}19, WMH^{+}18, Wor19]. GTP-binding [LHY^{+}19]. GTP-tubulin [Wor19]. GTPase [ALLA18, DBS18, FLS^{+}16, GGC^{+}17, GKC^{+}17, JRH^{+}16, LR18, MXV^{+}16, MF18, NVP17, OOT^{+}18, RLJ^{+}17, RS16, TF16, TJF18, WHL17, YSW^{+}15, YHG^{+}17]. GTPase-1 [OOT^{+}18]. GTPases [GWL^{+}19, HKK^{+}19, LT19b, MP17a, MF16b, OFP^{+}19, RGMM18]. GTSE1 [BRH^{+}16, Sho16i, TWD^{+}17]. guanine [ZTR^{+}17]. guanosine [MOM^{+}18]. guardian [CE16]. guidance [BJL^{+}18, DKM^{+}15, GKGK16]. guide [CSM17, GTMZ^{+}15, Sho16c, vS15]. Gulp1 [GGL^{+}19]. Günter [Tra18]. Gustavo [O’D19e]. gut [RMB^{+}18, Sho15t, Sho17g, SLG^{+}18]. gyrations [Pow16c].

H [MPMP16, KML^{+}15, LLW^{+}15]. H1 [FDR^{+}16, AIS^{+}18, ANM^{+}19, HGA^{+}17, IZBH^{+}17, MH15]. H2AX [CQB^{+}19]. H2B [EMRS^{+}18, SKW^{+}19]. H3 [YTGA16]. H4K20me2 [CR18]. Hair [PGMM^{+}19, GI19, HBWY18, LMdM^{+}16, PCM16]. Hall [Mar15]. hand [Sho15-71]. handle [NO19]. Hansenula [SKVvdK15]. HAP2 [VML^{+}17]. HAP2/GCS1 [VML^{+}17]. haploinsufficient [BFS^{+}19]. hard [Sho15-34]. Hari [Jor16f]. harm [Les17]. Haspin [YTGA16]. Haynes [Pow15a]. HDAC6 [KKC^{+}19, Van19, WTC^{+}19]. Head [DB15a, CGD^{+}18]. Head-to-tail [DB15a]. headed [TBK^{+}16]. heal [Sho15g]. healing [Les15-29, MCGM15a, MCGM15b, ZPT^{+}15]. health [VV17b]. Healthy [LM19]. heart [FTAB^{+}15, GGF^{+}19, Sho16-33]. Heat [DMC^{+}17, AB18, Can17, DBS18, OI18b]. heat-damaged [DBS18]. Hec1 [DMB^{+}18]. Hedgehog [LLK^{+}17, Sho15t, CIS^{+}17, HGD^{+}15, Les15j, TSJ^{+}15]. Heike [Jor16g]. helical [VAB^{+}18]. helicase [DBS18, DKS15, PMHB17, UDH^{+}16, VLZ15]. helicases [CNA^{+}17]. helix [CWCG19, GLC^{+}19, HGF^{+}18, TSFP^{+}15]. helix-distorting [TSFP^{+}15]. help [MB17a, MGSO^{+}18]. helper [HH18]. Helping [SLH17]. helps [Sho15-32, Sho16e, Sho16s, Sho16x, Sho17a, Sho17d, Sho17k, Sho18d]. hematopoiesis [DMD19]. hematopoietic [SCP^{+}17]. hemodynamic [VCD^{+}15]. hemorrhagic [BLO^{+}16]. heparan [HGD^{+}15]. hepatic [QJP^{+}17]. hepatocyte [SSC^{+}19]. hepatocytes [SWS^{+}19]. hepatocytic [LDM17]. hereditary [AEP^{+}17, BLO^{+}16, XTT^{+}18]. herpes [Nie19, TGK^{+}19]. herpesvirus [CSF^{+}17, CSF^{+}18]. herpesviruses [SZ17b]. Heterochromatic [DMG^{+}19, SWD^{+}19]. heterochromatin [MS19a, MS19b, NKH^{+}19]. heterodimer [MF18, dlFEvW^{+}15]. Heterodimeric [AKD^{+}17]. heterogeneity [BFPD19, MOJ16]. heterogeneous [JPD^{+}16]. heterologous [PUY^{+}19]. heterotrimeric [FKO^{+}18]. heterotypic [HNF^{+}18]. hexagonal [KKD^{+}16]. Hexokinase [ALY^{+}17]. hexose [HDA^{+}17]. HflX [DBS18]. HGF [RSC^{+}19]. HGF-induced [RSC^{+}19]. HHIP1 [HGD^{+}15]. hi [O’D16a]. Hierarchical [BMW^{+}18, KSL^{+}17]. High [BCG^{+}19, PCF^{+}19, BZG^{+}17, CDF^{+}18, DWH^{+}17a, FGR^{+}18, GPD^{+}19, Pow16d, XIZ^{+}18]. high-avidity [GPD^{+}19]. high-probability [BZG^{+}17]. High-resolution [PCF^{+}19]. High-throughput [BCG^{+}19, Pow16d]. Higher [WZR19, WGHE^{+}18]. Higher-order [WZR19]. highlights [MP17a]. highways [SS16]. hijack [Pow15h]. hinders [MWB^{+}19]. hinge [KCB^{+}16]. hinge-loop [KCB^{+}16]. HIPK2 [DCB^{+}15]. Hippo [KG15, MBF17, MpDN^{+}17, Sho15y]. Hippo/YAP [MpDN^{+}17]. hippocampal [AMS^{+}17, SVD^{+}15, SQC^{+}16]. hippocampus [BLZ^{+}15]. HIRA [LJ17b]. histamine [MPW^{+}19]. Histone [GCA^{+}17, LJ17b, MGA19, UBBSM15, AIS^{+}18, CDF^{+}18, GCVAGS^{+}18, HGA^{+}17, MH15, Pri17, RMTR17, TTC^{+}16, YTGA16]. historical [SB17]. hitch [Sho16x]. hitchhike [GSB^{+}15]. hitchhiking [SERP16]. Hitting [O’D16a]. HIV [DLBMA^{+}15, GCJ^{+}15, HBS^{+}15, Sho15-48]. HIV-1 [DLBMA^{+}15, GCJ^{+}15, HBS^{+}15, Sho15-48]. HK2 [gXNG^{+}15, gXNG^{+}16]. HMGB2 [AIK^{+}16, GG16]. hnRNP [CAI^{+}15]. HNRNPL [MYT^{+}16]. hold [Kay16, Sed15k]. holds [GG16, Sho16-28]. home [CSM17]. homeodomain [HGA^{+}17]. homeostasis [CHP^{+}17, CRK^{+}17, FTAB^{+}15, GCH15, HSB^{+}19, HCC^{+}17, KO19, LZC^{+}15, QJP^{+}17, VTG^{+}16, ZWW^{+}19]. Homeostatic [GBB^{+}19, AWS^{+}18, CL19, PNE^{+}19]. homodimer [BKG^{+}15]. homodimers [WIS^{+}17]. homogeneity [LLS^{+}16]. homologous [BLL15, LTC^{+}16, LCD^{+}17, QSZ^{+}17a, QSZ^{+}17b, VML^{+}17]. homologue [LPHH16, RAS^{+}19, ZWB^{+}19]. homologues [TF16]. homotypic [DOA^{+}17, LKM^{+}15a, NAFM^{+}17, YSW^{+}15]. Hook1 [ODH19]. Hook2 [DKR^{+}19a, DKR^{+}19b]. Hook3 [KDR^{+}19, SV16]. HOPS [Juh16]. HORMA [RC15]. Horne [Pow15j]. Horwitz [Sed15u]. host [CSM17, DV16, RNP^{+}17, SPH^{+}19, SKL^{+}18, TB16]. host-cell [TB16]. hosts [Sho15-52]. hot [BP19a, BP19b, HHT^{+}16, Sed15g, SHO^{+}15-74, WLC^{+}17]. hot-wiring [WLC^{+}17]. hotspots [Sho15a]. Hrr25 [GSD^{+}15, WDM^{+}15]. Hrr25/CK1 [GSD^{+}15, WDM^{+}15]. HRS [MBS^{+}18]. Hsc70 [DKM^{+}15, GHD^{+}17, Les15m]. HSF1 [Can17, QJP^{+}17]. Hsh155 [MTM^{+}17]. Hsp104 [OCS15]. Hsp104-Hsp110 [OCS15]. Hsp110 [OCS15]. Hsp40 [JLB^{+}18, PXN18]. Hsp40/ [PXN18]. Hsp42 [GUM^{+}18]. Hsp70 [JLB^{+}18]. Hsp72 [OSR^{+}15, Sho15w]. HSP90 [MCM^{+}17, zLSSS^{+}18, AB18]. HSP90/R2TP [MCM^{+}17]. HSV [JNW15]. HSV-1 [JNW15]. Hu [Pow15f]. hub [GYS18, KSGL19, VZFG^{+}18]. Human [NAFM^{+}17, BSK^{+}19, CS16b, CEM^{+}15, DSC^{+}18, GKG^{+}18, HV17, IZZ^{+}18, JNW15, KY15, KWB^{+}15, LBG^{+}17, LBV^{+}17, MB15, MJSB16, MWF^{+}15, NTT^{+}15, NWP^{+}16, NPC17, PTK16, PBS^{+}16, QYY^{+}16, RBC^{+}17, RDN^{+}19, SRT^{+}18, TST^{+}17, THG19, TBL^{+}15, UDH^{+}16, WHP^{+}18, WPA^{+}18, YSR^{+}18, ZDM^{+}15, ZGDS^{+}16, ZCH^{+}18]. Humanin [GTD^{+}18]. humans [SZ17a]. hunger [Jor16a]. Huntington [SS19]. hydrolases [DBG^{+}15]. hydrolysis [WMH^{+}18]. hydroxyglutarate [HGM^{+}19]. hyperactivate [HV17]. hyperexcitability [LRD19]. hypertrophy [NWW17]. hypothesis [MT19, Sho17h]. hypoxia [DZL^{+}15].

I/II [CWL^{+}16]. ICAM [CLBB15]. ICAM-1 [CLBB15]. ICMT [CAA^{+}17]. Identification [EBMW^{+}18, KBB^{+}17]. identified [BPH^{+}19, OG16, SZ17a, VRM^{+}19]. identifies [MHI^{+}18, NDL17, QPZ^{+}17, RFG19, SIBM17, WHB^{+}18]. identify [AHS^{+}15]. identity [GWL^{+}19, Ham18, MLR^{+}16, SCP^{+}15]. IFT [CHH^{+}15, FKO^{+}18]. IFT20 [SSV^{+}18]. Igaki [O’D18f]. IGF1R [XTT^{+}18]. IgG [PBG18]. II [ABGG16, CWL^{+}16, EJK^{+}16, FLN^{+}10, FLN^{+}16, FKO^{+}18, HLST19, ITN^{+}17, JGCAC^{+}15, LRS^{+}17, MSE^{+}17, NHA^{+}19, OKY^{+}16, Sho16-36, SOW^{+}17, YTGA16, ZYA^{+}17, ZAT^{+}19]. IIA [SAT^{+}17, GCVAGS^{+}18]. IIB [FB15, SRI^{+}19, SAT^{+}17, TYD^{+}15]. III [CWI^{+}19, CWL^{+}16, ISK^{+}15, JJB^{+}19, LMdM^{+}16, MHS^{+}18]. IKK [Hu15]. IKK-NF- [Hu15]. IL-1 [NNK^{+}15]. IL-6 [dVGO^{+}16]. illuminates [CGT16, MPA^{+}16]. illustrated [OSL^{+}19]. image [GSC^{+}16]. Imaging [BYMS^{+}19, CDF^{+}18, BYUJ17, BPS^{+}15, EGY^{+}19, NHA^{+}19, NIdG^{+}18, OSL^{+}19, PCF^{+}19, SK16a, SLD^{+}15, WS18, XRH^{+}18a, XRH^{+}18b]. immediate [AIS^{+}18]. Immune [Hui19, JNS^{+}19, Gek17, Inf18b, LAMACE^{+}17, MHY^{+}16, OSW^{+}17, SDI^{+}19, SAK^{+}18, TCWM18]. immunity [KJC^{+}15, O’D17c, O’D18c]. immunization [TCP^{+}18]. immunogenic [VRK^{+}17]. immunological [CBB15, NKP^{+}15, Sho15b]. impact [AEP^{+}17, BBMM^{+}16]. impacts [LTRW15, PCF^{+}19]. impair [EW17, XTT^{+}18]. Impaired [GWF17, PMP^{+}17, YBZ^{+}18]. impairing [HGG^{+}17]. impairs [DLBMA^{+}15, FWL^{+}17, ZWW^{+}19, vDMR^{+}19]. implantation [THG19]. implications [SF15, Sch17b, VZ17]. import [APK^{+}18, ATRG19, BHB^{+}18, KdBKvdK15, Mok16, RPMC^{+}16, RDN^{+}19, RBR19, WXFS17]. importance [Blu15b, MSV16, OM19]. important [GKK16a, GKK16b, GGL^{+}19, WWTF17]. Importin [CHS^{+}17, CNN^{+}17, Les17]. Importin- [CHS^{+}17]. Importin-11 [CNN^{+}17, Les17]. improvement [CSM17]. improves [CEM^{+}15, Sho15-68]. inactivates [FKW^{+}17]. inactivation [HPE^{+}19]. inactive [SHVO^{+}18]. INCENP [FTDC17, WWTF17]. INCENP/Sli15 [FTDC17]. includes [VGA^{+}15]. inclusion [BBK16, DS16a]. inclusions [YCSJ^{+}17]. incorporation [BGH18]. incorrect [Sho17e]. increase [HHH^{+}19, HGM^{+}19, PBL^{+}16]. Increased [UGHB^{+}16, CMTH^{+}15, JPC^{+}17]. increases [DOH^{+}17, LRBB15, LMC^{+}18]. increasing [WCY^{+}16a, WCY^{+}16b]. Incredibly [SZ17a]. Independent [IKRMN16, DBC^{+}15, FVF^{+}16, GKGK16, Gra16, JKD^{+}19, KML^{+}15, KJON^{+}17, MMW^{+}19, MBC^{+}19, SDHC17, WG16, ZSdO^{+}15, MP17a, PAC^{+}15]. independently [HHS^{+}16, LDMW^{+}15, MJN^{+}18, RHJW18, SPH^{+}19, iYJF^{+}16]. individual [DB15b]. induce [WFOA15]. induced [ATS19, ACG^{+}15, BSL^{+}15, CYH^{+}16, DKM^{+}15, FDR^{+}16, FVF^{+}16, FCLoS19, Gek17, GKG^{+}18, HH16, JBE^{+}17, KMRD^{+}16, KNQ^{+}19, NWW17, NLH^{+}19, RSC^{+}19, TMFR^{+}19, WXC^{+}18, XSJ18, XWZ^{+}15, ZGDS^{+}16, GAS^{+}15, HNF^{+}18, LYO15, VXF^{+}15]. inducer [ASM^{+}15]. induces [AGGSF^{+}16, BJB^{+}18, DMC^{+}17, FCB^{+}09, FCB^{+}19, KDV^{+}15, KPGG^{+}19, MOS^{+}18, NNK^{+}15, Sho15t, TTU^{+}17, WG16]. Inducible [PABM16, LCTP17, MTC17]. inducing [MWB^{+}19, THA^{+}16]. induction [SSRG18]. inequality [Sho16w]. INF2 [CJS^{+}18]. INF2-mediated [CJS^{+}18]. infected [PMW18, HGG^{+}17]. infection [DAG^{+}15, IZZ^{+}18, OBY^{+}15, SD16b]. infiltration [LWZ^{+}18]. infinite [Gar15b]. inflammasome [SK16b, dlRHM^{+}18]. inflammasomes [SK16b]. Inflammation [Sho15x, FDR^{+}16, HS16, NS18]. inflammatory [HSZ^{+}18, KKP^{+}17]. influence [PKS^{+}19, Pas16, SOII18, TT19]. influences [MN17]. influx [MWB^{+}19]. infrastructure [PC17]. ing [SJ16, FC16]. ingression [FLN^{+}10, FLN^{+}16, RCS^{+}19, SOW^{+}17]. inheritance [BCH^{+}17, CMA19, SZF^{+}15]. inhibit [FLG^{+}15, FLG^{+}19, GPAA^{+}18, SBP^{+}16]. inhibiting [BRH^{+}16, KDM^{+}18, MRGWB^{+}16, YCSJ^{+}17]. Inhibition [CSG^{+}15, DWB^{+}17, MKD^{+}18, CMMB^{+}15, CRS^{+}17, FKL^{+}18a, FKL^{+}18b, GMTL18, KKC^{+}19, LM15, LTC^{+}16, LLS^{+}16, MXV^{+}16, SZSS18, SSRG18, SID^{+}18]. inhibitions [Sho15-42]. inhibitor [CEM^{+}15, GSKL^{+}18, Les15j, MWW^{+}16, NWW17, RYS^{+}15, SAF^{+}19, SNGO16]. inhibitors [Hui19]. inhibitory [Bro16, LBV^{+}17, SCL^{+}16]. inhibits [CHZ^{+}17, GGC^{+}17, HSZ^{+}18, QZY^{+}19, SQB^{+}15, WWZ^{+}17]. initial [AWS^{+}16, BVR^{+}17, HR16, VGY^{+}17]. initiate [FBPN^{+}18, PSL^{+}17, SSR^{+}17]. initiates [GSCIL^{+}15, JNW15, TST^{+}17, WV18a]. initiation [ALLA18, BCS^{+}17, GJW^{+}17, JKA^{+}15, LRH^{+}15, LLY^{+}19]. injured [GSCIL^{+}15]. Injury [TSJ^{+}15, GCZ^{+}19, JNS^{+}19, MpDN^{+}17]. Injury-stimulated [TSJ^{+}15]. innate [KJC^{+}15, OSW^{+}17]. inner [AFT^{+}19, BPS^{+}15, BHS^{+}19, CJS^{+}18, EJK^{+}16, KJTY19, SKG^{+}16, UKHK15, WYoS17, WLJ18, WF15]. inner-membrane [WLJ18]. innovations [Pow15g]. innovator [Pow15d]. Inositol [RHC^{+}16, DZB^{+}18, NMN^{+}15]. INPP5E [DCF^{+}17]. INPP5F [NMN^{+}15]. INPP5K [DZB^{+}18]. INs [GS18]. insertion [CGY^{+}19, IZZ^{+}18, SHO^{+}18g]. insight [QYY^{+}16, SID^{+}18, YSW^{+}15]. insights [Gli17, KTK^{+}18, LT19a, PIA16, TGCO15, WHP^{+}18, vGWC^{+}18, NO19, PCM16]. instability [CNA^{+}17, CYMS^{+}19, LUC^{+}15, YYM^{+}18]. instruct [RRM^{+}17]. Insulin [KOR^{+}19, Sho16m, BBC^{+}16, FWL^{+}17, GKG^{+}18, NGX^{+}19, RHJW18, Sho18e, TMK18]. insulin-stimulated [BBC^{+}16]. integral [BhHS^{+}17]. Integrated [GSKL^{+}18]. integrates [MPW^{+}19, SIO^{+}16]. integration [GGA^{+}17, JDG16]. Integrin [Sho15y, SGF16, WWZ^{+}18, ACG^{+}17, BBSA^{+}16, BAGM17, FSB^{+}15, FVF^{+}16, GLJ^{+}17, GLC^{+}19, Les15u, LLC^{+}17, PAC^{+}15, Sed15x, SLG^{+}18, WWZ^{+}17, ZT15]. integrin- [ACG^{+}17]. Integrin-mediated [SGF16]. Integrins [JCK^{+}19, LMPG^{+}15, LS18, Pow15e, SHVO^{+}18]. integrity [CLV17, KL19, LCZ^{+}16, LM19, PSP^{+}15, RLM^{+}15, SCK^{+}19, SCK^{+}23, UOT^{+}16]. intensities [BDAW15]. intentional [Gar15b]. interact [iNLM^{+}19, PHA^{+}17]. interacting [AHS^{+}15, LRBB15, HBS^{+}15]. Interaction [FKO^{+}18, ARB^{+}19, CSG^{+}15, CKS^{+}15, DSvNA^{+}15a, DSvNA^{+}15b, DLZ^{+}15, EG19, KBT^{+}15, LHY^{+}19, LSS^{+}15, MKA^{+}17, PMRM17, PKKB17, SLW^{+}18, SCG17, Van19]. interactions [ACG^{+}15, BG18, DNMB16, FRP^{+}17, HQW15, KSM^{+}17, MBR19, QZY^{+}19, RLS18a, RLS18b, Sho18a, SEMP15, XLW^{+}18]. interactome [AR15, JSB^{+}18]. interacts [HGD^{+}15, MCH^{+}18]. intercellular [AGL^{+}15, HVH^{+}19, KTM19, SDW^{+}19, SZR^{+}15]. intercellularly [FSF^{+}15]. interchangeable [FFÁTC15]. Interchromosomal [MBR19, FMS^{+}19, RHCS^{+}16]. interdependence [VZB19]. interface [GNM16, NDC^{+}19, UFT^{+}15]. interfaces [PHA^{+}17]. interferes [MCGC^{+}15]. interferon [CHZ^{+}17, HGG^{+}17, WBL^{+}15]. Interleukin [TCP^{+}15]. Interleukin-4 [TCP^{+}15]. interlocks [MGSO^{+}18]. Intermediate [DPGS^{+}18, FC15, GDV19, LH19, LEM17]. intermediates [KMBO^{+}15]. intermembrane [MRWM18]. internal [BHB^{+}18]. internalization [CMMB^{+}15]. internalized [JPD^{+}16]. interneuron [NYW^{+}17]. internodal [EVR^{+}19]. interorganelle [HZB^{+}15, Hen19]. Interphase [LJP^{+}15, PSP^{+}15]. interplay [NP15]. Interrogating [BOL17]. intestinal [CNC^{+}18, KPEJ17, RMB^{+}18, RRM^{+}17, Sho17g, TSJ^{+}15, YLND^{+}16]. intestine [LRH^{+}15]. intra [AFXS16]. Intracellular [SZL^{+}16, CWL^{+}17, CSM17, GYK^{+}17, KOR^{+}19, MSE^{+}17, SPH^{+}19, SiYM^{+}18, SDW^{+}19, UGHB^{+}16, WGHE^{+}18]. intracentromere [LVF^{+}15]. intraflagellar [BMF^{+}18, YSM^{+}17]. intrakinetochore [MHA^{+}16]. Intramembrane [CBH^{+}15, CKS^{+}15, HESKK15a, HESKK15b]. Intranuclear [BPW15]. intravacuolar [RNP^{+}17]. Intrinsic [MHG^{+}19, MSC19, MGT^{+}19, WTB^{+}19]. Introducing [NA17]. intronic [POTZ15]. Intronless [WWW^{+}18]. Inturned [YHS^{+}15]. invadopodia [ASM^{+}15, GLL^{+}18b, RHC^{+}16, Sho15m]. invadosomes [CLO^{+}19, POE^{+}16]. invagination [MFVS18, SK16a, UBBSM15]. invasion [ACG^{+}17, CC19, DCM^{+}17, FBPN^{+}18, GML16, HHBG17, LR18, MBS^{+}18, MCCL^{+}15, MBS^{+}17, PAC^{+}15, RHC^{+}16, TB16, TYD^{+}15, ZRDP19]. involved [BPH^{+}15, IdSCB^{+}16, SLM^{+}15]. involvement [CBF^{+}18]. involves [SMOO17]. ion [KO19, UFT^{+}15, Sed15g]. iPS [MT19]. IQGAP1 [BKH^{+}15]. IRE1 [BMM^{+}19, TCP^{+}18, TSK^{+}18, TSK^{+}19]. IRGM [KJF^{+}18]. iron [DNMB16]. irreversible [DSSF^{+}15]. ischemia [LWZ^{+}18]. ischemic [ZZW^{+}19]. islets [EWL16]. ISM1 [OWW^{+}19]. isn’t [Sed15q]. isoform [CRPSC^{+}19, CBF^{+}18, KNPC16, OKY^{+}16]. isoform-specific [CBF^{+}18]. isoforms [MSE^{+}17, PMRMS17, TJMM^{+}18]. isotropic [KST^{+}19]. isotype [Kaw17, PTK16]. isotype-specific [PTK16]. itch [GD16]. itself [CSF^{+}17, CSF^{+}18]. IV [CPB^{+}16, JCK^{+}19]. Ivaska [Pow15e].

JAM [ONT^{+}19]. Jan [Pow15d]. Janus [DK17]. Janus-faced [DK17]. JCB [NA17]. Jeffrey [Sed15o]. Jeremy [Pow16c]. Jerry [OI18a]. Jessica [Sed16a]. Jim [Cas16b, Sed16b]. JIP3 [GWF17]. JIP3-dependent [GWF17]. JIP3/ [MCCL^{+}15]. JMY [HM19]. JNK [CV19, SRF19]. Job [O’D16a]. Johan [Inf18c]. Johanna [Pow15e]. join [Les15s]. joined [MYT^{+}16]. joining [LTC^{+}16]. joins [Can19, Hen19]. Jonathan [O’D17c]. Jonikas [Pow16d]. Journal [Hal15]. journey [KM18b, ZZ19, CM16]. judgments [Sho16z]. juggles [Can17]. Julia [Sed15p]. Julie [CJ16]. Jump [Sch15]. Jump-starting [Sch15]. junction [BPH^{+}18, BPW15, GPAA^{+}18, KT15a, KT15b, ONT^{+}19, SOII18, SLM^{+}15]. Junctional [GFWG15, iNLM^{+}19, CRPSC^{+}19, SWPS^{+}19, TNK18, ZAAN17]. junctions [AHA^{+}19, CPP^{+}18, CCQ^{+}18, CCLL17, DJV^{+}16, ES18, Har16, JKD^{+}19, KDM^{+}18, KLS^{+}19, LPWK15, NIS^{+}16, SOII18, TE15, TCD^{+}15, WW16]. Junjie [Pow15f]. Junk [Sed15q].

K-fiber [OSR^{+}15]. K63 [KSM^{+}18]. Kadonaga [Sed16b]. Kagan [O’D17c]. KAMPs [Sho18d]. Kap122 [ATRG19]. Kaposi [CSF^{+}17, CSF^{+}18]. Kar1 [SER^{+}15]. Karyopherins [KHRL17]. Katja [O’D16b]. Kay [CJ16]. Kazuhiro [O’D19f]. KCC2 [LSS^{+}15]. KCP2 [SCG17]. KCTD10 [KSM^{+}18]. KDM3A [YSM^{+}17]. KDM4B [Les15n, UBBSM15]. KDM5A [GCA^{+}17, Pri17]. keep [FG16, GI19, Kon17]. Keeping [O’D19a, O’D19h]. keeps [Hu15, Les16b, Les17, Sho15u, Sho15-37, Sho15-67, Sho16t, ZB18]. kept [OO18]. Keratin [CYT^{+}18, WCL^{+}18, CAI^{+}15, FC15, KBB^{+}15, KBB^{+}16]. keratinocyte [WCL^{+}18]. keratinocytes [CAI^{+}15, FC15]. kernel5 [ZWB^{+}19]. key [Gek17, GG16, QPZ^{+}17, RFO^{+}16, YGW^{+}17]. KIAA1468 [SiYM^{+}18]. kicked [Ver16]. KIF15 [MDOS19]. KIF18A [MDOS19]. Kif18b [MGW18]. KIF1B [AGB^{+}19, XTT^{+}18]. KIF1C [KDR^{+}19]. KIF27 [YBZ^{+}18]. KIF2A [WKM^{+}15]. Kif4A [TWD^{+}17]. KIF7 [YBZ^{+}18]. kill [FD18]. killer [RM16]. killers [FD18, Les15h]. killing [HMC^{+}16, SPH^{+}19, WZC^{+}15]. Kinase [AHS^{+}15, ASZ^{+}18, AFT^{+}19, BSL^{+}15, BHS^{+}19, CRN^{+}19, CB16, DBC^{+}15, DMB^{+}18, GWZ^{+}19a, GMTL18, GAS^{+}15, HLLK19, JJB^{+}19, JPF^{+}16, KGN^{+}15, LRBB15, LDMW^{+}15, MCL^{+}15, RSvW^{+}15, SS18, TWD^{+}17, TGQ^{+}17, UMC^{+}15, UMC^{+}17, WDM^{+}15, YTGA16, LLL^{+}18]. kinase-dependent [ASZ^{+}18]. kinase-independent [DBC^{+}15]. Kinase-interacting [AHS^{+}15]. Kinases [Sho15z, CKKG17, GGWL^{+}19, TS15a, YWW17]. kindles [Sho15m]. kindlin [KBT^{+}19, BBSA^{+}16, BVR^{+}17, GCC^{+}18, WWZ^{+}17, Zha19]. kindlin- [KBT^{+}19]. Kindlin-2 [BBSA^{+}16, BVR^{+}17, GCC^{+}18, WWZ^{+}17]. Kindlin-3 [Zha19]. Kinesin [MDOS19, SMK^{+}18, EG19, HMM^{+}19, KYN^{+}18, MGW18, SMF^{+}15, SID^{+}18, SNGO16, YTTH^{+}17, YCSJ^{+}17, YBZ^{+}18, MDC^{+}16, SNGO16, FKO^{+}18]. kinesin-1 [HMM^{+}19, KYN^{+}18, YCSJ^{+}17, MDC^{+}16]. kinesin-12 [SNGO16]. kinesin-14 [SMF^{+}15, YTTH^{+}17]. kinesin-4 [YBZ^{+}18]. Kinesin-5 [SNGO16]. kinesin-8 [EG19, MGW18]. Kinesin-binding [MDOS19, SMK^{+}18]. Kinesins [IBFDB18, BDLB15, CKKG17]. kinetics [CHB^{+}16]. Kinetochore [KMLG^{+}15, WF15, ASZ^{+}18, BHS^{+}19, CRZ^{+}16, DW17, DRL^{+}19, DRMW17, DMB^{+}18, DSL^{+}17, DUL^{+}19, EG19, GCL^{+}15, GHS16a, GHS16b, HAPC^{+}19, LHB^{+}18, MHSD^{+}15, NDC^{+}19, RVS^{+}19, SPGB^{+}17, SSdLA^{+}15, SD17, Sho15k, Sho16o, VGY^{+}17, WHiO^{+}19, YAHH15, ZYA^{+}17, ZGZ^{+}15, KMLG^{+}16]. kinetochore-directed [DRL^{+}19]. Kinetochore-localized [KMLG^{+}15, KMLG^{+}16]. Kinetochores [Sho15-27, APHH^{+}19, ACRM17, CGT16, GPS^{+}17, IWM^{+}16, KY15, KWB^{+}15, KD19, MHA^{+}16, MF16a, MW17, MWF^{+}15, NHCB15, Sho15-60, SRT^{+}18]. kinetoplastid [LHB^{+}18]. King [Jor16h]. kink [Les15x]. KISS [Sho15z]. kissing [MBR19]. KKT4 [LHB^{+}18]. KLC1 [JERL^{+}15]. KLP [CSC^{+}15]. KLP-7 [CSC^{+}15]. Klp3a [KEV^{+}17]. KMN [KY15]. knockout [HKK^{+}19, Pow15d]. known [Zhu17]. knows [Sed15d]. kon [LPHH16]. kon-tiki [LPHH16]. Kornblihtt [Cas17a]. Kota [O’D19g]. Kraft [Jor16a]. KRAS [FdSR^{+}17, XWZ^{+}15]. KRS [Rab17]. Kuduk [DWH^{+}17b]. Kuzbanian [DCO^{+}12, DCO^{+}16].

L [HGM^{+}19]. L-2-hydroxyglutarate [HGM^{+}19]. L1 [PST18]. L1CAM [SEMP15]. L1CAM/ [SEMP15]. L2 [IZZ^{+}18]. labeled [DSC^{+}18, CXZ^{+}18]. labor [CG16]. lack [JERL^{+}15]. Lackner [O’D18d]. laden [YPY^{+}15]. lamellipodia [JKA^{+}15, YKKB17]. Lamellipodial [Sho15-28, HAK^{+}15, KS17, THA^{+}16]. Lamin [CMM^{+}15, Nie19, OBS^{+}17, EW17]. lamina [HLW^{+}15]. lamination [IKRMN16]. laminopathic [MHW19]. laminopathies [MT19]. laminopathy [BFS^{+}19, EW17]. lamins [HLW^{+}15, TGK^{+}19]. lamp [LLL^{+}15]. LAMP1 [CXZ^{+}18]. LAMP1-labeled [CXZ^{+}18]. LAMTOR [FdAV^{+}17]. LAMTOR/ [FdAV^{+}17]. landscape [AIK^{+}16, Hyr15, LS18, Roy16]. Lane [Pow15g]. language [May15]. Lano [CTI^{+}19]. large [MSCS19, SZSS18]. Laser [CRZ^{+}16, CGT16]. late [BDW19, CZL^{+}15, DOA^{+}17, FdAV^{+}17, KNQ^{+}19, Les16f, LJS^{+}16a, LJS^{+}16b]. lateral [KD17b, NIS^{+}16]. latrophilin [AMS^{+}17]. latrophilin-2 [AMS^{+}17]. lattice [OSL^{+}19]. Laura [O’D18d]. Lava [LLL^{+}15]. layered [SEMP15]. layers [Jor16g]. Laylin [O’D19h]. Lazarou [O’D19i]. LC3 [HM19, Mar16a, NPU^{+}16]. LC3/GABARAP [Mar16a, NPU^{+}16]. LC3C [LLW^{+}17]. LCMV [DAG^{+}15]. LD [XLW^{+}18, SLPW19]. Ldb16 [GBM^{+}15]. Ldo [TJMM^{+}18]. LDs [Hen19, XLW^{+}18]. lead [ACG^{+}17]. leading [DKS15, IBG^{+}15, PMG^{+}17, O’D18e]. leads [BNS^{+}17, JERL^{+}15, PCP17, PMP^{+}17, Sho15-35, Sho16i]. learning [GLS^{+}17, O’D18b, SVD^{+}15]. leave [Les15n]. leaves [Les16h]. left [OM19]. Legionella [AKTR18]. length [AKD^{+}17, FBX^{+}15, GDB^{+}17, GJFR16, GKG^{+}18, LDM15, LRS^{+}17, LDR^{+}19, SCK^{+}19, SCK^{+}23]. Lennon [Sil16a]. Lennon-Duménil [Sil16a]. lesion [GCW^{+}16]. lesions [TSFP^{+}15]. Lessons [SG19, SK16a]. let [Sho15-51, Sho16u, LWH^{+}18]. LET-413 [LWH^{+}18]. LET-413/ [LWH^{+}18]. leukemia [CHZ^{+}17]. leukocyte [LWZ^{+}18]. leupaxin [KBT^{+}19]. level [LLW^{+}15, MN17, SBM17]. levels [Far16, GCJ^{+}15, GSM^{+}15, LJS^{+}16a, LM16, LJS^{+}16b, MMB^{+}15, NHG^{+}18, NCV^{+}16, SVD^{+}15, Sho15-68, Sho17k, SKZ^{+}18a]. Leydig [GLL^{+}18a]. LFA [CWL^{+}17, CBB15]. LFA-1 [CWL^{+}17, CBB15]. Lgl [DZL^{+}15]. Lgr5 [CNC^{+}18]. Liaisons [CG17]. Liang [O’D18a]. licenses [Ott16]. licensing [Blo19, Col18]. life [O’D17g, Sch15, Zha19, Jor16i]. lifting [Sho16d]. ligand [GKG^{+}18, HGD^{+}15, WWZ^{+}18]. ligand-induced [GKG^{+}18]. ligands [LWF^{+}15]. ligase [BHS^{+}16, CHL^{+}19, HESKK15a, HESKK15b, HSN^{+}16, LKE15, PNE^{+}19, SZE19, SvZS^{+}16, SVD^{+}15, SMA^{+}19, WXFS17, XWZ^{+}15]. ligases [GCW^{+}16, SSV^{+}18]. Light [WS18, FGR^{+}18, FML^{+}17, GDV19, HYC16, LDMW^{+}15, MFVS18, OSL^{+}19, RMOG17]. light-sheet [FGR^{+}18, OSL^{+}19]. lighting [O’D19c]. like [CNA^{+}17, DMC^{+}17, DVS^{+}17, FFG^{+}18, GFWG15, GUM^{+}18, HKM^{+}15, HR16, ISL^{+}18, JERL^{+}15, KJZ^{+}19, KGN^{+}15, KD19, LgYL^{+}18, OG16, SS19, VZ17, YSW^{+}15, ZQZ19]. Lillian [O’D19h]. LIM [BPH^{+}18]. limit [CSC^{+}15, MCD^{+}19]. Limited [SLAR^{+}16]. limiting [CTN^{+}19, LSMG18, LMdM^{+}16, dVGO^{+}16]. limits [Sho15-61, Sho16r]. LIMK1 [LZD^{+}16]. LIMK1/cofilin [LZD^{+}16]. LIMK1/cofilin-mediated [LZD^{+}16]. LIN [DMG^{+}19, Sed15q]. LIN-65 [DMG^{+}19]. lin28a [MSK^{+}19]. LINC [SJ16, CWG15, CGY^{+}19, DWH^{+}17b, LTC^{+}16, SR17b, WSP^{+}18]. LINC-ing [SJ16]. Lindquist [Bev17]. line [SHW^{+}17, OO18]. lineage [CSS^{+}18, MLR^{+}16, SRF19]. lineage-specific [CSS^{+}18]. lineages [ZGDS^{+}16]. Ling [O’D17d]. Ling-Ling [O’D17d]. link [GCVAGS^{+}18, NS18, Sho15-60]. linkages [PUY^{+}19, SZR^{+}15]. linked [EBMW^{+}18, CGBD^{+}17]. linker [AHS^{+}18, AIS^{+}18, KQM^{+}19, SERP16]. linking [ZAT^{+}17, BBW16]. links [CSO^{+}19, DN17, HPE^{+}19, OLT^{+}19, PTMP^{+}15, Qi17, SLW^{+}18, VMP16, WZC^{+}15, YHS^{+}15]. Lipid [FW16, MCGC^{+}15, SWS^{+}19, VTG^{+}16, AFO^{+}16, Boh18, CWI^{+}19, COGP15, DLH^{+}19, DJV^{+}16, EBMW^{+}18, GBK^{+}17, GY18, GLC^{+}19, GSB^{+}15, HSB^{+}19, HCC^{+}17, KTK^{+}18, KBB^{+}15, KBB^{+}16, KLHC^{+}18, KOR^{+}19, Mes16, NO19, OPP^{+}18, OKY^{+}16, PKC^{+}16, SOII18, Sho15-60, SMA^{+}19, SAB^{+}18, TJMM^{+}18, VKT^{+}15, XLW^{+}18, GBM^{+}15, NO19]. lipid-dependent [GLC^{+}19]. lipids [Kti19, TG15, VYB^{+}19]. Lipodystrophic [EW17]. lipodystrophy [OBS^{+}17]. lipodystrophy-causing [OBS^{+}17]. lipolysis [SWS^{+}19]. lipophagy [SWS^{+}19]. Lipoprotein [Pfe16]. lipotoxicity [LCTP17, TCZ^{+}16]. liquid [KKD^{+}16]. LIR [Too18]. LIS1 [QZX19]. LIS1-mediated [LM15]. Listening [Blu15a]. Listeria [SPH^{+}19, ZB19].

LITE [FGR^{+}18]. little [BH15]. Live [BPS^{+}15, EGY^{+}19, HBS^{+}15, CDF^{+}18, MSE^{+}17, PBS^{+}16, RZS^{+}15, XRH^{+}18a, XRH^{+}18b, SK16a]. Live-cell [HBS^{+}15]. lived [DRL^{+}19, TALR^{+}19]. liver [KOR^{+}19, ZSH17]. lives [Pow15c]. living [DSL^{+}17, DMH^{+}15, FWH^{+}16, FJ17, LT18, MTN^{+}16, Pow16e, SKG^{+}16, VM19]. LMNA [MT19]. lncRNA [NWW17]. lncRNAs [PST18]. Lnx1 [LZH^{+}18]. load [ACRM17]. loading [MWW^{+}16, PMHB17]. lobular [SEMP15]. Local [WG16, KJ16, WWT18]. Localization [ESS^{+}17, AOL^{+}18, GDB^{+}15, HAPC^{+}19, HGD^{+}15, HBDW^{+}15, KYN^{+}18, Lac19, LTS17, NHCB15, SHR17, THA^{+}16, WHiO^{+}19, YCSJ^{+}17, YVM18, ZRDP19, ZGZ^{+}15]. localized [BGJ^{+}16, KMLG^{+}15, KLHC^{+}18, WKW^{+}15, KMLG^{+}16, MST^{+}15]. localizes [KMK^{+}17a, KMK^{+}17b]. localizing [SKG^{+}16]. locate [ZNR^{+}18]. locations [GY18]. loci [AIK^{+}16, BLL15]. locus [OBS^{+}17, TTC^{+}16]. long [BH15, CIS^{+}17, DRL^{+}19, FTAB^{+}15, Les15a, Les16e, McM19, MS19a, MS19b, TALR^{+}19]. long-lived [DRL^{+}19, TALR^{+}19]. long-range [MS19a, MS19b]. long-term [FTAB^{+}15]. longevity [LCZ^{+}16, LTRW15, TM18]. look [Jor16f, NA17, Ric18, SA19]. loop [DKMV15, Gar15b, HBM^{+}19, Hu15, KCB^{+}16, LRM^{+}19, Les15x, WHB^{+}18, CNA^{+}17, WZC^{+}15]. loops [LVF^{+}15, Sho15-65]. loose [NHA^{+}19]. loses [O’D19c]. Loss [EKP^{+}19, KL19, KMRD^{+}16, LCM^{+}16, RMS^{+}18, ARV^{+}18, DKA^{+}16, GHKW^{+}19, LDU^{+}16, LSMZ^{+}18, LMC^{+}18, MAK^{+}16, PCP17, PSP^{+}15, RCS^{+}19, Sho15-34, SOW^{+}17, TTU^{+}17]. love [MBR19]. loves [Wor19]. Loving [O’D18a]. low [FGR^{+}18]. Lrp4 [XJG^{+}17]. Lrrk [LLL^{+}15]. LRRK2 [GWL^{+}19]. LSR [Sho15-29, SLM^{+}15]. LSR/ [SLM^{+}15]. Ltc1 [MST^{+}15, Sho15-30]. LTK [CRN^{+}19]. LTP [BNB^{+}15]. LTP-triggered [BNB^{+}15]. Ltv1 [CGD^{+}18, GSD^{+}15]. lumenogenesis [KJZ^{+}19, RMOG17]. luminal [JOJG16]. lung [Sho17e]. Lymphatic [BJL^{+}18]. lymphocyte [MPH^{+}15]. Lymphocytes [Sho16n, HSK^{+}16, MCGC^{+}15, PLD17]. lymphoid [SLG^{+}18]. lysine [LH19, ZWW^{+}19]. lysosomal [CXZ^{+}18, CJ17, HPW^{+}17, KDM^{+}18, LKE15, LCZ^{+}16, LE16, LTB^{+}17, OSK^{+}15, STR^{+}18, WZG^{+}17, ZSdO^{+}15]. Lysosome [TCZ^{+}16, AEP^{+}17, DS16b, EEE^{+}16, GWZ^{+}19a, GWF17, JJW17, LCZ^{+}16, LZ16, MSV^{+}19, NPU^{+}16, PKKB17, SDI^{+}19, Sho15-44, SE18, TCP^{+}15, WZG^{+}17, WTC^{+}19]. Lysosomes [GF16, Kon17, BMM^{+}19, GSS^{+}17, GX16, JOJG16, KM18b, MAJ^{+}17, MJN^{+}18, MF18, MSV^{+}19, SZL^{+}16, Sho15-47, VZ17, VBL^{+}18, YDM^{+}18, ZZ19, Zhu17]. lysyl [KKP^{+}17]. lysyl-tRNA [KKP^{+}17]. lytic [HMC^{+}16].

M [HHCK19]. MACF1 [OLT^{+}19]. Machine [GLS^{+}17, SSC^{+}19, WTB^{+}19]. machineries [EMRS^{+}18]. machinery [DW17, MSvO17, OCS15, PA19, PSCS16]. machines [Sil17]. macroautophagy [WDM^{+}15]. macroH2A [SWD^{+}19]. Macrophages [ED17, CKM^{+}16, CHC^{+}18, HGG^{+}17, LBG^{+}17, LBV^{+}17, YSR^{+}18]. macropinocytosis [PPK^{+}16]. Macropinosome [CHC^{+}18, DS16b]. macropinosomes [YPY^{+}15]. MAD [ZY16]. MAD1 [APHH^{+}19]. MAD1-dependent [APHH^{+}19]. Mad2 [NHCB15]. Maeshima [O’D19f]. magic [Sho17f]. Mahak [Inf19a]. maintain [Can19, CSF^{+}17, CSF^{+}18, ES18, EMRS^{+}18, KNQ^{+}19, MGT^{+}19, MMB^{+}15, OLT^{+}19, WRGB^{+}15]. maintaining [CMM^{+}15]. maintains [BDZ^{+}15, CNN^{+}17, CHP^{+}17, DER^{+}18, FBX^{+}15, GBB^{+}19, HSB^{+}19, KDM^{+}18, LCZ^{+}16, MLR^{+}16, RDH^{+}19, SHO^{+}18g, SCP^{+}15, ZLZD16]. maintenance [HKG^{+}18, HCC^{+}17, NIN^{+}19, RHPH^{+}18, Sed15r, SSR^{+}17, WFS15]. Maize [ZWB^{+}19]. major [UKHK15]. make [MS19a, MS19b]. makes [Kel16, Les15-32]. Making [Les16f, AS17, BOL17, GY18, Mar15, MW17]. male [VPD^{+}16]. malignancies [HOH^{+}16]. malignant [DMD19]. Mallik [Sil17]. Mammalian [KD19, MXS17, BPH^{+}19, BPS^{+}15, BCS^{+}17, DSA15, HGL^{+}17, Hyr15, KJF^{+}18, NOS^{+}15, PIA16, YLW^{+}15, ZZMC^{+}15]. mammals [DCO^{+}12, DCO^{+}16, QPZ^{+}17, YYM^{+}18]. Mammary [CSS^{+}18, FG15, RSCR15, Sho15-50, TZC^{+}15]. Man [LSMG18]. manage [Sho17k]. Management [Yel18]. manages [Sch19]. manner [FVF^{+}16, ITN^{+}17, KD19, MBC^{+}19, SJL^{+}19, TBK^{+}16]. Mannose [Les15o, OSK^{+}15]. map [WHP^{+}18, GAS^{+}15]. MAP7 [HMM^{+}19]. MAPK [AS17, DKA^{+}16, HSZ^{+}18, vDMR^{+}19, KDM^{+}18]. MAPK-dependent [DKA^{+}16]. Mapping [CZW^{+}18, Sho15-31, Sho16o]. MAPs [SID^{+}18]. MARCH1 [OPP^{+}18]. Marcos [Blu15a]. Margot [O’D17e]. Maria [Sil16a]. mark [Les15n]. MARK2 [ZCH^{+}18]. MARK2-mediated [ZCH^{+}18]. marked [EGY^{+}19]. markers [DSC^{+}18]. marks [NGG^{+}16, RMTR17]. marries [Wor19]. Martin [Pow16d]. mass [SKG17, SKO^{+}15, ZZMC^{+}15]. Mast [CPCtR^{+}15, Les15p, MDC^{+}16]. master [BP19c]. Mastering [Pow16e]. materials [Jor16b]. Mating [WTB^{+}19, ADBST^{+}15, ML15b, MSW^{+}07, MSW^{+}17, MKD^{+}18]. matrices [ECAB^{+}16]. matricryptins [SCL^{+}16]. matrix [DCM^{+}17, DN17, FVF^{+}16, GML16, KdBKvdK15, LCM^{+}16, MB15, SSL^{+}17, WSDY17, WXFS17, WCL^{+}18]. matrix-bound [FVF^{+}16]. matrix-degrading [MB15]. matter [BG19, Gra16]. Matthew [Sed15r]. Maturation [CPBG19, BPH^{+}15, CW17, CBB15, DS16b, DLBMA^{+}15, FC16, KT15a, KT15b, KSGL19, LZH^{+}18, LLL^{+}18, MCD^{+}19, SSH^{+}15, SSR^{+}17, VRK^{+}17, WQD^{+}18, ZZ19, vHGD^{+}15]. Maturation-driven [CPBG19]. mature [CST^{+}17, Nie19, NWD^{+}19]. matures [Sho17h]. Maya [Sed16c]. MCAK [BRH^{+}16]. MCL1 [WTSA17]. McLaughlin [IO18]. Mcp1 [PHA^{+}17]. Mcp5 [CMA19]. mDia1 [QYC^{+}17]. mDia2 [LM16]. Mdm1 [HSB^{+}19, HZB^{+}15, Sho15-32]. Mdm1/ [HZB^{+}15]. Mdm12 [KTK^{+}18]. Mdm2 [CBF^{+}18, KR18]. Mdm35 [AFO^{+}16, MWT^{+}16]. measure [PCM16]. measured [SPJ^{+}15]. measurements [SKO^{+}15, ZZMC^{+}15]. Measuring [Nel17]. Mecak [Spe17b]. Mechanical [LDG18, vDMR^{+}19, DK16, Nel17, PBL^{+}16, WXC^{+}18]. mechanics [CPP^{+}18, Cas16b, CC19, FK17, HF15, Pow16e, TB16]. Mechanism [KJF^{+}18, SKZ^{+}18b, ASZ^{+}18, BBS^{+}17, CHH^{+}15, DSA15, GPS^{+}17, JKD^{+}19, LHA^{+}15, LPHH16, MRM18, OI18b, PHKY17, SM18, ZCH^{+}18]. mechanisms [BPSK^{+}16, CLH^{+}18, Gar15a, HKH16, JDG16, KY15, VGY^{+}17, VY18, XS16, YEM^{+}19, SK16b]. Mechanistic [PIA16]. mechano [GWZ^{+}19b, XPZ^{+}19]. mechano-regulated [XPZ^{+}19]. mechano-regulation [GWZ^{+}19b]. mechanocatalysis [ZAT^{+}19]. mechanopathology [MHW19]. mechanoprotection [LNH^{+}15]. mechanosensation [GJW^{+}17]. Mechanosensing [CID17, PPR^{+}19]. mechanosensitive [GYS18, SRI^{+}19]. mechanosensitivity [KOV^{+}16a, KOV^{+}16b]. mechanosensory [CBH^{+}15, KDA^{+}18]. mechanotransducer [TSB^{+}18]. Mechanotransduction [BBHBFSF18, WSP^{+}18, BLO^{+}16, CSO^{+}19, FSB^{+}15, GAS^{+}18, Nie16, SGF16]. Meckel [RDO^{+}15]. medial [SWC^{+}17]. mediate [ABGG16, CCH^{+}17, GKGK16, KPA^{+}16, KPA^{+}20, KJON^{+}17, MAK^{+}16, RHH^{+}18, SCG17]. mediated [ATH^{+}19, ARV^{+}18, BDZ^{+}15, BLPV^{+}17, CJS^{+}18, CHI^{+}15, DSC^{+}18, DGS^{+}18, DLT^{+}18, FML^{+}17, FTS^{+}19, FWH^{+}16, FVF^{+}16, FCB^{+}09, FCB^{+}19, GDL^{+}15, GSKL^{+}18, GTD^{+}18, GFWG15, HLW^{+}15, HPE^{+}19, IM16, JCF^{+}17, JJW17, JPF^{+}16, KSL^{+}17, KJC^{+}15, KQM^{+}19, LKM^{+}15a, MBS^{+}18, MWB^{+}19, NDL17, PhHS^{+}16, PD19, QZX19, RHCS^{+}16, SZE19, SKW^{+}19, SD19, SWC^{+}17, SGF16, TGQ^{+}17, VTG^{+}16, WTSA17, WWY^{+}18, WHiO^{+}19, WLC^{+}17, gXNG^{+}15, gXNG^{+}16, vGWC^{+}18, AWL18, ANM^{+}19, DKMV15, HGA^{+}17, JHF^{+}15, JIB^{+}19, JNS^{+}19, KML^{+}15, LM15, LHT^{+}19, LZD^{+}16, MvVV^{+}16, PBL^{+}16, VMP16, WZC^{+}15, WXFS17, YKKB17, YSR^{+}18, ZDM^{+}15, ZCH^{+}18, vDMR^{+}19]. mediates [AMS^{+}17, AIS^{+}18, BLO^{+}16, CRPSC^{+}19, CDT^{+}19, CMMB^{+}15, DQB^{+}16, DZL^{+}15, FCLoS19, GKK16a, GKK16b, GCW^{+}16, GWZ^{+}19b, HZH^{+}15, KKC^{+}19, KBT^{+}15, KST^{+}19, LDP^{+}15, MCH^{+}18, MBT16, NNH17, SBR^{+}15, SZSS18, SZL^{+}16, SDHC17, WYoS17, WB18]. mediators [SPH^{+}19]. meets [GGR15]. Megakaryocyte [NS15, NNK^{+}15]. Megan [Jor16h]. meiosis [BPSK^{+}16, BCM^{+}18, HHH^{+}19, KBKW19, Lac19, LSJY15, LTC^{+}18, LWZ^{+}19, MSŁK^{+}18, MGSO^{+}18, SJJ^{+}19, ZYA^{+}17]. Meiotic [BTV16, BNKB15, CSC^{+}15, CO19, Das17, DRMW17, FFÁTC15, HHCK19, KHA^{+}18, PTR^{+}19, PMRM17, PBG^{+}15, RO18, SCNTC^{+}18, YAHH15]. MEIS [HGA^{+}17]. melanogaster [FLG^{+}18, POTZ15, RGR^{+}18]. melanoma [vHGD^{+}15]. melanosomal [RHH^{+}18]. melanosome [FC16]. melanosomes [DMS^{+}15, DDAR^{+}16]. members [POE^{+}16, YTTH^{+}17]. Membrane [GWL^{+}19, JDG16, KFAMR17, LKE15, LBV^{+}17, PMG^{+}17, SHR17, SE18, AFO^{+}16, AFXS16, AWS^{+}16, AKTR18, BhHS^{+}17, BDK^{+}18, BGKL15, BPS^{+}15, BVR^{+}17, BJO^{+}16, CWCG19, CZZ^{+}15, CJS^{+}18, CC19, CMTH^{+}15, CSA19, CCH^{+}17, CPB^{+}16, DWH^{+}17a, DQB^{+}16, Dic17, DSA15, DWH^{+}17b, DS16b, DZL^{+}15, EEE^{+}16, FdSR^{+}17, GDD^{+}15, GFH^{+}16, GCJ^{+}15, GPD^{+}19, Ham18, HHBG17, iHMM^{+}17, IZZ^{+}18, IM16, JHC^{+}16, KJTY19, KBB^{+}17, LSPC16, Les16d, LXR^{+}15, LCZ^{+}16, LMM16, LDMW^{+}15, LDG^{+}15, MPH^{+}15, MSC19, McM19, MSE^{+}17, MCGC^{+}15, MHA^{+}19, MFP17, MGE^{+}15, MYN^{+}17, MP17b, NIS^{+}16, PLS^{+}15, PD19, PH18, PKC^{+}16, RS16, RLM^{+}15, RBM^{+}19, RHH^{+}18, RSvW^{+}15, SZE19, SSL^{+}17, SSC^{+}19, Sed15w, SOII18, Sho15-57, SKG^{+}16, SZK^{+}19, SKZ^{+}18a, SCL^{+}19, TBJ^{+}17, TG15, UKHK15, VKT^{+}15, WYoS17, WLJ18, Yud19, KS17]. Membrane-anchored [LKE15]. membrane-bending [TBJ^{+}17]. membrane-binding [GFH^{+}16]. membrane-shaping [JHC^{+}16]. membranes [GYK^{+}17, JCK^{+}19, KTK^{+}18, LTG^{+}18, NHG^{+}18, Pow15h, Pow15i, SKN19, WIS^{+}17]. Memoriam [AD18]. memory [BSL^{+}15, EWL16, GML16, OBY^{+}15, Sho15-68]. Mena [RHC^{+}16]. menin [XMJ^{+}19]. mentor [Bev17]. mere [Sch17a]. merge [BHS18]. Merkel [WRGB^{+}15]. Merlin [CMMB^{+}15, MpDN^{+}17, Sho15-33, Sho17f]. merotelic [CGT16]. mESC [CSYB^{+}17]. mesenchymal [GCC^{+}18, SXT16, VWM^{+}18, ZDSM^{+}18]. Mesp1 [CLL^{+}16, Kel16]. message [Col19, Les15f, Sho15n]. messenger [CS16b]. MET-2 [DMG^{+}19]. MET-2/SETDB1 [DMG^{+}19]. Metabolic [MC16, STF18, Can17, Les15-27, gXNG^{+}15, gXNG^{+}16, YKO^{+}16]. metabolically [TJMM^{+}18]. metabolism [BS18, CSO^{+}19, DN17, LZ16, MSV16, O’D18c, RGOS^{+}16, SPMM^{+}17, Yel18]. Metalloprotease [SAF^{+}19]. Metalloproteinase [EWL16, GML16]. metaphase [CO19, DMB^{+}18, FBX^{+}15, LWZ^{+}19, OM19]. metastasis [DCM^{+}17, HOH^{+}16, Les15-28, Lin15, QCC^{+}19]. metastatic [IBG^{+}15]. metazoans [RGMM18]. Methionine [LHT^{+}19]. method [AHS^{+}15]. methods [JW19]. methylation [CAA^{+}17]. mevalonate [HOH^{+}16]. Mex67 [DMV^{+}19]. Mex67p [SLD^{+}15]. Mff [CRC^{+}15]. Mff-deficient [CRC^{+}15]. MglA [TLMG^{+}15]. MHCII [OPP^{+}18]. Mia2 [SNOBM16]. Mia2/cTAGE5 [SNOBM16]. Mic60 [TBJ^{+}17]. mice [FWL^{+}17, GHKW^{+}19, WRGB^{+}15]. Michael [O’D19i]. MICOS [AFO^{+}16, TBJ^{+}17]. microautophagy [MOS^{+}18, OMK^{+}17]. microbiome [Gar15b]. Microcephaly [IG15, KMC^{+}19]. microchannel [SKO^{+}15]. microcluster [MHY^{+}16]. microexon [RSC^{+}19]. microexon-spliced [RSC^{+}19]. Microglia [CBAP^{+}17, HHS18, Pow15c, WYHG17]. microglial [HSZ^{+}18]. micromanagers [ED17]. Micromanaging [Pow15c]. micrometer [CWCG19]. micrometer-scale [CWCG19]. micron [LMM16, BJO^{+}16]. micron-scale [LMM16, BJO^{+}16]. Micronuclei [Gek17]. Micronucleophagy [VR18]. MicroRNA [ZLG^{+}15, CRS^{+}17, HBWY18]. microRNA-200 [HBWY18]. MicroRNA-7 [ZLG^{+}15]. MicroRNA-7/ [ZLG^{+}15]. microRNA-dependent [CRS^{+}17]. MicroRNAs [PPK^{+}16, VZFG^{+}18]. MicroRNAs-103 [PPK^{+}16]. MicroRNAs-103/ [PPK^{+}16]. microscope [MHH18, O’D19f]. microscopy [BCG^{+}19, FGR^{+}18, HYC16, JW19, LW17, PABM16, PUY^{+}19, SBM^{+}19, SAB^{+}18, WMK^{+}16]. microsurgery [CRZ^{+}16]. Microtubule [BP19b, Con16, JERL^{+}15, JNW15, MGW18, MRM18, PSL^{+}17, SMF^{+}15, SFA^{+}19, ZAT^{+}17, AHS^{+}18, ASZ^{+}18, AATP17, Ava18, ANM^{+}19, BRH^{+}16, BMF^{+}18, Bro19, BS17b, CKX^{+}16, DMB^{+}18, DUL^{+}19, DLBMA^{+}15, EG19, EFM17, FFG^{+}18, FTS^{+}19, FK17, FBX^{+}15, GCL^{+}15, JBE^{+}17, JIB^{+}19, KDR^{+}19, KNPC16, KZW^{+}18, LCP^{+}15, LEM17, cLNF^{+}16, LNS^{+}19, LHB^{+}18, MNLB16, MG16, NLS^{+}18, OG16, OLT^{+}19, Ric18, RO18, RVS^{+}19, SXE^{+}19, SID^{+}18, Sle16, TKM16, TWD^{+}17, UFT^{+}15, WKM^{+}15, Wor19, WB18, YTTH^{+}17, YWdH^{+}17, YCSJ^{+}17, YVM18, YIT15, ZYA^{+}17, vdVFM^{+}17, BP19a, Sed16a]. microtubule- [ANM^{+}19, OLT^{+}19]. microtubule-associated [LEM17]. microtubule-based [KDR^{+}19]. microtubule-dependent [DLBMA^{+}15]. microtubule-induced [JBE^{+}17]. Microtubule-organizing [Sed16a]. microtubule-proximal [DUL^{+}19]. Microtubule-severing [MRM18]. Microtubules [KTM19, MOM^{+}18, SRT^{+}18, BNS^{+}17, BCM^{+}18, DRL^{+}19, FKG^{+}19, HGL^{+}17, IG15, KEV^{+}17, KD19, LVG^{+}18, MSV^{+}19, QYC^{+}17, RFO^{+}16, SMK^{+}18, SKZ^{+}18b, VGY^{+}17, Ver18, VGA^{+}15, WWTF17]. microvilli [LMdM^{+}16]. Mid1 [VMP16]. MiD49 [OMKM16]. MiD49/51 [OMKM16]. Midbody [Ott16, BRACA^{+}16, PP19]. midzone [IBFDB18, LDG^{+}15, PCF^{+}19]. Mierzwa [Inf18a]. migrating [Haw18, IBG^{+}15, MB17a, WKM^{+}15]. migration [BDAW15, BNS^{+}17, BBMM^{+}16, Bro16, BJL^{+}18, CAKL16, Cas16b, CLL^{+}16, DATI18, DPGS^{+}18, EAW^{+}17, FDR^{+}16, GGC^{+}17, GDV19, GAS^{+}18, GTMZ^{+}15, HKH16, JBE^{+}17, Kaw17, Kay16, KMJ^{+}18, LRM^{+}19, LR18, LL17, MBS^{+}17, MPMP16, NW19, NYW^{+}17, PVP^{+}19, PHKY17, PSL^{+}17, PBL^{+}16, PPR^{+}19, RSC^{+}19, SM16, SBC^{+}16a, SBC^{+}16b, SYK^{+}17, SW18, SMN^{+}16, TG17, WCL^{+}18, XPZ^{+}19, YEM^{+}19, ZTR^{+}17]. migratory [JhZbYmP15, JGCAC^{+}15]. Miller [O’D17a]. mimics [SKL^{+}18]. mind [Sho16f]. miniature [MTC17]. minimal [DS16a, FTDC17]. minimalist [Pow15i]. minimally [LSMZ^{+}18]. minus [Bro19, FTS^{+}19, IG15, NLS^{+}18, SMF^{+}15, SFA^{+}19, Wor19, YTTH^{+}17, YWdH^{+}17, YIT15]. minus-end [SFA^{+}19]. Mio [PTMP^{+}15]. miR [HZH^{+}15, LCM^{+}16, Sho15-34]. miR-203 [LCM^{+}16]. miR-501-3p [HZH^{+}15]. miR-7 [Sho15-34]. MIR335 [OBS^{+}17]. miRNA [KDA^{+}18]. Miro1 [OOT^{+}18, KKC^{+}19, Van19]. Misdirection [Sho15-35]. misfolded [BA18, GUM^{+}18, NOS^{+}15]. misfolding [BBK16, DS16a, NPC17]. mishap [ML15b]. Mismatch [Sho15-36]. mist [Hyr15]. mistargeted [SZE19]. MiT/ [NWFY15]. mito [MPA^{+}16]. Mitochondria [KL17, KBJ16, Sed15s, Sho17g, AFO^{+}16, BLG^{+}15, BJB^{+}18, BCH^{+}17, CMA19, DNMB16, FKW^{+}17, KMRD^{+}16, Les15b, MWT^{+}16, PHA^{+}17, PKC^{+}16, PSCS16, RXEB^{+}19, Sho15l, Sho17a, SK18b, WLJ16, ZZW^{+}19, MST^{+}15, PHA^{+}17, RGOS^{+}16, SJL^{+}19]. mitochondria-associated [WLJ16]. Mitochondria-driven [KL17]. Mitochondrial [BPW^{+}17, FR16, HGM^{+}19, JBMM16, Mes16, Mok16, Sho16p, TTU^{+}17, WLJ18, iYJF^{+}16, BHB^{+}18, BJB^{+}18, CJS^{+}18, CRC^{+}15, CCS^{+}19, CRS^{+}17, DWB^{+}17, GHKW^{+}19, HPE^{+}19, JLB^{+}18, Juh16, KKC^{+}19, KBB^{+}17, KBB^{+}15, KBB^{+}16, Lac19, LPWK15, LH19, LgYL^{+}18, LXR^{+}15, LTRW15, MRWM18, MLMF16, MPA^{+}16, MC16, MMB^{+}15, NS18, OI18a, OOT^{+}18, OMKM16, Pow15a, QYY^{+}16, Qi17, QPZ^{+}17, RPMC^{+}16, RLM^{+}15, SJJ^{+}19, SPMM^{+}17, SSL^{+}17, SG18a, SG18b, SK18b, VMP16, VKT^{+}15, WYV^{+}19, WEQ^{+}15, YKO^{+}16, ZYL^{+}16, ZWW^{+}19, ZGDS^{+}16]. mitochondrial-derived [Juh16]. Mitofusin [MMB^{+}15, Sho15-37, FR16, MRWM18, QYY^{+}16]. mitofusins [GS18]. mitophagy [BPW^{+}17, Gra16, LSMZ^{+}18, MPA^{+}16, NWFY15, NPU^{+}16, PSCS16, iYJF^{+}16]. Mitosis [ZS15, BG19, Bra16, EJK^{+}16, EGY^{+}19, GMTL18, HLEM^{+}18, LDU^{+}16, LVF^{+}15, LK17, Les15-32, MHA^{+}16, MBG^{+}18b, Nil19, PKN^{+}15, PSP^{+}15, SWD^{+}19, SPWM15, SKO^{+}15, SHC^{+}18, YTGA16, ZGZ^{+}15, ZZMC^{+}15]. Mitotic [BCMM^{+}19, FMS^{+}19, GHS16b, LDM15, Les15q, PS16, Sho16q, SMOO17, BHS^{+}16, BC19, BKR^{+}19, BS17b, CTS^{+}18, CANG^{+}17, CHS^{+}17, CHB^{+}16, DKR^{+}19a, DKR^{+}19b, FC19, HK15, KNPC16, KY15, LL19, Les15c, LLW^{+}15, LDG^{+}15, ML15a, MAK^{+}16, MWF^{+}15, OSR^{+}15, OM19, PDZ18, PTMP^{+}15, SLW^{+}18, SPK^{+}18, SKW^{+}19, Sho16-36, Sho18c, TNP^{+}15, TG19, WHP^{+}18, WZC^{+}15, WHiO^{+}19, WWTF17, YLW^{+}15, GHS16a]. mixture [O’D17g]. MKLP1 [MNLB16]. MLCK [Les15r]. MLKL [SPH^{+}19, ZB19]. MLL5 [Sho16r, ZLZD16]. Mmm1 [KTK^{+}18]. MMP [EWL16, Lin15, MCCL^{+}15, MHA^{+}19, QCC^{+}19]. MMSET [CR18]. MMSET-catalyzed [CR18]. mobile [DMV^{+}19, GSKL^{+}18]. mobility [CMMB^{+}15, CLBB15, RZS^{+}15]. model [ASPY^{+}16, BFS^{+}19, BYUJ17, CKS^{+}15, DV16, FGR^{+}18, LL19, NWP^{+}16, OSW^{+}17, Pug15, War15, WLM^{+}15]. modeling [BPS^{+}15, THG19]. models [HTLG18, MNL^{+}16]. moderation [NW19]. modes [IKRMN16, JCK^{+}19]. modifications [AZS^{+}15, LJ16, Pri17, STF18]. Modifiers [HCN^{+}15]. modifying [PLH18, Sed16e]. modular [CED^{+}15]. modularity [FLS^{+}16]. modulate [DBC^{+}15, vBMG^{+}15]. modulated [DNMB16, DUL^{+}19, LK17, VLP^{+}15]. modulates [BMS^{+}17, CCS^{+}19, DWH^{+}17b, EMB^{+}15, EPF16, FTAB^{+}15, GDB^{+}15, LFT^{+}16, LHY^{+}19, LKM^{+}15b, MCS^{+}15, MH15, SSV^{+}18, SAT^{+}17, WHiO^{+}19, YYZ^{+}15, YAHH15, ZYA^{+}17]. Modulating [PVP18, XSJ18]. Modulation [HHS^{+}16, vLvdKR18, PAM^{+}16]. modulator [DSvNA^{+}15a, DSvNA^{+}15b, RGOS^{+}16]. modulatory [VBJ^{+}18a, VBJ^{+}18b]. module [GGA^{+}17]. Molecular [BYUJ17, CLH^{+}18, DSSF^{+}15, GPS^{+}17, LPHH16, VGY^{+}17, VRM^{+}19, BPS^{+}15, CSS^{+}18, CST^{+}16, CO19, FCLoS19, GFH^{+}16, Gar15a, HGF^{+}18, HB18, JDG16, MRM18, MKA^{+}17, SPWM15, Sil17, SGB^{+}17, Tra18]. molecule [CPEE^{+}15, MSvO17, NIdG^{+}18, SPJ^{+}15]. molecules [Bea16, JPD^{+}16, LM19, MC15, Sho15-53]. monitor [BNKB15]. monoclonal [Ewe18]. monocyte [NLH^{+}19]. monomer [BKG^{+}15]. monomers [LMR^{+}17, Sch17a]. mononucleate [CV19]. monoubiquitinated [EMRS^{+}18]. Moreno [O’D18g]. morphine [CYH^{+}16, Sho16l]. morphine-induced [CYH^{+}16]. morphogenesis [AUTM16, BPH^{+}18, CRA^{+}19, DKMV15, ECAB^{+}16, GGF^{+}19, HBWY18, KS19, KQM^{+}19, MOJ16, MXV^{+}16, MG17, OWW^{+}19, PMRMS17, Pow15j, Pug15, RBZ18, SK18a, TST^{+}17, WSDY17]. morphology [BFPD19, HCS^{+}18, SBM^{+}19, TBJ^{+}17, YEM^{+}19, YWdH^{+}17]. Mosaic [Bea16, KHS^{+}16, SK19]. mosaicism [TALR^{+}19]. mossy [LZH^{+}18]. most [Sed16d, SRT^{+}18]. Mother [Ver16, BPSK^{+}16]. Mothers [Sed15b]. motif [CGBD^{+}17, JGCAC^{+}15]. motile [GSB^{+}15]. motility [BFPD19, EMB^{+}15, FLLM17, GHKW^{+}19, HBDW^{+}15, HV17, LM15, LE16, MCD^{+}19, MWB^{+}19, NIS^{+}16, SAT^{+}17, THA^{+}16, TB16, TG17, UFT^{+}15, YBZ^{+}18, ZDSM^{+}18]. motion [AWS^{+}16, JPD^{+}16, NTT^{+}15, TKM16]. motoneurons [MNLB16, MSS^{+}17]. Motor [FFG^{+}18, MGW18, YVM18, zLSSS^{+}18]. motors [Ava18, JERL^{+}15, KDR^{+}19, LEM17, SMF^{+}15, Sil17, YBZ^{+}18]. mounts [Sho15-65]. Mouse [PBG^{+}15, CSG^{+}15, PBG18, SHC^{+}18, SFZ^{+}17]. move [Jor16i, Kel16, LE16, SERP16]. movement [ABPS17, HLST19, MGSO^{+}18, MHY^{+}16, PCF^{+}19, SHW^{+}17, SJ16]. Moving [Gar15a, SR17b]. MOZART1 [cLNF^{+}16]. MPR [SDHC17]. MPS1 [HAPC^{+}19, HBM^{+}19, IWM^{+}16, MF16a]. Mps3 [KJTY19]. MRCK [GSP^{+}18]. MreB [TLMG^{+}15]. MRN [ABGG16]. mRNA [ALY^{+}17, ADBST^{+}15, ACG^{+}15, BMM^{+}19, BYMS^{+}19, BMS^{+}17, EMRS^{+}18, HCN^{+}15, NCV^{+}16, PBL^{+}19, PMP^{+}17, SBR^{+}15, SPJ^{+}15, SLD^{+}15, TTC^{+}16, XSJ18]. mRNAs [KNL^{+}17, PH16, SPMM^{+}17, Sed15n, Sho15-38, WWW^{+}18]. mRNP [KP18]. mRNPs [BPW15]. MRTF [FBBRCA^{+}18, HHBG17]. Msd1 [YIT15]. Msi [YLND^{+}16]. Msps [CKX^{+}16]. Msps-dependent [CKX^{+}16]. MST [TS15a]. Msx1 [VCD^{+}15]. MT [HLLK19, VM19]. MT1 [EWL16, Lin15, MCCL^{+}15, MHA^{+}19, QCC^{+}19]. MT1-MMP [EWL16, Lin15, MCCL^{+}15, MHA^{+}19, QCC^{+}19]. MTCBP [QCC^{+}19]. MTCBP-1 [QCC^{+}19]. mtDNA [MTGG18]. mTOR [BLG^{+}15, CIK^{+}17, PTMP^{+}15]. mTORC [MA17]. mTORC1 [CNRR^{+}17, HPW^{+}17, KNQ^{+}19, YPY^{+}15]. mTORC2 [ESS^{+}17, Sed15t, ZCL^{+}15]. mucociliary [GBRH15]. Muhire [O’D19c]. Müller [MRO^{+}15]. Mullins [Pow15b]. Multi [QPZ^{+}17]. Multi-omics [QPZ^{+}17]. Multiciliated [HTK^{+}16, WPA^{+}18, YHS^{+}15]. multifaceted [CJ17, RC15]. multifunction [LL17]. multifunctionality [CWG15]. Multilevel [HDA^{+}17]. multiorgan [dVGO^{+}16]. Multiple [KY15, YTTH^{+}17, DSS^{+}15, GY18, MCGM15a, MCGM15b]. multistep [OI18b]. Multitiered [MGA19]. multivariate [GSC^{+}16]. multivesicular [HAR^{+}15]. Munc13 [MWSM18, MWSM19, ZJM^{+}17]. Munc13-4 [MWSM18, MWSM19]. Munc18 [CST^{+}16, DR16, KCB^{+}16]. Munc18-1 [CST^{+}16, DR16, KCB^{+}16]. Muscle [BHS18, DV16, ST16a, DQB^{+}16, GCH15, LNH^{+}15, NWW17, NWP^{+}16, PLG^{+}15, Sho15o, SCP^{+}15, SFZ^{+}17, TYD^{+}15, WRV15, ST16a]. Muscle-on-chip [DV16]. muscles [DSS^{+}15, SRF19]. Muscling [O’D17e]. muscular [CKM^{+}16, NWP^{+}16, PPB^{+}15]. mushroom [SEMP15]. musical [SG17]. Mutant [LRM^{+}19, CGBD^{+}17, MGJ^{+}16, MT19, OBS^{+}17]. mutase [QSZ^{+}17a, QSZ^{+}17b]. mutated [RDO^{+}15]. Mutation [BNS^{+}17, EW17]. mutations [BBMM^{+}16, HV17, LPWK15, LYO15, TVG^{+}19, XTT^{+}18]. MVP [Sho16a]. My [PCM16, DR19]. MYC [TF19, MSK^{+}19, ZCL^{+}15]. Myc-dependent [ZCL^{+}15]. Mycobacterium [LBG^{+}17]. Myelin [Sho15-39, BBW16, GSCIL^{+}15, GCVAGS^{+}18, LM19, YKO^{+}16, vBMG^{+}15]. myelinating [PC17]. myelination [EVR^{+}19, MPN^{+}18]. Myelinophagy [TS15b, GSCIL^{+}15]. myeloid [CHZ^{+}17, DMC^{+}16]. Myo1E [Sho16d]. Myo5B [VKJ^{+}15]. myoblast [CLO^{+}19, DSS^{+}15, Sed15m]. myoblasts [SRF19]. Myoepithelial [GN18, SSE18]. myofibril [ARV^{+}18]. myomitokine [CRK^{+}17]. myonuclear [RAS^{+}19]. myonuclei [WSP^{+}18]. Myosin [IYP^{+}18, LDMW^{+}15, PLD^{+}15, RHH^{+}18, SRI^{+}19, SOW^{+}17, WW16, DB15a, EKP^{+}19, FLN^{+}10, FLN^{+}16, FB15, JGCAC^{+}15, LFK^{+}17a, MSE^{+}17, PUTM15, SLW^{+}18, SR17a, SWPS^{+}19, TBK^{+}16, THA^{+}16, TYD^{+}15, YWW17, ZAT^{+}19]. myosin-10 [SLW^{+}18]. Myosin-dependent [WW16]. myosins [LMdM^{+}16, PD19, SAT^{+}17]. mysterin [SMA^{+}19]. mysterious [O’D17g]. Myt1 [HHCK19].

N [LGH^{+}18, CVL^{+}19, KLS^{+}19, ZWS^{+}16]. N-cadherin [KLS^{+}19]. N-glycosylation [CVL^{+}19]. NAC [NP15, Sho15-49]. NAIPs [AMT^{+}15]. name [Sho15-45]. Nance [Pow16c]. Nanobodies [HCML15, Ewe18, PBG18]. nanoclusters [ED17, KG19, LBV^{+}17, SHVO^{+}18, SDP^{+}15a, SDP^{+}15b, THM^{+}19]. nanodomain [KCB^{+}16]. nanofibers [GTMZ^{+}15]. Nanometer [Tar15]. Nanometer-scale [Tar15]. nanoparticles [TCZ^{+}16]. nanoscale [SBM^{+}19]. Nanoscopic [AWS^{+}16]. nanotubes [VZ17]. Nap1 [MH15]. nascent [BG18, COGP15, Sho15-49]. Nasser [O’D17f]. native [CDF^{+}18]. Natural [FD18, RM16]. navigate [SW18]. Navigating [LW17]. navigation [AGB^{+}19, FFG^{+}18]. NBL1 [MBS^{+}17]. NBR1 [KSG^{+}16, Sho16s]. NCAM [FCB^{+}09, FCB^{+}19]. Ncd [BBS^{+}17, Das17]. Nck [DMH^{+}15, LRBB15]. Nck-dependent [DMH^{+}15]. Nck-interacting [LRBB15]. Ndc80 [CSC^{+}15]. Nde1 [WV18b]. Ndel1 [IGK^{+}16, Sho16t]. Ndj1 [LSJY15]. near [SRT^{+}18]. necessary [CRS^{+}17, FDR^{+}16, GSD^{+}15, LLY^{+}19, MSK^{+}19, TBJ^{+}17, UGHB^{+}16]. Necroptosis [SPH^{+}19, HGG^{+}17, PCK^{+}17, ZB19]. necrosis [TL17]. necrotic [LBG^{+}17]. nectins [KHS^{+}16]. need [Haw18, LW16b]. Needhi [Sed16d]. needs [NS15, NNK^{+}15]. Negative [LJS^{+}16b, FdAV^{+}17, LJS^{+}16a]. negatively [GLJ^{+}17, RSvW^{+}15, SLPW19]. neighbor [NF19]. Nek5 [PSP^{+}15]. Nek6 [OSR^{+}15]. nematodes [MOJ16]. neocentromeres [NKH^{+}19, MS19a, MS19b]. neocortex [SCL^{+}16]. neogenesis [DVS^{+}17]. neonatal [EKP^{+}19]. neoplasia [LMC^{+}18]. nerve [GCZ^{+}19, Sho17f, SCL^{+}16, vBMG^{+}15]. nerves [GSCIL^{+}15, KO19]. nervous [LPHH16, MPN^{+}18, OFP^{+}19]. Nesprin [SFZ^{+}17, WRV15]. netrin [DKM^{+}15]. netrin-1 [DKM^{+}15]. network [ASM^{+}15, BPH^{+}15, CRS^{+}17, CBM^{+}16, CPB^{+}16, DLZ^{+}15, GTW^{+}15, GKC^{+}17, HKT^{+}17, IB19a, IB19b, KS17, KWB^{+}15, KQM^{+}19, KST^{+}19, LEM17, OG16, RRM^{+}17, RNP^{+}17, Sho16e, YHS^{+}15]. networks [ES18, FK17, HVH^{+}19, NHA^{+}19]. neural [HB18, LJ17b, MBS^{+}17, PCP17, SXT16, SZF^{+}15, Sho16v, SR17a, SMN^{+}16, WCY^{+}16a, WCY^{+}16b, WYHG17, WHB^{+}18, YVlMS18]. Neuralized [PMRMS17]. Neurexin [KG19, THM^{+}19]. neurexin-1 [THM^{+}19]. neurite [FLS^{+}16, NWD^{+}19]. neurites [Les16a]. Neurl [LTS17]. Neurl-4 [LTS17]. neuroblast [SOW^{+}17]. neurodegeneration [KMRD^{+}16, NS18]. neurodegenerative [KM17, KM18a, VZ17]. neuroectodermal [ZGDS^{+}16]. neuroendocrine [ZJM^{+}17]. neuroepithelia [YEM^{+}19]. neuroepithelial [FC19]. neurogenesis [GDV19, LJ17b, XSJ18]. Neuroglian [SEMP15]. Neuroligin [LZD^{+}16, NL16]. neuromuscular [OLT^{+}19]. neuron [CKJ^{+}15, GHKW^{+}19, GJW^{+}17]. Neuronal [KM18b, AIS^{+}18, BNS^{+}17, CED^{+}15, DWB^{+}17, HCS^{+}18, HGM^{+}19, KDA^{+}18, Kaw17, KJ16, LRD19, LXJ^{+}17, MNL^{+}16, OG16, RFO^{+}16, Sho17d, TVG^{+}19, UGG18, WFOA15, WZR19, Yel18]. neuronal/axonal [WFOA15]. Neurons [CL19, FV17, Les15s, Sho16u, CF15, CXZ^{+}18, DGS^{+}18, IO18, JhZbYmP15, Kay16, MSC19, NWD^{+}19, ODH19, SZL^{+}16, SQC^{+}16, Sho16b, Van19, ZZW^{+}19]. neuropathy [XTT^{+}18]. neuroprotective [FKW^{+}17]. neuroscience [Sil16b]. neurotransmitter [BZG^{+}17, PNE^{+}19]. neurotrophin [KD17a]. neurotrophins [FAH^{+}17]. Neutrophil [Nie16, BLG^{+}15, FDR^{+}16, SAO^{+}17]. neutrophils [BRY^{+}19]. newly [OG16, SZ17a]. Nexin [WZR19]. NF [Hu15, LAMACE^{+}17, MCS^{+}15, ZLG^{+}15, dVGO^{+}16]. NF- [LAMACE^{+}17, MCS^{+}15, ZLG^{+}15]. NF2 [CMMB^{+}15]. NF2/ [CMMB^{+}15]. NG2 [LPHH16]. NGF [Sho15-39]. niche [LZC^{+}15, LWF^{+}15, Sho15-62, VZFG^{+}18]. niche-associated [LZC^{+}15]. niches [CAKL16, LLK^{+}17]. Nick [Pow15g]. nidulans [SMOO17]. Niemann [Col19, TVG^{+}19]. Nihal [Pow15h]. NIK [LRBB15, Les15t]. NIMA [CKKG17]. NIMA-related [CKKG17]. Nine [Pol17]. Ninein [RAS^{+}19]. NK [HMC^{+}16, SAK^{+}18, vHGD^{+}15]. NLRC5 [Hu15, MCS^{+}15]. NLRP3 [dlRHM^{+}18]. No [Les15u, Mar16a, OM19, KL19]. NOD [YVM18]. NODAL [OWW^{+}19, LFT^{+}16]. nodes [AOL^{+}18, CPEE^{+}15]. Nogales [Sed15i]. Non [DGS^{+}18, TYD^{+}15, GGC^{+}17, KDA^{+}18, NOS^{+}15, WYHG17, YYM^{+}18]. Non-canonical [DGS^{+}18]. non-cell [WYHG17]. non-channel [GGC^{+}17]. non-diploid [YYM^{+}18]. non-glycoprotein [NOS^{+}15]. Non-muscle [TYD^{+}15]. non-neuronal [KDA^{+}18]. non-self-distinction [LDP^{+}15]. nonautophagic [CD18]. Noncanonical [CSYB^{+}17, CGY^{+}19, SD16b]. noncatalytic [EJK^{+}16]. noncoding [Cas17b, GHS16a, GHS16b]. nondegradative [CXZ^{+}18]. nonhomologous [LTC^{+}16]. nonmuscle [JGCAC^{+}15, MSE^{+}17, SAT^{+}17]. nonreceptor [LMPG^{+}15]. nonselective [LgYL^{+}18]. NORE1A [DCB^{+}15, Sho15-40]. normal [DMD19, HCN^{+}15, TBJ^{+}17, ZPT^{+}15]. nose [Bea16]. Notch [Sho15-41, CAA^{+}17, DCO^{+}12, DCO^{+}16, EPF16, GLSS^{+}15b, GLSS^{+}15a, KDA^{+}18, KR18, LKM^{+}15b, WHB^{+}18]. Notch/p53 [KR18]. nothing [O’D19c]. Novel [BRACA^{+}16, FML^{+}17, LLZ^{+}19, AHS^{+}15, BDLB15, BLZ^{+}15, CWL^{+}17, CPP^{+}18, DER^{+}18, EMB^{+}15, GRU18, GLS^{+}15, HZB^{+}15, KG19, KOV^{+}16a, KOV^{+}16b, MWF^{+}15, PLS^{+}15, Pow15k, RXEB^{+}19, RO18, SERP16, SHR17, WXFS17, XWZ^{+}15]. novo [CG17, MPN^{+}18]. NPC [Sho15-43]. NPCs [BYMS^{+}19]. NPHP4 [YHS^{+}15]. NPM1 [APS^{+}17]. NRF1 [CCS^{+}19, PTR^{+}19]. NRF1/Ehmt1 [PTR^{+}19]. Nrp1 [HKH16]. NRZ [XLW^{+}18]. nuage [ABF^{+}16]. Nuclear [BSP^{+}17, HH16, KHA^{+}18, LW16a, MSW^{+}07, MJSB16, OI18b, SD16a, SPJ^{+}15, SZR^{+}15, XIZ^{+}18, APK^{+}18, ATRG19, BPS^{+}15, BRY^{+}19, CNN^{+}17, CZW^{+}18, CTN^{+}19, CGY^{+}19, DWH^{+}17b, FW16, GFvA^{+}15, GCH15, HSK^{+}19, HLW^{+}15, KHRL17, KL19, KJTY19, KPGG^{+}19, LDM15, LPRW17, LDG18, LTRW15, MTN^{+}16, MYT^{+}16, MBG^{+}18b, MWB^{+}19, MCOGD^{+}17, NGG^{+}16, OKY^{+}16, PVP^{+}19, PHKY17, PXN18, RHCS^{+}16, RFG19, RND^{+}17, SBR^{+}15, SHW^{+}17, SMF^{+}15, SG19, SM18, SER^{+}15, Sho15c, Sho15-31, SLD^{+}15, SKG^{+}16, SHC^{+}18, SFZ^{+}17, SMOO17, TYD^{+}15, TRM^{+}16, TGK^{+}19, UKHK15, WWW^{+}18, WMK^{+}16, YEM^{+}19, ZNR^{+}18, vdVFM^{+}17, MSW^{+}17]. nuclease [BSK^{+}19]. nucleate [BHDK17]. nucleated [BCM^{+}18, DBG^{+}15]. nucleates [JBE^{+}17]. nucleation [AATP17, cLNF^{+}16, RO18, VM19]. nuclei [Sho15q, SR17b, TALR^{+}19, WRV15]. nucleic [PCK^{+}17]. Nucleocytoplasmic [LDG^{+}15]. Nucleolar [OR17, BMW^{+}18, MRK^{+}18]. nucleolus [APS^{+}17, Pow15k]. nucleophagy [MRK^{+}18]. Nucleoplasmin [CTN^{+}19]. nucleoporin [DMV^{+}19, GCH15, SPGB^{+}17]. nucleoporins [LPRW17, SK19]. nucleosome [ARB^{+}19, NHA^{+}19, NAFM^{+}17]. nucleotide [CR18, GCW^{+}16, MF18, RLJ^{+}17, ZTR^{+}17]. nucleus [CID17, GAS^{+}18, HH16, Haw18, LW16a, PLH18, SZR^{+}15]. Num1 [LM15, PKC^{+}16, Sho15-42]. NuMA [CHS^{+}17, LSMG18]. NUMB [ZDSM^{+}18, CBF^{+}18, FG15, TZC^{+}15]. Numb/p53 [TZC^{+}15]. number [CSC^{+}15, FAH^{+}17]. numbers [ZSH17]. Numb’ing [KR18]. Nup132 [YAHH15]. Nup2 [SMOO17]. Nup210 [GCH15]. Nup60 [Les16h]. Nup82 [GFvA^{+}15, Sho15-43]. NuRD [GCA^{+}17]. nutrient [CCBC19, TM18]. nutrient-sensing [TM18]. nutrients [LK17]. NXF1 [BYMS^{+}19, BPW15].

o [Inf18c, BH15, KML^{+}15, LLW^{+}15]. Obelus [VLZ15]. obligatory [KMBO^{+}15]. observation [HBS^{+}15, VM19]. Occluding [CPP^{+}18]. occurs [HHS^{+}16, KD17b, MSW^{+}07, MSW^{+}17, iYJF^{+}16]. octameric [NAFM^{+}17]. off [Sed16a, SFA^{+}19]. off-rate [SFA^{+}19]. offers [FGR^{+}18]. oh [PCM16]. old [SD17, SK19]. olfactory [KHS^{+}16]. oligodendrocyte [dlFEvW^{+}15, vBMG^{+}15]. oligodendrocytes [EVR^{+}19]. oligomeric [RZS^{+}15]. Oligomerization [GPPJ^{+}18, CBM^{+}16, JCF^{+}17]. oligomers [HBS^{+}15]. oligosaccharyltransferase [SCG17]. Om45 [WLJ18]. OMA1 [KMRD^{+}16]. omics [QPZ^{+}17]. on-site [PAM^{+}16]. Oncogene [MTC17]. Oncogene-inducible [MTC17]. Oncogenic [RRM^{+}17, GI19, PGMM^{+}19, YGMR^{+}17]. One [BA18, MSvO17, Pow16b]. only [ISL^{+}18, NA16, MA17]. onset [HBS^{+}15, KBKW19, KMLG^{+}15, KMLG^{+}16]. oocyte [BCM^{+}18, CSC^{+}15, KBKW19, LTC^{+}18, LWZ^{+}19, LLS^{+}18, PMRM17, Ver18]. oocytes [BBS^{+}17, BTV16, BPSK^{+}16, BCS^{+}17, CO19, DRMW17, GCL^{+}15, HHH^{+}19, LJ17a, RO18]. oogenesis [O’D17e]. Ooplasmic [LLS^{+}18]. OPA1 [KMRD^{+}16, Les16e]. Open [BDAW15, Ewe18, UBR^{+}17, VAB^{+}18]. Open-source [Ewe18, UBR^{+}17]. Opening [THG19]. operated [CCQ^{+}18, RYS^{+}15, SBP^{+}16, WWT18]. Opi1 [HGF^{+}18]. Opportunities [HTLG18]. Opposing [KOIT^{+}16, Pow15e]. opposite [KDR^{+}19]. opposite-polarity [KDR^{+}19]. optic [CED^{+}15, Les15l]. Optical [ZZMC^{+}15]. optimal [LT19b]. optogenetic [AHS^{+}18, BOL17]. optogenetics [GGA^{+}17]. Orai1 [SBP^{+}16]. orange [PCK^{+}17]. orchestrate [BLG^{+}15]. orchestrates [AIK^{+}16, TSFP^{+}15, XSJ18]. order [BDW19, WZR19]. ordered [DB15b, KP18, MBG^{+}18a, SDP^{+}15a, SDP^{+}15b]. Org [SRF19]. Org-1 [SRF19]. organ [OWW^{+}19, SK18a]. organelle [BDK^{+}18, SJJ^{+}19, SPWM15, YWW17]. organelle-exclusion [SPWM15]. organelle-specific [BDK^{+}18]. organelles [MPMP16]. organism [FWH^{+}16]. organismal [SBM17]. organisms [FGR^{+}18, War15]. Organization [HR16, AGL^{+}15, BDZ^{+}15, CZW^{+}18, CBB15, Con16, DZB^{+}18, FC15, FZD^{+}19, GFH^{+}16, KQM^{+}19, LJP^{+}15, NLBA^{+}15, SSM^{+}18, UGG18, WMK^{+}16, YWdH^{+}17]. organize [AFO^{+}16, MKS17]. organized [TYK19]. organizes [BZG^{+}17, CKJ^{+}15, LFK^{+}17b, MXV^{+}16, RDO^{+}15]. organizing [TST^{+}17, Sed16a]. Organoids [SB17, HTLG18, MTC17, RRM^{+}17]. organs [KDA^{+}18]. orientation [EMB^{+}15, FC19, JKA^{+}15, LDM17, LSMG18]. origami [O’D16b]. origin [Blo19, SCP^{+}17, TGJ^{+}17, Yud19]. origins [Hyr15, KD17a]. orofaciodigital [RDO^{+}15]. ORP5 [SKZ^{+}18a]. ORP5/8 [SKZ^{+}18a]. orthologue [MCH^{+}18]. oscillation [NCV^{+}16]. osin [PCM16]. Osteoblastic [XJG^{+}17]. osteocalcin [FLG^{+}15, FLG^{+}19]. osteoclastogenesis [XJG^{+}17]. osteopetrosis [ZT15]. Osteopontin [CKM^{+}16]. other [ES18, O’D19c, TMFR^{+}19, War15]. otic [UBBSM15]. our [Inf19b]. outcompete [Pow16b]. Outer [DWH^{+}17b, BHS^{+}19, DW17, DUL^{+}19, KBB^{+}17, LXR^{+}15, SPE^{+}17a, SD17, VKT^{+}15, WLJ18, WF15]. outer-membrane [WLJ18]. outgrowth [FLS^{+}16, KBT^{+}15]. outposts [KYN^{+}18, LLL^{+}15]. output [CMTH^{+}15, JRH^{+}16]. OUTs [GS18]. outside-in [BBSA^{+}16]. outside-the-cell [She15]. ovarian [HHH^{+}19, LWF^{+}15]. ovary [LZC^{+}15]. overcoming [QZX19]. Oxidative [WFOA15, CYH^{+}16, CF15, KML^{+}15, TSK^{+}18, TSK^{+}19, Yel18]. OXPHOS [KM17, KM18a]. oxygen [Sch19].

P [CCLL17, CWZ^{+}15, Dic17, GPD^{+}19, HQW15, MBC^{+}19, PBL^{+}16, RSC^{+}19, DR19, SKZ^{+}18a, Yud19, GX16, KSGL19, NHG^{+}18, NGX^{+}19, Sho15-43, Sho15-45]. P-cadherin [PBL^{+}16]. p130 [WQD^{+}18]. p190 [FKL^{+}18a, FKL^{+}18b]. p190RhoGAP [BBMM^{+}16]. P2 [WXC^{+}18]. p21 [AGGSF^{+}16, YWW17, LDU^{+}16]. p21-activated [YWW17]. P2X4 [CZZ^{+}15, Sho15-44]. p37 [LSMG18]. p37/UBXN2B [LSMG18]. p38 [GAS^{+}15, HSZ^{+}18, DKA^{+}16]. p53 [DCB^{+}15, FG15, HPB19, KR18, LRM^{+}19, LUC^{+}15, LDU^{+}16, Les15v, LMC^{+}18, MAK^{+}16, HOH^{+}16, TZC^{+}15]. P53- [HOH^{+}16]. p62 [LLW^{+}17, SSRG18, Sho16v, WCY^{+}16a, WCY^{+}16b]. p62/SQSTM1 [WCY^{+}16a, WCY^{+}16b]. p63 [CE16]. p75 [DGS^{+}18]. p75-mediated [DGS^{+}18]. Pac1 [LM15]. Pac1/LIS1 [LM15]. Pac1/LIS1-mediated [LM15]. pace [GB18, Sho15-58]. Pace4 [BMC15]. pacemaker [ZB18]. packing [KKD^{+}16, Sho16m]. painful [GX16]. pains [Sho15-39]. pair [UFT^{+}15]. pairing [ABPS17, BLL15]. PAK2 [CSO^{+}19]. PAK4 [DBC^{+}15]. Palade [MPW^{+}19]. palmitoylation [MBF17]. palmitoyltransferase [MBF17]. Pancreatic [PW19, QCC^{+}19, ALY^{+}17, CIK^{+}17, DVS^{+}17, FWL^{+}17, KOIT^{+}16, TSK^{+}18, TSK^{+}19, VWM^{+}18]. PAPC [LDP^{+}15]. papillomavirus [IZZ^{+}18]. Par-1 [JH19]. PAR-4 [PUTM15]. paracrine [LLC^{+}17]. Paragons [CN15]. parallel [CWL^{+}16, SMF^{+}15]. paralogs [SG18a, SG18b]. paranodal [EVR^{+}19]. paraplegia [AEP^{+}17]. parasite [RNP^{+}17, TB16]. paraspeckle [ATS19, HYC16, WMK^{+}16]. Paraspeckles [CN15, HKM^{+}15]. Parkin [KPEJ17, LSMZ^{+}18, NS18, NWFY15, NPU^{+}16, OKK^{+}15, Sho15l]. Parkin-dependent [MLMF16]. parkinsonism [MGJ^{+}16]. PARP [MWW^{+}16]. PARP1 [AIS^{+}18, HGA^{+}17]. PARP1-dependent [AIS^{+}18]. PARP1/ARTD1 [HGA^{+}17]. PARP1/ARTD1-mediated [HGA^{+}17]. part [FB15, LTC^{+}16]. participate [SSM^{+}18]. participates [DZB^{+}18]. particle [KDV^{+}15, SLD^{+}15]. partition [LTG^{+}18, NIdG^{+}18]. partner [DK17]. partners [CB16, MP17b]. partying [Sed15c]. pass [SZ17b, Sho17a, SLH17]. passage [YNN18]. passenger [ARB^{+}19, FTDC17, IBFDB18]. Passing [O’D18d]. passive [TRM^{+}16]. patch [DPS^{+}18, Sho15-36]. pathogenesis [JERL^{+}15]. pathogenic [GLSS^{+}15b, GLSS^{+}15a]. pathology [GWF17]. pathway [BHS^{+}16, HOH^{+}16, IGK^{+}16, KVK^{+}17, KP18, KG15, KBT^{+}19, MBF17, MNLB16, MSW^{+}07, MSW^{+}17, MWSM18, MWSM19, MBC^{+}19, NMN^{+}15, NOS^{+}15, OG16, PYO^{+}18, RO18, SYK^{+}17, SB19, TJF18, TSK^{+}18, TSK^{+}19, vLvdKR18]. pathway-driven [HOH^{+}16]. Pathways [KJH18, AS17, ANM^{+}19, BSP16, CD18, CANG^{+}17, CID17, FA16, FC16, LH15, LKM^{+}15b, SMF^{+}15, Sho17b, TM18]. Patronin [FTS^{+}19, Bro19]. Patronin-mediated [FTS^{+}19]. pattern [KHS^{+}16, MBG^{+}18a, TYK19]. Patterned [WSDY17]. patterning [Bea16, HF15, YVlMS18]. patterns [GKC^{+}17, HTK^{+}16]. Pavarotti [MNLB16]. Pavarotti/MKLP1 [MNLB16]. Pax4 [DVS^{+}17]. paxillin [BVR^{+}17, KBT^{+}19, POE^{+}16]. PC7 [BMC15]. pcd [GHKW^{+}19]. PcG [CMM^{+}15]. PCNA [HSN^{+}16]. PCNA-binding [HSN^{+}16]. PDGFR [PLH18, SSV^{+}18]. PDK1 [PLS^{+}15]. PDK2 [GDL^{+}15]. PDK2-mediated [GDL^{+}15]. Pdr6 [ATRG19, RFG19]. Pdr6/Kap122 [ATRG19]. Peaks [Hyr15]. Pearce [O’D18c]. Pederson [Pow15k]. Peeling [Jor16g]. peerless [Bev17]. perception [BOL17]. perform [KGN^{+}15]. Peri [Pow15c]. pericentrosomal [LLW^{+}15]. periciliary [SHO^{+}15-74]. pericyte [NIN^{+}19]. pericytes [NLH^{+}19]. perilipin [GBK^{+}17]. perinuclear [SBR^{+}15, SHW^{+}17]. peripheral [KO19, MSC19, MPN^{+}18, OFP^{+}19, Sho17f]. Peripherin [MG17, SPD^{+}17]. peripherin-dependent [SPD^{+}17]. periphery [RHCS^{+}16, SM18]. Perlecan [CPEE^{+}15, NIN^{+}19]. permeability [iHMM^{+}17, KLS^{+}19]. permissive [HAPC^{+}19]. permits [DLZ^{+}15]. peroxiredoxin [LLW^{+}15]. peroxisomal [AFXS16, iHMM^{+}17, JHC^{+}16, KdBKvdK15, WXFS17]. Peroxisome [RBR19, SKVvdK15, CLV17, HCC^{+}17, MGE^{+}15, RDH^{+}19, Sho16-29, SZ17a]. Peroxisomes [GSB^{+}15, SERP16, CWI^{+}19, CLV17, CCH^{+}17, Hen19, HCC^{+}17, MHS^{+}18, OOT^{+}18, Sho16x]. perpendicular [SWC^{+}17]. persist [NF19]. Persistent [CNRR^{+}17, Bro19, MCD^{+}19, NW19, WB18]. Personal [O’D19b, Inf18b]. perspective [SB17, TG17]. perturbations [GMTL18]. perturbs [BNS^{+}17, GLSS^{+}15b, GLSS^{+}15a]. Pet10p [GBK^{+}17]. Petra [Pow15i]. pex1} [KdBKvdK15, KdBKvdK15, MGE^{+}15]. PEX2 [SvZS^{+}16]. Pex25 [SKVvdK15]. PEX5 [WXFS17]. PEX5-mediated [WXFS17]. pexophagy [RDH^{+}19, SvZS^{+}16]. Pgam5 [BJB^{+}18]. PGAP6 [LFT^{+}16]. pH [HHH^{+}19, JOJG16, NHG^{+}18, UGHB^{+}16, WGHE^{+}18]. phagocytes [Log17]. phagocytic [GWZ^{+}19a]. phagocytosis [OSW^{+}17]. phagophore [TLH^{+}19]. phagophores [GSRG^{+}18]. phagosomal [CWZ^{+}15]. phagosome [DLBMA^{+}15, LLL^{+}18]. phagosomes [JERL^{+}15, KSGL19, SLH17]. Phase [Woo18, BG19, CNC^{+}18, MGT^{+}19, SH17]. phases [MG16, RCS^{+}19]. phenotype [AIK^{+}16, CKM^{+}16, GG16, RFO^{+}16]. phenotypes [RRM^{+}17]. phenotypic [GLS^{+}17]. Pheromone [ADBST^{+}15, WTB^{+}19, vDMR^{+}19]. Pheromone-encoding [ADBST^{+}15]. pheromone-gradient [WTB^{+}19]. PHLPP2 [NKW^{+}19, TF19]. phosphatase [DJV^{+}16, DZB^{+}18, JKD^{+}19, MBC^{+}19, NMN^{+}15, NGX^{+}19, NKW^{+}19, RHC^{+}16]. phosphatase-independent [JKD^{+}19, MBC^{+}19]. phosphatases [Nil19]. phosphate [HHM15, Les15o, LJS^{+}16a, LJS^{+}16b, VMR^{+}19]. phosphatidic [HGF^{+}18, VKT^{+}15]. phosphatidylinositol [GCJ^{+}15, HHM15, JJB^{+}19, LJS^{+}16a, LJS^{+}16b]. Phosphatidylserine [MWT^{+}16]. phospho [WF15]. phospho-dependent [WF15]. phosphofructokinase [WDW^{+}17, ZSH17]. phosphofructokinase-1 [WDW^{+}17]. Phosphoglycerate [QSZ^{+}17b, QSZ^{+}17a]. Phosphoinositide [LMR^{+}17, SWC^{+}17, CW17, DCF^{+}17, WIS^{+}17]. phosphoinositide-activated [WIS^{+}17]. Phosphoinositide-dependent [LMR^{+}17, DCF^{+}17]. Phosphoinositide-mediated [SWC^{+}17]. phosphoinositides [DWH^{+}17a, DJV^{+}16, O’D18e]. phospholipase [LFT^{+}16]. Phospholipid [Les15w, AFO^{+}16]. phosphoproteomics [BC19]. Phosphoregulation [PKH^{+}19]. phosphorylate [TAQ^{+}19]. Phosphorylated [OKK^{+}15, BSP^{+}17, HBDW^{+}15]. phosphorylates [DMB^{+}18, HHCK19, TNP^{+}15, UMC^{+}15, UMC^{+}17]. phosphorylating [HSZ^{+}18, LRBB15]. Phosphorylation [LLY^{+}19, PKS^{+}19, SCNTC^{+}18, TCP^{+}18, TT19, CDF^{+}18, FBX^{+}15, GWL^{+}19, IKK^{+}18, JPF^{+}16, LSPC16, MFVS18, MNL^{+}16, MBG^{+}18a, MWW^{+}16, OSK^{+}15, SKW^{+}19, Sho16b, TGQ^{+}17, TTU^{+}17, WFOA15, WTSA17, WWY^{+}18, WKM^{+}15, WHiO^{+}19, WV18b, XMJ^{+}19, Yel18, YTGA16]. photoactivated [TCZ^{+}16]. photobleaching [FGR^{+}18]. Photoreceptor [Pug15, SPD^{+}17, DER^{+}18, HKG^{+}18, MSC19, MG17, SPE^{+}17a]. photoreceptor-specific [DER^{+}18]. photoreceptors [BK19, DSA15]. photosynthesis [Pow16d]. Physical [CKS^{+}15, ST17]. physiological [DTW^{+}16, PPB^{+}15]. PI [CCLL17, Dic17, GPD^{+}19, HQW15, KSGL19, NGX^{+}19, RSC^{+}19, Sho15-45, SKZ^{+}18a, WXC^{+}18, Yud19]. PI3 [JJW17, LLL^{+}18]. PI3-kinase [LLL^{+}18]. PI3K [GPD^{+}19, KG15, MDC^{+}16]. PI3K-dependent [MDC^{+}16]. PI4P [SE19, SKZ^{+}18a]. Pick [Col19, TVG^{+}19]. PICK1 [FRP^{+}17]. Piddini [Pow16b]. PIDDosome [APS^{+}17]. PIDDosome-dependent [APS^{+}17]. pieces [Sho18b]. Piecing [Sed15w]. PIK3CA [RRM^{+}17]. PINCH [GWZ^{+}19b]. PINCH-1 [GWZ^{+}19b]. PINK1 [MLMF16, NPU^{+}16, PSCS16, VGB^{+}17, KPEJ17, LSMZ^{+}18, NS18]. PINK1/Parkin [MLMF16, NPU^{+}16, NS18]. PINK1/Parkin-dependent [MLMF16]. Pioneer [Tra18, KS19, Pow15d]. pipeline [GSC^{+}16, UBR^{+}17]. PIPs [Yud19]. piston [PHKY17]. pit [MFVS18]. PIWI [ABF^{+}16]. PIX [LSS^{+}15]. PKA [IKK^{+}18]. PKA-RII [IKK^{+}18]. Pkc1 [vDMR^{+}19]. Pkc1-mediated [vDMR^{+}19]. Pkl1 [YIT15]. placement [PKH^{+}19]. places [MRWM18]. Plakophilin [DKA^{+}16]. Plakophilin-2 [DKA^{+}16]. planar [KZW^{+}18, iNLM^{+}19]. plant [BLPV^{+}17, NLS^{+}18, YTTH^{+}17]. plants [DMC^{+}17]. plaque [GWF17]. Plasma [RS16, AKTR18, BJO^{+}16, DQB^{+}16, Dic17, DZL^{+}15, EEE^{+}16, GDD^{+}15, GCJ^{+}15, GPD^{+}19, HHBG17, MPH^{+}15, MCGC^{+}15, MHA^{+}19, PD19, PKC^{+}16, RBM^{+}19, RSvW^{+}15, SKZ^{+}18a, TG15, Yud19]. plastic [Sil16b]. plasticity [BSL^{+}15, BLZ^{+}15, DSL^{+}17, FAH^{+}17, LMR^{+}17, LZD^{+}16, NGG^{+}16, PNE^{+}19, SVD^{+}15, TVG^{+}19, WBL^{+}15]. plasticity-induced [BSL^{+}15]. Plastin [DOH^{+}17, KKD^{+}16]. platelet [NS15, NNK^{+}15]. platform [Gek17, KY15, MTC17, Sho15-48]. play [KBJ16, VHB18]. player [KO19]. plays [ABP^{+}19, DKR^{+}19a, DKR^{+}19b, GAS^{+}15, Kti19, LCD^{+}17, LNH^{+}15, MHSD^{+}15, OKY^{+}16, Sho15-59, YGW^{+}17, Zhu17]. PLC [MBC^{+}19]. PLCXD [DR19]. PLEKHM1 [MAJ^{+}17]. PLK [TNP^{+}15]. PLK-1 [TNP^{+}15]. PLK1 [ABP^{+}19, Sho16r, ZLZD16, PTMP^{+}15, ZGZ^{+}15]. Plk4 [BCS^{+}17, GB18, MBG^{+}18a, MCL^{+}15, Sho15e]. ploidy [SHO^{+}18g]. PLP [LJP^{+}15]. pluripotency [CSYB^{+}17]. pluripotent [CEM^{+}15, TST^{+}17, ZGDS^{+}16]. plus [AHS^{+}18, FFG^{+}18, JNW15, KNPC16, LNS^{+}19, Wor19, YVM18]. PlxnD1 [HKH16]. PM [CCQ^{+}18, CCLL17, DJV^{+}16, SKZ^{+}18a]. PML [OKY^{+}16]. podocalyxin [MF16b]. Podosome [RLJ^{+}17, EWL16, Gen17, KBT^{+}19]. podosomes [Zha19]. point [Sho18c]. pointed [BHDK17]. points [NAFM^{+}17]. Pol [YGMR^{+}17]. polar [MSK^{+}18]. Polarity [GJW^{+}17, WKW^{+}15, BCMG19, BP19c, CLL^{+}16, CTI^{+}19, FTS^{+}19, FB15, GPAA^{+}18, GPPJ^{+}18, GAS^{+}18, IM16, JDZ^{+}16, JGCAC^{+}15, KDR^{+}19, KZW^{+}18, LCP^{+}15, LL17, MKA^{+}19, NLBA^{+}15, ONT^{+}19, PVP18, PP19, SWPS^{+}19, VLZ15, WTB^{+}19, WQD^{+}18, YHS^{+}15]. polarity-dependent [BCMG19]. polarization [BDZ^{+}15, CM18, DOH^{+}17, LWZ^{+}18, LEM17, LDMW^{+}15, RMOG17, ZAAN17]. Polarized [HLHFG15, BRACA^{+}16, GBRH15, NiYT^{+}16, PBL^{+}19, SMK^{+}18, SHO^{+}15-74, WB18]. polarizes [SAT^{+}17]. pole [CHC^{+}18, CSC^{+}15, IG15, PH18, RND^{+}17, SZF^{+}15, Sho15h, SHO^{+}18g, YIT15]. poles [UMC^{+}15, UMC^{+}17]. POLG [MTGG18]. Polo [KGN^{+}15]. Polo-like [KGN^{+}15]. polyadenylated [ACG^{+}15]. polybasic [DZL^{+}15]. Polycomb [DMD19]. polymerase [HLST19, MTGG18, NHA^{+}19, YGMR^{+}17]. polymerization [CJS^{+}18, CDT^{+}19, DWH^{+}17a, DMH^{+}15, DN16, DLT^{+}18, MG16, PTK16, PPR^{+}19, ZSH17]. polymorpha [SKVvdK15]. polyploidy [SF15]. polyposis [JBE^{+}17]. PolyQ [LOG15]. PolyQ-dependent [LOG15]. polysomes [PBS^{+}16, VLP^{+}15]. polySUMOylation [LHA^{+}15]. polySUMOylation-driven [LHA^{+}15]. polyubiquitinated [PAM^{+}16]. Pom1 [UMC^{+}15, UMC^{+}17]. pombe [RCS^{+}19]. pool [CWL^{+}17, PAM^{+}16, QSZ^{+}17a, QSZ^{+}17b]. pools [Sch17a]. population [WRGB^{+}15]. populations [OBY^{+}15, SFG^{+}17]. pore [BYMS^{+}19, GFvA^{+}15, KHRL17, KPGG^{+}19, LPRW17, LTRW15, NGG^{+}16, RND^{+}17, Sho15c, Sho15-31, SMOO17, TRM^{+}16, ZNR^{+}18]. pores [Sed15m]. porosity [LLZ^{+}19]. portraits [CSS^{+}18]. POSH [Sho15-46, WLM^{+}15]. position [AvdH16, JOJG16, PUTM15, WG16]. Positioning [LW16b, EEE^{+}16, FdAV^{+}17, HPW^{+}17, KNPC16, LDG18, MGW18, PMG^{+}17, PKKB17, RHCS^{+}16, RAS^{+}19, SFG^{+}17, SFZ^{+}17, ZAT^{+}17]. positively [Log17]. possible [MB15, POTZ15]. Post [SL19, THG19]. post-implantation [THG19]. Post-translational [SL19]. posterior [LLS^{+}18]. postmeiotic [PBG^{+}15]. postmitotic [PP19, TALR^{+}19]. Postnatal [AGGSF^{+}16, SQC^{+}16, NYW^{+}17, SCP^{+}15, SFZ^{+}17]. Postsynaptic [AMS^{+}17, LZH^{+}18, RKK^{+}18]. Posttranscriptional [HH18]. Posttranslational [NGG^{+}16]. potassium [Zhu17]. potent [ASM^{+}15]. potential [SSL^{+}17, SM18, ZGDS^{+}16]. potential-dependent [SSL^{+}17]. power [Les15-29, vV17a]. powerhouse [OI18a]. PP1 [LSMG18, RVS^{+}19]. PP1/Repo [LSMG18]. PP2A [CHB^{+}16, HBM^{+}19, JRH^{+}16, MBG^{+}18b, NNH17, PS16]. PP2A-B55 [MBG^{+}18b]. Ppz [LHT^{+}19]. Ppz-mediated [LHT^{+}19]. praise [War15]. Prdx4 [KML^{+}15]. Prdx4-mediated [KML^{+}15]. pre [CS16b, GSD^{+}15, JHC^{+}16, BMW^{+}18, SNOBM16, TTC^{+}16]. pre-40S [GSD^{+}15]. pre-60S [BMW^{+}18]. pre-chylomicrons [SNOBM16]. pre-chylomicrons/VLDLs [SNOBM16]. pre-mRNA [TTC^{+}16]. pre-peroxisomal [JHC^{+}16]. preassembly [zLSSS^{+}18]. precedes [CGPB17, IKK^{+}18]. precision [KJC^{+}15, PA19, WLC^{+}17]. precursor [LDR^{+}19]. precursors [LLK^{+}17, NYW^{+}17]. predicts [NTT^{+}15]. predisposes [GLSS^{+}15b, GLSS^{+}15a]. predominantly [GPD^{+}19]. prefer [Les16c]. preferential [BDLB15]. prefigure [NC18]. premature [SM18]. premRNPs [BPW15]. premRNPs/mRNPs [BPW15]. prepares [KHA^{+}18]. Prepatterning [NVP17]. Preprotein [CST^{+}17, BHB^{+}18]. preribosome [BPH^{+}15]. Presenilin [EPF16]. Presenilin-2 [EPF16]. presentation [YYZ^{+}15]. presenting [ST17]. presents [DJV^{+}16]. presequence [RDN^{+}19]. preserve [ACG^{+}15]. preserves [CRS^{+}17, FKW^{+}17]. pressure [Sed15o]. presynaptic [KG19, PAM^{+}16, VV17b]. prevails [Pug15, ZSH17]. prevent [ACRM17, GHKW^{+}19, HMC^{+}16, UMC^{+}15, UMC^{+}17, VAKB^{+}18, dVGO^{+}16]. preventing [KMRD^{+}16, ZLZD16]. prevents [DPS^{+}18, PMRM17, RSG^{+}15, VXF^{+}15]. Primary [VAKB^{+}18, ALLA18, ASPY^{+}16, BRACA^{+}16, GDB^{+}17, JNW15, LTG^{+}18, MFP17, OTG^{+}18, Ott16, PhHS^{+}16, PSL^{+}17, PM15, WHC^{+}19]. primes [Ger18, ZGDS^{+}16]. priming [FSB^{+}15, KCB^{+}16, KMK^{+}17a, KMK^{+}17b]. primordial [CAKL16]. principles [LDG18]. Prion [HKM^{+}15, OCS15, GUM^{+}18, HKG17, VZ17]. Prion-like [HKM^{+}15, GUM^{+}18, VZ17]. prions [Sed15c]. PRMT [AZS^{+}15]. PRMT-1-dependent [AZS^{+}15]. pro [CKM^{+}16, SG17]. Pro- [SG17]. pro-regenerative [CKM^{+}16]. Proactive [GML16]. probability [BZG^{+}17]. probe [GM18]. Probing [Pow16a]. problem [Mar16a, Sho16v, KL19]. Problems [MB15]. process [DB15b, GCZ^{+}19]. processes [BSP^{+}17]. processing [AS17, CYT^{+}18, GHS16a, GHS16b, KMRD^{+}16, LFT^{+}16, LGH^{+}18, RS19, Sho15-49, TTC^{+}16, TBL^{+}15]. Processive [CDT^{+}19, MGW18]. procollagen [GYK^{+}17, MSCS19]. prodegenerative [FKW^{+}17]. product [GPD^{+}19]. production [KML^{+}15, SLPW19, TCP^{+}18, dVGO^{+}16]. profile [LLL^{+}15]. profiling [BFPD19, FLS^{+}16, GLS^{+}17]. progenitor [CSS^{+}18, CLL^{+}16, FG15, LJ17b, SCP^{+}17, YVlMS18, dlFEvW^{+}15]. progenitors [TGJ^{+}17, WYHG17, WRGB^{+}15]. program [DAG^{+}15, GCVAGS^{+}18, PS16, Sho16w]. programmed [CLH^{+}18]. programs [HCS^{+}18]. Progranulin [CB16, Sho15-47, NBG^{+}16, ZSdO^{+}15]. progression [BS18, BBHBFSF18, DKR^{+}19a, DKR^{+}19b, HGC^{+}19, JPC^{+}17, LL19, Les15c, LLW^{+}15, NKW^{+}19, OSR^{+}15, SLW^{+}18, SENL^{+}15, SNB^{+}18, WWTF17]. progressively [ACRM17]. proinsulin [FWL^{+}17, TSK^{+}18, TSK^{+}19]. projection [ADBST^{+}15]. proliferating [IGK^{+}16]. proliferation [BP19c, CAKL16, CM18, CRA^{+}19, Col18, FC19, HBWY18, KPEJ17, LJ17b, MLR^{+}16, PLG^{+}15, PP19, SBS^{+}18, TSJ^{+}15]. Prolonged [GMTL18, LDU^{+}16, LH15]. prometaphase [SRT^{+}18]. promote [BVR^{+}17, Bro16, BJL^{+}18, CZZ^{+}15, CQB^{+}19, CLO^{+}19, DSS^{+}15, DS16b, DCO^{+}12, DCO^{+}16, EAW^{+}17, GLC^{+}19, GCA^{+}17, HMC^{+}16, JDZ^{+}16, KTM19, LTC^{+}16, LHT^{+}19, LZC^{+}15, LLW^{+}17, LWH^{+}18, MAJ^{+}17, MNLB16, MCL^{+}15, NLS^{+}18, OSR^{+}15, OFP^{+}19, Pri17, Sho17l, SWPS^{+}19, SKL^{+}18, SWC^{+}17, SCL^{+}16, VYB^{+}19, WTSA17, WTC^{+}19, WIS^{+}17, ZNR^{+}18, ZGZ^{+}15]. promoted [HLEM^{+}18]. promoter [GCL^{+}15, WWY^{+}18]. promotes [APHH^{+}19, ALY^{+}17, ACG^{+}15, ARV^{+}18, BKH^{+}15, BSL^{+}15, BAGM17, BS17b, CIS^{+}17, CLBB15, DPS^{+}18, DBC^{+}15, DLH^{+}19, DKA^{+}16, EKP^{+}19, FdSR^{+}17, FKL^{+}18a, FKL^{+}18b, GDD^{+}15, GBK^{+}17, GHKW^{+}19, GWF17, GHS16a, GHS16b, HSZ^{+}18, IZZ^{+}18, IBG^{+}15, IYP^{+}18, ISL^{+}18, KBKW19, KMLG^{+}15, KMLG^{+}16, KMC^{+}19, LAMACE^{+}17, LFK^{+}17a, LNS^{+}19, LLL^{+}18, LTB^{+}17, MMW^{+}19, MBF17, MBS^{+}17, MBG^{+}18b, NHCB15, OBY^{+}15, PNE^{+}19, PAC^{+}15, PBL^{+}16, PSP^{+}15, QSZ^{+}17a, QSZ^{+}17b, QZY^{+}19, RGM^{+}16, RMB^{+}18, RKK^{+}18, SCNTC^{+}18, SXT16, SSRG18, SQC^{+}16, Sho15-44, SOW^{+}17, SHH^{+}16, SCP^{+}15, SMN^{+}16, THA^{+}16, TSJ^{+}15, UDH^{+}16, VKT^{+}15, VGB^{+}17, WMB^{+}15, WMH^{+}18, Woo18, gXNG^{+}15, gXNG^{+}16, XJG^{+}17, XLW^{+}18, YTL15, YKO^{+}16, YHG^{+}17, ZCL^{+}15]. promoting [FFÁTC15, FMS^{+}19, KJTY19, RSG^{+}15, SHO^{+}18g, WEQ^{+}15, YYZ^{+}15, vdVFM^{+}17]. proofreading [CYMS^{+}19]. propagation [IYP^{+}18]. propel [MHY^{+}16]. proper [BG19, BDW19, YLW^{+}15, ZGZ^{+}15]. properties [ECAB^{+}16, GFH^{+}16, HNF^{+}18, PTK16, ST17, WRV15]. prophase [PTR^{+}19]. Prosaposin [ZSdO^{+}15]. prostaglandin [CBAP^{+}17]. prostate [Les15-30, NKW^{+}19, TF19]. Protease [SBS^{+}18, RDN^{+}19, WLJ18, RFG19]. proteases [BMC15, CC19]. proteasomal [CYT^{+}18, SPGB^{+}17, UOT^{+}16]. proteasome [PAM^{+}16, SSRG18, Sho18d]. protect [AMT^{+}15, BCMM^{+}19, CMTH^{+}15, VZFG^{+}18]. Protecting [IO18, OR17]. protection [JJW17, RS19, Sho15-64]. protects [BMM^{+}19, DWB^{+}17, HSN^{+}16, LUC^{+}15, NLS^{+}18, OPP^{+}18, PVP^{+}19, Sho16b, Sho16-27, WW16]. Protein [BSL^{+}15, BBK16, GNM16, HPB19, Nil19, RM19, RSvW^{+}15, SM18, SS18, SB19, AWS^{+}16, AKD^{+}17, BPH^{+}19, BPH^{+}18, BCH^{+}17, Boh18, BSP16, BG18, BK19, BPW^{+}17, Can17, CJ16, CMM^{+}15, CMA19, CMB^{+}18, CNN^{+}17, DMC^{+}16, DWH^{+}17b, DLZ^{+}15, DLBMA^{+}15, FdSR^{+}17, FKG^{+}19, FCLoS19, GPPJ^{+}18, GY18, GGWL^{+}19, GDB^{+}15, GLS^{+}15, Gli17, HKG17, HZB^{+}15, IdSCB^{+}16, IZZ^{+}18, IG15, ISK^{+}15, JHF^{+}15, KMK^{+}17a, KMK^{+}17b, KJTY19, KMC^{+}19, KML^{+}15, LPWK15, LHY^{+}19, Les15z, Les16i, LMPG^{+}15, LSJY15, LCZ^{+}16, LLW^{+}17, LGH^{+}18, LHB^{+}18, MDOS19, MLJ^{+}16, Mok16, MFP17, MWF^{+}15, NCV^{+}16, NEW^{+}17, NDRJ15, NPÖ^{+}17, NP15, PhHS^{+}16, PDZ18, PBL^{+}19, PXN18, PMP^{+}17, QCC^{+}19, QJP^{+}17, RPMC^{+}16, RBZ18, RLM^{+}15, RBR19, RLS18a, RLS18b, SERP16, SCNTC^{+}18, SSL^{+}17, SV16, SG18a, SG18b, SS19, SiYM^{+}18, SLM^{+}15, SHO^{+}15-74, SAB^{+}18, SE18, SCL^{+}19, TCP^{+}15, UMC^{+}15, UKHK15]. protein [VML^{+}17, VAB^{+}18, VMR^{+}19, VKT^{+}15, VGA^{+}15, WYoS17, WXFS17, WHS^{+}19, WYV^{+}19, WQD^{+}18, XSJ18, YHS^{+}15, YKO^{+}16, ZJM^{+}17, ZGDS^{+}16, LOG15, TLMG^{+}15]. Proteinopathies [KM17, KM18a]. Proteins [LVG^{+}18, AFXS16, AGL^{+}15, ABF^{+}16, BhHS^{+}17, BGJ^{+}16, BBS^{+}17, BMW^{+}18, BA18, BNKB15, BMS^{+}17, CD18, CGPB17, COGP15, DW17, DATI18, GUM^{+}18, HKG^{+}18, HR17, HGA^{+}17, HKM^{+}15, HMM^{+}19, JDZ^{+}16, JLB^{+}18, JHC^{+}16, KdBKvdK15, KPGG^{+}19, KJF^{+}18, KLHC^{+}18, KJON^{+}17, Les15g, Les15q, LMPG^{+}15, LKE15, MMW^{+}19, Mar16a, MSC19, MGJ^{+}16, MCOGD^{+}17, MRK^{+}18, MGE^{+}15, iNLM^{+}19, NPU^{+}16, NIdG^{+}18, NPÖ^{+}17, OLT^{+}19, SSM^{+}18, SZE19, SAF^{+}19, Sed16e, SG18a, SG18b, SD17, SMC^{+}15, Sho15k, Sho15-35, Sho17j, SLAR^{+}16, SWPS^{+}19, SKG^{+}16, SD16b, SJL^{+}19, TALR^{+}19, TG15, VZ17, WLJ18, YLND^{+}16]. proteoglycans [GSM^{+}15]. Proteolipid [YKO^{+}16]. proteolysis [CKS^{+}15, SXT16].

Proteome [ZHP^{+}19, MLJ^{+}16, SSdLA^{+}15]. proteomics [SKG17]. proteostasis [KJH18, SBM17]. proteotoxicity [OPP^{+}18]. Protocadherins [CED^{+}15]. protofilaments [MOM^{+}18]. protrusion [BBMM^{+}16, SS19]. protrusion-targeting [BBMM^{+}16]. protrusions [BDAW15, BVR^{+}17, ZDSM^{+}18]. Protrusive [GTMZ^{+}15]. prove [Sho17j]. provide [Gek17, MRO^{+}15, QYY^{+}16, YSW^{+}15]. provides [CYMS^{+}19, RMS^{+}18, SK16a, Sho15-36, WHP^{+}18, WRV15, WHL17, vGWC^{+}18]. proximal [DUL^{+}19]. PRPF8 [MCM^{+}17]. prunes [Les15w, Sho15s]. PS [SKZ^{+}18a]. pseudopod [DATI18, FLLM17]. pseudopod- [DATI18]. pseudopod-based [FLLM17]. PtdIns [CWZ^{+}15, MBC^{+}19, NHG^{+}18, DR19]. PtdIns3P [CWZ^{+}15, GWZ^{+}19a, HPW^{+}17]. PTEN [CNN^{+}17, Les17, MBC^{+}19, DR19]. PTP [JKD^{+}19]. publishing [Mar19]. Pucadyil [Sed15w]. Puertollano [Blu15b]. pulmonary [YGW^{+}17]. pulsed [MRMM18]. Pulses [Wu17]. pulsing [HPB19]. pump [Ger18]. Purkinje [GHKW^{+}19]. push [ES18, Les15s]. puts [Les15-30, Les16d, Sho15-29, Sho15-43]. Putting [Les15x, LW16a]. Puzzling [MP17b]. PxdA [SERP16, Sho16x]. Pyd [CRPSC^{+}19]. Pyd/ZO [CRPSC^{+}19]. Pyd/ZO-1 [CRPSC^{+}19]. Pyk2 [GLL^{+}18b]. pyruvate [DWB^{+}17].

Q [MMB^{+}15, Sho15-37, SJL^{+}19]. Q&A [Mar19]. QC [MPA^{+}16]. Qin [ABF^{+}16]. Quality [RLM^{+}15, Can17, PXN18, SZE19, SLAR^{+}16, SB19]. Quantification [AB18, BDAW15]. quantify [LLZ^{+}19]. Quantifying [VBJ^{+}18b, VBJ^{+}18a]. Quantitative [BC19, CVL^{+}19, CTS^{+}18, DSC^{+}18, KS19, WHP^{+}18, RLS18a, RLS18b]. question [HR17]. questions [Pol17]. quiescence [LCP^{+}15, MHG^{+}19, YLND^{+}16]. Quinlan [O’D17e].

R [CNA^{+}17, WZC^{+}15, XJG^{+}17]. R-loop [CNA^{+}17, WZC^{+}15]. R2TP [zLSSS^{+}18, MCM^{+}17]. Rab [IM16, LWH^{+}18, GWL^{+}19, HKK^{+}19, MF16b, RGMM18, RNP^{+}17]. RAB-10 [LWH^{+}18]. RAB-5 [LWH^{+}18]. Rab-mediated [IM16]. Rab1 [TJF18, WDM^{+}15]. RAB10 [BBC^{+}16]. Rab11 [SiYM^{+}18, TF16]. Rab11-binding [SiYM^{+}18]. Rab13 [IBG^{+}15]. Rab18 [XLW^{+}18]. Rab2 [LTB^{+}17, YHG^{+}17]. Rab27a [GCJ^{+}15, Sho15-48]. Rab27b [MDC^{+}16]. Rab3a [EEE^{+}16, RS16]. Rab3a-dependent [EEE^{+}16]. Rab46 [MPW^{+}19]. Rab5 [LT19b, ZWZ^{+}19]. Rab5-dependent [ZWZ^{+}19]. Rab5a [KSGL19]. RAB6 [FKG^{+}19]. RAB7 [CAA^{+}17, KNQ^{+}19, CW17, LXJ^{+}17, MAJ^{+}17, WHC^{+}19, YDM^{+}18]. Rab7-dependent [YDM^{+}18]. RAB8 [CAA^{+}17, NiYT^{+}16, WLM^{+}15]. rabbit [PBG18]. RabGEF [MVJ^{+}19]. Rabs [BDLB15]. Rac [LCM^{+}16, GKK16a, GKK16b, GGA^{+}17, Sho16y]. Rac1 [DSvNA^{+}15a, DSvNA^{+}15b, GFWG15, MRGWB^{+}16, MOJ16, RKK^{+}18]. Rac1/ [MOJ16]. Rac3 [DCM^{+}17]. race [ST16b]. RAD18 [YGMR^{+}17].

RAD51 [MWW^{+}16, PMHB17, RZS^{+}15, ZDM^{+}15]. Rad51-mediated [ZDM^{+}15]. Radial [Kay16, OFP^{+}19]. RADical [LS16]. Raft [KSM^{+}17, MCGC^{+}15, OPP^{+}18]. Raft-based [KSM^{+}17]. raft-dependent [MCGC^{+}15]. Rag [MP17a, MF18]. Ragulator [CJ17, FdAV^{+}17, PKKB17]. raises [SE19]. Ral [OFP^{+}19, HAR^{+}15]. RAL-1 [HAR^{+}15]. RalGTPases [GCZ^{+}19]. range [GRB19, MS19a, MS19b]. RanGTP [JHF^{+}15]. Ranvier [CPEE^{+}15]. Rap1 [GGC^{+}17, GLC^{+}19, OSL^{+}19]. Rapid [BLL15, LJ16, SSRG18, MOS^{+}18]. Rapsyn [OLT^{+}19]. RAS [Mar15, DCB^{+}15, KDM^{+}18, MKD^{+}18, PGMM^{+}19, ZWS^{+}16]. Ras-driven [PGMM^{+}19]. Ras/MAPK [KDM^{+}18]. Rasband [Sed15r]. RASSF4 [CCLL17, Dic17]. ratcheting [MVJ^{+}19]. rate [RCS^{+}19, Sho15d, SFA^{+}19]. rather [MHA^{+}16]. RBM3 [XSJ18]. Rbx1 [KSM^{+}18]. rDNA [MRK^{+}18]. re [MHG^{+}19]. re-entry [MHG^{+}19]. reach [Sho15-27]. Reaching [FR16]. reaction [Sho15t]. real [FJ17, SPJ^{+}15]. real-time [FJ17, SPJ^{+}15]. rearrangements [GGL^{+}19]. rearward [SHW^{+}17]. reason [Sho16z]. Reassessing [TB16]. RecA [AWL18]. RecA-mediated [AWL18]. recapitulated [RBR19]. recaptures [Les15g]. Receptor [JCF^{+}17, TGQ^{+}17, WBL^{+}15, APK^{+}18, BKH^{+}15, BSP16, BNB^{+}15, CRN^{+}19, CNN^{+}17, CB16, FdSR^{+}17, FRP^{+}17, FCB^{+}09, FCB^{+}19, FCLoS19, FTAB^{+}15, GPAA^{+}18, GKK16a, GKK16b, GKG^{+}18, HZH^{+}15, IdSCB^{+}16, KDV^{+}15, MBT16, MNLB16, MHY^{+}16, NBG^{+}16, NDL17, OKK^{+}15, PhHS^{+}16, RKK^{+}18, RFG19, STR^{+}18, SMC^{+}15, Sho18e, WV18a, dlFEvW^{+}15, OG16]. receptor-FoxO [MNLB16]. Receptor-mediated [JCF^{+}17, NDL17]. receptors [LKM^{+}15b, cLNF^{+}16, MOS^{+}18, Pas16, Pas19, SBP^{+}16, YYZ^{+}15]. Reciprocal [Sch17b]. reckoned [Jor16h]. recognition [AMS^{+}17, CHB^{+}16, FCLoS19, HGF^{+}18, KDV^{+}15, MTN^{+}16, SDHC17, WWTF17]. recognize [LMM16]. recognized [BJO^{+}16]. recognizes [HESKK15a, HESKK15b]. Recognizing [MA17]. recombinant [Ewe18, HCML15]. recombination [KHA^{+}18, LTC^{+}16, LCD^{+}17, QSZ^{+}17a, QSZ^{+}17b]. recombinational [LS16, Lov18]. reconstitute [MG16]. Reconstitution [BPL^{+}18]. reconstitutional [Sle16]. recovery [MRGWB^{+}16, SGB^{+}17]. RECQ [CNA^{+}17]. RECQ-like [CNA^{+}17]. RECQ5 [UDH^{+}16]. recruit [BDK^{+}18, CZL^{+}15, KNL^{+}17, Sho15-50, SWPS^{+}19]. recruited [CPEE^{+}15, WIS^{+}17]. recruiters [CWL^{+}16]. recruiting [KMC^{+}19]. recruitment [APHH^{+}19, BBK16, BMW^{+}18, EJK^{+}16, GPS^{+}17, HMM^{+}19, IB19a, IB19b, JCK^{+}19, KJF^{+}18, LSMG18, LCD^{+}17, MF18, MHSD^{+}15, MDC^{+}16, SKZ^{+}18a, TLH^{+}19, WTC^{+}19, WMB^{+}15]. recruits [BVR^{+}17, CR18, MKS17, MHA^{+}19, RHC^{+}16]. recycle [Les15h, Les16j]. recycling [Blu15b, BSP16, CPBG19, CYL^{+}18, DMS^{+}15, FC16, GLSS^{+}15b, GLSS^{+}15a, HHM15, LRM^{+}19, LWH^{+}18, MP17a, MBS^{+}18, MPH^{+}15, PAC^{+}15, Sho15-45, SE18, SCL^{+}19, ZDSM^{+}18]. redistribution [BNB^{+}15]. redox [MRWM18]. redox-regulated [MRWM18]. reduce [HSK^{+}16, PKN^{+}15, Sho16w]. reduced [LJ17a, PCP17]. reduces [MBC^{+}19]. reduction [SXE^{+}19, VXF^{+}15]. redundantly [CTI^{+}19]. reemergence [EWL16]. reestablishment [LCP^{+}15]. Reevaluating [BKG^{+}15]. Reevaluation [MGE^{+}15]. refilling [WZG^{+}17]. reformation [GWZ^{+}19a, MBG^{+}18b]. regeneration [GBRH15, LRH^{+}15, MSK^{+}19, NWD^{+}19, PGMM^{+}19, Sho16-33, ZYL^{+}16]. regenerative [CKM^{+}16, TSJ^{+}15]. region [BBMM^{+}16, MTC^{+}19, SPWM15]. regions [BA18, EGY^{+}19, NKH^{+}19, NPÖ^{+}17, PBL^{+}19, SER^{+}15]. regression [VAKB^{+}18]. regulate [AGL^{+}15, ATS19, BMC15, BNKB15, CNA^{+}17, CWZ^{+}15, CTI^{+}19, CL19, DMB^{+}18, DDAR^{+}16, DJV^{+}16, DCO^{+}12, DCO^{+}16, GLL^{+}18b, HKG^{+}18, HMM^{+}19, HBDW^{+}15, HM19, KHRL17, KPEJ17, KNL^{+}17, LGH^{+}18, LWF^{+}15, MCCL^{+}15, MPW^{+}19, MCOGD^{+}17, MYN^{+}17, NiYT^{+}16, NLBA^{+}15, ONT^{+}19, PUTM15, PPK^{+}16, PGRY^{+}19, RHCS^{+}16, SAF^{+}19, Sch17a, Sho15-39, WLM^{+}15, YWdH^{+}17, YSM^{+}17, ZWS^{+}16]. regulated [AFT^{+}19, CKS^{+}15, DB15b, EFM17, JAHH18, LHY^{+}19, MRWM18, MWSM18, MWSM19, ST17, TL17, THM^{+}19, XPZ^{+}19, zLSSS^{+}18]. regulates [AHA^{+}19, BPH^{+}18, BSK^{+}19, BSL^{+}15, BBSA^{+}16, CRN^{+}19, CANG^{+}17, CCQ^{+}18, CEM^{+}15, DMC^{+}16, DCM^{+}17, DCF^{+}17, FSB^{+}15, FRP^{+}17, FLG^{+}18, GLL^{+}18a, GLJ^{+}17, GLS^{+}15, GCA^{+}17, GCC^{+}18, HBWY18, HGD^{+}15, HHH^{+}19, iHMM^{+}17, HHM15, IKK^{+}18, JHC^{+}16, JGCAC^{+}15, KNPC16, KG15, KZW^{+}18, KBT^{+}19, KJTY19, KBB^{+}15, KBB^{+}16, LWZ^{+}18, LRM^{+}19, LCM^{+}16, LSMG18, LMR^{+}17, LSJY15, LJ17b, LLL^{+}15, LM16, LZD^{+}16, LSS^{+}15, Log17, LDMW^{+}15, MFVS18, MPMP16, MWW^{+}16, NKP^{+}15, NIN^{+}19, NLH^{+}19, OG16, OWW^{+}19, PTR^{+}19, PLH18, PMHB17, PMRMS17, PDZ18, PM15, QZX19, QSZ^{+}17a, QSZ^{+}17b, RBZ18, RHJW18, RSCR15, RSvW^{+}15, SPMM^{+}17, SBS^{+}18, SENL^{+}15, SR17b, SLPW19, TCP^{+}18, TF16, TWD^{+}17, TNK18, TSK^{+}18, TSK^{+}19, UOT^{+}16, UBBSM15, VTG^{+}16, VLZ15, WTSA17, WDM^{+}15, WWT18, WCL^{+}18, WSP^{+}18, WHS^{+}19, WHC^{+}19, WKM^{+}15, XWZ^{+}15, YTGA16, ZRDP19, ZPT^{+}15, ZLG^{+}15, ZGZ^{+}15, ZAAN17, dlFEvW^{+}15]. Regulating [ABPS17, FG15, Har16, PP19, FBX^{+}15, GCJ^{+}15, GCH15, HSB^{+}19, HQW15, IGK^{+}16, IYP^{+}18, KGN^{+}15, LLK^{+}17, LRD19, LTS17, MpDN^{+}17, RDH^{+}19, SSV^{+}18, SVD^{+}15, WYHG17, WBNH18, XJG^{+}17, YAHH15, ZTR^{+}17]. Regulation [ATH^{+}19, CYH^{+}16, CHS^{+}17, CAI^{+}15, CAA^{+}17, Far16, HLLK19, KSL^{+}17, LEM17, MF16b, PC17, RMMS^{+}17, TJMM^{+}18, TM18, ZAT^{+}19, ABP^{+}19, AOL^{+}18, AKD^{+}17, BBS^{+}17, Can17, CCLL17, DZL^{+}15, DB15a, FAH^{+}17, GCZ^{+}19, GP17, GWZ^{+}19b, HBM^{+}19, HH18, HDA^{+}17, HZH^{+}15, JNS^{+}19, KQM^{+}19, LAMACE^{+}17, LJ16, LJS^{+}16a, LJS^{+}16b, LZH^{+}18, MGA19, MC16, MSK^{+}19, MSL16, NNH17, Nil19, NPÖ^{+}17, OBS^{+}17, PTMP^{+}15, PMP^{+}17, RVS^{+}19, SJJ^{+}19, SXT16, Sch17b, STF18, SK16b, SK18a, SL19, YWW17, ZQZ19, ZCL^{+}15, vDMR^{+}19, vGWC^{+}18, CLV17]. regulator [CJ17, EMB^{+}15, FG15, FdAV^{+}17, KBJ16, PPB^{+}15, QJP^{+}17, QPZ^{+}17, SLW^{+}18, SIO^{+}16, SIBM17, VWM^{+}18, VMP16, WPA^{+}18, Dic17]. regulators [BCMM^{+}19, CPP^{+}18, KJC^{+}15, MCGM15a, MCGM15b, WZC^{+}15]. regulatory [DUL^{+}19, JGCAC^{+}15, POTZ15, RND^{+}17, SM18, SBC^{+}16a, SBC^{+}16b, VZB19, WBL^{+}15, WHB^{+}18, ZLG^{+}15]. reinnervation [SCP^{+}15]. reintroduction [KdBKvdK15, SKVvdK15]. rejection [vHGD^{+}15]. related [BMC15, CD18, DMC^{+}16, MGE^{+}15, VAB^{+}18, CKKG17]. relates [LLZ^{+}19]. relations [BFPD19]. relationships [CID17, TKM16]. relative [CZW^{+}18]. relax [GF16]. relaxation [GBD^{+}18]. relaxes [EW17]. RELCH [SiYM^{+}18]. RELCH/KIAA1468 [SiYM^{+}18]. release [BLZ^{+}15, BZG^{+}17, CZZ^{+}15, DNMB16, Das17, GSD^{+}15, KOR^{+}19, MWSM18, MWSM19, MPW^{+}19, MG17, NGX^{+}19, Nie19, NLH^{+}19, PNE^{+}19, PMW18, Rab17, SPD^{+}17, SZSS18, SZL^{+}16, SK18b, XJG^{+}17, YWW17]. released [BJB^{+}18]. releasing [KMJ^{+}18]. reliant [ZCH^{+}18]. relies [FFG^{+}18]. relieves [Sho15-42]. relieving [LM15]. relocalize [IBFDB18]. remain [LJ17a]. remnant [Ott16]. remodel [LLAC18a, LLAC18b]. remodeled [CO19]. remodeler [GCA^{+}17]. Remodeling [CAP^{+}16, BPH^{+}15, CRPSC^{+}19, GCW^{+}16, HLHFG15, KFAMR17, MSE^{+}17, OMKM16, PLH18, SJJ^{+}19, Sho16p, VCD^{+}15, WW16, Lac19]. remodels [AKTR18, BMW^{+}18]. removal [PhHS^{+}16, PM15]. remove [MGSO^{+}18, YNN18]. remyelination [Sho15-72]. renal [DSH^{+}18, LAMACE^{+}17, SQ15]. renewal [TZC^{+}15]. reorganization [CYT^{+}18, DTW^{+}16, LZD^{+}16]. reorientation [BP19a, BP19b, LNS^{+}19, NLS^{+}18]. Repair [HSK^{+}19, AWL18, ABGG16, BLL15, Can19, CPP^{+}18, CR18, DQB^{+}16, EEE^{+}16, EMRS^{+}18, GCZ^{+}19, GCA^{+}17, GRB19, GCW^{+}16, HLHFG15, LWZ^{+}18, LS16, LCD^{+}17, LPHH16, Lov18, MCGC^{+}15, MpDN^{+}17, NVP17, NIN^{+}19, PLG^{+}15, PMHB17, Pri17, QSZ^{+}17a, QSZ^{+}17b, SG19, Sho15-36, Sho17f, SOP^{+}16, SJ16, XIZ^{+}18, YGMR^{+}17, vV17a]. repairs [RS16]. Replicate [Góm17]. replicates [LBG^{+}17]. replicating [CST^{+}16]. Replication [OO18, ATH^{+}19, AWL18, BG18, BHS18, BCMM^{+}19, CQB^{+}19, CRS^{+}17, Col18, DKS15, Ger15, GRB19, HSN^{+}16, Hyr15, LL19, Les16g, Lov18, MLJ^{+}16, MN17, PST18, RS19, RLS18a, RLS18b, SPH^{+}19, SD16b, TBL^{+}15, UDH^{+}16, WSP^{+}18, XRH^{+}18a, XRH^{+}18b, YGMR^{+}17, ZDM^{+}15]. replication-associated [Lov18]. replication-dependent [AWL18]. replicative [TZC^{+}15]. Reply [TT19]. Repo [LSMG18]. reporter [Yud19]. repositioning [KDV^{+}15, Sho16-34]. repositions [BMM^{+}19]. represses [MNLB16]. repressing [FKL^{+}18a, FKL^{+}18b]. repression [CMM^{+}15, DMG^{+}19]. repressor [KDA^{+}18, OBY^{+}15]. Reproducibility [YH15, MB15]. reproduction [CM18, TM18]. reprogramming [IZBH^{+}17, LT19a, SRF19]. repulsion [GKK16a, GKK16b, MvVV^{+}16, PLD^{+}15]. require [HOH^{+}16, HBM^{+}19]. required [AKD^{+}17, BBC^{+}16, DOA^{+}17, EMRS^{+}18, FTS^{+}19, FBPN^{+}18, GJFR16, GHD^{+}17, GWL^{+}19, HHBG17, LJP^{+}15, LKE15, LWZ^{+}19, cLNF^{+}16, LXJ^{+}17, LT19b, LBB^{+}15, MHS^{+}18, MCGM15a, MCGM15b, MJN^{+}18, MGW18, MBG^{+}18a, MJSB16, MFP17, MHI^{+}18, MMB^{+}15, NCV^{+}16, ODH19, PLG^{+}15, PKN^{+}15, RLM^{+}15, RDN^{+}19, SPGB^{+}17, SvZS^{+}16, SZF^{+}15, SKW^{+}19, SHR17, SDW^{+}19, WRH^{+}16, WZG^{+}17, WYV^{+}19, XTS^{+}15, YLW^{+}15, YSR^{+}18, ZWB^{+}19, dlRHM^{+}18]. requirement [CM18, MvVV^{+}16]. requirements [AFXS16, BPS^{+}15, DS16a, WFS15]. requires [BLZ^{+}15, CAA^{+}17, HSK^{+}19, HB16, LLS^{+}16, PPR^{+}19, SKVvdK15, Too18, VMR^{+}19, WKW^{+}15, YDM^{+}18]. Rescue [XPZ^{+}19, Sho15j, VGB^{+}17]. rescues [CRC^{+}15, STR^{+}18]. rescuing [ZYL^{+}16]. research [Inf18c, MB15, O’D18b, O’D18f, O’D18g, O’D19a, O’D19i]. resection [SG17]. resegregation [BLL15]. reserve [YLND^{+}16]. reside [CNC^{+}18]. resident [CRN^{+}19]. resist [Nie19]. Resistance [AZ19, BS18, BAGM17, HOH^{+}16, SNGO16]. resists [SWC^{+}17]. resolution [CS16a, CDF^{+}18, EGY^{+}19, FGR^{+}18, HYC16, LLC^{+}17, PCF^{+}19, SNB^{+}18, UDH^{+}16, WMK^{+}16, XRH^{+}18a, XRH^{+}18b]. resolved [CSA19, WHB^{+}18]. Resonant [SKO^{+}15]. respiration [MWT^{+}16, VMP16]. respiration-active [MWT^{+}16]. Respiratory [HPE^{+}19, Sho15-37]. response [AB18, BSP^{+}17, CYT^{+}18, CAP^{+}16, CHZ^{+}17, CCBC19, DAG^{+}15, FCB^{+}09, FCB^{+}19, Gek17, HPB19, KH19, LAMACE^{+}17, LDM15, LH15, LCM^{+}16, MF18, NHCB15, NNK^{+}15, PKKB17, QJP^{+}17, QPZ^{+}17, RCS^{+}19, Sho17c, TCP^{+}15, VCD^{+}15, XTS^{+}15, ZDM^{+}15, ZZ16]. responses [HSZ^{+}18, HGG^{+}17, TSFP^{+}15]. restart [TBL^{+}15]. restores [BRY^{+}19, DBS18, TCZ^{+}16]. restraining [MBS^{+}17, NIS^{+}16]. restraint [Sho15-62]. restrict [AGB^{+}19, FKG^{+}19, KLS^{+}19, ZB19]. restricted [BMF^{+}18, QCC^{+}19, ZDSM^{+}18]. restricting [CYL^{+}18, DPGS^{+}18]. restricts [CCQ^{+}18]. results [CNRR^{+}17, DVS^{+}17, GMTL18]. resume [HHH^{+}19]. Ret [DGS^{+}18]. retardation [HPE^{+}19]. Retargeting [SWD^{+}19]. retention [LGH^{+}18, MF16a, UKHK15, XIZ^{+}18]. reticulum [GSRG^{+}18, GSB^{+}15, HSB^{+}19, JCF^{+}17, KML^{+}15, LPGB16, LLAC18a, LLAC18b, LGH^{+}18, MHS^{+}18, NDRJ15, PYO^{+}18, Pow15f, SNOBM16]. retina [HVH^{+}19, IKRMN16, MRO^{+}15, MSK^{+}19]. retinal [HKG^{+}18, TGJ^{+}17, Kay16]. retinoid [dlFEvW^{+}15]. retinoschisin [HVH^{+}19]. retracted [Sho16t]. Retraction [UMC^{+}17, HAK^{+}15]. retrieve [SMC^{+}15]. Retrograde [LZH^{+}18, CZL^{+}15, JNW15, KJON^{+}17, SHW^{+}17, SMC^{+}15, Sør17]. Retromer [CR17, CCY^{+}19, KNQ^{+}19, CCBC19, KJON^{+}17, MGJ^{+}16, SDHC17, SCL^{+}19]. retromer-independent [SDHC17]. retrotransposon [PST18]. retrotransposons [CHZ^{+}17, RMTR17]. return [GMTL18]. Reuse [GML16]. reveal [CSS^{+}18, ECAB^{+}16, GAS^{+}18, LYO15, VAB^{+}18]. revealed [GGWL^{+}19, GGA^{+}17, HGL^{+}17, KSM^{+}17]. reveals [AB18, BPL^{+}18, BC19, BBS^{+}17, BYMS^{+}19, BDLB15, BPS^{+}15, CVL^{+}19, CRZ^{+}16, CBF^{+}18, DB15b, EGY^{+}19, FTDC17, FLS^{+}16, GRU18, GPD^{+}19, GSC^{+}16, HBS^{+}15, JPC^{+}17, JPD^{+}16, KP18, KOV^{+}16a, KOV^{+}16b, MLJ^{+}16, NHA^{+}19, NIdG^{+}18, NKH^{+}19, NP15, PCF^{+}19, PBS^{+}16, PUY^{+}19, RND^{+}17, TALR^{+}19, VBJ^{+}18a, VBJ^{+}18b, XPZ^{+}19, XRH^{+}18a, XRH^{+}18b, Yud19]. reversal [ZDM^{+}15]. reverse [SYK^{+}17]. reversed [TBL^{+}15]. reverses [Les15z]. Reversible [MCS^{+}15, DSvNA^{+}15a, DSvNA^{+}15b]. revert [SRF19]. revisited [CR17]. rewire [CSM17]. Rewired [KR18]. rheostat [HB18]. Rhes [SS19]. RHO [Mar15, FLS^{+}16, GKC^{+}17, LR18, LW16b, LDMW^{+}15, NVP17, O’D18g, OOT^{+}18, SZSS18, Sho16y, TY16, TCWM18, TAQ^{+}19]. Rho1 [JRH^{+}16, Les16b, SKVvdK15]. RhoA [MOJ16, MXV^{+}16, MRMM18, MWB^{+}19, RMMS^{+}17, TGQ^{+}17, WG16]. RhoA-mediated [MWB^{+}19]. RhoB [KSM^{+}18, MRGWB^{+}16, ZRDP19]. RhoGAP [FKL^{+}18a, FKL^{+}18b]. RhoGEF [SWPS^{+}19]. RhoGEFs [NVP17]. RhoU [DBC^{+}15]. RI [LBV^{+}17]. ribosomal [BMW^{+}18, PUY^{+}19, SG18a, SG18b]. Ribosome [CGD^{+}18, KDV^{+}15, FCLoS19, GSD^{+}15, NP15, VLP^{+}15]. ribosome-induced [FCLoS19]. Ribosomes [Sho15-49, BMW^{+}18, DBS18, GSD^{+}15, SG18a, SG18b]. Richard [Jor16i]. Rick [Sed15u]. RIDD [TCP^{+}18]. Ride [Sør17, Kay16, Sho16x]. rideshare [Sho15f]. rificed [Sho15-45]. right [AvdH16, Jor16b, Les15o, LBD18, Sho17c]. rigidity [CLO^{+}19]. rigorous [JW19]. RII [IKK^{+}18]. ring [CHP^{+}17, GSP^{+}18, MSK^{+}18, Mar16b, Sho15-63, SOP^{+}16, SWC^{+}17, SKZ^{+}18b, SHO^{+}15-74, WMB^{+}15, XS16]. rings [Gli17, Mar16b, RBP^{+}17, Roy16, Sho16f]. RIP [Sho15-51]. RIPK3 [SPH^{+}19, ZB19]. RISC [KNL^{+}17]. Rivera [O’D19b]. Rme [GLSS^{+}15b, GLSS^{+}15a]. Rme-8 [GLSS^{+}15b, GLSS^{+}15a]. RNA [ACG^{+}15, BSK^{+}19, BSP^{+}17, CS16b, DMV^{+}19, DBS18, GHS16a, GHS16b, HBS^{+}15, HKM^{+}15, HLST19, JBMM16, LOG15, LJ16, NHA^{+}19, NIdG^{+}18, NPÖ^{+}17, O’D17d, PA19, PST18, SH17, YCSJ^{+}17, YLND^{+}16]. RNA-binding [NIdG^{+}18, NPÖ^{+}17, YLND^{+}16]. RNA-interacting [HBS^{+}15]. RNAi [NDL17]. RNAs [Cas17b, POTZ15]. RNase [MYT^{+}16]. RNase-sensitive [MYT^{+}16]. RNF11 [SIO^{+}16]. RNP [ADBST^{+}15, HCN^{+}15, Sho15-55]. ROBO [Sho16-27, MvVV^{+}16]. ROBO-cop [Sho16-27]. Robo-mediated [MvVV^{+}16]. ROBO1 [LCM^{+}16]. ROBO1/Rac/FAK [LCM^{+}16]. Robust [SOP^{+}16, DOH^{+}17, JDZ^{+}16]. ROCK1 [NLBA^{+}15]. rod [DSA15, GPS^{+}17]. role [AATP17, BYMS^{+}19, BRACA^{+}16, CAP^{+}16, FC15, FB15, FTDC17, GRU18, GAS^{+}15, JGCAC^{+}15, JIB^{+}19, LCD^{+}17, LNH^{+}15, MP17a, MHSD^{+}15, MGE^{+}15, MWF^{+}15, NYW^{+}17, NP15, OKY^{+}16, RND^{+}17, Sho16g, SR17a, SLD^{+}15, VBJ^{+}18a, VBJ^{+}18b, YGW^{+}17, Zhu17, CYH^{+}16]. Roles [POE^{+}16, ABP^{+}19, CR17, DKR^{+}19a, DKR^{+}19b, FFÁTC15, GGWL^{+}19, GDV19, GAS^{+}18, KGN^{+}15, KOIT^{+}16, LYO15, MSS^{+}17, POTZ15, Pow15k, RGMM18, RC15, SD16b, WBNH18, ZTR^{+}17]. Roll [Spe17b]. ROMO1 [RDN^{+}19, LgYL^{+}18]. Roop [Sil17]. root [SZL^{+}16]. rootlet [CKJ^{+}15]. Rootletin [CKJ^{+}15]. roots [Les15c, TG17, TL17]. Röper [O’D16b]. Ror [Sho15-50]. Ror2 [RSCR15]. ROS [SAO^{+}17, WBNH18]. Rosa [Blu15b]. Rotating [ST16b]. rotation [SBC^{+}16a, SBC^{+}16b, ZCH^{+}18]. route [Sho15-47, ZDSM^{+}18]. RPE [JERL^{+}15]. rRNA [BPH^{+}15]. RSC [SHO^{+}18g]. RSG1 [ALLA18]. Rsp5 [SZE19]. ruffles [PH18]. ruffling [CHC^{+}18]. ruler [CZW^{+}18]. rules [TRM^{+}16]. run [Sed15g, Sed15i, TCP^{+}15]. runs [Gli17]. rupture [CMTH^{+}15, HH16, LW16a, NS15, NNK^{+}15, PVP^{+}19, XIZ^{+}18]. ruptures [HSK^{+}19]. Rusan [O’D17f]. Rusty [Sil16b]. RZZ [MKA^{+}17].

S [BG19, MGT^{+}19]. S-phase [MGT^{+}19]. S. [LTRW15, RCS^{+}19, YTL15]. S149 [TT19, PKS^{+}19]. S729 [TCP^{+}18]. SAC [CYL^{+}18, Sho15-45]. SAC-1 [CYL^{+}18]. Sac-rificed [Sho15-45]. Sac1 [VMR^{+}19]. Sac2 [HHM15, NMN^{+}15, NGX^{+}19]. Sac2/ [NMN^{+}15]. Saccharomyces [LKM^{+}15a]. Saccharopine [LH19, ZWW^{+}19]. safe [Les17]. safeguards [MGT^{+}19]. Sag [XWZ^{+}15]. Sag- [XWZ^{+}15]. SAGA [EMRS^{+}18, Sed16e]. sail [Sho16d]. Saito [O’D19g]. Sall4 [Les15y, XTS^{+}15]. Sally [Pow15j]. Salmonella [SKL^{+}18, HGG^{+}17]. SAM [WEQ^{+}15]. Sam37 [WEQ^{+}15]. SAM68 [PPB^{+}15, MYT^{+}16]. Sanz [O’D18g]. saps [Les15-29]. sarcoma [CSF^{+}17, CSF^{+}18]. sarcoma-associated [CSF^{+}17, CSF^{+}18]. satellite [RHPH^{+}18]. satisfaction [KD17b]. Sbf [MVJ^{+}19]. scaffold [BP19c, CANG^{+}17, FLG^{+}18, GFvA^{+}15, HVH^{+}19, KDR^{+}19, KBB^{+}15, KBB^{+}16, LJP^{+}15, vDMR^{+}19]. scaffolds [GBB^{+}19, SZK^{+}19]. scale [BJO^{+}16, CWCG19, LMM16, SWD^{+}19, Tar15]. scales [LDM15]. scaling [CTN^{+}19, CID17, CL19, GBB^{+}19]. scan [Sho15-38]. scanning [SBR^{+}15]. scar [RMMS^{+}17, FLLM17]. SCAV [LCZ^{+}16]. SCAV-3 [LCZ^{+}16]. scenic [Roy16]. scents [Bea16]. schizophrenia [Sho16g]. Schuldiner [Sed16c]. Schwann [GCZ^{+}19, GSCIL^{+}15, GCVAGS^{+}18, MpDN^{+}17, MPN^{+}18, OFP^{+}19, TS15b]. Schwille [Pow15i]. science [CGT16, Inf18a, IO18]. scientific [O’D16a, O’D18d]. scientist [Bev17]. scientists [O’D17a, O’D19b]. scissioning [MHS^{+}18]. Scoping [O’D17f]. screen [BPH^{+}19, MLJ^{+}16, MHI^{+}18, NDL17, SIBM17, Sho18a]. screening [AHS^{+}15, BCG^{+}19]. Scrib [SYK^{+}17]. Scribble [AHA^{+}19, BP19c, CTI^{+}19]. seal [Sho15-29]. sealing [CWZ^{+}15]. search [HK15, Sho15j]. Sec10 [vGWC^{+}18]. Sec16 [MKS17]. SEC16A [BBC^{+}16, Sho16m]. Sec4p [LSPC16]. SecA [WYoS17]. secondary [Ewe18, PBG18]. Secretase [IZZ^{+}18, Sho15-51, LDR^{+}19]. secretases [CKS^{+}15]. Secreted [HGD^{+}15, GGWL^{+}19]. secretion [CRN^{+}19, CST^{+}17, CGBD^{+}17, DCP^{+}19, FKG^{+}19, GM16, Gli17, HAR^{+}15, ISK^{+}15, KKP^{+}17, KOIT^{+}16, MKS17, Pfe16, SDI^{+}19, SHH^{+}16, VBJ^{+}18a, VBJ^{+}18b, VBL^{+}18]. secretion-based [ISK^{+}15]. Secretory [CBM^{+}16, AIK^{+}16, CPBG19, DB15b, GDD^{+}15, GG16, KCB^{+}16, KOK^{+}19, NPÖ^{+}17, SSH^{+}15, SZSS18, SB19, TJF18, VBL^{+}18, vLvdKR18]. secrets [Blu15a, Spe17b]. secures [ARB^{+}19]. security [Sho15-36]. See [Sed15i]. seeds [Inf18c]. Seeing [Sho18e]. segment [AWS^{+}16, HR16, SPE^{+}17a]. segregate [LBV^{+}17, PSCS16, SHVO^{+}18]. Segregation [CGPB17, BRH^{+}16, BTV16, BG19, DW17, DLM^{+}15, KEV^{+}17, MDOS19, OKN^{+}16, OLL^{+}17, QZY^{+}19, SCNTC^{+}18, SKW^{+}19, SMOO17]. Seipin [SLPW19, Boh18, EBMW^{+}18, GBM^{+}15, SAB^{+}18]. seipin-linked [EBMW^{+}18]. seizure [HS16]. Selective [MTM^{+}17, BPW^{+}17, CCY^{+}19, KJON^{+}17, LLAC18a, LLAC18b, MOS^{+}18, MPW^{+}19, MFP17, VR18, VKJ^{+}15]. selectively [VZFG^{+}18]. Self [SAT^{+}17, CD18, CST^{+}16, LDP^{+}15, Sho15-62, TYK19, TST^{+}17, TZC^{+}15]. self-eating [CD18]. self-organized [TYK19]. self-organizing [TST^{+}17]. self-renewal [TZC^{+}15]. self-replicating [CST^{+}16]. self-restraint [Sho15-62]. Self-sorting [SAT^{+}17]. self/non [LDP^{+}15]. self/non-self-distinction [LDP^{+}15]. Sema3d [HKH16]. Sema4A [SYK^{+}17]. sends [Sho16m]. Senescence [MG18, AIK^{+}16, CNRR^{+}17, DCB^{+}15, GG16, Sho15-40]. senescence-associated [AIK^{+}16, GG16]. Senescent [NF19, DCB^{+}15, TMFR^{+}19]. Sense [Sch19, CWCG19, Nie16, Sed15j, Sed15s]. sensibility [Sch19]. sensing [CNRR^{+}17, McM19, TM18]. sensitive [KD19, RBZ18, MYT^{+}16]. sensitivity [MWW^{+}16]. sensitize [VAKB^{+}18]. sensor [KOV^{+}16a, KOV^{+}16b, WGHE^{+}18]. sensory [CKJ^{+}15]. Sentin [GCL^{+}15]. separate [MRK^{+}18]. Separating [KJ16]. separation [BCMG19, BDW19, LSJY15, LDP^{+}15, RSG^{+}15, Woo18]. septation [RCS^{+}19]. septin [BJO^{+}16, GMTL18, RBC^{+}17]. Septins [DS16b, LMM16, PMW18, CWCG19, GFH^{+}16, McM19, RBC^{+}17, Sho16-32, Sed15h]. septum [RCS^{+}19]. Seq [CZW^{+}18]. Sequence [SDHC17, POTZ15, YVM18]. Sequence-dependent [SDHC17]. sequences [BHB^{+}18]. sequentially [LDR^{+}19]. sequester [BA18]. sequestering [YCSJ^{+}17]. sequesters [RNP^{+}17]. sequestration [DLM^{+}15, SM18]. serial [SAK^{+}18]. series [CQB^{+}19]. serine [MSV16, SKW^{+}19]. Serrano [O’D19b]. serve [PXN18]. serves [SMC^{+}15]. service [PM18, Sho15f]. set [GSM^{+}15, Sed15h, Sho16d, Sho18d, AFT^{+}19, QZY^{+}19]. SET/TAF1 [AFT^{+}19]. Set1C [KHA^{+}18]. SETDB1 [CHZ^{+}17, RMTR17, DMG^{+}19]. sets [AWS^{+}18, BMF^{+}18, MF16a, SHC^{+}18]. severe [ZT15]. severely [NOS^{+}15]. severing [LNS^{+}19, MRM18, NLS^{+}18]. sexual [CM18]. Sey1p [YSW^{+}15]. SFI1 [KMC^{+}19, SER^{+}15]. SFT [SSM^{+}18]. SFT-4 [SSM^{+}18]. SFT-4/ [SSM^{+}18]. SGEF [AHA^{+}19]. SGK [HHCK19, HHH^{+}19]. Sgo1 [QZY^{+}19]. Sgo2 [AFT^{+}19]. Sgs1 [CNA^{+}17]. Shade [Sed15x]. Shao [Sed15j]. Shape [TKM16, FG16, Jor16d, LCM^{+}16, LMdM^{+}16, O’D19d, PCM16, SPE^{+}17a, SK18a, UGG18, Pow15f]. Shape-shifting [Pow15f]. shapes [JJB^{+}19, MCS^{+}15, PA19]. Shaping [SBM17, JHC^{+}16, LDG18, O’D17a, O’D17d]. share [Sho15-54]. Sharma [Inf19a]. Sharon [Sed16e]. She1 [ZAT^{+}17]. shear [BLO^{+}16, FDR^{+}16, Nie16]. shed [FD18]. Shedding [HYC16, SAK^{+}18, LFT^{+}16]. sheds [RMOG17]. sheet [FGR^{+}18, OSL^{+}19, WS18]. shell [Ver18]. shift [WHL17]. shifting [CKM^{+}16, May15, Pow15f]. Shiga [SIBM17]. SHIP2 [RHC^{+}16]. shock [AB18, Can17]. Shootin1 [KBT^{+}15]. short [Les16e, McM19, vS15]. show [JhZbYmP15, Sho15-69, SKO^{+}15, ZZMC^{+}15]. showcases [Les16a]. shows [Les15d, Sho15-62]. shrink [Les15k, Sho16n]. Shroff [Jor16f]. Shutting [vV17a]. shuttle [HR17]. side [BYMS^{+}19, BBHBFSF18, Jor16e, Sho18f]. sides [Sho15b]. Sigma1 [SBP^{+}16]. signal [AS17, CHB^{+}16, GGA^{+}17, KBT^{+}15, KDV^{+}15, Sho17h, Sho17i, SCL^{+}19, ZQZ19]. signaling [AUTM16, APHH^{+}19, ASM^{+}15, BJB^{+}18, BBSA^{+}16, BSP16, BDZ^{+}15, CD18, CS16a, CAKL16, CIS^{+}17, CNRR^{+}17, CRA^{+}19, CJ17, CAA^{+}17, CKKG17, DGS^{+}18, DLT^{+}18, ED17, EPF16, FVF^{+}16, FG16, FC19, FLS^{+}16, GPAA^{+}18, GKK16a, GKK16b, GDB^{+}15, GP17, GLSS^{+}15b, GLSS^{+}15a, GCVAGS^{+}18, GKGK16, GPD^{+}19, GYS18, GKC^{+}17, GWZ^{+}19b, GKG^{+}18, HBM^{+}19, HPW^{+}17, HHS^{+}16, HB18, IdSCB^{+}16, JJW17, JKA^{+}15, JRH^{+}16, KD17a, KG15, KZW^{+}18, KBT^{+}19, KNL^{+}17, KLS^{+}19, KNQ^{+}19, LLK^{+}17, LDU^{+}16, LR18, LFT^{+}16, LKM^{+}15b, MBT16, MPW^{+}19, MYN^{+}17, OKN^{+}16, OWW^{+}19, PDZ18, PLD^{+}15, RHJW18, RC15, SBS^{+}18, Sch19, Sch17b, SSV^{+}18, Sed15x, SD19, Sho15y, Sho17b, SL19, SYK^{+}17, THA^{+}16, TCWM18, TSJ^{+}15, TVG^{+}19, VBJ^{+}18a, VBJ^{+}18b, WBNH18, WV18a, WBL^{+}15, XMJ^{+}19, XJG^{+}17, YYZ^{+}15, YPY^{+}15, YSR^{+}18, ZLG^{+}15, dlFEvW^{+}15, vDMR^{+}19, vdVFM^{+}17, ODH19]. signals [CF15, CST^{+}17, CPB^{+}16, KMJ^{+}18, LZC^{+}15]. signature [SGB^{+}17]. signatures [PCK^{+}17]. significantly [BKR^{+}19]. SIK3 [KO19, LRD19]. silence [RMTR17]. Silencing [CHZ^{+}17, LPRW17, LLS^{+}16, RVS^{+}19]. Silva [O’D19e]. Simple [TRM^{+}16]. simplex [TGK^{+}19]. simplicity [HK15]. Single [DCP^{+}19, NHA^{+}19, NIdG^{+}18, BYMS^{+}19, ISL^{+}18, PBS^{+}16, PH16, SPJ^{+}15, SLD^{+}15, TBK^{+}16, VBJ^{+}18a, VBJ^{+}18b]. single-headed [TBK^{+}16]. Single-molecule [NIdG^{+}18, SPJ^{+}15]. single-particle [SLD^{+}15]. singularity [BDZ^{+}15]. SipA [SKL^{+}18]. Sir4 [LPRW17]. SIRF [RLS18a, RLS18b]. SIRFing [BG18]. siRNA [MLJ^{+}16, SIBM17]. SIRP [LBV^{+}17]. sisRNAs [Sho15-52]. sister [SNB^{+}18, TH18]. sisters [Lov18]. site [CGY^{+}19, NP15, PKH^{+}19, PAM^{+}16, SSM^{+}18, SZ17a, SLPW19, WTB^{+}19]. sites [BDK^{+}18, CVL^{+}19, CR18, GSRG^{+}18, GCW^{+}16, GBM^{+}15, KHA^{+}18, KBJ16, KLHC^{+}18, LFK^{+}17b, MKS17, MSŁK^{+}18, MYN^{+}17, NC18, OLL^{+}17, PMHB17, PHA^{+}17, RBP^{+}17, SA19, SKZ^{+}18a, VMR^{+}19, VGA^{+}15, XIZ^{+}18, ZNR^{+}18]. situ [NDC^{+}19, RLS18a, RLS18b]. size [AOL^{+}18, AWS^{+}18, CTN^{+}19, GM18, LDM15, LK17, PCP17, SWS^{+}19]. size-dependent [AOL^{+}18]. Ska [ACRM17, RGM^{+}16]. SKAP [KNPC16]. skeletal [LNH^{+}15, PLG^{+}15, SAF^{+}19, SFZ^{+}17]. Ski2 [VLZ15]. Ski2-family [VLZ15]. skin [CAI^{+}15, FC15, Fuc15, KH19, Sho15-64, XWZ^{+}15]. SLAMF1 [YSR^{+}18]. SLC [GWZ^{+}19a]. SLC-36.1 [GWZ^{+}19a]. Sli15 [FTDC17]. slide [Sho16u]. slidin [Bra16]. Sliding [RFO^{+}16]. Slip [Bra16]. slippage [BHS^{+}16]. Slit [Sho16-28, CRPSC^{+}19, MvVV^{+}16]. slow [Bro19, CDT^{+}19, GHD^{+}17, SCP^{+}15]. Slp3 [MDC^{+}16]. Slp3/Rab27b [MDC^{+}16]. Slp4a [VKJ^{+}15]. Sm [PMP^{+}17]. Smad [FVF^{+}16]. small [ALLA18, CGD^{+}18, FB15, GGC^{+}17, MF16b, MC15, RS16, TLMG^{+}15, YHG^{+}17]. Smallish [BPH^{+}18]. SMN2 [PPB^{+}15]. Smoothened [PhHS^{+}16]. SMRTer [Sho15-53]. Smurf1 [GWZ^{+}19b, WWZ^{+}17]. Snail [ST16a, WW16]. Snail-dependent [WW16]. Snail1 [LDP^{+}15]. Snail1-dependent [LDP^{+}15]. SNAP [Sho16z]. SNAP23 [KOIT^{+}16]. SNARE [BPL^{+}18, DR16, GRU18, KCB^{+}16, MHA^{+}19, SKL^{+}18, XLW^{+}18]. SNAREing [Too18]. SNAREs [LKM^{+}15a, Sho15-54]. snRNP [MCM^{+}17]. SNX [KJON^{+}17, SDHC17]. Snx13 [HZB^{+}15]. Snx14 [DLH^{+}19]. SNX3 [KSGL19]. SOCE [CCLL17]. SOD1 [CGBD^{+}17]. Software [OSL^{+}19, BDAW15]. solid [HCN^{+}15]. solidify [Sho15i]. Solidifying [Sho15-55]. Soluble [BKR^{+}19, CS16a, SSM^{+}18]. solution [Con16]. solutions [MB15]. soma [KM18b]. somal [GDV19]. Somatic [LYO15, ZZW^{+}19, CAKL16, VML^{+}17, YYM^{+}18, YDM^{+}18]. Something [SS18]. Song [O’D18a]. Sophie [Pow16e]. sortilin [ZSdO^{+}15]. sortilin-independent [ZSdO^{+}15]. Sorting [Sed15p, Sho15-56, Sho16-29, VR18, AFXS16, BK19, CR17, CBM^{+}16, CCY^{+}19, GNM16, KMBO^{+}15, LKM^{+}15b, MGJ^{+}16, NEW^{+}17, OFP^{+}19, Sho15-41, SAT^{+}17, SDW^{+}19, SCL^{+}19, WZR19]. Sotomayor [Blu15a]. source [BDAW15, Ewe18, MB15, UBR^{+}17]. sources [IdSCB^{+}16]. Sowing [Inf18c]. Sox9 [CRA^{+}19]. space [BPW15, Jor16c, MRWM18]. span [Zha19]. spastic [AEP^{+}17]. Spastin [CWI^{+}19, Hen19]. SPAT [TNP^{+}15]. SPAT-1 [TNP^{+}15]. SPAT-1/Bora [TNP^{+}15]. SPATA7 [DER^{+}18]. Spatial [AS17, CS16a, WV18a, YWW17, BSP16, BDZ^{+}15, CBB15, OLL^{+}17, TYK19, WLC^{+}17, vdVFM^{+}17]. spatially [HSB^{+}19, LFK^{+}17b, MSV^{+}19]. Spatiotemporal [ANM^{+}19, HHM15, JhZbYmP15, UGG18, FLS^{+}16, NVP17]. SPB [SER^{+}15]. Spc105 [RVS^{+}19]. Spc105-bound [RVS^{+}19]. Special [Sho15-57]. Specialized [SG18b, SG18a]. specific [ADBST^{+}15, ASM^{+}15, BDK^{+}18, BDLB15, CRPSC^{+}19, CSS^{+}18, CBF^{+}18, DER^{+}18, FCB^{+}09, FCB^{+}19, GFH^{+}16, ITN^{+}17, KMK^{+}17a, KMK^{+}17b, KYN^{+}18, LL17, LHT^{+}19, MGJ^{+}16, PTK16, PMRMS17, SG18a, SG18b, WWZ^{+}18]. specificities [PKC^{+}16]. specificity [BS17b, Kaw17, SCL^{+}19, WWZ^{+}18]. specifies [BBK16, JRH^{+}16]. speck [BS17a, KST^{+}17a, KST^{+}17b]. speckle [PABM16]. speckles [Les16i, WWW^{+}18]. Spector [Jor16c]. spectral [PCK^{+}17]. spectraplakin [WRV15]. spectrins [FLG^{+}18]. spectrometry [SKG17]. spectrum [APK^{+}18]. speed [CLL^{+}16, O’D19h]. speeds [Sho16-34]. Spef1 [KZW^{+}18]. spell [Sho15-33]. sperm [EMB^{+}15, GJFR16, PMRM17]. spermatid [PBG^{+}15]. spermatogenesis [RGR^{+}18, RXEB^{+}19]. sphere [Pas16]. sphingolipid [SLPW19]. Sphingolipids [PYO^{+}18]. sphingomyelin [KSM^{+}17]. Spiliotis [Sed15h]. spin [Pow15j]. spinal [CBAP^{+}17, PPB^{+}15]. Spindle [ABF^{+}16, BNKB15, KD17b, Sho15-58, ZCH^{+}18, AvdH16, APHH^{+}19, BBS^{+}17, BTV16, BCS^{+}17, BS17b, CHS^{+}17, CSC^{+}15, CO19, FFÁTC15, FTDC17, FC19, FBX^{+}15, GHS16a, GHS16b, HAPC^{+}19, HK15, IBFDB18, IWM^{+}16, IG15, KNPC16, KY15, LSMG18, LDG^{+}15, ML15a, MSŁK^{+}18, MGW18, iNLM^{+}19, NHCB15, NDC^{+}19, OSR^{+}15, PUTM15, PCF^{+}19, PMRM17, PDZ18, PCP17, RO18, RVS^{+}19, RND^{+}17, SLW^{+}18, SFG^{+}17, SPWM15, Sho15h, Sho15u, Sho15-59, SHO^{+}18g, TWD^{+}17, UOT^{+}16, VGY^{+}17, WG16, Woo18, YLW^{+}15, YAHH15, YIT15, ZLZD16, ZAT^{+}17]. spindle-centering [ZCH^{+}18]. Spindle-E [ABF^{+}16]. spindles [Das17, GSC^{+}16, Sho15w, Sho18c]. Spindly [Sho15-60, GPS^{+}17, MKA^{+}17, MWF^{+}15]. spine [WQD^{+}18]. spines [BSL^{+}15, GSS^{+}17, LMR^{+}17, LZD^{+}16, LSS^{+}15, Sch17a]. spinocerebellar [MNL^{+}16]. spinogenesis [RKK^{+}18]. splice [Sed15l]. spliced [RSC^{+}19]. spliceosomal [PMP^{+}17]. Splicing [Les15z, Sho15-61, WZC^{+}15, Cas17a, CS16b, GDL^{+}15, MTM^{+}17, MCOGD^{+}17, OOT^{+}18, PPB^{+}15, RYS^{+}15, VLZ15]. sporulation [Sho16-32]. spot [Les16i, SHO^{+}15-74]. spotlight [BP19a, BP19b, Sed15v]. spots [BP19a, BP19b, HHT^{+}16, Sho15-70]. Spotting [Pow15k]. Spp1 [KHA^{+}18]. SPR2 [NLS^{+}18]. spread [Sho15-61, VZ17]. spreading [BVR^{+}17, FVF^{+}16]. sprouting [NWD^{+}19]. spurs [Sho16-32]. SQSTM1 [WCY^{+}16a, WCY^{+}16b]. squeeze [LW16a]. Squeezing [SR17a]. SR [BMS^{+}17, HR17]. Src [GBD^{+}18, KG15, TAQ^{+}19, ANM^{+}19, CEM^{+}15, DPS^{+}18, HHS^{+}16, Sho18f, WWY^{+}18]. Src- [GBD^{+}18]. Src-mediated [ANM^{+}19, WWY^{+}18]. Srf [RHPH^{+}18, FBBRCA^{+}18]. SRP [FCLoS19]. SRP-receptor [FCLoS19]. stability [ATH^{+}19, AGL^{+}15, BRH^{+}16, BRY^{+}19, Can19, CQB^{+}19, EFM17, HPB19, HSN^{+}16, KBT^{+}19, KGN^{+}15, LLL^{+}18, MGT^{+}19, MHSD^{+}15, PBG^{+}15, SSV^{+}18, SFA^{+}19, TF19, WGHE^{+}18, XSJ18, YSM^{+}17, ZAT^{+}17]. Stabilization [QYC^{+}17, BSL^{+}15, DBC^{+}15, LNS^{+}19, PSL^{+}17, SID^{+}18]. stabilize [BGJ^{+}16, DSvNA^{+}15a, DSvNA^{+}15b, KMC^{+}19, RHC^{+}16]. stabilizes [ASZ^{+}18, ALY^{+}17, DLZ^{+}15, EG19, GBK^{+}17, GBM^{+}15, JKD^{+}19, NIS^{+}16, SMA^{+}19]. stabilizing [NKP^{+}15]. Stable [POTZ15, BGH18, GCL^{+}15, KMBO^{+}15, LCP^{+}15]. stably [CSF^{+}17, CSF^{+}18]. stacks [Les15l]. stage [ITN^{+}17, WG16]. stages [NDL17]. stalled [Les15i, RS19]. Stardust [DK17, PMRMS17]. starfish [BPSK^{+}16, HHH^{+}19, Ver16]. start [Les15i]. starting [Sch15]. Starvation [MOS^{+}18, MTGG18, NPU^{+}16, SvZS^{+}16, VTG^{+}16, vLvdKR18]. Starving [Les15-27]. STAT3 [RMMS^{+}17]. state [BMS^{+}17, HLW^{+}15, HAPC^{+}19, SRI^{+}19]. states [JPD^{+}16, RRM^{+}17, VRK^{+}17, VWM^{+}18]. static [Nel17]. stationary [GSKL^{+}18]. status [MF18]. stay [FV17, Sho15-32, Sho16n]. Staying [SA19]. Ste5 [vDMR^{+}19]. STED [SBM^{+}19]. steering [WRH^{+}16]. steers [Les15o]. Stem [KF18, ASPY^{+}16, BSK^{+}19, CNC^{+}18, CSG^{+}15, CEM^{+}15, Col18, DSC^{+}18, FG15, GBRH15, GCC^{+}18, GWZ^{+}19b, KPEJ17, LLK^{+}17, Les15y, LWF^{+}15, PLG^{+}15, PA19, RMB^{+}18, SZF^{+}15, Sed15e, Sho15-62, Sho16c, Sho16v, Sho17g, SR17a, SCP^{+}17, TGJ^{+}17, TST^{+}17, TSJ^{+}15, UGHB^{+}16, VY18, VZFG^{+}18, WCY^{+}16a, WCY^{+}16b, WHB^{+}18, XTS^{+}15, YLND^{+}16, ZGDS^{+}16, THG19]. Stemness [XTS^{+}15, NTT^{+}15, PP19, Sed15b]. step [ALLA18, Mar16b, MSW^{+}07, MSW^{+}17, ZB18]. stepped [MVJ^{+}19]. steps [SSL^{+}17]. stereocilia [AKD^{+}17, KKD^{+}16, LMdM^{+}16]. stereotypical [HTK^{+}16]. Sterol [MYN^{+}17, MST^{+}15]. stick [BK19]. stiff [Sho16-27]. stiffness [FVF^{+}16, LCM^{+}16]. stiffness-independent [FVF^{+}16]. STIL [KMC^{+}19, MCL^{+}15]. STIM1 [CCQ^{+}18, SBP^{+}16, WWT18]. STIM2 [RYS^{+}15]. stimulated [BBC^{+}16, MWSM18, MWSM19, TSJ^{+}15]. stimulates [CJS^{+}18]. stimulating [CIS^{+}17]. stitch [Góm17]. stoichiometry [DUL^{+}19, DMH^{+}15]. stokes [Zha19]. stomach [Sho15-34]. stop [Les15-32]. stops [Les15r]. store [CCQ^{+}18, RYS^{+}15, SBP^{+}16, WWT18]. store-operated [CCQ^{+}18, RYS^{+}15, SBP^{+}16, WWT18]. stores [WZG^{+}17]. story [MBR19]. straighter [BNS^{+}17]. strand [BLL15, DKS15, PMHB17, SJ16]. stranded [BSP^{+}17]. strange [Sho16-28]. STRAP [HM19]. Strategies [RHCS^{+}16]. stratified [MLR^{+}16]. Strength [ZSH17, CL19, GBB^{+}19, MRO^{+}15]. strengthens [NL16]. Stress [ATS19, Les15-28, ACG^{+}15, BMM^{+}19, BLO^{+}16, BCMM^{+}19, CYH^{+}16, CF15, CIK^{+}17, CHL^{+}19, DMC^{+}17, FDR^{+}16, HKG17, HSN^{+}16, HGG^{+}17, HGM^{+}19, JJW17, KPA^{+}16, KPA^{+}20, KP18, KPEJ17, KMRD^{+}16, LH15, LFK^{+}17a, Les16g, MTM^{+}17, NS18, Nie16, NIdG^{+}18, OI18b, PIA16, PKS^{+}19, PYO^{+}18, QPZ^{+}17, Sch19, Sho16q, Sho16-31, Sho17a, Sho17c, Sho17k, SENL^{+}15, TT19, WFOA15, XSJ18, YGMR^{+}17, vDMR^{+}19]. stress-dependent [WFOA15]. stress-induced [ATS19, ACG^{+}15, KMRD^{+}16]. stressed [Sed15n]. stresses [Blo19]. stretch [GSM^{+}15]. stretch-activated [GSM^{+}15]. strike [CLV17]. STRIPAK [NNH17]. stripes [BS17a]. stromal [SCP^{+}17]. Structural [ATRG19, AATP17, DSL^{+}17, FZD^{+}19, GFvA^{+}15, HGL^{+}17, SID^{+}18, WMK^{+}16, GML16, MSŁK^{+}18]. Structure [KTK^{+}18, MKA^{+}17, Boh18, CJ16, HYC16, KG19, LLZ^{+}19, OCS15, RCS^{+}19, SAB^{+}18, XRH^{+}18a, XRH^{+}18b]. structured [PTK16]. Structures [QYY^{+}16, VAB^{+}18, YSW^{+}15, GM18]. STT3A [CVL^{+}19]. STT3A- [CVL^{+}19]. STT3B [CVL^{+}19]. STT3B-dependent [CVL^{+}19]. Stu2 [vdVFM^{+}17]. study [Sho15k]. studying [MXS17, O’D17b, Sed15a]. Stx17 [KJF^{+}18]. subapical [YHS^{+}15]. subcellular [GKC^{+}17, ZRDP19]. Subdiffractional [JPD^{+}16]. subdomains [JHC^{+}16]. subnuclear [HKM^{+}15]. subpopulation [EBMW^{+}18]. subpopulations [BDLB15]. subset [KMK^{+}17a, KMK^{+}17b, LPRW17]. substitute [Les15u]. substrate [AHS^{+}15, CHB^{+}16, SJL^{+}19, XWZ^{+}15]. substrate-dependent [SJL^{+}19]. substrates [AHS^{+}15, WLJ16]. substructures [MYT^{+}16]. subtleties [Les16a]. subtype [IZBH^{+}17]. subunit [CGD^{+}18, HZH^{+}15, IKK^{+}18, KHA^{+}18, LRBB15, NEW^{+}17]. subunits [KPA^{+}16, KPA^{+}20]. suburbs [GF16]. suffice [Sed15u]. Sugar [Les15-29, BH15]. suggest [DW17]. suggests [HVH^{+}19, RXEB^{+}19]. sulfate [HGD^{+}15]. summits [O’D18f]. SUMO [RFG19, Sho15-63]. SUMO/protease [RFG19]. SUMOylation [RBC^{+}17, YTGA16]. sun [Sed15x, SS18, Sho15-64]. super [HYC16, WMK^{+}16, XRH^{+}18a, XRH^{+}18b]. super-resolution [HYC16, WMK^{+}16, XRH^{+}18a, XRH^{+}18b]. supercomplex [FZD^{+}19, WEQ^{+}15]. Superoxide [WBNH18, WCY^{+}16a, WCY^{+}16b]. Superresolution [PUY^{+}19, BYUJ17, LW17]. supplement [Sho15w, Sho15-72]. supply [MGA19, Sch17a, vV17a]. support [LKM^{+}15a, SCP^{+}17]. suppress [PMW18, SPH^{+}19]. suppresses [IGK^{+}16, LRD19, MTM^{+}17, NWD^{+}19, PGMM^{+}19, XMJ^{+}19]. suppressing [CRA^{+}19]. Suppression [LLC^{+}17, KDA^{+}18, SPD^{+}17, TZC^{+}15]. suppressor [CNN^{+}17, DMC^{+}16, TAQ^{+}19, WWY^{+}18]. suppressors [iNLM^{+}19]. sure [Les16f]. surf [HHT^{+}16, SS16]. Surf4 [SSM^{+}18]. Surface [DSS^{+}15, Far16, PLD17, YYZ^{+}15]. surface-tethered [PLD17]. surgery [CGT16]. surveillance [PYO^{+}18]. surveyed [SBM^{+}19]. Survival [Sho16-30, CSO^{+}19, CIK^{+}17, GDL^{+}15, PLS^{+}15, Sed15t, SSRG18, TMFR^{+}19, VTG^{+}16, ZCL^{+}15]. Susan [Bev17]. susceptibility [HS16]. suspended [SKO^{+}15]. sustaining [DPGS^{+}18]. sustains [CMM^{+}15]. SUV420H2 [VWM^{+}18]. Suzanne [KS19]. sweet [Jor16e]. swell [SKO^{+}15, ZS15, ZZMC^{+}15]. switch [KD19, Les15t, MCS^{+}15, Sho16-31]. switch-like [KD19]. switches [GDL^{+}15, WWZ^{+}18]. Switching [Sho16-32, SMK^{+}18]. symmetry [LOG15]. sympathetic [DGS^{+}18]. symphony [CV19]. synapse [AMS^{+}17, CBB15, Kon17, LMR^{+}17, MHY^{+}16, NKP^{+}15, NL16, SDI^{+}19, Sho15b, SAK^{+}18]. synapses [CBAP^{+}17, CPCtR^{+}15, OLT^{+}19, Sho17d]. synapsin [GHD^{+}17]. synapsis [ABPS17, BNKB15]. Synaptic [THM^{+}19, BNB^{+}15, BLZ^{+}15, BZG^{+}17, CYH^{+}16, CL19, GBB^{+}19, JPD^{+}16, KMK^{+}17a, KMK^{+}17b, LZD^{+}16, NLBA^{+}15, SES^{+}19, SVD^{+}15, Sho15-46, Sho16l, WLM^{+}15]. synaptogenesis [BP19c, KJ16, RKK^{+}18, SQC^{+}16]. synaptonemal [LHA^{+}15, SCNTC^{+}18]. Synaptopodin [KT15b, KT15a]. synaptotoxicity [QYC^{+}17]. syncytial [AGL^{+}15, SRF19]. syncytium [CV19]. syndrome [OSW^{+}17, SPE^{+}17a]. syndromes [RDO^{+}15]. Synergistic [KHS^{+}16]. syntaxin [MJN^{+}18, Juh16, MLMF16, VKJ^{+}15, KCB^{+}16]. Syntaxin-17 [Juh16, MLMF16]. syntaxin-1A [KCB^{+}16]. Syntaxin8 [SKL^{+}18]. synthesis [AFO^{+}16, DKS15, GLL^{+}18a, MTGG18, MPN^{+}18, RLM^{+}15, Sho15-36, TSFP^{+}15]. synthesizing [KMJ^{+}18]. synthetase [KKP^{+}17]. Synthetic [ECAB^{+}16, MXS17, PBL^{+}19, Sch15]. synuclein [CST^{+}16, DR16]. SYP [SCNTC^{+}18]. SYP-1 [SCNTC^{+}18]. system [LNH^{+}15, LPHH16, MLMF16, MPN^{+}18, OFP^{+}19, RFG19, SZE19, VRM^{+}19, WFS15]. systematic [KBB^{+}17]. systemic [CRK^{+}17, KH19, RHJW18, TMK18, ZQZ19]. systems [Pow16e, Sed16c].

T [HB16, CWL^{+}17, CPCtR^{+}15, CLBB15, DAG^{+}15, DK16, HBM^{+}19, HH18, MPH^{+}15, MJSB16, MHY^{+}16, NKP^{+}15, O’D18c, OBY^{+}15, PSC^{+}15, vHGD^{+}15]. T-bet [DAG^{+}15, vHGD^{+}15]. T-loop [HBM^{+}19]. TAF1 [AFT^{+}19]. TAGLN2 [NKP^{+}15]. tail [DB15a, FLN^{+}10, FLN^{+}16, Les15f]. tails [LBB^{+}15]. TAK1 [WLM^{+}15]. take [Les15q, Sho16s]. takes [FB15, LS16, Pfe16, Sho15q, Sho15-47, Sho16-37]. Taking [Les16g, SA19, Jor16f, Pow15i, Pow15j]. TALE [Bob17]. TALI [SNOBM16]. talin [GLC^{+}19, GYS18, KOV^{+}16a, KOV^{+}16b]. talk [MBT16, SZR^{+}15, TF16, WB18]. tally [OG16]. Tamas [O’D18e]. TamB [ZWB^{+}19]. TAN [SHW^{+}17]. tandem [GLC^{+}19]. tangential [NYW^{+}17]. tango [Sho15q, Pfe16]. TANGO1 [Gli17, MKS17, RBP^{+}17, SNOBM16, LFK^{+}17b]. Tapping [DK16, Yud19]. target [AMS^{+}17, CMM^{+}15, KVK^{+}17, MTN^{+}16, SPMM^{+}17, SAK^{+}18, SCP^{+}17]. targeted [HHT^{+}16, NDL17, OSR^{+}15]. targeting [BHB^{+}18, BBS^{+}17, BBMM^{+}16, BPS^{+}15, CST^{+}17, DZL^{+}15, FLN^{+}10, FLN^{+}16, FCLoS19, HLW^{+}15, KPGG^{+}19, LZH^{+}18, LBB^{+}15, MB17b, MWF^{+}15, Sho16-35, UOT^{+}16, UKHK15, WYoS17]. targets [BHS^{+}16, DMS^{+}15, HPB19, IWM^{+}16, KJC^{+}15, SKZ^{+}18b, WLJ16]. TAT1 [FBBRCA^{+}18]. TATA [PLH18]. Tatsushi [O’D18f]. Tau [QYC^{+}17, SID^{+}18, VXF^{+}15]. Tau-dependent [QYC^{+}17]. taxol [MHA^{+}16]. taxol-treated [MHA^{+}16]. TAZ [MCD^{+}19, PGRY^{+}19, GCC^{+}18, NW19]. TBC1D5 [KNQ^{+}19]. TCAF1 [Les15-30]. TCF1 [CSG^{+}15]. TDP [FSF^{+}15, Les15-31]. TDP-43 [FSF^{+}15, Les15-31]. technology [vS15]. tectum [CED^{+}15, Les15l]. telangiectasia [BLO^{+}16]. tell [Sho15z]. telomerase [OLL^{+}17]. Telomere [MB17b, AGGSF^{+}16, CG17, DKS15, Ger15, LSJY15, RSG^{+}15]. telomere-associated [LSJY15]. Telomere-driven [MB17b]. telomere-targeting [MB17b]. Telomeres [FFÁTC15, Sho16-33, DKS15, Sho15-73]. template [cLNF^{+}16]. temporal [AS17, GGA^{+}17, HPB19, WLC^{+}17, YVlMS18]. temporally [BMP^{+}18]. tensile [KJZ^{+}19, MRO^{+}15]. Tensin [DN17, GLJ^{+}17]. tensin-dependent [GLJ^{+}17]. tension [CAP^{+}16, DSvNA^{+}15a, DSvNA^{+}15b, KS17, KD17b, KOV^{+}16a, KOV^{+}16b, LVF^{+}15, LDMW^{+}15, MHA^{+}16, PMG^{+}17, RBM^{+}19, Sho15-65, TCD^{+}15, TNK18]. tensions [BHS18]. tent [CHC^{+}18, PH18]. term [FTAB^{+}15]. Terminal [vHGD^{+}15, DAG^{+}15, GDV19, LZH^{+}18, OBY^{+}15, SCL^{+}16, NL16, SER^{+}15]. terminals [FSF^{+}15, KG19]. termination [IKK^{+}18]. terminus [MRWM18]. test [Sho16q, Sho17a, SLH17]. testosterone [GLL^{+}18a]. tether [FR16, HCC^{+}17, Lac19, LCTP17, SJJ^{+}19, SES^{+}19]. tethered [PLD17]. tethering [CMA19, HZB^{+}15, IB19a, IB19b, MGW18, NO19, QYY^{+}16, TCP^{+}15, XLW^{+}18]. tethers [AKTR18, CWI^{+}19]. tetramer [FKO^{+}18]. tetraspanin [OPP^{+}18]. tetraspanins [DCO^{+}12, DCO^{+}16]. TFE [NWFY15]. TFEB [CCBC19]. TGF [DKA^{+}16, YYZ^{+}15, ZQZ19]. TGF- [DKA^{+}16]. TGN [DOA^{+}17, VMR^{+}19]. their [Bea16, BCM^{+}18, DLZ^{+}15, EGY^{+}19, GBK^{+}17, HPB19, JOJG16, Les16a, Les16c, NHG^{+}18, Sed15l, SZ17b, Sho15-27, Sho15-52, Sho15-69, Sho16j, Sho17b, Sho17j, Sho17k, TMFR^{+}19, ZNR^{+}18]. themselves [Sho15-38]. theory [GGR15, TYK19]. therapeutic [MG18]. therapies [MB17b, TG19]. therapy [ASPY^{+}16]. There [Sho17l]. thereby [WEQ^{+}15]. things [Jor16d, O’D19d, Sed16d]. Thinking [MC15, She15, SB17]. thiol [RGOS^{+}16]. thiol-based [RGOS^{+}16]. Thirty [HK15]. Thomas [Sed15w]. Thoru [Pow15k]. thought [Yel18]. Three [FAH^{+}17, Sho16-34, VLP^{+}15, JhZbYmP15, MSW^{+}07, MSW^{+}17, MG16, Sho17i, SB17]. three-alarm [Sho17i]. three-step [MSW^{+}07, MSW^{+}17]. Three-tier [FAH^{+}17]. threshold [WXC^{+}18, XPZ^{+}19]. thrombopoiesis [NNK^{+}15]. Thrombospondin [RKK^{+}18]. throughout [MGA19]. throughput [BCG^{+}19, Pow16d]. thy [NF19, FSB^{+}15, Sed15x]. Thy-1 [FSB^{+}15, Sed15x]. thyself [Sho15g]. Tiam [GKK16a, GKK16b]. ticket [Sho15l]. tier [FAH^{+}17]. tight [ONT^{+}19, SOII18, SLM^{+}15, TE15]. tiki [LPHH16]. Tilted [FGR^{+}18]. TIM23 [RPMC^{+}16]. Time [CSA19, WHB^{+}18, BPW15, FJ17, Góm17, Jor16c, Les16f, MSvO17, O’D18a, SPJ^{+}15, SHC^{+}18]. Time-resolved [CSA19, WHB^{+}18]. timeline [Sho15-66]. timely [ABPS17, DOH^{+}17, SOP^{+}16]. timer [BMP^{+}18, MF16a]. timing [MN17, OO18, PST18]. TIMP [SAF^{+}19]. tip [Les15m, LHB^{+}18, NDRJ15, YVM18]. Tipping [AvdH16]. tips [MOM^{+}18, Sed15t, Sho15-71, Sho16l, THA^{+}16]. Tissue [HF15, LLS^{+}16, CPP^{+}18, Jan18, JNS^{+}19, KS19, KTM19, KQM^{+}19, LDP^{+}15, O’D16b, PLG^{+}15, SBM17, Sho17l, SBC^{+}16a, SBC^{+}16b, SLG^{+}18, TY16, YEM^{+}19]. tissues [FBPN^{+}18, Nel17]. Titration [CRC^{+}15]. Tks5 [CLO^{+}19]. Tkv [LWF^{+}15]. TLR3 [HCS^{+}18]. TLR4 [HS16, SQC^{+}16, YSR^{+}18]. TLR4-mediated [YSR^{+}18]. TLR7 [HCS^{+}18]. TLR8 [HCS^{+}18]. TMEM231 [RDO^{+}15, Sho15-67]. TMEM41B [MHI^{+}18]. TMX1 [RGOS^{+}16]. TNT [SS19]. TNT-like [SS19]. TOG [BS17b]. TOG-tubulin [BS17b]. together [MB17a, O’D18c, O’D18g, Sed15k, Sed15w, Sho18b, TH18]. tolerance [RMS^{+}18, YGMR^{+}17]. tolerogenic [VRK^{+}17]. Toll [MNLB16, OG16]. Toll-like [OG16]. Toll-tally [OG16]. Tolls [FAH^{+}17]. TOM [WEQ^{+}15]. Tom22 [WLJ18]. Tom70 [BHB^{+}18]. tool [ISK^{+}15, LLZ^{+}19]. toolbox [PBG18]. tools [BOL17, WLC^{+}17]. tooth [AUTM16, SK16a]. TOPBP1 [LS16, LCD^{+}17, MWW^{+}16, SG17, PKN^{+}15]. TOPII [MGSO^{+}18]. Topo [Sho16-36]. topoisomerase [ABGG16, EJK^{+}16, YTGA16, LRS^{+}17]. topology [GRU18, MRWM18, SLG^{+}18]. TORC1 [MP17a, MYN^{+}17, vdVFM^{+}17]. TORC1-independent [MP17a]. TORC2 [LT19b, MYN^{+}17, RBM^{+}19, VMP16]. TORC2-mediated [VMP16]. TorsinA [SHW^{+}17, SR17b]. totipotency [LT19a]. touch [SA19, Sho15-32]. touchy [Les15p]. touchy-feely [Les15p]. toxicity [MCH^{+}18]. toxin [LH19, SIBM17]. Toxoplasma [RNP^{+}17]. TPX2 [AATP17, BCMM^{+}19, Can19, FBX^{+}15]. TPXL [MSK^{+}18]. TPXL-1 [MSK^{+}18]. TR [GX16]. track [Ava18, EFM17, Sho15-53]. Tracking [DB15b, Sho16-35, JPD^{+}16, SPJ^{+}15, WTB^{+}19, PH16, Pow15h]. traction [DPGS^{+}18, JhZbYmP15]. traffic [DS16b, HHT^{+}16, Inf19a, Jor16e, MAJ^{+}17, MP17b, WDM^{+}15, WHS^{+}19]. Trafficking [dlRHM^{+}18, BhHS^{+}17, BKH^{+}15, BBC^{+}16, CWI^{+}19, CJ17, DCO^{+}12, DCO^{+}16, DLBMA^{+}15, EPF16, FWL^{+}17, Far16, FCB^{+}09, FCB^{+}19, GLS^{+}15, GSS^{+}17, IM16, MRGWB^{+}16, MSCS19, MFP17, MF16b, ODH19, SSC^{+}19, SIBM17, Sør17, ZSdO^{+}15, ZWS^{+}16]. trail [Pow15a, Sho15a]. trails [Roy16]. training [O’D19b]. trains [YNN18]. TRAIP [HSN^{+}16]. TRAM [YSR^{+}18]. TRAMM [Les15-32, MHSD^{+}15]. TRAMM/ [MHSD^{+}15]. tranquilizes [Sho15y]. trans [CBM^{+}16, DSvNA^{+}15a, DSvNA^{+}15b, GKK16a, GKK16b, IB19a, IB19b, Sed15p]. trans-endocytosis [GKK16a, GKK16b]. trans-Golgi [CBM^{+}16, IB19a, IB19b, Sed15p]. trans-interaction [DSvNA^{+}15a, DSvNA^{+}15b]. Transcription [EMRS^{+}18, BMP^{+}18, BGH18, Cas17a, CDF^{+}18, DAG^{+}15, Ger18, HHBG17, LT18, NWFY15, PBG^{+}15, RHCS^{+}16, SPK^{+}18, Sed16b, TSB^{+}18, UDH^{+}16]. Transcriptional [OBY^{+}15, SK18a, VRK^{+}17, Bob17, CMM^{+}15, CIK^{+}17, DMG^{+}19, FBBRCA^{+}18, GCVAGS^{+}18, HCS^{+}18, KF18, QJP^{+}17, SIO^{+}16, SXT16, XMJ^{+}19]. transcriptome [LJ16]. transcriptomics [WHB^{+}18]. transcytosis [NDL17]. transducer [ITN^{+}17]. transduces [VCD^{+}15]. Transducing [CF15]. transduction [KBT^{+}15]. transfer [AFO^{+}16, CPCtR^{+}15, GY18, KTK^{+}18]. transferrin [DNMB16]. transforming [KKD^{+}16]. Transient [LXR^{+}15]. transit [BGH18, WWW^{+}18]. Transition [VPD^{+}16, DCF^{+}17, HHCK19, LWZ^{+}19, RDO^{+}15, SXT16, SSPD15, YNN18]. translating [KP18, O’D19e]. Translation [PBL^{+}19, PBS^{+}16, PH16, SPMM^{+}17, SG18a, SG18b, VLP^{+}15]. Translational [NPÖ^{+}17, Inf18c, O’D19a, SL19]. translationally [SENL^{+}15]. translesion [Sho15-36, TSFP^{+}15]. translocase [CST^{+}17, RDN^{+}19]. translocates [PLH18]. translocation [CCQ^{+}18, GDV19, IKRMN16, SSL^{+}17, TYD^{+}15, WYoS17]. translocon [KDV^{+}15, SCG17]. Transmembrane [GSM^{+}15, GKG^{+}18, SLG^{+}18]. transmigration [NLH^{+}19]. Transmission [SLG^{+}18, KTM19, KOV^{+}16a, KOV^{+}16b, PKN^{+}15]. transmitted [FSF^{+}15]. transport [AGB^{+}19, BPH^{+}19, BMF^{+}18, BCH^{+}17, CPBG19, CDT^{+}19, CNN^{+}17, CGPB17, CZL^{+}15, CHH^{+}15, CCY^{+}19, DOA^{+}17, DDAR^{+}16, GHD^{+}17, GYK^{+}17, GWF17, JERL^{+}15, JNW15, KKC^{+}19, KHRL17, KLHC^{+}18, KOR^{+}19, KOK^{+}19, KJON^{+}17, LCTP17, LE16, LLS^{+}18, LDG^{+}15, Mes16, MHA^{+}19, MWT^{+}16, NiYT^{+}16, NNH17, OI18b, OOT^{+}18, RM19, RFG19, SIO^{+}16, SD19, SMK^{+}18, Sho15-35, SDHC17, VGB^{+}17, VXF^{+}15, XTT^{+}18, YTTH^{+}17, YDM^{+}18, YWW17, YSM^{+}17, ZYL^{+}16]. transported [ADBST^{+}15]. transporter [GWZ^{+}19a, HDA^{+}17, MST^{+}15]. transporters [MYN^{+}17]. transports [SS19, TBK^{+}16, VYB^{+}19]. trap [SAO^{+}17]. TRAPP [RGMM18]. TrappC12 [MHSD^{+}15]. TRAPPII [TF16]. TRAPPIII [TJF18]. trapping [GGC^{+}17]. traps [WWT18]. trash [Blu15b, VR18]. travels [SS19]. TrCP [XWZ^{+}15]. Tre1 [LL17, TCWM18]. treated [MHA^{+}16]. treatment [BS18]. treatments [ZDM^{+}15]. TREX [EMRS^{+}18]. TREX-2 [EMRS^{+}18]. triage [Sed15n]. tricellular [SLM^{+}15]. trichoplein [IGK^{+}16]. TRIF [YSR^{+}18]. TRIF-dependent [YSR^{+}18]. trigger [HHCK19, TNP^{+}15]. triggered [BNB^{+}15, SMK^{+}18]. triggering [DK16, ZB19]. triggers [LHT^{+}19]. triglycerides [KOR^{+}19]. trilayered [KJZ^{+}19]. TRIM [KJC^{+}15]. TRIM-mediated [KJC^{+}15]. TRIM3 [SVD^{+}15]. TRIM37 [WXFS17]. TRIM’d [Sho15x]. trimer [KJON^{+}17]. trimeric [LMPG^{+}15]. TRIMming [Sho15-68]. Trio [DKM^{+}15, KLS^{+}19]. trip [Les15m]. TRIP13 [NHCB15]. Tripathi [O’D19a]. triphosphate [MOM^{+}18]. TrkA [FKW^{+}17]. TrkB.T1 [FTAB^{+}15]. tRNA [KKP^{+}17]. trogocytosis [GGL^{+}19]. tropomyosin [PKH^{+}19, TBK^{+}16]. tropomyosin-dependent [TBK^{+}16]. trouble [Sed15l, Van19]. TRP [GLS^{+}15]. TRPA1 [SZL^{+}16]. TRPM8 [GGC^{+}17, GLS^{+}15]. TRRAP [WPA^{+}18]. truncated [FTAB^{+}15]. Trypanosome [DW17, BMF^{+}18, DSSF^{+}15]. Trypanosomes [Sho15-69]. TSA [CZW^{+}18]. TSC [RHJW18]. TspanC8 [DCO^{+}12, DCO^{+}16]. TTBK2 [LLY^{+}19, WKM^{+}15]. Tuba1a [BNS^{+}17, Kaw17]. Tuba8 [Kaw17]. Tubby [BhHS^{+}17, Sho17j]. tube [SCK^{+}19, SCK^{+}23]. tuberculosis [LBG^{+}17]. tubular [DDAR^{+}16, OG16]. tubulation [WZR19]. tubule [RHH^{+}18, RGR^{+}18]. tubules [DMS^{+}15, MCCL^{+}15]. Tubulin [CHH^{+}15, GDB^{+}17, Kaw17, BKR^{+}19, BNS^{+}17, BS17b, FBBRCA^{+}18, cLNF^{+}16, MOM^{+}18, MSL16, PTK16, SKZ^{+}18b, Spe17b, SFA^{+}19, Wor19]. tug [SR17a]. tumor [ACG^{+}17, BBMM^{+}16, BBHBFSF18, CNN^{+}17, CAI^{+}15, LRH^{+}15, iNLM^{+}19, PHKY17, QCC^{+}19, SENL^{+}15, TZC^{+}15, TAQ^{+}19, WZC^{+}15, WWY^{+}18]. tumorigenesis [AMT^{+}15, LMC^{+}18, Sho15-40, XWZ^{+}15]. tumors [Les15-27]. TUNEL [Sho15-70]. tunes [BRH^{+}16, TF19]. tuning [MBT16, NCV^{+}16]. tunneling [VZ17]. tunnels [RM19]. turn [cLNF^{+}16]. turnover [CHP^{+}17, JHF^{+}15, JBE^{+}17, JIB^{+}19, KSG^{+}16, LKE15, PKH^{+}19]. turns [Jan18, Sho18f]. TuSC [cLNF^{+}16]. twist [PH18]. Two [KGN^{+}15, RCS^{+}19, SSL^{+}17, SFG^{+}17, Bea16, BMF^{+}18, MYT^{+}16, Mar16b, Pfe16, PKC^{+}16, RGMM18, Sho15q, VGA^{+}15, FA16]. two-step [Mar16b]. Type [HGG^{+}17, PD19, Col19, CGBD^{+}17, HLW^{+}15, ITN^{+}17, ISK^{+}15, JCK^{+}19, LBD18, MNL^{+}16, PW19, SPE^{+}17a, TVG^{+}19, TGK^{+}19, WBL^{+}15]. typecasts [Sho15o]. Typhimurium [HGG^{+}17]. tyrosine [CRN^{+}19, CB16, JPF^{+}16, TGQ^{+}17].

U5 [MCM^{+}17]. ubiquinone [VGB^{+}17]. ubiquitin [BHS^{+}16, CYT^{+}18, GP17, HESKK15a, HESKK15b, HSN^{+}16, KJ16, LKE15, O’D19e, OKK^{+}15, PNE^{+}19, SvZS^{+}16, SSV^{+}18, SL19, DS16a, GAS^{+}15, SZE19, SVD^{+}15, SMA^{+}19]. ubiquitinated [WLJ16]. Ubiquitination [LLL^{+}18, CHL^{+}19, Hu15, KSM^{+}18, MCS^{+}15, WWZ^{+}17]. Ubiquitylation [Les16h]. Ubr5 [JHF^{+}15]. Ubr5-mediated [JHF^{+}15]. UBXN2B [LSMG18]. Ugo1 [VKT^{+}15]. ULK1 [HSZ^{+}18, NCV^{+}16, ZZ16]. ultrastructure [BCG^{+}19, CSA19, WYV^{+}19]. unanswered [Pol17].

Unattached [MHA^{+}16, NHCB15, RFO^{+}16]. UNC [LFK^{+}17a, Sho17k, SJ16]. UNC-45a [LFK^{+}17a, Sho17k]. UNC-84 [SJ16]. UNC50 [SIBM17]. uncanny [Pow16e]. uncontrolled [MBS^{+}17]. Unconventional [VBL^{+}18, CGBD^{+}17, DCP^{+}19, LHB^{+}18, MLJ^{+}16]. Uncoordinated [YYM^{+}18]. uncoupled [LJ17a]. underlies [DKM^{+}15, SPWM15, YYM^{+}18]. understanding [Inf19b, Jor16a, LS18, RS19, Mar17]. Unearthing [Pow15g]. unequal [LLS^{+}16]. unexpected [Mok16]. unfolding [Sed16e]. Unidirectional [OKN^{+}16]. uniparental [CMA19]. Unipotent [WRGB^{+}15]. Unique [PCK^{+}17, VAB^{+}18]. universally [DBS18]. unlicensed [CNC^{+}18]. unpredictability [LH15]. unscheduled [Les15-32]. unstructured [DMG^{+}19]. Untangling [Sho16-36]. up-regulation [ZCL^{+}15]. uPARAP [JNS^{+}19]. uPARAP-mediated [JNS^{+}19]. upon [KdBKvdK15, LCP^{+}15, LBV^{+}17, LMC^{+}18, MTM^{+}17, SSRG18]. UPR [ITN^{+}17]. upregulated [CHH^{+}15]. ups [ZZ16]. Ups2 [AFO^{+}16, MWT^{+}16]. upstream [FG15, IB19a, IB19b]. uptake [CJS^{+}18, GLL^{+}18a, JNS^{+}19, LLZ^{+}19]. use [Ava18, PLD17, Sho15f, WTB^{+}19]. Usher [SPE^{+}17a]. Using [Les16i, AHS^{+}18, BCG^{+}19, CZW^{+}18, ISL^{+}18, Juh16, SERP16, TCWM18, BA18]. UsnRNP [PMP^{+}17]. USP10 [KPA^{+}20, KPA^{+}16]. Usp16 [ZGZ^{+}15]. USP28 [LDU^{+}16, MAK^{+}16]. USP30 [RDH^{+}19]. USP9X [KMC^{+}19]. utilizing [AKTR18].

v [ZT15, DB15a, TBK^{+}16, YWW17, WHB^{+}18]. v-ATPase [WHB^{+}18]. v-ATPase/ [WHB^{+}18]. vaccinia [PMW18]. vacuole [AKTR18, BPL^{+}18, JJW17, PHA^{+}17, SZE19, Sho15-32, SE18, MST^{+}15]. vacuole/lysosome [JJW17, SE18]. vacuoles [GRU18]. vagaries [Pow16a]. Validating [JW19]. valve [GGF^{+}19, WBL^{+}15]. VAMP7 [DDAR^{+}16, VKJ^{+}15]. VAMP8 [MPH^{+}15]. VAMP8-dependent [MPH^{+}15]. VAPB [CCH^{+}17]. VAPs [HCC^{+}17]. variant [RSC^{+}19]. variants [OOT^{+}18]. varicosity [GJW^{+}17]. varying [OI18b]. Vascular [YGW^{+}17, LWZ^{+}18, VAKB^{+}18]. vasculature [Sho17e]. VASP [LWZ^{+}18]. Vav3 [HNF^{+}18]. Vav3-induced [HNF^{+}18]. Vb [EKP^{+}19]. VCAM [LAMACE^{+}17]. VCAM-1 [LAMACE^{+}17]. VDAC2 [iHMM^{+}17]. VE [CBH^{+}15, DSvNA^{+}15a, DSvNA^{+}15b, GDB^{+}15, JKD^{+}19, Sho15v]. VE-cadherin [CBH^{+}15, DSvNA^{+}15a, DSvNA^{+}15b, GDB^{+}15, JKD^{+}19, Sho15v]. VE-PTP [JKD^{+}19]. VEGF [Sho15-71]. VEGFR2 [CBH^{+}15]. VEGFR3 [CBH^{+}15]. velocity [AGB^{+}19]. Ventura [Cas17b]. versatile [GSC^{+}16]. versus [DS16a, ST16a]. vertebrate [HKT^{+}17, SSdLA^{+}15]. vertebrates [OTG^{+}18]. vesicle [BLPV^{+}17, JHC^{+}16, Juh16, KCB^{+}16, KMK^{+}17a, KMK^{+}17b, Sho15-35, TF16, UBBSM15]. vesicles [BZG^{+}17, DB15b, JPD^{+}16, KMJ^{+}18, MLMF16, NNH17, RM19, RNP^{+}17, SES^{+}19, SZSS18, SKN19, SCP^{+}17, SW18, ZJM^{+}17]. vessels [FG16, Pow16a, TCWM18]. VHL [LAMACE^{+}17]. VI [RHH^{+}18]. via [ASZ^{+}18, ASM^{+}15, BKH^{+}15, BJB^{+}18, CBAP^{+}17, CAKL16, CMMB^{+}15, CRS^{+}17, CCY^{+}19, DWH^{+}17a, DDAR^{+}16, DCB^{+}15, FBBRCA^{+}18, FRP^{+}17, GCZ^{+}19, GGC^{+}17, HKH16, HLLK19, HGM^{+}19, JKD^{+}19, KG15, KSM^{+}18, KLS^{+}19, KD17b, KJF^{+}18, LAMACE^{+}17, LSMG18, LHY^{+}19, LMPG^{+}15, LJ17b, LZD^{+}16, LZH^{+}18, LSS^{+}15, MFVS18, MBF17, MBC^{+}19, NHA^{+}19, OMKM16, PLH18, PLS^{+}15, PAC^{+}15, PMRMS17, PKC^{+}16, PBL^{+}16, RKK^{+}18, SS19, SOII18, SKZ^{+}18a, SK18b, SYK^{+}17, TTU^{+}17, UBBSM15, WWZ^{+}18, WSP^{+}18, XSJ18, YVM18, YKO^{+}16, vDMR^{+}19]. viability [BRY^{+}19, SFZ^{+}17]. Victoria [O’D18g]. view [BGKL15, O’D17c, Sho15-55]. views [CJ16]. Vimentin [PVP^{+}19]. viral [DAG^{+}15, NPÖ^{+}17, SD16b, VML^{+}17]. viroporin [LgYL^{+}18]. viroporin-like [LgYL^{+}18]. virulence [BHDK17]. virus [IZZ^{+}18, MLJ^{+}16, Nie19, PMW18, TGK^{+}19]. viruses [Pow15h]. viscoelastic [CRZ^{+}16]. viscoelasticity [CGT16]. vision [FLS^{+}16, GLS^{+}17, Sho15-70]. Visionary [Bev17]. visits [GM18]. Visualization [GKG^{+}18, KOK^{+}19, PBS^{+}16, TALR^{+}19, DCP^{+}19]. Visualizing [LT18]. Vitamin [dlFEvW^{+}15, Sho15-72]. vitro [ASPY^{+}16, DV16, GRU18, NWP^{+}16]. vivo [BMP^{+}18, BS17a, BPW15, DB15a, KST^{+}17a, KST^{+}17b, MMW^{+}19, MPA^{+}16, MPMP16, QZX19, SLD^{+}15, SMN^{+}16]. VKORC1 [FLG^{+}15, FLG^{+}19]. VLDLs [SNOBM16]. voltage [RPMC^{+}16, WZG^{+}17]. voltage-dependent [WZG^{+}17]. volume [CTN^{+}19, KO19, LJ17a, MWB^{+}19, PGRY^{+}19, SKO^{+}15, ZZMC^{+}15]. VopF [BHDK17]. VopL [BHDK17, VQ17]. VopL/F [VQ17]. Vpr [DLBMA^{+}15]. VPS [LLL^{+}18]. VPS-34 [LLL^{+}18]. Vps1 [VAB^{+}18]. VPS13 [GY18, BDK^{+}18, LPWK15, MP17b, PHA^{+}17]. Vps13-Mcp1 [PHA^{+}17]. VPS13A [KLHC^{+}18]. VPS13C [KLHC^{+}18]. Vps13p [DOA^{+}17]. Vps34 [NHG^{+}18]. VPS35 [ZWS^{+}16]. VPS37A [TLH^{+}19]. vulnerabilities [CSS^{+}18].

wall [DBG^{+}15, LLZ^{+}19]. Wallerian [WTSA17]. Wang [Inf19b]. Want [Sho17l]. war [SR17a]. WASH [MBS^{+}18]. WASP [DATI18, FLLM17]. Watching [FJ17, MSvO17, SK16a, Lov18]. water [LRD19]. wave [IYP^{+}18, Sør17, NIS^{+}16, SBC^{+}16a, SBC^{+}16b, CRA^{+}19]. Waves [Roy16, DD18, GTMZ^{+}15, Wu17]. way [BH15, Les15-28, O’D18e, Sed15e, Sed15d, Sed15l, Sed15s, Sho15-35, Sho15-53]. wayward [Les15g]. Wdr1-deficient [BRY^{+}19]. Wdr8 [YIT15]. WDR81 [LLW^{+}17]. WDR91 [LXJ^{+}17]. Wdr92 [zLSSS^{+}18]. weaves [Sho17f]. Weaving [Inf19a]. web [OPP^{+}18]. Wee1 [AOL^{+}18, SLW^{+}18]. Weibel [MPW^{+}19]. well [Zhu17]. well-known [Zhu17]. WH2 [ISL^{+}18]. Which [VQ17]. whips [ST16b]. Who [Kti19]. Whole [SSdLA^{+}15]. Whole-proteome [SSdLA^{+}15]. wide [BPH^{+}19, MHI^{+}18, PST18, SIBM17]. widens [KKD^{+}16]. widespread [HKG17, LSMZ^{+}18]. width [AKD^{+}17]. wild [CGBD^{+}17]. wild-type [CGBD^{+}17]. wing [FLG^{+}18, TNK18]. wins [Sho16a, ST16a]. wires [NPC17]. wiring [WLC^{+}17]. WISp39 [HBDW^{+}15]. wither [SD17]. within [BYMS^{+}19, FZD^{+}19, Ger15, HHT^{+}16, IG15, JOJG16, KWB^{+}15, KML^{+}15, KOK^{+}19, LBG^{+}17, PST18, RSCR15, RNP^{+}17, TALR^{+}19]. without [GM16, ZB19]. Wnt [BJB^{+}18, CAKL16, CRA^{+}19, HB18, LRH^{+}15, LWF^{+}15]. Wnt-dependent [LRH^{+}15]. women [IO18]. won’t [Les15y]. Words [Sed15u]. works [Sho15b]. wound [CPP^{+}18, HLHFG15, LLC^{+}17, MCGM15a, MCGM15b, NVP17, ZPT^{+}15]. wounds [TCWM18]. Wrapping [Boh18].

X [CAI^{+}15, OSW^{+}17, dlFEvW^{+}15]. XBP1 [TSK^{+}18, TSK^{+}19]. Xenopus [GHS16a, GHS16b, GSC^{+}16, KZW^{+}18, MH15, RBR19]. Xiaochen [Inf19b]. XLF [CQB^{+}19, GRB19]. XPA [CR18]. Xpo7 [APK^{+}18]. Xrp1 [MCH^{+}18].

YAP [FKL^{+}18a, FKL^{+}18b, GDB^{+}15, MCD^{+}19, NW19, PGRY^{+}19, XSJ18, MpDN^{+}17]. YAP/TAZ [NW19]. YAP1 [GCC^{+}18]. YAP1/TAZ [GCC^{+}18]. Yaron [O’D17g]. YB [SENL^{+}15]. YB-1 [SENL^{+}15]. years [Hal15, HK15]. Yeast [CGY^{+}19, KdBKvdK15, LPRW17, Sho16-37, ADBST^{+}15, BYUJ17, DBG^{+}15, DTW^{+}16, GBK^{+}17, HESKK15a, HESKK15b, HGL^{+}17, LCP^{+}15, LK17, LHA^{+}15, LSJY15, LDG^{+}15, ML15b, MSW^{+}07, MSW^{+}17, MKD^{+}18, MKA^{+}19, NDRJ15, OCS15, SPGB^{+}17, SBR^{+}15, SPK^{+}18, Sho15q, SHO^{+}18g, SLD^{+}15, TBK^{+}16, WTB^{+}19, YSW^{+}15, YAHH15, YIT15]. yield [AZ19]. Yki [CV19]. YKT6 [MJN^{+}18, BPL^{+}18, GRU18]. Yme1 [WLJ18]. YME1L [RDN^{+}19]. Yorkie [SRF19, SCK^{+}19, SCK^{+}23]. young [SK19]. Ypt1 [TJF18, WDM^{+}15]. Ypt1/Rab1 [WDM^{+}15]. Yurt [GPPJ^{+}18, PVP18].

Z [Har16]. Zas1 [SPK^{+}18]. Zds1 [JRH^{+}16]. Zds1/ [JRH^{+}16]. Zds2 [JRH^{+}16]. ZEB2 [DAG^{+}15, OBY^{+}15, vHGD^{+}15]. Zebrafish [BS17a, CED^{+}15, GGF^{+}19, IKRMN16, MSK^{+}19]. Zen [Sed15r]. ZMYND8 [GCA^{+}17]. ZNRF1 [WFOA15]. ZO [TCD^{+}15]. ZO-1 [CRPSC^{+}19, TCD^{+}15]. zone [DER^{+}18, DCF^{+}17, GBB^{+}19, RDO^{+}15, SES^{+}19, SSPD15, Sho15-67, THM^{+}19, VPD^{+}16, YNN18]. zones [CL19, KMK^{+}17a, KMK^{+}17b]. zonula [CAP^{+}16]. ZRF1 [GCW^{+}16]. Zurzolo [Sed15d]. Zw10 [GPS^{+}17]. Zwilch [GPS^{+}17]. zygote [Ver16].

References

Alfaro-Aco:2017:SAR

[AATP17] Raymundo Alfaro-Aco, Akanksha Thawani, and Sabine Petry. Structural analysis of the role of TPX2 in branching microtubule nucleation. Journal of Cell Biology, 216(4):983–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/4/
983.

Alford:2018:QHA

[AB18] Brian D. Alford and Onn Brandman. Quantification of Hsp90 availability reveals differential coupling to the heat shock response. Journal of Cell Biology, 217(11):3809–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/
11/3809.

Andress:2016:SCB

[ABF^{+}16] Arlise Andress, Yanxia Bei, Bryan R. Fonslow, Ritika Giri, Yilong Wu, John R. Yates, and Richard W. Carthew. Spindle-e cycling between nuage and cytoplasm is controlled by Qin and PIWI proteins. Journal of Cell Biology, 213(2):201–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/213/2/
201.

Aparicio:2016:MCB

[ABGG16] Tomas Aparicio, Richard Baer, Max Gottesman, and Jean Gautier. MRN, CtIP, and BRCA1 mediate repair of topoisomerase II–DNA adducts. Journal of Cell Biology, 212(4): 399–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/4/399.

Adriaans:2019:PPD

[ABP^{+}19] Ingrid E. Adriaans, Angika Basant, Bas Ponsioen, Michael Glotzer, and Susanne M. A. Lens. PLK1 plays dual roles in centralspindlin regulation during cytokinesis. Journal of Cell Biology, 218(4):1250–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/4/1250.

Alleva:2017:RCM

[ABPS17] Benjamin Alleva, Nathan Balukoff, Amy Peiper, and Sarit Smolikove. Regulating chromosomal movement by the cochaperone FKB-6 ensures timely pairing and synapsis. Journal of Cell Biology, 216(2):393–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/2/393.

Aulas:2015:GPS

[ACG^{+}15] Anaïs Aulas, Guillaume Caron, Christos G. Gkogkas, Nguyen-Vi Mohamed, Laurie Destroismaisons, Nahum Sonenberg, Nicole Leclerc, J. Alex Parker, and Christine Vande Velde. G3BP1 promotes stress-induced RNA granule interactions to preserve polyadenylated mRNA. Journal of Cell Biology, 209 (1):73–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/1/73.

Attieh:2017:CAF

[ACG^{+}17] Youmna Attieh, Andrew G. Clark, Carina Grass, Sophie Richon, Marc Pocard, Pascale Mariani, Nadia Elkhatib, Timo Betz, Basile Gurchenkov, and Danijela Matic Vignjevic. Cancer-associated fibroblasts lead tumor invasion through integrin-β3-dependent fibronectin assembly. Journal of Cell Biology, 216(11):3509–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/11/3509.

Auckland:2017:CKP

[ACRM17] Philip Auckland, Nicholas I. Clarke, Stephen J. Royle, and Andrew D. McAinsh. Congressing kinetochores progressively load Ska complexes to prevent force-dependent detachment. Journal of Cell Biology, 216(6):1623–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/6/1623.

Allen:2018:MBB

[AD18] Nicola J. Allen and Richard Daneman. In memoriam: Ben Barres. Journal of Cell Biology, 217(2):435–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/217/2/435.

Aronov:2015:PEM

[ADBST^{+}15] Stella Aronov, Saray Dover-Biterman, Edith Suss-Toby, Michael Shmoish, Lea Duek, and Mordechai Choder. Pheromoneencoding mRNA is transported to the yeast mating projection by specific RNP granules. Journal of Cell Biology, 209 (6):829–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/6/829.

Allison:2017:DEC

[AEP^{+}17] Rachel Allison, James R. Edgar, Guy Pearson, Tania Rizo, Timothy Newton, Sven Günther, Fiamma Berner, Jennifer Hague, James W. Connell, Jürgen Winkler, Jennifer Lippincott-Schwartz, Christian Beetz, Beate Winner, and Evan Reid. Defects in ER–endosome contacts impact lysosome function in hereditary spastic paraplegia. Journal of Cell Biology, 216(5):1337–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/5/1337.

Aaltonen:2016:MPT

[AFO^{+}16] Mari J. Aaltonen, Jonathan R. Friedman, Christof Osman, Bénédicte Salin, Jean-Paul di Rago, Jodi Nunnari, Thomas Langer, and Takashi Tatsuta. MICOS and phospholipid transfer by Ups2–Mdm35 organize membrane lipid synthesis in mitochondria. Journal of Cell Biology, 213(5):525–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/213/5/525.

Asai:2019:ABK

[AFT^{+}19] Yuichiro Asai, Koh Fukuchi, Yuji Tanno, Saki Koitabashi-Kiyozuka, Tatsuyuki Kiyozuka, Yuko Noda, Rieko Matsumura, Tetsuo Koizumi, Atsushi Watanabe, Kyosuke Nagata, Yoshinori Watanabe, and Yasuhiko Terada. Aurora B kinase activity is regulated by SET/TAF1 on Sgo2 at the inner centromere. Journal of Cell Biology, 218(10):3223–??, October 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/10/3223.

Agrawal:2016:DRI

[AFXS16] Gaurav Agrawal, Scott N. Fassas, Zhi-Jie Xia, and Suresh Subramani. Distinct requirements for intra–ER sorting and budding of peroxisomal membrane proteins from the ER. Journal of Cell Biology, 212(3):335–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/3/335.

Atkins:2019:FAK

[AGB^{+}19] Melody Atkins, Laïla Gasmi, Valérie Bercier, Céline Revenu, Filippo Del Bene, Jamilé Hazan, and Coralie Fassier. FIGNL1 associates with KIF1Bβ and BICD1 to restrict dynein transport velocity during axon navigation. Journal of Cell Biology, 218(10):3290–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/10/3290.

Aix:2016:PTD

[AGGSF^{+}16] Esther Aix, Óscar Gutiérrez-Gutiérrez, Carlota Sánchez-Ferrer, Tania Aguado, and Ignacio Flores. Postnatal telomere dysfunction induces cardiomyocyte cell-cycle arrest through p21 activation. Journal of Cell Biology, 213(5):571–??, June 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/213/5/571.

Amini:2015:CEA

[AGL^{+}15] Rana Amini, Eugénie Goupil, Sara Labella, Monique Zetka, Amy S. Maddox, Jean-Claude Labbé, and Nicolas T. Chartier. C. elegans Anillin proteins regulate intercellular bridge stability and germline syncytial organization. Journal of Cell Biology, 209(3):467–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/3/467.

Awadia:2019:SFC

[AHA^{+}19] Sahezeel Awadia, Farah Huq, Torey R. Arnold, Silvia M. Goicoechea, Young Joo Sun, Titus Hou, Gabriel Kreider-Letterman, Paola Massimi, Lawrence Banks, Ernesto J. Fuentes, Ann L. Miller, and Rafael Garcia-Mata. SGEF forms a complex with Scribble and Dlg1 and regulates epithelial junctions and contractility. Journal of Cell Biology, 218(8):2699–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/8/2699.

Amano:2015:KIS

[AHS^{+}15] Mutsuki Amano, Tomonari Hamaguchi, Md. Hasanuzzaman Shohag, Kei Kozawa, Katsuhiro Kato, Xinjian Zhang, Yoshimitsu Yura, Yoshiharu Matsuura, Chikako Kataoka, Tomoki Nishioka, and Kozo Kaibuchi. Kinase-interacting substrate screening is a novel method to identify kinase substrates. Journal of Cell Biology, 209(6):895–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/209/6/895.

Adikes:2018:CMD

[AHS^{+}18] Rebecca C. Adikes, Ryan A. Hallett, Brian F. Saway, Brian Kuhlman, and Kevin C. Slep. Control of microtubule dynamics using an optogenetic microtubule plus end–F-actin cross-linker. Journal of Cell Biology, 217(2):779–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/2/779.

Aird:2016:HOC

[AIK^{+}16] Katherine M. Aird, Osamu Iwasaki, Andrew V. Kossenkov, Hideki Tanizawa, Nail Fatkhutdinov, Benjamin G. Bitler, Linh Le, Gretchen Alicea, Ting-Lin Yang, F. Brad Johnson, Ken ichi Noma, and Rugang Zhang. HMGB2 orchestrates the chromatin landscape of senescence-associated secretory phenotype gene loci. Journal of Cell Biology, 215(3):325–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/215/3/
325.

Azad:2018:PDE

[AIS^{+}18] Gajendra Kumar Azad, Kenji Ito, Badi Sri Sailaja, Alva Biran, Malka Nissim-Rafinia, Yasuhiro Yamada, David T. Brown, Takumi Takizawa, and Eran Meshorer. PARP1-dependent eviction of the linker histone H1 mediates immediate early gene expression during neuronal activation. Journal of Cell Biology, 217(2):473–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/2/473.

Avenarius:2017:HCP

[AKD^{+}17] Matthew R. Avenarius, Jocelyn F. Krey, Rachel A. Dumont, Clive P. Morgan, Connor B. Benson, Sarath Vijayakumar, Christopher L. Cunningham, Deborah I. Scheffer, David P. Corey, Ulrich Müller, Sherri M. Jones, and Peter G. Barr-Gillespie. Heterodimeric capping protein is required for stereocilia length and width regulation. Journal of Cell Biology, 216(11):3861–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/11/3861.

Arasaki:2018:LRP

[AKTR18] Kohei Arasaki, Hana Kimura, Mitsuo Tagaya, and Craig R. Roy. Legionella remodels the plasma membrane–derived vacuole by utilizing exocyst components as tethers. Journal of Cell Biology, 217(11):3863–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/11/3863.

Agbu:2018:SGR

[ALLA18] Stephanie O. Agbu, Yinwen Liang, Aimin Liu, and Kathryn V. Anderson. The small GTPase RSG1 controls a final step in primary cilia initiation. Journal of Cell Biology, 217(1):413–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/1/413.

An:2017:BDS

[ALY^{+}17] Ming-Xin An, Si Li, Han-Bing Yao, Chao Li, Jia-Mei Wang, Jia Sun, Xin-Yu Li, Xiao-Na Meng, and Hua-Qin Wang. BAG3 directly stabilizes Hexokinase 2 mRNA and promotes aerobic glycolysis in pancreatic cancer cells. Journal of Cell Biology, 216(12):4091–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/12/4091.

Anderson:2017:PAG

[AMS^{+}17] Garret R. Anderson, Stephan Maxeiner, Richard Sando, Theodoros Tsetsenis, Robert C. Malenka, and Thomas C. Südhof. Postsynaptic adhesion GPCR latrophilin-2 mediates target recognition in entorhinal-hippocampal synapse assembly. Journal of Cell Biology, 216(11):3831–??, November 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/11/3831.

Allam:2015:ENP

[AMT^{+}15] Ramanjaneyulu Allam, Michel H. Maillard, Aubry Tardivel, Vijaykumar Chennupati, Hristina Bega, Chi Wang Yu, Dominique Velin, Pascal Schneider, and Kendle M. Maslowski. Epithelial NAIPs protect against colonic tumorigenesis. Journal of Cell Biology, 208(6):??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/6/2086OIA28.

Azoitei:2019:SDG

[ANM^{+}19] Mihai L. Azoitei, Jungsik Noh, Daniel J. Marston, Philippe Roudot, Christopher B. Marshall, Timothy A. Daugird, Sidney L. Lisanza, María-José Sandí, Mitsu Ikura, John Sondek, Robert Rottapel, Klaus M. Hahn, and Gaudenz Danuser. Spatiotemporal dynamics of GEF–H1 activation controlled by microtubule- and Src-mediated pathways. Journal of Cell Biology, 218(9):3077–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/9/3077.

Allard:2018:CSD

[AOL^{+}18] Corey A. H. Allard, Hannah E. Opalko, Ko-Wei Liu, Uche Medoh, and James B. Moseley. Cell size-dependent regulation of Wee1 localization by Cdr2 cortical nodes. Journal of Cell Biology, 217(5):1589–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/5/1589.

Alfonso-Perez:2019:MDR

[APHH^{+}19] Tatiana Alfonso-Pérez, Daniel Hayward, James Holder, Ulrike Gruneberg, and Francis A. Barr. MAD1-dependent recruitment of CDK1-CCNB1 to kinetochores promotes spindle checkpoint signaling. Journal of Cell Biology, 218(4):1108–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/4/
1108.

Aksu:2018:XBS

[APK^{+}18] Metin Aksu, Tino Pleiner, Samir Karaca, Christin Kappert, Heinz-Jürgen Dehne, Katharina Seibel, Henning Urlaub, Markus T. Bohnsack, and Dirk Görlich. Xpo7 is a broadspectrum exportin and a nuclear import receptor. Journal of Cell Biology, 217(7):2329–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/7/2329.

Ando:2017:NDP

[APS^{+}17] Kiyohiro Ando, Melissa J. Parsons, Richa B. Shah, Chloé I. Charendoff, Sheré L. Paris, Peter H. Liu, Sara R. Fassio, Brittany A. Rohrman, Ruth Thompson, Andrew Oberst, Samuel Sidi, and Lisa Bouchier-Hayes. NPM1 directs PIDDosomedependent caspase-2 activation in the nucleolus. Journal of Cell Biology, 216(6):1795–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/6/1795.

Aitchison:2015:IC

[AR15] John D. Aitchison and Michael P. Rout. The interactome challenge. Journal of Cell Biology, 211(4):729–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/211/4/729.

Abad:2019:BNI

[ARB^{+}19] Maria A. Abad, Jan G. Ruppert, Lana Buzuk, Martin Wear, Juan Zou, Kim M. Webb, David A. Kelly, Philipp Voigt, Juri Rappsilber, William C. Earnshaw, and A. Arockia Jeyaprakash. Borealin–nucleosome interaction secures chromosome association of the chromosomal passenger complex. Journal of Cell Biology, 218(12):3912–3925, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/218/12/3912/132542/Borealin-
nucleosome-interaction-secures-chromosome.

Aweida:2018:GPC

[ARV^{+}18] Dina Aweida, Inga Rudesky, Alexandra Volodin, Eitan Shimko, and Shenhav Cohen. GSK3-β promotes calpain-1–mediated desmin filament depolymerization and myofibril loss in atrophy. Journal of Cell Biology, 217(10):3698–??, October 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/10/3698.

Atay:2017:STS

[AS17] Oguzhan Atay and Jan M. Skotheim. Spatial and temporal signal processing and decision making by MAPK pathways. Journal of Cell Biology, 216(2):317–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/2/317.

Artym:2015:DFC

[ASM^{+}15] Vira V. Artym, Stephen Swatkoski, Kazue Matsumoto, Catherine B. Campbell, Ryan J. Petrie, Emilios K. Dimitriadis, Xin Li, Susette C. Mueller, Thomas H. Bugge, Marjan Gucek, and Kenneth M. Yamada. Dense fibrillar collagen is a potent inducer of invadopodia via a specific signaling network. Journal of Cell Biology, 208(3):331–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/3/331.

Aratyn-Schaus:2016:CPS

[ASPY^{+}16] Yvonne Aratyn-Schaus, Francesco S. Pasqualini, Hongyan Yuan, Megan L. McCain, George J. C. Ye, Sean P. Sheehy, Patrick H. Campbell, and Kevin Kit Parker. Coupling primary and stem cell-derived cardiomyocytes in an in vitro model of cardiac cell therapy. Journal of Cell Biology, 212(4): 389–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/4/389.

Agarwal:2018:CSK

[ASZ^{+}18] Shivangi Agarwal, Kyle Paul Smith, Yizhuo Zhou, Aussie Suzuki, Richard J. McKenney, and Dileep Varma. Cdt1 stabilizes kinetochore–microtubule attachments via an Aurora B kinase-dependent mechanism. Journal of Cell Biology, 217(10): 3446–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/10/3446.

Achuthankutty:2019:REM

[ATH^{+}19] Divya Achuthankutty, Roshan Singh Thakur, Peter Haahr, Saskia Hoffmann, Alexandros P. Drainas, Anna H. Bizard, Joachim Weischenfeldt, Ian D. Hickson, and Niels Mailand. Regulation of ETAA1-mediated ATR activation couples DNA replication fidelity and genome stability. Journal of Cell Biology, 218(12):3943–3953, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

https://rupress.org/jcb/article/218/12/3943/132516/
Regulation-of-ETAA1-mediated-ATR-activation.

Aksu:2019:SBN

[ATRG19] Metin Aksu, Sergei Trakhanov, Arturo Vera Rodriguez, and Dirk Görlich. Structural basis for the nuclear import and export functions of the biportin Pdr6/Kap122. Journal of Cell Biology, 218(6):1839–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/6/1839.

An:2019:SGR

[ATS19] Haiyan An, Jing Tong Tan, and Tatyana A. Shelkovnikova. Stress granules regulate stress-induced paraspeckle assembly. Journal of Cell Biology, 218(12):4127–4140, December 2019.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/218/12/
4127/132514/Stress-granules-regulate-stress-induced.

Ahtiainen:2016:EES

[AUTM16] Laura Ahtiainen, Isa Uski, Irma Thesleff, and Marja L. Mikkola. Early epithelial signaling center governs tooth budding morphogenesis. Journal of Cell Biology, 214(6):753– ??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/214/6/753.

Avasthi:2018:CMM

[Ava18] Prachee Avasthi. Can microtubule motors use every available track? Journal of Cell Biology, 217(12):4055–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/
12/4055.

Akhmanova:2016:TSR

[AvdH16] Anna Akhmanova and Sander van den Heuvel. Tipping the spindle into the right position. Journal of Cell Biology, 213 (3):293–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/3/293.

Amarh:2018:DRM

[AWL18] Vincent Amarh, Martin A. White, and David R. F. Leach. Dynamics of RecA-mediated repair of replication-dependent DNA breaks. Journal of Cell Biology, 217(7):2299–??, July 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/7/2299.

Albrecht:2016:NCM

[AWS^{+}16] David Albrecht, Christian M. Winterflood, Mohsen Sadeghi, Thomas Tschager, Frank Noé, and Helge Ewers. Nanoscopic compartmentalization of membrane protein motion at the axon initial segment. Journal of Cell Biology, 215(1):37–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/215/1/

Aydogan:2018:HCS

[AWS^{+}18] Mustafa G. Aydogan, Alan Wainman, Saroj Saurya, Thomas L. Steinacker, Anna Caballe, Zsofia A. Novak, Janina Baumbach, Nadine Muschalik, and Jordan W. Raff. A homeostatic clock sets daughter centriole size in flies. Journal of Cell Biology, 217 (4):1233–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/4/1233.

Alper:2019:RFC

[AZ19] Joshua Alper and Marija Zanic. Resistance is futile: Centering forces yield for asymmetric cell division. Journal of Cell Biology, 218(3):727–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/3/727.

Albrecht:2015:GPD

[AZS^{+}15] Lauren V. Albrecht, Lichao Zhang, Jeffrey Shabanowitz, Enkhsaikhan Purevjav, Jeffrey A. Towbin, Donald F. Hunt, and Kathleen J. Green. GSK3- and PRMT-1-dependent modifications of desmoplakin control desmoplakin–cytoskeleton dynamics. Journal of Cell Biology, 208(5):597–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/208/5/597.

Boczek:2018:ODF

[BA18] Edgar E. Boczek and Simon Alberti. One domain fits all: Using disordered regions to sequester misfolded proteins. Journal of Cell Biology, 217(4):1173–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/4/1173.

Brown:2017:IPR

[BAGM17] Caitlin W. Brown, John J. Amante, Hira Lal Goel, and Arthur M. Mercurio. The α6β4 integrin promotes resistance to ferroptosis. Journal of Cell Biology, 216(12):4287–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/12/4287.

Bruno:2016:SRE

[BBC^{+}16] Joanne Bruno, Alexandria Brumfield, Natasha Chaudhary, David Iaea, and Timothy E. McGraw. SEC16A is a RAB10 effector required for insulin-stimulated GLUT4 trafficking in adipocytes. Journal of Cell Biology, 214(1):61–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/214/1/61.

Broders-Bondon:2018:MTP

[BBHBFSF18] Florence Broders-Bondon, Thanh Huong Nguyen Ho-Bouldoires, Maria-Elena Fernandez-Sanchez, and Emmanuel Farge. Mechanotransduction in tumor progression: The dark side of the force. Journal of Cell Biology, 217(5):1571–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/5/1571.

Bersuker:2016:PMS

[BBK16] Kirill Bersuker, Michael Brandeis, and Ron R. Kopito. Protein misfolding specifies recruitment to cytoplasmic inclusion bodies. Journal of Cell Biology, 213(2):229–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/213/2/229.

Biname:2016:CAM

[BBMM^{+}16] Fabien Binamé, Aurélien Bidaud-Meynard, Laure Magnan, Léo Piquet, Bertille Montibus, Anne Chabadel, Frédéric Saltel, Valérie Lagrée, and Violaine Moreau. Cancer-associated mutations in the protrusion-targeting region of p190RhoGAP impact tumor cell migration. Journal of Cell Biology, 214(7): 859–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/214/7/859.

Beaven:2017:RNR

[BBS^{+}17] Robin Beaven, Ricardo Nunes Bastos, Christos Spanos, Pierre Romé, C. Fiona Cullen, Juri Rappsilber, Régis Giet, Gohta Goshima, and Hiroyuki Ohkura. 14-3-3 regulation of Ncd reveals a new mechanism for targeting proteins to the spindle in oocytes. Journal of Cell Biology, 216(10):3029–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/10/3029.

Bledzka:2016:KDB

[BBSA^{+}16] Kamila Bledzka, Katarzyna Bialkowska, Khalid Sossey-Alaoui, Julia Vaynberg, Elzbieta Pluskota, Jun Qin, and Edward F.

Plow. Kindlin-2 directly binds actin and regulates integrin outside-in signaling. Journal of Cell Biology, 213(1):97–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
213/1/97.

Beirowski:2016:ADL

[BBW16] Bogdan Beirowski, Elisabetta Babetto, and Lawrence Wrabetz. Axon degeneration: Linking axonal bioenergetics to myelin. Journal of Cell Biology, 215(4):437–??, November 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/215/4/437.

Bass:2019:QPR

[BC19] Thomas E. Bass and David Cortez. Quantitative phosphoproteomics reveals mitotic function of the ATR activator ETAA1. Journal of Cell Biology, 218(4):1235–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/4/1235.

Bykov:2019:HTU

[BCG^{+}19] Yury S. Bykov, Nir Cohen, Natalia Gabrielli, Hetty Manenschijn, Sonja Welsch, Petr Chlanda, Wanda Kukulski, Kiran R. Patil, Maya Schuldiner, and John A. G. Briggs. Highthroughput ultrastructure screening using electron microscopy and fluorescent barcoding. Journal of Cell Biology, 218(8): 2797–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/8/2797.

Bockler:2017:FFT

[BCH^{+}17] Stefan Böckler, Xenia Chelius, Nadine Hock, Till Klecker, Madita Wolter, Matthias Weiss, Ralf J. Braun, and Benedikt Westermann. Fusion, fission, and transport control asymmetric inheritance of mitochondria and protein aggregates. Journal of Cell Biology, 216(8):2481–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/8/2481.

Burdyniuk:2018:FAN

[BCM^{+}18] Mariia Burdyniuk, Andrea Callegari, Masashi Mori, François Nédélec, and Péter Lénárt. F-actin nucleated on chromosomes coordinates their capture by microtubules in oocyte meiosis.

Journal of Cell Biology, 217(8):2661–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/8/2661.

Bondaz:2019:CPD

[BCMG19] Alexandra Bondaz, Luca Cirillo, Patrick Meraldi, and Monica Gotta. Cell polarity-dependent centrosome separation in the C. elegans embryo. Journal of Cell Biology, 218(12):4112–4126, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/218/12/4112/132530/Cell-polarity-dependent-
centrosome-separation-in.

Byrum:2019:MRT

[BCMM^{+}19] Andrea K. Byrum, Denisse Carvajal-Maldonado, Miranda C. Mudge, David Valle-Garcia, Mona C. Majid, Romil Patel, Mathew E. Sowa, Steven P. Gygi, J. Wade Harper, Yang Shi, Alessandro Vindigni, and Nima Mosammaparast. Mitotic regulators TPX2 and Aurora A protect DNA forks during replication stress by counteracting 53BP1 function. Journal of Cell Biology, 218(2):422–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/2/422.

Bury:2017:PAC

[BCS^{+}17] Leah Bury, Paula A. Coelho, Angela Simeone, Samantha Ferries, Claire E. Eyers, Patrick A. Eyers, Magdalena Zernicka-Goetz, and David M. Glover. Plk4 and Aurora A cooperate in the initiation of acentriolar spindle assembly in mammalian oocytes. Journal of Cell Biology, 216(11):3571–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/
11/3571.

Barry:2015:OSS

[BDAW15] David J. Barry, Charlotte H. Durkin, Jasmine V. Abella, and Michael Way. Open source software for quantification of cell migration, protrusions, and fluorescence intensities. Journal of Cell Biology, 209(1):163–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/209/1/163.

Bean:2018:COS

[BDK^{+}18] Björn D. M. Bean, Samantha K. Dziurdzik, Kathleen L. Kolehmainen, Claire M. S. Fowler, Waldan K. Kwong, Leslie I. Grad, Michael Davey, Cayetana Schluter, and Elizabeth Conibear. Competitive organelle-specific adaptors recruit Vps13 to membrane contact sites. Journal of Cell Biology, 217(10):3593–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/10/3593.

Bentley:2015:NAR

[BDLB15] Marvin Bentley, Helena Decker, Julie Luisi, and Gary Banker. A novel assay reveals preferential binding between Rabs, kinesins, and specific endosomal subpopulations. Journal of Cell Biology, 208(3):273–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/208/3/273.

Brace:2019:CSC

[BDW19] Jennifer L. Brace, Matthew D. Doerfler, and Eric L. Weiss. A cell separation checkpoint that enforces the proper order of late cytokinetic events. Journal of Cell Biology, 218(1):150–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/1/150.

Bruurs:2015:AMS

[BDZ^{+}15] Lucas J. M. Bruurs, Lisa Donker, Susan Zwakenberg, Fried J. Zwartkruis, Harry Begthel, A. S. Knisely, George Posthuma, Stan F. J. van de Graaf, Coen C. Paulusma, and Johannes L. Bos. ATP8B1-mediated spatial organization of Cdc42 signaling maintains singularity during enterocyte polarization. Journal of Cell Biology, 210(7):1055–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/210/7/1055.

Beaudoin:2016:MCP

[Bea16] Gerard M. J. Beaudoin. Mosaic cellular patterning in the nose: Adhesion molecules give their two scents. Journal of Cell Biology, 212(5):495–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/5/495.

Bevis:2017:SLV

[Bev17] Brooke J. Bevis. Susan Lindquist: Visionary scientist and peerless mentor. Journal of Cell Biology, 216(1):5–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/216/1/5.

Bagonis:2019:APG

[BFPD19] Maria M. Bagonis, Ludovico Fusco, Olivier Pertz, and Gaudenz Danuser. Automated profiling of growth cone heterogeneity defines relations between morphology and motility. Journal of Cell Biology, 218(1):350–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/1/350.

Bertero:2019:CCD

[BFS^{+}19] Alessandro Bertero, Paul A. Fields, Alec S. T. Smith, Andrea Leonard, Kevin Beussman, Nathan J. Sniadecki, Deok-Ho Kim, Hung-Fat Tse, Lil Pabon, Jay Shendure, William S. Noble, and Charles E. Murry. Chromatin compartment dynamics in a haploinsufficient model of cardiac laminopathy. Journal of Cell Biology, 218(9):2919–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/9/2919.

Branzei:2018:SRF

[BG18] Dana Branzei and Michele Giannattasio. SIRFing the replication fork: Assessing protein interactions with nascent DNA. Journal of Cell Biology, 217(4):1177–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/4/1177.

Besteiro:2019:EEP

[BG19] Marina Alejandra González Besteiro and Vanesa Gottifredi. ETAA1 ensures proper chromosome segregation: a matter of S phase or mitosis? Journal of Cell Biology, 218(12):3883– 3884, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/218/12/3883/132508/ETAA1-ensures-proper-
chromosome-segregation-A.

Bobkov:2018:CTA

[BGH18] Georg O. M. Bobkov, Nick Gilbert, and Patrick Heun. Centromere transcription allows CENP–A to transit from chromatin association to stable incorporation. Journal of Cell Biology, 217(6):1957–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/6/1957.

Bayless:2016:ALP

[BGJ^{+}16] Brian A. Bayless, Domenico F. Galati, Anthony D. Junker, Chelsea B. Backer, Jacek Gaertig, and Chad G. Pearson. Asymmetrically localized proteins stabilize basal bodies against ciliary beating forces. Journal of Cell Biology, 215(4):457– ??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/215/4/457.

Bezanilla:2015:CDV

[BGKL15] Magdalena Bezanilla, Amy S. Gladfelter, David R. Kovar, and Wei-Lih Lee. Cytoskeletal dynamics: a view from the membrane. Journal of Cell Biology, 209(3):329–??, May 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/209/3/329.

Bond:2015:LSG

[BH15] Michelle R. Bond and John A. Hanover. A little sugar goes a long way: The cell biology of O–GlcNAc. Journal of Cell Biology, 208(7):869–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/7/869.

Backes:2018:TEM

[BHB^{+}18] Sandra Backes, Steffen Hess, Felix Boos, Michael W. Woellhaf, Sabrina Gödel, Martin Jung, Timo Mühlhaus, and Johannes M. Herrmann. Tom70 enhances mitochondrial preprotein import efficiency by binding to internal targeting sequences. Journal of Cell Biology, 217(4):1369–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/4/1369.

Burke:2017:BVF

[BHDK17] Thomas A. Burke, Alyssa J. Harker, Roberto Dominguez, and David R. Kovar. The bacterial virulence factors VopL and VopF nucleate actin from the pointed end. Journal of Cell Biology, 216(5):1267–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/5/1267.

Badgandi:2017:TFP

[BhHS^{+}17] Hemant B. Badgandi, Sun hee Hwang, Issei S. Shimada, Evan Loriot, and Saikat Mukhopadhyay. Tubby family proteins are adapters for ciliary trafficking of integral membrane proteins. Journal of Cell Biology, 216(3):743–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/3/743.

Balachandran:2016:ULC

[BHS^{+}16] Riju S. Balachandran, Cassandra S. Heighington, Natalia G. Starostina, James W. Anderson, David L. Owen, Srividya Vasudevan, and Edward T. Kipreos. The ubiquitin ligase CRL2^{ZYG11} targets cyclin B1 for degradation in a conserved pathway that facilitates mitotic slippage. Journal of Cell Biology, 215(2):151–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/2/151.

Brayson:2018:MTM

[BHS18] Daniel Brayson, Chin Yee Ho, and Catherine M. Shanahan. Muscle tensions merge to cause a DNA replication crisis. Journal of Cell Biology, 217(6):1891–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/6/1891.

Bonner:2019:EAB

[BHS^{+}19] Mary Kate Bonner, Julian Haase, Jason Swinderman, Hyunmi Halas, Lisa M. Miller Jenkins, and Alexander E. Kelly. Enrichment of Aurora B kinase at the inner kinetochore controls outer kinetochore assembly. Journal of Cell Biology, 218(10): 3237–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/10/3237.

Bernkopf:2018:PRD

[BJB^{+}18] Dominic B. Bernkopf, Kowcee Jalal, Martina Brückner, Karl X. Knaup, Marc Gentzel, Alexandra Schambony, and Jürgen Behrens. Pgam5 released from damaged mitochondria induces mitochondrial biogenesis via Wnt signaling. Journal of Cell Biology, 217(4):1383–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/4/1383.

Brown:2018:LEP

[BJL^{+}18] Markus Brown, Louise A. Johnson, Dario A. Leone, Peter Majek, Kari Vaahtomeri, Daniel Senfter, Nora Bukosza, Helga Schachner, Gabriele Asfour, Brigitte Langer, Robert Hauschild, Katja Parapatics, Young-Kwon Hong, Keiryn L. Bennett, Renate Kain, Michael Detmar, Michael Sixt, David G. Jackson, and Dontscho Kerjaschki. Lymphatic exosomes promote dendritic cell migration along guidance cues. Journal of Cell Biology, 217(6):2205–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/6/2205.

Bridges:2016:MSP

[BJO^{+}16] Andrew A. Bridges, Maximilian S. Jentzsch, Patrick W. Oakes, Patricia Occhipinti, and Amy S. Gladfelter. Micron-scale plasma membrane curvature is recognized by the septin cytoskeleton. Journal of Cell Biology, 213(1):23–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/213/1/23.

Brown:2019:DFS

[BK19] Aidan I. Brown and Elena F. Koslover. Drive, filter, and stick: a protein sorting conspiracy in photoreceptors. Journal of Cell Biology, 218(11):3533–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/11/3533.

Bianchini:2015:RCM

[BKG^{+}15] Julie M. Bianchini, Khameeka N. Kitt, Martijn Gloerich, Sabine Pokutta, William I. Weis, and W. James Nelson. Reevaluating αE-catenin monomer and homodimer functions by characterizing E-cadherin/αE-catenin chimeras. Journal of Cell Biology, 210(7):1065–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/7/1065.

Bamidele:2015:IPC

[BKH^{+}15] Adebowale O. Bamidele, Kimberly N. Kremer, Petra Hirsova, Ian C. Clift, Gregory J. Gores, Daniel D. Billadeau, and Karen E. Hedin. IQGAP1 promotes CXCR4 chemokine receptor function and trafficking via EEA-1^{+} endosomes. Journal of Cell Biology, 210(2):257–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/2/257.

Baumgart:2019:STS

[BKR^{+}19] Johannes Baumgart, Marcel Kirchner, Stefanie Redemann, Alec Bond, Jeffrey Woodruff, Jean-Marc Verbavatz, Frank Jülicher, Thomas Müller-Reichert, Anthony A. Hyman, and Jan Brugués. Soluble tubulin is significantly enriched at mitotic centrosomes. Journal of Cell Biology, 218(12):3977–3985, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
218/12/3977/132532/Soluble-tubulin-is-significantly-
enriched-at.

Bao:2015:MDA

[BLG^{+}15] Yi Bao, Carola Ledderose, Amelie F. Graf, Bianca Brix, Theresa Birsak, Albert Lee, Jingping Zhang, and Wolfgang G. Junger. mTOR and differential activation of mitochondria orchestrate neutrophil chemotaxis. Journal of Cell Biology, 210(7):1153–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/7/1153.

Badrinarayanan:2015:RPR

[BLL15] Anjana Badrinarayanan, Tung B. K. Le, and Michael T. Laub. Rapid pairing and resegregation of distant homologous loci enables double-strand break repair in bacteria. Journal of Cell Biology, 210(3):385–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/3/385.

Baeyens:2016:DFS

[BLO^{+}16] Nicolas Baeyens, Bruno Larrivée, Roxana Ola, Brielle Hayward-Piatkowskyi, Alexandre Dubrac, Billy Huang, Tyler D. Ross, Brian G. Coon, Elizabeth Min, Maya Tsarfati, Haibin Tong, Anne Eichmann, and Martin A. Schwartz. Defective fluid shear stress mechanotransduction mediates hereditary hemorrhagic telangiectasia. Journal of Cell Biology, 214(7):807–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/214/7/
807.

Blow:2019:DOL

[Blo19] J. Julian Blow. Defects in the origin licensing checkpoint stresses cells exiting G0. Journal of Cell Biology, 218(7):2080– ??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/7/2080.

Buono:2017:EMV

[BLPV^{+}17] Rafael Andrade Buono, André Leier, Julio Paez-Valencia, Janice Pennington, Kaija Goodman, Nathan Miller, Paul Ahlquist, Tatiana T. Marquez-Lago, and Marisa S. Otegui. ESCRTmediated vesicle concatenation in plant endosomes. Journal of Cell Biology, 216(7):2167–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/7/2167.

Blum:2015:MSL

[Blu15a] Karen Blum. Marcos Sotomayor: Listening in on the cadherin family’s secrets. Journal of Cell Biology, 211(5):938–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/5/
938.

Blum:2015:RPI

[Blu15b] Karen Blum. Rosa Puertollano: The importance of recycling cellular trash. Journal of Cell Biology, 211(6):1100–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/6/
1100.

Briz:2015:NFS

[BLZ^{+}15] Victor Briz, Yan Liu, Guoqi Zhu, Xiaoning Bi, and Michel Baudry. A novel form of synaptic plasticity in field CA3 of hippocampus requires GPER1 activation and BDNF release. Journal of Cell Biology, 210(7):1225–??, September 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/210/7/1225.

Bessonnard:2015:PRP

[BMC15] Sylvain Bessonnard, Daniel Mesnard, and Daniel B. Constam. PC7 and the related proteases Furin and Pace4 regulate Ecadherin function during blastocyst formation. Journal of Cell Biology, 210(7):1185–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/7/1185.

Bertiaux:2018:BIT

[BMF^{+}18] Eloïse Bertiaux, Adeline Mallet, Cécile Fort, Thierry Blisnick, Serge Bonnefoy, Jamin Jung, Moara Lemos, Sergio Marco, Sue Vaughan, Sylvain Trépout, Jean-Yves Tinevez, and Philippe Bastin. Bidirectional intraflagellar transport is restricted to two sets of microtubule doublets in the trypanosome flagellum. Journal of Cell Biology, 217(12):4284–??, December 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/12/4284.

Bae:2019:DBM

[BMM^{+}19] Donghwi Bae, Kristin A. Moore, Jessica M. Mella, Samantha Y. Hayashi, and Julie Hollien. Degradation of Blos1 mRNA by IRE1 repositions lysosomes and protects cells from stress. Journal of Cell Biology, 218(4):1118–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/4/1118.

Bending:2018:TAT

[BMP^{+}18] David Bending, Paz Prieto Martín, Alina Paduraru, Catherine Ducker, Erik Marzaganov, Marie Laviron, Satsuki Kitano, Hitoshi Miyachi, Tessa Crompton, and Masahiro Ono. A timer for analyzing temporally dynamic changes in transcription during differentiation in vivo. Journal of Cell Biology, 217(8):2931–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/8/2931.

Botti:2017:CDS

[BMS^{+}17] Valentina Botti, François McNicoll, Michaela C. Steiner, Florian M. Richter, Anfisa Solovyeva, Marius Wegener, Oliver D. Schwich, Ina Poser, Kathi Zarnack, Ilka Wittig, Karla M. Neugebauer, and Michaela Müller-McNicoll. Cellular differentiation state modulates the mRNA export activity of SR proteins. Journal of Cell Biology, 216(7):1993–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/7/1993.

Biedka:2018:HRR

[BMW^{+}18] Stephanie Biedka, Jelena Micic, Daniel Wilson, Hailey Brown, Luke Diorio-Toth, and John L. Woolford. Hierarchical recruitment of ribosomal proteins and assembly factors remodels nucleolar pre-60S ribosomes. Journal of Cell Biology, 217(7):2503– ??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/7/2503.

Brachet:2015:LTC

[BNB^{+}15] Anna Brachet, Stephanie Norwood, Jos F. Brouwers, Ernest Palomer, J. Bernd Helms, Carlos G. Dotti, and José A. Esteban. LTP-triggered cholesterol redistribution activates Cdc42 and drives AMPA receptor synaptic delivery. Journal of Cell Biology, 208(6):791–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/6/791.

Bohr:2015:SAC

[BNKB15] Tisha Bohr, Christian R. Nelson, Erin Klee, and Needhi Bhalla. Spindle assembly checkpoint proteins regulate and monitor meiotic synapsis in C. elegans. Journal of Cell Biology, 211(2): 233–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/2/233.

Belvindrah:2017:MTT

[BNS^{+}17] Richard Belvindrah, Kathiresan Natarajan, Preety Shabajee, Elodie Bruel-Jungerman, Jennifer Bernard, Marie Goutierre, Imane Moutkine, Xavier H. Jaglin, Mythili Savariradjane, Theano Irinopoulou, Jean-Christophe Poncer, Carsten Janke, and Fiona Francis. Mutation of the α-tubulin Tuba1a leads to straighter microtubules and perturbs neuronal migration. Journal of Cell Biology, 216(8):2443–??, August 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/8/2443.

Bobola:2017:DBT

[Bob17] Nicoletta Bobola. From DNA binding to transcriptional activation: Is the TALE complete? Journal of Cell Biology, 216(9): 2603–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/9/2603.

Bohnert:2018:WFS

[Boh18] Maria Bohnert. Wrapping up the fats — a structure of the lipid droplet biogenesis protein seipin. Journal of Cell Biology, 217(12):4053–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/12/4053.

Bugaj:2017:ICP

[BOL17] Lukasz J. Bugaj, Geoff P. O’Donoghue, and Wendell A. Lim. Interrogating cellular perception and decision making with optogenetic tools. Journal of Cell Biology, 216(1):25–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/1/
25.

Benoit:2019:CMR

[BP19a] Béatrice Benoit and Christian Poüs. Correction: Microtubule reorientation in the blue spotlight: Cutting and CLASPing at dynamic hot spots. Journal of Cell Biology, 218(2): 723–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/2/723. See [BP19b].

Benoit:2019:MRB

[BP19b] Béatrice Benoit and Christian Poüs. Microtubule reorientation in the blue spotlight: Cutting and CLASPing at dynamic hot spots. Journal of Cell Biology, 218(1):8–??, January 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/1/8. See cor-

rection [BP19a].

Bonello:2019:SMS

[BP19c] Teresa T. Bonello and Mark Peifer. Scribble: a master scaffold in polarity, adhesion, synaptogenesis, and proliferation. Journal of Cell Biology, 218(3):742–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/3/742.

Bassler:2015:NAF

[BPH^{+}15] Jochen Baßler, Helge Paternoga, Iris Holdermann, Matthias Thoms, Sander Granneman, Clara Barrio-Garcia, Afua Nyarko, Woonghee Lee, Gunter Stier, Sarah A. Clark, Daniel Schraivogel, Martina Kallas, Roland Beckmann, David Tollervey, Elisar Barbar, Irmi Sinning, and Ed Hurt. A network of assembly factors is involved in remodeling rRNA elements during preribosome maturation. Journal of Cell Biology, 210(1):169–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
210/1/169.

Beati:2018:AJA

[BPH^{+}18] Hamze Beati, Irina Peek, Paulina Hordowska, Mona Honemann-Capito, Jade Glashauser, Fabian A. Renschler, Parisa Kakanj, Andreas Ramrath, Maria Leptin, Stefan Luschnig, Silke Wiesner, and Andreas Wodarz. The adherens junction–associated LIM domain protein Smallish regulates epithelial morphogenesis. Journal of Cell Biology, 217(3):1079–??, March 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/3/1079.

Bassaganyas:2019:NFP

[BPH^{+}19] Laia Bassaganyas, Stephanie J. Popa, Max Horlbeck, Claudia Puri, Sarah E. Stewart, Felix Campelo, Anupama Ashok, Cristian M. Butnaru, Nathalie Brouwers, Kartoosh Heydari, Jean Ripoche, Jonathan Weissman, David C. Rubinsztein, Randy Schekman, Vivek Malhotra, Kevin Moreau, and Julien Villeneuve. New factors for protein transport identified by a genome-wide CRISPRi screen in mammalian cells. Journal of Cell Biology, 218(11):3861–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/11/3861.

Bas:2018:RRY

[BPL^{+}18] Levent Bas, Daniel Papinski, Mariya Licheva, Raffaela Torggler, Sabrina Rohringer, Martina Schuschnig, and Claudine Kraft. Reconstitution reveals Ykt6 as the autophagosomal SNARE in autophagosome–vacuole fusion. Journal of Cell Biology, 217(10):3656–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/10/3656.

Boni:2015:LIM

[BPS^{+}15] Andrea Boni, Antonio Z. Politi, Petr Strnad, Wanqing Xiang, M. Julius Hossain, and Jan Ellenberg. Live imaging and modeling of inner nuclear membrane targeting reveals its molecular requirements in mammalian cells. Journal of Cell Biology, 209 (5):705–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/5/705.

Borrego-Pinto:2016:DME

[BPSK^{+}16] Joana Borrego-Pinto, Kálmán Somogyi, Matthia A. Karreman, Julia König, Thomas Müller-Reichert, Mónica Bettencourt-Dias, Pierre Gönczy, Yannick Schwab, and Péter Lénárt. Distinct mechanisms eliminate mother and daughter centrioles in meiosis of starfish oocytes. Journal of Cell Biology, 212 (7):815–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/7/815.

Bjork:2015:IBS

[BPW15] Petra Björk, Jan-Olov Persson, and Lars Wieslander. Intranuclear binding in space and time of exon junction complex and NXF1 to premRNPs/mRNPs in vivo. Journal of Cell Biology, 211(1):63–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/1/63.

Burman:2017:MFF

[BPW^{+}17] Jonathon L. Burman, Sarah Pickles, Chunxin Wang, Shiori Sekine, Jose Norberto S. Vargas, Zhe Zhang, Alice M. Youle, Catherine L. Nezich, Xufeng Wu, John A. Hammer, and Richard J. Youle. Mitochondrial fission facilitates the selective mitophagy of protein aggregates. Journal of Cell Biology, 216(10):3231–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/10/3231.

Brandeis:2016:SSM

[Bra16] Michael Brandeis. Slip slidin’ away of mitosis with CRL2^{Zyg11}. Journal of Cell Biology, 215(2):143–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/215/2/143.

Bernabe-Rubio:2016:NRM

[BRACA^{+}16] Miguel Bernabé-Rubio, Germán Andrés, Javier Casares-Arias, Jaime Fernández-Barrera, Laura Rangel, Natalia Reglero-Real, David C. Gershlick, José J. Fernández, Jaime Millán, Isabel Correas, David G. Miguez, and Miguel A. Alonso. Novel role for the midbody in primary ciliogenesis by polarized epithelial cells. Journal of Cell Biology, 214(3):259–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/3/259.

Bendre:2016:GTM

[BRH^{+}16] Shweta Bendre, Arnaud Rondelet, Conrad Hall, Nadine Schmidt, Yu-Chih Lin, Gary J. Brouhard, and Alexander W. Bird. GTSE1 tunes microtubule stability for chromosome alignment and segregation by inhibiting the microtubule depolymerase MCAK. Journal of Cell Biology, 215(5):631–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/215/5/
631.

Bronner:2016:HIC

[Bro16] Marianne E. Bronner. How inhibitory cues can both constrain and promote cell migration. Journal of Cell Biology, 213(5):505–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/5/505.

Broihier:2019:EPS

[Bro19] Heather T. Broihier. Expecto Patronin for slow and persistent minus end microtubule growth in dendrites. Journal of Cell Biology, 218(7):2084–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/7/2084.

Bowes:2019:CDR

[BRY^{+}19] Charnese Bowes, Michael Redd, Malika Yousfi, Muriel Tauzin, Emi Murayama, and Philippe Herbomel. Coronin 1A depletion restores the nuclear stability and viability of Aip1/Wdr1deficient neutrophils. Journal of Cell Biology, 218(10):3258–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/10/3258.

Bezbradica:2017:ZES

[BS17a] Jelena S. Bezbradica and Kate Schroder. Zebrafish earns its stripes for in vivo ASC speck dynamics. Journal of Cell Biology, 216(9):2615–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/9/2615.

Byrnes:2017:TTB

[BS17b] Amy E. Byrnes and Kevin C. Slep. TOG-tubulin binding specificity promotes microtubule dynamics and mitotic spindle formation. Journal of Cell Biology, 216(6):1641–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/6/
1641.

Bansal:2018:GMC

[BS18] Ankita Bansal and M. Celeste Simon. Glutathione metabolism in cancer progression and treatment resistance. Journal of Cell Biology, 217(7):2291–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/7/2291.

Belair:2019:REN

[BSK^{+}19] Cedric Belair, Soyeong Sim, Kun-Yong Kim, Yoshiaki Tanaka, In-Hyun Park, and Sandra L. Wolin. The RNA exosome nuclease complex regulates human embryonic stem cell differentiation. Journal of Cell Biology, 218(8):2564–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/8/
2564.

Bencsik:2015:PKD

[BSL^{+}15] Norbert Bencsik, Zsófia Szíber, Hanna Liliom, Krisztián Tárnok, Sándor Borbély, Márton Gulyás, Anikó Rátkai, Attila Szűcs, Diána Hazai-Novák, Kornelia Ellwanger, Bence Rácz, Klaus Pfizenmaier, Angelika Hausser, and Katalin Schlett. Protein kinase D promotes plasticity-induced F-actin stabilization in dendritic spines and regulates memory formation. Journal of Cell Biology, 210(5):771–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/5/771.

Bowman:2016:DGP

[BSP16] Shanna Lynn Bowman, Daniel John Shiwarski, and Manojkumar A. Puthenveedu. Distinct G protein–coupled receptor recycling pathways allow spatial control of downstream G protein signaling. Journal of Cell Biology, 214(7):797–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/214/7/

Burger:2017:NPD

[BSP^{+}17] Kaspar Burger, Margarita Schlackow, Martin Potts, Svenja Hester, Shabaz Mohammed, and Monika Gullerova. Nuclear phosphorylated Dicer processes double-stranded RNA in response to DNA damage. Journal of Cell Biology, 216(8): 2373–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/8/2373.

Bennabi:2016:MSA

[BTV16] Isma Bennabi, Marie-Emilie Terret, and Marie-Hélène Verlhac. Meiotic spindle assembly and chromosome segregation in oocytes. Journal of Cell Biology, 215(5):611–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/215/5/
611.

Brukman:2019:HCF

[BUPC19] Nicolas G. Brukman, Berna Uygur, Benjamin Podbilewicz, and Leonid V. Chernomordik. How cells fuse. Journal of Cell Biology, 218(5):1436–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/5/1436.

Bottcher:2017:KRP

[BVR^{+}17] Ralph T. Böttcher, Maik Veelders, Pascaline Rombaut, Jan Faix, Marina Theodosiou, Theresa E. Stradal, Klemens Rottner, Roy Zent, Franz Herzog, and Reinhard Fässler. Kindlin-2 recruits paxillin and Arp2/3 to promote membrane protrusions during initial cell spreading. Journal of Cell Biology, 216(11): 3785–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/11/3785.

Ben-Yishay:2019:IWS

[BYMS^{+}19] Rakefet Ben-Yishay, Amir Mor, Amit Shraga, Asaf Ashkenazy-Titelman, Noa Kinor, Avital Schwed-Gross, Avi Jacob, Noga Kozer, Pramod Kumar, Yuval Garini, and Yaron Shav-Tal. Imaging within single NPCs reveals NXF1’s role in mRNA export on the cytoplasmic side of the pore. Journal of Cell Biology, 218(9):2962–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/9/2962.

Bestul:2017:MMF

[BYUJ17] Andrew J. Bestul, Zulin Yu, Jay R. Unruh, and Sue L. Jaspersen. Molecular model of fission yeast centrosome assembly determined by superresolution imaging. Journal of Cell Biology, 216(8):2409–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/8/2409.

Bruckner:2017:FOS

[BZG^{+}17] Joseph J. Bruckner, Hong Zhan, Scott J. Gratz, Monica Rao, Fiona Ukken, Gregory Zilberg, and Kate M. O’Connor-Giles. Fife organizes synaptic vesicles and calcium channels for highprobability neurotransmitter release. Journal of Cell Biology, 216(1):231–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/1/231.

Court:2017:RNS

[CAA^{+}17] Helen Court, Ian M. Ahearn, Marc Amoyel, Erika A. Bach, and Mark R. Philips. Regulation of NOTCH signaling by RAB7 and RAB8 requires carboxyl methylation by ICMT. Journal of Cell Biology, 216(12):4165–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/12/4165.

Chung:2015:RCX

[CAI^{+}15] Byung Min Chung, Artem Arutyunov, Erika Ilagan, Nu Yao, Marsha Wills-Karp, and Pierre A. Coulombe. Regulation of C–X–C chemokine gene expression by keratin 17 and hnRNP K in skin tumor keratinocytes. Journal of Cell Biology, 208 (5):613–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/208/5/613.

Cantu:2016:DSN

[CAKL16] Andrea V. Cantú, Svetlana Altshuler-Keylin, and Diana J. Laird. Discrete somatic niches coordinate proliferation and migration of primordial germ cells via Wnt signaling. Journal of Cell Biology, 214(2):215–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/214/2/215.

Canto:2017:HSF

[Can17] Carles Cantó. The heat shock factor HSF1 juggles protein quality control and metabolic regulation. Journal of Cell Biology, 216(3):551–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/3/551.

Cantor:2019:TJM

[Can19] Sharon Cantor. TPX2 joins 53BP1 to maintain DNA repair and fork stability. Journal of Cell Biology, 218(2):383–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/2/
383.

Chan:2017:DBC

[CANG^{+}17] Kuan Yoow Chan, Marisa Alonso-Nuñez, Agnes Grallert, Kayoko Tanaka, Yvonne Connolly, Duncan L. Smith, and Iain M. Hagan. Dialogue between centrosomal entrance and exit scaffold pathways regulates mitotic commitment. Journal of Cell Biology, 216(9):2795–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/9/2795.

Choi:2016:RZA

[CAP^{+}16] Wangsun Choi, Bipul R. Acharya, Grégoire Peyret, Marc-Antoine Fardin, René-Marc Mège, Benoit Ladoux, Alpha S. Yap, Alan S. Fanning, and Mark Peifer. Remodeling the zonula adherens in response to tension and the role of afadin in this response. Journal of Cell Biology, 213(2):243–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/213/2/243.

Casadio:2016:FGE

[Cas16a] Melina Casadio. Fanni Gergely: Exploring centrosome biology. Journal of Cell Biology, 215(3):294–??, November 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/215/3/294.

Casadio:2016:JBD

[Cas16b] Melina Casadio. Jim Bear: Delineating the mechanics of cell migration. Journal of Cell Biology, 215(5):594–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/215/5/
594.

Casadio:2017:AKC

[Cas17a] Melina Casadio. Alberto Kornblihtt: Coupling alternative splicing with transcription. Journal of Cell Biology, 216(2): 284–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/2/284.

Casadio:2017:AVD

[Cas17b] Melina Casadio. Andrea Ventura: Decrypting noncoding RNAs. Journal of Cell Biology, 216(7):1866–??, July 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/7/1866.

Chitramuthu:2016:PRT

[CB16] Babykumari Chitramuthu and Andrew Bateman. Progranulin and the receptor tyrosine kinase EphA2, partners in crime? Journal of Cell Biology, 215(5):603–??, December 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/215/5/603.

Cantaut-Belarif:2017:MCG

[CBAP^{+}17] Yasmine Cantaut-Belarif, Myriam Antri, Rocco Pizzarelli, Sabrina Colasse, Ilaria Vaccari, Sylvia Soares, Marianne Renner, Radhouane Dallel, Antoine Triller, and Alain Bessis. Microglia control the glycinergic but not the GABAergic synapses via prostaglandin E2 in the spinal cord. Journal of Cell Biology, 216(9):2979–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/9/2979.

Comrie:2015:FAF

[CBB15] William A. Comrie, Alexander Babich, and Janis K. Burkhardt. F-actin flow drives affinity maturation and spatial organization of LFA-1 at the immunological synapse. Journal of Cell Biology, 208(4):475–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/4/475.

Colaluca:2018:NMF

[CBF^{+}18] Ivan Nicola Colaluca, Andrea Basile, Lee Freiburger, Veronica D’Uva, Davide Disalvatore, Manuela Vecchi, Stefano Confalonieri, Daniela Tosoni, Valentina Cecatiello, Maria Grazia Malabarba, Chun-Jiun Yang, Masatsune Kainosho, Michael Sattler, Marina Mapelli, Salvatore Pece, and Pier Paolo Di Fiore. A Numb–Mdm2 fuzzy complex reveals an isoformspecific involvement of Numb in breast cancer. Journal of Cell Biology, 217(2):745–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/2/745.

Coon:2015:IBV

[CBH^{+}15] Brian G. Coon, Nicolas Baeyens, Jinah Han, Madhusudhan Budatha, Tyler D. Ross, Jennifer S. Fang, Sanguk Yun, Jeon-Leon Thomas, and Martin A. Schwartz. Intramembrane binding of VE-cadherin to VEGFR2 and VEGFR3 assembles the endothelial mechanosensory complex. Journal of Cell Biology, 208 (7):975–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/208/7/975.

Crevenna:2016:SCS

[CBM^{+}16] Alvaro H. Crevenna, Birgit Blank, Andreas Maiser, Derya Emin, Jens Prescher, Gisela Beck, Christine Kienzle, Kira Bartnik, Bianca Habermann, Mehrshad Pakdel, Heinrich Leonhardt, Don C. Lamb, and Julia von Blume. Secretory cargo sorting by Ca^{2+}-dependent Cab45 oligomerization at the trans-Golgi network. Journal of Cell Biology, 213(3):305–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/213/3/
305.

Chang:2019:BPB

[CC19] Julie Chang and Ovijit Chaudhuri. Beyond proteases: Basement membrane mechanics and cancer invasion. Journal of Cell Biology, 218(8):2456–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/8/2456.

Curnock:2019:TCR

[CCBC19] Rachel Curnock, Alessia Calcagni, Andrea Ballabio, and Peter J. Cullen. TFEB controls retromer expression in response to nutrient availability. Journal of Cell Biology, 218(12):3954– 3966, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/218/12/3954/132541/TFEB-controls-retromer-
expression-in-response-to.

Costello:2017:AVM

[CCH^{+}17] Joseph L. Costello, Inês G. Castro, Christian Hacker, Tina A. Schrader, Jeremy Metz, Dagmar Zeuschner, Afsoon S. Azadi, Luis F. Godinho, Victor Costina, Peter Findeisen, Andreas Manner, Markus Islinger, and Michael Schrader. ACBD5 and VAPB mediate membrane associations between peroxisomes and the ER. Journal of Cell Biology, 216(2):331–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/2/
331.

Chen:2017:RCS

[CCLL17] Yu-Ju Chen, Chi-Lun Chang, Wan-Ru Lee, and Jen Liou. RASSF4 controls SOCE and ER–PM junctions through regulation of PI(4,5)P_{2}. Journal of Cell Biology, 216(7):2011–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/7/2011.

Chang:2018:EBR

[CCQ^{+}18] Chi-Lun Chang, Yu-Ju Chen, Carlo Giovanni Quintanilla, Ting-Sung Hsieh, and Jen Liou. EB1 binding restricts STIM1 translocation to ER–PM junctions and regulates store-operated Ca^{2+} entry. Journal of Cell Biology, 217(6):2047–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/6/
2047.

Chow:2019:AAA

[CCS^{+}19] Hei-Man Chow, Aifang Cheng, Xuan Song, Mavis R. Swerdel, Ronald P. Hart, and Karl Herrup. ATM is activated by ATP depletion and modulates mitochondrial function through NRF1. Journal of Cell Biology, 218(3):909–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/3/909.

Cui:2019:RSF

[CCY^{+}19] Yi Cui, Julian M. Carosi, Zhe Yang, Nicholas Ariotti, Markus C. Kerr, Robert G. Parton, Timothy J. Sargeant, and Rohan D. Teasdale. Retromer has a selective function in cargo sorting via endosome transport carriers. Journal of Cell Biology, 218(2):615–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/2/615.

Cadwell:2018:BSE

[CD18] Ken Cadwell and Jayanta Debnath. Beyond self-eating: The control of nonautophagic functions and signaling pathways by autophagy-related proteins. Journal of Cell Biology, 217(3): 813–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/3/813.

Conic:2018:INT

[CDF^{+}18] Sascha Conic, Dominique Desplancq, Alexia Ferrand, Veronique Fischer, Vincent Heyer, Bernardo Reina San Martin, Julien Pontabry, Mustapha Oulad-Abdelghani, Kishore Babu N., Graham D. Wright, Nacho Molina, Etienne Weiss, and László Tora. Imaging of native transcription factors and histone phosphorylation at high resolution in live cells. Journal of Cell Biology, 217(4):1537–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/4/1537.

Chakrabarty:2019:PFB

[CDT^{+}19] Nilaj Chakrabarty, Pankaj Dubey, Yong Tang, Archan Ganguly, Kelsey Ladt, Christophe Leterrier, Peter Jung, and Subhojit Roy. Processive flow by biased polymerization mediates the slow axonal transport of actin. Journal of Cell Biology, 218 (1):112–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/1/112.

Cohen:2016:CNG

[CE16] Idan Cohen and Elena Ezhkova. Cbx4: a new guardian of p63’s domain of epidermal control. Journal of Cell Biology, 212 (1):9–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/1/9.

Cooper:2015:PCM

[CED^{+}15] Sharon R. Cooper, Michelle R. Emond, Phan Q. Duy, Brandon G. Liebau, Marc A. Wolman, and James D. Jontes. Protocadherins control the modular assembly of neuronal columns in the zebrafish optic tectum. Journal of Cell Biology, 211(4): 807–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/4/807.

Chetty:2015:SIR

[CEM^{+}15] Sundari Chetty, Elise N. Engquist, Elie Mehanna, Kathy O. Lui, Alexander M. Tsankov, and Douglas A. Melton. A Src inhibitor regulates the cell cycle of human pluripotent stem cells and improves directed differentiation. Journal of Cell Biology, 210(7):1257–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/7/1257.

Cao:2015:TOS

[CF15] Xu Cao and Yanshan Fang. Transducing oxidative stress to death signals in neurons. Journal of Cell Biology, 211(4): 741–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/4/741.

Catlett:2016:DLG

[CG16] Timothy S. Catlett and Timothy M. Gomez. Division of labor in the growth cone by DSCR1. Journal of Cell Biology, 213(4):407–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/4/407.

Churikov:2017:NTA

[CG17] Dmitri Churikov and Vincent Géli. De novo telomere addition at chromosome breaks: Dangerous Liaisons. Journal of Cell Biology, 216(8):2243–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/8/2243.

Cruz-Garcia:2017:DMD

[CGBD^{+}17] David Cruz-Garcia, Nathalie Brouwers, Juan M. Duran, Gabriel Mora, Amy J. Curwin, and Vivek Malhotra. A diacidic motif determines unconventional secretion of wild-type and ALS-linked mutant SOD1. Journal of Cell Biology, 216(9): 2691–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/9/2691.

Collins:2018:RBF

[CGD^{+}18] Jason C. Collins, Homa Ghalei, Joanne R. Doherty, Haina Huang, Rebecca N. Culver, and Katrin Karbstein. Ribosome biogenesis factor Ltv1 chaperones the assembly of the small subunit head. Journal of Cell Biology, 217(12):4141–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/12/4141.

Chen:2017:SGC

[CGPB17] Yu Chen, David C. Gershlick, Sang Yoon Park, and Juan S. Bonifacino. Segregation in the Golgi complex precedes export of endolysosomal proteins in distinct transport carriers. Journal of Cell Biology, 216(12):4141–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/12/4141.

Cabello:2016:CES

[CGT16] Simon Cabello, Yannick Gachet, and Sylvie Tournier. Cutting edge science: Laser surgery illuminates viscoelasticity of merotelic kinetochores. Journal of Cell Biology, 212(7):747–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
212/7/747.

Chen:2019:YCC

[CGY^{+}19] Jingjing Chen, Jennifer M. Gardner, Zulin Yu, Sarah E. Smith, Sean McKinney, Brian D. Slaughter, Jay R. Unruh, and Sue L. Jaspersen. Yeast centrosome components form a noncanonical LINC complex at the nuclear envelope insertion site. Journal of Cell Biology, 218(5):1478–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/5/1478.

Cundell:2016:PBR

[CHB^{+}16] Michael J. Cundell, Lukas H. Hutter, Ricardo Nunes Bastos, Elena Poser, James Holder, Shabaz Mohammed, Bela Novak, and Francis A. Barr. A PP2A–B55 recognition signal controls substrate dephosphorylation kinetics during mitotic exit. Journal of Cell Biology, 214(5):539–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/5/539.

Condon:2018:MFT

[CHC^{+}18] Nicholas D. Condon, John M. Heddleston, Teng-Leong Chew, Lin Luo, Peter S. McPherson, Maria S. Ioannou, Louis Hodgson, Jennifer L. Stow, and Adam A. Wall. Macropinosome formation by tent pole ruffling in macrophages. Journal of Cell Biology, 217(11):3873–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/11/3873.

Craft:2015:TTI

[CHH^{+}15] Julie M. Craft, J. Aaron Harris, Sebastian Hyman, Peter Kner, and Karl F. Lechtreck. Tubulin transport by IFT is upregulated during ciliary growth by a cilium-autonomous mechanism. Journal of Cell Biology, 208(2):223–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/2/223.

Chen:2015:CMC

[CHI^{+}15] Chi-Shuo Chen, Soonjin Hong, Indrajyoti Indra, Alina P. Sergeeva, Regina B. Troyanovsky, Lawrence Shapiro, Barry Honig, and Sergey M. Troyanovsky. α-catenin–mediated cadherin clustering couples cadherin and actin dynamics. Journal of Cell Biology, 210(4):647–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/4/647.

Chen:2019:BUE

[CHL^{+}19] Fei-Yun Chen, Min-Yu Huang, Yu-Min Lin, Chi-Huan Ho, Shu-Yu Lin, Hsin-Yi Chen, Mien-Chie Hung, and Ruey-Hwa Chen. BIK ubiquitination by the E3 ligase Cul5-ASB11 determines cell fate during cellular stress. Journal of Cell Biology, 218(9): 3002–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/9/3002.

Chew:2017:ATM

[CHP^{+}17] Ting Gang Chew, Junqi Huang, Saravanan Palani, Ruth Sommese, Anton Kamnev, Tomoyuki Hatano, Ying Gu, Snezhana Oliferenko, Sivaraj Sivaramakrishnan, and Mohan K. Balasubramanian. Actin turnover maintains actin filament homeostasis during cytokinetic ring contraction. Journal of Cell Biology, 216(9):2657–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/9/2657.

Chang:2017:RMS

[CHS^{+}17] Chih-Chia Chang, Tzu-Lun Huang, Yuta Shimamoto, Su-Yi Tsai, and Kuo-Chiang Hsia. Regulation of mitotic spindle assembly factor NuMA by Importin-β. Journal of Cell Biology, 216(11):3453–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/11/3453.

Cuellar:2017:SRS

[CHZ^{+}17] Trinna L. Cuellar, Anna-Maria Herzner, Xiaotian Zhang, Yogesh Goyal, Colin Watanabe, Brad A. Friedman, Vasantharajan Janakiraman, Steffen Durinck, Jeremy Stinson, David Arnott, Tommy K. Cheung, Subhra Chaudhuri, Zora Modrusan, Jonas Martin Doerr, Marie Classon, and Benjamin Haley. Silencing of retrotransposons by SETDB1 inhibits the interferon response in acute myeloid leukemia. Journal of Cell Biology, 216(11):3535–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/11/3535.

Cho:2017:MNP

[CID17] Sangkyun Cho, Jerome Irianto, and Dennis E. Discher. Mechanosensing by the nucleus: From pathways to scaling relationships. Journal of Cell Biology, 216(2):305–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/2/
305.

Chau:2017:MCC

[CIK^{+}17] Gia Cac Chau, Dong Uk Im, Tong Mook Kang, Jeong Mo Bae, Won Kim, Suhkneung Pyo, Eun-Yi Moon, and Sung Hee Um. mTOR controls ChREBP transcriptional activity and pancreatic β cell survival under diabetic stress. Journal of Cell Biology, 216(7):2091–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/7/2091.

Capurro:2017:GPG

[CIS^{+}17] Mariana Capurro, Tomomi Izumikawa, Philippe Suarez, Wen Shi, Marzena Cydzik, Tomoyuki Kaneiwa, Jean Gariepy, Luisa Bonafe, and Jorge Filmus. Glypican-6 promotes the growth of developing long bones by stimulating Hedgehog signaling. Journal of Cell Biology, 216(9):2911–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/9/2911.

Casadio:2016:JFK

[CJ16] Melina Casadio and Shawn Jordan. Julie Forman–Kay: Dynamic views on protein structure. Journal of Cell Biology, 214(6):638–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/6/638.

Colaco:2017:RMR

[CJ17] Alexandria Colaço and Marja Jäättelä. Ragulator — a multifaceted regulator of lysosomal signaling and trafficking. Journal of Cell Biology, 216(12):3895–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/12/3895.

Chakrabarti:2018:IMA

[CJS^{+}18] Rajarshi Chakrabarti, Wei-Ke Ji, Radu V. Stan, Jaime de Juan Sanz, Timothy A. Ryan, and Henry N. Higgs. INF2mediated actin polymerization at the ER stimulates mitochondrial calcium uptake, inner membrane constriction, and division. Journal of Cell Biology, 217(1):251–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/217/1/251.

Chen:2015:ROC

[CKJ^{+}15] Jieyan V. Chen, Ling-Rong Kao, Swadhin C. Jana, Elena Sivan-Loukianova, Susana Mendonça, Oscar A. Cabrera, Priyanka Singh, Clemens Cabernard, Daniel F. Eberl, Monica Bettencourt-Dias, and Timothy L. Megraw. Rootletin organizes the ciliary rootlet to achieve neuron sensory function in Drosophila. Journal of Cell Biology, 211(2):435–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/2/
435.

Cullati:2017:BSC

[CKKG17] Sierra N. Cullati, Lilian Kabeche, Arminja N. Kettenbach, and Scott A. Gerber. A bifurcated signaling cascade of NIMArelated kinases controls distinct kinesins in anaphase. Journal of Cell Biology, 216(8):2339–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/8/2339.

Capote:2016:OAA

[CKM^{+}16] Joana Capote, Irina Kramerova, Leonel Martinez, Sylvia Vetrone, Elisabeth R. Barton, H. Lee Sweeney, M. Carrie Miceli, and Melissa J. Spencer. Osteopontin ablation ameliorates muscular dystrophy by shifting macrophages to a proregenerative phenotype. Journal of Cell Biology, 213(2):275–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
213/2/275.

Chen:2015:PFI

[CKS^{+}15] Allen C. Chen, Sumin Kim, Nina Shepardson, Sarvagna Patel, Soyon Hong, and Dennis J. Selkoe. Physical and functional interaction between the α- and γ-secretases: a new model of regulated intramembrane proteolysis. Journal of Cell Biology, 211(6):1157–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/6/1157.

Chen:2016:AMD

[CKX^{+}16] Keng Chen, Chwee Tat Koe, Zhanyuan Benny Xing, Xiaolin Tian, Fabrizio Rossi, Cheng Wang, Quan Tang, Wenhui Zong, Wan Jin Hong, Reshma Taneja, Fengwei Yu, Cayetano Gonzalez, Chunlai Wu, Sharyn Endow, and Hongyan Wang. Arl2and Msps-dependent microtubule growth governs asymmetric division. Journal of Cell Biology, 212(6):661–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/212/6/661.

Cunningham:2019:NRS

[CL19] Karen L. Cunningham and J. Troy Littleton. Neurons regulate synaptic strength through homeostatic scaling of active zones. Journal of Cell Biology, 218(5):1434–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/5/1434.

Comrie:2015:DCC

[CLBB15] William A. Comrie, Shuixing Li, Sarah Boyle, and Janis K. Burkhardt. The dendritic cell cytoskeleton promotes T cell adhesion and activation by constraining ICAM-1 mobility. Journal of Cell Biology, 208(4):457–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/4/457.

Csizmadia:2018:MMD

[CLH^{+}18] Tamás Csizmadia, Péter Lőrincz, Krisztina Hegedűs, Szilvia Széplaki, Péter Lőw, and Gábor Juhász. Molecular mechanisms of developmentally programmed crinophagy in Drosophila. Journal of Cell Biology, 217(1):361–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/1/361.

Chiapparo:2016:MCS

[CLL^{+}16] Giuseppe Chiapparo, Xionghui Lin, Fabienne Lescroart, Samira Chabab, Catherine Paulissen, Lorenzo Pitisci, Antoine Bondue, and Cédric Blanpain. Mesp1 controls the speed, polarity, and directionality of cardiovascular progenitor migration. Journal of Cell Biology, 213(4):463–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/213/4/463.

Lin:2016:MTC

[cLNF^{+}16] Tien chen Lin, Annett Neuner, Dirk Flemming, Peng Liu, Takumi Chinen, Ursula Jäkle, Robert Arkowitz, and Elmar Schiebel. MOZART1 and γ-tubulin complex receptors are both required to turn γ-TuSC into an active microtubule nucleation template. Journal of Cell Biology, 215(6):823–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/215/6/

Chuang:2019:TDE

[CLO^{+}19] Mei-Chun Chuang, Shan-Shan Lin, Ryosuke L. Ohniwa, Gang-Hui Lee, You-An Su, Yu-Chen Chang, Ming-Jer Tang, and Ya-Wen Liu. Tks5 and Dynamin-2 enhance actin bundle rigidity in invadosomes to promote myoblast fusion. Journal of Cell Biology, 218(5):1670–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/5/1670.

Chipuk:2017:PSB

[CLV17] Jerry Edward Chipuk and Mark P. A. Luna-Vargas. The peroxisomes strike BAK: Regulation of peroxisome integrity by the Bcl-2 family. Journal of Cell Biology, 216(3):547–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/3/
547.

Chen:2016:CAJ

[CM16] Chin-Chi Chen and Barbara G. Mellone. Chromatin assembly: Journey to the CENter of the chromosome. Journal of Cell Biology, 214(1):13–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/1/13.

Castro:2018:DGR

[CM18] Daniela Gallo Castro and Sophie G. Martin. Differential GAP requirement for Cdc42-GTP polarization during proliferation and sexual reproduction. Journal of Cell Biology, 217(12): 4215–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/12/4215.

Chacko:2019:CTM

[CMA19] Leeba Ann Chacko, Kritika Mehta, and Vaishnavi Ananthanarayanan. Cortical tethering of mitochondria by the anchor protein Mcp5 enables uniparental inheritance. Journal of Cell Biology, 218(11):3560–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/11/3560.

Chang:2018:CGE

[CMB^{+}18] Hsin-Fang Chang, Stefanie Mannebach, Andreas Beck, Keerthana Ravichandran, Elmar Krause, Katja Frohnweiler, Claudia Fecher-Trost, Claudia Schirra, Varsha Pattu, Veit Flockerzi, and Jens Rettig. Cytotoxic granule endocytosis depends on the Flower protein. Journal of Cell Biology, 217(2):667–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/2/
667.

Cesarini:2015:LCS

[CMM^{+}15] Elisa Cesarini, Chiara Mozzetta, Fabrizia Marullo, Francesco Gregoretti, Annagiusi Gargiulo, Marta Columbaro, Alice Cortesi, Laura Antonelli, Simona Di Pelino, Stefano Squarzoni, Daniela Palacios, Alessio Zippo, Beatrice Bodega, Gennaro Oliva, and Chiara Lanzuolo. Lamin A/C sustains PcG protein architecture, maintaining transcriptional repression at target genes. Journal of Cell Biology, 211(3):533–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/3/
533.

Chiasson-MacKenzie:2015:NMM

[CMMB^{+}15] Christine Chiasson-MacKenzie, Zachary S. Morris, Quentin Baca, Brett Morris, Joanna K. Coker, Rossen Mirchev, Anne E. Jensen, Thomas Carey, Shannon L. Stott, David E. Golan, and Andrea I. McClatchey. NF2/ Merlin mediates contactdependent inhibition of EGFR mobility and internalization via cortical actomyosin. Journal of Cell Biology, 211(2):391–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
211/2/391.

Cheng:2015:CPE

[CMTH^{+}15] Jade P. X. Cheng, Carolina Mendoza-Topaz, Gillian Howard, Jessica Chadwick, Elena Shvets, Andrew S. Cowburn, Benjamin J. Dunmore, Alexi Crosby, Nicholas W. Morrell, and Benjamin J. Nichols. Caveolae protect endothelial cells from membrane rupture during increased cardiac output. Journal of Cell Biology, 211(1):53–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/211/1/53.

Courchaine:2015:PPF

[CN15] Edward Courchaine and Karla M. Neugebauer. Paraspeckles: Paragons of functional aggregation. Journal of Cell Biology, 210 (4):527–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/4/527.

Chang:2017:RLH

[CNA^{+}17] Emily Yun-Chia Chang, Carolina A. Novoa, Maria J. Aristizabal, Yan Coulombe, Romulo Segovia, Richa Chaturvedi, Yaoqing Shen, Christelle Keong, Annie S. Tam, Steven J. M. Jones, Jean-Yves Masson, Michael S. Kobor, and Peter C. Stirling. RECQ-like helicases Sgs1 and BLM regulate R-loop– associated genome instability. Journal of Cell Biology, 216(12): 3991–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/12/3991.

Carroll:2018:LIS

[CNC^{+}18] Thomas D. Carroll, Ian P. Newton, Yu Chen, J. Julian Blow, and Inke Näthke. Lgr5^{+} intestinal stem cells reside in an unlicensed G_{1} phase. Journal of Cell Biology, 217(5):1667–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/5/
1667.

Chen:2017:NTR

[CNN^{+}17] Muhan Chen, Dawid G. Nowak, Navneet Narula, Brian Robinson, Kaitlin Watrud, Alexandra Ambrico, Tali M. Herzka, Martha E. Zeeman, Matthias Minderer, Wu Zheng, Saya H. Ebbesen, Kendra S. Plafker, Carlos Stahlhut, Victoria M. Y. Wang, Lorna Wills, Abu Nasar, Mireia Castillo-Martin, Carlos Cordon-Cardo, John E. Wilkinson, Scott Powers, Raffaella Sordella, Nasser K. Altorki, Vivek Mittal, Brendon M. Stiles, Scott M. Plafker, and Lloyd C. Trotman. The nuclear transport receptor Importin-11 is a tumor suppressor that maintains PTEN protein. Journal of Cell Biology, 216(3):641–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/3/
641.

Carroll:2017:PMS

[CNRR^{+}17] Bernadette Carroll, Glyn Nelson, Yoana Rabanal-Ruiz, Olena Kucheryavenko, Natasha A. Dunhill-Turner, Charlotte C. Chesterman, Qabil Zahari, Tong Zhang, Sarah E. Conduit, Christina A. Mitchell, Oliver D. K. Maddocks, Penny Lovat, Thomas von Zglinicki, and Viktor I. Korolchuk. Persistent mTORC1 signaling in cell senescence results from defects in amino acid and growth factor sensing. Journal of Cell Biology, 216(7):1949–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/7/1949.

Costa:2019:MAM

[CO19] Mariana F. A. Costa and Hiroyuki Ohkura. The molecular architecture of the meiotic spindle is remodeled during metaphase arrest in oocytes. Journal of Cell Biology, 218(9): 2854–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/9/2854.

Choudhary:2015:CFP

[COGP15] Vineet Choudhary, Namrata Ojha, Andy Golden, and William A. Prinz. A conserved family of proteins facilitates nascent lipid droplet budding from the ER. Journal of Cell Biology, 211(2):261–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/2/261.

Coller:2018:DRL

[Col18] Hilary A. Coller. DNA replication licensing in stem cells: Gatekeeping the commitment to proliferation. Journal of Cell Biology, 217(5):1563–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/5/1563.

Cologna:2019:CMN

[Col19] Stephanie M. Cologna. A calcium message for Niemann–Pick type C. Journal of Cell Biology, 218(12):3890–3891, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/218/
12/3890/132510/A-calcium-message-for-Niemann-Pick-type-
CA-calcium.

Conduit:2016:MOC

[Con16] Paul T. Conduit. Microtubule organization: a complex solution. Journal of Cell Biology, 213(6):609–??, June 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/213/6/609.

Cummings:2016:ECS

[CPB^{+}16] Christopher F. Cummings, Vadim Pedchenko, Kyle L. Brown, Selene Colon, Mohamed Rafi, Celestial Jones-Paris, Elena Pokydeshava, Min Liu, Jose C. Pastor-Pareja, Cody Stothers, Isi A. Ero-Tolliver, A. Scott McCall, Roberto Vanacore, Gautam Bhave, Samuel Santoro, Timothy S. Blackwell, Roy Zent, Ambra Pozzi, and Billy G. Hudson. Extracellular chloride signals collagen IV network assembly during basement membrane formation. Journal of Cell Biology, 213(4):479–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/213/4/479.

Casler:2019:MDT

[CPBG19] Jason C. Casler, Effrosyni Papanikou, Juan J. Barrero, and Benjamin S. Glick. Maturation-driven transport and AP-1-dependent recycling of a secretory cargo in the Golgi. Journal of Cell Biology, 218(5):1582–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/5/1582.

Carroll-Portillo:2015:MCD

[CPCtR^{+}15] Amanda Carroll-Portillo, Judy L. Cannon, Joost te Riet, Anna Holmes, Yuko Kawakami, Toshiaki Kawakami, Alessandra Cambi, and Diane S. Lidke. Mast cells and dendritic cells form synapses that facilitate antigen transfer for T cell activation. Journal of Cell Biology, 210(5):851–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/210/5/851.

Colombelli:2015:PRD

[CPEE^{+}15] Cristina Colombelli, Marilena Palmisano, Yael Eshed-Eisenbach, Desirée Zambroni, Ernesto Pavoni, Cinzia Ferri, Stefania Saccucci, Sophie Nicole, Raija Soininen, Karen K. McKee, Peter D. Yurchenco, Elior Peles, Lawrence Wrabetz, and M. Laura Feltri. Perlecan is recruited by dystroglycan to nodes of Ranvier and binds the clustering molecule gliomedin. Journal of Cell Biology, 208(3):313–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/3/313.

Carvalho:2018:OJN

[CPP^{+}18] Lara Carvalho, Pedro Patricio, Susana Ponte, Carl-Philipp Heisenberg, Luis Almeida, André S. Nunes, Nuno A. M. Araújo, and Antonio Jacinto. Occluding junctions as novel regulators of tissue mechanics during wound repair. Journal of Cell Biology, 217(12):4267–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/12/4267.

Chen:2019:XHF

[CQB^{+}19] Bo-Ruei Chen, Annabel Quinet, Andrea K. Byrum, Jessica Jackson, Matteo Berti, Saravanabhavan Thangavel, Andrea L. Bredemeyer, Issa Hindi, Nima Mosammaparast, Jessica K. Tyler, Alessandro Vindigni, and Barry P. Sleckman. XLF and H2AX function in series to promote replication fork stability. Journal of Cell Biology, 218(7):2113–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/7/2113.

Chamberland:2017:RRE

[CR17] John P. Chamberland and Brigitte Ritter. Retromer revisited: Evolving roles for retromer in endosomal sorting. Journal of Cell Biology, 216(11):3433–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/11/3433.

Chitale:2018:DMC

[CR18] Shalaka Chitale and Holger Richly. DICER- and MMSETcatalyzed H4K20me2 recruits the nucleotide excision repair factor XPA to DNA damage sites. Journal of Cell Biology, 217 (2):527–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/2/527.

Cohen:2019:WCC

[CRA^{+}19] Jonathan Cohen, Shaul Raviv, Orit Adir, Krishnanand Padmanabhan, Arad Soffer, and Chen Luxenburg. The Wave complex controls epidermal morphogenesis and proliferation by suppressing Wnt–Sox9 signaling. Journal of Cell Biology, 218 (4):1390–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/4/1390.

Chen:2015:TMF

[CRC^{+}15] Hsiuchen Chen, Shuxun Ren, Clary Clish, Mohit Jain, Vamsi Mootha, J. Michael McCaffery, and David C. Chan. Titration of mitochondrial fusion rescues Mff-deficient cardiomyopathy. Journal of Cell Biology, 211(4):795–??, November 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/211/4/795.

Chung:2017:GDF

[CRK^{+}17] Hyo Kyun Chung, Dongryeol Ryu, Koon Soon Kim, Joon Young Chang, Yong Kyung Kim, Hyon-Seung Yi, Seul Gi Kang, Min Jeong Choi, Seong Eun Lee, Saet-Byel Jung, Min Jeong Ryu, Soung Jung Kim, Gi Ryang Kweon, Hail Kim, Jung Hwan Hwang, Chul-Ho Lee, Se-Jin Lee, Christopher E. Wall, Michael Downes, Ronald M. Evans, Johan Auwerx, and Minho Shong. Growth differentiation factor 15 is a myomitokine governing systemic energy homeostasis. Journal of Cell Biology, 216(1): 149–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/1/149.

Centonze:2019:LRR

[CRN^{+}19] Federica G. Centonze, Veronika Reiterer, Karsten Nalbach, Kota Saito, Krzysztof Pawlowski, Christian Behrends, and Hesso Farhan. LTK is an ER-resident receptor tyrosine kinase that regulates secretion. Journal of Cell Biology, 218(8): 2470–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/8/2470.

Carrasco-Rando:2019:SIP

[CRPSC^{+}19] Marta Carrasco-Rando, Silvia Prieto-Sánchez, Joaquim Culi, Antonio S. Tutor, and Mar Ruiz-Gómez. A specific isoform of Pyd/ZO-1 mediates junctional remodeling and formation of slit diaphragms. Journal of Cell Biology, 218(7):2294–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/7/
2294.

Chowdhury:2017:CPM

[CRS^{+}17] Suvagata Roy Chowdhury, Anastasija Reimer, Malvika Sharan, Vera Kozjak-Pavlovic, Ana Eulalio, Bhupesh K. Prusty, Martin Fraunholz, Karthika Karunakaran, and Thomas Rudel. Chlamydia preserves the mitochondrial network necessary for replication via microRNA-dependent inhibition of fission. Journal of Cell Biology, 216(4):1071–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/4/1071.

Cojoc:2016:LMR

[CRZ^{+}16] Gheorghe Cojoc, Emanuele Roscioli, Lijuan Zhang, Alfonso García-Ulloa, Jagesh V. Shah, Michael W. Berns, Nenad Pavin, Daniela Cimini, Iva M. Tolić, and Juraj Gregan. Laser microsurgery reveals conserved viscoelastic behavior of the kinetochore. Journal of Cell Biology, 212(7):767–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/212/7/767.

Caldieri:2016:SRC

[CS16a] Giusi Caldieri and Sara Sigismund. Spatial resolution of cAMP signaling by soluble adenylyl cyclase. Journal of Cell Biology, 214(2):125–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/214/2/125.

Chabot:2016:DCP

[CS16b] Benoit Chabot and Lulzim Shkreta. Defective control of pre– messenger RNA splicing in human disease. Journal of Cell Biology, 212(1):13–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/1/13.

Chikina:2019:TRU

[CSA19] Aleksandra S. Chikina, Tatyana M. Svitkina, and Antonina Y. Alexandrova. Time-resolved ultrastructure of the cortical actin cytoskeleton in dynamic membrane blebs. Journal of Cell Biology, 218(2):445–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/2/445.

Connolly:2015:KAT

[CSC^{+}15] Amy A. Connolly, Kenji Sugioka, Chien-Hui Chuang, Joshua B. Lowry, and Bruce Bowerman. KLP-7 acts through the Ndc80 complex to limit pole number in C. elegans oocyte meiotic spindle assembly. Journal of Cell Biology, 210(6):917–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/210/6/
917.

Chiu:2017:KSA

[CSF^{+}17] Ya-Fang Chiu, Arthur U. Sugden, Kathryn Fox, Mitchell Hayes, and Bill Sugden. Kaposi’s sarcoma-associated herpesvirus stably clusters its genomes across generations to maintain itself extrachromosomally. Journal of Cell Biology, 216(9):2745– ??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress. org/content/216/9/2745. See correction [CSF^{+}18].

Chiu:2018:CKS

[CSF^{+}18] Ya-Fang Chiu, Arthur U. Sugden, Kathryn Fox, Mitchell Hayes, and Bill Sugden. Correction: Kaposi’s sarcoma-associated herpesvirus stably clusters its genomes across generations to maintain itself extrachromosomally. Journal of Cell Biology, 217(10):3766–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/10/3766. See [CSF^{+}17].

Chatterjee:2015:ICT

[CSG^{+}15] Sujash S. Chatterjee, Abil Saj, Tenzin Gocha, Matthew Murphy, Foster C. Gonsalves, Xiaoqian Zhang, Penelope Hayward, Betül Akgöl Oksuz, Steven S. Shen, Aviv Madar, Alfonso Martinez Arias, and Ramanuj DasGupta. Inhibition of β–catenin– TCF1 interaction delays differentiation of mouse embryonic stem cells. Journal of Cell Biology, 211(1):39–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/211/1/39.

Cornejo:2017:HRH

[CSM17] Elias Cornejo, Philipp Schlaermann, and Shaeri Mukherjee. How to rewire the host cell: a home improvement guide for intracellular bacteria. Journal of Cell Biology, 216(12):3931– ??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/12/3931.

Campbell:2019:PLC

[CSO^{+}19] Hannah K. Campbell, Alicia M. Salvi, Timothy O’Brien, Richard Superfine, and Kris A. DeMali. PAK2 links cell survival to mechanotransduction and metabolism. Journal of Cell Biology, 218(6):1958–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/6/1958.

Casey:2018:MMP

[CSS^{+}18] Alison E. Casey, Ankit Sinha, Rajat Singhania, Julie Livingstone, Paul Waterhouse, Pirashaanthy Tharmapalan, Jennifer Cruickshank, Mona Shehata, Erik Drysdale, Hui Fang, Hyeyeon Kim, Ruth Isserlin, Swneke Bailey, Tiago Medina, Genevieve Deblois, Yu-Jia Shiah, Dalia Barsyte-Lovejoy, Stefan Hofer, Gary Bader, Mathieu Lupien, Cheryl Arrowsmith, Stefan Knapp, Daniel De Carvalho, Hal Berman, Paul C. Boutros, Thomas Kislinger, and Rama Khokha. Mammary molecular portraits reveal lineage-specific features and progenitor cell vulnerabilities. Journal of Cell Biology, 217(8):2951–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/8/
2951.

Chai:2016:MMC

[CST^{+}16] Ye Jin Chai, Emma Sierecki, Vanesa M. Tomatis, Rachel S. Gormal, Nichole Giles, Isabel C. Morrow, Di Xia, Jürgen Götz, Robert G. Parton, Brett M. Collins, Yann Gambin, and Frédéric A. Meunier. Munc18-1 is a molecular chaperone for α-synuclein, controlling its self-replicating aggregation. Journal of Cell Biology, 214(6):705–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/6/705.

Chatzi:2017:PMD

[CST^{+}17] Katerina E. Chatzi, Marios Frantzeskos Sardis, Alexandra Tsirigotaki, Marina Koukaki, Nikolina Šoštarić, Albert Konijnenberg, Frank Sobott, Charalampos G. Kalodimos, Spyridoula Karamanou, and Anastassios Economou. Preprotein mature domains contain translocase targeting signals that are essential for secretion. Journal of Cell Biology, 216(5):1357–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/5/
1357.

Cirera-Salinas:2017:NFD

[CSYB^{+}17] Daniel Cirera-Salinas, Jian Yu, Maxime Bodak, Richard P. Ngondo, Kristina M. Herbert, and Constance Ciaudo. Noncanonical function of DGCR8 controls mESC exit from pluripotency. Journal of Cell Biology, 216(2):355–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/216/2/355.

Choi:2019:SEL

[CTI^{+}19] Jongho Choi, Regina B. Troyanovsky, Indrajyoti Indra, Brian J. Mitchell, and Sergey M. Troyanovsky. Scribble, Erbin, and Lano redundantly regulate epithelial polarity and apical adhesion complex. Journal of Cell Biology, 218(7):2277–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/7/
2277.

Chen:2019:NLC

[CTN^{+}19] Pan Chen, Miroslav Tomschik, Katherine M. Nelson, John Oakey, Jesse C. Gatlin, and Daniel L. Levy. Nucleoplasmin is a limiting component in the scaling of nuclear size with cytoplasmic volume. Journal of Cell Biology, 218(12): 4063–4078, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/218/12/4063/132540/Nucleoplasmin-is-
a-limiting-component-in-the.

Carvalhal:2018:QAC

[CTS^{+}18] Sara Carvalhal, Alexandra Tavares, Mariana B. Santos, Mihailo Mirkovic, and Raquel A. Oliveira. A quantitative analysis of cohesin decay in mitotic fidelity. Journal of Cell Biology, 217(10):3343–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/10/3343.

Chaturvedi:2019:SMC

[CV19] Dhananjay Chaturvedi and K. VijayRaghavan. From a syncytium to mononucleate cells and back: Yki and JNK in symphony. Journal of Cell Biology, 218(11):3531–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/
11/3531.

Cherepanova:2019:QGR

[CVL^{+}19] Natalia A. Cherepanova, Sergey V. Venev, John D. Leszyk, Scott A. Shaffer, and Reid Gilmore. Quantitative glycoproteomics reveals new classes of STT3A- and STT3B-dependent N-glycosylation sites. Journal of Cell Biology, 218(8):2782–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/8/2782.

Casanova:2017:NRE

[CW17] James E. Casanova and Bettina Winckler. A new Rab7 effector controls phosphoinositide conversion in endosome maturation. Journal of Cell Biology, 216(10):2995–??, October 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/10/2995.

Cannon:2019:AHE

[CWCG19] Kevin S. Cannon, Benjamin L. Woods, John M. Crutchley, and Amy S. Gladfelter. An amphipathic helix enables septins to sense micrometer-scale membrane curvature. Journal of Cell Biology, 218(4):1128–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/4/1128.

Chang:2015:AAL

[CWG15] Wakam Chang, Howard J. Worman, and Gregg G. Gundersen. Accessorizing and anchoring the LINC complex for multifunctionality. Journal of Cell Biology, 208(1):11–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/208/1/11.

Chang:2019:STL

[CWI^{+}19] Chi-Lun Chang, Aubrey V. Weigel, Maria S. Ioannou, H. Amalia Pasolli, C. Shan Xu, David R. Peale, Gleb Shtengel, Melanie Freeman, Harald F. Hess, Craig Blackstone, and Jennifer Lippincott-Schwartz. Spastin tethers lipid droplets to peroxisomes and directs fatty acid trafficking through ESCRT– III. Journal of Cell Biology, 218(8):2583–??, August 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/8/2583.

Christ:2016:AEI

[CWL^{+}16] Liliane Christ, Eva M. Wenzel, Knut Liestøl, Camilla Raiborg, Coen Campsteijn, and Harald Stenmark. ALIX and ESCRT-I/ II function as parallel ESCRT-III recruiters in cytokinetic abscission. Journal of Cell Biology, 212(5):499–??, February 2016.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/212/5/499.

Capece:2017:NIP

[CWL^{+}17] Tara Capece, Brandon L. Walling, Kihong Lim, Kyun-Do Kim, Seyeon Bae, Hung-Li Chung, David J. Topham, and Minsoo Kim. A novel intracellular pool of LFA-1 is critical for asymmetric CD8^{+} T cell activation and differentiation. Journal of Cell Biology, 216(11):3817–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/11/3817.

Cheng:2015:PPP

[CWZ^{+}15] Shiya Cheng, Kun Wang, Wei Zou, Rui Miao, Yaling Huang, Haibin Wang, and Xiaochen Wang. PtdIns(4,5)P_{2} and PtdIns3P coordinate to regulate phagosomal sealing for apoptotic cell clearance. Journal of Cell Biology, 210(3):485–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/210/3/
485.

Cheng:2018:CLL

[CXZ^{+}18] Xiu-Tang Cheng, Yu-Xiang Xie, Bing Zhou, Ning Huang, Tamar Farfel-Becker, and Zu-Hang Sheng. Characterization of LAMP1-labeled nondegradative lysosomal and endocytic compartments in neurons. Journal of Cell Biology, 217(9):3127– ??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/9/3127.

Cai:2016:RMI

[CYH^{+}16] Yu Cai, Lu Yang, Guoku Hu, Xufeng Chen, Fang Niu, Li Yuan, Han Liu, Huangui Xiong, Jyothi Arikkath, and Shilpa Buch. Regulation of morphine-induced synaptic alterations: Role of oxidative stress, ER stress, and autophagy. Journal of Cell Biology, 215(2):245–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/2/245.

Chen:2018:SEE

[CYL^{+}18] Dan Chen, Chao Yang, Sha Liu, Weijian Hang, Xianghong Wang, Juan Chen, and Anbing Shi. SAC-1 ensures epithelial endocytic recycling by restricting ARF-6 activity. Journal of Cell Biology, 217(6):2121–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/6/2121.

Chen:2019:DIC

[CYMS^{+}19] Yan Chen, Jeffery Yong, Antonio Martínez-Sánchez, Yang Yang, Yumei Wu, Pietro De Camilli, Rubén Fernández-Busnadiego, and Min Wu. Dynamic instability of clathrin assembly provides proofreading control for endocytosis. Journal of Cell Biology, 218(10):3200–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/10/3200.

Chan:2018:KRU

[CYT^{+}18] Jonathan K. L. Chan, Don Yuen, Priscilla Hiu-Mei Too, Yan Sun, Belinda Willard, David Man, and Connie Tam. Keratin 6a reorganization for ubiquitin–proteasomal processing is a direct antimicrobial response. Journal of Cell Biology, 217(2): 731–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/2/731.

Cheng:2015:AAR

[CZL^{+}15] Xiu-Tang Cheng, Bing Zhou, Mei-Yao Lin, Qian Cai, and Zu-Hang Sheng. Axonal autophagosomes recruit dynein for retrograde transport through fusion with late endosomes. Journal of Cell Biology, 209(3):377–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/209/3/377.

Chaix:2016:CCC

[CZP16] Amandine Chaix, Amir Zarrinpar, and Satchidananda Panda. The circadian coordination of cell biology. Journal of Cell Biology, 215(1):15–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/1/15.

Chen:2018:MGO

[CZW^{+}18] Yu Chen, Yang Zhang, Yuchuan Wang, Liguo Zhang, Eva K. Brinkman, Stephen A. Adam, Robert Goldman, Bas van Steensel, Jian Ma, and Andrew S. Belmont. Mapping 3D genome organization relative to nuclear compartments using TSA–Seq as a cytological ruler. Journal of Cell Biology, 217(11):4025–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/11/4025.

Cao:2015:CRT

[CZZ^{+}15] Qi Cao, Xi Zoë Zhong, Yuanjie Zou, Ruth Murrell-Lagnado, Michael X. Zhu, and Xian-Ping Dong. Calcium release through P2X4 activates calmodulin to promote endolysosomal membrane fusion. Journal of Cell Biology, 209(6):879–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/209/6/879.

Dominguez:2015:TFZ

[DAG^{+}15] Claudia X. Dominguez, Robert A. Amezquita, Tianxia Guan, Heather D. Marshall, Nikhil S. Joshi, Steven H. Kleinstein, and Susan M. Kaech. The transcription factors ZEB2 and T-bet cooperate to program cytotoxic T cell terminal differentiation in response to LCMV viral infection. Journal of Cell Biology, 211 (3):??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/3/2113OIA258.

Dasso:2017:CRC

[Das17] Mary Dasso. Catch and release: 14-3-3 controls Ncd in meiotic spindles. Journal of Cell Biology, 216(10):3003–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/10/3003.

Davidson:2018:WFP

[DATI18] Andrew J. Davidson, Clelia Amato, Peter A. Thomason, and Robert H. Insall. WASP family proteins and formins compete in pseudopod- and bleb-based migration. Journal of Cell Biology, 217(2):701–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/2/701.

Donovan:2015:HTR

[DB15a] Kirk W. Donovan and Anthony Bretscher. Head-to-tail regulation is critical for the in vivo function of myosin V. Journal of Cell Biology, 209(3):359–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/209/3/359.

Donovan:2015:TIS

[DB15b] Kirk W. Donovan and Anthony Bretscher. Tracking individual secretory vesicles during exocytosis reveals an ordered and regulated process. Journal of Cell Biology, 210(2):181–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/210/2/
181.

Dart:2015:PPK

[DBC^{+}15] Anna E. Dart, Gary M. Box, William Court, Madeline E. Gale, John P. Brown, Sarah E. Pinder, Suzanne A. Eccles, and Claire M. Wells. PAK4 promotes kinase-independent stabilization of RhoU to modulate cell adhesion. Journal of Cell Biology, 211(4):863–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/4/863.

Dudin:2015:FNA

[DBG^{+}15] Omaya Dudin, Felipe O. Bendezú, Raphael Groux, Thierry Laroche, Arne Seitz, and Sophie G. Martin. A formin-nucleated actin aster concentrates cell wall hydrolases for cell fusion in fission yeast. Journal of Cell Biology, 208(7):897–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/208/7/897.

Dey:2018:UCG

[DBS18] Sandip Dey, Chiranjit Biswas, and Jayati Sengupta. The universally conserved GTPase HflX is an RNA helicase that restores heat-damaged Escherichia coli ribosomes. Journal of Cell Biology, 217(7):2519–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/7/2519.

Donninger:2015:NRS

[DCB^{+}15] Howard Donninger, Diego F. Calvisi, Thibaut Barnoud, Jennifer Clark, M. Lee Schmidt, Michele D. Vos, and Geoffrey J. Clark. NORE1A is a Ras senescence effector that controls the apoptotic/senescent balance of p53 via HIPK2. Journal of Cell Biology, 208(6):777–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/6/777.

Dyson:2017:IRP

[DCF^{+}17] Jennifer M. Dyson, Sarah E. Conduit, Sandra J. Feeney, Sandra Hakim, Tia DiTommaso, Alex J. Fulcher, Absorn Sriratana, Georg Ramm, Kristy A. Horan, Rajendra Gurung, Carol Wicking, Ian Smyth, and Christina A. Mitchell. INPP5E regulates phosphoinositide-dependent cilia transition zone function. Journal of Cell Biology, 216(1):247–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/1/247.

Donnelly:2017:RRB

[DCM^{+}17] Sara K. Donnelly, Ramon Cabrera, Serena P. H. Mao, John R. Christin, Bin Wu, Wenjun Guo, Jose Javier Bravo-Cordero, John S. Condeelis, Jeffrey E. Segall, and Louis Hodgson. Rac3 regulates breast cancer invasion and metastasis by controlling adhesion and matrix degradation. Journal of Cell Biology, 216(12):4331–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/12/4331.

Dornier:2012:TTR

[DCO^{+}12] Emmanuel Dornier, Franck Coumailleau, Jean-François Ottavi, Julien Moretti, Claude Boucheix, Philippe Mauduit, François Schweisguth, and Eric Rubinstein. TspanC8 tetraspanins regulate ADAM10/Kuzbanian trafficking and promote Notch activation in flies and mammals. Journal of Cell Biology, 199 (3):481–??, October 2012. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress. org/content/199/3/481. See correction [DCO^{+}16].

Dornier:2016:CTT

[DCO^{+}16] Emmanuel Dornier, Franck Coumailleau, Jean-François Ottavi, Julien Moretti, Claude Boucheix, Philippe Mauduit, François Schweisguth, and Eric Rubinstein. Correction: TspanC8 tetraspanins regulate ADAM10/Kuzbanian trafficking and promote Notch activation in flies and mammals. Journal of Cell Biology, 213(4):495–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/213/4/495. See [DCO^{+}12].

Dimou:2019:SEV

[DCP^{+}19] Eleni Dimou, Katia Cosentino, Evgenia Platonova, Uris Ros, Mohsen Sadeghi, Purba Kashyap, Taxiarchis Katsinelos, Sabine Wegehingel, Frank Noé, Ana J. García-Sáez, Helge Ewers, and Walter Nickel. Single event visualization of unconventional secretion of FGF2. Journal of Cell Biology, 218(2):683–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/2/
683.

Deneke:2018:CWC

[DD18] Victoria E. Deneke and Stefano Di Talia. Chemical waves in cell and developmental biology. Journal of Cell Biology, 217 (4):1193–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/4/1193.

Dennis:2016:BBR

[DDAR^{+}16] Megan K. Dennis, Cédric Delevoye, Amanda Acosta-Ruiz, Ilse Hurbain, Maryse Romao, Geoffrey G. Hesketh, Philip S. Goff, Elena V. Sviderskaya, Dorothy C. Bennett, J. Paul Luzio, Thierry Galli, David J. Owen, Graça Raposo, and Michael S. Marks. BLOC-1 and BLOC-3 regulate VAMP7 cycling to and from melanosomes via distinct tubular transport carriers. Journal of Cell Biology, 214(3):293–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/3/293.

Dharmat:2018:SMN

[DER^{+}18] Rachayata Dharmat, Aiden Eblimit, Michael A. Robichaux, Zhixian Zhang, Thanh-Minh T. Nguyen, Sung Yun Jung, Feng He, Antrix Jain, Yumei Li, Jun Qin, Paul Overbeek, Ronald Roepman, Graeme Mardon, Theodore G. Wensel, and Rui Chen. SPATA7 maintains a novel photoreceptor-specific zone in the distal connecting cilium. Journal of Cell Biology, 217(8): 2851–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/8/2851.

Donnelly:2018:NCR

[DGS^{+}18] Christopher R. Donnelly, Nicole A. Gabreski, Esther B. Suh, Monzurul Chowdhury, and Brian A. Pierchala. Non-canonical Ret signaling augments p75-mediated cell death in developing sympathetic neurons. Journal of Cell Biology, 217(9): 3237–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/9/3237.

Dickson:2017:RRP

[Dic17] Eamonn J. Dickson. RASSF4: Regulator of plasma membrane PI(4,5)P_{2}. Journal of Cell Biology, 216(7):1879–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/7/
1879.

Dickson:2016:DFP

[DJV^{+}16] Eamonn J. Dickson, Jill B. Jensen, Oscar Vivas, Martin Kruse, Alexis E. Traynor-Kaplan, and Bertil Hille. Dynamic formation of ER–PM junctions presents a lipid phosphatase to regulate phosphoinositides. Journal of Cell Biology, 213(1):33–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/213/1/
33.

Dustin:2016:TMC

[DK16] Michael L. Dustin and Lance C. Kam. Tapping out a mechanical code for T cell triggering. Journal of Cell Biology, 213 (5):501–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/5/501.

Das:2017:SJF

[DK17] Shradha Das and Elisabeth Knust. Stardust, the Janus-faced partner of Crumbs. Journal of Cell Biology, 216(5):1219–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/5/1219.

Dubash:2016:PLP

[DKA^{+}16] Adi D. Dubash, Chen Y. Kam, Brian A. Aguado, Dipal M. Patel, Mario Delmar, Lonnie D. Shea, and Kathleen J. Green. Plakophilin-2 loss promotes TGF-β1/p38 MAPK-dependent fibrotic gene expression in cardiomyocytes. Journal of Cell Biology, 212(4):425–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/4/425.

DeGeer:2015:HCA

[DKM^{+}15] Jonathan DeGeer, Andrew Kaplan, Pierre Mattar, Morgane Morabito, Ursula Stochaj, Timothy E. Kennedy, Anne Debant, Michel Cayouette, Alyson E. Fournier, and Nathalie Lamarche-Vane. Hsc70 chaperone activity underlies Trio GEF function in axon growth and guidance induced by netrin-1. Journal of Cell Biology, 210(5):817–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/5/817.

Ducuing:2015:DMF

[DKMV15] Antoine Ducuing, Charlotte Keeley, Bertrand Mollereau, and Stéphane Vincent. A DPP-mediated feed-forward loop canalizes morphogenesis during Drosophila dorsal closure. Journal of Cell Biology, 208(2):239–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/208/2/239.

Dwivedi:2019:CDA

[DKR^{+}19a] Devashish Dwivedi, Amrita Kumari, Siddhi Rathi, Sivaram V. S. Mylavarapu, and Mahak Sharma. Correction: The dynein adaptor Hook2 plays essential roles in mitotic progression and cytokinesis. Journal of Cell Biology, 218(10):3526–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/

218/10/3526. See [DKR^{+}19b].

Dwivedi:2019:DAH

[DKR^{+}19b] Devashish Dwivedi, Amrita Kumari, Siddhi Rathi, Sivaram V. S. Mylavarapu, and Mahak Sharma. The dynein adaptor Hook2 plays essential roles in mitotic progression and cytokinesis. Journal of Cell Biology, 218(3):871–??, March 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/3/871. See cor-

rection [DKR^{+}19a].

Drosopoulos:2015:BHF

[DKS15] William C. Drosopoulos, Settapong T. Kosiyatrakul, and Carl L. Schildkraut. BLM helicase facilitates telomere replication during leading strand synthesis of telomeres. Journal of Cell Biology, 210(2):191–??, July 2015. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/2/191.

Dumas:2015:HPV

[DLBMA^{+}15] Audrey Dumas, Gabrielle Lê-Bury, Florence Marie-Anaïs, Floriane Herit, Julie Mazzolini, Thomas Guilbert, Pierre Bourdoncle, David G. Russell, Serge Benichou, Ahmed Zahraoui, and Florence Niedergang. The HIV-1 protein Vpr impairs phagosome maturation by controlling microtubule-dependent trafficking. Journal of Cell Biology, 211(2):359–??, October 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/211/2/359.

delaFuente:2015:VDR

[dlFEvW^{+}15] Alerie Guzman de la Fuente, Oihana Errea, Peter van Wijngaarden, Ginez A. Gonzalez, Christophe Kerninon, Andrew A. Jarjour, Hilary J. Lewis, Clare A. Jones, Brahim Nait-Oumesmar, Chao Zhao, Jeffrey K. Huang, Charles ffrench Constant, and Robin J. M. Franklin. Vitamin D receptor–retinoid X receptor heterodimer signaling regulates oligodendrocyte progenitor cell differentiation. Journal of Cell Biology, 211(5): 975–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/5/975.

Datta:2019:CAD

[DLH^{+}19] Sanchari Datta, Yang Liu, Hanaa Hariri, Jade Bowerman, and W. Mike Henne. Cerebellar ataxia disease–associated Snx14 promotes lipid droplet growth at ER–droplet contacts. Journal of Cell Biology, 218(4):1335–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/4/1335.

Derive:2015:BBD

[DLM^{+}15] Nicolas Derive, Cedric Landmann, Emilie Montembault, Marie-Charlotte Claverie, Priscillia Pierre-Elies, Damien Goutte-Gattat, Nabila Founounou, Derek McCusker, and Anne Royou. Bub3–BubR1-dependent sequestration of Cdc20^{Fizzy} at DNA breaks facilitates the correct segregation of broken chromosomes. Journal of Cell Biology, 211(3):517–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/211/3/517.

delaRoche:2018:TCR

[dlRHM^{+}18] Marianne de la Roche, Claire Hamilton, Rebecca Mortensen, A. Arockia Jeyaprakash, Sanjay Ghosh, and Paras K. Anand. Trafficking of cholesterol to the ER is required for NLRP3 inflammasome activation. Journal of Cell Biology, 217(10): 3560–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/10/3560.

Drummond:2018:APC

[DLT^{+}18] Michael L. Drummond, Mischa Li, Eric Tarapore, Tuyen T. L. Nguyen, Baina J. Barouni, Shaun Cruz, Kevin C. Tan, Anthony E. Oro, and Scott X. Atwood. Actin polymerization controls cilia-mediated signaling. Journal of Cell Biology, 217(9): 3255–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/9/3255.

Draheim:2015:CCI

[DLZ^{+}15] Kyle M. Draheim, Xiaofeng Li, Rong Zhang, Oriana S. Fisher, Giulia Villari, Titus J. Boggon, and David A. Calderwood. CCM2–CCM3 interaction stabilizes their protein expression and permits endothelial network formation. Journal of Cell Biology, 208(7):987–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/7/987.

DeLuca:2018:AKP

[DMB^{+}18] Keith F. DeLuca, Amanda Meppelink, Amanda J. Broad, Jeanne E. Mick, Olve B. Peersen, Sibel Pektas, Susanne M. A. Lens, and Jennifer G. DeLuca. Aurora A kinase phosphorylates Hec1 to regulate metaphase kinetochore–microtubule dynamics. Journal of Cell Biology, 217(1):163–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/217/1/163.

DAmico:2016:DRP

[DMC^{+}16] Lucia D’Amico, Sahil Mahajan, Aude-Helene Capietto, Zhengfeng Yang, Ali Zamani, Biancamaria Ricci, David B. Bumpass, Melissa A. Meyer, Xinming Su, Andrea Wang-Gillam, Kathy N. Weilbaecher, Sheila Ann Stewart, David DeNardo, and Roberta Faccio. Dickkopf-related protein 1 (Dkk1) regulates the accumulation and function of myeloid derived suppressor cells in cancer. Journal of Cell Biology, 213(1):??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/1/2131OIA66.

Distefano:2017:HSI

[DMC^{+}17] Ayelén Mariana Distéfano, María Victoria Martin, Juan Pablo Córdoba, Andrés Martín Bellido, Sebastián D’Ippólito, Silvana Lorena Colman, Débora Soto, Juan Alfredo Roldán, Carlos Guillermo Bartoli, Eduardo Julián Zabaleta, Diego Fernando Fiol, Brent R. Stockwell, Scott J. Dixon, and Gabriela Carolina Pagnussat. Heat stress induces ferroptosislike cell death in plants. Journal of Cell Biology, 216(2): 463–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/2/463.

DiCarlo:2019:PCN

[DMD19] Valerio Di Carlo, Ivano Mocavini, and Luciano Di Croce. Polycomb complexes in normal and malignant hematopoiesis. Journal of Cell Biology, 218(1):55–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/1/55.

Delaney:2019:HFT

[DMG^{+}19] Colin E. Delaney, Stephen P. Methot, Micol Guidi, Iskra Katic, Susan M. Gasser, and Jan Padeken. Heterochromatic foci and transcriptional repression by an unstructured MET-2/SETDB1 co-factor LIN-65. Journal of Cell Biology, 218(3):820–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/3/
820.

Ditlev:2015:CSN

[DMH^{+}15] Jonathon A. Ditlev, Paul J. Michalski, Greg Huber, Gonzalo M. Rivera, William A. Mohler, Leslie M. Loew, and Bruce J. Mayer. Correction: Stoichiometry of Nck-dependent actin polymerization in living cells. Journal of Cell Biology, 211(5): 1095–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/5/1095.

Dennis:2015:BTR

[DMS^{+}15] Megan K. Dennis, Adriana R. Mantegazza, Olivia L. Snir, Danièle Tenza, Amanda Acosta-Ruiz, Cédric Delevoye, Richard Zorger, Anand Sitaram, Wilfredo de Jesus-Rojas, Keerthana Ravichandran, John Rux, Elena V. Sviderskaya, Dorothy C. Bennett, Graça Raposo, Michael S. Marks, and Subba Rao Gangi Setty. BLOC-2 targets recycling endosomal tubules to melanosomes for cargo delivery. Journal of Cell Biology, 209(4):563–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/4/563.

Derrer:2019:REF

[DMV^{+}19] Carina Patrizia Derrer, Roberta Mancini, Pascal Vallotton, Sébastien Huet, Karsten Weis, and Elisa Dultz. The RNA export factor Mex67 functions as a mobile nucleoporin. Journal of Cell Biology, 218(12):3967–3976, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

https://rupress.org/jcb/article/218/12/3967/132517/
The-RNA-export-factor-Mex67-functions-as-a-mobile.

Dmitrieff:2016:AAP

[DN16] Serge Dmitrieff and François Nédélec. Amplification of actin polymerization forces. Journal of Cell Biology, 212(7):763–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
212/7/763.

Dornier:2017:TLE

[DN17] Emmanuel Dornier and Jim C. Norman. Tensin links energy metabolism to extracellular matrix assembly. Journal of Cell Biology, 216(4):867–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/4/867.

Das:2016:EMI

[DNMB16] Anupam Das, Sagarika Nag, Anne B. Mason, and Margarida M. Barroso. Endosome–mitochondria interactions are modulated by iron release from transferrin. Journal of Cell Biology, 214(7): 831–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/214/7/831.

De:2017:VCC

[DOA^{+}17] Mithu De, Austin N. Oleskie, Mariam Ayyash, Somnath Dutta, Liliya Mancour, Mohamed E. Abazeed, Eddy J. Brace, Georgios Skiniotis, and Robert S. Fuller. The Vps13p–Cdc31p complex is directly required for TGN late endosome transport and TGN homotypic fusion. Journal of Cell Biology, 216(2): 425–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/2/425.

Ding:2017:PIC

[DOH^{+}17] Wei Yung Ding, Hui Ting Ong, Yusuke Hara, Jantana Wongsantichon, Yusuke Toyama, Robert C. Robinson, François Nédélec, and Ronen Zaidel-Bar. Plastin increases cortical connectivity to facilitate robust polarization and timely cytokinesis. Journal of Cell Biology, 216(5):1371–??, May 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/5/1371.

DePascalis:2018:IFC

[DPGS^{+}18] Chiara De Pascalis, Carlos Pérez-González, Shailaja Seetharaman, Batiste Boëda, Benoit Vianay, Mithila Burute, Cécile Leduc, Nicolas Borghi, Xavier Trepat, and Sandrine Etienne-Manneville. Intermediate filaments control collective migration by restricting traction forces and sustaining cell–cell contacts. Journal of Cell Biology, 217(9):3031–??, September 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/9/3031.

Dandoulaki:2018:SAC

[DPS^{+}18] Maria Dandoulaki, Eleni Petsalaki, David Sumpton, Sara Zanivan, and George Zachos. Src activation by Chk1 promotes actin patch formation and prevents chromatin bridge breakage in cytokinesis. Journal of Cell Biology, 217(9):3071–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/9/
3071.

Demonbreun:2016:ADA

[DQB^{+}16] Alexis R. Demonbreun, Mattia Quattrocelli, David Y. Barefield, Madison V. Allen, Kaitlin E. Swanson, and Elizabeth M. McNally. An actin-dependent annexin complex mediates plasma membrane repair in muscle. Journal of Cell Biology, 213(6):705–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/213/6/705.

Deshpande:2016:BSM

[DR16] Mugdha Deshpande and Avital A. Rodal. Beyond the SNARE: Munc18-1 chaperones α-synuclein. Journal of Cell Biology, 214(6):641–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/6/641.

Signore:2019:EME

[DR19] Steven J. Del Signore and Avital A. Rodal. The enemy of my enemy: PTEN and PLCXD collude to fight endosomal PtdIns(4,5)P_{2}. Journal of Cell Biology, 218(7):2082–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/7/
2082.

David:2019:AAL

[DRL^{+}19] Ana F. David, Philippe Roudot, Wesley R. Legant, Eric Betzig, Gaudenz Danuser, and Daniel W. Gerlich. Augmin accumulation on long-lived microtubules drives amplification and kinetochore-directed growth. Journal of Cell Biology, 218 (7):2150–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/7/2150.

Davis-Roca:2017:CEO

[DRMW17] Amanda C. Davis-Roca, Christina C. Muscat, and Sarah M. Wignall. Caenorhabditis elegans oocytes detect meiotic errors in the absence of canonical end-on kinetochore attachments. Journal of Cell Biology, 216(5):1243–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/5/1243.

Dantuma:2016:UVM

[DS16a] Nico P. Dantuma and Florian A. Salomons. Ubiquitin versus misfolding: The minimal requirements for inclusion body formation. Journal of Cell Biology, 213(2):147–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/213/2/147.

Dolat:2016:SPM

[DS16b] Lee Dolat and Elias T. Spiliotis. Septins promote macropinosome maturation and traffic to the lysosome by facilitating membrane fusion. Journal of Cell Biology, 214(5):517–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/214/5/
517.

Ding:2015:DMR

[DSA15] Jin-Dong Ding, Raquel Y. Salinas, and Vadim Y. Arshavsky. Discs of mammalian rod photoreceptors form through the membrane evagination mechanism. Journal of Cell Biology, 211(3): 495–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/3/495.

Dambournet:2018:GEH

[DSC^{+}18] Daphné Dambournet, Kem A. Sochacki, Aaron T. Cheng, Matthew Akamatsu, Justin W. Taraska, Dirk Hockemeyer, and David G. Drubin. Genome-edited human stem cells expressing fluorescently labeled endocytic markers allow quantitative analysis of clathrin-mediated endocytosis during differentiation. Journal of Cell Biology, 217(9):3301–??, September 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/9/3301.

Dionne:2018:CAD

[DSH^{+}18] Lai Kuan Dionne, Kyuhwan Shim, Masato Hoshi, Tao Cheng, Jinzhi Wang, Veronique Marthiens, Amanda Knoten, Renata Basto, Sanjay Jain, and Moe R. Mahjoub. Centrosome amplification disrupts renal development and causes cystogenesis. Journal of Cell Biology, 217(7):2485–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/7/2485.

Dhatchinamoorthy:2017:SPL

[DSL^{+}17] Karthik Dhatchinamoorthy, Manjunatha Shivaraju, Jeffrey J. Lange, Boris Rubinstein, Jay R. Unruh, Brian D. Slaughter, and Jennifer L. Gerton. Structural plasticity of the living kinetochore. Journal of Cell Biology, 216(11):3551–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/
11/3551.

Dhanyasi:2015:SAM

[DSS^{+}15] Nagaraju Dhanyasi, Dagan Segal, Eyal Shimoni, Vera Shinder, Ben-Zion Shilo, K. VijayRaghavan, and Eyal D. Schejter. Surface apposition and multiple cell contacts promote myoblast fusion in Drosophila flight muscles. Journal of Cell Biology, 211 (1):191–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/1/191.

Domingo-Sananes:2015:MCI

[DSSF^{+}15] Maria Rosa Domingo-Sananes, Balazs Szöőr, Michael A. J. Ferguson, Michael D. Urbaniak, and Keith R. Matthews. Molecular control of irreversible bistability during trypanosome developmental commitment. Journal of Cell Biology, 211(2):455–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
211/2/455.

Daneshjou:2015:RFRa

[DSvNA^{+}15a] Nazila Daneshjou, Nathan Sieracki, Geerten P. van Nieuw Amerongen, Daniel E. Conway, Martin A. Schwartz, Yulia A. Komarova, and Asrar B. Malik. Rac1 functions as a reversible tension modulator to stabilize VE-cadherin trans-interaction. Journal of Cell Biology, 208(1):23–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/1/23.

Daneshjou:2015:RFRb

[DSvNA^{+}15b] Nazila Daneshjou, Nathan Sieracki, Geerten P. van Nieuw Amerongen, Daniel E. Conway, Martin A. Schwartz, Yulia A. Komarova, and Asrar B. Malik. Rac1 functions as a reversible tension modulator to stabilize VE-cadherin trans-interaction. Journal of Cell Biology, 209(1):181–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/209/1/181.

Dultz:2016:GRB

[DTW^{+}16] Elisa Dultz, Harianto Tjong, Elodie Weider, Mareike Herzog, Barry Young, Christiane Brune, Daniel Müllner, Christopher Loewen, Frank Alber, and Karsten Weis. Global reorganization of budding yeast chromosome conformation in different physiological conditions. Journal of Cell Biology, 212(3):

321–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/3/321.

Dhatchinamoorthy:2019:SOK

[DUL^{+}19] Karthik Dhatchinamoorthy, Jay R. Unruh, Jeffrey J. Lange, Michaella Levy, Brian D. Slaughter, and Jennifer L. Gerton. The stoichiometry of the outer kinetochore is modulated by microtubule-proximal regulatory factors. Journal of Cell Biology, 218(7):2124–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/7/2124.

Dierickx:2016:MCV

[DV16] Pieterjan Dierickx and Linda W. Van Laake. Muscle-on-chip: an in vitro model for donor–host cardiomyocyte coupling. Journal of Cell Biology, 212(4):371–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/212/4/371.

deValle:2016:NEP

[dVGO^{+}16] Elisha de Valle, George Grigoriadis, Lorraine A. O’Reilly, Simon N. Willis, Mhairi J. Maxwell, Lynn M. Corcoran, Evelyn Tsantikos, Jasper K. S. Cornish, Kirsten A. Fairfax, Ajithkumar Vasanthakumar, Mark A. Febbraio, Margaret L. Hibbs, Marc Pellegrini, Ashish Banerjee, Philip D. Hodgkin, Axel Kallies, Fabienne Mackay, Andreas Strasser, Steve Gerondakis, and Raffi Gugasyan. NFκB1 is essential to prevent the development of multiorgan autoimmunity by limiting IL-6 production in follicular B cells. Journal of Cell Biology, 213(1):??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/213/1/
2131OIA67.

Druelle:2017:EEP

[DVS^{+}17] Noémie Druelle, Andhira Vieira, Aidin Shabro, Monica Courtney, Magali Mondin, Samah Rekima, Tiziana Napolitano, Serena Silvano, Sergi Navarro-Sanz, Biljana Hadzic, Fabio Avolio, Minoo Rassoulzadegan, Herbert A. Schmid, Ahmed Mansouri, and Patrick Collombat. Ectopic expression of Pax4 in pancreatic δ cells results in β-like cell neogenesis. Journal of Cell Biology, 216(12):4299–??, December 2017. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/12/4299.

DArchivio:2017:TOK

[DW17] Simon D’Archivio and Bill Wickstead. Trypanosome outer kinetochore proteins suggest conservation of chromosome segregation machinery across eukaryotes. Journal of Cell Biology, 216 (2):379–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/2/379.

Divakaruni:2017:IMP

[DWB^{+}17] Ajit S. Divakaruni, Martina Wallace, Caodu Buren, Kelly Martyniuk, Alexander Y. Andreyev, Edward Li, Jerel A. Fields, Thekla Cordes, Ian J. Reynolds, Brenda L. Bloodgood, Lynn A. Raymond, Christian M. Metallo, and Anne N. Murphy. Inhibition of the mitochondrial pyruvate carrier protects from excitotoxic neuronal death. Journal of Cell Biology, 216(4):1091–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/4/1091.

Daste:2017:CAP

[DWH^{+}17a] Frederic Daste, Astrid Walrant, Mikkel R. Holst, Jonathan R. Gadsby, Julia Mason, Ji-Eun Lee, Daniel Brook, Marcel Mettlen, Elin Larsson, Steven F. Lee, Richard Lundmark, and Jennifer L. Gallop. Control of actin polymerization via the coincidence of phosphoinositides and high membrane curvature. Journal of Cell Biology, 216(11):3745–??, November 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/11/3745.

Ding:2017:ONM

[DWH^{+}17b] Zhao-Ying Ding, Ying-Hsuan Wang, Yu-Cheng Huang, Myong-Chol Lee, Min-Jen Tseng, Ya-Hui Chi, and Min-Lang Huang. Outer nuclear membrane protein Kuduk modulates the LINC complex and nuclear envelope architecture. Journal of Cell Biology, 216(9):2827–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/9/2827.

Dong:2018:IPI

[DZB^{+}18] Rui Dong, Ting Zhu, Lorena Benedetti, Swetha Gowrishankar, Huichao Deng, Yiying Cai, Xiangming Wang, Kang Shen, and Pietro De Camilli. The inositol 5-phosphatase INPP5K participates in the fine control of ER organization. Journal of Cell Biology, 217(10):3577–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/217/10/3577.

Dong:2015:CPD

[DZL^{+}15] Wei Dong, Xuejing Zhang, Weijie Liu, Yi jiun Chen, Juan Huang, Erin Austin, Alicia M. Celotto, Wendy Z. Jiang, Michael J. Palladino, Yu Jiang, Gerald R. V. Hammond, and Yang Hong. A conserved polybasic domain mediates plasma membrane targeting of Lgl and its regulation by hypoxia. Journal of Cell Biology, 211(2):273–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/211/2/273.

Erdogan:2017:CAF

[EAW^{+}17] Begum Erdogan, Mingfang Ao, Lauren M. White, Anna L. Means, Bryson M. Brewer, Lijie Yang, M. Kay Washington, Chanjuan Shi, Omar E. Franco, Alissa M. Weaver, Simon W. Hayward, Deyu Li, and Donna J. Webb. Cancer-associated fibroblasts promote directional cancer cell migration by aligning fibronectin. Journal of Cell Biology, 216(11):3799–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/11/3799.

Eisenberg-Bord:2018:ISL

[EBMW^{+}18] Michal Eisenberg-Bord, Muriel Mari, Uri Weill, Eden Rosenfeld-Gur, Ofer Moldavski, Inês G. Castro, Krishnakant G. Soni, Nofar Harpaz, Tim P. Levine, Anthony H. Futerman, Fulvio Reggiori, Vytas A. Bankaitis, Maya Schuldiner, and Maria Bohnert. Identification of seipin-linked factors that act as determinants of a lipid droplet subpopulation. Journal of Cell Biology, 217 (1):269–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/1/269.

Enemchukwu:2016:SMR

[ECAB^{+}16] Nduka O. Enemchukwu, Ricardo Cruz-Acuña, Tom Bongiorno, Christopher T. Johnson, José R. García, Todd Sulchek, and Andrés J. García. Synthetic matrices reveal contributions of ECM biophysical and biochemical properties to epithelial morphogenesis. Journal of Cell Biology, 212(1):113–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/212/1/
113.

Eggeling:2017:MMA

[ED17] Christian Eggeling and Simon J. Davis. Macrophages: micromanagers of antagonistic signaling nanoclusters. Journal of Cell Biology, 216(4):871–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/4/871.

Encarnacao:2016:RDC

[EEE^{+}16] Marisa Encarnação, Lília Espada, Cristina Escrevente, Denisa Mateus, José Ramalho, Xavier Michelet, Inês Santarino, Victor W. Hsu, Michael B. Brenner, Duarte C. Barral, and Otília V. Vieira. A Rab3a-dependent complex essential for lysosome positioning and plasma membrane repair. Journal of Cell Biology, 213(6):631–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/213/6/631.

Estrem:2017:DRS

[EFM17] Cassi Estrem, Colby P. Fees, and Jeffrey K. Moore. Dynein is regulated by the stability of its microtubule track. Journal of Cell Biology, 216(7):2047–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/7/2047.

Edzuka:2019:DKS

[EG19] Tomoya Edzuka and Gohta Goshima. Drosophila kinesin-8 stabilizes the kinetochore–microtubule interaction. Journal of Cell Biology, 218(2):474–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/2/474.

Eykelenboom:2019:LIM

[EGY^{+}19] John K. Eykelenboom, Marek Gierliński, Zuojun Yue, Nadia Hegarat, Hilary Pollard, Tatsuo Fukagawa, Helfrid Hochegger, and Tomoyuki U. Tanaka. Live imaging of marked chromosome regions reveals their dynamic resolution and compaction in mitosis. Journal of Cell Biology, 218(5):1531–??, May 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/5/1531.

Edgerton:2016:NFT

[EJK^{+}16] Heather Edgerton, Marnie Johansson, Daniel Keifenheim, Soumya Mukherjee, Jeremy M. Chacón, Jeff Bachant, Melissa K. Gardner, and Duncan J. Clarke. A noncatalytic function of the topoisomerase II CTD in Aurora B recruitment to inner centromeres during mitosis. Journal of Cell Biology, 213(6):651–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
213/6/651.

Engevik:2019:LMV

[EKP^{+}19] Amy C. Engevik, Izumi Kaji, Meagan M. Postema, James J. Faust, Anne R. Meyer, Janice A. Williams, Gillian N. Fitz, Matthew J. Tyska, Jean M. Wilson, and James R. Goldenring. Loss of myosin Vb promotes apical bulk endocytosis in neonatal enterocytes. Journal of Cell Biology, 218(11):3647–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/11/3647.

Ernesto:2015:CNC

[EMB^{+}15] Juan I. Ernesto, Mariana Weigel Muñoz, María A. Battistone, Gustavo Vasen, Pablo Martínez-López, Gerardo Orta, Dulce Figueiras-Fierro, José L. De la Vega-Beltran, Ignacio A. Moreno, Héctor A. Guidobaldi, Laura Giojalas, Alberto Darszon, Débora J. Cohen, and Patricia S. Cuasnicú. CRISP1 as a novel CatSper regulator that modulates sperm motility and orientation during fertilization. Journal of Cell Biology, 210(7):1213–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/7/1213.

Evangelista:2018:TME

[EMRS^{+}18] Federica M. Evangelista, Anne Maglott-Roth, Matthieu Stierle, Laurent Brino, Evi Soutoglou, and László Tora. Transcription and mRNA export machineries SAGA and TREX-2 maintain monoubiquitinated H2B balance required for DNA repair. Journal of Cell Biology, 217(10):3382–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/10/3382.

Ezratty:2016:PAT

[EPF16] Ellen J. Ezratty, H. Amalia Pasolli, and Elaine Fuchs. A Presenilin-2–ARF4 trafficking axis modulates Notch signaling during epidermal differentiation. Journal of Cell Biology, 214 (1):89–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/214/1/89.

Efimova:2018:BAN

[ES18] Nadia Efimova and Tatyana M. Svitkina. Branched actin networks push against each other at adherens junctions to maintain cell–cell adhesion. Journal of Cell Biology, 217(5):1827–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/5/1827.

Ebner:2017:LMA

[ESS^{+}17] Michael Ebner, Benjamin Sinkovics, Magdalena Szczygieł, Daniela Wolfschoon Ribeiro, and Ivan Yudushkin. Localization of mTORC2 activity inside cells. Journal of Cell Biology, 216(2):343–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/2/343.

Elazar:2019:CIP

[EVR^{+}19] Nimrod Elazar, Anya Vainshtein, Katya Rechav, Michael Tsoory, Yael Eshed-Eisenbach, and Elior Peles. Coordinated internodal and paranodal adhesion controls accurate myelination by oligodendrocytes. Journal of Cell Biology, 218(9): 2887–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/9/2887.

Elzeneini:2017:LLL

[EW17] Eman Elzeneini and Sara A. Wickström. Lipodystrophic laminopathy: Lamin A mutation relaxes chromatin architecture to impair adipogenesis. Journal of Cell Biology, 216(9): 2607–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/9/2607.

Ewers:2018:OSR

[Ewe18] Helge Ewers. Open-source recombinant monoclonal secondary nanobodies. Journal of Cell Biology, 217(3):809–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/3/
809.

ElAzzouzi:2016:MMM

[EWL16] Karim El Azzouzi, Christiane Wiesner, and Stefan Linder. Metalloproteinase MT1-MMP islets act as memory devices for podosome reemergence. Journal of Cell Biology, 213(1):109–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
213/1/109.

Frankel:2016:BCB

[FA16] E. B. Frankel and Anjon Audhya. Burning cellular bridges: Two pathways to the big breakup. Journal of Cell Biology, 212 (5):491–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/5/491.

Foldi:2017:TTR

[FAH^{+}17] Istvan Foldi, Niki Anthoney, Neale Harrison, Monique Gangloff, Brett Verstak, Mohanakarthik Ponnadai Nallasivan, Samaher AlAhmed, Bangfu Zhu, Mark Phizacklea, Maria Losada-Perez, Marta Moreira, Nicholas J. Gay, and Alicia Hidalgo. Three-tier regulation of cell number plasticity by neurotrophins and Tolls in Drosophila. Journal of Cell Biology, 216(5):1421–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/5/

Farhan:2016:RES

[Far16] Hesso Farhan. Regulation of EGFR surface levels by COPIIdependent trafficking. Journal of Cell Biology, 215(4):441– ??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/215/4/441.

Fenix:2015:SPM

[FB15] Aidan M. Fenix and Dylan T. Burnette. A small part of myosin IIB takes on a big role in cell polarity. Journal of Cell Biology, 209(1):11–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/1/11.

Fernandez-Barrera:2018:AMS

[FBBRCA^{+}18] Jaime Fernández-Barrera, Miguel Bernabé-Rubio, Javier Casares-Arias, Laura Rangel, Laura Fernández-Martín, Isabel Correas, and Miguel A. Alonso. The actin–MRTF–SRF transcriptional circuit controls tubulin acetylation via α-TAT1 gene expression. Journal of Cell Biology, 217(3):929–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/217/3/929.

Fessenden:2018:DDA

[FBPN^{+}18] Tim B. Fessenden, Yvonne Beckham, Mathew Perez-Neut, Guillermina Ramirez-San Juan, Aparajita H. Chourasia, Kay F. Macleod, Patrick W. Oakes, and Margaret L. Gardel. Dia1-dependent adhesions are required by epithelial tissues to initiate invasion. Journal of Cell Biology, 217(4):1485–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/4/
1485.

Fu:2015:TPM

[FBX^{+}15] Jingyan Fu, Minglei Bian, Guangwei Xin, Zhaoxuan Deng, Jia Luo, Xiao Guo, Hao Chen, Yao Wang, Qing Jiang, and Chuanmao Zhang. TPX2 phosphorylation maintains metaphase spindle length by regulating microtubule flux. Journal of Cell Biology, 210(3):373–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/3/373.

Feng:2015:RDB

[FC15] Xia Feng and Pierre A. Coulombe. A role for disulfide bonding in keratin intermediate filament organization and dynamics in skin keratinocytes. Journal of Cell Biology, 209(1):59–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/209/1/
59.

Futter:2016:CGB

[FC16] Clare E. Futter and Daniel F. Cutler. Coming or going? Un– BLOC–ing delivery and recycling pathways during melanosome maturation. Journal of Cell Biology, 214(3):245–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/214/3/
245.

Franco:2019:ESC

[FC19] Maribel Franco and Ana Carmena. Eph signaling controls mitotic spindle orientation and cell proliferation in neuroepithelial cells. Journal of Cell Biology, 218(4):1200–??, April 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/4/1200.

Francavilla:2009:BNF

[FCB^{+}09] Chiara Francavilla, Paola Cattaneo, Vladimir Berezin, Elisabeth Bock, Diletta Ami, Ario de Marco, Gerhard Christofori, and Ugo Cavallaro. The binding of NCAM to FGFR1 induces a specific cellular response mediated by receptor trafficking. Journal of Cell Biology, 187(7):1101–??, December 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/187/7/1101. See correc-

tion [FCB^{+}19].

Francavilla:2019:CBN

[FCB^{+}19] Chiara Francavilla, Paola Cattaneo, Vladimir Berezin, Elisabeth Bock, Diletta Ami, Ario de Marco, Gerhard Christofori, and Ugo Cavallaro. Correction: The binding of NCAM to FGFR1 induces a specific cellular response mediated by receptor trafficking. Journal of Cell Biology, 218(4):1422–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/4/

1422. See [FCB^{+}09].

Fu:2019:MRF

[FCLoS19] Yu-Hsien Hwang Fu, Sowmya Chandrasekar, Jae Ho Lee, and Shu ou Shan. A molecular recognition feature mediates ribosome-induced SRP-receptor assembly during protein targeting. Journal of Cell Biology, 218(10):3307–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/
10/3307.

Felce:2018:NKS

[FD18] James H. Felce and Michael L. Dustin. Natural killers shed attachments to kill again. Journal of Cell Biology, 217(9): 2983–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/9/2983.

Filipek:2017:LRN

[FdAV^{+}17] Przemyslaw A. Filipek, Mariana E. G. de Araujo, Georg F. Vogel, Cedric H. De Smet, Daniela Eberharter, Manuele Rebsamen, Elena L. Rudashevskaya, Leopold Kremser, Teodor Yordanov, Philipp Tschaikner, Barbara G. Fürnrohr, Stefan Lechner, Theresia Dunzendorfer-Matt, Klaus Scheffzek, Keiryn L. Bennett, Giulio Superti-Furga, Herbert H. Lindner, Taras Stasyk, and Lukas A. Huber. LAMTOR/ ragulator is a negative regulator of Arl8b- and BORC-dependent late endosomal positioning. Journal of Cell Biology, 216(12):4199–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/12/4199.

Fine:2016:GHN

[FDR^{+}16] Noah Fine, Ioannis D. Dimitriou, Jacob Rullo, María José Sandí, Björn Petri, Jack Haitsma, Hisham Ibrahim, Jose La Rose, Michael Glogauer, Paul Kubes, Myron Cybulsky, and Robert Rottapel. GEF-h1 is necessary for neutrophil shear stress–induced migration during inflammation. Journal of Cell Biology, 215(1):107–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/1/107.

Fehrenbacher:2017:GPC

[FdSR^{+}17] Nicole Fehrenbacher, Israel Tojal da Silva, Craig Ramirez, Yong Zhou, Kwang-Jin Cho, Shafi Kuchay, Jie Shi, Susan Thomas, Michele Pagano, John F. Hancock, Dafna Bar-Sagi, and Mark R. Philips. The G protein–coupled receptor GPR31 promotes membrane association of KRAS. Journal of Cell Biology, 216(8):2329–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/8/2329.

Fennell:2015:TCI

[FFÁTC15] Alex Fennell, Alfonso Fernández-Álvarez, Kazunori Tomita, and Julia Promisel Cooper. Telomeres and centromeres have interchangeable roles in promoting meiotic spindle formation. Journal of Cell Biology, 208(4):415–??, February 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/208/4/415.

Fassier:2018:MAN

[FFG^{+}18] Coralie Fassier, Amélie Fréal, Laïla Gasmi, Christian Delphin, Daniel Ten Martin, Stéphanie De Gois, Monica Tambalo, Christophe Bosc, Philippe Mailly, Céline Revenu, Leticia Peris, Susanne Bolte, Sylvie Schneider-Maunoury, Corinne Houart, Fatiha Nothias, Jean-Christophe Larcher, Annie Andrieux, and Jamilé Hazan. Motor axon navigation relies on Fidgetin-like 1–driven microtubule plus end dynamics. Journal of Cell Biology, 217(5):1719–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/5/1719.

Faraldo:2015:RRN

[FG15] Marisa M. Faraldo and Marina A. Glukhova. Regulating the regulator: Numb acts upstream of p53 to control mammary stem and progenitor cell. Journal of Cell Biology, 211(4): 737–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/4/737.

Franco:2016:BFB

[FG16] Claudio A. Franco and Holger Gerhardt. Blood flow boosts BMP signaling to keep vessels in shape. Journal of Cell Biology, 214(7):793–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/7/793.

Fadero:2018:LMT

[FGR^{+}18] Tanner C. Fadero, Therese M. Gerbich, Kishan Rana, Aussie Suzuki, Matthew DiSalvo, Kristina N. Schaefer, Jennifer K. Heppert, Thomas C. Boothby, Bob Goldstein, Mark Peifer, Nancy L. Allbritton, Amy S. Gladfelter, Amy S. Maddox, and Paul S. Maddox. LITE microscopy: Tilted light-sheet excitation of model organisms offers high resolution and low photobleaching. Journal of Cell Biology, 217(5):1869–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/5/
1869.

Frazier:2017:WRT

[FJ17] Meredith N. Frazier and Lauren P. Jackson. Watching real-time endocytosis in living cells. Journal of Cell Biology, 216(1):9–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/1/9.

Forth:2017:MMN

[FK17] Scott Forth and Tarun M. Kapoor. The mechanics of microtubule networks in cell division. Journal of Cell Biology, 216 (6):1525–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/6/1525.

Fourriere:2019:RMR

[FKG^{+}19] Lou Fourriere, Amal Kasri, Nelly Gareil, Sabine Bardin, Hugo Bousquet, David Pereira, Franck Perez, Bruno Goud, Gaelle Boncompain, and Stéphanie Miserey-Lenkei. RAB6 and microtubules restrict protein secretion to focal adhesions. Journal of Cell Biology, 218(7):2215–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/7/2215.

Frank:2018:CPR

[FKL^{+}18a] Scott R. Frank, Clemens P. Köllmann, Phi Luong, Giorgio G. Galli, Lihua Zou, André Bernards, Gad Getz, Raffaele A. Calogero, Morten Frödin, and Steen H. Hansen. Correction: p190 RhoGAP promotes contact inhibition in epithelial cells by repressing YAP activity. Journal of Cell Biology, 217(9): 3313–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/9/3313. See [FKL^{+}18b].

Frank:2018:PRP

[FKL^{+}18b] Scott R. Frank, Clemens P. Köllmann, Phi Luong, Giorgio G. Galli, Lihua Zou, André Bernards, Gad Getz, Raffaele A. Calogero, Morten Frödin, and Steen H. Hansen. p190 RhoGAP promotes contact inhibition in epithelial cells by repressing YAP activity. Journal of Cell Biology, 217(9):3183–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/9/

3183. See correction [FKL^{+}18a].

Funabashi:2018:IHK

[FKO^{+}18] Teruki Funabashi, Yohei Katoh, Misato Okazaki, Maho Sugawa, and Kazuhisa Nakayama. Interaction of heterotrimeric kinesin–II with IFT–B–connecting tetramer is crucial for ciliogenesis. Journal of Cell Biology, 217(8):2867–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/8/
2867.

Feinberg:2017:NAI

[FKW^{+}17] Konstantin Feinberg, Adelaida Kolaj, Chen Wu, Natalie Grinshtein, Jonathan R. Krieger, Michael F. Moran, Lee L. Rubin, Freda D. Miller, and David R. Kaplan. A neuroprotective agent that inactivates prodegenerative TrkA and preserves mitochondria. Journal of Cell Biology, 216(11):3655–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/11/3655.

Ferron:2015:GVI

[FLG^{+}15] Mathieu Ferron, Julie Lacombe, Amélie Germain, Franck Oury, and Gérard Karsenty. GGCX and VKORC1 inhibit osteocalcin endocrine functions. Journal of Cell Biology, 208(6):761–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/

208/6/761. See correction [FLG^{+}19].

Forest:2018:ASB

[FLG^{+}18] Elodie Forest, Rémi Logeay, Charles Géminard, Diala Kantar, Florence Frayssinoux, Lisa Heron-Milhavet, and Alexandre Djiane. The apical scaffold big bang binds to spectrins and regulates the growth of Drosophila melanogaster wing discs. Journal of Cell Biology, 217(3):1047–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/3/1047.

Ferron:2019:CGV

[FLG^{+}19] Mathieu Ferron, Julie Lacombe, Amélie Germain, Franck Oury, and Gérard Karsenty. Correction: GGCX and VKORC1 inhibit osteocalcin endocrine functions. Journal of Cell Biology, 218 (6):2071–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/6/2071. See [FLG^{+}15].

Fritz-Laylin:2017:WSE

[FLLM17] Lillian K. Fritz-Laylin, Samuel J. Lord, and R. Dyche Mullins. WASP and SCAR are evolutionarily conserved in actin-filled pseudopod-based motility. Journal of Cell Biology, 216(6): 1673–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/6/1673.

Fang:2010:BTC

[FLN^{+}10] Xiaodong Fang, Jianying Luo, Ryuichi Nishihama, Carsten Wloka, Christopher Dravis, Mirko Travaglia, Masayuki Iwase, Elizabeth A. Vallen, and Erfei Bi. Biphasic targeting and cleavage furrow ingression directed by the tail of a myosin II. Journal of Cell Biology, 191(7):1333–??, December 2010. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/191/7/1333. See correc-

tion [FLN^{+}16].

Fang:2016:CBT

[FLN^{+}16] Xiaodong Fang, Jianying Luo, Ryuichi Nishihama, Carsten Wloka, Christopher Dravis, Mirko Travaglia, Masayuki Iwase, Elizabeth A. Vallen, and Erfei Bi. Correction: Biphasic targeting and cleavage furrow ingression directed by the tail of a myosin II. Journal of Cell Biology, 215(4):591–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/215/4/

591. See [FLN^{+}10].

Fusco:2016:CVP

[FLS^{+}16] Ludovico Fusco, Riwal Lefort, Kevin Smith, Fethallah Benmansour, German Gonzalez, Caterina Barillari, Bernd Rinn, François Fleuret, Pascal Fua, and Olivier Pertz. Computer vision profiling of neurite outgrowth dynamics reveals spatiotemporal modularity of Rho GTPase signaling. Journal of Cell Biology, 212(1):91–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/1/91.

Farrell:2017:NFD

[FML^{+}17] Kristen B. Farrell, Seth McDonald, Andrew K. Lamb, Colette Worcester, Olve B. Peersen, and Santiago M. Di Pietro. Novel function of a dynein light chain in actin assembly during clathrin-mediated endocytosis. Journal of Cell Biology, 216(8): 2565–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/8/2565.

Fonseca:2019:MCA

[FMS^{+}19] Cindy L. Fonseca, Heidi L. H. Malaby, Leslie A. Sepaniac, Whitney Martin, Candice Byers, Anne Czechanski, Dana Messinger, Mary Tang, Ryoma Ohi, Laura G. Reinholdt, and Jason Stumpff. Mitotic chromosome alignment ensures mitotic fidelity by promoting interchromosomal compaction during anaphase. Journal of Cell Biology, 218(4):1148–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/4/
1148.

Formosa:2016:MFR

[FR16] Luke E. Formosa and Michael T. Ryan. Mitochondrial fusion: Reaching the end of mitofusin’s tether. Journal of Cell Biology, 215(5):597–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/5/597.

Fiuza:2017:PRA

[FRP^{+}17] Maria Fiuza, Christine M. Rostosky, Gabrielle T. Parkinson, Alexei M. Bygrave, Nagaraj Halemani, Marcio Baptista, Ira Milosevic, and Jonathan G. Hanley. PICK1 regulates AMPA receptor endocytosis via direct interactions with AP2 α-appendage and dynamin. Journal of Cell Biology, 216(10): 3323–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/10/3323.

Fiore:2015:CCI

[FSB^{+}15] Vincent F. Fiore, Patrick W. Strane, Anton V. Bryksin, Eric S. White, James S. Hagood, and Thomas H. Barker. Conformational coupling of integrin and Thy-1 regulates Fyn priming and fibroblast mechanotransduction. Journal of Cell Biology, 211 (1):173–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/1/173.

Feiler:2015:TIT

[FSF^{+}15] Marisa S. Feiler, Benjamin Strobel, Axel Freischmidt, Anika M. Helferich, Julia Kappel, Bryson M. Brewer, Deyu Li, Dietmar R. Thal, Paul Walther, Albert C. Ludolph, Karin M. Danzer, and Jochen H. Weishaupt. TDP-43 is intercellularly transmitted across axon terminals. Journal of Cell Biology, 211 (4):897–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/4/897.

Fulgenzi:2015:BMH

[FTAB^{+}15] Gianluca Fulgenzi, Francesco Tomassoni-Ardori, Lucia Babini, Jodi Becker, Colleen Barrick, Sandrine Puverel, and Lino Tessarollo. BDNF modulates heart contraction force and long-term homeostasis through truncated TrkB.T1 receptor activation. Journal of Cell Biology, 210(6):1003–??, September 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/210/6/1003.

Fink:2017:EMC

[FTDC17] Sarah Fink, Kira Turnbull, Arshad Desai, and Christopher S. Campbell. An engineered minimal chromosomal passenger complex reveals a role for INCENP/Sli15 spindle association in chromosome biorientation. Journal of Cell Biology, 216 (4):911–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/4/911.

Feng:2019:PMM

[FTS^{+}19] Chengye Feng, Pankajam Thyagarajan, Matthew Shorey, Dylan Y. Seebold, Alexis T. Weiner, Richard M. Albertson, Kavitha S. Rao, Alvaro Sagasti, Daniel J. Goetschius, and Melissa M. Rolls. Patronin-mediated minus end growth is required for dendritic microtubule polarity. Journal of Cell Biology, 218(7):2309–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/7/2309.

Fuchs:2015:CBM

[Fuc15] Elaine Fuchs. Cell biology: More than skin deep. Journal of Cell Biology, 209(5):629–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/209/5/629.

Fendt:2017:NEG

[FV17] Sarah-Maria Fendt and Patrik Verstreken. Neurons eat glutamate to stay alive. Journal of Cell Biology, 216(4):863–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/4/863.

Fourel:2016:IMS

[FVF^{+}16] Laure Fourel, Anne Valat, Eva Faurobert, Raphael Guillot, Ingrid Bourrin-Reynard, Kefeng Ren, Laurence Lafanechère, Emmanuelle Planus, Catherine Picart, and Corinne Albiges-Rizo. β3 integrin–mediated spreading induced by matrix-bound BMP-2 controls Smad signaling in a stiffness-independent manner. Journal of Cell Biology, 212(6):693–??, March 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/212/6/693.

Farese:2016:LDG

[FW16] Robert V. Farese and Tobias C. Walther. Lipid droplets go nuclear. Journal of Cell Biology, 212(1):7–??, January 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/212/1/7.

Ferguson:2016:DDC

[FWH^{+}16] Joshua P. Ferguson, Nathan M. Willy, Spencer P. Heidotting, Scott D. Huber, Matthew J. Webber, and Comert Kural. Deciphering dynamics of clathrin-mediated endocytosis in a living organism. Journal of Cell Biology, 214(3):347–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/214/3/347.

Fan:2017:CDP

[FWL^{+}17] Junwan Fan, Yaqing Wang, Liang Liu, Hongsheng Zhang, Feng Zhang, Lei Shi, Mei Yu, Fei Gao, and Zhiheng Xu. cTAGE5 deletion in pancreatic β cells impairs proinsulin trafficking and insulin biogenesis in mice. Journal of Cell Biology, 216(12): 4153–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/12/4153.

Fu:2019:SOC

[FZD^{+}19] Gang Fu, Lei Zhao, Erin Dymek, Yuqing Hou, Kangkang Song, Nhan Phan, Zhiguo Shang, Elizabeth F. Smith, George B. Witman, and Daniela Nicastro. Structural organization of the C1a-e-c supercomplex within the ciliary central apparatus. Journal of Cell Biology, 218(12):4236–4251, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/218/12/
4236/132536/Structural-organization-of-the-C1a-e-c.

Gardel:2015:MBM

[Gar15a] Margaret L. Gardel. Moving beyond molecular mechanisms. Journal of Cell Biology, 208(2):143–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/2/143.

Garrett:2015:CBM

[Gar15b] Wendy S. Garrett. From cell biology to the microbiome: an intentional infinite loop. Journal of Cell Biology, 210(1):7–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
210/1/7.

Grimsey:2015:UPA

[GAS^{+}15] Neil J. Grimsey, Berenice Aguilar, Thomas H. Smith, Phillip Le, Amanda L. Soohoo, Manojkumar A. Puthenveedu, Victor Nizet, and JoAnn Trejo. Ubiquitin plays an atypical role in GPCR-induced p38 MAP kinase activation on endosomes. Journal of Cell Biology, 210(7):1117–??, September 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/210/7/1117.

Graham:2018:ECR

[GAS^{+}18] David M. Graham, Tomas Andersen, Lisa Sharek, Gunes Uzer, Katheryn Rothenberg, Brenton D. Hoffman, Janet Rubin, Martial Balland, James E. Bear, and Keith Burridge. Enucleated cells reveal differential roles of the nucleus in cell migration, polarity, and mechanotransduction. Journal of Cell Biology, 217 (3):895–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/3/895.

Gemble:2018:FFL

[GB18] Simon Gemble and Renata Basto. Fast and furious … or not, Plk4 dictates the pace. Journal of Cell Biology, 217(4):1169–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/4/1169.

Goel:2019:HSA

[GBB^{+}19] Pragya Goel, Dominique Dufour Bergeron, Mathias A. Böhme, Luke Nunnelly, Martin Lehmann, Christopher Buser, Alexander M. Walter, Stephan J. Sigrist, and Dion Dickman. Homeostatic scaling of active zone scaffolds maintains global synaptic strength. Journal of Cell Biology, 218(5):1706–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/5/
1706.

Gayrard:2018:SCD

[GBD^{+}18] Charlène Gayrard, Clément Bernaudin, Théophile Déjardin, Cynthia Seiler, and Nicolas Borghi. Src- and confinementdependent FAK activation causes E-cadherin relaxation and βcatenin activity. Journal of Cell Biology, 217(3):1063–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/3/
1063.

Gao:2017:PYP

[GBK^{+}17] Qiang Gao, Derk D. Binns, Lisa N. Kinch, Nick V. Grishin, Natalie Ortiz, Xiao Chen, and Joel M. Goodman. Pet10p is a yeast perilipin that stabilizes lipid droplets and promotes their assembly. Journal of Cell Biology, 216(10):3199–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/10/3199.

Grippa:2015:SCF

[GBM^{+}15] Alexandra Grippa, Laura Buxó, Gabriel Mora, Charlotta Funaya, Fatima-Zahra Idrissi, Francesco Mancuso, Raul Gomez, Júlia Muntanyà, Eduard Sabidó, and Pedro Carvalho. The seipin complex Fld1/Ldb16 stabilizes ER–lipid droplet contact sites. Journal of Cell Biology, 211(4):829–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/211/4/829.

Gao:2015:GCR

[GBRH15] Xia Gao, Aman S. Bali, Scott H. Randell, and Brigid L. M. Hogan. GRHL2 coordinates regeneration of a polarized mucociliary epithelium from basal stem cells. Journal of Cell Biology, 211(3):669–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/3/669.

Gong:2017:HDK

[GCA^{+}17] Fade Gong, Thomas Clouaire, Marion Aguirrebengoa, Gaëlle Legube, and Kyle M. Miller. Histone demethylase KDM5A regulates the ZMYND8–NuRD chromatin remodeler to promote DNA repair. Journal of Cell Biology, 216(7):1959–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/7/1959.

Guo:2018:KRM

[GCC^{+}18] Ling Guo, Ting Cai, Keng Chen, Rong Wang, Jiaxin Wang, Chunhong Cui, Jifan Yuan, Kuo Zhang, Zhongzhen Liu, Yi Deng, Guozhi Xiao, and Chuanyue Wu. Kindlin-2 regulates mesenchymal stem cell differentiation through control of YAP1/ TAZ. Journal of Cell Biology, 217(4):1431–??, April 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/4/1431.

Gomez-Cavazos:2015:NGN

[GCH15] J. Sebastian Gomez-Cavazos and Martin W. Hetzer. The nucleoporin gp210/Nup210 controls muscle differentiation by regulating nuclear envelope/ER homeostasis. Journal of Cell Biology, 208(6):671–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/6/671.

Gerber:2015:RCH

[GCJ^{+}15] Pehuén Pereyra Gerber, Mercedes Cabrini, Carolina Jancic, Luciana Paoletti, Claudia Banchio, Catalina von Bilderling, Lorena Sigaut, Lía I. Pietrasanta, Gabriel Duette, Eric O. Freed, Genevieve de Saint Basile, Catarina Ferreira Moita, Luis Ferreira Moita, Sebastian Amigorena, Philippe Benaroch, Jorge Geffner, and Matías Ostrowski. Rab27a controls HIV-1 assembly by regulating plasma membrane levels of phosphatidylinositol 4,5-bisphosphate. Journal of Cell Biology, 209 (3):435–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/3/435.

Gluszek:2015:MCP

[GCL^{+}15] A. Agata Głuszek, C. Fiona Cullen, Wenjing Li, Rachel A. Battaglia, Sarah J. Radford, Mariana F. Costa, Kim S. McKim, Gohta Goshima, and Hiroyuki Ohkura. The microtubule catastrophe promoter Sentin delays stable kinetochore–microtubule attachment in oocytes. Journal of Cell Biology, 211(6):1113– ??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/6/1113.

Gomis-Coloma:2018:CIH

[GCVAGS^{+}18] Clara Gomis-Coloma, Sergio Velasco-Aviles, Jose A. Gomez-Sanchez, Angeles Casillas-Bajo, Johannes Backs, and Hugo Cabedo. Class IIa histone deacetylases link cAMP signaling to the myelin transcriptional program of Schwann cells. Journal of Cell Biology, 217(4):1249–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/4/1249.

Gracheva:2016:ZMR

[GCW^{+}16] Ekaterina Gracheva, Shalaka Chitale, Thomas Wilhelm, Alexander Rapp, Jonathan Byrne, Jens Stadler, Rebeca Medina, M. Cristina Cardoso, and Holger Richly. ZRF1 mediates remodeling of E3 ligases at DNA lesion sites during nucleotide excision repair. Journal of Cell Biology, 213(2):185–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/213/2/
185.

Galino:2019:RCS

[GCZ^{+}19] Jorge Galino, Ilaria Cervellini, Ning Zhu, Nina Stöberl, Meike Hütte, Florence R. Fricker, Garrett Lee, Lucy McDermott, Giovanna Lalli, and David L. H. Bennett. RalGTPases contribute to Schwann cell repair after nerve injury via regulation of process formation. Journal of Cell Biology, 218(7):2370–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/7/
2370.

Green:2016:CBA

[GD16] Dustin Green and Xinzhong Dong. The cell biology of acute itch. Journal of Cell Biology, 213(2):155–??, April 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/213/2/155.

Giampietro:2015:ABP

[GDB^{+}15] Costanza Giampietro, Andrea Disanza, Luca Bravi, Miriam Barrios-Rodiles, Monica Corada, Emanuela Frittoli, Cecilia Savorani, Maria Grazia Lampugnani, Barbara Boggetti, Carien Niessen, Jeff L. Wrana, Giorgio Scita, and Elisabetta Dejana. The actin-binding protein EPS8 binds VE-cadherin and modulates YAP localization and signaling. Journal of Cell Biology, 211(6):1177–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/6/1177.

Gadadhar:2017:TGC

[GDB^{+}17] Sudarshan Gadadhar, Hala Dadi, Satish Bodakuntla, Anne Schnitzler, Ivan Bièche, Filippo Rusconi, and Carsten Janke. Tubulin glycylation controls primary cilia length. Journal of Cell Biology, 216(9):2701–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/9/2701.

Gabel:2015:AAD

[GDD^{+}15] Marion Gabel, Franck Delavoie, Valérie Demais, Cathy Royer, Yannick Bailly, Nicolas Vitale, Marie-France Bader, and Sylvette Chasserot-Golaz. Annexin A2-dependent actin bundling promotes secretory granule docking to the plasma membrane and exocytosis. Journal of Cell Biology, 210(5):785–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/210/5/
785.

Gang:2015:PMA

[GDL^{+}15] Hongying Gang, Rimpy Dhingra, Junjun Lin, Yan Hai, Yaron Aviv, Victoria Margulets, Mohammad Hamedani, Thatchawan Thanasupawat, Etienne Leygue, Thomas Klonisch, James R. Davie, and Lorrie A. Kirshenbaum. PDK2-mediated alternative splicing switches Bnip3 from cell death to cell survival. Journal of Cell Biology, 210(7):1101–??, September 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/210/7/1101.

Goncalves:2019:DRD

[GDV19] João Carlos Gonçalves, Tiago J. Dantas, and Richard B. Vallee. Distinct roles for dynein light intermediate chains in neurogenesis, migration, and terminal somal translocation. Journal of Cell Biology, 218(3):808–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/3/808.

Gekara:2017:DDI

[Gek17] Nelson O. Gekara. DNA damage-induced immune response: Micronuclei provide key platform. Journal of Cell Biology, 216(10):2999–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/10/2999.

Genot:2017:ACP

[Gen17] Elisabeth Genot. ARF1 at the crossroads of podosome construction and function. Journal of Cell Biology, 216(1):13–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/1/13.

Gerbi:2015:BED

[Ger15] Susan A. Gerbi. Beginning at the end: DNA replication within the telomere. Journal of Cell Biology, 210(2):177–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/210/2/177.

Gerton:2018:TFP

[Ger18] Jennifer L. Gerton. A transcription factor primes the condensin pump. Journal of Cell Biology, 217(7):2233–??, July 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/7/2233.

Gowrishankar:2016:LRC

[GF16] Swetha Gowrishankar and Shawn M. Ferguson. Lysosomes relax in the cellular suburbs. Journal of Cell Biology, 212(6):617–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
212/6/617.

Garcia:2016:AMO

[GFH^{+}16] Galo Garcia, Gregory C. Finnigan, Lydia R. Heasley, Sarah M. Sterling, Adeeti Aggarwal, Chad G. Pearson, Eva Nogales, Michael A. McMurray, and Jeremy Thorner. Assembly, molecular organization, and membrane-binding properties of development-specific septins. Journal of Cell Biology, 212(5): 515–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/5/515.

Gaik:2015:SBA

[GFvA^{+}15] Monika Gaik, Dirk Flemming, Alexander von Appen, Panagiotis Kastritis, Norbert Mücke, Jessica Fischer, Philipp Stelter, Alessandro Ori, Khanh Huy Bui, Jochen Baßler, Elisar Barbar, Martin Beck, and Ed Hurt. Structural basis for assembly and function of the Nup82 complex in the nuclear pore scaffold. Journal of Cell Biology, 208(3):283–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/3/283.

Grikscheit:2015:JAA

[GFWG15] Katharina Grikscheit, Tanja Frank, Ying Wang, and Robert Grosse. Junctional actin assembly is mediated by Formin-like 2 downstream of Rac1. Journal of Cell Biology, 209(3):367–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
209/3/367.

Guerrero:2016:HHK

[GG16] Ana Guerrero and Jesús Gil. HMGB2 holds the key to the senescence-associated secretory phenotype. Journal of Cell Biology, 215(3):297–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/3/297.

Graziano:2017:MRT

[GGA^{+}17] Brian R. Graziano, Delquin Gong, Karen E. Anderson, Anne Pipathsouk, Anna R. Goldberg, and Orion D. Weiner. A module for Rac temporal signal integration revealed with optogenetics. Journal of Cell Biology, 216(8):2515–??, August 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/8/2515.

Genova:2017:TIE

[GGC^{+}17] Tullio Genova, Guillaume P. Grolez, Chiara Camillo, Michela Bernardini, Alexandre Bokhobza, Elodie Richard, Marco Scianna, Loic Lemonnier, Donatella Valdembri, Luca Munaron, Mark R. Philips, Virginie Mattot, Guido Serini, Natalia Prevarskaya, Dimitra Gkika, and Alessandra Fiorio Pla. TRPM8 inhibits endothelial cell migration via a non-channel function by trapping the small GTPase Rap1. Journal of Cell Biology, 216 (7):2107–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/7/2107.

Gunawan:2019:FAE

[GGF^{+}19] Felix Gunawan, Alessandra Gentile, Ryuichi Fukuda, Ayele Taddese Tsedeke, Vanesa Jiménez-Amilburu, Radhan Ramadass, Atsuo Iida, Atsuko Sehara-Fujisawa, and Didier Y. R. Stainier. Focal adhesions are essential to drive zebrafish heart valve morphogenesis. Journal of Cell Biology, 218(3):1039–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/3/
1039.

Gong:2019:GCE

[GGL^{+}19] Jingyi Gong, Thomas N. Gaitanos, Olivia Luu, Yunyun Huang, Louise Gaitanos, Jana Lindner, Rudolf Winklbauer, and Rüdiger Klein. Gulp1 controls Eph/ ephrin trogocytosis and is important for cell rearrangements during development.

Journal of Cell Biology, 218(10):3455–??, October 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/10/3455.

Gonzalez-Gaitan:2015:WCB

[GGR15] Marcos Gonzalez-Gaitan and Aurélien Roux. When cell biology meets theory. Journal of Cell Biology, 210(7):1041–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/210/7/
1041.

Gerson-Gurwitz:2019:ARF

[GGWL^{+}19] Adina Gerson-Gurwitz, Carolyn A. Worby, Kian-Yong Lee, Renat Khaliullin, Jeff Bouffard, Dhanya Cheerambathur, Karen Oegema, Erin J. Cram, Jack E. Dixon, and Arshad Desai. Ancestral roles of the Fam20C family of secreted protein kinases revealed in C. elegans. Journal of Cell Biology, 218(11): 3795–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/11/3795.

Ganguly:2017:HCA

[GHD^{+}17] Archan Ganguly, Xuemei Han, Utpal Das, Lina Wang, Jonathan Loi, Jichao Sun, Daniel Gitler, Ghislaine Caillol, Christophe Leterrier, John R. Yates, and Subhojit Roy. Hsc70 chaperone activity is required for the cytosolic slow axonal transport of synapsin. Journal of Cell Biology, 216(7):2059–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/7/2059.

Gilmore-Hall:2019:CPM

[GHKW^{+}19] Stephen Gilmore-Hall, Jennifer Kuo, Jacqueline M. Ward, Rabaab Zahra, Richard S. Morrison, Guy Perkins, and Albert R. La Spada. CCP1 promotes mitochondrial fusion and motility to prevent Purkinje cell neuron loss in pcd mice. Journal of Cell Biology, 218(1):206–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/1/206.

Grenfell:2016:CMN

[GHS16a] Andrew W. Grenfell, Rebecca Heald, and Magdalena Strzelecka. Correction: Mitotic noncoding RNA processing promotes kinetochore and spindle assembly in Xenopus. Journal of Cell Biology, 214(6):783–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/6/783.

Grenfell:2016:MNR

[GHS16b] Andrew W. Grenfell, Rebecca Heald, and Magdalena Strzelecka. Mitotic noncoding RNA processing promotes kinetochore and spindle assembly in Xenopus. Journal of Cell Biology, 214(2):133–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/214/2/133.

Gay:2019:AHF

[GI19] Denise Gay and Mayumi Ito. Adult hair follicles keep oncogenic growth in check. Journal of Cell Biology, 218(10):3163–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/10/3163.

Galletta:2016:ARC

[GJFR16] Brian J. Galletta, Katherine C. Jacobs, Carey J. Fagerstrom, and Nasser M. Rusan. Asterless is required for centriole length control and sperm development. Journal of Cell Biology, 213 (4):435–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/4/435.

Gu:2017:PVI

[GJW^{+}17] Yuanzheng Gu, Peter Jukkola, Qian Wang, Thomas Esparza, Yi Zhao, David Brody, and Chen Gu. Polarity of varicosity initiation in central neuron mechanosensation. Journal of Cell Biology, 216(7):2179–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/7/2179.

Graessl:2017:ERG

[GKC^{+}17] Melanie Graessl, Johannes Koch, Abram Calderon, Dominic Kamps, Soumya Banerjee, Tomáš Mazel, Nina Schulze, Jana Kathrin Jungkurth, Rutuja Patwardhan, Djamschid Solouk, Nico Hampe, Bernd Hoffmann, Leif Dehmelt, and Perihan Nalbant. An excitable Rho GTPase signaling network generates dynamic subcellular contraction patterns. Journal of Cell Biology, 216(12):4271–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/12/4271.

Gutmann:2018:VLI

[GKG^{+}18] Theresia Gutmann, Kelly H. Kim, Michal Grzybek, Thomas Walz, and Ünal Coskun. Visualization of ligand-induced transmembrane signaling in the full-length human insulin receptor. Journal of Cell Biology, 217(5):1643–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/5/1643.

Gong:2016:EMC

[GKGK16] Jingyi Gong, Roman Körner, Louise Gaitanos, and Rüdiger Klein. Exosomes mediate cell contact–independent ephrin– Eph signaling during axon guidance. Journal of Cell Biology, 214(1):35–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/214/1/35.

Gaitanos:2016:CTR

[GKK16a] Thomas N. Gaitanos, Jorg Koerner, and Ruediger Klein. Correction: Tiam–Rac signaling mediates trans-endocytosis of ephrin receptor EphB2 and is important for cell repulsion. Journal of Cell Biology, 215(3):431–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/215/3/431.

Gaitanos:2016:TRS

[GKK16b] Thomas N. Gaitanos, Jorg Koerner, and Ruediger Klein. Tiam– Rac signaling mediates trans-endocytosis of ephrin receptor EphB2 and is important for cell repulsion. Journal of Cell Biology, 214(6):735–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/6/735.

Gingras:2019:RBL

[GLC^{+}19] Alexandre R. Gingras, Frederic Lagarrigue, Monica N. Cuevas, Andrew J. Valadez, Marcus Zorovich, Wilma McLaughlin, Miguel Alejandro Lopez-Ramirez, Nicolas Seban, Klaus Ley, William B. Kiosses, and Mark H. Ginsberg. Rap1 binding and a lipid-dependent helix in talin F1 domain promote integrin activation in tandem. Journal of Cell Biology, 218(6):1799–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/6/1799.

Glick:2017:NIP

[Gli17] Benjamin S. Glick. New insights into protein secretion: TANGO1 runs rings around the COPII coat. Journal of Cell Biology, 216(4):859–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/4/859.

Georgiadou:2017:ANR

[GLJ^{+}17] Maria Georgiadou, Johanna Lilja, Guillaume Jacquemet, Camilo Guzmán, Maria Rafaeva, Charlotte Alibert, Yan Yan, Pranshu Sahgal, Martina Lerche, Jean-Baptiste Manneville, Tomi P. Mäkelä, and Johanna Ivaska. AMPK negatively regulates tensin-dependent integrin activity. Journal of Cell Biology, 216(4):1107–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/4/1107.

Gao:2018:ART

[GLL^{+}18a] Fengyi Gao, Guoping Li, Chao Liu, Hui Gao, Hao Wang, Weixiao Liu, Min Chen, Yongliang Shang, Lina Wang, Jian Shi, Wenlong Xia, Jianwei Jiao, Fei Gao, Jian Li, Liang Chen, and Wei Li. Autophagy regulates testosterone synthesis by facilitating cholesterol uptake in Leydig cells. Journal of Cell Biology, 217(6):2103–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/6/2103.

Genna:2018:PFD

[GLL^{+}18b] Alessandro Genna, Stefanie Lapetina, Nikola Lukic, Shams Twafra, Tomer Meirson, Ved P. Sharma, John S. Condeelis, and Hava Gil-Henn. Pyk2 and FAK differentially regulate invadopodia formation and function in breast cancer cells. Journal of Cell Biology, 217(1):375–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/1/375.

Gkika:2015:TCA

[GLS^{+}15] Dimitra Gkika, Loic Lemonnier, George Shapovalov, Dmitri Gordienko, Céline Poux, Michela Bernardini, Alexandre Bokhobza, Gabriel Bidaux, Cindy Degerny, Kathye Verreman, Basma Guarmit, Mohamed Benahmed, Yvan de Launoit, Rene J. M. Bindels, Alessandra Fiorio Pla, and Natalia Prevarskaya. TRP channel–associated factors are a novel protein family that regulates TRPM8 trafficking and activity. Journal of Cell Biology, 208(1):89–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/1/89.

Grys:2017:MLC

[GLS^{+}17] Ben T. Grys, Dara S. Lo, Nil Sahin, Oren Z. Kraus, Quaid Morris, Charles Boone, and Brenda J. Andrews. Machine learning and computer vision approaches for phenotypic profiling. Journal of Cell Biology, 216(1):65–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/1/65.

Gomez-Lamarca:2015:RDPb

[GLSS^{+}15a] Maria Gomez-Lamarca, Laura Amy Snowdon, Ekatarina Seib, Thomas Klein, and Sarah Bray. Rme-8 depletion perturbs Notch recycling and predisposes to pathogenic signaling. Journal of Cell Biology, 210(3):517–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/210/3/517.

Gomez-Lamarca:2015:RDPa

[GLSS^{+}15b] Maria J. Gomez-Lamarca, Laura A. Snowdon, Ekatarina Seib, Thomas Klein, and Sarah J. Bray. Rme-8 depletion perturbs Notch recycling and predisposes to pathogenic signaling. Journal of Cell Biology, 210(2):303–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/210/2/303.

Gangoda:2016:CEE

[GM16] Lahiru Gangoda and Suresh Mathivanan. Cortactin enhances exosome secretion without altering cargo. Journal of Cell Biology, 214(2):129–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/2/129.

Gerganova:2018:DVC

[GM18] Veneta Gerganova and Sophie G. Martin. Dynamic visits of cortical structures probe for cell size. Journal of Cell Biology, 217 (5):1559–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/5/1559.

Gucciardo:2016:PIR

[GML16] Erika Gucciardo, Mohammad Mobashir, and Kaisa Lehti. Proactive for invasion: Reuse of matrix metalloproteinase for structural memory. Journal of Cell Biology, 213(1):11–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/213/1/
11.

Gihana:2018:PCD

[GMTL18] Gabriel M. Gihana, Tiffany R. Musser, Oscar Thompson, and Soni Lacefield. Prolonged cyclin-dependent kinase inhibition results in septin perturbations during return to growth and mitosis. Journal of Cell Biology, 217(7):2429–??, July 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/7/2429.

Goodwin:2018:MCC

[GN18] Katharine Goodwin and Celeste M. Nelson. Myoepithelial crowd control of cancer cells. Journal of Cell Biology, 217(10): 3319–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/10/3319.

Gomez-Navarro:2016:PSG

[GNM16] Natalia Gomez-Navarro and Elizabeth Miller. Protein sorting at the ER–Golgi interface. Journal of Cell Biology, 215(6): 769–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/215/6/769.

Gomez:2017:STR

[Góm17] María Gómez. A stitch in time: Replicate early and escape dosage compensation to express more. Journal of Cell Biology, 216(7):1869–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/7/1869.

Gilberto:2017:DUS

[GP17] Samuel Gilberto and Matthias Peter. Dynamic ubiquitin signaling in cell cycle regulation. Journal of Cell Biology, 216(8): 2259–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/8/2259.

Gagnoux-Palacios:2018:CPA

[GPAA^{+}18] Laurent Gagnoux-Palacios, Hala Awina, Stéphane Audebert, Aurélie Rossin, Magali Mondin, Franck Borgese, Carlota Planas-Botey, Amel Mettouchi, Jean-Paul Borg, and Anne-Odile Hueber. Cell polarity and adherens junction formation inhibit epithelial Fas cell death receptor signaling. Journal of Cell Biology, 217(11):3839–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/11/3839.

Goulden:2019:HAB

[GPD^{+}19] Brady D. Goulden, Jonathan Pacheco, Allyson Dull, James P. Zewe, Alexander Deiters, and Gerald R. V. Hammond. A highavidity biosensor reveals plasma membrane PI(3,4)P_{2} is predominantly a class I PI3K signaling product. Journal of Cell Biology, 218(3):1066–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/3/1066.

Gamblin:2018:OFF

[GPPJ^{+}18] Clémence L. Gamblin, Frédérique Parent-Prévost, Kévin Jacquet, Cornélia Biehler, Alexandra Jetté, and Patrick Laprise. Oligomerization of the FERM–FA protein Yurt controls epithelial cell polarity. Journal of Cell Biology, 217(11): 3853–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/11/3853.

Gama:2017:MMD

[GPS^{+}17] José B. Gama, Cláudia Pereira, Patrícia A. Simões, Ricardo Celestino, Rita M. Reis, Daniel J. Barbosa, Helena R. Pires, Cátia Carvalho, João Amorim, Ana X. Carvalho, Dhanya K. Cheerambathur, and Reto Gassmann. Molecular mechanism of dynein recruitment to kinetochores by the Rod–Zw10–Zwilch complex and Spindly. Journal of Cell Biology, 216(4):943–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/4/943.

Graef:2016:DMD

[Gra16] Martin Graef. A dividing matter: Drp1/Dnm1-independent mitophagy. Journal of Cell Biology, 215(5):599–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/215/5/
599.

Gonzalez-Rodriguez:2019:XER

[GRB19] Yanira Gonzalez-Rodriguez and Samuel F. Bunting. XLF extends its range from DNA repair to replication. Journal of Cell Biology, 218(7):2075–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/7/2075.

Gao:2018:NVA

[GRU18] Jieqiong Gao, Fulvio Reggiori, and Christian Ungermann. A novel in vitro assay reveals SNARE topology and the role of Ykt6 in autophagosome fusion with vacuoles. Journal of Cell Biology, 217(10):3670–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/217/10/3670.

Giacomello:2018:IOM

[GS18] Marta Giacomello and Luca Scorrano. The INs and OUTs of mitofusins. Journal of Cell Biology, 217(2):439–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/2/
439.

Guimaraes:2015:PLD

[GSB^{+}15] Sofia C. Guimaraes, Martin Schuster, Ewa Bielska, Gulay Dagdas, Sreedhar Kilaru, Ben R. A. Meadows, Michael Schrader, and Gero Steinberg. Peroxisomes, lipid droplets, and endoplasmic reticulum “hitchhike” on motile early endosomes. Journal of Cell Biology, 211(5):945–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/211/5/945.

Grenfell:2016:VMI

[GSC^{+}16] Andrew W. Grenfell, Magdalena Strzelecka, Marina E. Crowder, Kara J. Helmke, Anne-Lore Schlaitz, and Rebecca Heald. A versatile multivariate image analysis pipeline reveals features of Xenopus extract spindles. Journal of Cell Biology, 213 (1):127–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/1/127.

Gomez-Sanchez:2015:SCA

[GSCIL^{+}15] Jose A. Gomez-Sanchez, Lucy Carty, Marta Iruarrizaga-Lejarreta, Marta Palomo-Irigoyen, Marta Varela-Rey, Megan Griffith, Janina Hantke, Nuria Macias-Camara, Mikel Azkargorta, Igor Aurrekoetxea, Virginia Gutiérrez De Juan, Harold B. J. Jefferies, Patricia Aspichueta, Félix Elortza, Ana M. Aransay, María L. Martínez-Chantar, Frank Baas, José M. Mato, Rhona Mirsky, Ashwin Woodhoo, and Kristján R. Jessen. Schwann cell autophagy, myelinophagy, initiates myelin clearance from injured nerves. Journal of Cell Biology, 210 (1):153–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/1/153.

Ghalei:2015:HCD

[GSD^{+}15] Homa Ghalei, Franz X. Schaub, Joanne R. Doherty, Yoshihiko Noguchi, William R. Roush, John L. Cleveland, M. Elizabeth Stroupe, and Katrin Karbstein. Hrr25/CK1δ-directed release of Ltv1 from pre-40S ribosomes is necessary for ribosome assembly and cell growth. Journal of Cell Biology, 208(6):745–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/208/6/
745.

Garabedian:2018:ICF

[GSKL^{+}18] Mikael V. Garabedian, Tatiana Stanishneva-Konovalova, Chenyu Lou, Thomas J. Rands, Luther W. Pollard, Olga S. Sokolova, and Bruce L. Goode. Integrated control of forminmediated actin assembly by a stationary inhibitor and a mobile activator. Journal of Cell Biology, 217(10):3512–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/
10/3512.

Gopal:2015:TPC

[GSM^{+}15] Sandeep Gopal, Pernille Søgaard, Hinke A. B. Multhaupt, Csilla Pataki, Elena Okina, Xiaojie Xian, Mikael E. Pedersen, Troy Stevens, Oliver Griesbeck, Pyong Woo Park, Roger Pocock, and John R. Couchman. Transmembrane proteoglycans control stretch-activated channels to set cytosolic calcium levels. Journal of Cell Biology, 210(7):1199–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/210/7/
1199.

Gagliardi:2018:MAC

[GSP^{+}18] Paolo Armando Gagliardi, Desiana Somale, Alberto Puliafito, Giulia Chiaverina, Laura di Blasio, Michele Oneto, Paolo Bianchini, Federico Bussolino, and Luca Primo. MRCKα is activated by caspase cleavage to assemble an apical actin ring for epithelial cell extrusion. Journal of Cell Biology, 217(1):231–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/1/231.

Gomez-Sanchez:2018:AEA

[GSRG^{+}18] Rubén Gómez-Sánchez, Jaqueline Rose, Rodrigo Guimarães, Muriel Mari, Daniel Papinski, Ester Rieter, Willie J. Geerts, Ralph Hardenberg, Claudine Kraft, Christian Ungermann, and Fulvio Reggiori. Atg9 establishes Atg2-dependent contact sites between the endoplasmic reticulum and phagophores. Journal of Cell Biology, 217(8):2743–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/8/2743.

Goo:2017:ADT

[GSS^{+}17] Marisa S. Goo, Laura Sancho, Natalia Slepak, Daniela Boassa, Thomas J. Deerinck, Mark H. Ellisman, Brenda L. Bloodgood, and Gentry N. Patrick. Activity-dependent trafficking of lysosomes in dendrites and dendritic spines. Journal of Cell Biology, 216(8):2499–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/8/2499.

Gong:2018:HEA

[GTD^{+}18] Zhenwei Gong, Inmaculada Tasset, Antonio Diaz, Jaime Anguiano, Emir Tas, Lingguang Cui, Regina Kuliawat, Honghai Liu, Bernhard Kühn, Ana Maria Cuervo, and Radhika Muzumdar. Humanin is an endogenous activator of chaperonemediated autophagy. Journal of Cell Biology, 217(2):635– ??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/2/635.

Guetta-Terrier:2015:PWG

[GTMZ^{+}15] Charlotte Guetta-Terrier, Pascale Monzo, Jie Zhu, Hongyan Long, Lakshmi Venkatraman, Yue Zhou, PeiPei Wang, Sing Yian Chew, Alexander Mogilner, Benoit Ladoux, and Nils C. Gauthier. Protrusive waves guide 3D cell migration along nanofibers. Journal of Cell Biology, 211(3):683– ??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/3/683.

Ganguly:2015:DFD

[GTW^{+}15] Archan Ganguly, Yong Tang, Lina Wang, Kelsey Ladt, Jonathan Loi, Bénédicte Dargent, Christophe Leterrier, and Subhojit Roy. A dynamic formin-dependent deep F-actin network in axons. Journal of Cell Biology, 210(3):401–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/210/3/
401.

Grousl:2018:PLD

[GUM^{+}18] Tomas Grousl, Sophia Ungelenk, Stephanie Miller, Chi-Ting Ho, Maria Khokhrina, Matthias P. Mayer, Bernd Bukau, and Axel Mogk. A prion-like domain in Hsp42 drives chaperonefacilitated aggregation of misfolded proteins. Journal of Cell Biology, 217(4):1269–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/4/1269.

Gowrishankar:2017:IJD

[GWF17] Swetha Gowrishankar, Yumei Wu, and Shawn M. Ferguson. Impaired JIP3-dependent axonal lysosome transport promotes amyloid plaque pathology. Journal of Cell Biology, 216(10): 3291–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/10/3291.

Gomez:2019:MAI

[GWL^{+}19] Rachel C. Gomez, Paulina Wawro, Pawel Lis, Dario R. Alessi, and Suzanne R. Pfeffer. Membrane association but not identity is required for LRRK2 activation and phosphorylation of Rab GTPases. Journal of Cell Biology, 218(12):4157– 4170, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/218/12/4157/132547/Membrane-association-
but-not-identity-is-required.

Gan:2019:AAT

[GWZ^{+}19a] Qiwen Gan, Xin Wang, Qian Zhang, Qiuyuan Yin, Youli Jian, Yubing Liu, Nan Xuan, Jinglin Li, Junxiang Zhou, Kai Liu, Yudong Jing, Xiaochen Wang, and Chonglin Yang. The amino acid transporter SLC-36.1 cooperates with PtdIns3P 5-kinase to control phagocytic lysosome reformation. Journal of Cell Biology, 218(8):2619–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/8/2619.

Guo:2019:PSS

[GWZ^{+}19b] Ling Guo, Rong Wang, Kuo Zhang, Jifan Yuan, Jiaxin Wang, Xiaoxia Wang, Jianfei Ma, and Chuanyue Wu. A PINCH-1– Smurf1 signaling axis mediates mechano-regulation of BMPR2 and stem cell differentiation. Journal of Cell Biology, 218(11): 3773–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/11/3773.

Gu:2016:PTP

[GX16] Mingxue Gu and Haoxing Xu. A painful TR(i)P to lysosomes. Journal of Cell Biology, 215(3):309–??, November 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/215/3/309.

Xia:2015:DHC

[gXNG^{+}15] Hong guang Xia, Ayaz Najafov, Jiefei Geng, Lorena Galan-Acosta, Xuemei Han, Yuan Guo, Bing Shan, Yaoyang Zhang, Erik Norberg, Tao Zhang, Lifeng Pan, Junli Liu, Jonathan L. Coloff, Dimitry Ofengeim, Hong Zhu, Kejia Wu, Yu Cai, John R. Yates, Zhengjiang Zhu, Junying Yuan, and Helin Vakifahmetoglu-Norberg. Degradation of HK2 by chaperonemediated autophagy promotes metabolic catastrophe and cell death. Journal of Cell Biology, 210(5):705–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/210/5/705.

Xia:2016:CDH

[gXNG^{+}16] Hong guang Xia, Ayaz Najafov, Jiefei Geng, Lorena Galan-Acosta, Xuemei Han, Yuan Guo, Bing Shan, Yaoyang Zhang, Erik Norberg, Tao Zhang, Lifeng Pan, Junli Liu, Jonathan L. Coloff, Dimitry Ofengeim, Hong Zhu, Kejia Wu, Yu Cai, John R. Yates, Zhengjiang Zhu, Junying Yuan, and Helin Vakifahmetoglu-Norberg. Correction: Degradation of HK2 by chaperone-mediated autophagy promotes metabolic catastrophe and cell death. Journal of Cell Biology, 212(7):881–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
212/7/881.

Gao:2018:VLT

[GY18] Mingming Gao and Hongyuan Yang. VPS13: a lipid transfer protein making contacts at multiple cellular locations. Journal of Cell Biology, 217(10):3322–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/10/3322.

Gorur:2017:CCM

[GYK^{+}17] Amita Gorur, Lin Yuan, Samuel J. Kenny, Satoshi Baba, Ke Xu, and Randy Schekman. COPII-coated membranes function as transport carriers of intracellular procollagen I. Journal of Cell Biology, 216(6):1745–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/6/1745.

Goult:2018:TMS

[GYS18] Benjamin T. Goult, Jie Yan, and Martin A. Schwartz. Talin as a mechanosensitive signaling hub. Journal of Cell Biology, 217(11):3776–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/11/3776.

Haynes:2015:GCB

[HAK^{+}15] Elizabeth M. Haynes, Sreeja B. Asokan, Samantha J. King, Heath E. Johnson, Jason M. Haugh, and James E. Bear. GMFβ controls branched actin content and lamellipodial retraction in fibroblasts. Journal of Cell Biology, 209(6):803–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/209/6/803.

Hall:2015:CFY

[Hal15] Alan Hall. Celebrating the first 60 years of The Journal of Cell Biology. Journal of Cell Biology, 208(1):9–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/208/1/
9.

Hammond:2018:DMI

[Ham18] Gerald R. V. Hammond. DepHining membrane identity. Journal of Cell Biology, 217(1):19–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/1/19.

Hayward:2019:CCC

[HAPC^{+}19] Daniel Hayward, Tatiana Alfonso-Pérez, Michael J. Cundell, Michael Hopkins, James Holder, James Bancroft, Lukas H. Hutter, Bela Novak, Francis A. Barr, and Ulrike Gruneberg. CDK1-CCNB1 creates a spindle checkpoint–permissive state by enabling MPS1 kinetochore localization. Journal of Cell Biology, 218(4):1182–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/4/1182.

Hyenne:2015:RCM

[HAR^{+}15] Vincent Hyenne, Ahmet Apaydin, David Rodriguez, Coralie Spiegelhalter, Sarah Hoff-Yoessle, Maxime Diem, Saurabh Tak, Olivier Lefebvre, Yannick Schwab, Jacky G. Goetz, and Michel Labouesse. RAL-1 controls multivesicular body biogenesis and exosome secretion. Journal of Cell Biology, 211(1):27–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/1/

Hardin:2016:RCC

[Har16] Jeff Hardin. Regulating cell–cell junctions from A to Z. Journal of Cell Biology, 213(2):151–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/213/2/151.

Hawkins:2018:DMC

[Haw18] Rhoda J. Hawkins. Do migrating cells need a nucleus? Journal of Cell Biology, 217(3):799–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/3/799.

Hu:2016:CAR

[HB16] Kenneth H. Hu and Manish J. Butte. T cell activation requires force generation. Journal of Cell Biology, 213(5):535–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/213/5/
535.

Hutchins:2018:DAM

[HB18] Erica J. Hutchins and Marianne E. Bronner. Draxin acts as a molecular rheostat of canonical Wnt signaling to control cranial neural crest EMT. Journal of Cell Biology, 217(10):3683–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/10/3683.

Howell:2015:WBP

[HBDW^{+}15] Michael Howell, Howard Brickner, Violaine D. Delorme-Walker, Justin Choi, Jean-Michel Saffin, Daniel Miller, Andreas Panopoulos, Céline DerMardirossian, Arun Fotedar, Robert L. Margolis, and Rati Fotedar. WISp39 binds phosphorylated Coronin 1B to regulate Arp2/3 localization and Cofilindependent motility. Journal of Cell Biology, 208(7):961–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
208/7/961.

Hayward:2019:CSE

[HBM^{+}19] Daniel Hayward, James Bancroft, Davinderpreet Mangat, Tatiana Alfonso-Pérez, Sholto Dugdale, Julia McCarthy, Francis A. Barr, and Ulrike Gruneberg. Checkpoint signaling and error correction require regulation of the MPS1 T-loop by PP2A–B56. Journal of Cell Biology, 218(10):3188–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/10/3188.

Hendrix:2015:LCO

[HBS^{+}15] Jelle Hendrix, Viola Baumgärtel, Waldemar Schrimpf, Sergey Ivanchenko, Michelle A. Digman, Enrico Gratton, Hans-Georg Kräusslich, Barbara Müller, and Don C. Lamb. Live-cell observation of cytosolic HIV-1 assembly onset reveals RNAinteracting Gag oligomers. Journal of Cell Biology, 210(4): 629–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/4/629.

Hoefert:2018:MFC

[HBWY18] Jaimee E. Hoefert, Glen A. Bjerke, Dongmei Wang, and Rui Yi. The microRNA-200 family coordinately regulates cell adhesion and proliferation in hair morphogenesis. Journal of Cell Biology, 217(6):2185–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/6/2185.

Hua:2017:VAT

[HCC^{+}17] Rong Hua, Derrick Cheng, Étienne Coyaud, Spencer Freeman, Erminia Di Pietro, Yuqing Wang, Adriano Vissa, Christopher M. Yip, Gregory D. Fairn, Nancy Braverman, John H. Brumell, William S. Trimble, Brian Raught, and Peter K. Kim. VAPs and ACBD5 tether peroxisomes to the ER for peroxisome maintenance and lipid homeostasis. Journal of Cell Biology, 216(2):367–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/2/367.

Helma:2015:NRB

[HCML15] Jonas Helma, M. Cristina Cardoso, Serge Muyldermans, and Heinrich Leonhardt. Nanobodies and recombinant binders in cell biology. Journal of Cell Biology, 209(5):633–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/209/5/633.

Hubstenberger:2015:MSR

[HCN^{+}15] Arnaud Hubstenberger, Cristiana Cameron, Scott L. Noble, Sean Keenan, and Thomas C. Evans. Modifiers of solid RNP granules control normal RNP dynamics and mRNA activity in early development. Journal of Cell Biology, 211(3): 703–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/3/703.

Hung:2018:ETT

[HCS^{+}18] Yun-Fen Hung, Chiung-Ya Chen, Yi-Chun Shih, Hsin-Yu Liu, Chiao-Ming Huang, and Yi-Ping Hsueh. Endosomal TLR3, TLR7, and TLR8 control neuronal morphology through different transcriptional programs. Journal of Cell Biology, 217(8): 2727–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/8/2727.

Hovsepian:2017:MRA

[HDA^{+}17] Junie Hovsepian, Quentin Defenouillère, Véronique Albanèse, Libuše Váchová, Camille Garcia, Zdena Palková, and Sébastien Léon. Multilevel regulation of an α-arrestin by glucose depletion controls hexose transporter endocytosis. Journal of Cell Biology, 216(6):1811–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/6/1811.

Henne:2019:SJL

[Hen19] W. Mike Henne. Spastin joins LDs and peroxisomes in the interorganelle contact ballet. Journal of Cell Biology, 218(8): 2439–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/8/2439.

Habeck:2015:YECa

[HESKK15a] Gregor Habeck, Felix A. Ebner, Hiroko Shimada-Kreft, and Stefan G. Kreft. The yeast ERAD–C ubiquitin ligase Doa10 recognizes an intramembrane degron. Journal of Cell Biology, 209(2):261–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/2/261.

Habeck:2015:YECb

[HESKK15b] Gregor Habeck, Felix A. Ebner, Hiroko Shimada-Kreft, and Stefan G. Kreft. The yeast ERAD–C ubiquitin ligase Doa10 recognizes an intramembrane degron. Journal of Cell Biology, 209(4):621–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/4/621.

Heller:2015:TPC

[HF15] Evan Heller and Elaine Fuchs. Tissue patterning and cellular mechanics. Journal of Cell Biology, 211(2):219–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/2/
219.

Hau:2017:MHP

[HGA^{+}17] Ann-Christin Hau, Britta Moyo Grebbin, Zsuzsa Agoston, Marie Anders-Maurer, Tamara Müller, Anja Groß, Jasmine Kolb, Julian D. Langer, Claudia Döring, and Dorothea Schulte. MEIS homeodomain proteins facilitate PARP1/ ARTD1-mediated eviction of histone H1. Journal of Cell Biology, 216(9):2715–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/9/2715.

Hached:2019:EAD

[HGC^{+}19] Khaled Hached, Perrine Goguet, Sophie Charrasse, Suzanne Vigneron, Maria P. Sacristan, Thierry Lorca, and Anna Castro. ENSA and ARPP19 differentially control cell cycle progression and development. Journal of Cell Biology, 218(2): 541–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/2/541.

Holtz:2015:SHI

[HGD^{+}15] Alexander M. Holtz, Samuel C. Griffiths, Samantha J. Davis, Benjamin Bishop, Christian Siebold, and Benjamin L. Allen. Secreted HHIP1 interacts with heparan sulfate and regulates Hedgehog ligand localization and function. Journal of Cell Biology, 209(5):739–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/5/739.

Hofbauer:2018:MRP

[HGF^{+}18] Harald F. Hofbauer, Michael Gecht, Sabine C. Fischer, Anja Seybert, Achilleas S. Frangakis, Ernst H. K. Stelzer, Roberto Covino, Gerhard Hummer, and Robert Ernst. The molecular recognition of phosphatidic acid by an amphipathic helix in Opi1. Journal of Cell Biology, 217(9):3109–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/9/
3109.

Hos:2017:TIE

[HGG^{+}17] Nina Judith Hos, Raja Ganesan, Saray Gutiérrez, Deniz Hos, Jennifer Klimek, Zeinab Abdullah, Martin Krönke, and Nirmal Robinson. Type I interferon enhances necroptosis of Salmonella Typhimurium–infected macrophages by impairing antioxidative stress responses. Journal of Cell Biology, 216(12):4107–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/12/4107.

Howes:2017:SDB

[HGL^{+}17] Stuart C. Howes, Elisabeth A. Geyer, Benjamin LaFrance, Rui Zhang, Elizabeth H. Kellogg, Stefan Westermann, Luke M. Rice, and Eva Nogales. Structural differences between yeast and mammalian microtubules revealed by cryo–EM. Journal of Cell Biology, 216(9):2669–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/9/2669.

Hunt:2019:MSC

[HGM^{+}19] Rachel J. Hunt, Lucy Granat, Gregory S. McElroy, Ramya Ranganathan, Navdeep S. Chandel, and Joseph M. Bateman. Mitochondrial stress causes neuronal dysfunction via an ATF4-dependent increase in L-2-hydroxyglutarate. Journal of Cell Biology, 218(12):4007–4016, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

https://rupress.org/jcb/article/218/12/4007/132528/
Mitochondrial-stress-causes-neuronal-dysfunction.

Hatch:2016:NER

[HH16] Emily M. Hatch and Martin W. Hetzer. Nuclear envelope rupture is induced by actin-based nucleus confinement. Journal of Cell Biology, 215(1):27–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/215/1/27.

Hoefig:2018:PRH

[HH18] Kai P. Hoefig and Vigo Heissmeyer. Posttranscriptional regulation of T helper cell fate decisions. Journal of Cell Biology, 217(8):2615–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/8/2615.

Hinojosa:2017:MTE

[HHBG17] Laura Soto Hinojosa, Manuel Holst, Christian Baarlink, and Robert Grosse. MRTF transcription and Ezrin-dependent plasma membrane blebbing are required for entotic invasion. Journal of Cell Biology, 216(10):3087–??, October 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/10/3087.

Hiraoka:2019:SPC

[HHCK19] Daisaku Hiraoka, Enako Hosoda, Kazuyoshi Chiba, and Takeo Kishimoto. SGK phosphorylates Cdc25 and Myt1 to trigger cyclin B–Cdk1 activation at the meiotic G2/M transition. Journal of Cell Biology, 218(11):3597–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/11/3597.

Hosoda:2019:SRP

[HHH^{+}19] Enako Hosoda, Daisaku Hiraoka, Noritaka Hirohashi, Saki Omi, Takeo Kishimoto, and Kazuyoshi Chiba. SGK regulates pH increase and cyclin B–Cdk1 activation to resume meiosis in starfish ovarian oocytes. Journal of Cell Biology, 218(11): 3612–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/11/3612.

Hsu:2015:SCP

[HHM15] FoSheng Hsu, Fenghua Hu, and Yuxin Mao. Spatiotemporal control of phosphatidylinositol 4-phosphate by Sac2 regulates endocytic recycling. Journal of Cell Biology, 209(1):97–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
209/1/97.

Horton:2016:MFS

[HHS^{+}16] Edward R. Horton, Jonathan D. Humphries, Ben Stutchbury, Guillaume Jacquemet, Christoph Ballestrem, Simon T. Barry, and Martin J. Humphries. Modulation of FAK and Src adhesion signaling occurs independently of adhesion complex composition. Journal of Cell Biology, 212(3):349–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/212/3/349.

Hansen:2018:MAD

[HHS18] David V. Hansen, Jesse E. Hanson, and Morgan Sheng. Microglia in Alzheimer’s disease. Journal of Cell Biology, 217(2): 459–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/2/459.

Heusermann:2016:ESF

[HHT^{+}16] Wolf Heusermann, Justin Hean, Dominic Trojer, Emmanuelle Steib, Stefan von Bueren, Alexandra Graff-Meyer, Christel Genoud, Katrin Martin, Nicolas Pizzato, Johannes Voshol, David V. Morrissey, Samir E. L. Andaloussi, Matthew J. Wood, and Nicole C. Meisner-Kober. Exosomes surf on filopodia to enter cells at endocytic hot spots, traffic within endosomes, and are targeted to the ER. Journal of Cell Biology, 213(2):173–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
213/2/173.

Heald:2015:TYS

[HK15] Rebecca Heald and Alexey Khodjakov. Thirty years of search and capture: The complex simplicity of mitotic spindle assembly. Journal of Cell Biology, 211(6):1103–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/6/
1103.

Hamdan:2017:SCW

[HKG17] Norfadilah Hamdan, Paraskevi Kritsiligkou, and Chris M. Grant. ER stress causes widespread protein aggregation and prion formation. Journal of Cell Biology, 216(8):2295–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/8/2295.

Hagiwara:2018:CPC

[HKG^{+}18] Akari Hagiwara, Yosuke Kitahara, Chad Paul Grabner, Christian Vogl, Manabu Abe, Ryo Kitta, Keisuke Ohta, Keiichiro Nakamura, Kenji Sakimura, Tobias Moser, Akinori Nishi, and Toshihisa Ohtsuka. Cytomatrix proteins CAST and ELKS regulate retinal photoreceptor development and maintenance. Journal of Cell Biology, 217(11):3993–??, November 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/11/3993.

Hamm:2016:SCC

[HKH16] Mailin Julia Hamm, Bettina Carmen Kirchmaier, and Wiebke Herzog. Sema3d controls collective endothelial cell migration by distinct mechanisms via Nrp1 and PlxnD1. Journal of Cell Biology, 215(3):415–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/3/415.

Homma:2019:CKA

[HKK^{+}19] Yuta Homma, Riko Kinoshita, Yoshihiko Kuchitsu, Paulina S. Wawro, Soujiro Marubashi, Mai E. Oguchi, Morié Ishida, Naonobu Fujita, and Mitsunori Fukuda. Comprehensive knockout analysis of the Rab family GTPases in epithelial cells. Journal of Cell Biology, 218(6):2035–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/6/2035.

Hennig:2015:PLD

[HKM^{+}15] Sven Hennig, Geraldine Kong, Taro Mannen, Agata Sadowska, Simon Kobelke, Amanda Blythe, Gavin J. Knott, K. Swaminathan Iyer, Diwei Ho, Estella A. Newcombe, Kana Hosoki, Naoki Goshima, Tetsuya Kawaguchi, Danny Hatters, Laura Trinkle-Mulcahy, Tetsuro Hirose, Charles S. Bond, and Archa H. Fox. Prion-like domains in RNA binding proteins are essential for building subnuclear paraspeckles. Journal of Cell Biology, 210(4):529–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/4/529.

Hori:2017:CCA

[HKT^{+}17] Tetsuya Hori, Naoko Kagawa, Atsushi Toyoda, Asao Fujiyama, Sadahiko Misu, Norikazu Monma, Fumiaki Makino, Kazuho Ikeo, and Tatsuo Fukagawa. Constitutive centromere-associated network controls centromere drift in vertebrate cells. Journal of Cell Biology, 216(1):101–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/1/101.

Huang:2018:AAM

[HLEM^{+}18] Yuejia Huang, Teng Li, Stephanie C. Ems-McClung, Claire E. Walczak, Claude Prigent, Xueliang Zhu, Xuemin Zhang, and Yixian Zheng. Aurora A activation in mitosis promoted by BuGZ. Journal of Cell Biology, 217(1):107–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/217/1/107.

Hunter:2015:PCE

[HLHFG15] Miranda V. Hunter, Donghoon M. Lee, Tony J. C. Harris, and Rodrigo Fernandez-Gonzalez. Polarized E-cadherin endocytosis directs actomyosin remodeling during embryonic wound repair. Journal of Cell Biology, 210(5):801–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/210/5/801.

Hu:2019:RMD

[HLLK19] Yuhan Hu, Wanqing Lyu, Laura Anne Lowery, and Anthony J. Koleske. Regulation of MT dynamics via direct binding of an Abl family kinase. Journal of Cell Biology, 218(12): 3986–3997, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/218/12/3986/132520/Regulation-of-MT-
dynamics-via-direct-binding-of-an.

Hochberg-Laufer:2019:ARP

[HLST19] Hodaya Hochberg-Laufer and Yaron Shav-Tal. Active RNA polymerase II curbs chromatin movement. Journal of Cell Biology, 218(5):1427–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/5/1427.

Harr:2015:DTC

[HLW^{+}15] Jennifer C. Harr, Teresa Romeo Luperchio, Xianrong Wong, Erez Cohen, Sarah J. Wheelan, and Karen L. Reddy. Directed targeting of chromatin to the nuclear lamina is mediated by chromatin state and A-type lamins. Journal of Cell Biology, 208(1):33–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/1/33.

Hu:2019:LSR

[HM19] Xiaohua Hu and R. Dyche Mullins. LC3 and STRAP regulate actin filament assembly by JMY during autophagosome formation. Journal of Cell Biology, 218(1):251–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/218/1/251.

Hsu:2016:NCC

[HMC^{+}16] Hsiang-Ting Hsu, Emily M. Mace, Alexandre F. Carisey, Dixita I. Viswanath, Athanasia E. Christakou, Martin Wiklund, Björn Önfelt, and Jordan S. Orange. NK cells converge lytic granules to promote cytotoxicity and prevent bystander killing. Journal of Cell Biology, 215(6):875–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/215/6/875.

Hooikaas:2019:MFP

[HMM^{+}19] Peter Jan Hooikaas, Maud Martin, Tobias Mühlethaler, Gert-Jan Kuijntjes, Cathelijn A. E. Peeters, Eugene A. Katrukha, Luca Ferrari, Riccardo Stucchi, Daan G. F. Verhagen, Wilhelmina E. van Riel, Ilya Grigoriev, A. F. Maarten Altelaar, Casper C. Hoogenraad, Stefan G. D. Rüdiger, Michel O. Steinmetz, Lukas C. Kapitein, and Anna Akhmanova. MAP7 family proteins regulate kinesin-1 recruitment and activation. Journal of Cell Biology, 218(4):1298–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/4/1298.

Hilfenhaus:2018:VIC

[HNF^{+}18] Georg Hilfenhaus, Dai Phuong Nguyen, Jonathan Freshman, Divya Prajapati, Feiyang Ma, Dana Song, Safiyyah Ziyad, Myriam Cuadrado, Matteo Pellegrini, Xosé R. Bustelo, and M. Luisa Iruela-Arispe. Vav3-induced cytoskeletal dynamics contribute to heterotypic properties of endothelial barriers. Journal of Cell Biology, 217(8):2813–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/8/2813.

Hashimoto:2016:PMP

[HOH^{+}16] Ari Hashimoto, Tsukasa Oikawa, Shigeru Hashimoto, Hirokazu Sugino, Ayumu Yoshikawa, Yutaro Otsuka, Haruka Handa, Yasuhito Onodera, Jin-Min Nam, Chitose Oneyama, Masato Okada, Mitsunori Fukuda, and Hisataka Sabe. P53- and mevalonate pathway-driven malignancies require Arf6 for metastasis and drug resistance. Journal of Cell Biology, 213(1):81–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
213/1/81.

Hanson:2019:PSP

[HPB19] Ryan L. Hanson, Joshua R. Porter, and Eric Batchelor. Protein stability of p53 targets determines their temporal expression dynamics in response to p53 pulsing. Journal of Cell Biology, 218(4):1282–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/4/1282.

Holzer:2019:RCI

[HPE^{+}19] Tatjana Holzer, Kristina Probst, Julia Etich, Markus Auler, Veronika S. Georgieva, Björn Bluhm, Christian Frie, Juliane Heilig, Anja Niehoff, Julian Nüchel, Markus Plomann, Jens M. Seeger, Hamid Kashkar, Olivier R. Baris, Rudolf J. Wiesner, and Bent Brachvogel. Respiratory chain inactivation links cartilage-mediated growth retardation to mitochondrial diseases. Journal of Cell Biology, 218(6):1853–??, June 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/6/1853.

Hong:2017:PCM

[HPW^{+}17] Zhi Hong, Nina Marie Pedersen, Ling Wang, Maria Lyngaas Torgersen, Harald Stenmark, and Camilla Raiborg. PtdIns3P controls mTORC1 signaling through lysosomal positioning. Journal of Cell Biology, 216(12):4217–??, December 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/12/4217.

Hong:2015:PPC

[HQW15] Nan Hyung Hong, Aidong Qi, and Alissa M. Weaver. PI(3,5)P_{2} controls endosomal branched actin dynamics by regulating cortactin–actin interactions. Journal of Cell Biology, 210(5):

753–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/5/753.

Huang:2016:OAI

[HR16] Yu-Mei Huang and Matthew N. Rasband. Organization of the axon initial segment: Actin like a fence. Journal of Cell Biology, 215(1):9–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/1/9.

Hammarskjold:2017:SPS

[HR17] Marie-Louise Hammarskjold and David Rekosh. SR proteins: To shuttle or not to shuttle, that is the question. Journal of Cell Biology, 216(7):1875–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/7/1875.

Henneberger:2016:ATC

[HS16] Christian Henneberger and Christian Steinhäuser. Astrocytic TLR4 at the crossroads of inflammation and seizure susceptibility. Journal of Cell Biology, 215(5):607–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/215/5/607.

Hariri:2019:MME

[HSB^{+}19] Hanaa Hariri, Natalie Speer, Jade Bowerman, Sean Rogers, Gang Fu, Evan Reetz, Sanchari Datta, J. Ryan Feathers, Rupali Ugrankar, Daniela Nicastro, and W. Mike Henne. Mdm1 maintains endoplasmic reticulum homeostasis by spatially regulating lipid droplet biogenesis. Journal of Cell Biology, 218 (4):1319–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/4/1319.

Hecht:2016:BCR

[HSK^{+}16] Vivian C. Hecht, Lucas B. Sullivan, Robert J. Kimmerling, Dong-Hwee Kim, Aaron M. Hosios, Max A. Stockslager, Mark M. Stevens, Joon Ho Kang, Denis Wirtz, Matthew G. Vander Heiden, and Scott R. Manalis. Biophysical changes reduce energetic demand in growth factor–deprived lymphocytes. Journal of Cell Biology, 212(4):439–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/212/4/439.

Halfmann:2019:RNR

[HSK^{+}19] Charles T. Halfmann, Rhiannon M. Sears, Aditya Katiyar, Brook W. Busselman, London K. Aman, Qiao Zhang, Christopher S. O’Bryan, Thomas E. Angelini, Tanmay P. Lele, and Kyle J. Roux. Repair of nuclear ruptures requires barrier-to-autointegration factor. Journal of Cell Biology, 218(7):2136–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/7/2136.

Hoffmann:2016:TPB

[HSN^{+}16] Saskia Hoffmann, Stine Smedegaard, Kyosuke Nakamura, Gulnahar B. Mortuza, Markus Räschle, Alain Ibañez de Opakua, Yasuyoshi Oka, Yunpeng Feng, Francisco J. Blanco, Matthias Mann, Guillermo Montoya, Anja Groth, Simon Bekker-Jensen, and Niels Mailand. TRAIP is a PCNA-binding ubiquitin ligase that protects genome stability after replication stress. Journal of Cell Biology, 212(1):63–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/1/63.

He:2018:PMI

[HSZ^{+}18] Yingli He, Hua She, Ting Zhang, Haidong Xu, Lihong Cheng, Manuel Yepes, Yingren Zhao, and Zixu Mao. p38 MAPK inhibits autophagy and promotes microglial inflammatory responses by phosphorylating ULK1. Journal of Cell Biology, 217 (1):315–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/1/315.

Herawati:2016:MCB

[HTK^{+}16] Elisa Herawati, Daisuke Taniguchi, Hatsuho Kanoh, Kazuhiro Tateishi, Shuji Ishihara, and Sachiko Tsukita. Multiciliated cell basal bodies align in stereotypical patterns coordinated by the apical cytoskeleton. Journal of Cell Biology, 214(5):571–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
214/5/571.

Hu:2018:OON

[HTLG18] Jennifer L. Hu, Michael E. Todhunter, Mark A. LaBarge, and Zev J. Gartner. Opportunities for organoids as new models of aging. Journal of Cell Biology, 217(1):39–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/217/1/39.

Hu:2015:FLN

[Hu15] Yinling Hu. A feedforward loop of NLRC5 (de)ubiquitination keeps IKK-NF-κB in check. Journal of Cell Biology, 211(5): 941–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/5/941.

Hui:2019:ICI

[Hui19] Enfu Hui. Immune checkpoint inhibitors. Journal of Cell Biology, 218(3):740–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/3/740.

Huynh:2017:DAM

[HV17] Walter Huynh and Ronald D. Vale. Disease-associated mutations in human BICD2 hyperactivate motility of dynein– dynactin. Journal of Cell Biology, 216(10):3051–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/
10/3051.

Heymann:2019:CER

[HVH^{+}19] J. Bernard Heymann, Camasamudram Vijayasarathy, Rick K. Huang, Altaira D. Dearborn, Paul A. Sieving, and Alasdair C. Steven. Cryo-EM of retinoschisin branched networks suggests an intercellular adhesive scaffold in the retina. Journal of Cell Biology, 218(3):1027–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/3/1027.

Hu:2016:SLP

[HYC16] Shi-Bin Hu, Run-Wen Yao, and Ling-Ling Chen. Shedding light on paraspeckle structure by super-resolution microscopy. Journal of Cell Biology, 214(7):789–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/7/789.

Hyrien:2015:PCM

[Hyr15] Olivier Hyrien. Peaks cloaked in the mist: The landscape of mammalian replication origins. Journal of Cell Biology, 208 (2):147–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/208/2/147.

Henne:2015:MSN

[HZB^{+}15] W. Mike Henne, Lu Zhu, Zsolt Balogi, Christopher Stefan, Jeffrey A. Pleiss, and Scott D. Emr. Mdm1/ snx13 is a novel ER–endolysosomal interorganelle tethering protein. Journal of Cell Biology, 210(4):541–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/4/541.

Hu:2015:MMA

[HZH^{+}15] Zhonghua Hu, Jun Zhao, Tianyi Hu, Yan Luo, Jun Zhu, and Zheng Li. miR-501-3p mediates the activity-dependent regulation of the expression of AMPA receptor subunit GluA1. Journal of Cell Biology, 208(7):949–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/7/949.

Ishida:2019:AFU

[IB19a] Morié Ishida and Juan S. Bonifacino. ARFRP1 functions upstream of ARL1 and ARL5 to coordinate recruitment of distinct tethering factors to the trans-Golgi network. Journal of Cell Biology, 218(11):3681–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/11/3681. See correc-

tion [IB19b].

Ishida:2019:CAF

[IB19b] Morié Ishida and Juan S. Bonifacino. Correction: ARFRP1 functions upstream of ARL1 and ARL5 to coordinate recruitment of tethering factors to the trans-Golgi network. Journal of Cell Biology, 218(11):3880–??, November 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/11/3880. See

[IB19a].

Ibarlucea-Benitez:2018:KRC

[IBFDB18] Itziar Ibarlucea-Benitez, Luke S. Ferro, David G. Drubin, and Georjana Barnes. Kinesins relocalize the chromosomal passenger complex to the midzone for spindle disassembly. Journal of Cell Biology, 217(5):1687–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/5/1687.

Ioannou:2015:DAR

[IBG^{+}15] Maria S. Ioannou, Emily S. Bell, Martine Girard, Mathilde Chaineau, Jason N. R. Hamlin, Mark Daubaras, Anie Monast, Morag Park, Louis Hodgson, and Peter S. McPherson. DENND2B activates Rab13 at the leading edge of migrating cells and promotes metastatic behavior. Journal of Cell Biology, 208(5):629–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/5/629.

Inda:2016:DCS

[IdSCB^{+}16] Carolina Inda, Paula A. dos Santos Claro, Juan J. Bonfiglio, Sergio A. Senin, Giuseppina Maccarrone, Christoph W. Turck, and Susana Silberstein. Different cAMP sources are critically involved in G protein–coupled receptor CRHR1 signaling. Journal of Cell Biology, 214(2):181–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/2/181.

Ito:2015:MPA

[IG15] Ami Ito and Gohta Goshima. Microcephaly protein Asp focuses the minus ends of spindle microtubules at the pole and within the spindle. Journal of Cell Biology, 211(5):999–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/5/
999.

Inaba:2016:NSC

[IGK^{+}16] Hironori Inaba, Hidemasa Goto, Kousuke Kasahara, Kanako Kumamoto, Shigenobu Yonemura, Akihito Inoko, Shotaro Yamano, Hideki Wanibuchi, Dongwei He, Naoki Goshima, Tohru Kiyono, Shinji Hirotsune, and Masaki Inagaki. Ndel1 suppresses ciliogenesis in proliferating cells by regulating the trichoplein–Aurora A pathway. Journal of Cell Biology, 212 (4):409–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/4/409.

Hosoi:2017:VBA

[iHMM^{+}17] Ken ichiro Hosoi, Non Miyata, Satoru Mukai, Satomi Furuki, Kanji Okumoto, Emily H. Cheng, and Yukio Fujiki. The VDAC2–BAK axis regulates peroxisomal membrane permeability. Journal of Cell Biology, 216(3):709–??, March 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/3/709.

Isensee:2018:PRS

[IKK^{+}18] Jörg Isensee, Melanie Kaufholz, Matthias J. Knape, Jan Hasenauer, Hanna Hammerich, Humberto Gonczarowska-Jorge, René P. Zahedi, Frank Schwede, Friedrich W. Herberg, and Tim Hucho. PKA-RII subunit phosphorylation precedes activation by cAMP and regulates activity termination. Journal of Cell Biology, 217(6):2167–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/6/2167.

Icha:2016:IMG

[IKRMN16] Jaroslav Icha, Christiane Kunath, Mauricio Rocha-Martins, and Caren Norden. Independent modes of ganglion cell translocation ensure correct lamination of the zebrafish retina. Journal of Cell Biology, 215(2):259–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/215/2/259.

Ioannou:2016:RMM

[IM16] Maria S. Ioannou and Peter S. McPherson. Rab-mediated membrane trafficking and the control of epithelial cell polarity. Journal of Cell Biology, 213(3):301–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/213/3/301.

Infarinato:2018:BMB

[Inf18a] Nicole Infarinato. Beata Mierzwa: Bridging the divide between science and art. Journal of Cell Biology, 217(12):4051–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/12/4051.

Infarinato:2018:DDC

[Inf18b] Nicole Infarinato. Dan Davis: Up close and personal with immune cells. Journal of Cell Biology, 217(9):2975–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/9/
2975.

Infarinato:2018:JAS

[Inf18c] Nicole Infarinato. Johan Auwerx:o Sowing the seeds of translational research. Journal of Cell Biology, 217(5):1557–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/5/
1557.

Infarinato:2019:MSW

[Inf19a] Nicole Infarinato. Mahak Sharma: Weaving through traffic. Journal of Cell Biology, 218(3):725–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/3/725.

Infarinato:2019:XWB

[Inf19b] Nicole Infarinato. Xiaochen Wang: Building up our understanding of breaking down. Journal of Cell Biology, 218(2): 381–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/2/381.

Nakajima:2019:JTS

[iNLM^{+}19] Yu ichiro Nakajima, Zachary T. Lee, Sean A. McKinney, Selene K. Swanson, Laurence Florens, and Matthew C. Gibson. Junctional tumor suppressors interact with 14-3-3 proteins to control planar spindle alignment. Journal of Cell Biology, 218 (6):1824–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/6/1824.

Infarinato:2018:BMP

[IO18] Nicole Infarinato and Marie Anne O’Donnell. BethAnn McLaughlin: Protecting neurons and women in science. Journal of Cell Biology, 217(11):3769–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/11/3769.

Ittig:2015:BTI

[ISK^{+}15] Simon J. Ittig, Christoph Schmutz, Christoph A. Kasper, Marlise Amstutz, Alexander Schmidt, Loïc Sauteur, M. Alessandra Vigano, Shyan Huey Low, Markus Affolter, Guy R. Cornelis, Erich A. Nigg, and Cécile Arrieumerlou. A bacterial type III secretion-based protein delivery tool for broad applications in cell biology. Journal of Cell Biology, 211(4):913–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/4/
913.

Izadi:2018:CLP

[ISL^{+}18] Maryam Izadi, Dirk Schlobinski, Maria Lahr, Lukas Schwintzer, Britta Qualmann, and Michael M. Kessels. Cobl-like promotes actin filament formation and dendritic branching using only a single WH2 domain. Journal of Cell Biology, 217(1):211–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/1/211.

Ishikawa:2017:UTB

[ITN^{+}17] Tokiro Ishikawa, Takuya Toyama, Yuki Nakamura, Kentaro Tamada, Hitomi Shimizu, Satoshi Ninagawa, Tetsuya Okada, Yasuhiro Kamei, Tomoko Ishikawa-Fujiwara, Takeshi Todo, Eriko Aoyama, Masaharu Takigawa, Akihiro Harada, and Kazutoshi Mori. UPR transducer BBF2H7 allows export of type II collagen in a cargo- and developmental stage– specific manner. Journal of Cell Biology, 216(6):1761–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/6/
1761.

Isokane:2016:AES

[IWM^{+}16] Mayumi Isokane, Thomas Walter, Robert Mahen, Bianca Nijmeijer, Jean-Karim Hériché, Kota Miura, Stefano Maffini, Miroslav Penchev Ivanov, Tomoya S. Kitajima, Jan-Michael Peters, and Jan Ellenberg. ARHGEF17 is an essential spindle assembly checkpoint factor that targets Mps1 to kinetochores. Journal of Cell Biology, 212(6):647–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/212/6/647.

Yamashita:2016:MDO

[iYJF^{+}16] Shun ichi Yamashita, Xiulian Jin, Kentaro Furukawa, Maho Hamasaki, Akiko Nezu, Hidenori Otera, Tetsu Saigusa, Tamotsu Yoshimori, Yasuyoshi Sakai, Katsuyoshi Mihara, and Tomotake Kanki. Mitochondrial division occurs concurrently with autophagosome formation but independently of Drp1 during mitophagy. Journal of Cell Biology, 215(5):649–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/215/5/
649.

Iuliano:2018:MPA

[IYP^{+}18] Olga Iuliano, Azumi Yoshimura, Marie-Thérèse Prospéri, René Martin, Hans-Joachim Knölker, and Evelyne Coudrier. Myosin 1b promotes axon formation by regulating actin wave propagation and growth cone dynamics. Journal of Cell Biology, 217 (6):2033–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/6/2033.

Izzo:2017:DCH

[IZBH^{+}17] Annalisa Izzo, Céline Ziegler-Birling, Peter W. S. Hill, Lydia Brondani, Petra Hajkova, Maria-Elena Torres-Padilla, and Robert Schneider. Dynamic changes in H1 subtype composition during epigenetic reprogramming. Journal of Cell Biology, 216(10):3017–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/10/3017.

Inoue:2018:SPM

[IZZ^{+}18] Takamasa Inoue, Pengwei Zhang, Wei Zhang, Kylia Goodner-Bingham, Allison Dupzyk, Daniel DiMaio, and Billy Tsai. γ-secretase promotes membrane insertion of the human papillomavirus L2 capsid protein during virus infection. Journal of Cell Biology, 217(10):3545–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/10/3545.

Jones:2018:CAR

[JAHH18] Matthew C. Jones, Janet A. Askari, Jonathan D. Humphries, and Martin J. Humphries. Cell adhesion is regulated by CDK1 during the cell cycle. Journal of Cell Biology, 217(9):3203– ??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/9/3203.

Janody:2018:BBT

[Jan18] Florence Janody. The Big Bang of tissue growth: Apical cell constriction turns into tissue expansion. Journal of Cell Biology, 217(3):807–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/3/807.

Juanes:2017:APC

[JBE^{+}17] M. Angeles Juanes, Habib Bouguenina, Julian A. Eskin, Richa Jaiswal, Ali Badache, and Bruce L. Goode. Adenomatous polyposis coli nucleates actin assembly to drive cell migration and microtubule-induced focal adhesion turnover. Journal of Cell Biology, 216(9):2859–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/9/2859.

Jourdain:2016:MRG

[JBMM16] Alexis A. Jourdain, Erik Boehm, Kinsey Maundrell, and Jean-Claude Martinou. Mitochondrial RNA granules: Compartmentalizing mitochondrial gene expression. Journal of Cell Biology, 212(6):611–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/6/611.

Ji:2017:RMD

[JCF^{+}17] Wei-Ke Ji, Rajarshi Chakrabarti, Xintao Fan, Lori Schoenfeld, Stefan Strack, and Henry N. Higgs. Receptor-mediated Drp1 oligomerization on endoplasmic reticulum. Journal of Cell Biology, 216(12):4123–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/12/4123.

Jayadev:2019:TID

[JCK^{+}19] Ranjay Jayadev, Qiuyi Chi, Daniel P. Keeley, Eric L. Hastie, Laura C. Kelley, and David R. Sherwood. α-integrins dictate distinct modes of type IV collagen recruitment to basement membranes. Journal of Cell Biology, 218(9):3098–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/9/

Jarsch:2016:MCC

[JDG16] Iris K. Jarsch, Frederic Daste, and Jennifer L. Gallop. Membrane curvature in cell biology: an integration of molecular mechanisms. Journal of Cell Biology, 214(4):375–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/214/4/
375.

Jordan:2016:CPP

[JDZ^{+}16] Shawn N. Jordan, Tim Davies, Yelena Zhuravlev, Julien Dumont, Mimi Shirasu-Hiza, and Julie C. Canman. Cortical PAR polarity proteins promote robust cytokinesis during asymmetric cell division. Journal of Cell Biology, 212(1):39–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/212/1/
39.

Jiang:2015:MMT

[JERL^{+}15] Mei Jiang, Julian Esteve-Rudd, Vanda S. Lopes, Tanja Diemer, Concepción Lillo, Agrani Rump, and David S. Williams. Microtubule motors transport phagosomes in the RPE, and lack of KLC1 leads to AMD-like pathogenesis. Journal of Cell Biology, 210(4):595–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/4/595.

Juanes-Garcia:2015:RMN

[JGCAC^{+}15] Alba Juanes-Garcia, Jessica R. Chapman, Rocio Aguilar-Cuenca, Cristina Delgado-Arevalo, Jennifer Hodges, Leanna A. Whitmore, Jeffrey Shabanowitz, Donald F. Hunt, Alan Rick Horwitz, and Miguel Vicente-Manzanares. A regulatory motif in nonmuscle myosin II–B regulates its role in migratory front– back polarity. Journal of Cell Biology, 209(1):23–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/209/1/23.

Jiang:2019:PCC

[JH19] Tao Jiang and Tony J. C. Harris. Par-1 controls the composition and growth of cortical actin caps during Drosophila embryo cleavage. Journal of Cell Biology, 218(12):4195–4214, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/218/

12/4195/132548/Par-1-controls-the-composition-and-growth-
of.

Joshi:2016:FMS

[JHC^{+}16] Amit S. Joshi, Xiaofang Huang, Vineet Choudhary, Tim P. Levine, Junjie Hu, and William A. Prinz. A family of membrane-shaping proteins at ER subdomains regulates preperoxisomal vesicle biogenesis. Journal of Cell Biology, 215(4): 515–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/215/4/515.

Jiang:2015:RAA

[JHF^{+}15] Hao Jiang, Xiaonan He, Di Feng, Xueliang Zhu, and Yixian Zheng. RanGTP aids anaphase entry through Ubr5-mediated protein turnover. Journal of Cell Biology, 211(1):7–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/1/
7.

Jiang:2015:SDT

[JhZbYmP15] Jian Jiang, Zheng hong Zhang, Xiao bin Yuan, and Mu ming Poo. Spatiotemporal dynamics of traction forces show three contraction centers in migratory neurons. Journal of Cell Biology, 209(5):759–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/5/759.

Juanes:2019:RAM

[JIB^{+}19] M. Angeles Juanes, Daniel Isnardon, Ali Badache, Sophie Brasselet, Manos Mavrakis, and Bruce L. Goode. The role of APCmediated actin assembly in microtubule capture and focal adhesion turnover. Journal of Cell Biology, 218(10):3415–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/10/3415.

Judith:2019:ASF

[JJB^{+}19] Delphine Judith, Harold B. J. Jefferies, Stefan Boeing, David Frith, Ambrosius P. Snijders, and Sharon A. Tooze. ATG9A shapes the forming autophagosome through Arfaptin 2 and phosphatidylinositol 4-kinase IIIβ. Journal of Cell Biology, 218 (5):1634–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/5/1634.

Jin:2017:EPS

[JJW17] Natsuko Jin, Yui Jin, and Lois S. Weisman. Early protection to stress mediated by CDK-dependent PI3,5P_{2} signaling from the vacuole/lysosome. Journal of Cell Biology, 216(7):2075–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/7/2075.

Johnson:2015:FAB

[JKA^{+}15] Heath E. Johnson, Samantha J. King, Sreeja B. Asokan, Jeremy D. Rotty, James E. Bear, and Jason M. Haugh. Factin bundles direct the initiation and orientation of lamellipodia through adhesion-based signaling. Journal of Cell Biology, 208(4):443–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/4/443.

Juettner:2019:VPS

[JKD^{+}19] Vanessa V. Juettner, Kevin Kruse, Arkaprava Dan, Vinh H. Vu, Yousaf Khan, Jonathan Le, Deborah Leckband, Yulia Komarova, and Asrar B. Malik. VE-PTP stabilizes VE-cadherin junctions and the endothelial barrier via a phosphataseindependent mechanism. Journal of Cell Biology, 218(5):1725– ??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/5/1725.

Jores:2018:CHH

[JLB^{+}18] Tobias Jores, Jannis Lawatscheck, Viktor Beke, Mirita Franz-Wachtel, Kaori Yunoki, Julia C. Fitzgerald, Boris Macek, Toshiya Endo, Hubert Kalbacher, Johannes Buchner, and Doron Rapaport. Cytosolic Hsp70 and Hsp40 chaperones enable the biogenesis of mitochondrial β-barrel proteins. Journal of Cell Biology, 217(9):3091–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/9/3091.

Jurgensen:2019:IRF

[JNS^{+}19] Henrik J. Jürgensen, Kirstine S. Nørregaard, Megan M. Sibree, Eric Santoni-Rugiu, Daniel H. Madsen, Katharina Wassilew, Dorrit Krustrup, Peter Garred, Thomas H. Bugge, Lars H. Engelholm, and Niels Behrendt. Immune regulation by fibroblasts in tissue injury depends on uPARAP-mediated uptake of collectins. Journal of Cell Biology, 218(1):333–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/218/1/333.

Jovasevic:2015:MPE

[JNW15] Vladimir Jovasevic, Mojgan H. Naghavi, and Derek Walsh. Microtubule plus end–associated CLIP-170 initiates HSV-1 retrograde transport in primary human cells. Journal of Cell Biology, 211(2):323–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/2/323.

Johnson:2016:PLW

[JOJG16] Danielle E. Johnson, Philip Ostrowski, Valentin Jaumouillé, and Sergio Grinstein. The position of lysosomes within the cell determines their luminal pH. Journal of Cell Biology, 212 (6):677–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/6/677.

Jordan:2016:CKH

[Jor16a] Shawn Jordan. Claudine Kraft: a hunger for understanding. Journal of Cell Biology, 213(6):602–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/213/6/602.

Jordan:2016:CBA

[Jor16b] Shawn Jordan. Cliff Brangwynne: All the right materials. Journal of Cell Biology, 214(2):122–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/2/122.

Jordan:2016:DSC

[Jor16c] Shawn Jordan. David Spector: Coordinating gene expression in space and time. Journal of Cell Biology, 214(7):786–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/214/7/

Jordan:2016:FCS

[Jor16d] Shawn Jordan. Fred Chang: The shape of things to come. Journal of Cell Biology, 213(5):498–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/213/5/498.

Jordan:2016:FBS

[Jor16e] Shawn Jordan. Frederic Bard: The sweet side of traffic. Journal of Cell Biology, 214(3):234–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/214/3/234.

Jordan:2016:HST

[Jor16f] Shawn Jordan. Hari Shroff: Taking a closer look. Journal of Cell Biology, 214(4):360–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/214/4/360.

Jordan:2016:HFP

[Jor16g] Shawn Jordan. Heike Folsch: Peeling back the layers. Journal of Cell Biology, 213(4):405–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/213/4/405.

Jordan:2016:MKF

[Jor16h] Shawn Jordan. Megan King: a force to be reckoned with. Journal of Cell Biology, 214(1):2–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/214/1/2.

Jordan:2016:RCL

[Jor16i] Shawn Jordan. Richard Cheney: Life on the move. Journal of Cell Biology, 214(5):492–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/214/5/492.

Jacquemet:2017:FRI

[JPC^{+}17] Guillaume Jacquemet, Ilkka Paatero, Alexandre F. Carisey, Artur Padzik, Jordan S. Orange, Hellyeh Hamidi, and Johanna Ivaska. FiloQuant reveals increased filopodia density during breast cancer progression. Journal of Cell Biology, 216(10): 3387–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/10/3387.

Joensuu:2016:STI

[JPD^{+}16] Merja Joensuu, Pranesh Padmanabhan, Nela Durisic, Adekunle T. D. Bademosi, Elizabeth Cooper-Williams, Isabel C. Morrow, Callista B. Harper, WooRam Jung, Robert G. Parton, Geoffrey J. Goodhill, Andreas Papadopulos, and Frédéric A. Meunier. Subdiffractional tracking of internalized molecules reveals heterogeneous motion states of synaptic vesicles. Journal of Cell Biology, 215(2):277–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/215/2/277.

Jungas:2016:EMT

[JPF^{+}16] Thomas Jungas, Renaud T. Perchey, Mohamad Fawal, Caroline Callot, Carine Froment, Odile Burlet-Schiltz, Arnaud Besson, and Alice Davy. Eph-mediated tyrosine phosphorylation of citron kinase controls abscission. Journal of Cell Biology, 214 (5):555–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/214/5/555.

Jonasson:2016:ZZP

[JRH^{+}16] Erin M. Jonasson, Valentina Rossio, Riko Hatakeyama, Mitsuhiro Abe, Yoshikazu Ohya, and Satoshi Yoshida. Zds1/ zds2– PP2A^{Cdc55} complex specifies signaling output from Rho1 GT-Pase. Journal of Cell Biology, 212(1):51–??, January 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/212/1/51.

Jung:2018:CFF

[JSB^{+}18] WooRam Jung, Emma Sierecki, Michele Bastiani, Ailis O’Carroll, Kirill Alexandrov, James Rae, Wayne Johnston, Dominic J. B. Hunter, Charles Ferguson, Yann Gambin, Nicholas Ariotti, and Robert G. Parton. Cell-free formation and interactome analysis of caveolae. Journal of Cell Biology, 217 (6):2141–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/6/2141.

Juhasz:2016:MDV

[Juh16] Gábor Juhász. A mitochondrial-derived vesicle HOPS to endolysosomes using Syntaxin-17. Journal of Cell Biology, 214 (3):241–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/214/3/241.

Jost:2019:DRM

[JW19] Anna Payne-Tobin Jost and Jennifer C. Waters. Designing a rigorous microscopy experiment: Validating methods and avoiding bias. Journal of Cell Biology, 218(5):1452–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/5/
1452.

Kawauchi:2017:TIS

[Kaw17] Takeshi Kawauchi. Tubulin isotype specificity in neuronal migration: Tuba8 can’t fill in for Tuba1a. Journal of Cell Biology, 216(8):2247–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/8/2247.

Kay:2016:RMR

[Kay16] Jeremy N. Kay. Radial migration: Retinal neurons hold on for the ride. Journal of Cell Biology, 215(2):147–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/215/2/
147.

Kumar:2015:KSR

[KBB^{+}15] Vinod Kumar, Jamal-Eddine Bouameur, Janina Bär, Robert H. Rice, Hue-Tran Hornig-Do, Dennis R. Roop, Nicole Schwarz, Susanne Brodesser, Sören Thiering, Rudolf E. Leube, Rudolf J. Wiesner, Preethi Vijayaraj, Christina B. Brazel, Sandra Heller, Hans Binder, Henry Löffler-Wirth, Peter Seibel, and Thomas M. Magin. A keratin scaffold regulates epidermal barrier formation, mitochondrial lipid composition, and activity. Journal of Cell Biology, 211(5):1057–??, December 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/211/5/1057.

Kumar:2016:CKS

[KBB^{+}16] Vinod Kumar, Jamal-Eddine Bouameur, Janina Bär, Robert H. Rice, Hue-Tran Hornig-Do, Dennis R. Roop, Nicole Schwarz, Susanne Brodesser, Sören Thiering, Rudolf E. Leube, Rudolf J. Wiesner, Preethi Vijayaraj, Christina B. Brazel, Sandra Heller, Hans Binder, Henry Löffler-Wirth, Peter Seibel, and Thomas M. Magin. Correction: A keratin scaffold regulates epidermal barrier formation, mitochondrial lipid composition, and activity. Journal of Cell Biology, 212(7):877–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/212/7/877.

Kruger:2017:INC

[KBB^{+}17] Vivien Krüger, Thomas Becker, Lars Becker, Malayko Montilla-Martinez, Lars Ellenrieder, F.-Nora Vögtle, Helmut E. Meyer, Michael T. Ryan, Nils Wiedemann, Bettina Warscheid, Nikolaus Pfanner, Richard Wagner, and Chris Meisinger. Identification of new channels by systematic analysis of the mitochondrial outer membrane. Journal of Cell Biology, 216(11): 3485–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/11/3485.

Krols:2016:MCS

[KBJ16] Michiel Krols, Geert Bultynck, and Sophie Janssens. ER– mitochondria contact sites: a new regulator of cellular calcium flux comes into play. Journal of Cell Biology, 214(4):367–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
214/4/367.

Karasu:2019:CBP

[KBKW19] Mehmet E. Karasu, Nora Bouftas, Scott Keeney, and Katja Wassmann. Cyclin B3 promotes anaphase I onset in oocyte meiosis. Journal of Cell Biology, 218(4):1265–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/4/
1265.

Kubo:2015:SCI

[KBT^{+}15] Yusuke Kubo, Kentarou Baba, Michinori Toriyama, Takunori Minegishi, Tadao Sugiura, Satoshi Kozawa, Kazushi Ikeda, and Naoyuki Inagaki. Shootin1–cortactin interaction mediates signal–force transduction for axon outgrowth. Journal of Cell Biology, 210(4):663–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/4/663.

Klapproth:2019:KLP

[KBT^{+}19] Sarah Klapproth, Thomas Bromberger, Clara Türk, Marcus Krüger, and Markus Moser. A kindlin-3–leupaxin–paxillin signaling pathway regulates podosome stability. Journal of Cell Biology, 218(10):3436–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/218/10/3436.

Kasula:2016:MDH

[KCB^{+}16] Ravikiran Kasula, Ye Jin Chai, Adekunle T. Bademosi, Callista B. Harper, Rachel S. Gormal, Isabel C. Morrow, Eric Hosy, Brett M. Collins, Daniel Choquet, Andreas Papadopulos, and Frédéric A. Meunier. The Munc18-1 domain 3a hinge-loop controls syntaxin-1A nanodomain assembly and engagement with the SNARE complex during secretory vesicle priming. Journal of Cell Biology, 214(7):847–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/7/847.

Keeler:2017:EOA

[KD17a] Austin B. Keeler and Christopher D. Deppmann. The evolutionary origins of antagonistic neurotrophin signaling. Journal of Cell Biology, 216(5):1223–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/5/1223.

Kuhn:2017:SAC

[KD17b] Jonathan Kuhn and Sophie Dumont. Spindle assembly checkpoint satisfaction occurs via end-on but not lateral attachments under tension. Journal of Cell Biology, 216(6):1533–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/6/
1533.

Kuhn:2019:MKC

[KD19] Jonathan Kuhn and Sophie Dumont. Mammalian kinetochores count attached microtubules in a sensitive and switchlike manner. Journal of Cell Biology, 218(11):3583–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/11/3583.

Kavaler:2018:MSN

[KDA^{+}18] Joshua Kavaler, Hong Duan, Rajaguru Aradhya, Luis F. de Navas, Brian Joseph, Boris Shklyar, and Eric C. Lai. miRNA suppression of a Notch repressor directs non-neuronal fate in Drosophila mechanosensory organs. Journal of Cell Biology, 217(2):571–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/2/571.

Knoops:2015:YPC

[KdBKvdK15] Kèvin Knoops, Rinse de Boer, Anita Kram, and Ida J. van der Klei. Yeast pex1 cells contain peroxisomal ghosts that import matrix proteins upon reintroduction of Pex1. Journal of Cell Biology, 211(5):955–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/211/5/955.

Kam:2018:DMG

[KDM^{+}18] Chen Yuan Kam, Adi D. Dubash, Elisa Magistrati, Simona Polo, Karla J. F. Satchell, Farah Sheikh, Paul D. Lampe, and Kathleen J. Green. Desmoplakin maintains gap junctions by inhibiting Ras/MAPK and lysosomal degradation of connexin-43. Journal of Cell Biology, 217(9):3219–??, September 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/9/3219.

Kendrick:2019:HSO

[KDR^{+}19] Agnieszka A. Kendrick, Andrea M. Dickey, William B. Redwine, Phuoc Tien Tran, Laura Pontano Vaites, Monika Dzieciatkowska, J. Wade Harper, and Samara L. Reck-Peterson. Hook3 is a scaffold for the opposite-polarity microtubule-based motors cytoplasmic dynein-1 and KIF1C. Journal of Cell Biology, 218(9):2982–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/9/2982.

Kuhn:2015:RBI

[KDV^{+}15] Patrick Kuhn, Albena Draycheva, Andreas Vogt, Narcis-Adrian Petriman, Lukas Sturm, Friedel Drepper, Bettina Warscheid, Wolfgang Wintermeyer, and Hans-Georg Koch. Ribosome binding induces repositioning of the signal recognition particle receptor on the translocon. Journal of Cell Biology, 211(1):91–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
211/1/91.

Kelly:2016:HMM

[Kel16] Robert G. Kelly. How Mesp1 makes a move. Journal of Cell Biology, 213(4):411–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/213/4/411.

Karg:2017:CKM

[KEV^{+}17] Travis Karg, Mary Williard Elting, Hannah Vicars, Sophie Dumont, and William Sullivan. The chromokinesin Klp3a and microtubules facilitate acentric chromosome segregation. Journal of Cell Biology, 216(6):1597–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/6/1597.

Keyes:2018:SCA

[KF18] Brice E. Keyes and Elaine Fuchs. Stem cells: Aging and transcriptional fingerprints. Journal of Cell Biology, 217(1):79–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/1/79.

Konig:2017:MRD

[KFAMR17] Julia König, E. B. Frankel, Anjon Audhya, and Thomas Müller-Reichert. Membrane remodeling during embryonic abscission in Caenorhabditis elegans. Journal of Cell Biology, 216(5):1277– ??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/5/1277.

Kim:2015:AFR

[KG15] Nam-Gyun Kim and Barry M. Gumbiner. Adhesion to fibronectin regulates Hippo signaling via the FAK–Src–PI3K pathway. Journal of Cell Biology, 210(3):503–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/210/3/503.

Kawaguchi:2019:NNN

[KG19] Daichi Kawaguchi and Yukiko Gotoh. Neurexin nanoclusters: a novel structure at presynaptic terminals. Journal of Cell Biology, 218(8):2442–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/8/2442.

Klebba:2015:TPL

[KGN^{+}15] Joseph E. Klebba, Brian J. Galletta, Jonathan Nye, Karen M. Plevock, Daniel W. Buster, Natalie A. Hollingsworth, Kevin C. Slep, Nasser M. Rusan, and Gregory C. Rogers. Two Polo-like kinase 4 binding domains in Asterless perform distinct roles in regulating kinase stability. Journal of Cell Biology, 208(4): 401–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/208/4/401.

Kumsta:2019:GUS

[KH19] Caroline Kumsta and Malene Hansen. Getting under the skin: Cuticle damage elicits systemic autophagy response in C. elegans. Journal of Cell Biology, 218(12): 3885–3887, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/218/12/3885/132509/Getting-
under-the-skin-Cuticle-damage-elicits.

Karanyi:2018:NDS

[KHA^{+}18] Zsolt Karányi, László Halász, Laurent Acquaviva, Dávid Jónás, Szabolcs Hetey, Beáta Boros-Oláh, Feng Peng, Doris Chen, Franz Klein, Vincent Géli, and Lóránt Székvölgyi. Nuclear dynamics of the Set1C subunit Spp1 prepares meiotic recombination sites for break formation. Journal of Cell Biology, 217(10): 3398–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/10/3398.

Kapinos:2017:KRN

[KHRL17] Larisa E. Kapinos, Binlu Huang, Chantal Rencurel, and Roderick Y. H. Lim. Karyopherins regulate nuclear pore complex barrier and transport function. Journal of Cell Biology, 216(11): 3609–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/11/3609.

Katsunuma:2016:SAN

[KHS^{+}16] Sayaka Katsunuma, Hisao Honda, Tomoyasu Shinoda, Yukitaka Ishimoto, Takaki Miyata, Hiroshi Kiyonari, Takaya Abe, Ken ichi Nibu, Yoshimi Takai, and Hideru Togashi. Synergistic action of nectins and cadherins generates the mosaic cellular pattern of the olfactory epithelium. Journal of Cell Biology, 212(5):561–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/5/561.

Kim:2016:SNC

[KJ16] Ju Youn Kim and Samie R. Jaffrey. Separating neuronal compartments gives clues as to local effect of ubiquitin conjugates in synaptogenesis. Journal of Cell Biology, 212(7):751–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/212/7/
751.

Kimura:2015:TMP

[KJC^{+}15] Tomonori Kimura, Ashish Jain, Seong Won Choi, Michael A. Mandell, Kate Schroder, Terje Johansen, and Vojo Deretic. TRIM-mediated precision autophagy targets cytoplasmic regulators of innate immunity. Journal of Cell Biology, 210(6): 973–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/6/973.

Kumar:2018:MSR

[KJF^{+}18] Suresh Kumar, Ashish Jain, Farzin Farzam, Jingyue Jia, Yuexi Gu, Seong Won Choi, Michal H. Mudd, Aurore Claude-Taupin, Michael J. Wester, Keith A. Lidke, Tor-Erik Rusten, and Vojo Deretic. Mechanism of Stx17 recruitment to autophagosomes via IRGM and mammalian Atg8 proteins. Journal of Cell Biology, 217(3):997–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/3/997.

Klaips:2018:PCP

[KJH18] Courtney L. Klaips, Gopal Gunanathan Jayaraj, and F. Ulrich Hartl. Pathways of cellular proteostasis in aging and disease. Journal of Cell Biology, 217(1):51–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/1/51.

Kvainickas:2017:CSS

[KJON^{+}17] Arunas Kvainickas, Ana Jimenez-Orgaz, Heike Nägele, Zehan Hu, Jörn Dengjel, and Florian Steinberg. Cargo-selective SNX– BAR proteins mediate retromer trimer independent retrograde transport. Journal of Cell Biology, 216(11):3677–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/11/3677.

Koch:2019:APC

[KJTY19] Bailey A. Koch, Hui Jin, Robert J. Tomko, and Hong-Guo Yu. The anaphase-promoting complex regulates the degradation of the inner nuclear membrane protein Mps3. Journal of Cell Biology, 218(3):839–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/3/839.

Khan:2019:TTC

[KJZ^{+}19] Liakot A. Khan, Gholamali Jafari, Nan Zhang, Edward Membreno, Siyang Yan, Hongjie Zhang, and Verena Gobel. A tensile trilayered cytoskeletal endotube drives capillary-like lumenogenesis. Journal of Cell Biology, 218(7):2403–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/7/
2403.

Kalinski:2019:DMH

[KKC^{+}19] Ashley L. Kalinski, Amar N. Kar, John Craver, Andrew P. Tosolini, James N. Sleigh, Seung Joon Lee, Alicia Hawthorne, Paul Brito-Vargas, Sharmina Miller-Randolph, Ryan Passino, Liang Shi, Victor S. C. Wong, Cristina Picci, Deanna S. Smith, Dianna E. Willis, Leif A. Havton, Giampietro Schiavo, Roman J. Giger, Brett Langley, and Jeffery L. Twiss. Deacetylation of Miro1 by HDAC6 blocks mitochondrial transport and mediates axon growth inhibition. Journal of Cell Biology, 218 (6):1871–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/6/1871.

Krey:2016:PWS

[KKD^{+}16] Jocelyn F. Krey, Evan S. Krystofiak, Rachel A. Dumont, Sarath Vijayakumar, Dongseok Choi, Francisco Rivero, Bechara Kachar, Sherri M. Jones, and Peter G. Barr-Gillespie. Plastin 1 widens stereocilia by transforming actin filament packing from hexagonal to liquid. Journal of Cell Biology, 215(4): 467–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/215/4/467.

Kim:2017:CCS

[KKP^{+}17] Sang Bum Kim, Hye Rim Kim, Min Chul Park, Seongmin Cho, Peter C. Goughnour, Daeyoung Han, Ina Yoon, YounHa Kim, Taehee Kang, Eunjoo Song, Pilhan Kim, Hyosun Choi, Ji Young Mun, Chihong Song, Sangmin Lee, Hyun Suk Jung, and Sunghoon Kim. Caspase-8 controls the secretion of inflammatory lysyl-tRNA synthetase in exosomes from cancer cells. Journal of Cell Biology, 216(7):2201–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/7/2201.

Kraft:2017:MDA

[KL17] Lauren M. Kraft and Laura L. Lackner. Mitochondria-driven assembly of a cortical anchor for mitochondria and dynein. Journal of Cell Biology, 216(10):3061–??, October 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/10/3061.

King:2019:LNE

[KL19] Megan C. King and C. Patrick Lusk. Loss of nuclear envelope integrity? No probLEM — BAF has it covered. Journal of Cell Biology, 218(7):2077–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/7/2077.

Kumar:2018:VVL

[KLHC^{+}18] Nikit Kumar, Marianna Leonzino, William Hancock-Cerutti, Florian A. Horenkamp, PeiQi Li, Joshua A. Lees, Heather Wheeler, Karin M. Reinisch, and Pietro De Camilli. VPS13A and VPS13C are lipid transport proteins differentially localized at ER contact sites. Journal of Cell Biology, 217(10):3625–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/10/3625.

Kruse:2019:CST

[KLS^{+}19] Kevin Kruse, Quinn S. Lee, Ying Sun, Jeff Klomp, Xiaoyan Yang, Fei Huang, Mitchell Y. Sun, Shuangping Zhao, Zhigang Hong, Stephen M. Vogel, Jae-Won Shin, Deborah E. Leckband, Leon M. Tai, Asrar B. Malik, and Yulia A. Komarova. Ncadherin signaling via Trio assembles adherens junctions to restrict endothelial permeability. Journal of Cell Biology, 218 (1):299–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/1/299.

Kawamata:2017:POD

[KM17] Hibiki Kawamata and Giovanni Manfredi. Proteinopathies and OXPHOS dysfunction in neurodegenerative diseases. Journal of Cell Biology, 216(12):3917–??, December 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/12/3917. See

[KM18a].

Kawamata:2018:CPO

[KM18a] Hibiki Kawamata and Giovanni Manfredi. Correction: Proteinopathies and OXPHOS dysfunction in neurodegenerative diseases. Journal of Cell Biology, 217(1):429–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/217/1/429.

See [KM17].

Kulkarni:2018:NEL

[KM18b] Vineet Vinay Kulkarni and Sandra Maday. Neuronal endosomes to lysosomes: a journey to the soma. Journal of Cell Biology, 217(9):2977–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/9/2977.

Kalaidzidis:2015:AEO

[KMBO^{+}15] Inna Kalaidzidis, Marta Miaczynska, Marta Brewińska-Olchowik, Anna Hupalowska, Charles Ferguson, Robert G. Parton, Yannis Kalaidzidis, and Marino Zerial. APPL endosomes are not obligatory endocytic intermediates but act as stable cargo-sorting compartments. Journal of Cell Biology, 211(1): 123–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/1/123.

Kodani:2019:SPC

[KMC^{+}19] Andrew Kodani, Tyler Moyer, Allen Chen, Andrew Holland, Christopher A. Walsh, and Jeremy F. Reiter. SFI1 promotes centriole duplication by recruiting USP9X to stabilize the microcephaly protein STIL. Journal of Cell Biology, 218(7):2185– ??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/7/2185.

Kriebel:2018:EVD

[KMJ^{+}18] Paul W. Kriebel, Ritankar Majumdar, Lisa M. Jenkins, Hiroshi Senoo, Weiye Wang, Sonia Ammu, Song Chen, Kedar Narayan, Miho Iijima, and Carole A. Parent. Extracellular vesicles direct migration by synthesizing and releasing chemotactic signals. Journal of Cell Biology, 217(8):2891–??, August 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/8/2891.

Kawabe:2017:CEL

[KMK^{+}17a] Hiroshi Kawabe, Miso Mitkovski, Pascal S. Kaeser, Johannes Hirrlinger, Felipe Opazo, Dennis Nestvogel, Stefan Kalla, Anna Fejtova, Sophie E. Verrier, Simon R. Bungers, Benjamin H. Cooper, Frederique Varoqueaux, Yun Wang, Ralf B. Nehring, Eckart D. Gundelfinger, Christian Rosenmund, Silvio O. Rizzoli, Thomas C. Südhof, Jeong-Seop Rhee, and Nils Brose. Correction: ELKS1 localizes the synaptic vesicle priming protein bMunc13-2 to a specific subset of active zones. Journal of Cell Biology, 216(4):1205–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/4/1205. See [KMK^{+}17b].

Kawabe:2017:ELS

[KMK^{+}17b] Hiroshi Kawabe, Miso Mitkovski, Pascal S. Kaeser, Johannes Hirrlinger, Felipe Opazo, Dennis Nestvogel, Stefan Kalla, Anna Fejtova, Sophie E. Verrier, Simon R. Bungers, Benjamin H. Cooper, Frederique Varoqueaux, Yun Wang, Ralf B. Nehring, Eckart D. Gundelfinger, Christian Rosenmund, Silvio O. Rizzoli, Thomas C. Südhof, Jeong-Seop Rhee, and Nils Brose. ELKS1 localizes the synaptic vesicle priming protein bMunc13-2 to a specific subset of active zones. Journal of Cell Biology, 216(4):1143–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress. org/content/216/4/1143. See correction [KMK^{+}17a].

Konno:2015:EIP

[KML^{+}15] Tasuku Konno, Eduardo Pinho Melo, Carlos Lopes, Ilir Mehmeti, Sigurd Lenzen, David Ron, and Edward Avezov. ERO1-independent production of H_{2} O_{2} within the endoplasmic reticulum fuels Prdx4-mediated oxidative protein folding. Journal of Cell Biology, 211(2):253–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/211/2/253.

Kim:2015:KLB

[KMLG^{+}15] Taekyung Kim, Mark W. Moyle, Pablo Lara-Gonzalez, Christian De Groot, Karen Oegema, and Arshad Desai. Kinetochorelocalized BUB-1/BUB-3 complex promotes anaphase onset in C. elegans. Journal of Cell Biology, 209(4):507–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/209/4/507.

See correction [KMLG^{+}16].

Kim:2016:CKL

[KMLG^{+}16] Taekyung Kim, Mark W. Moyle, Pablo Lara-Gonzalez, Christian De Groot, Karen Oegema, and Arshad Desai. Correction: Kinetochore-localized BUB-1/BUB-3 complex promotes anaphase onset in C. elegans. Journal of Cell Biology, 212 (6):737–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/6/737. See [KMLG^{+}15].

Korwitz:2016:LOD

[KMRD^{+}16] Anne Korwitz, Carsten Merkwirth, Ricarda Richter-Dennerlein, Simon E. Tröder, Hans-Georg Sprenger, Pedro M. Quirós, Carlos López-Otín, Elena I. Rugarli, and Thomas Langer. Loss of OMA1 delays neurodegeneration by preventing stress-induced OPA1 processing in mitochondria. Journal of Cell Biology, 212 (2):157–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/2/157.

Kourtidis:2017:CCR

[KNL^{+}17] Antonis Kourtidis, Brian Necela, Wan-Hsin Lin, Ruifeng Lu, Ryan W. Feathers, Yan W. Asmann, E. Aubrey Thompson, and Panos Z. Anastasiadis. Cadherin complexes recruit mRNAs and RISC to regulate epithelial cell signaling. Journal of Cell Biology, 216(10):3073–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/216/10/3073.

Kern:2016:MSI

[KNPC16] David M. Kern, Peter K. Nicholls, David C. Page, and Iain M. Cheeseman. A mitotic SKAP isoform regulates spindle positioning at astral microtubule plus ends. Journal of Cell Biology, 213(3):315–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/3/315.

Kvainickas:2019:RTM

[KNQ^{+}19] Arunas Kvainickas, Heike Nägele, Wenjing Qi, Ladislav Dokládal, Ana Jimenez-Orgaz, Luca Stehl, Dipak Gangurde, Qian Zhao, Zehan Hu, Jörn Dengjel, Claudio De Virgilio, Ralf Baumeister, and Florian Steinberg. Retromer and TBC1D5 maintain late endosomal RAB7 domains to enable amino acid– induced mTORC1 signaling. Journal of Cell Biology, 218(9): 3019–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/9/3019.

Kahanovitch:2019:GSC

[KO19] Uri Kahanovitch and Michelle L. Olsen. Glial SIK3: a central player in ion and volume homeostasis in Drosophila peripheral nerves. Journal of Cell Biology, 218(12):3888– 3889, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/218/12/3888/132529/Glial-SIK3-A-central-
player-in-ion-and-volume.

Kunii:2016:ORS

[KOIT^{+}16] Masataka Kunii, Mica Ohara-Imaizumi, Noriko Takahashi, Masaki Kobayashi, Ryosuke Kawakami, Yasumitsu Kondoh, Takeshi Shimizu, Siro Simizu, Bangzhong Lin, Kazuto Nunomura, Kyota Aoyagi, Mitsuyo Ohno, Masaki Ohmuraya, Takashi Sato, Shin ichiro Yoshimura, Ken Sato, Reiko Harada, Yoon-Jeong Kim, Hiroyuki Osada, Tomomi Nemoto, Haruo Kasai, Tadahiro Kitamura, Shinya Nagamatsu, and Akihiro Harada. Opposing roles for SNAP23 in secretion in exocrine and endocrine pancreatic cells. Journal of Cell Biology, 215 (1):121–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/215/1/121.

Kurokawa:2019:VSC

[KOK^{+}19] Kazuo Kurokawa, Hiroko Osakada, Tomoko Kojidani, Miho Waga, Yasuyuki Suda, Haruhiko Asakawa, Tokuko Haraguchi, and Akihiko Nakano. Visualization of secretory cargo transport within the Golgi apparatus. Journal of Cell Biology, 218 (5):1602–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/5/1602.

Kononenko:2017:LCK

[Kon17] Natalia L. Kononenko. Lysosomes convene to keep the synapse clean. Journal of Cell Biology, 216(8):2251–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/8/
2251.

Kumar:2019:IAI

[KOR^{+}19] Mukesh Kumar, Srikant Ojha, Priyanka Rai, Alaumy Joshi, Siddhesh S. Kamat, and Roop Mallik. Insulin activates intracellular transport of lipid droplets to release triglycerides from the liver. Journal of Cell Biology, 218(11):3697–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/11/3697.

Kumar:2016:CTT

[KOV^{+}16a] Abhishek Kumar, Mingxing Ouyang, Koen Van den Dries, Ewan James McGhee, Keiichiro Tanaka, Marie D. Anderson, Alexander Groisman, Benjamin T. Goult, Kurt I. Anderson, and Martin A. Schwartz. Correction: Talin tension sensor reveals novel features of focal adhesion force transmission and mechanosensitivity. Journal of Cell Biology, 214(2):231–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
214/2/231.

Kumar:2016:TTS

[KOV^{+}16b] Abhishek Kumar, Mingxing Ouyang, Koen Van den Dries, Ewan James McGhee, Keiichiro Tanaka, Marie D. Anderson, Alexander Groisman, Benjamin T. Goult, Kurt I. Anderson, and Martin A. Schwartz. Talin tension sensor reveals novel features of focal adhesion force transmission and mechanosensitivity. Journal of Cell Biology, 213(3):371–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/213/3/371.

Khong:2018:MAT

[KP18] Anthony Khong and Roy Parker. mRNP architecture in translating and stress conditions reveals an ordered pathway of mRNP compaction. Journal of Cell Biology, 217(12):4124– ??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/12/4124.

Kedersha:2016:GCU

[KPA^{+}16] Nancy Kedersha, Marc D. Panas, Christopher A. Achorn, Shawn Lyons, Sarah Tisdale, Tyler Hickman, Marshall Thomas, Judy Lieberman, Gerald M. McInerney, Pavel Ivanov, and Paul Anderson. G3BP–caprin1–USP10 complexes mediate stress granule condensation and associate with 40S subunits. Journal of Cell Biology, 212(7):845–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/212/7/845. See correc-

tion [KPA^{+}20].

Kedersha:2020:CGC

[KPA^{+}20] Nancy Kedersha, Marc D. Panas, Christopher A. Achorn, Shawn Lyons, Sarah Tisdale, Tyler Hickman, Marshall Thomas, Judy Lieberman, Gerald M. McInerney, Pavel Ivanov, and Paul Anderson. Correction: G3BP Caprin1 USP10 complexes mediate stress granule condensation and associate with 40S subunits. Journal of Cell Biology, 219(1):e20150802809202019c, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). See [KPA^{+}16].

Koehler:2017:PPR

[KPEJ17] Christopher L. Koehler, Guy A. Perkins, Mark H. Ellisman, and D. Leanne Jones. Pink1 and Parkin regulate Drosophila intestinal stem cell proliferation during stress and aging. Journal of Cell Biology, 216(8):2315–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/8/2315.

Kuhn:2019:CTN

[KPGG^{+}19] Terra M. Kuhn, Pau Pascual-Garcia, Alejandro Gozalo, Shawn C. Little, and Maya Capelson. Chromatin targeting of nuclear pore proteins induces chromatin decondensation. Journal of Cell Biology, 218(9):2945–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/9/2945.

Krueger:2019:CLM

[KQM^{+}19] Daniel Krueger, Theresa Quinkler, Simon Arnold Mortensen, Carsten Sachse, and Stefano De Renzis. Cross-linker–mediated regulation of actin network organization controls tissue morphogenesis. Journal of Cell Biology, 218(8):2743–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/8/
2743.

Kim:2018:RNP

[KR18] Hyungsoo Kim and Ze’ev A. Ronai. Rewired Notch/p53 by Numb’ing Mdm2. Journal of Cell Biology, 217(2):445– ??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/2/445.

Keren:2017:BMT

[KS17] Kinneret Keren and Tom Shemesh. Buckle up: Membrane tension drives lamellipodial network compression and adhesion deposition. Journal of Cell Biology, 216(9):2619–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/9/
2619.

Knust:2019:SEP

[KS19] Elisabeth Knust and Kai Simons. Suzanne Eaton (1959–2019): a pioneer in quantitative tissue morphogenesis. Journal of Cell Biology, 218(9):2819–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/9/2819.

Kenific:2016:NEA

[KSG^{+}16] Candia M. Kenific, Samantha J. Stehbens, Juliet Goldsmith, Andrew M. Leidal, Nathalie Faure, Jordan Ye, Torsten Wittmann, and Jayanta Debnath. NBR1 enables autophagydependent focal adhesion turnover. Journal of Cell Biology, 212 (5):577–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/5/577.

Klose:2019:SDM

[KSGL19] Matthias Klose, Johann E. Salloum, Hannes Gonschior, and Stefan Linder. SNX3 drives maturation of Borrelia phagosomes by forming a hub for PI(3)P, Rab5a, and galectin-9. Journal of Cell Biology, 218(9):3039–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/9/3039.

Kadlecova:2017:RCM

[KSL^{+}17] Zuzana Kadlecova, Stephanie J. Spielman, Dinah Loerke, Aparna Mohanakrishnan, Dana Kim Reed, and Sandra L. Schmid. Regulation of clathrin-mediated endocytosis by hierarchical allosteric activation of AP2. Journal of Cell Biology, 216(1):167–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/1/167.

Kinoshita:2017:RBS

[KSM^{+}17] Masanao Kinoshita, Kenichi G. N. Suzuki, Nobuaki Matsumori, Misa Takada, Hikaru Ano, Kenichi Morigaki, Mitsuhiro Abe, Asami Makino, Toshihide Kobayashi, Koichiro M. Hirosawa, Takahiro K. Fujiwara, Akihiro Kusumi, and Michio Murata. Raft-based sphingomyelin interactions revealed by new fluorescent sphingomyelin analogs. Journal of Cell Biology, 216 (4):1183–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/4/1183.

Kovacevic:2018:CRK

[KSM^{+}18] Igor Kovačević, Tomohisa Sakaue, Jisca Majoleé, Manon C. Pronk, Masashi Maekawa, Dirk Geerts, Mar Fernandez-Borja, Shigeki Higashiyama, and Peter L. Hordijk. The Cullin-3– Rbx1–KCTD10 complex controls endothelial barrier function via K63 ubiquitination of RhoB. Journal of Cell Biology, 217 (3):1015–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/3/1015.

Kuri:2017:CDV

[KST^{+}17a] Paola Kuri, Nicole L. Schieber, Thomas Thumberger, Joachim Wittbrodt, Yannick Schwab, and Maria Leptin. Correction: Dynamics of in vivo ASC speck formation. Journal of Cell Biology, 216(10):3423–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/216/10/3423. See [KST^{+}17a].

Kuri:2017:DVA

[KST^{+}17b] Paola Kuri, Nicole L. Schieber, Thomas Thumberger, Joachim Wittbrodt, Yannick Schwab, and Maria Leptin. Dynamics of in vivo ASC speck formation. Journal of Cell Biology, 216(9): 2891–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress. org/content/216/9/2891. See correction [KST^{+}17a].

Kumar:2019:FMI

[KST^{+}19] Abhishek Kumar, Maria S. Shutova, Keiichiro Tanaka, Daniel V. Iwamoto, David A. Calderwood, Tatyana M. Svitkina, and Martin A. Schwartz. Filamin A mediates isotropic distribution of applied force across the actin network. Journal of Cell Biology, 218(8):2481–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/8/2481.

Kannan:2015:CSC

[KT15a] Nivetha Kannan and Vivian W. Tang. Correction: Synaptopodin couples epithelial contractility to α-actinin-4-dependent junction maturation. Journal of Cell Biology, 211(4):933– ??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/4/933.

Kannan:2015:SCE

[KT15b] Nivetha Kannan and Vivian W. Tang. Synaptopodin couples epithelial contractility to α-actinin-4-dependent junction maturation. Journal of Cell Biology, 211(2):407–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/211/2/407.

Ktistakis:2019:WPF

[Kti19] Nicholas T. Ktistakis. Who plays the ferryman: ATG2 channels lipids into the forming autophagosome. Journal of Cell Biology, 218(6):1767–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/6/1767.

Kawano:2018:SFI

[KTK^{+}18] Shin Kawano, Yasushi Tamura, Rieko Kojima, Siqin Bala, Eri Asai, Agnès H. Michel, Benoît Kornmann, Isabelle Riezman, Howard Riezman, Yoshitake Sakae, Yuko Okamoto, and Toshiya Endo. Structure–function insights into direct lipid transfer between membranes by Mmm1–Mdm12 of ERMES. Journal of Cell Biology, 217(3):959–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/3/959.

Ko:2019:MPI

[KTM19] Clint S. Ko, Vardges Tserunyan, and Adam C. Martin. Microtubules promote intercellular contractile force transmission during tissue folding. Journal of Cell Biology, 218(8):2726–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/8/2726.

Kang:2017:BTG

[KVK^{+}17] Min-Ji Kang, Deepika Vasudevan, Kwonyoon Kang, Kyunggon Kim, Jung-Eun Park, Nan Zhang, Xiaomei Zeng, Thomas A. Neubert, Michael T. Marr, and Hyung Don Ryoo. 4e-BP is a target of the GCN2–ATF4 pathway during Drosophila development and aging. Journal of Cell Biology, 216(1):115–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/1/115.

Klare:2015:CCB

[KWB^{+}15] Kerstin Klare, John R. Weir, Federica Basilico, Tomasz Zimniak, Lucia Massimiliano, Nina Ludwigs, Franz Herzog, and Andrea Musacchio. CENP-c is a blueprint for constitutive centromere–associated network assembly within human kinetochores. Journal of Cell Biology, 210(1):11–??, July 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/210/1/11.

Kim:2015:MAM

[KY15] Soonjoung Kim and Hongtao Yu. Multiple assembly mechanisms anchor the KMN spindle checkpoint platform at human mitotic kinetochores. Journal of Cell Biology, 208(2):181–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
208/2/181.

Kelliher:2018:AKE

[KYN^{+}18] Michael T. Kelliher, Yang Yue, Ashley Ng, Daichi Kamiyama, Bo Huang, Kristen J. Verhey, and Jill Wildonger. Autoinhibition of kinesin-1 is essential to the dendrite-specific localization of Golgi outposts. Journal of Cell Biology, 217(7):2531–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/7/
2531.

Kim:2018:CSR

[KZW^{+}18] Sun K. Kim, Siwei Zhang, Michael E. Werner, Eva J. Brotslaw, Jennifer W. Mitchell, Mohamed M. Altabbaa, and Brian J. Mitchell. CLAMP/ spef1 regulates planar cell polarity signaling and asymmetric microtubule accumulation in the Xenopus ciliated epithelia. Journal of Cell Biology, 217(5):1633–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/5/
1633.

Lacefield:2019:DTR

[Lac19] Soni Lacefield. Detaching the tether: Remodeling mitochondrial localization during meiosis. Journal of Cell Biology, 218 (2):389–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/2/389.

Labrousse-Arias:2017:VPI

[LAMACE^{+}17] David Labrousse-Arias, Emma Martínez-Alonso, María Corral-Escariz, Raquel Bienes-Martínez, Jaime Berridy, Leticia Serrano-Oviedo, Elisa Conde, María-Laura García-Bermejo, José M. Giménez-Bachs, Antonio S. Salinas-Sánchez, Ricardo Sánchez-Prieto, Masahiro Yao, Marina Lasa, and María J. Calzada. VHL promotes immune response against renal cell carcinoma via NF-κB-dependent regulation of VCAM-1. Journal of Cell Biology, 216(3):835–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/3/835.

Logsdon:2015:BTC

[LBB^{+}15] Glennis A. Logsdon, Evelyne J. Barrey, Emily A. Bassett, Jamie E. DeNizio, Lucie Y. Guo, Tanya Panchenko, Jennine M. Dawicki-McKenna, Patrick Heun, and Ben E. Black. Both tails and the centromere targeting domain of CENP–A are required for centromere establishment. Journal of Cell Biology, 208 (5):521–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/208/5/521.

Loncarek:2018:BRC

[LBD18] Jadranka Loncarek and Mónica Bettencourt-Dias. Building the right centriole for each cell type. Journal of Cell Biology, 217 (3):823–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/3/823.

Lerner:2017:MTR

[LBG^{+}17] Thomas R. Lerner, Sophie Borel, Daniel J. Greenwood, Urska Repnik, Matthew R. G. Russell, Susanne Herbst, Martin L. Jones, Lucy M. Collinson, Gareth Griffiths, and Maximiliano G. Gutierrez. Mycobacterium tuberculosis replicates within necrotic human macrophages. Journal of Cell Biology, 216(3):583–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/3/583.

Lock:2019:CCA

[LBJ^{+}19] John G. Lock, Francesco Baschieri, Matthew C. Jones, Jonathan D. Humphries, Guillaume Montagnac, Staffan Strömblad, and Martin J. Humphries. Clathrin-containing adhesion complexes. Journal of Cell Biology, 218(7):2086–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/7/
2086.

Lopes:2017:MNF

[LBV^{+}17] Filipa B. Lopes, Štefan Bálint, Salvatore Valvo, James H. Felce, Edith M. Hessel, Michael L. Dustin, and Daniel M. Davis. Membrane nanoclusters of FcγRI segregate from inhibitory SIRPα upon activation of human macrophages. Journal of Cell Biology, 216(4):1123–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/4/1123.

Liu:2017:TPC

[LCD^{+}17] Yi Liu, José Renato Cussiol, Diego Dibitetto, Jennie Rae Sims, Shyam Twayana, Robert Samuel Weiss, Raimundo Freire, Federica Marini, Achille Pellicioli, and Marcus Bustamante Smolka. TOPBP1^{Dpb11} plays a conserved role in homologous recombination DNA repair through the coordinated recruitment of 53BP1^{Rad9}. Journal of Cell Biology, 216(3):623–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/3/
623.

Le:2016:LMR

[LCM^{+}16] Lily Thao-Nhi Le, Oscar Cazares, Janna K. Mouw, Sharmila Chatterjee, Hector Macias, Angel Moran, Jillian Ramos, Patricia J. Keely, Valerie M. Weaver, and Lindsay Hinck. Loss of miR-203 regulates cell shape and matrix adhesion through ROBO1/Rac/FAK in response to stiffness. Journal of Cell Biology, 212(6):707–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/6/707.

Laporte:2015:SMA

[LCP^{+}15] Damien Laporte, Fabien Courtout, Benoît Pinson, Jim Dompierre, Bénédicte Salin, Lysiane Brocard, and Isabelle Sagot. A stable microtubule array drives fission yeast polarity reestablishment upon quiescence exit. Journal of Cell Biology, 210 (1):99–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/1/99.

Liu:2017:IGT

[LCTP17] Li-Ka Liu, Vineet Choudhary, Alexandre Toulmay, and William A. Prinz. An inducible ER–Golgi tether facilitates ceramide transport to alleviate lipotoxicity. Journal of Cell Biology, 216(1):131–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/1/131.

Li:2016:LMP

[LCZ^{+}16] Yuan Li, Baohui Chen, Wei Zou, Xin Wang, Yanwei Wu, Dongfeng Zhao, Yanan Sun, Yubing Liu, Lianwan Chen, Long Miao, Chonglin Yang, and Xiaochen Wang. The lysosomal membrane protein SCAV-3 maintains lysosome integrity and adult longevity. Journal of Cell Biology, 215(2):167–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/215/2/
167.

Lucena:2015:NTM

[LDG^{+}15] Rafael Lucena, Noah Dephoure, Steve P. Gygi, Douglas R. Kellogg, Victor A. Tallada, Rafael R. Daga, and Juan Jimenez. Nucleocytoplasmic transport in the midzone membrane domain controls yeast mitotic spindle disassembly. Journal of Cell Biology, 209(3):387–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/3/387.

Lele:2018:MPN

[LDG18] Tanmay P. Lele, Richard B. Dickinson, and Gregg G. Gundersen. Mechanical principles of nuclear shaping and positioning. Journal of Cell Biology, 217(10):3330–??, October 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/10/3330.

Ladouceur:2015:MCL

[LDM15] Anne-Marie Ladouceur, Jonas F. Dorn, and Paul S. Maddox. Mitotic chromosome length scales in response to both cell and nuclear size. Journal of Cell Biology, 209(5):645–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/209/5/645.

Lazaro-Dieguez:2017:CCA

[LDM17] Francisco Lázaro-Diéguez and Anne Müsch. Cell–cell adhesion accounts for the different orientation of columnar and hepatocytic cell divisions. Journal of Cell Biology, 216(11): 3847–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/11/3847.

Lou:2015:MLC

[LDMW^{+}15] Sunny S. Lou, Alba Diz-Muñoz, Orion D. Weiner, Daniel A. Fletcher, and Julie A. Theriot. Myosin light chain kinase regulates cell polarization independently of membrane tension or Rho kinase. Journal of Cell Biology, 209(2):275–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/209/2/275.

Luu:2015:PMS

[LDP^{+}15] Olivia Luu, Erich W. Damm, Serge E. Parent, Debanjan Barua, Tamara H. L. Smith, Jason W. H. Wen, Stephanie E. Lepage, Martina Nagel, Hady Ibrahim-Gawel, Yunyun Huang, Ashley E. E. Bruce, and Rudolf Winklbauer. PAPC mediates self/ non-self-distinction during Snail1-dependent tissue separation. Journal of Cell Biology, 208(6):839–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/6/839.

Liu:2019:CCB

[LDR^{+}19] Lei Liu, Li Ding, Matteo Rovere, Michael S. Wolfe, and Dennis J. Selkoe. A cellular complex of BACE1 and γ-secretase sequentially generates Aβ from its full-length precursor. Journal of Cell Biology, 218(2):644–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/2/644.

Lambrus:2016:UPP

[LDU^{+}16] Bramwell G. Lambrus, Vikas Daggubati, Yumi Uetake, Phillip M. Scott, Kevin M. Clutario, Greenfield Sluder, and Andrew J. Holland. A USP28–53BP1–p53–p21 signaling axis arrests growth after centrosome loss or prolonged mitosis. Journal of Cell Biology, 214(2):143–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/214/2/143.

Lloyd-Evans:2016:MLC

[LE16] Emyr Lloyd-Evans. On the move, lysosomal CAX drives Ca^{2+} transport and motility. Journal of Cell Biology, 212(7):755–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
212/7/755.

Leduc:2017:RMA

[LEM17] Cécile Leduc and Sandrine Etienne-Manneville. Regulation of microtubule-associated motors drives intermediate filament network polarization. Journal of Cell Biology, 216(6):1689–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/6/1689.

Leslie:2015:AGL

[Les15a] Mitch Leslie. Actin is good at long division. Journal of Cell Biology, 208(1):??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/1/2.2.

Leslie:2015:BMB

[Les15b] Mitch Leslie. Balanced mitochondria behave better. Journal of Cell Biology, 211(4):??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/211/4/718.1.

Leslie:2015:BRM

[Les15c] Mitch Leslie. Bleach at the roots of mitotic progression. Journal of Cell Biology, 210(1):??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/1/2.1.

Leslie:2015:BSG

[Les15d] Mitch Leslie. BRWD1 shows its gender bias. Journal of Cell Biology, 208(1):3–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/1/3.

Leslie:2015:CDA

[Les15e] Mitch Leslie. Cilia drop anchor. Journal of Cell Biology, 210(1):3–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/1/3.

Leslie:2015:DTG

[Les15f] Mitch Leslie. Desmoplakin’s tail gets the message. Journal of Cell Biology, 208(5):??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/5/494.2.

Leslie:2015:EEC

[Les15g] Mitch Leslie. Erv41–Erv46 complex recaptures wayward ER proteins. Journal of Cell Biology, 208(2):??, January 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/208/2/138.3.

Leslie:2015:EKR

[Les15h] Mitch Leslie. Even killers recycle. Journal of Cell Biology, 210(1):??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/1/2.3.

Leslie:2015:FSS

[Les15i] Mitch Leslie. A fresh start for stalled forks. Journal of Cell Biology, 208(5):495–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/5/495.

Leslie:2015:HIG

[Les15j] Mitch Leslie. A Hedgehog inhibitor gets around. Journal of Cell Biology, 209(5):??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/5/624.2.

Leslie:2015:HCS

[Les15k] Mitch Leslie. How chromosomes shrink to fit. Journal of Cell Biology, 209(5):625–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/5/625.

Leslie:2015:HOT

[Les15l] Mitch Leslie. How the optic tectum stacks up. Journal of Cell Biology, 211(4):719–??, November 2015. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/211/4/719.

Leslie:2015:HTT

[Les15m] Mitch Leslie. Hsc70’s trip to the tip. Journal of Cell Biology, 210(5):??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/5/682.3.

Leslie:2015:KDL

[Les15n] Mitch Leslie. KDM4B doesn’t leave a mark. Journal of Cell Biology, 211(4):??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/4/718.2.

Leslie:2015:MPS

[Les15o] Mitch Leslie. Mannose 6-phosphate steers B cells on the right course. Journal of Cell Biology, 208(2):??, January 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/208/2/138.2.

Leslie:2015:MCG

[Les15p] Mitch Leslie. Mast cells get all touchy-feely. Journal of Cell Biology, 210(5):683–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/5/683.

Leslie:2015:MPT

[Les15q] Mitch Leslie. Mitotic proteins take on editorial duties. Journal of Cell Biology, 209(2):??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/2/184.2.

Leslie:2015:MSC

[Les15r] Mitch Leslie. MLCK stops cells from going full frontal. Journal of Cell Biology, 209(2):??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/2/184.3.

Leslie:2015:NJF

[Les15s] Mitch Leslie. Neurons join forces to push ahead. Journal of Cell Biology, 209(5):??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/5/624.3.

Leslie:2015:NFS

[Les15t] Mitch Leslie. NIK flips the switch on Arp2/3. Journal of Cell Biology, 208(2):??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/2/138.1.

Leslie:2015:NSI

[Les15u] Mitch Leslie. No substitute for β3 integrin. Journal of Cell Biology, 208(1):??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/1/2.3.

Leslie:2015:PCC

[Les15v] Mitch Leslie. p53 censuses centrosomes. Journal of Cell Biology, 210(1):??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/1/2.2.

Leslie:2015:PPA

[Les15w] Mitch Leslie. Phospholipid prunes actin branches. Journal of Cell Biology, 210(5):??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/5/682.2.

Leslie:2015:PKF

[Les15x] Mitch Leslie. Putting a kink in a familiar control loop. Journal of Cell Biology, 208(2):139–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/2/139.

Leslie:2015:SWG

[Les15y] Mitch Leslie. Sall4 won’t give stem cells a break. Journal of Cell Biology, 208(5):??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/5/494.1.

Leslie:2015:SRP

[Les15z] Mitch Leslie. Splicing reverses protein’s function. Journal of Cell Biology, 209(5):??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/5/624.1.

Leslie:2015:STE

[Les15-27] Mitch Leslie. Starving tumors eat metabolic enzyme. Journal of Cell Biology, 210(5):??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/5/682.1.

Leslie:2015:SGE

[Les15-28] Mitch Leslie. Stress granules ease the way for metastasis. Journal of Cell Biology, 208(7):??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/208/7/860.2.

Leslie:2015:SSF

[Les15-29] Mitch Leslie. Sugar saps FOXO1’s healing power. Journal of Cell Biology, 209(2):185–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/209/2/185.

Leslie:2015:TPC

[Les15-30] Mitch Leslie. TCAF1 puts a chill on prostate cancer. Journal of Cell Biology, 208(1):??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/208/1/2.1.

Leslie:2015:TCD

[Les15-31] Mitch Leslie. TDP-43 crosses the divide. Journal of Cell Biology, 211(4):??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/4/718.3.

Leslie:2015:TMU

[Les15-32] Mitch Leslie. TRAMM makes an unscheduled stop during mitosis. Journal of Cell Biology, 209(2):??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/209/2/184.1.

Leslie:2016:AAC

[Les16a] Mitch Leslie. All-access coverage for neurites showcases their subtleties. Journal of Cell Biology, 212(1):??, January 2016.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/212/1/2.2.

Leslie:2016:CKR

[Les16b] Mitch Leslie. Cdc55 keeps Rho1 focused on growth. Journal of Cell Biology, 212(1):??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/1/2.1.

Leslie:2016:EPT

[Les16c] Mitch Leslie. Epithelia prefer their ECM firm and easily digested. Journal of Cell Biology, 212(1):??, January 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/212/1/2.3.

Leslie:2016:MPB

[Les16d] Mitch Leslie. ER membrane puts up barriers in C. elegans. Journal of Cell Biology, 214(6):637–??, September 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/214/6/637.

Leslie:2016:LSO

[Les16e] Mitch Leslie. The long and short of OPA1. Journal of Cell Biology, 212(2):??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/2/126.1.

Leslie:2016:MSL

[Les16f] Mitch Leslie. Making sure late endosomes are on time. Journal of Cell Biology, 212(2):127–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/212/2/127.

Leslie:2016:TSR

[Les16g] Mitch Leslie. Taking the stress out of replication. Journal of Cell Biology, 212(1):3–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/1/3.

Leslie:2016:ULN

[Les16h] Mitch Leslie. Ubiquitylation leaves Nup60 a basket case. Journal of Cell Biology, 212(2):??, January 2016. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/2/126.2.

Leslie:2016:USS

[Les16i] Mitch Leslie. Using speckles to spot protein dynamics. Journal of Cell Biology, 212(2):??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/2/126.3.

Leslie:2016:WER

[Les16j] Mitch Leslie. Why endosomes recycle GPCRs. Journal of Cell Biology, 214(7):785–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/214/7/785.

Leslie:2017:IKP

[Les17] Nick R. Leslie. Importin-11 keeps PTEN safe from harm. Journal of Cell Biology, 216(3):539–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/3/539.

Lehtimaki:2017:UPM

[LFK^{+}17a] Jaakko I. Lehtimäki, Aidan M. Fenix, Tommi M. Kotila, Giuseppe Balistreri, Lassi Paavolainen, Markku Varjosalo, Dylan T. Burnette, and Pekka Lappalainen. UNC-45a promotes myosin folding and stress fiber assembly. Journal of Cell Biology, 216(12):4053–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/12/4053.

Liu:2017:TSO

[LFK^{+}17b] Min Liu, Zhi Feng, Hongmei Ke, Ying Liu, Tianhui Sun, Jianli Dai, Wenhong Cui, and José Carlos Pastor-Pareja. Tango1 spatially organizes ER exit sites to control ER export. Journal of Cell Biology, 216(4):1035–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/4/1035.

Lee:2016:GPP

[LFT^{+}16] Gun-Hee Lee, Morihisa Fujita, Katsuyoshi Takaoka, Yoshiko Murakami, Yoshitaka Fujihara, Noriyuki Kanzawa, Kei ichi Murakami, Eriko Kajikawa, Yoko Takada, Kazunobu Saito, Masahito Ikawa, Hiroshi Hamada, Yusuke Maeda, and Taroh Kinoshita. A GPI processing phospholipase A2, PGAP6, modulates Nodal signaling in embryos by shedding CRIPTO. Journal of Cell Biology, 215(5):705–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/215/5/705.

Liu:2018:GDP

[LGH^{+}18] Yi-Shi Liu, Xin-Yu Guo, Tetsuya Hirata, Yao Rong, Daisuke Motooka, Toshihiko Kitajima, Yoshiko Murakami, Xiao-Dong Gao, Shota Nakamura, Taroh Kinoshita, and Morihisa Fujita. N -Glycan-dependent protein folding and endoplasmic reticulum retention regulate GPI-anchor processing. Journal of Cell Biology, 217(2):585–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/2/585.

Lee:2018:RMN

[LgYL^{+}18] Gi Young Lee, Deok gyun You, Hye-Ra Lee, Sun Wook Hwang, C. Justin Lee, and Young Do Yoo. Romo1 is a mitochondrial nonselective cation channel with viroporin-like characteristics. Journal of Cell Biology, 217(6):2059–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/6/2059.

Lamech:2015:UPA

[LH15] Lilian T. Lamech and Cole M. Haynes. The unpredictability of prolonged activation of stress response pathways. Journal of Cell Biology, 209(6):781–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/209/6/781.

Leandro:2019:SLD

[LH19] João Leandro and Sander M. Houten. Saccharopine, a lysine degradation intermediate, is a mitochondrial toxin. Journal of Cell Biology, 218(2):391–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/2/391.

Leung:2015:SCA

[LHA^{+}15] Wing-Kit Leung, Neil Humphryes, Negar Afshar, Bilge Argunhan, Yaroslav Terentyev, Tomomi Tsubouchi, and Hideo Tsubouchi. The synaptonemal complex is assembled by a polySUMOylation-driven feedback mechanism in yeast. Journal of Cell Biology, 211(4):785–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/211/4/785.

Llauro:2018:KKP

[LHB^{+}18] Aida Llauró, Hanako Hayashi, Megan E. Bailey, Alex Wilson, Patryk Ludzia, Charles L. Asbury, and Bungo Akiyoshi. The kinetoplastid kinetochore protein KKT4 is an unconventional microtubule tip–coupling protein. Journal of Cell Biology, 217(11):3886–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/11/3886.

Lee:2019:MTP

[LHT^{+}19] Sora Lee, Hsuan-Chung Ho, Jessica M. Tumolo, Pi-Chiang Hsu, and Jason A. MacGurn. Methionine triggers Ppz-mediated dephosphorylation of Art1 to promote cargo-specific endocytosis. Journal of Cell Biology, 218(3):977–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/3/977.

Lee:2019:DRG

[LHY^{+}19] Moonsup Lee, Yoo-Seok Hwang, Jaeho Yoon, Jian Sun, Adam Harned, Kunio Nagashima, and Ira O. Daar. Developmentally regulated GTP-binding protein 1 modulates ciliogenesis via an interaction with Dishevelled. Journal of Cell Biology, 218(8): 2659–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/8/2659.

Linder:2015:MME

[Lin15] Stefan Linder. MT1-MMP: Endosomal delivery drives breast cancer metastasis. Journal of Cell Biology, 211(2):215–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
211/2/215.

Licht:2016:RDT

[LJ16] Konstantin Licht and Michael F. Jantsch. Rapid and dynamic transcriptome regulation by RNA editing and RNA modifications. Journal of Cell Biology, 213(1):15–??, April 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/213/1/15.

Lane:2017:CBA

[LJ17a] Simon I. R. Lane and Keith T. Jones. Chromosome biorientation and APC activity remain uncoupled in oocytes with reduced volume. Journal of Cell Biology, 216(12):3949–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/12/3949.

Li:2017:HCH

[LJ17b] Yanxin Li and Jianwei Jiao. Histone chaperone HIRA regulates neural progenitor cell proliferation and neurogenesis via β-catenin. Journal of Cell Biology, 216(7):1975–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/7/
1975.

Lerit:2015:ICO

[LJP^{+}15] Dorothy A. Lerit, Holly A. Jordan, John S. Poulton, Carey J. Fagerstrom, Brian J. Galletta, Mark Peifer, and Nasser M. Rusan. Interphase centrosome organization by the PLP–Cnn scaffold is required for centrosome function. Journal of Cell Biology, 210(1):79–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/1/79.

Liu:2016:CNR

[LJS^{+}16a] Kai Liu, Youli Jian, Xiaojuan Sun, Chengkui Yang, Zhiyang Gao, Zhili Zhang, Xuezhao Liu, Yang Li, Jing Xu, Yudong Jing, Shohei Mitani, Sudan He, and Chonglin Yang. Correction: Negative regulation of phosphatidylinositol 3-phosphate levels in early-to-late endosome conversion. Journal of Cell Biology, 212(6):739–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/6/739.

Liu:2016:NRP

[LJS^{+}16b] Kai Liu, Youli Jian, Xiaojuan Sun, Chengkui Yang, Zhiyang Gao, Zhili Zhang, Xuezhao Liu, Yang Li, Jing Xu, Yudong Jing, Shohei Mitani, Sudan He, and Chonglin Yang. Negative regulation of phosphatidylinositol 3-phosphate levels in early-to-late endosome conversion. Journal of Cell Biology, 212(2):181–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
212/2/181.

Leitao:2017:DMD

[LK17] Ricardo M. Leitao and Douglas R. Kellogg. The duration of mitosis and daughter cell size are modulated by nutrients in budding yeast. Journal of Cell Biology, 216(11):3463–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/11/3463.

Li:2015:MAU

[LKE15] Ming Li, Tatsuhiro Koshi, and Scott D. Emr. Membraneanchored ubiquitin ligase complex is required for the turnover of lysosomal membrane proteins. Journal of Cell Biology, 211(3): 639–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/3/639.

Lee:2015:SSA

[LKM^{+}15a] Miriam Lee, Young-Joon Ko, Yeojin Moon, Minsoo Han, Hyung-Wook Kim, Sung Haeng Lee, KyeongJin Kang, and Youngsoo Jun. SNAREs support atlastin-mediated homotypic ER fusion in Saccharomyces cerevisiae. Journal of Cell Biology, 210(3):451–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/3/451.

Li:2015:ESN

[LKM^{+}15b] Haiying Li, Yeon Koo, Xicheng Mao, Luis Sifuentes-Dominguez, Lindsey L. Morris, Da Jia, Naoteru Miyata, Rebecca A. Faulkner, Jan M. van Deursen, Marc Vooijs, Daniel D. Billadeau, Bart van de Sluis, Ondine Cleaver, and Ezra Burstein. Endosomal sorting of Notch receptors through COMMD9dependent pathways modulates Notch signaling. Journal of Cell Biology, 211(3):605–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/211/3/605.

LeBlanc:2017:DSC

[LL17] Michelle G. LeBlanc and Ruth Lehmann. Domain-specific control of germ cell polarity and migration by multifunction Tre1 GPCR. Journal of Cell Biology, 216(9):2945–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/9/
2945.

Lemmens:2019:DRM

[LL19] Bennie Lemmens and Arne Lindqvist. DNA replication and mitotic entry: a brake model for cell cycle progression. Journal of Cell Biology, 218(12):3892–3902, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

https://rupress.org/jcb/article/218/12/3892/132524/
DNA-replication-and-mitotic-entry-A-brake-model.

Liang:2018:ARE

[LLAC18a] Jin Rui Liang, Emily Lingeman, Saba Ahmed, and Jacob E. Corn. Atlastins remodel the endoplasmic reticulum for selective autophagy. Journal of Cell Biology, 217(10):3354–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/

217/10/3354. See correction [LLAC18b].

Liang:2018:CAR

[LLAC18b] Jin Rui Liang, Emily Lingeman, Saba Ahmed, and Jacob E. Corn. Correction: Atlastins remodel the endoplasmic reticulum for selective autophagy. Journal of Cell Biology, 217(11): 4049–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/11/4049. See [LLAC18a].

Longmate:2017:SIE

[LLC^{+}17] Whitney M. Longmate, Scott P. Lyons, Sridar V. Chittur, Kevin M. Pumiglia, Livingston Van De Water, and C. Michael DiPersio. Suppression of integrin α3β1 by α9β1 in the epidermis controls the paracrine resolution of wound angiogenesis. Journal of Cell Biology, 216(5):1473–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/5/1473.

Lai:2017:HSE

[LLK^{+}17] Chun-Ming Lai, Kun-Yang Lin, Shih-Han Kao, Yi-Ning Chen, Fu Huang, and Hwei-Jan Hsu. Hedgehog signaling establishes precursors for germline stem cell niches by regulating cell adhesion. Journal of Cell Biology, 216(5):1439–??, May 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/5/1439.

Lin:2015:LRD

[LLL^{+}15] Chin-Hsien Lin, Hsun Li, Yi-Nan Lee, Ying-Ju Cheng, Ruey-Meei Wu, and Cheng-Ting Chien. Lrrk regulates the dynamic profile of dendritic Golgi outposts through the golgin Lava lamp. Journal of Cell Biology, 210(3):471–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/210/3/471.

Liu:2018:UPK

[LLL^{+}18] Jinchao Liu, Meijiao Li, Lin Li, She Chen, and Xiaochen Wang. Ubiquitination of the PI3-kinase VPS-34 promotes VPS-34 stability and phagosome maturation. Journal of Cell Biology, 217 (1):347–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/1/347.

Le:2016:THR

[LLS^{+}16] Hai H. Le, Monika Looney, Benjamin Strauss, Michael Bloodgood, and Antony M. Jose. Tissue homogeneity requires inhibition of unequal gene silencing during development. Journal of Cell Biology, 214(3):319–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/214/3/319.

Lu:2018:OFC

[LLS^{+}18] Wen Lu, Margot Lakonishok, Anna S. Serpinskaya, David Kirchenbüechler, Shuo-Chien Ling, and Vladimir I. Gelfand. Ooplasmic flow cooperates with transport and anchorage in Drosophila oocyte posterior determination. Journal of Cell Biology, 217(10):3497–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/10/3497.

Lim:2015:CPH

[LLW^{+}15] Jung Mi Lim, Kyung S. Lee, Hyun Ae Woo, Dongmin Kang, and Sue Goo Rhee. Control of the pericentrosomal H_{2} O_{2} level by peroxiredoxin I is critical for mitotic progression. Journal of Cell Biology, 210(1):23–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/1/23.

Liu:2017:BCP

[LLW^{+}17] Xuezhao Liu, Yang Li, Xin Wang, Ruxiao Xing, Kai Liu, Qiwen Gan, Changyong Tang, Zhiyang Gao, Youli Jian, Shouqing Luo, Weixiang Guo, and Chonglin Yang. The BEACHcontaining protein WDR81 coordinates p62 and LC3C to promote aggrephagy. Journal of Cell Biology, 216(5):1301–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/5/
1301.

Lo:2019:PCT

[LLY^{+}19] Chien-Hui Lo, I-Hsuan Lin, T. Tony Yang, Yen-Chun Huang, Barbara E. Tanos, Po-Chun Chou, Chih-Wei Chang, Yeou-Guang Tsay, Jung-Chi Liao, and Won-Jing Wang. Phosphorylation of CEP83 by TTBK2 is necessary for cilia initiation. Journal of Cell Biology, 218(10):3489–??, October 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/10/3489.

Liu:2019:NTQ

[LLZ^{+}19] Xiaohui Liu, Jiazhou Li, Heyu Zhao, Boyang Liu, Thomas Günther-Pomorski, Shaolin Chen, and Johannes Liesche. Novel tool to quantify cell wall porosity relates wall structure to cell growth and drug uptake. Journal of Cell Biology, 218 (4):1408–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/4/1408.

Lammers:2015:DCA

[LM15] Lindsay G. Lammers and Steven M. Markus. The dynein cortical anchor Num1 activates dynein motility by relieving Pac1/LIS1-mediated inhibition. Journal of Cell Biology, 211 (2):309–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/2/309.

Liu:2016:DFM

[LM16] Chenshu Liu and Yinghui Mao. Diaphanous formin mDia2 regulates CENP–A levels at centromeres. Journal of Cell Biology, 213(4):415–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/4/415.

Li:2019:HAC

[LM19] Jiaxing Li and Kelly R. Monk. Healthy attachments: Cell adhesion molecules collectively control myelin integrity. Journal of Cell Biology, 218(9):2824–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/9/2824.

Lopes:2018:CAA

[LMC^{+}18] Carla A. M. Lopes, Marta Mesquita, Ana Isabel Cunha, Joana Cardoso, Sara Carapeta, Cátia Laranjeira, António E. Pinto, José B. Pereira-Leal, António Dias-Pereira, Mónica Bettencourt-Dias, and Paula Chaves. Centrosome amplification arises before neoplasia and increases upon p53 loss in tumorigenesis. Journal of Cell Biology, 217(7):2353–??, July 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/7/2353.

Lelli:2016:CIM

[LMdM^{+}16] Andrea Lelli, Vincent Michel, Jacques Boutet de Monvel, Matteo Cortese, Montserrat Bosch-Grau, Asadollah Aghaie, Isabelle Perfettini, Typhaine Dupont, Paul Avan, Aziz El-Amraoui, and Christine Petit. Class III myosins shape the auditory hair bundles by limiting microvilli and stereocilia growth. Journal of Cell Biology, 212(2):231–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/212/2/231.

Lobato-Marquez:2016:SRM

[LMM16] Damián Lobato-Márquez and Serge Mostowy. Septins recognize micron-scale membrane curvature. Journal of Cell Biology, 213(1):5–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/1/5.

Leyme:2015:IAT

[LMPG^{+}15] Anthony Leyme, Arthur Marivin, Lorena Perez-Gutierrez, Lien T. Nguyen, and Mikel Garcia-Marcos. Integrins activate trimeric G proteins via the nonreceptor protein GIV/Girdin. Journal of Cell Biology, 210(7):1165–??, September 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/210/7/1165.

Lei:2017:PDE

[LMR^{+}17] Wenliang Lei, Kenneth R. Myers, Yanfang Rui, Siarhei Hladyshau, Denis Tsygankov, and James Q. Zheng. Phosphoinositidedependent enrichment of actin monomers in dendritic spines regulates synapse development and plasticity. Journal of Cell Biology, 216(8):2551–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/8/2551.

Lo:2015:CCS

[LNH^{+}15] Harriet P. Lo, Susan J. Nixon, Thomas E. Hall, Belinda S. Cowling, Charles Ferguson, Garry P. Morgan, Nicole L. Schieber, Manuel A. Fernandez-Rojo, Michele Bastiani, Matthias Floetenmeyer, Nick Martel, Jocelyn Laporte, Paul F. Pilch, and Robert G. Parton. The caveolin–cavin system plays a conserved and critical role in mechanoprotection of skeletal muscle. Journal of Cell Biology, 210(5):833–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/210/5/833.

Lindeboom:2019:CSP

[LNS^{+}19] Jelmer J. Lindeboom, Masayoshi Nakamura, Marco Saltini, Anneke Hibbel, Ankit Walia, Tijs Ketelaar, Anne Mie C. Emons, John C. Sedbrook, Viktor Kirik, Bela M. Mulder, and David W. Ehrhardt. CLASP stabilization of plus ends created by severing promotes microtubule creation and reorientation. Journal of Cell Biology, 218(1):190–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/1/190.

Lee:2015:PDR

[LOG15] ChangHwan Lee, Patricia Occhipinti, and Amy S. Gladfelter. PolyQ-dependent RNA–protein assemblies control symmetry breaking. Journal of Cell Biology, 208(5):533–??, March 2015.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/208/5/533.

Logan:2017:FPF

[Log17] Mary A. Logan. Fragile phagocytes: FMRP positively regulates engulfment activity. Journal of Cell Biology, 216(3):531–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/3/531.

Lovett:2018:BSW

[Lov18] Susan T. Lovett. Between sisters: Watching replicationassociated recombinational DNA repair. Journal of Cell Biology, 217(7):2225–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/7/2225.

Lee:2016:CER

[LPGB16] Zuo Yen Lee, Manoël Prouteau, Monica Gotta, and Yves Barral. Compartmentalization of the endoplasmic reticulum in the early C. elegans embryos. Journal of Cell Biology, 214(6): 665–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/214/6/665.

Losada-Perez:2016:MMC

[LPHH16] Maria Losada-Perez, Neale Harrison, and Alicia Hidalgo. Molecular mechanism of central nervous system repair by the Drosophila NG2 homologue kon-tiki. Journal of Cell Biology, 214(5):587–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/5/587.

Lapetina:2017:YSF

[LPRW17] Diego L. Lapetina, Christopher Ptak, Ulyss K. Roesner, and Richard W. Wozniak. Yeast silencing factor Sir4 and a subset of nucleoporins form a complex distinct from nuclear pore complexes. Journal of Cell Biology, 216(10):3145–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/10/3145.

Lang:2015:MJC

[LPWK15] Alexander B. Lang, Arun T. John Peter, Peter Walter, and Benoît Kornmann. ER-mitochondrial junctions can be bypassed by dominant mutations in the endosomal protein Vps13. Journal of Cell Biology, 210(6):883–??, September 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/210/6/883.

Lawson:2018:RGS

[LR18] Campbell D. Lawson and Anne J. Ridley. Rho GTPase signaling complexes in cell migration and invasion. Journal of Cell Biology, 217(2):447–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/2/447.

LeClaire:2015:NIK

[LRBB15] Lawrence L. LeClaire, Manish Rana, Martin Baumgartner, and Diane L. Barber. The Nck-interacting kinase NIK increases Arp2/3 complex activity by phosphorylating the Arp2 subunit. Journal of Cell Biology, 208(2):161–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/2/161.

Li:2019:SSN

[LRD19] Hailun Li, Alexandra Russo, and Aaron DiAntonio. SIK3 suppresses neuronal hyperexcitability by regulating the glial capacity to buffer K^{+} and water. Journal of Cell Biology, 218(12): 4017–4029, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/218/12/4017/132538/SIK3-suppresses-
neuronal-hyperexcitability-by.

Ladang:2015:EDW

[LRH^{+}15] Aurélie Ladang, Francesca Rapino, Lukas C. Heukamp, Lars Tharun, Kateryna Shostak, Damien Hermand, Sylvain Delaunay, Iva Klevernic, Zheshen Jiang, Nicolas Jacques, Diane Jamart, Valérie Migeot, Alexandra Florin, Serkan Göktuna, Brigitte Malgrange, Owen J. Sansom, Laurent Nguyen, Reinhard Büttner, Pierre Close, and Alain Chariot. Elp3 drives Wnt-dependent tumor initiation and regeneration in the intestine. Journal of Cell Biology, 211(3):??, November 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/211/3/2113OIA257.

Lakoduk:2019:MPA

[LRM^{+}19] Ashley M. Lakoduk, Philippe Roudot, Marcel Mettlen, Heather M. Grossman, Sandra L. Schmid, and Ping-Hung Chen. Mutant p53 amplifies a dynamin-1/APPL1 endosome feedback loop that regulates recycling and migration. Journal of Cell Biology, 218(6):1928–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/6/1928.

Ladouceur:2017:CTI

[LRS^{+}17] A.-M. Ladouceur, Rajesh Ranjan, Lydia Smith, Tanner Fadero, Jennifer Heppert, Bob Goldstein, Amy Shaub Maddox, and Paul S. Maddox. CENP-A and topoisomerase–II antagonistically affect chromosome length. Journal of Cell Biology, 216(9): 2645–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/9/2645.

Liu:2016:TTR

[LS16] Yi Liu and Marcus B. Smolka. TOPBP1 takes RADical command in recombinational DNA repair. Journal of Cell Biology, 212(3):263–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/3/263.

Li:2018:ELD

[LS18] Jing Li and Timothy A. Springer. Energy landscape differences among integrins establish the framework for understanding activation. Journal of Cell Biology, 217(1):397–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/217/1/397.

Li:2015:NTA

[LSJY15] Ping Li, Yize Shao, Hui Jin, and Hong-Guo Yu. Ndj1, a telomere-associated protein, regulates centrosome separation in budding yeast meiosis. Journal of Cell Biology, 209(2):247–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
209/2/247.

Lee:2018:PUR

[LSMG18] Byung Ho Lee, Françoise Schwager, Patrick Meraldi, and Monica Gotta. p37/UBXN2B regulates spindle orientation by limiting cortical NuMA recruitment via PP1/Repo–Man. Journal of Cell Biology, 217(2):483–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/2/483.

Lee:2018:BMW

[LSMZ^{+}18] Juliette J. Lee, Alvaro Sanchez-Martinez, Aitor Martinez Zarate, Cristiane Benincá, Ugo Mayor, Michael J. Clague, and Alexander J. Whitworth. Basal mitophagy is widespread in Drosophila but minimally affected by loss of Pink1 or Parkin. Journal of Cell Biology, 217(5):1613–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/5/1613.

Lepore:2016:CCD

[LSPC16] Dante Lepore, Olya Spassibojko, Gabrielle Pinto, and Ruth N. Collins. Cell cycle-dependent phosphorylation of Sec4p controls membrane deposition during cytokinesis. Journal of Cell Biology, 214(6):691–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/6/691.

Llano:2015:KRA

[LSS^{+}15] Olaya Llano, Sergey Smirnov, Shetal Soni, Andrey Golubtsov, Isabelle Guillemin, Pirta Hotulainen, Igor Medina, Hans Gerd Nothwang, Claudio Rivera, and Anastasia Ludwig. KCC2 regulates actin dynamics in dendritic spines via interaction with β-PIX. Journal of Cell Biology, 209(5):671–??, June 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/209/5/671.

Liu:2018:VTF

[LT18] Zhe Liu and Robert Tjian. Visualizing transcription factor dynamics in living cells. Journal of Cell Biology, 217(4):1181–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/4/1181.

Ladstatter:2019:GIC

[LT19a] Sabrina Ladstätter and Kikuë Tachibana. Genomic insights into chromatin reprogramming to totipotency in embryos. Journal of Cell Biology, 218(1):70–??, January 2019. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/1/70.

Locke:2019:RGR

[LT19b] Melissa N. Locke and Jeremy Thorner. Rab5 GTPases are required for optimal TORC2 function. Journal of Cell Biology, 218(3):961–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/3/961.

Lorincz:2017:RPA

[LTB^{+}17] Péter Lőrincz, Sarolta Tóth, Péter Benkő, Zsolt Lakatos, Attila Boda, Gábor Glatz, Martina Zobel, Sara Bisi, Krisztina Hegedűs, Szabolcs Takáts, Giorgio Scita, and Gábor Juhász. Rab2 promotes autophagic and endocytic lysosomal degradation. Journal of Cell Biology, 216(7):1937–??, July 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/7/1937.

Lawrence:2016:LCP

[LTC^{+}16] Katherine S. Lawrence, Erin C. Tapley, Victor E. Cruz, Qianyan Li, Kayla Aung, Kevin C. Hart, Thomas U. Schwartz, Daniel A. Starr, and JoAnne Engebrecht. LINC complexes promote homologous recombination in part through inhibition of nonhomologous end joining. Journal of Cell Biology, 215(6): 801–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/215/6/801.

Li:2018:CBC

[LTC^{+}18] Jian Li, Ji-Xin Tang, Jin-Mei Cheng, Bian Hu, Yu-Qian Wang, Batool Aalia, Xiao-Yu Li, Cheng Jin, Xiu-Xia Wang, Shou-Long Deng, Yan Zhang, Su-Ren Chen, Wei-Ping Qian, Qing-Yuan Sun, Xing-Xu Huang, and Yi-Xun Liu. Cyclin B2 can compensate for Cyclin B1 in oocyte meiosis I. Journal of Cell Biology, 217(11):3901–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/11/3901.

Lee:2018:AFP

[LTG^{+}18] Sungsu Lee, Han Yen Tan, Ivayla I. Geneva, Aleksandr Kruglov, and Peter D. Calvert. Actin filaments partition primary cilia membranes into distinct fluid corrals. Journal of Cell Biology, 217(8):2831–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/8/2831.

Lord:2015:ANP

[LTRW15] Christopher L. Lord, Benjamin L. Timney, Michael P. Rout, and Susan R. Wente. Altering nuclear pore complex function impacts longevity and mitochondrial function in S. cerevisiae. Journal of Cell Biology, 208(6):729–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/6/729.

Loukil:2017:DCC

[LTS17] Abdelhalim Loukil, Kati Tormanen, and Christine Sütterlin. The daughter centriole controls ciliogenesis by regulating Neurl-4 localization at the centrosome. Journal of Cell Biology, 216 (5):1287–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/5/1287.

Lambrus:2015:PPA

[LUC^{+}15] Bramwell G. Lambrus, Yumi Uetake, Kevin M. Clutario, Vikas Daggubati, Michael Snyder, Greenfield Sluder, and Andrew J. Holland. p53 protects against genome instability following centriole duplication failure. Journal of Cell Biology, 210(1):63–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
210/1/63.

Lawrimore:2015:DLG

[LVF^{+}15] Josh Lawrimore, Paula A. Vasquez, Michael R. Falvo, Russell M. Taylor, Leandra Vicci, Elaine Yeh, M. Gregory Forest, and Kerry Bloom. DNA loops generate intracentromere tension in mitosis. Journal of Cell Biology, 210(4):553–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/210/4/
553.

Louka:2018:PCG

[LVG^{+}18] Panagiota Louka, Krishna Kumar Vasudevan, Mayukh Guha, Ewa Joachimiak, Dorota Wloga, Raphaël F.-X. Tomasi, Charles N. Baroud, Pascale Dupuis-Williams, Domenico F.

Galati, Chad G. Pearson, Luke M. Rice, James J. Moresco, John R. Yates, Yu-Yang Jiang, Karl Lechtreck, William Dentler, and Jacek Gaertig. Proteins that control the geometry of microtubules at the ends of cilia. Journal of Cell Biology, 217(12):4298–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/12/4298.

Lammerding:2016:NER

[LW16a] Jan Lammerding and Katarina Wolf. Nuclear envelope rupture: Actin fibers are putting the squeeze on the nucleus. Journal of Cell Biology, 215(1):5–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/215/1/5.

Liu:2016:PCF

[LW16b] Zairan Liu and Orion D. Weiner. Positioning the cleavage furrow: All you need is Rho. Journal of Cell Biology, 213(6):605– ??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
213/6/605.

Lambert:2017:NCA

[LW17] Talley J. Lambert and Jennifer C. Waters. Navigating challenges in the application of superresolution microscopy. Journal of Cell Biology, 216(1):53–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/1/53.

Luo:2015:WLR

[LWF^{+}15] Lichao Luo, Huashan Wang, Chao Fan, Sen Liu, and Yu Cai. Wnt ligands regulate Tkv expression to constrain Dpp activity in the Drosophila ovarian stem cell niche. Journal of Cell Biology, 209(4):595–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/4/595.

Liu:2018:LEA

[LWH^{+}18] Hang Liu, Shimin Wang, Weijian Hang, Jinghu Gao, Wenjuan Zhang, Zihang Cheng, Chao Yang, Jun He, Jie Zhou, Juan Chen, and Anbing Shi. LET-413/ erbin acts as a RAB-5 effector to promote RAB-10 activation during endocytic recycling. Journal of Cell Biology, 217(1):299–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/1/299.

Laban:2018:VRL

[LWZ^{+}18] Hebatullah Laban, Andreas Weigert, Joana Zink, Amro Elgheznawy, Christoph Schürmann, Lea Günther, Randa Abdel Malik, Sabrina Bothur, Susanne Wingert, Rolf Bremer, Michael A. Rieger, Bernhard Brüne, Ralf P. Brandes, Ingrid Fleming, and Peter M. Benz. VASP regulates leukocyte infiltration, polarization, and vascular repair after ischemia. Journal of Cell Biology, 217(4):1503–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/4/1503.

Li:2019:CBR

[LWZ^{+}19] Yufei Li, Leyun Wang, Linlin Zhang, Zhengquan He, Guihai Feng, Hao Sun, Jiaqiang Wang, Zhikun Li, Chao Liu, Jiabao Han, Junjie Mao, Pengcheng Li, Xuewei Yuan, Liyuan Jiang, Ying Zhang, Qi Zhou, and Wei Li. Cyclin B3 is required for metaphase to anaphase transition in oocyte meiosis I. Journal of Cell Biology, 218(5):1553–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/5/1553.

Liu:2017:WRE

[LXJ^{+}17] Kai Liu, Ruxiao Xing, Youli Jian, Zhiyang Gao, Xinli Ma, Xiaojuan Sun, Yang Li, Meng Xu, Xin Wang, Yudong Jing, Weixiang Guo, and Chonglin Yang. WDR91 is a Rab7 effector required for neuronal development. Journal of Cell Biology, 216(10):3307–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/10/3307.

Li:2015:TAF

[LXR^{+}15] Sunan Li, Shan Xu, Brian A. Roelofs, Liron Boyman, W. Jonathan Lederer, Hiromi Sesaki, and Mariusz Karbowski. Transient assembly of F-actin on the outer mitochondrial membrane contributes to mitochondrial fission. Journal of Cell Biology, 208(1):109–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/1/109.

Li:2015:SCC

[LYO15] Wenjing Li, Peishan Yi, and Guangshuo Ou. Somatic CRISPR– Cas9–induced mutations reveal roles of embryonically essential dynein chains in Caenorhabditis elegans cilia. Journal of Cell Biology, 208(6):683–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/6/683.

Lim:2016:LCC

[LZ16] Chun-Yan Lim and Roberto Zoncu. The lysosome as a command-and-control center for cellular metabolism. Journal of Cell Biology, 214(6):653–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/6/653.

Liu:2015:CNA

[LZC^{+}15] Zhong Liu, Guohua Zhong, Phing Chian Chai, Lichao Luo, Sen Liu, Ying Yang, Gyeong-Hun Baeg, and Yu Cai. Coordinated niche-associated signals promote germline homeostasis in the Drosophila ovary. Journal of Cell Biology, 211(2):469–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/2/
469.

Liu:2016:NRS

[LZD^{+}16] An Liu, Zikai Zhou, Rui Dang, Yuehua Zhu, Junxia Qi, Guiqin He, Celeste Leung, Daniel Pak, Zhengping Jia, and Wei Xie. Neuroligin 1 regulates spines and synaptic plasticity via LIMK1/cofilin-mediated actin reorganization. Journal of Cell Biology, 212(4):449–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/4/449.

Liu:2018:RRM

[LZH^{+}18] Xian-Dong Liu, Xiao-Na Zhu, Michael M. Halford, Tian-Le Xu, Mark Henkemeyer, and Nan-Jie Xu. Retrograde regulation of mossy fiber axon targeting and terminal maturation via postsynaptic Lnx1. Journal of Cell Biology, 217(11):4007–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/11/4007.

Maedler:2017:MCM

[MA17] Kathrin Maedler and Amin Ardestani. mTORC in β cells: more Than Only Recognizing Comestibles. Journal of Cell Biology, 216(7):1883–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/7/1883.

Marwaha:2017:REP

[MAJ^{+}17] Rituraj Marwaha, Subhash B. Arya, Divya Jagga, Harmeet Kaur, Amit Tuli, and Mahak Sharma. The Rab7 effector PLEKHM1 binds Arl8b to promote cargo traffic to lysosomes. Journal of Cell Biology, 216(4):1051–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/4/1051.

Meitinger:2016:UMP

[MAK^{+}16] Franz Meitinger, John V. Anzola, Manuel Kaulich, Amelia Richardson, Joshua D. Stender, Christopher Benner, Christopher K. Glass, Steven F. Dowdy, Arshad Desai, Andrew K. Shiau, and Karen Oegema. 53BP1 and USP28 mediate p53 activation and G1 arrest after centrosome loss or extended mitotic duration. Journal of Cell Biology, 214(2):155–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/214/2/155.

Marshall:2015:RRM

[Mar15] Chris Marshall. From RAS to RHO: The making of the great cell biologist Alan Hall (1952–2015). Journal of Cell Biology, 209(4):481–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/4/481.

Martens:2016:NAN

[Mar16a] Sascha Martens. No ATG8s, no problem? How LC3/ GABARAP proteins contribute to autophagy. Journal of Cell Biology, 215(6):761–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/215/6/761.

Martin:2016:ERC

[Mar16b] Adam C. Martin. Embryonic ring closure: Actomyosin rings do the two-step. Journal of Cell Biology, 215(3):301–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/215/3/
301.

Marat:2017:AEU

[Mar17] Andrea Marat. Andrew Ewald: Understanding cellular cooperation. Journal of Cell Biology, 216(1):2–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/216/1/2.

Marat:2019:ECA

[Mar19] Andrea Marat. Early Career Advisory Board: Q&A on career and publishing. Journal of Cell Biology, 218(9):2815– ??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/9/2815.

Mayor:2015:SGL

[May15] Satyajit Mayor. The shifting geography and language of cell biology. Journal of Cell Biology, 209(3):323–??, May 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/209/3/323.

Madsen:2015:SMD

[MB15] Daniel H. Madsen and Thomas H. Bugge. The source of matrixdegrading enzymes in human cancer: Problems of research reproducibility and possible solutions. Journal of Cell Biology, 209(2):195–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/2/195.

Machesky:2017:FYC

[MB17a] Laura Machesky and Vania M. M. Braga. So far, yet so close: α-Catenin dimers help migrating cells get together. Journal of Cell Biology, 216(11):3437–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/11/3437.

Martinez:2017:TDD

[MB17b] Paula Martínez and Maria A. Blasco. Telomere-driven diseases and telomere-targeting therapies. Journal of Cell Biology, 216 (4):875–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/4/875.

Mondin:2019:PRE

[MBC^{+}19] Virginie E. Mondin, Khaled Ben El Kadhi, Clothilde Cauvin, Anthony Jackson-Crawford, Emilie Bélanger, Barbara Decelle, Rémi Salomon, Martin Lowe, Arnaud Echard, and Sébastien Carréno. PTEN reduces endosomal PtdIns(4,5)P_{2} in a phosphatase-independent manner via a PLC pathway. Journal of Cell Biology, 218(7):2198–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/7/2198.

Matakatsu:2017:PAP

[MBF17] Hitoshi Matakatsu, Seth S. Blair, and Richard G. Fehon. The palmitoyltransferase Approximated promotes growth via the Hippo pathway by palmitoylation of Fat. Journal of Cell Biology, 216(1):265–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/1/265.

McLamarrah:2018:OPA

[MBG^{+}18a] Tiffany A. McLamarrah, Daniel W. Buster, Brian J. Galletta, Cody J. Boese, John M. Ryniawec, Natalie Ann Hollingsworth, Amy E. Byrnes, Christopher W. Brownlee, Kevin C. Slep, Nasser M. Rusan, and Gregory C. Rogers. An ordered pattern of Ana2 phosphorylation by Plk4 is required for centriole assembly. Journal of Cell Biology, 217(4):1217–??, April 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/4/1217.

Mehsen:2018:PBP

[MBG^{+}18b] Haytham Mehsen, Vincent Boudreau, Damien Garrido, Mohammed Bourouh, Myreille Larouche, Paul S. Maddox, Andrew Swan, and Vincent Archambault. PP2A-b55 promotes nuclear envelope reformation after mitosis in Drosophila. Journal of Cell Biology, 217(12):4106–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/12/4106.

Maass:2019:IIG

[MBR19] Philipp G. Maass, A. Rasim Barutcu, and John L. Rinn. Interchromosomal interactions: a genomic love story of kissing chromosomes. Journal of Cell Biology, 218(1):27–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/218/1/27.

McLennan:2017:DNP

[MBS^{+}17] Rebecca McLennan, Caleb M. Bailey, Linus J. Schumacher, Jessica M. Teddy, Jason A. Morrison, Jennifer C. Kasemeier-Kulesa, Lauren A. Wolfe, Madeline M. Gogol, Ruth E. Baker, Philip K. Maini, and Paul M. Kulesa. DAN (NBL1) promotes collective neural crest migration by restraining uncontrolled invasion. Journal of Cell Biology, 216(10):3339–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/10/3339.

MacDonald:2018:HWA

[MBS^{+}18] Ewan MacDonald, Louise Brown, Arnaud Selvais, Han Liu, Thomas Waring, Daniel Newman, Jessica Bithell, Douglas Grimes, Sylvie Urbé, Michael J. Clague, and Tobias Zech. HRS– WASH axis governs actin-mediated endosomal recycling and cell invasion. Journal of Cell Biology, 217(7):2549–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/7/
2549.

Mattila:2016:DAC

[MBT16] Pieta K. Mattila, Facundo D. Batista, and Bebhinn Treanor. Dynamics of the actin cytoskeleton mediates receptor cross talk: an emerging concept in tuning receptor signaling. Journal of Cell Biology, 212(3):267–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/3/267.

Mruk:2015:TBS

[MC15] Karen Mruk and James K. Chen. Thinking big with small molecules. Journal of Cell Biology, 209(1):7–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/209/1/7.

Mishra:2016:MRM

[MC16] Prashant Mishra and David C. Chan. Metabolic regulation of mitochondrial dynamics. Journal of Cell Biology, 212(4): 379–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/4/379.

Marchesin:2015:AJR

[MCCL^{+}15] Valentina Marchesin, Antonio Castro-Castro, Catalina Lodillinsky, Alessia Castagnino, Joanna Cyrta, Hélène Bonsang-Kitzis, Laetitia Fuhrmann, Marie Irondelle, Elvira Infante, Guillaume Montagnac, Fabien Reyal, Anne Vincent-Salomon, and Philippe Chavrier. ARF6-JIP3/ 4 regulate endosomal tubules for MT1-MMP exocytosis in cancer invasion. Journal of Cell Biology, 211(2):339–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/2/339.

Mason:2019:YTL

[MCD^{+}19] Devon E. Mason, Joseph M. Collins, James H. Dawahare, Trung Dung Nguyen, Yang Lin, Sherry L. Voytik-Harbin, Pinar Zorlutuna, Mervin C. Yoder, and Joel D. Boerckel. YAP and TAZ limit cytoskeletal and focal adhesion maturation to enable persistent cell motility. Journal of Cell Biology, 218(4):1369–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/4/1369.

Miller:2015:LRD

[MCGC^{+}15] Heather Miller, Thiago Castro-Gomes, Matthias Corrotte, Christina Tam, Timothy K. Maugel, Norma W. Andrews, and Wenxia Song. Lipid raft-dependent plasma membrane repair interferes with the activation of B lymphocytes. Journal of Cell Biology, 211(6):1193–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/211/6/1193.

Matsubayashi:2015:EDCa

[MCGM15a] Yutaka Matsubayashi, Camilla Coulson-Gilmer, and Tom H. Millard. Endocytosis-dependent coordination of multiple actin regulators is required for wound healing. Journal of Cell Biology, 210(3):419–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/3/419.

Matsubayashi:2015:EDCb

[MCGM15b] Yutaka Matsubayashi, Camilla Coulson-Gilmer, and Tom H. Millard. Endocytosis-dependent coordination of multiple actin regulators is required for wound healing. Journal of Cell Biology, 210(4):677–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/4/677.

Mallik:2018:XGI

[MCH^{+}18] Moushami Mallik, Marica Catinozzi, Clemens B. Hug, Li Zhang, Marina Wagner, Julia Bussmann, Jonas Bittern, Sina Mersmann, Christian Klämbt, Hannes C. A. Drexler, Martijn A. Huynen, Juan M. Vaquerizas, and Erik Storkebaum. Xrp1 genetically interacts with the ALS-associated FUS orthologue caz and mediates its toxicity. Journal of Cell Biology, 217(11):3947–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/11/3947.

Moyer:2015:BSP

[MCL^{+}15] Tyler C. Moyer, Kevin M. Clutario, Bramwell G. Lambrus, Vikas Daggubati, and Andrew J. Holland. Binding of STIL to Plk4 activates kinase activity to promote centriole assembly. Journal of Cell Biology, 209(6):863–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/209/6/863.

Malinova:2017:AUS

[MCM^{+}17] Anna Malinová, Zuzana Cvačková, Daniel Matěju, Zuzana Hořejší, Claire Abéza, Franck Vandermoere, Edouard Bertrand, David Staněk, and Céline Verheggen. Assembly of the U5 snRNP component PRPF8 is controlled by the HSP90/R2TP chaperones. Journal of Cell Biology, 216(6):1579–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/6/
1579.

McMurray:2019:LSM

[McM19] Michael A. McMurray. The long and short of membrane curvature sensing by septins. Journal of Cell Biology, 218(4):1083–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/4/1083.

Moriel-Carretero:2017:FAF

[MCOGD^{+}17] María Moriel-Carretero, Sara Ovejero, Marie Gérus-Durand, Dimos Vryzas, and Angelos Constantinou. Fanconi anemia FANCD2 and FANCI proteins regulate the nuclear dynamics of splicing factors. Journal of Cell Biology, 216(12):4007–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/12/4007.

Meng:2015:RUS

[MCS^{+}15] Qingcai Meng, Chunmei Cai, Tingzhe Sun, Qianliang Wang, Weihong Xie, Rongfu Wang, and Jun Cui. Reversible ubiquitination shapes NLRC5 function and modulates NF-κB activation switch. Journal of Cell Biology, 211(5):1025–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/5/
1025.

Munoz:2016:KCM

[MDC^{+}16] Isabelle Munoz, Luca Danelli, Julien Claver, Nicolas Goudin, Mathieu Kurowska, Iris Karina Madera-Salcedo, Jian-Dong Huang, Alain Fischer, Claudia González-Espinosa, Geneviéve de Saint Basile, Ulrich Blank, and Gaël Ménasché. Kinesin-1 controls mast cell degranulation and anaphylaxis through PI3K-dependent recruitment to the granular Slp3/Rab27b complex. Journal of Cell Biology, 215(2):203–??, October 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/215/2/203.

Malaby:2019:KBP

[MDOS19] Heidi L. H. Malaby, Megan E. Dumas, Ryoma Ohi, and Jason Stumpff. Kinesin-binding protein ensures accurate chromosome segregation by buffering KIF18A and KIF15. Journal of Cell Biology, 218(4):1218–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/4/1218.

Mesmin:2016:MLT

[Mes16] Bruno Mesmin. Mitochondrial lipid transport and biosynthesis: a complex balance. Journal of Cell Biology, 214(1):9–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/214/1/

Marquardt:2016:AST

[MF16a] Joseph R. Marquardt and Harold A. Fisk. ARHGEF17 sets the timer for retention of Mps1 at kinetochores. Journal of Cell Biology, 212(6):615–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/6/615.

Mrozowska:2016:RPT

[MF16b] Paulina S. Mrozowska and Mitsunori Fukuda. Regulation of podocalyxin trafficking by Rab small GTPases in 2D and 3D epithelial cell cultures. Journal of Cell Biology, 213(3):355–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
213/3/355.

Meng:2018:GDR

[MF18] Jin Meng and Shawn M. Ferguson. GATOR1-dependent recruitment of FLCN–FNIP to lysosomes coordinates Rag GT-Pase heterodimer nucleotide status in response to amino acids. Journal of Cell Biology, 217(8):2765–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/8/2765.

Monis:2017:BRS

[MFP17] William J. Monis, Victor Faundez, and Gregory J. Pazour. BLOC-1 is required for selective membrane protein trafficking from endosomes to primary cilia. Journal of Cell Biology, 216 (7):2131–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/7/2131.

Maib:2018:CRC

[MFVS18] Hannes Maib, Filipe Ferreira, Stéphane Vassilopoulos, and Elizabeth Smythe. Cargo regulates clathrin-coated pit invagination via clathrin light chain phosphorylation. Journal of Cell Biology, 217(12):4253–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/12/4253.

Moriwaki:2016:FFC

[MG16] Takashi Moriwaki and Gohta Goshima. Five factors can reconstitute all three phases of microtubule polymerization dynamics. Journal of Cell Biology, 215(3):357–??, November 2016.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/215/3/357.

Molday:2017:PDC

[MG17] Robert S. Molday and Andrew F. X. Goldberg. Peripherin diverts ciliary ectosome release to photoreceptor disc morphogenesis. Journal of Cell Biology, 216(5):1227–??, May 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/5/1227.

McHugh:2018:SAC

[MG18] Domhnall McHugh and Jesús Gil. Senescence and aging: Causes, consequences, and therapeutic avenues. Journal of Cell Biology, 217(1):65–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/1/65.

Mendiratta:2019:HSM

[MGA19] Shweta Mendiratta, Alberto Gatto, and Genevieve Almouzni. Histone supply: Multitiered regulation ensures chromatin dynamics throughout the cell cycle. Journal of Cell Biology, 218 (1):39–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/1/39.

Motley:2015:RRP

[MGE^{+}15] Alison M. Motley, Paul C. Galvin, Lakhan Ekal, James M. Nuttall, and Ewald H. Hettema. Reevaluation of the role of Pex1 and dynamin-related proteins in peroxisome membrane biogenesis. Journal of Cell Biology, 211(5):1041–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/5/
1041.

McMillan:2016:APA

[MGJ^{+}16] Kirsty J. McMillan, Matthew Gallon, Adam P. Jellett, Thomas Clairfeuille, Frances C. Tilley, Ian McGough, Chris M. Danson, Kate J. Heesom, Kevin A. Wilkinson, Brett M. Collins, and Peter J. Cullen. Atypical parkinsonism–associated retromer mutant alters endosomal sorting of specific cargo proteins. Journal of Cell Biology, 214(4):389–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/214/4/389.

Martinez-Garcia:2018:TCM

[MGSO^{+}18] Marina Martinez-Garcia, Veit Schubert, Kim Osman, Alice Darbyshire, Eugenio Sanchez-Moran, and F. Chris H. Franklin. TOPII and chromosome movement help remove interlocks between entangled chromosomes during meiosis. Journal of Cell Biology, 217(12):4070–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/12/4070.

Michelena:2019:BCA

[MGT^{+}19] Jone Michelena, Marco Gatti, Federico Teloni, Ralph Imhof, and Matthias Altmeyer. Basal CHK1 activity safeguards its stability to maintain intrinsic S-phase checkpoint functions. Journal of Cell Biology, 218(9):2865–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/9/2865.

McHugh:2018:MET

[MGW18] Toni McHugh, Agata A. Gluszek, and Julie P. I. Welburn. Microtubule end tethering of a processive kinesin-8 motor Kif18b is required for spindle positioning. Journal of Cell Biology, 217 (7):2403–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/7/2403.

Miller:2015:GNM

[MH15] Kelly E. Miller and Rebecca Heald. Glutamylation of Nap1 modulates histone H1 dynamics and chromosome condensation in Xenopus. Journal of Cell Biology, 209(2):211–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/209/2/211.

Magidson:2016:UKR

[MHA^{+}16] Valentin Magidson, Jie He, Jeffrey G. Ault, Christopher B. O’Connell, Nachen Yang, Irina Tikhonenko, Bruce F. McEwen, Haixin Sui, and Alexey Khodjakov. Unattached kinetochores rather than intrakinetochore tension arrest mitosis in taxoltreated cells. Journal of Cell Biology, 212(3):307–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/212/3/

Miyagawa:2019:MMR

[MHA^{+}19] Takuya Miyagawa, Kana Hasegawa, Yoko Aoki, Takuya Watanabe, Yuka Otagiri, Kohei Arasaki, Yuichi Wakana, Kenichi Asano, Masato Tanaka, Hideki Yamaguchi, Mitsuo Tagaya, and Hiroki Inoue. MT1-MMP recruits the ER–Golgi SNARE Bet1 for efficient MT1-MMP transport to the plasma membrane. Journal of Cell Biology, 218(10):3355–??, October 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/10/3355.

Matson:2019:ICD

[MHG^{+}19] Jacob Peter Matson, Amy M. House, Gavin D. Grant, Huaitong Wu, Joanna Perez, and Jeanette Gowen Cook. Intrinsic checkpoint deficiency during cell cycle re-entry from quiescence. Journal of Cell Biology, 218(7):2169–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/7/2169.

Maeshima:2018:CUM

[MHH18] Kazuhiro Maeshima, Kayo Hibino, and Damien F. Hudson. Condensins under the microscope. Journal of Cell Biology, 217 (7):2229–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/7/2229.

Morita:2018:GWC

[MHI^{+}18] Keigo Morita, Yutaro Hama, Tamaki Izume, Norito Tamura, Toshihide Ueno, Yoshihiro Yamashita, Yuriko Sakamaki, Kaito Mimura, Hideaki Morishita, Wataru Shihoya, Osamu Nureki, Hiroyuki Mano, and Noboru Mizushima. Genome-wide CRISPR screen identifies TMEM41B as a gene required for autophagosome formation. Journal of Cell Biology, 217(11): 3817–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/11/3817.

Mast:2018:EIR

[MHS^{+}18] Fred D. Mast, Thurston Herricks, Kathleen M. Strehler, Leslie R. Miller, Ramsey A. Saleem, Richard A. Rachubinski, and John D. Aitchison. ESCRT-III is required for scissioning new peroxisomes from the endoplasmic reticulum. Journal of Cell Biology, 217(6):2087–??, June 2018. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/6/2087.

Milev:2015:TTP

[MHSD^{+}15] Miroslav P. Milev, Benedeta Hasaj, Djenann Saint-Dic, Sary Snounou, Qingchuan Zhao, and Michael Sacher. TRAMM/ TrappC12 plays a role in chromosome congression, kinetochore stability, and CENP–E recruitment. Journal of Cell Biology, 209(2):221–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/2/221.

Miroshnikova:2019:CBM

[MHW19] Yekaterina A. Miroshnikova, Tim Hammesfahr, and Sara A. Wickström. Cell biology and mechanopathology of laminopathic cardiomyopathies. Journal of Cell Biology, 218(2): 393–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/2/393.

Murugesan:2016:FGA

[MHY^{+}16] Sricharan Murugesan, Jinsung Hong, Jason Yi, Dong Li, Jordan R. Beach, Lin Shao, John Meinhardt, Grey Madison, Xufeng Wu, Eric Betzig, and John A. Hammer. Formingenerated actomyosin arcs propel T cell receptor microcluster movement at the immune synapse. Journal of Cell Biology, 215(3):383–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/3/383.

Matsui:2018:AYR

[MJN^{+}18] Takahide Matsui, Peidu Jiang, Saori Nakano, Yuriko Sakamaki, Hayashi Yamamoto, and Noboru Mizushima. Autophagosomal YKT6 is required for fusion with lysosomes independently of syntaxin 17. Journal of Cell Biology, 217(8):2633–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/8/
2633.

Monaco:2016:NCR

[MJSB16] Sara Monaco, Beate Jahraus, Yvonne Samstag, and Hilmar Bading. Nuclear calcium is required for human T cell activation. Journal of Cell Biology, 215(2):231–??, October 2016.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/215/2/231.

Mosalaganti:2017:SRC

[MKA^{+}17] Shyamal Mosalaganti, Jenny Keller, Anika Altenfeld, Michael Winzker, Pascaline Rombaut, Michael Saur, Arsen Petrovic, Annemarie Wehenkel, Sabine Wohlgemuth, Franziska Müller, Stefano Maffini, Tanja Bange, Franz Herzog, Herbert Waldmann, Stefan Raunser, and Andrea Musacchio. Structure of the RZZ complex and molecular basis of its interaction with Spindly. Journal of Cell Biology, 216(4):961–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/216/4/961.

Moran:2019:CCC

[MKA^{+}19] Kyle D. Moran, Hui Kang, Ana V. Araujo, Trevin R. Zyla, Koji Saito, Denis Tsygankov, and Daniel J. Lew. Cell-cycle control of cell polarity in yeast. Journal of Cell Biology, 218(1):171–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/1/171.

Merlini:2018:IRA

[MKD^{+}18] Laura Merlini, Bita Khalili, Omaya Dudin, Laetitia Michon, Vincent Vincenzetti, and Sophie G. Martin. Inhibition of Ras activity coordinates cell fusion with cell–cell contact during yeast mating. Journal of Cell Biology, 217(4):1467–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/4/
1467.

Maeda:2017:TRS

[MKS17] Miharu Maeda, Toshiaki Katada, and Kota Saito. TANGO1 recruits Sec16 to coordinately organize ER exit sites for efficient secretion. Journal of Cell Biology, 216(6):1731–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/6/
1731.

Maddox:2015:CMS

[ML15a] Paul S. Maddox and Anne-Marie Ladouceur. Concentrating on the mitotic spindle. Journal of Cell Biology, 210(5):691–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
210/5/691.

McClure:2015:AMM

[ML15b] Allison W. McClure and Daniel J. Lew. To avoid a mating mishap, yeast focus and communicate. Journal of Cell Biology, 208(7):867–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/7/867.

Mauthe:2016:SSA

[MLJ^{+}16] Mario Mauthe, Martijn Langereis, Jennifer Jung, Xingdong Zhou, Alex Jones, Wienand Omta, Sharon A. Tooze, Björn Stork, Søren Riis Paludan, Tero Ahola, Dave Egan, Christian Behrends, Michal Mokry, Cornelis de Haan, Frank van Kuppeveld, and Fulvio Reggiori. An siRNA screen for ATG protein depletion reveals the extent of the unconventional functions of the autophagy proteome in virus replication. Journal of Cell Biology, 214(5):619–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/5/619.

McLelland:2016:SDP

[MLMF16] Gian-Luca McLelland, Sydney A. Lee, Heidi M. McBride, and Edward A. Fon. Syntaxin-17 delivers PINK1/Parkin-dependent mitochondrial vesicles to the endolysosomal system. Journal of Cell Biology, 214(3):275–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/214/3/275.

Mardaryev:2016:CME

[MLR^{+}16] Andrei N. Mardaryev, Bo Liu, Valentina Rapisarda, Krzysztof Poterlowicz, Igor Malashchuk, Jana Rudolf, Andrey A. Sharov, Colin A. Jahoda, Michael Y. Fessing, Salvador A. Benitah, Guo-Liang Xu, and Vladimir A. Botchkarev. Cbx4 maintains the epithelial lineage identity and cell proliferation in the developing stratified epithelium. Journal of Cell Biology, 212(1):77–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
212/1/77.

Mourier:2015:MRM

[MMB^{+}15] Arnaud Mourier, Elisa Motori, Tobias Brandt, Marie Lagouge, Ilian Atanassov, Anne Galinier, Gunter Rappl, Susanne Brodesser, Kjell Hultenby, Christoph Dieterich, and Nils-Göran Larsson. Mitofusin 2 is required to maintain mitochondrial coenzyme Q levels. Journal of Cell Biology, 208(4): 429–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/208/4/429.

Marivin:2019:GIA

[MMW^{+}19] Arthur Marivin, Veronika Morozova, Isha Walawalkar, Anthony Leyme, Dmitry A. Kretov, Daniel Cifuentes, Isabel Dominguez, and Mikel Garcia-Marcos. GPCR-independent activation of G proteins promotes apical cell constriction in vivo. Journal of Cell Biology, 218(5):1743–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/5/1743.

Muller:2017:DRT

[MN17] Carolin A. Müller and Conrad A. Nieduszynski. DNA replication timing influences gene expression level. Journal of Cell Biology, 216(7):1907–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/7/1907.

Matos:2016:APD

[MNL^{+}16] Carlos A. Matos, Clévio Nóbrega, Susana R. Louros, Bruno Almeida, Elisabete Ferreiro, Jorge Valero, Luís Pereira de Almeida, Sandra Macedo-Ribeiro, and Ana Luísa Carvalho. Ataxin-3 phosphorylation decreases neuronal defects in spinocerebellar ataxia type 3 models. Journal of Cell Biology, 212(4):465–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/4/465.

McLaughlin:2016:TRF

[MNLB16] Colleen N. McLaughlin, Inna V. Nechipurenko, Nan Liu, and Heather T. Broihier. A Toll receptor-FoxO pathway represses Pavarotti/MKLP1 to promote microtubule dynamics in motoneurons. Journal of Cell Biology, 214(4):459–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/214/4/459.

Martin:2016:RRA

[MOJ16] Emmanuel Martin, Marie-Hélène Ouellette, and Sarah Jenna. Rac1/ RhoA antagonism defines cell-to-cell heterogeneity during epidermal morphogenesis in nematodes. Journal of Cell Biology, 215(4):483–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/4/483.

Mokranjac:2016:MPI

[Mok16] Dejana Mokranjac. Mitochondrial protein import: an unexpected disulfide bond. Journal of Cell Biology, 214(4):363–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
214/4/363.

McIntosh:2018:MGA

[MOM^{+}18] J. Richard McIntosh, Eileen O’Toole, Garry Morgan, Jotham Austin, Evgeniy Ulyanov, Fazoil Ataullakhanov, and Nikita Gudimchuk. Microtubules grow by the addition of bent guanosine triphosphate tubulin to the tips of curved protofilaments. Journal of Cell Biology, 217(8):2691–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/8/2691.

Mejlvang:2018:SIR

[MOS^{+}18] Jakob Mejlvang, Hallvard Olsvik, Steingrim Svenning, Jack-Ansgar Bruun, Yakubu Princely Abudu, Kenneth Bowitz Larsen, Andreas Brech, Tom E. Hansen, Hanne Brenne, Terkel Hansen, Harald Stenmark, and Terje Johansen. Starvation induces rapid degradation of selective autophagy receptors by endosomal microautophagy. Journal of Cell Biology, 217(10): 3640–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/10/3640.

MacDonald:2017:GDE

[MP17a] Chris MacDonald and Robert C. Piper. Genetic dissection of early endosomal recycling highlights a TORC1-independent role for Rag GTPases. Journal of Cell Biology, 216(10):3275–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/10/3275.

Myers:2017:VCC

[MP17b] Margaret D. Myers and Gregory S. Payne. Vps13 and Cdc31/ centrin: Puzzling partners in membrane traffic. Journal of Cell Biology, 216(2):299–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/2/299.

McWilliams:2016:MQI

[MPA^{+}16] Thomas G. McWilliams, Alan R. Prescott, George F. G. Allen, Jevgenia Tamjar, Michael J. Munson, Calum Thomson, Miratul M. K. Muqit, and Ian G. Ganley. mito -QC illuminates mitophagy and mitochondrial architecture in vivo. Journal of Cell Biology, 214(3):333–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/214/3/333.

Mindos:2017:MCR

[MpDN^{+}17] Thomas Mindos, Xin peng Dun, Katherine North, Robin D. S. Doddrell, Alexander Schulz, Philip Edwards, James Russell, Bethany Gray, Sheridan L. Roberts, Aditya Shivane, Georgina Mortimer, Melissa Pirie, Nailing Zhang, Duojia Pan, Helen Morrison, and David B. Parkinson. Merlin controls the repair capacity of Schwann cells after injury by regulating Hippo/YAP activity. Journal of Cell Biology, 216(2):495–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/216/2/495.

Marshall:2015:VDF

[MPH^{+}15] Misty R. Marshall, Varsha Pattu, Mahantappa Halimani, Monika Maier-Peuschel, Martha-Lena Müller, Ute Becherer, Wanjin Hong, Markus Hoth, Thomas Tschernig, Yenan T. Bryceson, and Jens Rettig. VAMP8-dependent fusion of recycling endosomes with the plasma membrane facilitates T lymphocyte cytotoxicity. Journal of Cell Biology, 210(1):135–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
210/1/135.

Melchionda:2016:CHE

[MPMP16] Manuela Melchionda, Jon K. Pittman, Roberto Mayor, and Sandip Patel. Ca^{2+} /H^{+} exchange by acidic organelles regulates cell migration in vivo. Journal of Cell Biology, 212(7):803–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
212/7/803.

Montani:2018:NFA

[MPN^{+}18] Laura Montani, Jorge A. Pereira, Camilla Norrmén, Hartmut B. F. Pohl, Elisa Tinelli, Martin Trötzmüller, Gianluca Figlia, Penelope Dimas, Belinda von Niederhäusern, Rachel Schwager, Sebastian Jessberger, Clay F. Semenkovich, Harald C. Köfeler, and Ueli Suter. De novo fatty acid synthesis by Schwann cells is essential for peripheral nervous system myelination. Journal of Cell Biology, 217(4):1353–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/4/1353.

Miteva:2019:RIC

[MPW^{+}19] Katarina T. Miteva, Lucia Pedicini, Lesley A. Wilson, Izzy Jayasinghe, Raphael G. Slip, Katarzyna Marszalek, Hannah J. Gaunt, Fiona Bartoli, Shruthi Deivasigamani, Diego Sobradillo, David J. Beech, and Lynn McKeown. Rab46 integrates Ca^{2+} and histamine signaling to regulate selective cargo release from Weibel–Palade bodies. Journal of Cell Biology, 218(7):2232–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/7/2232.

Marcos-Ramiro:2016:RCE

[MRGWB^{+}16] Beatriz Marcos-Ramiro, Diego García-Weber, Susana Barroso, Jorge Feito, María C. Ortega, Eva Cernuda-Morollón, Natalia Reglero-Real, Laura Fernández-Martín, Maria C. Durán, Miguel A. Alonso, Isabel Correas, Susan Cox, Anne J. Ridley, and Jaime Millán. RhoB controls endothelial barrier recovery by inhibiting Rac1 trafficking to the cell border. Journal of Cell Biology, 213(3):385–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/213/3/385.

Mostofa:2018:CCS

[MRK^{+}18] Md. Golam Mostofa, Muhammad Arifur Rahman, Naoki Koike, Akter M. S. T. Yeasmin, Nafisa Islam, Talukdar Muhammad Waliullah, Shun Hosoyamada, Mitsugu Shimobayashi, Takehiko Kobayashi, Michael N. Hall, and Takashi Ushimaru. CLIP and cohibin separate rDNA from nucleolar proteins destined for degradation by nucleophagy. Journal of Cell Biology, 217(8): 2675–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/8/2675.

McNally:2018:MSE

[MRM18] Francis J. McNally and Antonina Roll-Mecak. Microtubulesevering enzymes: From cellular functions to molecular mechanism. Journal of Cell Biology, 217(12):4057–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/
12/4057.

Michaux:2018:ERD

[MRMM18] Jonathan B. Michaux, François B. Robin, William M. McFadden, and Edwin M. Munro. Excitable RhoA dynamics drive pulsed contractions in the early C. elegans embryo. Journal of Cell Biology, 217(12):4230–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/12/4230.

MacDonald:2015:MGP

[MRO^{+}15] Ryan B. MacDonald, Owen Randlett, Julia Oswald, Takeshi Yoshimatsu, Kristian Franze, and William A. Harris. Müller glia provide essential tensile strength to the developing retina. Journal of Cell Biology, 210(7):1075–??, September 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/210/7/1075.

Mattie:2018:NMT

[MRWM18] Sevan Mattie, Jan Riemer, Jeremy G. Wideman, and Heidi M. McBride. A new mitofusin topology places the redox-regulated C terminus in the mitochondrial intermembrane space. Journal of Cell Biology, 217(2):507–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/2/507.

McNulty:2019:CGD

[MS19a] Shannon M. McNulty and Beth A. Sullivan. Correction: Going the distance: Neocentromeres make long-range contacts with heterochromatin. Journal of Cell Biology, 218(2): 722–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/2/722. See [MS19b].

McNulty:2019:GDN

[MS19b] Shannon M. McNulty and Beth A. Sullivan. Going the distance: Neocentromeres make long-range contacts with heterochromatin. Journal of Cell Biology, 218(1):5–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/1/

5. See correction [MS19a].

Maza:2019:ICC

[MSC19] Nycole A. Maza, William E. Schiesser, and Peter D. Calvert. An intrinsic compartmentalization code for peripheral membrane proteins in photoreceptor neurons. Journal of Cell Biology, 218(11):3753–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/11/3753.

McCaughey:2019:GTP

[MSCS19] Janine McCaughey, Nicola L. Stevenson, Stephen Cross, and David J. Stephens. ER-to-Golgi trafficking of procollagen in the absence of large carriers. Journal of Cell Biology, 218(3):929–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/3/929.

Milberg:2017:CAD

[MSE^{+}17] Oleg Milberg, Akiko Shitara, Seham Ebrahim, Andrius Masedunskas, Muhibullah Tora, Duy T. Tran, Yun Chen, Mary Anne Conti, Robert S. Adelstein, Kelly G. Ten Hagen, and Roberto Weigert. Concerted actions of distinct nonmuscle myosin II isoforms drive intracellular membrane remodeling in live animals. Journal of Cell Biology, 216(7):1925–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/7/1925.

Mangal:2018:TAA

[MSK^{+}18] Sriyash Mangal, Jennifer Sacher, Taekyung Kim, Daniel Sampaio Osório, Fumio Motegi, Ana Xavier Carvalho, Karen Oegema, and Esther Zanin. TPXL-1 activates Aurora A to clear contractile ring components from the polar cortex during cytokinesis. Journal of Cell Biology, 217(3):837–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/3/
837.

Mitra:2019:DRL

[MSK^{+}19] Soumitra Mitra, Poonam Sharma, Simran Kaur, Mohammad Anwar Khursheed, Shivangi Gupta, Mansi Chaudhary, Akshai J. Kurup, and Rajesh Ramachandran. Dual regulation of lin28a by Myc is necessary during zebrafish retina regeneration. Journal of Cell Biology, 218(2):489–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/218/2/489.

Muroyama:2016:DRF

[MSL16] Andrew Muroyama, Lindsey Seldin, and Terry Lechler. Divergent regulation of functionally distinct γ-tubulin complexes during differentiation. Journal of Cell Biology, 213(6):679–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
213/6/679.

Manil-Segalen:2018:CSA

[MSŁK^{+}18] Marion Manil-Ségalen, Małgorzata Łuksza, Joanne Kanaan, Véronique Marthiens, Simon I. R. Lane, Keith T. Jones, Marie-Emilie Terret, Renata Basto, and Marie-Hélène Verlhac. Chromosome structural anomalies due to aberrant spindle forces exerted at gene editing sites in meiosis. Journal of Cell Biology, 217(10):3416–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/10/3416.

Moradi:2017:DRA

[MSS^{+}17] Mehri Moradi, Rajeeve Sivadasan, Lena Saal, Patrick Lüningschrör, Benjamin Dombert, Reena Jagdish Rathod, Daniela C. Dieterich, Robert Blum, and Michael Sendtner. Differential roles of α-, β-, and γ-actin in axon growth and collateral branch formation in motoneurons. Journal of Cell Biology, 216(3):793–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/3/793.

Murley:2015:LLS

[MST^{+}15] Andrew Murley, Reta D. Sarsam, Alexandre Toulmay, Justin Yamada, William A. Prinz, and Jodi Nunnari. Ltc1 is an ER-localized sterol transporter and a component of ER– mitochondria and ER–vacuole contacts. Journal of Cell Biology, 209(4):539–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/4/539.

Mattaini:2016:ISM

[MSV16] Katherine R. Mattaini, Mark R. Sullivan, and Matthew G. Vander Heiden. The importance of serine metabolism in cancer. Journal of Cell Biology, 214(3):249–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/3/249.

Mohan:2019:DMS

[MSV^{+}19] Nitin Mohan, Elena M. Sorokina, Ione Vilanova Verdeny, Angel Sandoval Alvarez, and Melike Lakadamyali. Detyrosinated microtubules spatially constrain lysosomes facilitating lysosome–autophagosome fusion. Journal of Cell Biology, 218 (2):632–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/2/632.

Monachino:2017:WCM

[MSvO17] Enrico Monachino, Lisanne M. Spenkelink, and Antoine M. van Oijen. Watching cellular machinery in action, one molecule at a time. Journal of Cell Biology, 216(1):41–??, January 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/1/41.

Melloy:2007:NFD

[MSW^{+}07] Patricia Melloy, Shu Shen, Erin White, J. Richard McIntosh, and Mark D. Rose. Nuclear fusion during yeast mating occurs by a three-step pathway. Journal of Cell Biology, 179(4): 659–??, November 2007. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress. org/content/179/4/659. See correction [MSW^{+}17].

Melloy:2017:CNF

[MSW^{+}17] Patricia Melloy, Shu Shen, Erin White, J. Richard McIntosh, and Mark D. Rose. Correction: Nuclear fusion during yeast mating occurs by a three-step pathway. Journal of Cell Biology, 216(11):3883–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/11/3883. See [MSW^{+}07].

Mozzetta:2019:CCH

[MT19] Chiara Mozzetta and Francesco Saverio Tedesco. Challenging the “chromatin hypothesis” of cardiac laminopathies with LMNA mutant iPS cells. Journal of Cell Biology, 218(9): 2826–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/9/2826.

Mizutani:2017:OIO

[MTC17] Tomohiro Mizutani, Yoshiyuki Tsukamoto, and Hans Clevers. Oncogene-inducible organoids as a miniature platform to assess cancer characteristics. Journal of Cell Biology, 216(6):1505–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/6/1505.

Ma:2019:CCG

[MTC^{+}19] Hanhui Ma, Li-Chun Tu, Yu-Chieh Chung, Ardalan Naseri, David Grunwald, Shaojie Zhang, and Thoru Pederson. Cell cycle– and genomic distance-dependent dynamics of a discrete chromosomal region. Journal of Cell Biology, 218(5):1467–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/5/1467.

Medeiros:2018:ABM

[MTGG18] Tânia Catarina Medeiros, Ryan Lee Thomas, Ruben Ghillebert, and Martin Graef. Autophagy balances mtDNA synthesis and degradation by DNA polymerase POLG during starvation. Journal of Cell Biology, 217(5):1601–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/5/1601.

Mathew:2017:SAS

[MTM^{+}17] Veena Mathew, Annie S. Tam, Karissa L. Milbury, Analise K. Hofmann, Christopher S. Hughes, Gregg B. Morin, Christopher J. R. Loewen, and Peter C. Stirling. Selective aggregation of the splicing factor Hsh155 suppresses splicing upon genotoxic stress. Journal of Cell Biology, 216(12):4027–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/12/4027.

Ma:2016:CCN

[MTN^{+}16] Hanhui Ma, Li-Chun Tu, Ardalan Naseri, Maximiliaan Huisman, Shaojie Zhang, David Grunwald, and Thoru Pederson. CRISPR-cas9 nuclear dynamics and target recognition in living cells. Journal of Cell Biology, 214(5):529–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/214/5/529.

Miao:2019:CRT

[MVJ^{+}19] Hui Miao, Timothy E. Vanderleest, Cayla E. Jewett, Dinah Loerke, and J. Todd Blankenship. Cell ratcheting through the Sbf RabGEF directs force balancing and stepped apical constriction. Journal of Cell Biology, 218(11):3845–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/11/3845.

McConnell:2016:RFE

[MvVV^{+}16] Russell E. McConnell, J. Edward van Veen, Marina Vidaki, Adam V. Kwiatkowski, Aaron S. Meyer, and Frank B. Gertler. A requirement for filopodia extension toward Slit during Robomediated axon repulsion. Journal of Cell Biology, 213(2): 261–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/2/261.

McHugh:2017:DKM

[MW17] Toni McHugh and Julie P. I. Welburn. Dynein at kinetochores: Making the connection. Journal of Cell Biology, 216(4):855–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/4/855.

Mistriotis:2019:CHM

[MWB^{+}19] Panagiotis Mistriotis, Emily O. Wisniewski, Kaustav Bera, Jeremy Keys, Yizeng Li, Soontorn Tuntithavornwat, Robert A. Law, Nicolas A. Perez-Gonzalez, Eda Erdogmus, Yuqi Zhang, Runchen Zhao, Sean X. Sun, Petr Kalab, Jan Lammerding, and Konstantinos Konstantopoulos. Confinement hinders motility by inducing RhoA-mediated nuclear influx, volume expansion, and blebbing. Journal of Cell Biology, 218(12):4093–4111, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
218/12/4093/132522/Confinement-hinders-motility-by-
inducing-RhoA.

Moudgil:2015:NRF

[MWF^{+}15] Devinderjit K. Moudgil, Nathan Westcott, Jakub K. Famulski, Kinjal Patel, Dawn Macdonald, Howard Hang, and Gordon K. T. Chan. A novel role of farnesylation in targeting a mitotic checkpoint protein, human Spindly, to kinetochores. Journal of Cell Biology, 208(7):881–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/208/7/881.

Messenger:2018:CSE

[MWSM18] Scott W. Messenger, Sang Su Woo, Zhongze Sun, and Thomas F. J. Martin. A Ca^{2+}-stimulated exosome release pathway in cancer cells is regulated by Munc13-4. Journal of Cell Biology, 217(8):2877–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/217/8/2877. See correction

[MWSM19].

Messenger:2019:CCS

[MWSM19] Scott W. Messenger, Sang Su Woo, Zhongze Sun, and Thomas F. J. Martin. Correction: A Ca^{2+}-stimulated exosome release pathway in cancer cells is regulated by Munc13-4. Journal of Cell Biology, 218(4):1423–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/4/1423. See [MWSM18].

Miyata:2016:PTU

[MWT^{+}16] Non Miyata, Yasunori Watanabe, Yasushi Tamura, Toshiya Endo, and Osamu Kuge. Phosphatidylserine transport by Ups2–Mdm35 in respiration-active mitochondria. Journal of Cell Biology, 214(1):77–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/1/77.

Moudry:2016:TRR

[MWW^{+}16] Pavel Moudry, Kenji Watanabe, Kamila M. Wolanin, Jirina Bartkova, Isabel E. Wassing, Sugiko Watanabe, Robert Strauss, Rune Troelsgaard Pedersen, Vibe H. Oestergaard, Michael Lisby, Miguel Andújar-Sánchez, Apolinar Maya-Mendoza, Fumiko Esashi, Jiri Lukas, and Jiri Bartek. TOPBP1 regulates RAD51 phosphorylation and chromatin loading and determines PARP inhibitor sensitivity. Journal of Cell Biology, 212(3): 281–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/3/281.

Mathur:2017:MSB

[MXS17] Melina Mathur, Joy S. Xiang, and Christina D. Smolke. Mammalian synthetic biology for studying the cell. Journal of Cell Biology, 216(1):73–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/1/73.

Mason:2016:RGI

[MXV^{+}16] Frank M. Mason, Shicong Xie, Claudia G. Vasquez, Michael Tworoger, and Adam C. Martin. RhoA GTPase inhibition organizes contraction during epithelial morphogenesis. Journal of Cell Biology, 214(5):603–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/214/5/603.

Murley:2017:STM

[MYN^{+}17] Andrew Murley, Justin Yamada, Bradley J. Niles, Alexandre Toulmay, William A. Prinz, Ted Powers, and Jodi Nunnari. Sterol transporters at membrane contact sites regulate TORC1 and TORC2 signaling. Journal of Cell Biology, 216(9): 2679–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/9/2679.

Mannen:2016:SNB

[MYT^{+}16] Taro Mannen, Seisuke Yamashita, Kozo Tomita, Naoki Goshima, and Tetsuro Hirose. The Sam68 nuclear body is composed of two RNase-sensitive substructures joined by the adaptor HNRNPL. Journal of Cell Biology, 214(1):45–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/214/1/

Nunnari:2016:DWW

[NA16] Jodi Nunnari and Rebecca Alvania. Doing what we do best, only better. Journal of Cell Biology, 213(2):141–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/213/2/141.

Nunnari:2017:INL

[NA17] Jodi Nunnari and Rebecca Alvania. Introducing a new look for JCB. Journal of Cell Biology, 216(12):3885–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/12/3885.

Nechemia-Arbely:2017:HCC

[NAFM^{+}17] Yael Nechemia-Arbely, Daniele Fachinetti, Karen H. Miga, Nikolina Sekulic, Gautam V. Soni, Dong Hyun Kim, Adeline K. Wong, Ah Young Lee, Kristen Nguyen, Cees Dekker, Bing Ren, Ben E. Black, and Don W. Cleveland. Human centromeric CENP–A chromatin is a homotypic, octameric nucleosome at all cell cycle points. Journal of Cell Biology, 216(3):607–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/3/607.

Neill:2016:EFR

[NBG^{+}16] Thomas Neill, Simone Buraschi, Atul Goyal, Catherine Sharpe, Elizabeth Natkanski, Liliana Schaefer, Andrea Morrione, and Renato V. Iozzo. EphA2 is a functional receptor for the growth factor progranulin. Journal of Cell Biology, 215(5):687–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
215/5/687.

Nithianandam:2018:ABP

[NC18] Vanitha Nithianandam and Cheng-Ting Chien. Actin blobs prefigure dendrite branching sites. Journal of Cell Biology, 217(10):3731–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/10/3731.

Nazio:2016:FTU

[NCV^{+}16] Francesca Nazio, Marianna Carinci, Cristina Valacca, Pamela Bielli, Flavie Strappazzon, Manuela Antonioli, Fabiola Ciccosanti, Carlo Rodolfo, Silvia Campello, Gian Maria Fimia, Claudio Sette, Paolo Bonaldo, and Francesco Cecconi. Finetuning of ULK1 mRNA and protein levels is required for autophagy oscillation. Journal of Cell Biology, 215(6):841–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
215/6/841.

Ng:2019:ECA

[NDC^{+}19] Cai Tong Ng, Li Deng, Chen Chen, Hong Hwa Lim, Jian Shi, Uttam Surana, and Lu Gan. Electron cryotomography analysis of Dam1C/DASH at the kinetochore–spindle interface in situ. Journal of Cell Biology, 218(2):455–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/2/455.

Nelms:2017:TRS

[NDL17] Bradlee Nelms, Natasha Furtado Dalomba, and Wayne Lencer. A targeted RNAi screen identifies factors affecting diverse stages of receptor-mediated transcytosis. Journal of Cell Biology, 216(2):511–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/2/511.

Neller:2015:PCC

[NDRJ15] Joachim Neller, Alexander Dünkler, Reinhild Rösler, and Nils Johnsson. A protein complex containing Epo1p anchors the cortical endoplasmic reticulum to the yeast bud tip. Journal of Cell Biology, 208(1):71–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/208/1/71.

Nelson:2017:SAM

[Nel17] Celeste M. Nelson. From static to animated: Measuring mechanical forces in tissues. Journal of Cell Biology, 216(1):29–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/1/29.

Negredo:2017:CCA

[NEW^{+}17] Paloma Navarro Negredo, James R. Edgar, Antoni G. Wrobel, Nathan R. Zaccai, Robin Antrobus, David J. Owen, and Margaret S. Robinson. Contribution of the clathrin adaptor AP-1 subunit μ1 to acidic cluster protein sorting. Journal of Cell Biology, 216(9):2927–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/9/2927.

Napoli:2019:BTN

[NF19] Marco Napoli and Elsa R. Flores. Beware of thy neighbor: Senescent cancer cells feast on adjacent cells to persist. Journal of Cell Biology, 218(11):3535–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/11/3535.

Nino:2016:PMC

[NGG^{+}16] Carlos A. Niño, David Guet, Alexandre Gay, Sergine Brutus, Frédéric Jourquin, Shweta Mendiratta, Jean Salamero, Vincent Géli, and Catherine Dargemont. Posttranslational marks control architectural and functional plasticity of the nuclear pore complex basket. Journal of Cell Biology, 212(2):167–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
212/2/167.

Nguyen:2019:PPP

[NGX^{+}19] Phuoc My Nguyen, Nikhil R. Gandasi, Beichen Xie, Sari Sugahara, Yingke Xu, and Olof Idevall-Hagren. The PI(4)P phosphatase Sac2 controls insulin granule docking and release. Journal of Cell Biology, 218(11):3714–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/11/3714.

Nagashima:2019:SNI

[NHA^{+}19] Ryosuke Nagashima, Kayo Hibino, S. S. Ashwin, Michael Babokhov, Shin Fujishiro, Ryosuke Imai, Tadasu Nozaki, Sachiko Tamura, Tomomi Tani, Hiroshi Kimura, Michael Shribak, Masato T. Kanemaki, Masaki Sasai, and Kazuhiro Maeshima. Single nucleosome imaging reveals loose genome chromatin networks via active RNA polymerase II. Journal of Cell Biology, 218(5):1511–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/5/1511.

Nelson:2015:TPM

[NHCB15] Christian R. Nelson, Tom Hwang, Pin-Hsi Chen, and Needhi Bhalla. TRIP13^{PCH−2} promotes Mad2 localization to unattached kinetochores in the spindle checkpoint response. Journal of Cell Biology, 211(3):503–??, November 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/211/3/503.

Naufer:2018:PEC

[NHG^{+}18] Amriya Naufer, Victoria E. B. Hipolito, Suriakarthiga Ganesan, Akriti Prashar, Vanina Zaremberg, Roberto J. Botelho, and Mauricio R. Terebiznik. pH of endophagosomes controls association of their membranes with Vps34 and PtdIns(3)P levels. Journal of Cell Biology, 217(1):329–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/1/329.

Niewidok:2018:SMI

[NIdG^{+}18] Benedikt Niewidok, Maxim Igaev, Abel Pereira da Graca, Andre Strassner, Christine Lenzen, Christian P. Richter, Jacob Piehler, Rainer Kurre, and Roland Brandt. Single-molecule imaging reveals dynamic biphasic partition of RNA-binding proteins in stress granules. Journal of Cell Biology, 217(4): 1303–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/4/1303.

Niethammer:2016:NMG

[Nie16] Philipp Niethammer. Neutrophil mechanotransduction: a GEF to sense fluid shear stress. Journal of Cell Biology, 215(1):13–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
215/1/13.

Niedergang:2019:DCM

[Nie19] Florence Niedergang. Dendritic cells mature to resist lamin degradation and herpes virus release. Journal of Cell Biology, 218(2):387–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/2/387.

Nilsson:2019:PPR

[Nil19] Jakob Nilsson. Protein phosphatases in the regulation of mitosis. Journal of Cell Biology, 218(2):395–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/218/2/395.

Nakamura:2019:PRP

[NIN^{+}19] Kuniyuki Nakamura, Tomoko Ikeuchi, Kazuki Nara, Craig S. Rhodes, Peipei Zhang, Yuta Chiba, Saiko Kazuno, Yoshiki Miura, Tetsuro Ago, Eri Arikawa-Hirasawa, Yoh suke Mukouyama, and Yoshihiko Yamada. Perlecan regulates pericyte dynamics in the maintenance and repair of the blood–brain barrier. Journal of Cell Biology, 218(10):3506–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/
10/3506.

Nishimura:2016:DSE

[NIS^{+}16] Tamako Nishimura, Shoko Ito, Hiroko Saito, Sylvain Hiver, Kenta Shigetomi, Junichi Ikenouchi, and Masatoshi Takeichi. DAAM1 stabilizes epithelial junctions by restraining WAVE complex-dependent lateral membrane motility. Journal of Cell Biology, 215(4):559–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/215/4/559.

Nakajo:2016:ECR

[NiYT^{+}16] Atsuhiro Nakajo, Shin ichiro Yoshimura, Hiroko Togawa, Masataka Kunii, Tomohiko Iwano, Ayaka Izumi, Yuria Noguchi, Ayako Watanabe, Ayako Goto, Toshiro Sato, and Akihiro Harada. EHBP1L1 coordinates Rab8 and Bin1 to regulate apical-directed transport in polarized epithelial cells. Journal of Cell Biology, 212(3):297–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/3/297.

Nishimura:2019:GAR

[NKH^{+}19] Kohei Nishimura, Masataka Komiya, Tetsuya Hori, Takehiko Itoh, and Tatsuo Fukagawa. 3D genomic architecture reveals that neocentromeres associate with heterochromatin regions. Journal of Cell Biology, 218(1):134–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/1/134.

Na:2015:TRC

[NKP^{+}15] Bo-Ra Na, Hye-Ran Kim, Indre Piragyte, Hyun-Mee Oh, Min-Sung Kwon, Uroos Akber, Hyun-Su Lee, Do-Sim Park, Woo Keun Song, Zee-Yong Park, Sin-Hyeog Im, Mun-Chual Rho, Young-Min Hyun, Minsoo Kim, and Chang-Duk Jun. TAGLN2 regulates T cell activation by stabilizing the actin cytoskeleton at the immunological synapse. Journal of Cell Biology, 209(1):143–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/1/143.

Nowak:2019:PPD

[NKW^{+}19] Dawid G. Nowak, Ksenya Cohen Katsenelson, Kaitlin E. Watrud, Muhan Chen, Grinu Mathew, Vincent D. D’Andrea, Matthew F. Lee, Manojit Mosur Swamynathan, Irene Casanova-Salas, Megan C. Jibilian, Caroline L. Buckholtz, Alexandra J. Ambrico, Chun-Hao Pan, John E. Wilkinson, Alexandra C. Newton, and Lloyd C. Trotman. The PHLPP2 phosphatase is a druggable driver of prostate cancer progression. Journal of Cell Biology, 218(6):1943–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/6/1943.

Newell-Litwa:2016:BBN

[NL16] Karen A. Newell-Litwa. Breaking down to build up: Neuroligin’s C-terminal domain strengthens the synapse. Journal of Cell Biology, 212(4):375–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/4/375.

Newell-Litwa:2015:RDR

[NLBA^{+}15] Karen A. Newell-Litwa, Mathilde Badoual, Hannelore Asmussen, Heather Patel, Leanna Whitmore, and Alan Rick Horwitz. ROCK1 and 2 differentially regulate actomyosin organization to drive cell and synaptic polarity. Journal of Cell Biology, 210(2):225–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/2/225.

Niu:2019:CIR

[NLH^{+}19] Fang Niu, Ke Liao, Guoku Hu, Susmita Sil, Shannon Callen, Ming lei Guo, Lu Yang, and Shilpa Buch. Cocaine-induced release of CXCL10 from pericytes regulates monocyte transmigration into the CNS. Journal of Cell Biology, 218(2): 700–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/2/700.

Nakamura:2018:SPM

[NLS^{+}18] Masayoshi Nakamura, Jelmer J. Lindeboom, Marco Saltini, Bela M. Mulder, and David W. Ehrhardt. SPR2 protects minus ends to promote severing and reorientation of plant cortical microtubule arrays. Journal of Cell Biology, 217(3):915–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/3/915.

Nakatsu:2015:SII

[NMN^{+}15] Fubito Nakatsu, Mirko Messa, Ramiro Nández, Heather Czapla, Yixiao Zou, Stephen M. Strittmatter, and Pietro De Camilli. Sac2/ INPP5F is an inositol 4-phosphatase that functions in the endocytic pathway. Journal of Cell Biology, 209(1):85–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
209/1/85.

Neisch:2017:SCM

[NNH17] Amanda L. Neisch, Thomas P. Neufeld, and Thomas S. Hays. A STRIPAK complex mediates axonal transport of autophagosomes and dense core vesicles through PP2A regulation. Journal of Cell Biology, 216(2):441–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/2/441.

Nishimura:2015:ITT

[NNK^{+}15] Satoshi Nishimura, Mika Nagasaki, Shinji Kunishima, Akira Sawaguchi, Asuka Sakata, Hiroyasu Sakaguchi, Tsukasa Ohmori, Ichiro Manabe, Joseph E. Italiano, Tomiko Ryu, Naoya Takayama, Issei Komuro, Takashi Kadowaki, Koji Eto, and Ryozo Nagai. IL-1α induces thrombopoiesis through megakaryocyte rupture in response to acute platelet needs. Journal of Cell Biology, 209(3):453–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/209/3/453.

Nguyen:2019:GHL

[NO19] Truc B. Nguyen and James A. Olzmann. Getting a handle on lipid droplets: Insights into ER–lipid droplet tethering. Journal of Cell Biology, 218(4):1089–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/4/1089.

Ninagawa:2015:FDS

[NOS^{+}15] Satoshi Ninagawa, Tetsuya Okada, Yoshiki Sumitomo, Satoshi Horimoto, Takehiro Sugimoto, Tokiro Ishikawa, Shunichi Takeda, Takashi Yamamoto, Tadashi Suzuki, Yukiko Kamiya, Koichi Kato, and Kazutoshi Mori. Forcible destruction of severely misfolded mammalian glycoproteins by the nonglycoprotein ERAD pathway. Journal of Cell Biology, 211(4): 775–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/4/775.

Nyathi:2015:AIP

[NP15] Yvonne Nyathi and Martin R. Pool. Analysis of the interplay of protein biogenesis factors at the ribosome exit site reveals new role for NAC. Journal of Cell Biology, 210(2):287–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/210/2/
287.

Norton:2017:CWG

[NPC17] Heidi K. Norton and Jennifer E. Phillips-Cremins. Crossed wires: 3D genome misfolding in human disease. Journal of Cell Biology, 216(11):3441–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/11/3441.

Nordholm:2017:TRV

[NPÖ^{+}17] Johan Nordholm, Jeanne Petitou, Henrik Östbye, Diogo V. da Silva, Dan Dou, Hao Wang, and Robert Daniels. Translational regulation of viral secretory proteins by the 5’ coding regions and a viral RNA-binding protein. Journal of Cell Biology, 216(8):2283–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/8/2283.

Nguyen:2016:AFL

[NPU^{+}16] Thanh Ngoc Nguyen, Benjamin Scott Padman, Joanne Usher, Viola Oorschot, Georg Ramm, and Michael Lazarou. Atg8 family LC3/GABARAP proteins are crucial for autophagosome– lysosome fusion but not autophagosome formation during PINK1/Parkin mitophagy and starvation. Journal of Cell Biology, 215(6):857–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/6/857.

Nieswandt:2015:MRA

[NS15] Bernhard Nieswandt and Simon Stritt. Megakaryocyte rupture for acute platelet needs. Journal of Cell Biology, 209(3):327–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
209/3/327.

Newman:2018:PPL

[NS18] Laura E. Newman and Gerald S. Shadel. Pink1/Parkin link inflammation, mitochondrial stress, and neurodegeneration. Journal of Cell Biology, 217(10):3327–??, October 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/10/3327.

Nanba:2015:CMP

[NTT^{+}15] Daisuke Nanba, Fujio Toki, Sota Tate, Matome Imai, Natsuki Matsushita, Ken Shiraishi, Koji Sayama, Hiroshi Toki, Shigeki Higashiyama, and Yann Barrandon. Cell motion predicts human epidermal stemness. Journal of Cell Biology, 209 (2):305–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/2/305.

Nakamura:2017:PRG

[NVP17] Mitsutoshi Nakamura, Jeffrey M. Verboon, and Susan M. Parkhurst. Prepatterning by RhoGEFs governs Rho GTPase spatiotemporal dynamics during wound repair. Journal of Cell Biology, 216(12):3959–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/12/3959.

Nair:2019:EMM

[NW19] Praful R. Nair and Denis Wirtz. Enabling migration by moderation: YAP/TAZ are essential for persistent migration. Journal of Cell Biology, 218(4):1092–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/4/1092.

Noblett:2019:DSE

[NWD^{+}19] Nathaniel Noblett, Zilu Wu, Zhao Hua Ding, Seungmee Park, Tony Roenspies, Stephane Flibotte, Andrew D. Chisholm, Yishi Jin, and Antonio Colavita. DIP-2 suppresses ectopic neurite sprouting and axonal regeneration in mature neurons. Journal of Cell Biology, 218(1):125–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/1/125.

Nezich:2015:TTF

[NWFY15] Catherine L. Nezich, Chunxin Wang, Adam I. Fogel, and Richard J. Youle. MiT/ TFE transcription factors are activated during mitophagy downstream of Parkin and Atg5. Journal of Cell Biology, 210(3):435–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/3/435.

Nesmith:2016:HVM

[NWP^{+}16] Alexander P. Nesmith, Matthew A. Wagner, Francesco S. Pasqualini, Blakely B. O’Connor, Mark J. Pincus, Paul R. August, and Kevin Kit Parker. A human in vitro model of Duchenne muscular dystrophy muscle formation and contractility. Journal of Cell Biology, 215(1):47–??, October 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/215/1/47.

Neppl:2017:LCA

[NWW17] Ronald L. Neppl, Chia-Ling Wu, and Kenneth Walsh. lncRNA Chronos is an aging-induced inhibitor of muscle hypertrophy. Journal of Cell Biology, 216(11):3497–??, November 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/11/3497.

Nakamuta:2017:DRD

[NYW^{+}17] Shinichi Nakamuta, Yu-Ting Yang, Chia-Lin Wang, Nicholas B. Gallo, Jia-Ray Yu, Yilin Tai, and Linda Van Aelst. Dual role for DOCK7 in tangential migration of interneuron precursors in the postnatal forebrain. Journal of Cell Biology, 216(12): 4313–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/12/4313.

Oldenburg:2017:LCL

[OBS^{+}17] Anja Oldenburg, Nolwenn Briand, Anita L. Sørensen, Inswasti Cahyani, Akshay Shah, Jan Øivind Moskaug, and Philippe Collas. A lipodystrophy-causing lamin A mutant alters conformation and epigenetic regulation of the anti-adipogenic MIR335 locus. Journal of Cell Biology, 216(9):2731–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/9/
2731.

Omilusik:2015:TRZ

[OBY^{+}15] Kyla D. Omilusik, J. Adam Best, Bingfei Yu, Steven Goossens, Alexander Weidemann, Jessica V. Nguyen, Eve Seuntjens, Agata Stryjewska, Christiane Zweier, Rahul Roychoudhuri, Luca Gattinoni, Lynne M. Bird, Yujiro Higashi, Hisato Kondoh, Danny Huylebroeck, Jody Haigh, and Ananda W. Goldrath. Transcriptional repressor ZEB2 promotes terminal differentiation of CD8^{+} effector and memory T cell populations during infection. Journal of Cell Biology, 211(3):??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/3/
2113OIA259.

ODriscoll:2015:PAS

[OCS15] Jonathan O’Driscoll, Daniel Clare, and Helen Saibil. Prion aggregate structure in yeast cells is determined by the Hsp104-Hsp110 disaggregase machinery. Journal of Cell Biology, 211 (1):145–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/1/145.

ODonnell:2016:JDH

[O’D16a] Marie Anne O’Donnell. Job Dekker: Hitting the scientific hi– Cs. Journal of Cell Biology, 215(4):434–??, November 2016.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/215/4/434.

ODonnell:2016:KRD

[O’D16b] Marie Anne O’Donnell. Katja Röper: Deciphering tissue origami. Journal of Cell Biology, 215(2):140–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/215/2/140.

ODonnell:2017:AMS

[O’D17a] Marie Anne O’Donnell. Ann Miller: Shaping cells and scientists. Journal of Cell Biology, 216(8):2232–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/8/
2232.

ODonnell:2017:EGC

[O’D17b] Marie Anne O’Donnell. Erin Goley: Catching the bug for studying the cytoskeleton. Journal of Cell Biology, 216(3):528–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/3/528.

ODonnell:2017:JKC

[O’D17c] Marie Anne O’Donnell. Jonathan Kagan: a cell biologist’s view of immunity. Journal of Cell Biology, 216(6):1502–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/6/
1502.

ODonnell:2017:LLC

[O’D17d] Marie Anne O’Donnell. Ling-ling Chen: Shaping a career in RNA biology. Journal of Cell Biology, 216(11):3426–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/11/3426.

ODonnell:2017:MQM

[O’D17e] Marie Anne O’Donnell. Margot Quinlan: Muscling in on oogenesis. Journal of Cell Biology, 216(10):2992–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/
10/2992.

ODonnell:2017:NRS

[O’D17f] Marie Anne O’Donnell. Nasser Rusan: Scoping out centrosomes. Journal of Cell Biology, 216(12):3889–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/
12/3889.

ODonnell:2017:YFE

[O’D17g] Marie Anne O’Donnell. Yaron Fuchs: Exploring the mysterious mixture of life and death. Journal of Cell Biology, 216(9): 2600–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/9/2600.

ODonnell:2018:BLS

[O’D18a] Marie Anne O’Donnell. Bao-liang Song: Loving biology in the time of cholesterol. Journal of Cell Biology, 217(2): 433–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/2/433.

ODonnell:2018:DCE

[O’D18b] Marie Anne O’Donnell. Danica Chen: From early learning to aging research. Journal of Cell Biology, 217(1):07–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/1/
07.

ODonnell:2018:EPF

[O’D18c] Marie Anne O’Donnell. Erika Pearce: Fitting metabolism and immunity together, to a T. Journal of Cell Biology, 217 (7):2223–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/7/2223.

ODonnell:2018:LLP

[O’D18d] Marie Anne O’Donnell. Laura Lackner: Passing on the scientific baton. Journal of Cell Biology, 217(3):797–??, March 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/3/797.

ODonnell:2018:TBL

[O’D18e] Marie Anne O’Donnell. Tamas Balla: Leading the way with phosphoinositides. Journal of Cell Biology, 217(4):1161–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/4/1161.

ODonnell:2018:TIF

[O’D18f] Marie Anne O’Donnell. Tatsushi Igaki: Flying up the research summits. Journal of Cell Biology, 217(10):3317–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/
10/3317.

ODonnell:2018:VSM

[O’D18g] Marie Anne O’Donnell. Victoria Sanz–Moreno: Rho together for cancer research. Journal of Cell Biology, 217(6):1885–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/6/1885.

ODonnell:2019:BTK

[O’D19a] Marie Anne O’Donnell. Brajendra Tripathi: Keeping an eye out for translational research. Journal of Cell Biology, 218(10): 3161–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/10/3161.

ODonnell:2019:ERS

[O’D19b] Marie Anne O’Donnell. Efraín E. Rivera–Serrano: Personal training for scientists. Journal of Cell Biology, 218(8):2435–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/8/2435.

ODonnell:2019:GMG

[O’D19c] Marie Anne O’Donnell. Gabriel Muhire Gihana: a candle loses nothing by lighting other candles. Journal of Cell Biology, 218 (5):1425–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/5/1425.

ODonnell:2019:GAS

[O’D19d] Marie Anne O’Donnell. Greg Alushin: The shape of things to come. Journal of Cell Biology, 218(7):2073–??, July 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/7/2073.

ODonnell:2019:GST

[O’D19e] Marie Anne O’Donnell. Gustavo Silva: Translating the ubiquitin code. Journal of Cell Biology, 218(1):3–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/218/1/3.

ODonnell:2019:KME

[O’D19f] Marie Anne O’Donnell. Kazuhiro Maeshima: Excitement under the microscope. Journal of Cell Biology, 218(11):3529– ??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/11/3529.

ODonnell:2019:KSG

[O’D19g] Marie Anne O’Donnell. Kota Saito: Getting out and about. Journal of Cell Biology, 218(6):1765–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/6/1765.

ODonnell:2019:LFL

[O’D19h] Marie Anne O’Donnell. Lillian Fritz–Laylin: Keeping up to speed with evolutionary cell biology. Journal of Cell Biology, 218(4):1081–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/4/1081.

ODonnell:2019:MLB

[O’D19i] Marie Anne O’Donnell. Michael Lazarou: Building a body of research. Journal of Cell Biology, 218(12):3881–3882, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/218/12/3881/132519/Michael-Lazarou-Building-
a-body-of-researchMichael.

Olenick:2019:DAH

[ODH19] Mara A. Olenick, Roberto Dominguez, and Erika L. F. Holzbaur. Dynein activator Hook1 is required for trafficking of BDNF-signaling endosomes in neurons. Journal of Cell Biology, 218(1):220–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/1/220.

Ommer:2019:RGS

[OFP^{+}19] Andrea Ommer, Gianluca Figlia, Jorge A. Pereira, Anna Lena Datwyler, Joanne Gerber, Jonathan DeGeer, Giovanna Lalli, and Ueli Suter. Ral GTPases in Schwann cells promote radial axonal sorting in the peripheral nervous system. Journal of Cell Biology, 218(7):2350–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/7/2350.

OConnor-Giles:2016:TTT

[OG16] Kate O’Connor-Giles. Toll-tally tubular: a newly identified Toll-like receptor–FoxO pathway regulates dynamics of the neuronal microtubule network. Journal of Cell Biology, 214(4): 371–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/214/4/371.

ODonnell:2018:JEC

[OI18a] Marie Anne O’Donnell and Nicole Infarinato. Jerry Edward Chipuk: a powerhouse for mitochondrial biology. Journal of Cell Biology, 217(8):2599–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/8/2599.

Ogawa:2018:NTA

[OI18b] Yutaka Ogawa and Naoko Imamoto. Nuclear transport adapts to varying heat stress in a multistep mechanism. Journal of Cell Biology, 217(7):2341–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/7/2341.

Okatsu:2015:PUC

[OKK^{+}15] Kei Okatsu, Fumika Koyano, Mayumi Kimura, Hidetaka Kosako, Yasushi Saeki, Keiji Tanaka, and Noriyuki Matsuda. Phosphorylated ubiquitin chain is the genuine Parkin receptor. Journal of Cell Biology, 209(1):111–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/209/1/111.

ONeill:2016:UEE

[OKN^{+}16] Audrey K. O’Neill, Abigail A. Kindberg, Terren K. Niethamer, Andrew R. Larson, Hsin-Yi Henry Ho, Michael E. Greenberg, and Jeffrey O. Bush. Unidirectional Eph/ephrin signaling creates a cortical actomyosin differential to drive cell segregation. Journal of Cell Biology, 215(2):217–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/215/2/217.

Ohsaki:2016:PII

[OKY^{+}16] Yuki Ohsaki, Takeshi Kawai, Yukichika Yoshikawa, Jinglei Cheng, Eija Jokitalo, and Toyoshi Fujimoto. PML isoform II plays a critical role in nuclear lipid droplet formation. Journal of Cell Biology, 212(1):29–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/1/29.

Ouenzar:2017:CCD

[OLL^{+}17] Faissal Ouenzar, Maxime Lalonde, Hadrien Laprade, Geneviève Morin, Franck Gallardo, Samuel Tremblay-Belzile, and Pascal Chartrand. Cell cycle-dependent spatial segregation of telomerase from sites of DNA damage. Journal of Cell Biology, 216 (8):2355–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/8/2355.

Oury:2019:MLR

[OLT^{+}19] Julien Oury, Yun Liu, Ana Töpf, Slobodanka Todorovic, Esthelle Hoedt, Veeramani Preethish-Kumar, Thomas A. Neubert, Weichun Lin, Hanns Lochmüller, and Steven J. Burden. MACF1 links Rapsyn to microtubule- and actin-binding proteins to maintain neuromuscular synapses. Journal of Cell Biology, 218(5):1686–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/5/1686.

Orr:2019:NCL

[OM19] Bernardo Orr and Helder Maiato. No chromosome left behind: The importance of metaphase alignment for mitotic fidelity. Journal of Cell Biology, 218(4):1086–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/4/1086.

Oku:2017:EEC

[OMK^{+}17] Masahide Oku, Yuichiro Maeda, Yoko Kagohashi, Takeshi Kondo, Mai Yamada, Toyoshi Fujimoto, and Yasuyoshi Sakai. Evidence for ESCRT- and clathrin-dependent microautophagy. Journal of Cell Biology, 216(10):3263–??, October 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/10/3263.

Otera:2016:DDM

[OMKM16] Hidenori Otera, Non Miyata, Osamu Kuge, and Katsuyoshi Mihara. Drp1-dependent mitochondrial fission via MiD49/51 is essential for apoptotic cristae remodeling. Journal of Cell Biology, 212(5):531–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/5/531.

Otani:2019:CJC

[ONT^{+}19] Tetsuhisa Otani, Thanh Phuong Nguyen, Shinsaku Tokuda, Kei Sugihara, Taichi Sugawara, Kyoko Furuse, Takashi Miura, Klaus Ebnet, and Mikio Furuse. Claudins and JAM–A coordinately regulate tight junction formation and epithelial polarity. Journal of Cell Biology, 218(10):3372–??, October 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/10/3372.

ONeill:2018:RTK

[OO18] Rachel J. O’Neill and Michael J. O’Neill. Replication timing kept in LINE. Journal of Cell Biology, 217(2):441–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/2/
441.

Okumoto:2018:NSV

[OOT^{+}18] Kanji Okumoto, Tatsuaki Ono, Ryusuke Toyama, Ayako Shimomura, Aiko Nagata, and Yukio Fujiki. New splicing variants of mitochondrial Rho GTPase-1 (Miro1) transport peroxisomes. Journal of Cell Biology, 217(2):619–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/2/619.

Oh:2018:MPL

[OPP^{+}18] Jaehak Oh, Justin S. A. Perry, Heather Pua, Nicole Irgens-Möller, Satoshi Ishido, Chyi-Song Hsieh, and Jeoung-Sook Shin.

MARCH1 protects the lipid raft and tetraspanin web from MHCII proteotoxicity in dendritic cells. Journal of Cell Biology, 217(4):1395–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/4/1395.

OByrne:2017:NCP

[OR17] Kenneth J. O’Byrne and Derek J. Richard. Nucleolar caspase-2: Protecting us from DNA damage. Journal of Cell Biology, 216 (6):1521–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/6/1521.

Otomo:2015:MPL

[OSK^{+}15] Takanobu Otomo, Michaela Schweizer, Katrin Kollmann, Valéa Schumacher, Nicole Muschol, Eva Tolosa, Hans-Willi Mittrücker, and Thomas Braulke. Mannose 6 phosphorylation of lysosomal enzymes controls B cell functions. Journal of Cell Biology, 208(2):171–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/2/171.

OShaughnessy:2019:SLL

[OSL^{+}19] Ellen C. O’Shaughnessy, Orrin J. Stone, Paul K. LaFosse, Mihai L. Azoitei, Denis Tsygankov, John M. Heddleston, Wesley R. Legant, Erika S. Wittchen, Keith Burridge, Timothy C. Elston, Eric Betzig, Teng-Leong Chew, David Adalsteinsson, and Klaus M. Hahn. Software for lattice light-sheet imaging of FRET biosensors, illustrated with a new Rap1 biosensor. Journal of Cell Biology, 218(9):3153–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/9/3153.

ORegan:2015:HTM

[OSR^{+}15] Laura O’Regan, Josephina Sampson, Mark W. Richards, Axel Knebel, Daniel Roth, Fiona E. Hood, Anne Straube, Stephen J. Royle, Richard Bayliss, and Andrew M. Fry. Hsp72 is targeted to the mitotic spindle by Nek6 to promote K-fiber assembly and mitotic progression. Journal of Cell Biology, 209(3):349–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
209/3/349.

OConnor:2017:DMF

[OSW^{+}17] Reed M. O’Connor, Elizabeth F. Stone, Charlotte R. Wayne, Emily V. Marcinkevicius, Matt Ulgherait, Rebecca Delventhal, Meghan M. Pantalia, Vanessa M. Hill, Clarice G. Zhou, Sophie McAllister, Anna Chen, Jennifer S. Ziegenfuss, Wesley B. Grueber, Julie C. Canman, and Mimi M. Shirasu-Hiza. A Drosophila model of Fragile X syndrome exhibits defects in phagocytosis by innate immune cells. Journal of Cell Biology, 216(3):595–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/3/595.

Ogungbenro:2018:CCP

[OTG^{+}18] Yetunde Adesanya Ogungbenro, Teresa Casar Tena, David Gaboriau, Pierce Lalor, Peter Dockery, Melanie Philipp, and Ciaran G. Morrison. Centrobin controls primary ciliogenesis in vertebrates. Journal of Cell Biology, 217(4):1205–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/4/
1205.

Ott:2016:MRL

[Ott16] Carolyn M. Ott. Midbody remnant licenses primary cilia formation in epithelial cells. Journal of Cell Biology, 214(3):237–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
214/3/237.

Osorio:2019:IRN

[OWW^{+}19] Liliana Osório, Xuewei Wu, Linsheng Wang, Zhixin Jiang, Carlos Neideck, Guojun Sheng, and Zhongjun Zhou. ISM1 regulates NODAL signaling and asymmetric organ morphogenesis during development. Journal of Cell Biology, 218(7):2388–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/7/2388.

Pereira:2019:PRD

[PA19] Patrícia Pereira and Cecília M. Arraiano. A precision RNA degradation machinery shapes stem cell development. Journal of Cell Biology, 218(8):2437–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/8/2437.

Pereira:2016:IFS

[PABM16] António J. Pereira, Paulo Aguiar, Michael Belsley, and Helder Maiato. Inducible fluorescent speckle microscopy. Journal of Cell Biology, 212(2):245–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/2/245.

Paul:2015:IRP

[PAC^{+}15] Nikki R. Paul, Jennifer L. Allen, Anna Chapman, Maria Morlan-Mairal, Egor Zindy, Guillaume Jacquemet, Laura Fernandez del Ama, Nermina Ferizovic, David M. Green, Jonathan D. Howe, Elisabeth Ehler, Adam Hurlstone, and Patrick T. Caswell. α5β1 integrin recycling promotes Arp2/3independent cancer cell invasion via the formin FHOD3. Journal of Cell Biology, 210(6):1013–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/210/6/1013.

Pinto:2016:PCP

[PAM^{+}16] Maria J. Pinto, Pedro L. Alves, Luís Martins, Joana R. Pedro, Hyun R. Ryu, Noo Li Jeon, Anne M. Taylor, and Ramiro D. Almeida. The proteasome controls presynaptic differentiation through modulation of an on-site pool of polyubiquitinated conjugates. Journal of Cell Biology, 212(7):789–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/212/7/789.

Pasquale:2016:EES

[Pas16] Elena B. Pasquale. Exosomes expand the sphere of influence of Eph receptors and ephrins. Journal of Cell Biology, 214(1):5– ??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
214/1/5.

Pasquale:2019:ERE

[Pas19] Elena B. Pasquale. Eph receptors and ephrins engage in cellular cannibalism. Journal of Cell Biology, 218(10):3168–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/10/3168.

Pattabiraman:2015:MBC

[PBG^{+}15] Shrivatsav Pattabiraman, Claudia Baumann, Daniela Guisado, John J. Eppig, John C. Schimenti, and Rabindranath De La Fuente. Mouse BRWD1 is critical for spermatid postmeiotic transcription and female meiotic chromosome stability. Journal of Cell Biology, 208(1):53–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/208/1/53.

Pleiner:2018:TAM

[PBG18] Tino Pleiner, Mark Bates, and Dirk Görlich. A toolbox of antimouse and anti-rabbit IgG secondary nanobodies. Journal of Cell Biology, 217(3):1143–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/3/1143.

Plutoni:2016:PCP

[PBL^{+}16] Cédric Plutoni, Elsa Bazellieres, Maïlys Le Borgne-Rochet, Franck Comunale, Agusti Brugues, Martial Séveno, Damien Planchon, Sylvie Thuault, Nathalie Morin, Stéphane Bodin, Xavier Trepat, and Cécile Gauthier-Rouvière. P-cadherin promotes collective cell migration via a Cdc42-mediated increase in mechanical forces. Journal of Cell Biology, 212(2):199–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
212/2/199.

Pizzinga:2019:TFM

[PBL^{+}19] Mariavittoria Pizzinga, Christian Bates, Jennifer Lui, Gabriella Forte, Fabián Morales-Polanco, Emma Linney, Barbora Knotkova, Beverley Wilson, Clara A. Solari, Luke E. Berchowitz, Paula Portela, and Mark P. Ashe. Translation factor mRNA granules direct protein synthetic capacity to regions of polarized growth. Journal of Cell Biology, 218(5):1564–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/5/1564.

Pichon:2016:VSE

[PBS^{+}16] Xavier Pichon, Amandine Bastide, Adham Safieddine, Racha Chouaib, Aubin Samacoits, Eugenia Basyuk, Marion Peter, Florian Mueller, and Edouard Bertrand. Visualization of single endogenous polysomes reveals the dynamics of translation in live human cells. Journal of Cell Biology, 214(6): 769–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/214/6/769.

Pan:2017:RDA

[PC17] Simon Pan and Jonah R. Chan. Regulation and dysregulation of axon infrastructure by myelinating glia. Journal of Cell Biology, 216(12):3903–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/12/3903.

Pamula:2019:HRI

[PCF^{+}19] Melissa C. Pamula, Lina Carlini, Scott Forth, Priyanka Verma, Subbulakshmi Suresh, Wesley R. Legant, Alexey Khodjakov, Eric Betzig, and Tarun M. Kapoor. High-resolution imaging reveals how the spindle midzone impacts chromosome movement. Journal of Cell Biology, 218(8):2529–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/8/2529.

Plemel:2017:USS

[PCK^{+}17] Jason R. Plemel, Andrew V. Caprariello, Michael B. Keough, Tyler J. Henry, Shigeki Tsutsui, Tak H. Chu, Geert J. Schenk, Roel Klaver, V. Wee Yong, and Peter K. Stys. Unique spectral signatures of the nucleic acid dye acridine orange can distinguish cell death by apoptosis and necroptosis. Journal of Cell Biology, 216(4):1163–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/4/1163.

Pollock:2016:MOM

[PCM16] Lana M. Pollock, Shih-Wei Chou, and Brian M. McDermott. My oh my(osin): Insights into how auditory hair cells count, measure, and shape. Journal of Cell Biology, 212(2):135–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
212/2/135.

Poulton:2017:CSA

[PCP17] John S. Poulton, John C. Cuningham, and Mark Peifer. Centrosome and spindle assembly checkpoint loss leads to neural apoptosis and reduced brain size. Journal of Cell Biology, 216 (5):1255–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/5/1255.

Pedersen:2019:TMA

[PD19] Ross T. A. Pedersen and David G. Drubin. Type I myosins anchor actin assembly to the plasma membrane during clathrinmediated endocytosis. Journal of Cell Biology, 218(4):1138–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/4/1138.

Petsalaki:2018:EPC

[PDZ18] Eleni Petsalaki, Maria Dandoulaki, and George Zachos. The ESCRT protein Chmp4c regulates mitotic spindle checkpoint signaling. Journal of Cell Biology, 217(3):861–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/217/3/861.

Pfeffer:2016:LSI

[Pfe16] Suzanne R. Pfeffer. Lipoprotein secretion: It takes two to TANGO. Journal of Cell Biology, 213(3):297–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/213/3/297.

Pineda:2019:HFR

[PGMM^{+}19] Cristiana M. Pineda, David G. Gonzalez, Catherine Matte-Martone, Jonathan Boucher, Elizabeth Lathrop, Sara Gallini, Nathan R. Fons, Tianchi Xin, Karen Tai, Edward Marsh, Don X. Nguyen, Kathleen C. Suozzi, Slobodan Beronja, and Valentina Greco. Hair follicle regeneration suppresses Rasdriven oncogenic growth. Journal of Cell Biology, 218(10): 3212–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/10/3212.

Perez-Gonzalez:2019:YTR

[PGRY^{+}19] Nicolas A. Perez-Gonzalez, Nash D. Rochman, Kai Yao, Jiaxiang Tao, Minh-Tam Tran Le, Shannon Flanary, Lucia Sablich, Ben Toler, Eliana Crentsil, Felipe Takaesu, Bram Lambrus, Jessie Huang, Vivian Fu, Pragati Chengappa, Tia M. Jones, Andrew J. Holland, Steven An, Denis Wirtz, Ryan J. Petrie, Kun-Liang Guan, and Sean X. Sun. YAP and TAZ regulate cell volume. Journal of Cell Biology, 218(10):3472–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/10/3472.

Pingali:2016:BCT

[PH16] Hema V. Pingali and Angela K. Hilliker. A beacon in the cytoplasm: Tracking translation of single mRNAs. Journal of Cell Biology, 214(6):649–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/214/6/649.

Philip:2018:TPT

[PH18] Reuben V. Philip and Rene E. Harrison. A tent pole twist on membrane ruffles. Journal of Cell Biology, 217(11):3774– ??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/11/3774.

Peter:2017:VMI

[PHA^{+}17] Arun T. John Peter, Beatrice Herrmann, Diana Antunes, Doron Rapaport, Kai Stefan Dimmer, and Benoît Kornmann. Vps13-Mcp1 interact at vacuole–mitochondria interfaces and bypass ER–mitochondria contact sites. Journal of Cell Biology, 216(10):3219–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/10/3219.

Pal:2016:SDA

[PhHS^{+}16] Kasturi Pal, Sun hee Hwang, Bandarigoda Somatilaka, Hemant Badgandi, Peter K. Jackson, Kathryn DeFea, and Saikat Mukhopadhyay. Smoothened determines β-arrestin–mediated removal of the G protein–coupled receptor Gpr161 from the primary cilium. Journal of Cell Biology, 212(7):861–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/212/7/

Petrie:2017:ANP

[PHKY17] Ryan J. Petrie, Heather M. Harlin, Lulu I. T. Korsak, and Kenneth M. Yamada. Activating the nuclear piston mechanism of 3D migration in tumor cells. Journal of Cell Biology, 216 (1):93–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/1/93.

Panas:2016:MIM

[PIA16] Marc D. Panas, Pavel Ivanov, and Paul Anderson. Mechanistic insights into mammalian stress granule dynamics. Journal of Cell Biology, 215(3):313–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/215/3/313.

Ping:2016:NAM

[PKC^{+}16] Holly A. Ping, Lauren M. Kraft, WeiTing Chen, Amy E. Nilles, and Laura L. Lackner. Num1 anchors mitochondria to the plasma membrane via two domains with different lipid binding specificities. Journal of Cell Biology, 213(5):513–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/213/5/513.

Palani:2019:PTC

[PKH^{+}19] Saravanan Palani, Darius V. Köster, Tomoyuki Hatano, Anton Kamnev, Taishi Kanamaru, Holly R. Brooker, Juan Ramon Hernandez-Fernaud, Alexandra M. E. Jones, Jonathan B. A. Millar, Daniel P. Mulvihill, and Mohan K. Balasubramanian. Phosphoregulation of tropomyosin is crucial for actin cable turnover and division site placement. Journal of Cell Biology, 218(11):3548–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/11/3548.

Pu:2017:RBI

[PKKB17] Jing Pu, Tal Keren-Kaplan, and Juan S. Bonifacino. A Ragulator–BORC interaction controls lysosome positioning in response to amino acid availability. Journal of Cell Biology, 216(12):4183–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/12/4183.

Pedersen:2015:TRM

[PKN^{+}15] Rune Troelsgaard Pedersen, Thomas Kruse, Jakob Nilsson, Vibe H. Oestergaard, and Michael Lisby. TopBP1 is required at mitosis to reduce transmission of DNA damage to G1 daughter cells. Journal of Cell Biology, 210(4):565–??, August 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/210/4/565.

Panas:2019:PGS

[PKS^{+}19] Marc D. Panas, Nancy Kedersha, Tim Schulte, Rui M. Branca, Pavel Ivanov, and Paul Anderson. Phosphorylation of G3BP1-S149 does not influence stress granule assembly. Journal of Cell Biology, 218(7):2425–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/7/2425. See reply [TT19].

Prosperi:2015:MFE

[PLD^{+}15] Marie-Thérèse Prospéri, Priscilla Lépine, Florent Dingli, Perrine Paul-Gilloteaux, René Martin, Damarys Loew, Hans-Joachim Knölker, and Evelyne Coudrier. Myosin 1b functions as an effector of EphB signaling to control cell repulsion. Journal of Cell Biology, 210(2):347–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/210/2/347.

Pierobon:2017:UUF

[PLD17] Paolo Pierobon and Ana-Maria Lennon-Duménil. To use or not to use the force: How B lymphocytes extract surface-tethered antigens. Journal of Cell Biology, 216(1):17–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/216/1/17.

Parisi:2015:ARM

[PLG^{+}15] Alice Parisi, Floriane Lacour, Lorenzo Giordani, Sabine Colnot, Pascal Maire, and Fabien Le Grand. APC is required for muscle stem cell proliferation and skeletal muscle tissue repair. Journal of Cell Biology, 210(5):717–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/5/717.

Papadopoulos:2018:PTN

[PLH18] Natalia Papadopoulos, Johan Lennartsson, and Carl-Henrik Heldin. PDGFRβ translocates to the nucleus and regulates chromatin remodeling via TATA element–modifying factor 1. Journal of Cell Biology, 217(5):1701–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/5/1701.

Park:2015:ANM

[PLS^{+}15] Hyojin Park, Sungwoon Lee, Pravesh Shrestha, Jihye Kim, Jeong Ae Park, Yeongrim Ko, Young Ho Ban, Dae-Young Park, Sang-Jun Ha, Gou Young Koh, Victor Sukbong Hong, Naoki Mochizuki, Young-Myeong Kim, Weontae Lee, and Young-Guen Kwon. AMIGO2, a novel membrane anchor of PDK1, controls cell survival and angiogenesis via Akt activation. Journal of Cell Biology, 211(3):619–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/211/3/619.

Prosser:2015:CRC

[PM15] Suzanna L. Prosser and Ciaran G. Morrison. Centrin2 regulates CP110 removal in primary cilium formation. Journal of Cell Biology, 208(6):693–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/6/693.

Pai:2018:DAD

[PM18] Yun-Jin Pai and Adrian W. Moore. Dendritic actin delivery service. Journal of Cell Biology, 217(10):3325–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/10/3325.

Pontes:2017:MTC

[PMG^{+}17] Bruno Pontes, Pascale Monzo, Laurent Gole, Anabel-Lise Le Roux, Anita Joanna Kosmalska, Zhi Yang Tam, Weiwei Luo, Sophie Kan, Virgile Viasnoff, Pere Roca-Cusachs, Lisa Tucker-Kellogg, and Nils C. Gauthier. Membrane tension controls adhesion positioning at the leading edge of cells. Journal of Cell Biology, 216(9):2959–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/9/2959.

Patel:2017:BHR

[PMHB17] Dharm S. Patel, Sarah M. Misenko, Joonyoung Her, and Samuel F. Bunting. BLM helicase regulates DNA repair by counteracting RAD51 loading at DNA double-strand break sites. Journal of Cell Biology, 216(11):3521–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/
11/3521.

Prusty:2017:ISU

[PMP^{+}17] Archana Bairavasundaram Prusty, Rajyalakshmi Meduri, Bhupesh Kumar Prusty, Jens Vanselow, Andreas Schlosser, and Utz Fischer. Impaired spliceosomal UsnRNP assembly leads to Sm mRNA down-regulation and Sm protein degradation. Journal of Cell Biology, 216(8):2391–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/8/2391.

Panzica:2017:FAP

[PMRM17] Michelle T. Panzica, Harold C. Marin, Anne-Cecile Reymann, and Francis J. McNally. F-actin prevents interaction between sperm DNA and the oocyte meiotic spindle in C. elegans. Journal of Cell Biology, 216(8):2273–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/8/2273.

Perez-Mockus:2017:NRC

[PMRMS17] Gantas Perez-Mockus, Vanessa Roca, Khalil Mazouni, and François Schweisguth. Neuralized regulates Crumbs endocytosis and epithelium morphogenesis via specific Stardust isoforms. Journal of Cell Biology, 216(5):1405–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/5/1405.

Pfanzelter:2018:SSR

[PMW18] Julia Pfanzelter, Serge Mostowy, and Michael Way. Septins suppress the release of vaccinia virus from infected cells. Journal of Cell Biology, 217(8):2911–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/8/2911.

Patron:2019:CUL

[PNE^{+}19] Lilian A. Patrón, Kei Nagatomo, David Tyler Eves, Mays Imad, Kimberly Young, Meaghan Torvund, Xiufang Guo, Gregory C. Rogers, and Konrad E. Zinsmaier. Cul4 ubiquitin ligase cofactor DCAF12 promotes neurotransmitter release and homeostatic plasticity. Journal of Cell Biology, 218(3):993–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/3/
993.

Petropoulos:2016:RPF

[POE^{+}16] Christos Petropoulos, Christiane Oddou, Anouk Emadali, Edwige Hiriart-Bryant, Cyril Boyault, Eva Faurobert, Scott Vande Pol, Joo ri Kim-Kaneyama, Alexandra Kraut, Yohann Coute, Marc Block, Corinne Albiges-Rizo, and Olivier Destaing. Roles of paxillin family members in adhesion and ECM degradation coupling at invadosomes. Journal of Cell Biology, 213(5):585– ??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
213/5/585.

Pollard:2017:NUQ

[Pol17] Thomas D. Pollard. Nine unanswered questions about cytokinesis. Journal of Cell Biology, 216(10):3007–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/
10/3007.

Pek:2015:SIS

[POTZ15] Jun Wei Pek, Ismail Osman, Mandy Li-Ian Tay, and Ruther Teo Zheng. Stable intronic sequence RNAs have possible regulatory roles in Drosophila melanogaster. Journal of Cell Biology, 211(2):243–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/2/243.

Powell:2015:CHT

[Pow15a] Kendall Powell. Cole Haynes: On the trail of mitochondrial dysfunction. Journal of Cell Biology, 208(2):140–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/208/2/

Powell:2015:DMF

[Pow15b] Kendall Powell. Dyche Mullins: Finding filaments at the fringes. Journal of Cell Biology, 209(1):4–??, April 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/209/1/4.

Powell:2015:FPM

[Pow15c] Kendall Powell. Francesca Peri: Micromanaging the lives of microglia. Journal of Cell Biology, 211(2):208–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/211/2/208.

Powell:2015:JVD

[Pow15d] Kendall Powell. Jan van Deursen: From knockout pioneer to antiaging innovator. Journal of Cell Biology, 210(3):366–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
210/3/366.

Powell:2015:JIF

[Pow15e] Kendall Powell. Johanna Ivaska: Finding opposing forces in integrins. Journal of Cell Biology, 208(6):652–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/208/6/652.

Powell:2015:JHS

[Pow15f] Kendall Powell. Junjie Hu: Shape-shifting in the endoplasmic reticulum. Journal of Cell Biology, 209(5):626–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/209/5/626.

Powell:2015:NLU

[Pow15g] Kendall Powell. Nick Lane: Unearthing the first cellular innovations. Journal of Cell Biology, 210(5):684–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/210/5/684.

Powell:2015:NAB

[Pow15h] Kendall Powell. Nihal Altan–Bonnet: Tracking viruses that hijack membranes. Journal of Cell Biology, 210(7):1038– ??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/7/1038.

Powell:2015:PST

[Pow15i] Kendall Powell. Petra Schwille: Taking a minimalist approach to membranes. Journal of Cell Biology, 209(3):320–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/209/3/
320.

Powell:2015:SHB

[Pow15j] Kendall Powell. Sally Horne–Badovinac: Taking a spin around morphogenesis. Journal of Cell Biology, 210(1):4–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/210/1/4.

Powell:2015:TPS

[Pow15k] Kendall Powell. Thoru Pederson: Spotting novel roles for the nucleolus. Journal of Cell Biology, 208(4):384–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/208/4/
384.

Powell:2016:EDP

[Pow16a] Kendall Powell. Elisabetta Dejana: Probing the vagaries of vessels. Journal of Cell Biology, 212(4):368–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/212/4/368.

Powell:2016:EPC

[Pow16b] Kendall Powell. Eugenia Piddini: Chasing how cells outcompete one another. Journal of Cell Biology, 213(3):291–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/213/3/
291.

Powell:2016:JNC

[Pow16c] Kendall Powell. Jeremy Nance: Charting gastrulation’s gyrations. Journal of Cell Biology, 212(2):128–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/212/2/128.

Powell:2016:MJB

[Pow16d] Kendall Powell. Martin Jonikas: Bringing high-throughput genetics to photosynthesis. Journal of Cell Biology, 213(2): 139–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/2/139.

Powell:2016:SDM

[Pow16e] Kendall Powell. Sophie Dumont: Mastering the uncanny mechanics of living systems. Journal of Cell Biology, 212(7): 744–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/7/744.

Peterman:2019:PMR

[PP19] Eric Peterman and Rytis Prekeris. The postmitotic midbody: Regulating polarity, stemness, and proliferation. Journal of Cell Biology, 218(12):3903–3911, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

https://rupress.org/jcb/article/218/12/3903/132539/
The-postmitotic-midbody-Regulating-polarity.

Pagliarini:2015:SPR

[PPB^{+}15] Vittoria Pagliarini, Laura Pelosi, Maria Blaire Bustamante, Annalisa Nobili, Maria Grazia Berardinelli, Marcello D’Amelio, Antonio Musarò, and Claudio Sette. SAM68 is a physiological regulator of SMN2 splicing in spinal muscular atrophy. Journal of Cell Biology, 211(1):77–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/211/1/77.

Park:2016:MCR

[PPK^{+}16] Jong Kook Park, Han Peng, Julia Katsnelson, Wending Yang, Nihal Kaplan, Ying Dong, Joshua Z. Rappoport, CongCong He, and Robert M. Lavker. MicroRNAs-103/ 107 coordinately regulate macropinocytosis and autophagy. Journal of Cell Biology, 215(5):667–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/5/667.

Puleo:2019:MDD

[PPR^{+}19] Julieann I. Puleo, Sara S. Parker, Mackenzie R. Roman, Adam W. Watson, Kiarash Rahmani Eliato, Leilei Peng, Kathylynn Saboda, Denise J. Roe, Robert Ros, Frank B. Gertler, and Ghassan Mouneimne. Mechanosensing during directed cell migration requires dynamic actin polymerization at focal adhesions. Journal of Cell Biology, 218(12):4215–4235, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/218/12/4215/132545/Mechanosensing-during-directed-
cell-migration.

Price:2017:KDE

[Pri17] Brendan D. Price. KDM5A demethylase: Erasing histone modifications to promote repair of DNA breaks. Journal of Cell Biology, 216(7):1871–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/7/1871.

Pereira:2016:MED

[PS16] Gislene Pereira and Elmar Schiebel. Mitotic exit: Determining the PP2A dephosphorylation program. Journal of Cell Biology, 214(5):499–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/5/499.

Pham:2015:ACD

[PSC^{+}15] Kim Pham, Raz Shimoni, Mirren Charnley, Mandy J. Ludford-Menting, Edwin D. Hawkins, Kelly Ramsbottom, Jane Oliaro, David Izon, Stephen B. Ting, Joseph Reynolds, Grant Lythe, Carmen Molina-Paris, Heather Melichar, Ellen Robey, Patrick O. Humbert, Min Gu, and Sarah M. Russell. Asymmetric cell division during T cell development controls downstream fate. Journal of Cell Biology, 210(6):933–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/210/6/933.

Pryde:2016:PDA

[PSCS16] Kenneth R. Pryde, Heather L. Smith, Kai-Yin Chau, and Anthony H. V. Schapira. PINK1 disables the anti-fission machinery to segregate damaged mitochondria for mitophagy. Journal of Cell Biology, 213(2):163–??, April 2016. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/213/2/163.

Pitaval:2017:MSD

[PSL^{+}17] Amandine Pitaval, Fabrice Senger, Gaëlle Letort, Xavier Gidrol, Laurent Guyon, James Sillibourne, and Manuel Théry. Microtubule stabilization drives 3D centrosome migration to initiate primary ciliogenesis. Journal of Cell Biology, 216(11): 3713–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/11/3713.

Prosser:2015:NPC

[PSP^{+}15] Suzanna L. Prosser, Navdeep K. Sahota, Laurence Pelletier, Ciaran G. Morrison, and Andrew M. Fry. Nek5 promotes centrosome integrity in interphase and loss of centrosome cohesion in mitosis. Journal of Cell Biology, 209(3):339–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/209/3/339.

Platt:2018:LRA

[PST18] Emily J. Platt, Leslie Smith, and Mathew J. Thayer. L1 retrotransposon antisense RNA within ASAR lncRNAs controls chromosome-wide replication timing. Journal of Cell Biology, 217(2):541–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/2/541.

Pamula:2016:SCH

[PTK16] Melissa C. Pamula, Shih-Chieh Ti, and Tarun M. Kapoor. The structured core of human β tubulin confers isotype-specific polymerization properties. Journal of Cell Biology, 213(4):425– ??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
213/4/425.

Platani:2015:MDL

[PTMP^{+}15] Melpomeni Platani, Laura Trinkle-Mulcahy, Michael Porter, A. Arockia Jeyaprakash, and William C. Earnshaw. Mio depletion links mTOR regulation to Aurora A and Plk1 activation at mitotic centrosomes. Journal of Cell Biology, 210(1):45–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
210/1/45.

Palmer:2019:CRN

[PTR^{+}19] Nathan Palmer, S. Zakiah A. Talib, Chandrahas Koumar Ratnacaram, Diana Low, Xavier Bisteau, Joanna Hui Si Lee, Elisabeth Pfeiffenberger, Heike Wollmann, Joel Heng Loong Tan, Sheena Wee, Radoslaw Sobota, Jayantha Gunaratne, Daniel M. Messerschmidt, Ernesto Guccione, and Philipp Kaldis. CDK2 regulates the NRF1/Ehmt1 axis during meiotic prophase I. Journal of Cell Biology, 218(9):2896–??, September 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/9/2896.

Pugh:2015:PDM

[Pug15] Edward N. Pugh. Photoreceptor disc morphogenesis: The classical evagination model prevails. Journal of Cell Biology, 211 (3):491–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/3/491.

Pacquelet:2015:PAR

[PUTM15] Anne Pacquelet, Perrine Uhart, Jean-Pierre Tassan, and Grégoire Michaux. PAR-4 and anillin regulate myosin to coordinate spindle and furrow position during asymmetric division. Journal of Cell Biology, 210(7):1085–??, September 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/210/7/1085.

Potapova:2019:SMR

[PUY^{+}19] Tamara A. Potapova, Jay R. Unruh, Zulin Yu, Giulia Rancati, Hua Li, Martha R. Stampfer, and Jennifer L. Gerton. Superresolution microscopy reveals linkages between ribosomal DNA on heterologous chromosomes. Journal of Cell Biology, 218(8): 2492–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/8/2492.

Perez-Vale:2018:MAB

[PVP18] Kia Z. Perez-Vale and Mark Peifer. Modulating apical–basal polarity by building and deconstructing a Yurt. Journal of Cell Biology, 217(11):3772–??, November 2018. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/11/3772.

Patteson:2019:VPC

[PVP^{+}19] Alison E. Patteson, Amir Vahabikashi, Katarzyna Pogoda, Stephen A. Adam, Kalpana Mandal, Mark Kittisopikul, Suganya Sivagurunathan, Anne Goldman, Robert D. Goldman, and Paul A. Janmey. Vimentin protects cells against nuclear rupture and DNA damage during migration. Journal of Cell Biology, 218(12):4079–4092, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/218/12/4079/132537/Vimentin-
protects-cells-against-nuclear-rupture.

Park:2019:PCG

[PW19] Yoo Jin Park and Minna Woo. Pancreatic β cells: Gatekeepers of type 2 diabetes. Journal of Cell Biology, 218(4):1094–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/4/1094.

Prasad:2018:HCA

[PXN18] Rupali Prasad, Chengchao Xu, and Davis T. W. Ng. Hsp40/ 70/110 chaperones adapt nuclear protein quality control to serve cytosolic clients. Journal of Cell Biology, 217(6):2019–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/6/2019.

Pina:2018:SAE

[PYO^{+}18] Francisco Piña, Fumi Yagisawa, Keisuke Obara, J. D. Gregerson, Akio Kihara, and Maho Niwa. Sphingolipids activate the endoplasmic reticulum stress surveillance pathway. Journal of Cell Biology, 217(2):495–??, February 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/2/495.

Qiang:2019:PTC

[QCC^{+}19] Li Qiang, Hong Cao, Jing Chen, Shaun G. Weller, Eugene W. Krueger, Lizhi Zhang, Gina L. Razidlo, and Mark A. McNiven. Pancreatic tumor cell metastasis is restricted by MT1-MMP binding protein MTCBP-1. Journal of Cell Biology, 218(1): 317–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/1/317.

Qi:2017:ELM

[Qi17] Xin Qi. eIF2α links mitochondrial dysfunction to dendritic degeneration. Journal of Cell Biology, 216(3):555–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/3/
555.

Qiao:2017:TRC

[QJP^{+}17] Aijun Qiao, Xiongjie Jin, Junfeng Pang, Demetrius Moskophidis, and Nahid F. Mivechi. The transcriptional regulator of the chaperone response HSF1 controls hepatic bioenergetics and protein homeostasis. Journal of Cell Biology, 216(3):723–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/3/723.

Quiros:2017:MOA

[QPZ^{+}17] Pedro M. Quirós, Miguel A. Prado, Nicola Zamboni, Davide D’Amico, Robert W. Williams, Daniel Finley, Steven P. Gygi, and Johan Auwerx. Multi-omics analysis identifies ATF4 as a key regulator of the mitochondrial stress response in mammals. Journal of Cell Biology, 216(7):2027–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/7/2027.

Qu:2017:CPM

[QSZ^{+}17a] Jia Qu, Wenyi Sun, Jie Zhong, Hao Lv, Mingrui Zhu, Jun Xu, Nan Jin, Zuoquan Xie, Minjia Tan, Shu-Hai Lin, Meiyu Geng, Jian Ding, and Min Huang. Correction: Phosphoglycerate mutase 1 regulates dNTP pool and promotes homologous recombination repair in cancer cells. Journal of Cell Biology, 216(8): 2597–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/8/2597. See [QSZ^{+}17b].

Qu:2017:PMR

[QSZ^{+}17b] Jia Qu, Wenyi Sun, Jie Zhong, Hao Lv, Mingrui Zhu, Jun Xu, Nan Jin, Zuoquan Xie, Minjia Tan, Shu-Hai Lin, Meiyu Geng, Jian Ding, and Min Huang. Phosphoglycerate mutase 1 regulates dNTP pool and promotes homologous recombination repair in cancer cells. Journal of Cell Biology, 216(2): 409–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress. org/content/216/2/409. See correction [QSZ^{+}17a].

Qu:2017:SDM

[QYC^{+}17] Xiaoyi Qu, Feng Ning Yuan, Carlo Corona, Silvia Pasini, Maria Elena Pero, Gregg G. Gundersen, Michael L. Shelanski, and Francesca Bartolini. Stabilization of dynamic microtubules by mDia1 drives Tau-dependent Aβ_{1−42} synaptotoxicity. Journal of Cell Biology, 216(10):3161–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/10/3161.

Qi:2016:SHM

[QYY^{+}16] Yuanbo Qi, Liming Yan, Caiting Yu, Xiangyang Guo, Xin Zhou, Xiaoyu Hu, Xiaofang Huang, Zihe Rao, Zhiyong Lou, and Junjie Hu. Structures of human mitofusin 1 provide insight into mitochondrial tethering. Journal of Cell Biology, 215(5): 621–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/215/5/621.

Qiu:2019:LRC

[QZX19] Rongde Qiu, Jun Zhang, and Xin Xiang. LIS1 regulates cargo-adapter–mediated activation of dynein by overcoming its autoinhibition in vivo. Journal of Cell Biology, 218(11): 3630–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/11/3630.

Qu:2019:SBS

[QZY^{+}19] Qianhui Qu, Qian Zhang, Lu Yang, Yujue Chen, and Hong Liu. SET binding to Sgo1 inhibits Sgo1–cohesin interactions and promotes chromosome segregation. Journal of Cell Biology, 218(8):2514–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/8/2514.

Rabouille:2017:KCR

[Rab17] Catherine Rabouille. KRS: a cut away from release in exosomes. Journal of Cell Biology, 216(7):1891–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/7/1891.

Rosen:2019:DNH

[RAS^{+}19] Jonathan N. Rosen, Mafalda Azevedo, David B. Soffar, Vitaly P. Boyko, Matthew B. Brendel, Victoria K. Schulman, and Mary K. Baylies. The Drosophila Ninein homologue Bsg25D cooperates with Ensconsin in myonuclear positioning. Journal of Cell Biology, 218(2):524–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/2/524.

Ribet:2017:SHS

[RBC^{+}17] David Ribet, Serena Boscaini, Clothilde Cauvin, Martin Siguier, Serge Mostowy, Arnaud Echard, and Pascale Cossart. SUMOylation of human septins is critical for septin filament bundling and cytokinesis. Journal of Cell Biology, 216(12): 4041–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/12/4041.

Riggi:2019:TCE

[RBM^{+}19] Margot Riggi, Clélia Bourgoint, Mariano Macchione, Stefan Matile, Robbie Loewith, and Aurélien Roux. TORC2 controls endocytosis through plasma membrane tension. Journal of Cell Biology, 218(7):2265–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/7/2265.

Raote:2017:TAR

[RBP^{+}17] Ishier Raote, Maria Ortega Bellido, Marinella Pirozzi, Chong Zhang, David Melville, Seetharaman Parashuraman, Timo Zimmermann, and Vivek Malhotra. TANGO1 assembles into rings around COPII coats at ER exit sites. Journal of Cell Biology, 216(4):901–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/4/901.

Romano:2019:PPI

[RBR19] Fabian B. Romano, Neil B. Blok, and Tom A. Rapoport. Peroxisome protein import recapitulated in Xenopus egg extracts. Journal of Cell Biology, 218(6):2021–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/6/2021.

Razzell:2018:FSP

[RBZ18] William Razzell, Maria E. Bustillo, and Jennifer A. Zallen. The force-sensitive protein Ajuba regulates cell adhesion during epithelial morphogenesis. Journal of Cell Biology, 217(10): 3715–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/10/3715.

Rosenberg:2015:MRH

[RC15] Scott C. Rosenberg and Kevin D. Corbett. The multifaceted roles of the HORMA domain in cellular signaling. Journal of Cell Biology, 211(4):745–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/211/4/745.

Ramos:2019:TPS

[RCS^{+}19] Mariona Ramos, Juan Carlos G. Cortés, Mamiko Sato, Sergio A. Rincón, M. Belén Moreno, José Ángel Clemente-Ramos, Masako Osumi, Pilar Pérez, and Juan Carlos Ribas. Two S. pombe septation phases differ in ingression rate, septum structure, and response to F-actin loss. Journal of Cell Biology, 218(12):4171–4194, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/218/12/4171/132546/Two-S-
pombe-septation-phases-differ-in-ingression.

Riccio:2019:DEU

[RDH^{+}19] Victoria Riccio, Nicholas Demers, Rong Hua, Miluska Vissa, Derrick T. Cheng, Amy Wong Strilchuk, Yuqing Wang, G. Angus McQuibban, and Peter Kijun Kim. Deubiquitinating enzyme USP30 maintains basal peroxisome abundance by regulating pexophagy. Journal of Cell Biology, 218(3):798–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/3/
798.

Richter:2019:RCH

[RDN^{+}19] Frank Richter, Sven Dennerlein, Miroslav Nikolov, Daniel C. Jans, Nataliia Naumenko, Abhishek Aich, Thomas MacVicar, Andreas Linden, Stefan Jakobs, Henning Urlaub, Thomas Langer, and Peter Rehling. ROMO1 is a constituent of the human presequence translocase required for YME1L protease import. Journal of Cell Biology, 218(2):598–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/218/2/598.

Roberson:2015:TMO

[RDO^{+}15] Elle C. Roberson, William E. Dowdle, Aysegul Ozanturk, Francesc R. Garcia-Gonzalo, Chunmei Li, Jan Halbritter, Nadia Elkhartoufi, Jonathan D. Porath, Heidi Cope, Allison Ashley-Koch, Simon Gregory, Sophie Thomas, John A. Sayer, Sophie Saunier, Edgar A. Otto, Nicholas Katsanis, Erica E. Davis, Tania Attié-Bitach, Friedhelm Hildebrandt, Michel R. Leroux, and Jeremy F. Reiter. TMEM231, mutated in orofaciodigital and Meckel syndromes, organizes the ciliary transition zone. Journal of Cell Biology, 209(1):129–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/209/1/129.

Rodriguez:2019:ESP

[RFG19] Arturo Vera Rodriguez, Steffen Frey, and Dirk Görlich. Engineered SUMO/protease system identifies Pdr6 as a bidirectional nuclear transport receptor. Journal of Cell Biology, 218 (6):2006–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/6/2006.

Rao:2016:SCU

[RFO^{+}16] Anand N. Rao, Aditi Falnikar, Eileen T. O’Toole, Mary K. Morphew, Andreas Hoenger, Michael W. Davidson, Xiaobing Yuan, and Peter W. Baas. Sliding of centrosome-unattached microtubules defines key features of neuronal phenotype. Journal of Cell Biology, 213(3):329–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/213/3/329.

Redli:2016:SCP

[RGM^{+}16] Patrick M. Redli, Ivana Gasic, Patrick Meraldi, Erich A. Nigg, and Anna Santamaria. The Ska complex promotes Aurora B activity to ensure chromosome biorientation. Journal of Cell Biology, 215(1):77–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/1/77.

Riedel:2018:TTC

[RGMM18] Falko Riedel, Antonio Galindo, Nadine Muschalik, and Sean Munro. The two TRAPP complexes of metazoans have distinct roles and act on different Rab GTPases. Journal of Cell Biology, 217(2):601–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/2/601.

Raturi:2016:TDC

[RGOS^{+}16] Arun Raturi, Tomás Gutiérrez, Carolina Ortiz-Sandoval, Araya Ruangkittisakul, Maria Sol Herrera-Cruz, Jeremy P. Rockley, Kevin Gesson, Dimitar Ourdev, Phing-How Lou, Eliana Lucchinetti, Nasser Tahbaz, Michael Zaugg, Shairaz Baksh, Klaus Ballanyi, and Thomas Simmen. TMX1 determines cancer cell metabolism as a thiol-based modulator of ER–mitochondria Ca^{2+} flux. Journal of Cell Biology, 214(4):433–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/214/4/433.

Reina:2018:CEC

[RGR^{+}18] Jose Reina, Marco Gottardo, Maria G. Riparbelli, Salud Llamazares, Giuliano Callaini, and Cayetano Gonzalez. Centrobin is essential for C-tubule assembly and flagellum development in Drosophila melanogaster spermatogenesis. Journal of Cell Biology, 217(7):2365–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/7/2365.

Rajadurai:2016:IPS

[RHC^{+}16] Charles V. Rajadurai, Serhiy Havrylov, Paula P. Coelho, Colin D. H. Ratcliffe, Kossay Zaoui, Bruce H. Huang, Anie Monast, Naila Chughtai, Veena Sangwan, Frank B. Gertler, Peter M. Siegel, and Morag Park. 5’-inositol phosphatase SHIP2 recruits Mena to stabilize invadopodia for cancer cell invasion. Journal of Cell Biology, 214(6):719–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/6/719.

Randise-Hinchliff:2016:SRT

[RHCS^{+}16] Carlo Randise-Hinchliff, Robert Coukos, Varun Sood, Michael Chas Sumner, Stefan Zdraljevic, Lauren Meldi Sholl, Donna Garvey Brickner, Sara Ahmed, Lauren Watchmaker, and Jason H.

Brickner. Strategies to regulate transcription factor–mediated gene positioning and interchromosomal clustering at the nuclear periphery. Journal of Cell Biology, 212(6):633–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/212/6/
633.

Ripoll:2018:MVB

[RHH^{+}18] Léa Ripoll, Xavier Heiligenstein, Ilse Hurbain, Lia Domingues, Florent Figon, Karl J. Petersen, Megan K. Dennis, Anne Houdusse, Michael S. Marks, Graça Raposo, and Cédric Delevoye. Myosin VI and branched actin filaments mediate membrane constriction and fission of melanosomal tubule carriers. Journal of Cell Biology, 217(8):2709–??, August 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/8/2709.

Ren:2018:DRS

[RHJW18] Suxia Ren, Zengyi Huang, Yuqiang Jiang, and Tao Wang. dTBC1D7 regulates systemic growth independently of TSC through insulin signaling. Journal of Cell Biology, 217(2): 517–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/2/517.

Randrianarison-Huetz:2018:SCS

[RHPH^{+}18] Voahangy Randrianarison-Huetz, Aikaterini Papaefthymiou, Gaëlle Herledan, Chiara Noviello, Ulduz Faradova, Laura Collard, Alessandra Pincini, Emilie Schol, Jean François Decaux, Pascal Maire, Stéphane Vassilopoulos, and Athanassia Sotiropoulos. Srf controls satellite cell fusion through the maintenance of actin architecture. Journal of Cell Biology, 217(2): 685–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/2/685.

Rice:2018:NLG

[Ric18] Luke M. Rice. A new look for the growing microtubule end? Journal of Cell Biology, 217(8):2609–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/8/2609.

Risher:2018:TRP

[RKK^{+}18] W. Christopher Risher, Namsoo Kim, Sehwon Koh, Ji-Eun Choi, Petar Mitev, Erin F. Spence, Louis-Jan Pilaz, Dongqing Wang, Guoping Feng, Debra L. Silver, Scott H. Soderling, Henry H. Yin, and Cagla Eroglu. Thrombospondin receptor α2δ-1 promotes synaptogenesis and spinogenesis via postsynaptic Rac1. Journal of Cell Biology, 217(10):3747–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/10/3747.

Rafiq:2017:PAC

[RLJ^{+}17] Nisha Bte Mohd Rafiq, Zi Zhao Lieu, Tingting Jiang, Cheng han Yu, Paul Matsudaira, Gareth E. Jones, and Alexander D. Bershadsky. Podosome assembly is controlled by the GTPase ARF1 and its nucleotide exchange factor ARNO. Journal of Cell Biology, 216(1):181–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/1/181.

Richter:2015:QCM

[RLM^{+}15] Uwe Richter, Taina Lahtinen, Paula Marttinen, Fumi Suomi, and Brendan J. Battersby. Quality control of mitochondrial protein synthesis is required for membrane integrity and cell fitness. Journal of Cell Biology, 211(2):373–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/211/2/373.

Roy:2018:CSQ

[RLS18a] Sunetra Roy, Jessica W. Luzwick, and Katharina Schlacher. Correction: SIRF: Quantitative in situ analysis of protein interactions at DNA replication forks. Journal of Cell Biology, 217(4):1553–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/4/1553. See [RLS18b].

Roy:2018:SQS

[RLS18b] Sunetra Roy, Jessica W. Luzwick, and Katharina Schlacher. SIRF: Quantitative in situ analysis of protein interactions at DNA replication forks. Journal of Cell Biology, 217(4):1521–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/

217/4/1521. See correction [RLS18a].

Ritter:2016:NKC

[RM16] Alex T. Ritter and Ira Mellman. Natural killer cell granules converge to avoid collateral damage. Journal of Cell Biology, 215(6):765–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/6/765.

Raote:2019:PTV

[RM19] Ishier Raote and Vivek Malhotra. Protein transport by vesicles and tunnels. Journal of Cell Biology, 218(3):737–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/3/
737.

Resende:2018:AIS

[RMB^{+}18] Luís Pedro Resende, Augusta Monteiro, Rita Brás, Tatiana Lopes, and Claudio E. Sunkel. Aneuploidy in intestinal stem cells promotes gut dysplasia in Drosophila. Journal of Cell Biology, 217(11):3930–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/11/3930.

Renault-Mihara:2017:RRS

[RMMS^{+}17] François Renault-Mihara, Masahiko Mukaino, Munehisa Shinozaki, Hiromi Kumamaru, Satoshi Kawase, Matthieu Baudoux, Toshiki Ishibashi, Soya Kawabata, Yuichiro Nishiyama, Keiko Sugai, Kaori Yasutake, Seiji Okada, Masaya Nakamura, and Hideyuki Okano. Regulation of RhoA by STAT3 coordinates glial scar formation. Journal of Cell Biology, 216(8): 2533–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/8/2533.

Romero-Morales:2017:APL

[RMOG17] Alejandra I. Romero-Morales, Natalya A. Ortolano, and Vivian Gama. Apical polarization and lumenogenesis: The apicosome sheds new light. Journal of Cell Biology, 216(12):3891–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/12/3891.

Rhys:2018:LCP

[RMS^{+}18] Alexander D. Rhys, Pedro Monteiro, Christopher Smith, Malti Vaghela, Teresa Arnandis, Takuya Kato, Birgit Leitinger, Erik Sahai, Andrew McAinsh, Guillaume Charras, and Susana A. Godinho. Loss of E-cadherin provides tolerance to centrosome amplification in epithelial cancer cells. Journal of Cell Biology, 217(1):195–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/1/195.

Robbez-Masson:2017:CCY

[RMTR17] Luisa Robbez-Masson, Christopher H. C. Tie, and Helen M. Rowe. Cancer cells, on your histone marks, get SETDB1, silence retrotransposons, and go! Journal of Cell Biology, 216(11): 3429–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/11/3429.

Ruthnick:2017:CSP

[RND^{+}17] Diana Rüthnick, Annett Neuner, Franziska Dietrich, Daniel Kirrmaier, Ulrike Engel, Michael Knop, and Elmar Schiebel. Characterization of spindle pole body duplication reveals a regulatory role for nuclear pore complexes. Journal of Cell Biology, 216(8):2425–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/8/2425.

Romano:2017:PTS

[RNP^{+}17] Julia D. Romano, Sabrina J. Nolan, Corey Porter, Karen Ehrenman, Eric J. Hartman, Ru ching Hsia, and Isabelle Coppens. The parasite Toxoplasma sequesters diverse Rab host vesicles within an intravacuolar network. Journal of Cell Biology, 216(12):4235–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/12/4235.

Rome:2018:NMN

[RO18] Pierre Romé and Hiroyuki Ohkura. A novel microtubule nucleation pathway for meiotic spindle assembly in oocytes. Journal of Cell Biology, 217(10):3431–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/10/3431.

Roy:2016:WRT

[Roy16] Subhojit Roy. Waves, rings, and trails: The scenic landscape of axonal actin. Journal of Cell Biology, 212(2):131–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/212/2/
131.

Ramesh:2016:DBT

[RPMC^{+}16] Ajay Ramesh, Valentina Peleh, Sonia Martinez-Caballero, Florian Wollweber, Frederik Sommer, Martin van der Laan, Michael Schroda, R. Todd Alexander, María Luisa Campo, and Johannes M. Herrmann. A disulfide bond in the TIM23 complex is crucial for voltage gating and mitochondrial protein import. Journal of Cell Biology, 214(4):417–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/4/417.

Riemer:2017:OCP

[RRM^{+}17] Pamela Riemer, Mattias Rydenfelt, Matthias Marks, Karen van Eunen, Kathrin Thedieck, Bernhard G. Herrmann, Nils Blüthgen, Christine Sers, and Markus Morkel. Oncogenic βcatenin and PIK3CA instruct network states and cancer phenotypes in intestinal organoids. Journal of Cell Biology, 216 (6):1567–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/6/1567.

Raiborg:2016:PMR

[RS16] Camilla Raiborg and Harald Stenmark. Plasma membrane repairs by small GTPase Rab3a. Journal of Cell Biology, 213 (6):613–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/6/613.

Rickman:2019:AUD

[RS19] Kimberly Rickman and Agata Smogorzewska. Advances in understanding DNA processing and protection at stalled replication forks. Journal of Cell Biology, 218(4):1096–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/4/

Ratcliffe:2019:HIM

[RSC^{+}19] Colin D. H. Ratcliffe, Nadeem Siddiqui, Paula P. Coelho, Nancy Laterreur, Tumini N. Cookey, Nahum Sonenberg, and Morag Park. HGF-induced migration depends on the PI(3,4,5)P_{3}- binding microexon-spliced variant of the Arf6 exchange factor cytohesin-1. Journal of Cell Biology, 218(1):285–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/1/
285.

Roarty:2015:RRB

[RSCR15] Kevin Roarty, Amy N. Shore, Chad J. Creighton, and Jeffrey M. Rosen. Ror2 regulates branching, differentiation, and actincytoskeletal dynamics within the mammary epithelium. Journal of Cell Biology, 208(3):351–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/3/351.

Reyes:2015:ABP

[RSG^{+}15] Céline Reyes, Céline Serrurier, Tiphaine Gauthier, Yannick Gachet, and Sylvie Tournier. Aurora B prevents chromosome arm separation defects by promoting telomere dispersion and disjunction. Journal of Cell Biology, 208(6):713–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/208/6/713.

Rutkowski:2015:CBF

[RSS15] Joseph M. Rutkowski, Jennifer H. Stern, and Philipp E. Scherer. The cell biology of fat expansion. Journal of Cell Biology, 208(5):501–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/5/501.

Roelants:2015:PKG

[RSvW^{+}15] Françoise M. Roelants, Brooke M. Su, Joachim von Wulffen, Subramaniam Ramachandran, Elodie Sartorel, Amy E. Trott, and Jeremy Thorner. Protein kinase Gin4 negatively regulates flippase function and controls plasma membrane asymmetry. Journal of Cell Biology, 208(3):299–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/3/299.

Roy:2019:DCS

[RVS^{+}19] Babhrubahan Roy, Vikash Verma, Janice Sim, Adrienne Fontan, and Ajit P. Joglekar. Delineating the contribution of Spc105-bound PP1 to spindle checkpoint silencing and kinetochore microtubule attachment regulation. Journal of Cell Biology, 218(12):3926–3942, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

https://rupress.org/jcb/article/218/12/3926/132549/
Delineating-the-contribution-of-Spc105-bound-PP1.

Ren:2019:ECS

[RXEB^{+}19] Mindong Ren, Yang Xu, Hediye Erdjument-Bromage, Alec Donelian, Colin K. L. Phoon, Naohiro Terada, Douglas Strathdee, Thomas A. Neubert, and Michael Schlame. Extramitochondrial cardiolipin suggests a novel function of mitochondria in spermatogenesis. Journal of Cell Biology, 218 (5):1491–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/5/1491.

Rana:2015:ASC

[RYS^{+}15] Anshul Rana, Michelle Yen, Amir Masoud Sadaghiani, Seth Malmersjö, Chan Young Park, Ricardo E. Dolmetsch, and Richard S. Lewis. Alternative splicing converts STIM2 from an activator to an inhibitor of store-operated calcium channels. Journal of Cell Biology, 209(5):653–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/209/5/653.

Reuter:2015:BDO

[RZS^{+}15] Marcel Reuter, Alex Zelensky, Ihor Smal, Erik Meijering, Wiggert A. van Cappellen, H. Martijn de Gruiter, Gijsbert J. van Belle, Martin E. van Royen, Adriaan B. Houtsmuller, Jeroen Essers, Roland Kanaar, and Claire Wyman. BRCA2 diffuses as oligomeric clusters with RAD51 and changes mobility after DNA damage in live cells. Journal of Cell Biology, 208 (6):857–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/208/6/857.

Sassano:2019:STT

[SA19] Maria Livia Sassano and Patrizia Agostinis. Staying in touch: Taking a closer look at ER–Golgi contact sites. Journal of Cell Biology, 218(3):729–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/3/729.

Sui:2018:CEM

[SAB^{+}18] Xuewu Sui, Henning Arlt, Kelly P. Brock, Zon Weng Lai, Frank DiMaio, Debora S. Marks, Maofu Liao, Robert V. Farese, and Tobias C. Walther. Cryo–electron microscopy structure of the lipid droplet–formation protein seipin. Journal of Cell Biology, 217(12):4080–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/12/4080.

Saw:2019:MIT

[SAF^{+}19] Sanjay Saw, Alison Aiken, Hui Fang, Trevor D. McKee, Sarah Bregant, Otto Sanchez, Yan Chen, Ashley Weiss, Brendan C. Dickson, Bertrand Czarny, Ankit Sinha, Amanda Fosang, Vincent Dive, Paul D. Waterhouse, Thomas Kislinger, and Rama Khokha. Metalloprotease inhibitor TIMP proteins control FGF-2 bioavailability and regulate skeletal growth. Journal of Cell Biology, 218(9):3134–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/9/3134.

Srpan:2018:SCD

[SAK^{+}18] Katja Srpan, Ashley Ambrose, Alexandros Karampatzakis, Mezida Saeed, Adam N. R. Cartwright, Karolin Guldevall, Gabriela Dos Santos Cruz De Matos, Björn Önfelt, and Daniel M. Davis. Shedding of CD16 disassembles the NK cell immune synapse and boosts serial engagement of target cells. Journal of Cell Biology, 217(9):3267–??, September 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/9/3267.

Stojkov:2017:RGB

[SAO^{+}17] Darko Stojkov, Poorya Amini, Kevin Oberson, Christiane Sokollik, Andrea Duppenthaler, Hans-Uwe Simon, and Shida Yousefi. ROS and glutathionylation balance cytoskeletal dynamics in neutrophil extracellular trap formation. Journal of Cell Biology, 216(12):4073–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/12/4073.

Shutova:2017:SSN

[SAT^{+}17] Maria S. Shutova, Sreeja B. Asokan, Shefali Talwar, Richard K. Assoian, James E. Bear, and Tatyana M. Svitkina. Self-sorting of nonmuscle myosins IIA and IIB polarizes the cytoskeleton and modulates cell motility. Journal of Cell Biology, 216(9): 2877–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/9/2877.

Simian:2017:OHP

[SB17] Marina Simian and Mina J. Bissell. Organoids: a historical perspective of thinking in three dimensions. Journal of Cell Biology, 216(1):31–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/1/31.

Sun:2019:PQC

[SB19] Zhihao Sun and Jeffrey L. Brodsky. Protein quality control in the secretory pathway. Journal of Cell Biology, 218(10): 3171–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/10/3171.

Squarr:2016:CFA

[SBC^{+}16a] Anna Julia Squarr, Klaus Brinkmann, Baoyu Chen, Tim Steinbacher, Klaus Ebnet, Michael K. Rosen, and Sven Bogdan. Correction: Fat2 acts through the WAVE regulatory complex to drive collective cell migration during tissue rotation. Journal of Cell Biology, 212(7):883–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/7/883.

Squarr:2016:FAT

[SBC^{+}16b] Anna Julia Squarr, Klaus Brinkmann, Baoyu Chen, Tim Steinbacher, Klaus Ebnet, Michael K. Rosen, and Sven Bogdan. Fat2 acts through the WAVE regulatory complex to drive collective cell migration during tissue rotation. Journal of Cell Biology, 212(5):591–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/5/591.

Sala:2017:SPC

[SBM17] Ambre J. Sala, Laura C. Bott, and Richard I. Morimoto. Shaping proteostasis at the cellular, tissue, and organismal level. Journal of Cell Biology, 216(5):1231–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/5/1231.

Schroeder:2019:DNM

[SBM^{+}19] Lena K. Schroeder, Andrew E. S. Barentine, Holly Merta, Sarah Schweighofer, Yongdeng Zhang, David Baddeley, Joerg Bewersdorf, and Shirin Bahmanyar. Dynamic nanoscale morphology of the ER surveyed by STED microscopy. Journal of Cell Biology, 218(1):83–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/1/83.

Srivats:2016:SRI

[SBP^{+}16] Shyam Srivats, Dilshan Balasuriya, Mathias Pasche, Gerard Vistal, J. Michael Edwardson, Colin W. Taylor, and Ruth D. Murrell-Lagnado. Sigma1 receptors inhibit store-operated Ca^{2+} entry by attenuating coupling of STIM1 to Orai1. Journal of Cell Biology, 213(1):65–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/213/1/65.

Saroufim:2015:NBM

[SBR^{+}15] Mark-Albert Saroufim, Pierre Bensidoun, Pascal Raymond, Samir Rahman, Matthew R. Krause, Marlene Oeffinger, and Daniel Zenklusen. The nuclear basket mediates perinuclear mRNA scanning in budding yeast. Journal of Cell Biology, 211(6):1131–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/6/1131.

Schepis:2018:PSR

[SBS^{+}18] Antonino Schepis, Adrian Barker, Yoga Srinivasan, Eaman Balouch, Yaowu Zheng, Ian Lam, Hilary Clay, Chung-Der Hsiao, and Shaun R. Coughlin. Protease signaling regulates apical cell extrusion, cell contacts, and proliferation in epithelia. Journal of Cell Biology, 217(3):1097–??, March 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/3/1097.

Shrimal:2017:DKM

[SCG17] Shiteshu Shrimal, Natalia A. Cherepanova, and Reid Gilmore. DC2 and KCP2 mediate the interaction between the oligosaccharyltransferase and the ER translocon. Journal of Cell Biology, 216(11):3625–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/11/3625.

Schwille:2015:JSL

[Sch15] Petra Schwille. Jump-starting life? Fundamental aspects of synthetic biology. Journal of Cell Biology, 210(5):687–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
210/5/687.

Schlett:2017:MTM

[Sch17a] Katalin Schlett. More than a mere supply of monomers: G– Actin pools regulate actin dynamics in dendritic spines. Journal of Cell Biology, 216(8):2255–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/8/2255.

Schmid:2017:RRS

[Sch17b] Sandra L. Schmid. Reciprocal regulation of signaling and endocytosis: Implications for the evolving cancer cell. Journal of Cell Biology, 216(9):2623–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/9/2623.

Schlacher:2019:SSA

[Sch19] Katharina Schlacher. Sense and sensibility: ATM oxygen stress signaling manages brain cell energetics. Journal of Cell Biology, 218(3):732–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/3/732.

Skouloudaki:2019:YCT

[SCK^{+}19] Kassiani Skouloudaki, Ioannis Christodoulou, Dilan Khalili, Vasilios Tsarouhas, Christos Samakovlis, Pavel Tomancak, Elisabeth Knust, and Dimitrios K. Papadopoulos. Yorkie controls tube length and apical barrier integrity during airway development. Journal of Cell Biology, 218(8):2762–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/8/

2762. See correction [SCK^{+}23].

Skouloudaki:2023:CYC

[SCK^{+}23] Kassiani Skouloudaki, Ioannis Christodoulou, Dilan Khalili, Vasilios Tsarouhas, Christos Samakovlis, Pavel Tomancak, Elisabeth Knust, and Dimitrios K. Papadopoulos. Correction: Yorkie controls tube length and apical barrier integrity during airway development. Journal of Cell Biology, 222 (5):e20180912104072023c, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/222/5/e20180912104072023c/
214029/Correction-Yorkie-controls-tube-length-and-apical.

See [SCK^{+}19].

Su:2016:CDM

[SCL^{+}16] Jianmin Su, Jiang Chen, Kumiko Lippold, Aboozar Monavarfeshani, Gabriela Lizana Carrillo, Rachel Jenkins, and Michael A. Fox. Collagen-derived matricryptins promote inhibitory nerve terminal formation in the developing neocortex. Journal of Cell Biology, 212(6):721–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/6/721.

Suzuki:2019:BSS

[SCL^{+}19] Sho W. Suzuki, Ya-Shan Chuang, Ming Li, Matthew N. J. Seaman, and Scott D. Emr. A bipartite sorting signal ensures specificity of retromer complex in membrane protein recycling. Journal of Cell Biology, 218(9):2876–??, September 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/9/2876.

Sato-Carlton:2018:PSC

[SCNTC^{+}18] Aya Sato-Carlton, Chihiro Nakamura-Tabuchi, Stephane Kazuki Chartrand, Tomoki Uchino, and Peter Mark Carlton. Phosphorylation of the synaptonemal complex protein SYP-1 promotes meiotic chromosome segregation. Journal of Cell Biology, 217 (2):555–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/2/555.

Stark:2015:EAP

[SCP^{+}15] Danny A. Stark, Nathan J. Coffey, Hannah R. Pancoast, Laura L. Arnold, J. Peyton D. Walker, Joanne Vallée, Richard Robitaille, Michael L. Garcia, and D. D. W. Cornelison. Ephrin-A3 promotes and maintains slow muscle fiber identity during postnatal development and reinnervation. Journal of Cell Biology, 211(5):1077–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/5/1077.

Stik:2017:EVS

[SCP^{+}17] Gregoire Stik, Simon Crequit, Laurence Petit, Jennifer Durant, Pierre Charbord, Thierry Jaffredo, and Charles Durand. Extracellular vesicles of stromal origin target and support hematopoietic stem and progenitor cells. Journal of Cell Biology, 216 (7):2217–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/7/2217.

Sawyer:2016:NBB

[SD16a] Iain A. Sawyer and Miroslav Dundr. Nuclear bodies: Built to boost. Journal of Cell Biology, 213(5):509–??, June 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/213/5/509.

Solvik:2016:CAI

[SD16b] Tina Solvik and Jayanta Debnath. At the crossroads of autophagy and infection: Noncanonical roles for ATG proteins in viral replication. Journal of Cell Biology, 214(5):503–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
214/5/503.

Senaratne:2017:AOD

[SD17] Aruni P. Senaratne and Ines A. Drinnenberg. All that is old does not wither: Conservation of outer kinetochore proteins across all eukaryotes? Journal of Cell Biology, 216(2): 291–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/2/291.

Sharma:2019:CAC

[SD19] Mahak Sharma and Devashish Dwivedi. A CRACker of an adaptor connects dynein-mediated transport to calcium signaling. Journal of Cell Biology, 218(5):1429–??, May 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/5/1429.

Simonetti:2017:SDC

[SDHC17] Boris Simonetti, Chris M. Danson, Kate J. Heesom, and Peter J. Cullen. Sequence-dependent cargo recognition by SNX– BARs mediates retromer-independent transport of CI–MPR. Journal of Cell Biology, 216(11):3695–??, November 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/11/3695.

Saez:2019:ECL

[SDI^{+}19] Juan José Sáez, Jheimmy Diaz, Jorge Ibañez, Juan Pablo Bozo, Fernanda Cabrera Reyes, Martina Alamo, François-Xavier Gobert, Dorian Obino, María Rosa Bono, Ana-María Lennon-Duménil, Charles Yeaman, and María-Isabel Yuseff. The exocyst controls lysosome secretion and antigen extraction at the immune synapse of B cells. Journal of Cell Biology, 218(7):2247–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/7/2247.

Strale:2015:FONa

[SDP^{+}15a] Pierre-Olivier Strale, Laurence Duchesne, Grégoire Peyret, Lorraine Montel, Thao Nguyen, Evelyn Png, Robert Tampé, Sergey Troyanovsky, Sylvie Hénon, Benoit Ladoux, and René-Marc Mège. The formation of ordered nanoclusters controls cadherin anchoring to actin and cell–cell contact fluidity. Journal of Cell Biology, 210(2):333–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/210/2/333.

Strale:2015:FONb

[SDP^{+}15b] Pierre-Olivier Strale, Laurence Duchesne, Grégoire Peyret, Lorraine Montel, Thao Nguyen, Evelyn Png, Robert Tampé, Sergey Troyanovsky, Sylvie Hénon, Benoit Ladoux, and René-Marc Mège. The formation of ordered nanoclusters controls cadherin anchoring to actin and cell–cell contact fluidity. Journal of Cell Biology, 210(6):1033–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/210/6/1033.

Sohr:2019:DFC

[SDW^{+}19] Alex Sohr, Lijuan Du, Ruofan Wang, Li Lin, and Sougata Roy. Drosophila FGF cleavage is required for efficient intracellular sorting and intercellular dispersal. Journal of Cell Biology, 218 (5):1653–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/5/1653.

Suzuki:2018:MPR

[SE18] Sho W. Suzuki and Scott D. Emr. Membrane protein recycling from the vacuole/lysosome membrane. Journal of Cell Biology, 217(5):1623–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/5/1623.

Shatz:2019:ARB

[SE19] Oren Shatz and Zvulun Elazar. ATG9 raises the BAR for PI4P in autophagy. Journal of Cell Biology, 218(5):1432–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/5/
1432.

Sedwick:2015:ABG

[Sed15a] Caitlin Sedwick. Anne Brunet: Gracefully studying how we age. Journal of Cell Biology, 211(3):488–??, November 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/211/3/488.

Sedwick:2015:CGM

[Sed15b] Caitlin Sedwick. Cayetano González: Mothers, daughters, stemness, and cancer. Journal of Cell Biology, 208(3):254– ??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/208/3/254.

Sedwick:2015:CCP

[Sed15c] Caitlin Sedwick. Chaperones caught partying with prions. Journal of Cell Biology, 211(1):??, October 2015. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/211/1/2.2.

Sedwick:2015:CZG

[Sed15d] Caitlin Sedwick. Chiara Zurzolo: GPI knows the way to go. Journal of Cell Biology, 208(7):862–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/7/862.

Sedwick:2015:CWE

[Sed15e] Caitlin Sedwick. Clearing the way for efficient stem cell differentiation. Journal of Cell Biology, 210(7):??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/210/7/
1036.3.

Sedwick:2015:DAB

[Sed15f] Caitlin Sedwick. Demystifying Aurora B. Journal of Cell Biology, 209(2):189–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/2/189.

Sedwick:2015:EGI

[Sed15g] Caitlin Sedwick. Elena Gracheva: Ion channels run hot and cold. Journal of Cell Biology, 209(6):778–??, June 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/209/6/778.

Sedwick:2015:ESS

[Sed15h] Caitlin Sedwick. Elias Spiliotis: Septins set it up. Journal of Cell Biology, 210(4):524–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/4/524.

Sedwick:2015:ENS

[Sed15i] Caitlin Sedwick. Eva Nogales: See how they run. Journal of Cell Biology, 209(4):472–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/209/4/472.

Sedwick:2015:FSG

[Sed15j] Caitlin Sedwick. Feng Shao: Getting a sense for the defense. Journal of Cell Biology, 210(2):174–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/210/2/174.

Sedwick:2015:GHI

[Sed15k] Caitlin Sedwick. Glia hold it together. Journal of Cell Biology, 210(7):??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/7/1036.1.

Sedwick:2015:GCC

[Sed15l] Caitlin Sedwick. Glycolytic cancer cells splice their way out of trouble. Journal of Cell Biology, 210(7):1037–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/210/7/
1037.

Sedwick:2015:GGP

[Sed15m] Caitlin Sedwick. Growing up and growing pores in myoblast fusion. Journal of Cell Biology, 211(1):??, October 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/211/1/2.3.

Sedwick:2015:HSC

[Sed15n] Caitlin Sedwick. How stressed cells triage mRNAs. Journal of Cell Biology, 211(5):940–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/211/5/940.

Sedwick:2015:JFF

[Sed15o] Caitlin Sedwick. Jeffrey Fredberg: Flow under pressure. Journal of Cell Biology, 210(6):868–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/210/6/868.

Sedwick:2015:JBS

[Sed15p] Caitlin Sedwick. Julia von Blume: Sorting through the trans-Golgi. Journal of Cell Biology, 208(1):4–??, January 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/208/1/4.

Sedwick:2015:LJD

[Sed15q] Caitlin Sedwick. Lin He: “Junk” DNA isn’t. Journal of Cell Biology, 211(1):4–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/1/4.

Sedwick:2015:MRZ

[Sed15r] Caitlin Sedwick. Matthew Rasband: Zen and the art of axonal maintenance. Journal of Cell Biology, 209(2):186–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/209/2/
186.

Sedwick:2015:MSW

[Sed15s] Caitlin Sedwick. Mitochondria sense the way forward. Journal of Cell Biology, 210(7):??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/7/1036.2.

Sedwick:2015:MTB

[Sed15t] Caitlin Sedwick. mTORC2 tips the balance in cell survival. Journal of Cell Biology, 211(1):??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/211/1/2.1.

Sedwick:2015:RHW

[Sed15u] Caitlin Sedwick. Rick Horwitz: Words do not suffice. Journal of Cell Biology, 208(5):496–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/208/5/496.

Sedwick:2015:SCC

[Sed15v] Caitlin Sedwick. A spotlight on chromatin choreography. Journal of Cell Biology, 210(2):176–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/210/2/176.

Sedwick:2015:TPP

[Sed15w] Caitlin Sedwick. Thomas Pucadyil: Piecing together membrane fission. Journal of Cell Biology, 211(4):720–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/211/4/720.

Sedwick:2015:TSS

[Sed15x] Caitlin Sedwick. Thy-1: Shade and sun for integrin signaling. Journal of Cell Biology, 211(1):3–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/211/1/3.

Sedwick:2016:JFM

[Sed16a] Caitlin Sedwick. Jessica Feldman: Microtubule-organizing function dives off centrosomes. Journal of Cell Biology, 212 (5):484–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/5/484.

Sedwick:2016:JKE

[Sed16b] Caitlin Sedwick. Jim Kadonaga: Exploring transcription and chromatin. Journal of Cell Biology, 212(6):608–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/212/6/608.

Sedwick:2016:MSS

[Sed16c] Caitlin Sedwick. Maya Schuldiner: The systems that define us. Journal of Cell Biology, 213(1):3–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/213/1/3.

Sedwick:2016:NBC

[Sed16d] Caitlin Sedwick. Needhi Bhalla: Chromosomes do the most amazing things. Journal of Cell Biology, 212(3):260–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/212/3/
260.

Sedwick:2016:SDU

[Sed16e] Caitlin Sedwick. Sharon Dent: The unfolding SAGA of chromatin-modifying proteins. Journal of Cell Biology, 212 (1):4–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/1/4.

Siegenthaler:2015:LNC

[SEMP15] Dominique Siegenthaler, Eva-Maria Enneking, Eliza Moreno, and Jan Pielage. L1CAM/ neuroglian controls the axon–axon interactions establishing layered and lobular mushroom body architecture. Journal of Cell Biology, 208(7):1003–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/208/7/
1003.

Somasekharan:2015:YRS

[SENL^{+}15] Syam Prakash Somasekharan, Amal El-Naggar, Gabriel Leprivier, Hongwei Cheng, Shamil Hajee, Thomas G. P. Grunewald, Fan Zhang, Tony Ng, Olivier Delattre, Valentina Evdokimova, Yuzhuo Wang, Martin Gleave, and Poul H. Sorensen. YB-1 regulates stress granule formation and tumor progression by translationally activating G3BP1. Journal of Cell Biology, 208 (7):913–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/208/7/913.

Seybold:2015:KBS

[SER^{+}15] Christian Seybold, Menattallah Elserafy, Diana Rüthnick, Musa Ozboyaci, Annett Neuner, Benjamin Flottmann, Mike Heilemann, Rebecca C. Wade, and Elmar Schiebel. Kar1 binding to Sfi1 C-terminal regions anchors the SPB bridge to the nuclear envelope. Journal of Cell Biology, 209(6):843–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/209/6/
843.

Salogiannis:2016:PMH

[SERP16] John Salogiannis, Martin J. Egan, and Samara L. Reck-Peterson. Peroxisomes move by hitchhiking on early endosomes using the novel linker protein PxdA. Journal of Cell Biology, 212(3):289–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/3/289.

Scholz:2019:CCB

[SES^{+}19] Nicole Scholz, Nadine Ehmann, Divya Sachidanandan, Cordelia Imig, Benjamin H. Cooper, Olaf Jahn, Kerstin Reim, Nils Brose, Jutta Meyer, Marius Lamberty, Steffen Altrichter, Anne Bormann, Stefan Hallermann, Martin Pauli, Manfred Heckmann, Christian Stigloher, Tobias Langenhan, and Robert J. Kittel. Complexin cooperates with Bruchpilot to tether synaptic vesicles to the active zone cytomatrix. Journal of Cell Biology, 218(3):1011–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/3/1011.

Schoenfelder:2015:EIP

[SF15] Kevin P. Schoenfelder and Donald T. Fox. The expanding implications of polyploidy. Journal of Cell Biology, 209(4):485–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
209/4/485.

Strothman:2019:MME

[SFA^{+}19] Claire Strothman, Veronica Farmer, Göker Arpağ, Nicole Rodgers, Marija Podolski, Stephen Norris, Ryoma Ohi, and Marija Zanic. Microtubule minus-end stability is dictated by the tubulin off-rate. Journal of Cell Biology, 218(9):2841– ??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/9/2841.

Schmidt:2017:TPC

[SFG^{+}17] Ruben Schmidt, Lars-Eric Fielmich, Ilya Grigoriev, Eugene A. Katrukha, Anna Akhmanova, and Sander van den Heuvel. Two populations of cytoplasmic dynein contribute to spindle positioning in C. elegans embryos. Journal of Cell Biology, 216(9): 2777–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/9/2777.

Stroud:2017:NEM

[SFZ^{+}17] Matthew J. Stroud, Wei Feng, Jianlin Zhang, Jennifer Veevers, Xi Fang, Larry Gerace, and Ju Chen. Nesprin 1α2 is essential for mouse postnatal viability and nuclear positioning in skeletal muscle. Journal of Cell Biology, 216(7):1915–??, July 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/7/1915.

Shimada:2017:GMC

[SG17] Kenji Shimada and Susan M. Gasser. A game of musical chairs: Pro- and anti-resection factors compete for TOPBP1 binding after DNA damage. Journal of Cell Biology, 216(3):535–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/3/535.

Segev:2018:CSR

[SG18a] Nadav Segev and Jeffrey E. Gerst. Correction: Specialized ribosomes and specific ribosomal protein paralogs control translation of mitochondrial proteins. Journal of Cell Biology, 217 (3):1155–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/3/1155. See [SG18b].

Segev:2018:SRS

[SG18b] Nadav Segev and Jeffrey E. Gerst. Specialized ribosomes and specific ribosomal protein paralogs control translation of mitochondrial proteins. Journal of Cell Biology, 217(1):117–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/

217/1/117. See correction [SG18a].

Schrank:2019:AND

[SG19] Benjamin Schrank and Jean Gautier. Assembling nuclear domains: Lessons from DNA repair. Journal of Cell Biology, 218 (8):2444–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/8/2444.

Sun:2017:MSA

[SGB^{+}17] Gongping Sun, Elmer Guzman, Varuzhan Balasanyan, Christopher M. Conner, Kirsten Wong, Hongjun Robin Zhou, Kenneth S. Kosik, and Denise J. Montell. A molecular signature for anastasis, recovery from the brink of apoptotic cell death. Journal of Cell Biology, 216(10):3355–??, October 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/10/3355.

Sun:2016:IMM

[SGF16] Zhiqi Sun, Shengzhen S. Guo, and Reinhard Fässler. Integrinmediated mechanotransduction. Journal of Cell Biology, 215 (4):445–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/215/4/445.

Saha:2017:RGP

[SH17] Shambaditya Saha and Anthony A. Hyman. RNA gets in phase. Journal of Cell Biology, 216(8):2235–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/8/2235.

Strauss:2018:CBE

[SHC^{+}18] Bernhard Strauss, Andrew Harrison, Paula Almeida Coelho, Keiko Yata, Magdalena Zernicka-Goetz, and Jonathon Pines. Cyclin B1 is essential for mitosis in mouse embryos, and its nuclear export sets the time for mitosis. Journal of Cell Biology, 217(1):179–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/1/179.

Sherwood:2015:DBO

[She15] David R. Sherwood. A developmental biologist’s “outside-the-cell” thinking. Journal of Cell Biology, 210(3):369–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/210/3/
369.

Sinha:2016:CPE

[SHH^{+}16] Seema Sinha, Daisuke Hoshino, Nan Hyung Hong, Kellye C. Kirkbride, Nathan E. Grega-Larson, Motoharu Seiki, Matthew J. Tyska, and Alissa M. Weaver. Cortactin promotes exosome secretion by controlling branched actin dynamics. Journal of Cell Biology, 214(2):197–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/2/197.

Short:2015:AHB

[Sho15a] Ben Short. Actin hotspots blaze a trail along axons. Journal of Cell Biology, 210(3):365–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/3/365.

Short:2015:AWB

[Sho15b] Ben Short. Actin works both sides of the immunological synapse. Journal of Cell Biology, 208(4):383–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/208/4/383.

Short:2015:ACN

[Sho15c] Ben Short. Adding complexity to the nuclear pore complex. Journal of Cell Biology, 208(6):651–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/6/651.

Short:2015:AAG

[Sho15d] Ben Short. Assessing actin’s growth rate. Journal of Cell Biology, 208(5):499–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/5/499.

Short:2015:AAP

[Sho15e] Ben Short. Asterless amplifies Plk4. Journal of Cell Biology, 208(4):??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/208/4/382.1.

Short:2015:AAU

[Sho15f] Ben Short. Axonal autophagosomes use a rideshare service. Journal of Cell Biology, 209(3):??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/209/3/318.2.

Short:2015:BCH

[Sho15g] Ben Short. B cell, heal thyself. Journal of Cell Biology, 211 (6):??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/6/1098.3.

Short:2015:BBN

[Sho15h] Ben Short. Building a bridge to a new spindle pole body. Journal of Cell Biology, 209(6):??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/209/6/777.1.

Short:2015:CCS

[Sho15i] Ben Short. Cadherin clusters solidify adhesions. Journal of Cell Biology, 210(2):??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/2/172.3.

Short:2015:CSR

[Sho15j] Ben Short. Chromosome search and rescue. Journal of Cell Biology, 210(3):??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/3/364.1.

Short:2015:CSK

[Sho15k] Ben Short. A cohort study of kinetochore proteins. Journal of Cell Biology, 211(6):??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/211/6/1098.1.

Short:2015:DMG

[Sho15l] Ben Short. Damaged mitochondria get a Parkin ticket. Journal of Cell Biology, 208(7):865–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/208/7/865.

Short:2015:DCK

[Sho15m] Ben Short. Dense collagen kindles invadopodia formation. Journal of Cell Biology, 208(3):??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/3/252.3.

Short:2015:DDM

[Sho15n] Ben Short. Dimerization dictates the message. Journal of Cell Biology, 209(4):471–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/4/471.

Short:2015:EAT

[Sho15o] Ben Short. Ephrin-A3 typecasts muscle fibers. Journal of Cell Biology, 211(5):??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/5/936.3.

Short:2015:FBC

[Sho15p] Ben Short. The first buds of Cdc42. Journal of Cell Biology, 209(6):780–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/6/780.

Short:2015:FYN

[Sho15q] Ben Short. For fission yeast nuclei, it takes two to tango. Journal of Cell Biology, 209(1):??, April 2015. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/209/1/2.2.

Short:2015:GGA

[Sho15r] Ben Short. Gauging Gag assembly. Journal of Cell Biology, 210(4):??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/4/522.3.

Short:2015:GPA

[Sho15s] Ben Short. GMFβ prunes actin branches. Journal of Cell Biology, 209(6):776–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/6/776.

Short:2015:HIG

[Sho15t] Ben Short. Hedgehog induces a gut reaction. Journal of Cell Biology, 208(6):??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/6/650.3.

Short:2015:HSK

[Sho15u] Ben Short. How the spindle keeps its focus. Journal of Cell Biology, 211(5):937–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/5/937.

Short:2015:HVC

[Sho15v] Ben Short. How VE-cadherin goes with the flow. Journal of Cell Biology, 208(7):861–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/208/7/861.

Short:2015:HGS

[Sho15w] Ben Short. Hsp72 gives spindles a fiber supplement. Journal of Cell Biology, 209(3):??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/3/318.1.

Short:2015:IGT

[Sho15x] Ben Short. Inflammation gets TRIM’d back by autophagy. Journal of Cell Biology, 210(6):??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/210/6/866.2.

Short:2015:IST

[Sho15y] Ben Short. Integrin signaling tranquilizes Hippo. Journal of Cell Biology, 210(3):??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/3/364.3.

Short:2015:KKT

[Sho15z] Ben Short. Kinases KISS and tell. Journal of Cell Biology, 209(6):??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/6/777.3.

Short:2015:KET

[Sho15-27] Ben Short. Kinetochores expand their reach. Journal of Cell Biology, 210(6):867–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/6/867.

Short:2015:LAB

[Sho15-28] Ben Short. Lamellipodial actin branches out. Journal of Cell Biology, 211(1):6–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/1/6.

Short:2015:LPS

[Sho15-29] Ben Short. LSR puts its seal on brain endothelial cells. Journal of Cell Biology, 208(6):??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/6/650.1.

Short:2015:LGA

[Sho15-30] Ben Short. Ltc1 gets added to the ER’s contacts. Journal of Cell Biology, 209(4):??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/4/470.1.

Short:2015:MNP

[Sho15-31] Ben Short. Mapping out the nuclear pore complex. Journal of Cell Biology, 208(3):257–??, February 2015. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/208/3/257.

Short:2015:MHV

[Sho15-32] Ben Short. Mdm1 helps the ER and vacuole stay in touch. Journal of Cell Biology, 210(4):??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/4/522.1.

Short:2015:MCS

[Sho15-33] Ben Short. Merlin casts its spell on the cortical cytoskeleton. Journal of Cell Biology, 211(2):207–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/211/2/207.

Short:2015:MLH

[Sho15-34] Ben Short. miR-7 loss is hard to stomach. Journal of Cell Biology, 210(4):??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/4/522.2.

Short:2015:MLW

[Sho15-35] Ben Short. Misdirection leads the way to vesicle transport proteins. Journal of Cell Biology, 208(3):??, February 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/208/3/252.1.

Short:2015:MRP

[Sho15-36] Ben Short. Mismatch repair provides a security patch for translesion synthesis. Journal of Cell Biology, 209(1):??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/209/1/
2.1.

Short:2015:MKR

[Sho15-37] Ben Short. Mitofusin 2 keeps the respiratory chain on Q. Journal of Cell Biology, 208(4):??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/4/382.3.

Short:2015:MST

[Sho15-38] Ben Short. mRNAs scan themselves out. Journal of Cell Biology, 211(6):1099–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/6/1099.

Short:2015:MGG

[Sho15-39] Ben Short. Myelin goes to great pains to regulate NGF. Journal of Cell Biology, 210(6):??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/6/866.1.

Short:2015:NES

[Sho15-40] Ben Short. NORE1A erects a senescence barrier to tumorigenesis. Journal of Cell Biology, 208(6):??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/6/650.2.

Short:2015:NCC

[Sho15-41] Ben Short. Notch commandeers COMMD9 for endosomal sorting. Journal of Cell Biology, 211(3):??, November 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/211/3/486.3.

Short:2015:NRD

[Sho15-42] Ben Short. Num1 relieves dynein’s inhibitions. Journal of Cell Biology, 211(2):??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/2/206.2.

Short:2015:NCP

[Sho15-43] Ben Short. Nup82 complex puts the P in NPC. Journal of Cell Biology, 208(3):??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/3/252.2.

Short:2015:PPL

[Sho15-44] Ben Short. P2X4 promotes lysosome fusion. Journal of Cell Biology, 209(6):??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/6/777.2.

Short:2015:PPG

[Sho15-45] Ben Short. PI(4)P gets Sac-rificed in the name of endocytic recycling. Journal of Cell Biology, 209(1):3–??, April 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/209/1/3.

Short:2015:PAS

[Sho15-46] Ben Short. A POSH accent for synaptic growth. Journal of Cell Biology, 208(7):??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/7/860.3.

Short:2015:PTA

[Sho15-47] Ben Short. Progranulin takes an alternate route to lysosomes. Journal of Cell Biology, 210(6):??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/210/6/866.3.

Short:2015:RBP

[Sho15-48] Ben Short. Rab27a builds a platform for HIV-1. Journal of Cell Biology, 209(3):??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/3/318.3.

Short:2015:RNN

[Sho15-49] Ben Short. Ribosomes have a NAC for nascent chain processing. Journal of Cell Biology, 210(2):??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/210/2/172.2.

Short:2015:RRM

[Sho15-50] Ben Short. A Ror recruit to mammary gland development. Journal of Cell Biology, 208(3):253–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/3/253.

Short:2015:SCL

[Sho15-51] Ben Short. Secretase complexes let it RIP. Journal of Cell Biology, 211(6):??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/6/1098.2.

Short:2015:SET

[Sho15-52] Ben Short. sisRNAs engage with their hosts. Journal of Cell Biology, 211(2):??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/2/206.1.

Short:2015:SWT

[Sho15-53] Ben Short. A SMRTer way to track molecules. Journal of Cell Biology, 209(4):??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/4/470.3.

Short:2015:SSA

[Sho15-54] Ben Short. SNAREs share atlastin’s burden. Journal of Cell Biology, 210(3):??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/3/364.2.

Short:2015:SVR

[Sho15-55] Ben Short. Solidifying the view of RNP dynamics. Journal of Cell Biology, 211(3):487–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/211/3/487.

Short:2015:SEF

[Sho15-56] Ben Short. Sorting out endosome form and function. Journal of Cell Biology, 210(6):870–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/210/6/870.

Short:2015:SDA

[Sho15-57] Ben Short. Special delivery to the apical membrane. Journal of Cell Biology, 211(3):??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/211/3/486.2.

Short:2015:SAB

[Sho15-58] Ben Short. Spindle assembly at a BRISC pace. Journal of Cell Biology, 210(2):??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/2/172.1.

Short:2015:SPB

[Sho15-59] Ben Short. The spindle plays both ends. Journal of Cell Biology, 210(4):526–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/4/526.

Short:2015:SGL

[Sho15-60] Ben Short. Spindly gets a lipid link to kinetochores. Journal of Cell Biology, 208(7):??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/7/860.1.

Short:2015:SLS

[Sho15-61] Ben Short. Splicing limits the spread of Crumbs. Journal of Cell Biology, 211(5):??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/5/936.2.

Short:2015:SCN

[Sho15-62] Ben Short. A stem cell niche shows self-restraint. Journal of Cell Biology, 209(4):??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/4/470.2.

Short:2015:SER

[Sho15-63] Ben Short. SUMO enters the ring. Journal of Cell Biology, 211 (3):490–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/3/490.

Short:2015:SPS

[Sho15-64] Ben Short. SUN protection for the skin. Journal of Cell Biology, 209(3):319–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/3/319.

Short:2015:TMC

[Sho15-65] Ben Short. The tension mounts at centromeric loops. Journal of Cell Biology, 210(4):523–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/4/523.

Short:2015:TE

[Sho15-66] Ben Short. A timeline for exocytosis. Journal of Cell Biology, 210(2):173–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/2/173.

Short:2015:TKC

[Sho15-67] Ben Short. Tmem231 keeps cilia in the zone. Journal of Cell Biology, 209(1):??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/1/2.3.

Short:2015:TAL

[Sho15-68] Ben Short. TRIMming γ-actin levels improves the memory. Journal of Cell Biology, 211(3):??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/211/3/486.1.

Short:2015:TST

[Sho15-69] Ben Short. Trypanosomes show their commitment. Journal of Cell Biology, 211(2):??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/211/2/206.3.

Short:2015:TVS

[Sho15-70] Ben Short. TUNEL vision spots apoptotic cells. Journal of Cell Biology, 208(1):7–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/1/7.

Short:2015:VTH

[Sho15-71] Ben Short. VEGF tips its hand in angiogenesis. Journal of Cell Biology, 209(4):474–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/4/474.

Short:2015:VSR

[Sho15-72] Ben Short. A vitamin supplement for remyelination. Journal of Cell Biology, 211(5):??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/211/5/936.1.

Short:2015:WCF

[Sho15-73] Ben Short. When centromeres fill in for telomeres. Journal of Cell Biology, 208(4):??, February 2015. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/208/4/382.2.

Stoops:2015:PRP

[SHO^{+}15-74] Emily H. Stoops, Michael Hull, Christina Olesen, Kavita Mistry, Jennifer L. Harder, Felix Rivera-Molina, Derek Toomre, and Michael J. Caplan. The periciliary ring in polarized epithelial cells is a hot spot for delivery of the apical protein gp135. Journal of Cell Biology, 211(2):287–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/211/2/287.

Short:2016:AWM

[Sho16a] Ben Short. Arf6 wins the MVP award. Journal of Cell Biology, 213(1):2–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/1/2.

Short:2016:APP

[Sho16b] Ben Short. Ataxin-3 phosphorylation protects neurons. Journal of Cell Biology, 212(4):??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/4/366.3.

Short:2016:BGS

[Sho16c] Ben Short. A basic guide to stem cell differentiation. Journal of Cell Biology, 215(3):293–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/215/3/293.

Short:2016:CSS

[Sho16d] Ben Short. Cells set sail after lifting anchor from Myo1E. Journal of Cell Biology, 214(4):359–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/4/359.

Short:2016:CHC

[Sho16e] Ben Short. Chloride helps collagen build its network. Journal of Cell Biology, 213(4):404–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/213/4/404.

Short:2016:CRM

[Sho16f] Ben Short. Cytokinetic rings mind the gap. Journal of Cell Biology, 215(6):749–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/6/749.

Short:2016:DCR

[Sho16g] Ben Short. Decrypting a collagen’s role in schizophrenia. Journal of Cell Biology, 212(6):??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/6/606.3.

Short:2016:DDD

[Sho16h] Ben Short. Disrupted desmosomes drive fibrosis. Journal of Cell Biology, 212(4):??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/4/366.2.

Short:2016:GLC

[Sho16i] Ben Short. GTSE1 leads cancer cells into CIN. Journal of Cell Biology, 215(5):593–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/215/5/593.

Short:2016:HAF

[Sho16j] Ben Short. How asters find their center. Journal of Cell Biology, 212(7):743–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/7/743.

Short:2016:HGE

[Sho16k] Ben Short. How Gpr161 exits from cilia. Journal of Cell Biology, 212(7):??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/7/742.3.

Short:2016:HMT

[Sho16l] Ben Short. How morphine tips the synaptic balance. Journal of Cell Biology, 215(2):139–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/215/2/139.

Short:2016:ISS

[Sho16m] Ben Short. Insulin sends SEC16A packing. Journal of Cell Biology, 214(1):1–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/214/1/1.

Short:2016:LSS

[Sho16n] Ben Short. Lymphocytes shrink to stay fit. Journal of Cell Biology, 212(4):367–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/4/367.

Short:2016:MKD

[Sho16o] Ben Short. Mapping the kinetochore Delta. Journal of Cell Biology, 212(3):259–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/3/259.

Short:2016:MFC

[Sho16p] Ben Short. Mitochondrial fission is crucial for cristae remodeling. Journal of Cell Biology, 212(5):??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/212/5/482.1.

Short:2016:MCG

[Sho16q] Ben Short. Mitotic cells get a stress test. Journal of Cell Biology, 214(2):121–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/2/121.

Short:2016:MLP

[Sho16r] Ben Short. MLL5 limits PLK1 aggregation. Journal of Cell Biology, 212(7):??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/7/742.1.

Short:2016:NHA

[Sho16s] Ben Short. NBR1 helps autophagosomes take a bite out of focal adhesions. Journal of Cell Biology, 212(5):??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/212/5/
482.3.

Short:2016:NKC

[Sho16t] Ben Short. Ndel1 keeps cilia retracted. Journal of Cell Biology, 212(4):??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/4/366.1.

Short:2016:NLI

[Sho16u] Ben Short. Neurons let it slide. Journal of Cell Biology, 213 (3):290–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/3/290.

Short:2016:PAP

[Sho16v] Ben Short. p62 aggregation is a problem for neural stem cells. Journal of Cell Biology, 212(5):??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/212/5/482.2.

Short:2016:PRI

[Sho16w] Ben Short. A program to reduce inequality. Journal of Cell Biology, 214(3):233–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/3/233.

Short:2016:PHP

[Sho16x] Ben Short. PxdA helps peroxisomes hitch a ride. Journal of Cell Biology, 212(3):??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/3/258.1.

Short:2016:RRC

[Sho16y] Ben Short. Rac and Rho compete to cooperate. Journal of Cell Biology, 215(4):433–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/215/4/433.

Short:2016:RAS

[Sho16z] Ben Short. A reason to avoid SNAP judgments. Journal of Cell Biology, 215(1):1–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/1/1.

Short:2016:RCP

[Sho16-27] Ben Short. ROBO-cop protects cells from stiff environment. Journal of Cell Biology, 212(6):??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/212/6/606.2.

Short:2016:SHS

[Sho16-28] Ben Short. Slit holds a strange attraction for filopodia. Journal of Cell Biology, 213(2):138–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/213/2/138.

Short:2016:SPB

[Sho16-29] Ben Short. Sorting out peroxisome biogenesis. Journal of Cell Biology, 212(3):??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/3/258.3.

Short:2016:SF

[Sho16-30] Ben Short. Survival of the fattest. Journal of Cell Biology, 212 (6):607–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/6/607.

Short:2016:SSG

[Sho16-31] Ben Short. A switch for stress granule assembly. Journal of Cell Biology, 212(7):??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/7/742.2.

Short:2016:SSS

[Sho16-32] Ben Short. Switching septins spurs sporulation. Journal of Cell Biology, 212(5):483–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/212/5/483.

Short:2016:TCS

[Sho16-33] Ben Short. Telomeres come up short in heart regeneration. Journal of Cell Biology, 213(5):497–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/213/5/497.

Short:2016:TSG

[Sho16-34] Ben Short. Three speeds of gene repositioning. Journal of Cell Biology, 212(6):??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/212/6/606.1.

Short:2016:TCT

[Sho16-35] Ben Short. Tracking CRISPR targeting. Journal of Cell Biology, 214(5):491–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/5/491.

Short:2016:UTI

[Sho16-36] Ben Short. Untangling Topo II’s function at mitotic centromeres. Journal of Cell Biology, 213(6):601–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/213/6/601.

Short:2016:YCT

[Sho16-37] Ben Short. Yeast chromatin takes global action on carbon. Journal of Cell Biology, 212(3):??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/212/3/258.2.

Short:2017:AHM

[Sho17a] Ben Short. ATF4 helps mitochondria pass the stress test. Journal of Cell Biology, 216(7):1865–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/7/1865.

Short:2017:CSP

[Sho17b] Ben Short. Centrosome signaling pathways consult on their decision. Journal of Cell Biology, 216(9):2599–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/9/
2599.

Short:2017:CRR

[Sho17c] Ben Short. Choosing the right response to ER stress. Journal of Cell Biology, 216(6):1501–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/6/1501.

Short:2017:EHN

[Sho17d] Ben Short. ELKS1 helps neuronal synapses diversify. Journal of Cell Biology, 216(4):851–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/4/851.

Short:2017:HVD

[Sho17e] Ben Short. How the vasculature delivers lung epithelia from an incorrect fate. Journal of Cell Biology, 216(10):2991–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/10/2991.

Short:2017:MWM

[Sho17f] Ben Short. Merlin weaves its magic on peripheral nerve repair. Journal of Cell Biology, 216(2):283–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/2/283.

Short:2017:MDG

[Sho17g] Ben Short. Mitochondria deliver a gut check to intestinal stem cells. Journal of Cell Biology, 216(8):2231–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/8/
2231.

Short:2017:SHM

[Sho17h] Ben Short. The signal hypothesis matures with age. Journal of Cell Biology, 216(5):1207–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/5/1207.

Short:2017:TAS

[Sho17i] Ben Short. A three-alarm signal for endocytosis? Journal of Cell Biology, 216(11):3425–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/11/3425.

Short:2017:TPP

[Sho17j] Ben Short. Tubby proteins prove their adaptability. Journal of Cell Biology, 216(3):527–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/3/527.

Short:2017:UHC

[Sho17k] Ben Short. UNC-45a helps cells manage their stress levels. Journal of Cell Biology, 216(12):3887–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/12/3887.

Short:2017:WPT

[Sho17l] Ben Short. Want to promote tissue growth? There’s an app for that! Journal of Cell Biology, 216(1):1–??, January 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/1/1.

Short:2018:CFS

[Sho18a] Ben Short. A cell-free screen of caveolae interactions. Journal of Cell Biology, 217(6):1883–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/6/1883.

Short:2018:GTP

[Sho18b] Ben Short. Gluing together the pieces of crinophagy. Journal of Cell Biology, 217(1):05–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/1/05.

Short:2018:HMS

[Sho18c] Ben Short. How mitotic spindles point to the exit. Journal of Cell Biology, 217(3):795–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/3/795.

Short:2018:PHE

[Sho18d] Ben Short. The proteasome helps epithelial cells set up KAMPs. Journal of Cell Biology, 217(2):431–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/2/431.

Short:2018:SIR

[Sho18e] Ben Short. Seeing the insulin receptor in action. Journal of Cell Biology, 217(5):1555–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/5/1555.

Short:2018:STF

[Sho18f] Ben Short. Src turns FHL1 to the dark side. Journal of Cell Biology, 217(4):1159–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/4/1159.

Sing:2018:BYR

[SHO^{+}18g] Tina L. Sing, Minnie P. Hung, Shinsuke Ohnuki, Godai Suzuki, Bryan-Joseph San Luis, Melainia McClain, Jay R. Unruh, Zulin Yu, Jiongwen Ou, Jesse Marshall-Sheppard, Won-Ki Huh, Michael Costanzo, Charles Boone, Yoshikazu Ohya, Sue L. Jaspersen, and Grant W. Brown. The budding yeast RSC complex maintains ploidy by promoting spindle pole body insertion. Journal of Cell Biology, 217(7):2445–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/7/2445.

Smith:2017:MCD

[SHR17] Jean A. Smith, Allison E. Hall, and Mark D. Rose. Membrane curvature directs the localization of Cdc42p to novel foci required for cell–cell fusion. Journal of Cell Biology, 216(12): 3971–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/12/3971.

Spiess:2018:AII

[SHVO^{+}18] Matthias Spiess, Pablo Hernandez-Varas, Anna Oddone, Helene Olofsson, Hans Blom, Dominic Waithe, John G. Lock, Melike Lakadamyali, and Staffan Strömblad. Active and inactive β1 integrins segregate into distinct nanoclusters in focal adhesions. Journal of Cell Biology, 217(6):1929–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/6/1929.

Saunders:2017:TCT

[SHW^{+}17] Cosmo A. Saunders, Nathan J. Harris, Patrick T. Willey, Brian M. Woolums, Yuexia Wang, Alex J. McQuown, Amy Schoenhofen, Howard J. Worman, William T. Dauer, Gregg G. Gundersen, and G. W. Gant Luxton. TorsinA controls TAN line assembly and the retrograde flow of dorsal perinuclear actin cables during rearward nuclear movement. Journal of Cell Biology, 216(3):657–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/3/657.

Selyunin:2017:GWS

[SIBM17] Andrey S. Selyunin, Lakesla R. Iles, Geoffrey Bartholomeusz, and Somshuvra Mukhopadhyay. Genome-wide siRNA screen identifies UNC50 as a regulator of Shiga toxin 2 trafficking. Journal of Cell Biology, 216(10):3249–??, October 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/10/3249.

Shigematsu:2018:SIM

[SID^{+}18] Hideki Shigematsu, Tsuyoshi Imasaki, Chihiro Doki, Takuya Sumi, Mari Aoki, Tomomi Uchikubo-Kamo, Ayako Sakamoto, Kiyotaka Tokuraku, Mikako Shirouzu, and Ryo Nitta. Structural insight into microtubule stabilization and kinesin inhibition by Tau family MAPs. Journal of Cell Biology, 217(12): 4155–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/12/4155.

Siletti:2016:AML

[Sil16a] Kimberly Siletti. Ana-Maria Lennon-Duménil: a dynamic career. Journal of Cell Biology, 215(1):2–??, October 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/215/1/2.

Siletti:2016:RGP

[Sil16b] Kimberly Siletti. Rusty Gage: a plastic approach to neuroscience. Journal of Cell Biology, 215(6):750–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/215/6/750.

Siletti:2017:RMM

[Sil17] Kimberly Siletti. Roop Mallik: From machines to molecular motors. Journal of Cell Biology, 216(4):852–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/216/4/852.

Scharaw:2016:ETR

[SIO^{+}16] Sandra Scharaw, Murat Iskar, Alessandro Ori, Gaelle Boncompain, Vibor Laketa, Ina Poser, Emma Lundberg, Franck Perez, Martin Beck, Peer Bork, and Rainer Pepperkok. The endosomal transcriptional regulator RNF11 integrates degradation and transport of EGFR. Journal of Cell Biology, 215(4): 543–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/215/4/543.

Sobajima:2018:RBP

[SiYM^{+}18] Tomoaki Sobajima, Shin ichiro Yoshimura, Tomomi Maeda, Haruhiko Miyata, Eiji Miyoshi, and Akihiro Harada. The Rab11-binding protein RELCH/KIAA1468 controls intracellular cholesterol distribution. Journal of Cell Biology, 217 (5):1777–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/5/1777.

Silva:2016:ULI

[SJ16] Nicola Silva and Verena Jantsch. UNC-84: “LINC-ing” chromosome movement and double strand break repair. Journal of Cell Biology, 215(6):753–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/215/6/753.

Sawyer:2019:DRO

[SJJ^{+}19] Eric M. Sawyer, Pallavi R. Joshi, Victoria Jorgensen, Julius Yunus, Luke E. Berchowitz, and Elçin Ünal. Developmental regulation of an organelle tether coordinates mitochondrial remodeling in meiosis. Journal of Cell Biology, 218(2):559–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/2/
559.

Subramanian:2019:CQB

[SJL^{+}19] Kelly Subramanian, Adam Jochem, Maxence Le Vasseur, Samantha Lewis, Brett R. Paulson, Thiruchelvi R. Reddy, Jason D. Russell, Joshua J. Coon, David J. Pagliarini, and Jodi Nunnari. Coenzyme Q biosynthetic proteins assemble in a substrate-dependent manner into domains at ER– mitochondria contacts. Journal of Cell Biology, 218(4):1353–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/4/1353.

Sharir:2016:WDD

[SK16a] Amnon Sharir and Ophir D. Klein. Watching a deep dive: Live imaging provides lessons about tooth invagination. Journal of Cell Biology, 214(6):645–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/6/645.

Sharma:2016:CBI

[SK16b] Deepika Sharma and Thirumala-Devi Kanneganti. The cell biology of inflammasomes: Mechanisms of inflammasome activation and regulation. Journal of Cell Biology, 213(6):617–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
213/6/617.

Sivakumar:2018:TRC

[SK18a] Aravind Sivakumar and Natasza A. Kurpios. Transcriptional regulation of cell shape during organ morphogenesis. Journal of Cell Biology, 217(9):2987–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/9/2987.

Steffen:2018:MCA

[SK18b] Janos Steffen and Carla M. Koehler. ER-mitochondria contacts: Actin dynamics at the ER control mitochondrial fission via calcium release. Journal of Cell Biology, 217(1):15–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/1/
15.

Shimi:2019:MOY

[SK19] Takeshi Shimi and Hiroshi Kimura. A mosaic of old and young nucleoporins. Journal of Cell Biology, 218(2):385–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/2/
385.

Smoyer:2016:AMP

[SKG^{+}16] Christine J. Smoyer, Santharam S. Katta, Jennifer M. Gardner, Lynn Stoltz, Scott McCroskey, William D. Bradford, Melainia McClain, Sarah E. Smith, Brian D. Slaughter, Jay R. Unruh, and Sue L. Jaspersen. Analysis of membrane proteins localizing to the inner nuclear envelope in living cells. Journal of Cell Biology, 215(4):575–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/4/575.

Sidoli:2017:WPN

[SKG17] Simone Sidoli, Katarzyna Kulej, and Benjamin A. Garcia. Why proteomics is not the new genomics and the future of mass spectrometry in cell biology. Journal of Cell Biology, 216(1):21–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/1/21.

Singh:2018:SSM

[SKL^{+}18] Pawan Kishor Singh, Anjali Kapoor, Richa Madan Lomash, Kamal Kumar, Sukrut C. Kamerkar, Thomas J. Pucadyil, and Amitabha Mukhopadhyay. Salmonella SipA mimics a cognate SNARE for host Syntaxin8 to promote fusion with early endosomes. Journal of Cell Biology, 217(12):4199–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/12/4199.

Shima:2019:CVC

[SKN19] Takayuki Shima, Hiromi Kirisako, and Hitoshi Nakatogawa. COPII vesicles contribute to autophagosomal membranes. Journal of Cell Biology, 218(5):1503–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/5/1503.

Son:2015:RMV

[SKO^{+}15] Sungmin Son, Joon Ho Kang, Seungeun Oh, Marc W. Kirschner, T. J. Mitchison, and Scott Manalis. Resonant microchannel volume and mass measurements show that suspended cells swell during mitosis. Journal of Cell Biology, 211 (4):757–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/4/757.

Saraya:2015:PRH

[SKVvdK15] Ruchi Saraya, Arjen M. Krikken, Marten Veenhuis, and Ida J. van der Klei. Peroxisome reintroduction in Hansenula polymorpha requires Pex25 and Rho1. Journal of Cell Biology, 210 (3):519–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/3/519.

Seibert:2019:CMP

[SKW^{+}19] Markus Seibert, Marcus Krüger, Nikolaus A. Watson, Onur Sen, John R. Daum, Johan A. Slotman, Thomas Braun, Adriaan B. Houtsmuller, Gary J. Gorbsky, Ralf Jacob, Michael Kracht, Jonathan M. G. Higgins, and M. Lienhard Schmitz. CDK1-mediated phosphorylation at H2B serine 6 is required for mitotic chromosome segregation. Journal of Cell Biology, 218(4):1164–??, April 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/4/1164.

Sohn:2018:PPC

[SKZ^{+}18a] Mira Sohn, Marek Korzeniowski, James P. Zewe, Rachel C. Wills, Gerald R. V. Hammond, Jana Humpolickova, Lukas Vrzal, Dominika Chalupska, Vaclav Veverka, Gregory D. Fairn, Evzen Boura, and Tamas Balla. PI(4,5)P_{2} controls plasma membrane PI4P and PS levels via ORP5/8 recruitment to ER– PM contact sites. Journal of Cell Biology, 217(5):1797–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/5/
1797.

Song:2018:MHA

[SKZ^{+}18b] Jae-Geun Song, Matthew R. King, Rui Zhang, Rachel S. Kadzik, Akanksha Thawani, and Sabine Petry. Mechanism of how augmin directly targets the γ-tubulin ring complex to microtubules. Journal of Cell Biology, 217(7):2417–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/7/
2417.

Song:2019:PTR

[SL19] Lei Song and Zhao-Qing Luo. Post-translational regulation of ubiquitin signaling. Journal of Cell Biology, 218(6):1776–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/6/1776.

Sikorska:2016:LQC

[SLAR^{+}16] Natalia Sikorska, Leticia Lemus, Auxiliadora Aguilera-Romero, Javier Manzano-Lopez, Howard Riezman, Manuel Muñiz, and Veit Goder. Limited ER quality control for GPI-anchored proteins. Journal of Cell Biology, 213(6):693–??, June 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/213/6/693.

Smith:2015:VSP

[SLD^{+}15] Carlas Smith, Azra Lari, Carina Patrizia Derrer, Anette Ouwehand, Ammeret Rossouw, Maximiliaan Huisman, Thomas Dange, Mark Hopman, Aviva Joseph, Daniel Zenklusen, Karsten Weis, David Grunwald, and Ben Montpetit. In vivo single-particle imaging of nuclear mRNA export in budding yeast demonstrates an essential role for Mex67p. Journal of Cell Biology, 211(6):1121–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/211/6/1121.

Slep:2016:MDR

[Sle16] Kevin C. Slep. A microtubule dynamics reconstitutional convention. Journal of Cell Biology, 215(3):305–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/215/3/
305.

Sun:2018:TIT

[SLG^{+}18] Hao Sun, Frederic Lagarrigue, Alexandre R. Gingras, Zhichao Fan, Klaus Ley, and Mark H. Ginsberg. Transmission of integrin β7 transmembrane domain topology enables gut lymphoid tissue development. Journal of Cell Biology, 217(4):1453–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/4/1453.

Stuart:2017:GHP

[SLH17] Lynda M. Stuart and Adam Lacy-Hulbert. GOP-1: Helping phagosomes pass the acid test. Journal of Cell Biology, 216 (6):1517–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/6/1517.

Sohet:2015:LAT

[SLM^{+}15] Fabien Sohet, Christina Lin, Roeben N. Munji, Seo Yeon Lee, Nadine Ruderisch, Allison Soung, Thomas D. Arnold, Nikita Derugin, Zinaida S. Vexler, Frances T. Yen, and Richard Daneman. LSR/ angulin-1 is a tricellular tight junction protein involved in blood–brain barrier formation. Journal of Cell Biology, 208(6):703–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/6/703.

Su:2019:SNR

[SLPW19] Wei-Cheng Su, Yi-Hsiu Lin, Martin Pagac, and Chao-Wen Wang. Seipin negatively regulates sphingolipid production at the ER–LD contact site. Journal of Cell Biology, 218(11): 3663–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/11/3663.

Sandquist:2018:IBM

[SLW^{+}18] Joshua C. Sandquist, Matthew E. Larson, Sarah Woolner, Zhiwei Ding, and William M. Bement. An interaction between myosin-10 and the cell cycle regulator Wee1 links spindle dynamics to mitotic progression in epithelia. Journal of Cell Biology, 217(3):849–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/3/849.

Scarpa:2016:CCM

[SM16] Elena Scarpa and Roberto Mayor. Collective cell migration in development. Journal of Cell Biology, 212(2):143–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/212/2/
143.

Serebryannyy:2018:PSN

[SM18] Leonid Serebryannyy and Tom Misteli. Protein sequestration at the nuclear periphery as a potential regulatory mechanism in premature aging. Journal of Cell Biology, 217(1):21–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/1/21.

Sugihara:2019:AAU

[SMA^{+}19] Munechika Sugihara, Daisuke Morito, Shiori Ainuki, Yoshinobu Hirano, Kazutoyo Ogino, Akira Kitamura, Hiromi Hirata, and Kazuhiro Nagata. The AAA+ ATPase/ubiquitin ligase mysterin stabilizes cytoplasmic lipid droplets. Journal of Cell Biology, 218(3):??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/3/949.1.

Shibuya:2015:EEC

[SMC^{+}15] Aya Shibuya, Neil Margulis, Romain Christiano, Tobias C. Walther, and Charles Barlowe. The Erv41–Erv46 complex serves as a retrograde receptor to retrieve escaped ER proteins. Journal of Cell Biology, 208(2):197–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/2/197.

Scheffler:2015:MME

[SMF^{+}15] Kathleen Scheffler, Refael Minnes, Vincent Fraisier, Anne Paoletti, and Phong T. Tran. Microtubule minus end motors kinesin-14 and dynein drive nuclear congression in parallel pathways. Journal of Cell Biology, 209(1):47–??, April 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/209/1/47.

Shima:2018:KBT

[SMK^{+}18] Tomohiro Shima, Manatsu Morikawa, Junichi Kaneshiro, Taketoshi Kambara, Shinji Kamimura, Toshiki Yagi, Hiroyuki Iwamoto, Sotaro Uemura, Hideki Shigematsu, Mikako Shirouzu, Taro Ichimura, Tomonobu M. Watanabe, Ryo Nitta, Yasushi Okada, and Nobutaka Hirokawa. Kinesin-binding– triggered conformation switching of microtubules contributes to polarized transport. Journal of Cell Biology, 217(12):4164– ??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/12/4164.

Szabo:2016:VCP

[SMN^{+}16] András Szabó, Manuela Melchionda, Giancarlo Nastasi, Mae L. Woods, Salvatore Campo, Roberto Perris, and Roberto Mayor. In vivo confinement promotes collective migration of neural crest cells. Journal of Cell Biology, 213(5):543–??, June 2016.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/213/5/543.

Suresh:2017:MNP

[SMOO17] Subbulakshmi Suresh, Sarine Markossian, Aysha H. Osmani, and Stephen A. Osmani. Mitotic nuclear pore complex segregation involves Nup2 in Aspergillus nidulans. Journal of Cell Biology, 216(9):2813–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/9/2813.

Stanyte:2018:DSC

[SNB^{+}18] Rugile Stanyte, Johannes Nuebler, Claudia Blaukopf, Rudolf Hoefler, Roman Stocsits, Jan-Michael Peters, and Daniel W. Gerlich. Dynamics of sister chromatid resolution during cell cycle progression. Journal of Cell Biology, 217(6):1985–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/6/
1985.

Sturgill:2016:KIR

[SNGO16] Emma G. Sturgill, Stephen R. Norris, Yan Guo, and Ryoma Ohi. Kinesin-5 inhibitor resistance is driven by kinesin-12. Journal of Cell Biology, 213(2):213–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/213/2/213.

Santos:2016:TMC

[SNOBM16] António J. M. Santos, Cristina Nogueira, Maria Ortega-Bellido, and Vivek Malhotra. TANGO1 and Mia2/cTAGE5 (TALI) cooperate to export bulky pre-chylomicrons/VLDLs from the endoplasmic reticulum. Journal of Cell Biology, 213(3):343–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
213/3/343.

Shigetomi:2018:AJI

[SOII18] Kenta Shigetomi, Yumiko Ono, Tetsuichiro Inai, and Junichi Ikenouchi. Adherens junctions influence tight junction formation via changes in membrane lipid composition. Journal of Cell Biology, 217(7):2373–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/7/2373.

Silva:2016:RGR

[SOP^{+}16] Ana M. Silva, Daniel S. Osório, Antonio J. Pereira, Helder Maiato, Inês Mendes Pinto, Boris Rubinstein, Reto Gassmann, Ivo Andreas Telley, and Ana Xavier Carvalho. Robust gap repair in the contractile ring ensures timely completion of cytokinesis. Journal of Cell Biology, 215(6):789–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/215/6/789.

Sorensen:2017:RWR

[Sør17] Jakob B. Sørensen. Ride the wave: Retrograde trafficking becomes Ca^{2+} dependent with BAIAP3. Journal of Cell Biology, 216(7):1887–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/7/1887.

Simoes:2017:MIP

[SOW^{+}17] Sérgio Simões, Youjin Oh, Michael F. Z. Wang, Rodrigo Fernandez-Gonzalez, and Ulrich Tepass. Myosin II promotes the anisotropic loss of the apical domain during Drosophila neuroblast ingression. Journal of Cell Biology, 216(5):1387–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/5/
1387.

Salinas:2017:PDF

[SPD^{+}17] Raquel Y. Salinas, Jillian N. Pearring, Jin-Dong Ding, William J. Spencer, Ying Hao, and Vadim Y. Arshavsky. Photoreceptor discs form through peripherin-dependent suppression of ciliary ectosome release. Journal of Cell Biology, 216 (5):1489–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/5/1489.

Schietroma:2017:UST

[SPE^{+}17a] Cataldo Schietroma, Karine Parain, Amrit Estivalet, Asadollah Aghaie, Jacques Boutet de Monvel, Serge Picaud, José-Alain Sahel, Muriel Perron, Aziz El-Amraoui, and Christine Petit. Usher syndrome type 1–associated cadherins shape the photoreceptor outer segment. Journal of Cell Biology, 216(6):1849–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/6/1849.

Spencer:2017:ARM

[Spe17b] Timothy K. Spencer. Antonina Roll–Mecak: Decoding the secrets of tubulin complexity. Journal of Cell Biology, 216 (5):1208–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/5/1208.

Salas-Pino:2017:FYN

[SPGB^{+}17] Silvia Salas-Pino, Paola Gallardo, Ramón R. Barrales, Sigurd Braun, and Rafael R. Daga. The fission yeast nucleoporin Alm1 is required for proteasomal degradation of kinetochore components. Journal of Cell Biology, 216(11):3591–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/11/3591.

Sai:2019:NMR

[SPH^{+}19] Kazuhito Sai, Cameron Parsons, John S. House, Sophia Kathariou, and Jun Ninomiya-Tsuji. Necroptosis mediators RIPK3 and MLKL suppress intracellular Listeria replication independently of host cell killing. Journal of Cell Biology, 218(6):1994– ??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/6/1994.

Smith:2015:NAA

[SPJ^{+}15] Carlas S. Smith, Stephan Preibisch, Aviva Joseph, Sara Abrahamsson, Bernd Rieger, Eugene Myers, Robert H. Singer, and David Grunwald. Nuclear accessibility of β-actin mRNA is measured by 3D single-molecule real-time tracking. Journal of Cell Biology, 209(4):609–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/4/609.

Schiklenk:2018:CMC

[SPK^{+}18] Christoph Schiklenk, Boryana Petrova, Marc Kschonsak, Markus Hassler, Carlo Klein, Toby J. Gibson, and Christian H. Haering. Control of mitotic chromosome condensation by the fission yeast transcription factor Zas1. Journal of Cell Biology, 217(7):2383–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/7/2383.

Schatton:2017:CRM

[SPMM^{+}17] Désirée Schatton, David Pla-Martin, Marie-Charlotte Marx, Henriette Hansen, Arnaud Mourier, Ivan Nemazanyy, Alberto Pessia, Peter Zentis, Teresa Corona, Vangelis Kondylis, Esther Barth, Astrid C. Schauss, Vidya Velagapudi, and Elena I. Rugarli. CLUH regulates mitochondrial metabolism by controlling translation and decay of target mRNAs. Journal of Cell Biology, 216(3):675–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/3/675.

Schweizer:2015:OEE

[SPWM15] Nina Schweizer, Nisha Pawar, Matthias Weiss, and Helder Maiato. An organelle-exclusion envelope assists mitosis and underlies distinct molecular crowding in the spindle region. Journal of Cell Biology, 210(5):695–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/5/695.

Scott:2015:CBR

[SQ15] Rizaldy P. Scott and Susan E. Quaggin. The cell biology of renal filtration. Journal of Cell Biology, 209(2):199–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/209/2/
199.

Szabo:2015:CIC

[SQB^{+}15] Eva Szabo, Yuanyuan Qiu, Shairaz Baksh, Marek Michalak, and Michal Opas. Calreticulin inhibits commitment to adipocyte differentiation. Journal of Cell Biology, 208(2): 249–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/208/2/249.

Shen:2016:PAT

[SQC^{+}16] Yi Shen, Huaping Qin, Juan Chen, Lingyan Mou, Yang He, Yixiu Yan, Hang Zhou, Ya Lv, Zhong Chen, Junlu Wang, and Yu-Dong Zhou. Postnatal activation of TLR4 in astrocytes promotes excitatory synaptogenesis in hippocampal neurons. Journal of Cell Biology, 215(5):719–??, December 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/215/5/719.

Sidor:2017:STW

[SR17a] Clara Sidor and Katja Röper. Squeezing out in a “tug of war”: The role of myosin in neural stem cell delamination. Journal of Cell Biology, 216(5):1215–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/5/1215.

Starr:2017:TRL

[SR17b] Daniel A. Starr and Lesilee S. Rose. TorsinA regulates the LINC to moving nuclei. Journal of Cell Biology, 216(3):543–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/3/543.

Schaub:2019:YJR

[SRF19] Christoph Schaub, Marcel Rose, and Manfred Frasch. Yorkie and JNK revert syncytial muscles into myoblasts during Org-1–dependent lineage reprogramming. Journal of Cell Biology, 218(11):3572–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/11/3572.

Schiffhauer:2019:MIA

[SRI^{+}19] Eric S. Schiffhauer, Yixin Ren, Vicente A. Iglesias, Priyanka Kothari, Pablo A. Iglesias, and Douglas N. Robinson. Myosin IIB assembly state determines its mechanosensitive dynamics. Journal of Cell Biology, 218(3):895–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/3/895.

Sikirzhytski:2018:MAN

[SRT^{+}18] Vitali Sikirzhytski, Fioranna Renda, Irina Tikhonenko, Valentin Magidson, Bruce F. McEwen, and Alexey Khodjakov. Microtubules assemble near most kinetochores during early prometaphase in human cells. Journal of Cell Biology, 217 (8):2647–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/8/2647.

Schneider:2016:CFE

[SS16] Anja Schneider and Mikael Simons. Catching filopodia: Exosomes surf on fast highways to enter cells. Journal of Cell Biology, 213(2):143–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/213/2/143.

Smith:2018:PKA

[SS18] F. Donelson Smith and John D. Scott. Protein kinase A activation: Something new under the sun? Journal of Cell Biology, 217(6):1895–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/6/1895.

Sharma:2019:RTC

[SS19] Manish Sharma and Srinivasa Subramaniam. Rhes travels from cell to cell and transports Huntington disease protein via TNTlike protrusion. Journal of Cell Biology, 218(6):1972–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/6/
1972.

Schulze:2019:CBH

[SSC^{+}19] Ryan J. Schulze, Micah B. Schott, Carol A. Casey, Pamela L. Tuma, and Mark A. McNiven. The cell biology of the hepatocyte: a membrane trafficking machine. Journal of Cell Biology, 218(7):2096–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/7/2096.

Samejima:2015:WPG

[SSdLA^{+}15] Itaru Samejima, Christos Spanos, Flavia de Lima Alves, Tetsuya Hori, Marinela Perpelescu, Juan Zou, Juri Rappsilber, Tatsuo Fukagawa, and William C. Earnshaw. Whole-proteome genetic analysis of dependencies in assembly of a vertebrate kinetochore. Journal of Cell Biology, 211(6):1141–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/6/
1141.

Sirka:2018:MCD

[SSE18] Orit Katarina Sirka, Eliah R. Shamir, and Andrew J. Ewald. Myoepithelial cells are a dynamic barrier to epithelial dissemination. Journal of Cell Biology, 217(10):3368–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/
10/3368.

Schulz:2015:CDA

[SSH^{+}15] Anna M. Schulz, Susanne Stutte, Sebastian Hogl, Nancy Luckashenak, Diana Dudziak, Céline Leroy, Ignasi Forné, Axel Imhof, Stephan A. Müller, Cord H. Brakebusch, Stefan F. Lichtenthaler, and Thomas Brocker. Cdc42-dependent actin dynamics controls maturation and secretory activity of dendritic cells. Journal of Cell Biology, 211(3):553–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/211/3/553.

Schendzielorz:2017:TDM

[SSL^{+}17] Alexander Benjamin Schendzielorz, Christian Schulz, Oleksandr Lytovchenko, Anne Clancy, Bernard Guiard, Raffaele Ieva, Martin van der Laan, and Peter Rehling. Two distinct membrane potential-dependent steps drive mitochondrial matrix protein translocation. Journal of Cell Biology, 216(1): 83–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/1/83.

Saegusa:2018:SSC

[SSM^{+}18] Keiko Saegusa, Miyuki Sato, Nobukatsu Morooka, Taichi Hara, and Ken Sato. SFT-4/ surf4 control ER export of soluble cargo proteins and participate in ER exit site organization. Journal of Cell Biology, 217(6):2073–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/6/2073.

Schouteden:2015:CTZ

[SSPD15] Clementine Schouteden, Daniel Serwas, Mate Palfy, and Alexander Dammermann. The ciliary transition zone functions in cell adhesion but is dispensable for axoneme assembly in C. elegans. Journal of Cell Biology, 210(1):35–??, July 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/210/1/35.

Serwas:2017:CIC

[SSR^{+}17] Daniel Serwas, Tiffany Y. Su, Max Roessler, Shaohe Wang, and Alexander Dammermann. Centrioles initiate cilia assembly but are dispensable for maturation and maintenance in C. elegans. Journal of Cell Biology, 216(6):1659–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/6/1659.

Sha:2018:RIP

[SSRG18] Zhe Sha, Helena M. Schnell, Kerstin Ruoff, and Alfred Goldberg. Rapid induction of p62 and GABARAPL1 upon proteasome inhibition promotes survival before autophagy activation. Journal of Cell Biology, 217(5):1757–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/5/1757.

Schmid:2018:IMC

[SSV^{+}18] Fabian Marc Schmid, Kenneth Bødtker Schou, Martin Juel Vilhelm, Maria Schrøder Holm, Loretta Breslin, Pietro Farinelli, Lars Allan Larsen, Jens Skorstengaard Andersen, Lotte Bang Pedersen, and Søren Tvorup Christensen. IFT20 modulates ciliary PDGFRα signaling by regulating the stability of Cbl E3 ubiquitin ligases. Journal of Cell Biology, 217(1):151–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/1/151.

Simoes:2016:MVS

[ST16a] Sergio Simoes and Ulrich Tepass. Muscle versus Snail: Muscle wins. Journal of Cell Biology, 212(2):139–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/212/2/139.

Sturner:2016:REF

[ST16b] Tomke Stürner and Gaia Tavosanis. Rotating for elongation: Fat2 whips for the race. Journal of Cell Biology, 212(5): 487–??, February 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/5/487.

Spillane:2017:BCA

[ST17] Katelyn M. Spillane and Pavel Tolar. B cell antigen extraction is regulated by physical properties of antigen-presenting cells. Journal of Cell Biology, 216(1):217–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/1/217.

Schvartzman:2018:MRC

[STF18] Juan Manuel Schvartzman, Craig B. Thompson, and Lydia W. S. Finley. Metabolic regulation of chromatin modifications and gene expression. Journal of Cell Biology, 217(7):2247–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/7/2247.

Sampayo:2018:FRE

[STR^{+}18] Rocío G. Sampayo, Andrés M. Toscani, Matthew G. Rubashkin, Kate Thi, Luciano A. Masullo, Ianina L. Violi, Jonathon N. Lakins, Alfredo Cáceres, William C. Hines, Federico Coluccio Leskow, Fernando D. Stefani, Dante R. Chialvo, Mina J. Bissell, Valerie M. Weaver, and Marina Simian. Fibronectin rescues estrogen receptor α from lysosomal degradation in breast cancer cells. Journal of Cell Biology, 217(8):2777–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/8/
2777.

Schroeder:2016:AAD

[SV16] Courtney M. Schroeder and Ronald D. Vale. Assembly and activation of dynein–dynactin by the cargo adaptor protein Hook3. Journal of Cell Biology, 214(3):309–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/3/309.

Schreiber:2015:ULT

[SVD^{+}15] Joerg Schreiber, Marlene J. Végh, Julia Dawitz, Tim Kroon, Maarten Loos, Dorthe Labonté, Ka Wan Li, Pim Van Nierop, Michiel T. Van Diepen, Chris I. De Zeeuw, Matthias Kneussel, Rhiannon M. Meredith, August B. Smit, and Ronald E. Van Kesteren. Ubiquitin ligase TRIM3 controls hippocampal plasticity and learning by regulating synaptic γ-actin levels. Journal of Cell Biology, 211(3):569–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/211/3/569.

Sargent:2016:PEU

[SvZS^{+}16] Graeme Sargent, Tim van Zutphen, Tatiana Shatseva, Ling Zhang, Valeria Di Giovanni, Robert Bandsma, and Peter Kijun Kim. PEX2 is the E3 ubiquitin ligase required for pexophagy during starvation. Journal of Cell Biology, 214(6):

677–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/214/6/677.

Sung:2018:DMC

[SW18] Bong Hwan Sung and Alissa M. Weaver. Directed migration: Cells navigate by extracellular vesicles. Journal of Cell Biology, 217(8):2613–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/8/2613.

Snider:2017:PMR

[SWC^{+}17] Chloe E. Snider, Alaina H. Willet, Jun-Song Chen, Göker Arpağ, Marija Zanic, and Kathleen L. Gould. Phosphoinositidemediated ring anchoring resists perpendicular forces to promote medial cytokinesis. Journal of Cell Biology, 216(10):3041–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/10/3041.

Sato:2019:RMF

[SWD^{+}19] Hanae Sato, Bin Wu, Fabien Delahaye, Robert H. Singer, and John M. Greally. Retargeting of macroH2A following mitosis to cytogenetic-scale heterochromatic domains. Journal of Cell Biology, 218(6):1810–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/6/1810.

Silver:2019:APP

[SWPS^{+}19] Jordan T. Silver, Frederik Wirtz-Peitz, Sérgio Simões, Milena Pellikka, Dong Yan, Richard Binari, Takashi Nishimura, Yan Li, Tony J. C. Harris, Norbert Perrimon, and Ulrich Tepass. Apical polarity proteins recruit the RhoGEF Cysts to promote junctional myosin assembly. Journal of Cell Biology, 218(10): 3397–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/10/3397.

Schott:2019:LDS

[SWS^{+}19] Micah B. Schott, Shaun G. Weller, Ryan J. Schulze, Eugene W. Krueger, Kristina Drizyte-Miller, Carol A. Casey, and Mark A. McNiven. Lipid droplet size directs lipolysis and lipophagy catabolism in hepatocytes. Journal of Cell Biology, 218(10):

3320–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/10/3320.

Salle:2019:ADT

[SXE^{+}19] Jérémy Sallé, Jing Xie, Dmitry Ershov, Milan Lacassin, Serge Dmitrieff, and Nicolas Minc. Asymmetric division through a reduction of microtubule centering forces. Journal of Cell Biology, 218(3):771–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/3/771.

Schiffmacher:2016:CPP

[SXT16] Andrew T. Schiffmacher, Vivien Xie, and Lisa A. Taneyhill. Cadherin-6B proteolysis promotes the neural crest cell epithelial-to-mesenchymal transition through transcriptional regulation. Journal of Cell Biology, 215(5):735–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/215/5/
735.

Sun:2017:RSP

[SYK^{+}17] Tianliang Sun, Lida Yang, Harmandeep Kaur, Jenny Pestel, Mario Looso, Hendrik Nolte, Cornelius Krasel, Daniel Heil, Ramesh K. Krishnan, Marie-Josée Santoni, Jean-Paul Borg, Moritz Bünemann, Stefan Offermanns, Jakub M. Swiercz, and Thomas Worzfeld. A reverse signaling pathway downstream of Sema4A controls cell migration via Scrib. Journal of Cell Biology, 216(1):199–??, January 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/1/199.

Schuldiner:2017:ICN

[SZ17a] Maya Schuldiner and Einat Zalckvar. Incredibly close — A newly identified peroxisome–ER contact site in humans. Journal of Cell Biology, 216(2):287–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/2/287.

Serquina:2017:HHP

[SZ17b] Anna K. Serquiña and Joseph M. Ziegelbauer. How herpesviruses pass on their genomes. Journal of Cell Biology, 216(9):2611–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/9/2611.

Sardana:2019:RUL

[SZE19] Richa Sardana, Lu Zhu, and Scott D. Emr. Rsp5 Ubiquitin ligase–mediated quality control system clears membrane proteins mistargeted to the vacuole membrane. Journal of Cell Biology, 218(1):234–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/1/234.

Schoborg:2015:ACC

[SZF^{+}15] Todd Schoborg, Allison L. Zajac, Carey J. Fagerstrom, Rodrigo X. Guillen, and Nasser M. Rusan. An Asp–CaM complex is required for centrosome–pole cohesion and centrosome inheritance in neural stem cells. Journal of Cell Biology, 211(5): 987–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/5/987.

Snead:2019:BSD

[SZK^{+}19] Wilton T. Snead, Wade F. Zeno, Grace Kago, Ryan W. Perkins, J. Blair Richter, Chi Zhao, Eileen M. Lafer, and Jeanne C. Stachowiak. BAR scaffolds drive membrane fission by crowding disordered domains. Journal of Cell Biology, 218(2): 664–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/2/664.

Shang:2016:ITM

[SZL^{+}16] Shujiang Shang, Feipeng Zhu, Bin Liu, Zuying Chai, Qihui Wu, Meiqin Hu, Yuan Wang, Rong Huang, Xiaoyu Zhang, Xi Wu, Lei Sun, Yeshi Wang, Li Wang, Huadong Xu, Sasa Teng, Bing Liu, Lianghong Zheng, Chen Zhang, Fukang Zhang, Xinghua Feng, Desheng Zhu, Changhe Wang, Tao Liu, Michael X. Zhu, and Zhuan Zhou. Intracellular TRPA1 mediates Ca^{2+} release from lysosomes in dorsal root ganglion neurons. Journal of Cell Biology, 215(3):369–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/215/3/369.

Stewart:2015:NCL

[SZR^{+}15] Rachel M. Stewart, Amanda E. Zubek, Kathryn A. Rosowski, Sarah M. Schreiner, Valerie Horsley, and Megan C. King. Nuclear–cytoskeletal linkages facilitate cross talk between the nucleus and intercellular adhesions. Journal of Cell Biology, 209(3):403–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/209/3/403.

Segal:2018:FIA

[SZSS18] Dagan Segal, Assaf Zaritsky, Eyal D. Schejter, and Ben-Zion Shilo. Feedback inhibition of actin on Rho mediates content release from large secretory vesicles. Journal of Cell Biology, 217 (5):1815–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/5/1815.

Toyama:2019:VLL

[TALR^{+}19] Brandon H. Toyama, Rafael Arrojo e Drigo, Varda Lev-Ram, Ranjan Ramachandra, Thomas J. Deerinck, Claude Lechene, Mark H. Ellisman, and Martin W. Hetzer. Visualization of long-lived proteins reveals age mosaicism within nuclei of postmitotic cells. Journal of Cell Biology, 218(2):433–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/2/
433.

Tripathi:2019:SEC

[TAQ^{+}19] Brajendra K. Tripathi, Meghan F. Anderman, Xiaolan Qian, Ming Zhou, Dunrui Wang, Alex G. Papageorge, and Douglas R. Lowy. SRC and ERK cooperatively phosphorylate DLC1 and attenuate its Rho–GAP and tumor suppressor functions. Journal of Cell Biology, 218(9):3060–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/9/3060.

Taraska:2015:CBF

[Tar15] Justin W. Taraska. Cell biology of the future: Nanometer-scale cellular cartography. Journal of Cell Biology, 211(2):211–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
211/2/211.

Tardieux:2016:RMP

[TB16] Isabelle Tardieux and Jake Baum. Reassessing the mechanics of parasite motility and host-cell invasion. Journal of Cell Biology, 214(5):507–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/5/507.

Tarasenko:2017:MCM

[TBJ^{+}17] Daryna Tarasenko, Mariam Barbot, Daniel C. Jans, Benjamin Kroppen, Boguslawa Sadowski, Gudrun Heim, Wiebke Möbius, Stefan Jakobs, and Michael Meinecke. The MICOS component Mic60 displays a conserved membrane-bending activity that is necessary for normal cristae morphology. Journal of Cell Biology, 216(4):889–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/4/889.

Tang:2016:SHF

[TBK^{+}16] Qing Tang, Neil Billington, Elena B. Krementsova, Carol S. Bookwalter, Matthew Lord, and Kathleen M. Trybus. A single-headed fission yeast myosin V transports actin in a tropomyosin-dependent manner. Journal of Cell Biology, 214 (2):167–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/214/2/167.

Thangavel:2015:DDP

[TBL^{+}15] Saravanabhavan Thangavel, Matteo Berti, Maryna Levikova, Cosimo Pinto, Shivasankari Gomathinayagam, Marko Vujanovic, Ralph Zellweger, Hayley Moore, Eu Han Lee, Eric A. Hendrickson, Petr Cejka, Sheila Stewart, Massimo Lopes, and Alessandro Vindigni. DNA2 drives processing and restart of reversed replication forks in human cells. Journal of Cell Biology, 208(5):545–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/5/545.

Tornavaca:2015:ZCE

[TCD^{+}15] Olga Tornavaca, Minghao Chia, Neil Dufton, Lourdes Osuna Almagro, Daniel E. Conway, Anna M. Randi, Martin A. Schwartz, Karl Matter, and Maria S. Balda. ZO-1 controls endothelial adherens junctions, cell–cell tension, angiogenesis, and barrier formation. Journal of Cell Biology, 208(6):821–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
208/6/821.

Terawaki:2015:RFD

[TCP^{+}15] Seigo Terawaki, Voahirana Camosseto, Francesca Prete, Till Wenger, Alexia Papadopoulos, Christiane Rondeau, Alexis Combes, Christian Rodriguez Rodrigues, Thien-Phong Vu Manh, Mathieu Fallet, Luc English, Rodrigo Santamaria, Ana R. Soares, Tobias Weil, Hamida Hammad, Michel Desjardins, Jean-Pierre Gorvel, Manuel A. S. Santos, Evelina Gatti, and Philippe Pierre. RUN and FYVE domain-containing protein 4 enhances autophagy and lysosome tethering in response to Interleukin-4. Journal of Cell Biology, 210(7):1133– ??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/7/1133.

Tang:2018:PIS

[TCP^{+}18] Chih-Hang Anthony Tang, Shiun Chang, Adrienne W. Paton, James C. Paton, Dmitry I. Gabrilovich, Hidde L. Ploegh, Juan R. Del Valle, and Chih-Chi Andrew Hu. Phosphorylation of IRE1 at S729 regulates RIDD in B cells and antibody production after immunization. Journal of Cell Biology, 217 (5):1739–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/5/1739.

Thuma:2018:DIC

[TCWM18] Leila Thuma, Deborah Carter, Helen Weavers, and Paul Martin. Drosophila immune cells extravasate from vessels to wounds using Tre1 GPCR and Rho signaling. Journal of Cell Biology, 217(9):3045–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/9/3045.

Trudeau:2016:LAP

[TCZ^{+}16] Kyle M. Trudeau, Aaron H. Colby, Jialiu Zeng, Guy Las, Jiazuo H. Feng, Mark W. Grinstaff, and Orian S. Shirihai. Lysosome acidification by photoactivated nanoparticles restores autophagy under lipotoxicity. Journal of Cell Biology, 214(1):25– ??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
214/1/25.

Tietz:2015:BBC

[TE15] Silvia Tietz and Britta Engelhardt. Brain barriers: Crosstalk between complex tight junctions and adherens junctions. Journal of Cell Biology, 209(4):493–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/209/4/493.

Thomas:2016:GCT

[TF16] Laura L. Thomas and J. Christopher Fromme. GTPase cross talk regulates TRAPPII activation of Rab11 homologues during vesicle biogenesis. Journal of Cell Biology, 215(4):499– ??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/215/4/499.

Toivanen:2019:BAP

[TF19] Roxanne Toivanen and Luc Furic. A balancing act: PHLPP2 fine tunes AKT activity and MYC stability in prostate cancer. Journal of Cell Biology, 218(6):1771–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/6/1771.

Trimble:2015:BFD

[TG15] William S. Trimble and Sergio Grinstein. Barriers to the free diffusion of proteins and lipids in the plasma membrane. Journal of Cell Biology, 208(3):259–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/3/259.

Titus:2017:EPC

[TG17] Margaret A. Titus and Holly V. Goodson. An evolutionary perspective on cell migration: Digging for the roots of amoeboid motility. Journal of Cell Biology, 216(6):1509–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/6/
1509.

Tischer:2019:AMT

[TG19] Julia Tischer and Fanni Gergely. Anti-mitotic therapies in cancer. Journal of Cell Biology, 218(1):10–??, January 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/1/10.

Tabas:2015:RIC

[TGCO15] Ira Tabas, Guillermo García-Cardeña, and Gary K. Owens. Recent insights into the cellular biology of atherosclerosis. Journal of Cell Biology, 209(1):13–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/209/1/13.

Tang:2017:BPE

[TGJ^{+}17] Xia Tang, Jianan Gao, Xinling Jia, Wencao Zhao, Yijie Zhang, Weijun Pan, and Jie He. Bipotent progenitors as embryonic origin of retinal stem cells. Journal of Cell Biology, 216(6):1833– ??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/6/1833.

Turan:2019:ADL

[TGK^{+}19] Aykut Turan, Linda Grosche, Adalbert Krawczyk, Petra Mühl-Zürbes, Christina Drassner, Alexandra Düthorn, Mirko Kummer, Mike Hasenberg, Sylvia Voortmann, Holger Jastrow, Jan Dörrie, Niels Schaft, Max Kraner, Katinka Döhner, Beate Sodeik, Alexander Steinkasserer, and Christiane Silke Heilingloh. Autophagic degradation of lamins facilitates the nuclear egress of herpes simplex virus type 1. Journal of Cell Biology, 218(2):508–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/2/508.

Tripathi:2017:RTK

[TGQ^{+}17] Brajendra K. Tripathi, Tiera Grant, Xiaolan Qian, Ming Zhou, Philipp Mertins, Dunrui Wang, Alex G. Papageorge, Sergey G. Tarasov, Kent W. Hunter, Steven A. Carr, and Douglas R. Lowy. Receptor tyrosine kinase activation of RhoA is mediated by AKT phosphorylation of DLC1. Journal of Cell Biology, 216(12):4255–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/12/4255.

Takahashi:2018:DSC

[TH18] Motoko Takahashi and Toru Hirota. Dynamics of sister chromatids through the cell cycle: Together and apart. Journal of Cell Biology, 217(6):1887–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/6/1887.

Tanimura:2016:ESP

[THA^{+}16] Susumu Tanimura, Junya Hashizume, Naoya Arichika, Kazushi Watanabe, Kaname Ohyama, Kohsuke Takeda, and Michiaki Kohno. ERK signaling promotes cell motility by inducing the localization of myosin 1E to lamellipodial tips. Journal of Cell Biology, 214(4):475–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/214/4/475.

Taniguchi:2019:OBB

[THG19] Kenichiro Taniguchi, Idse Heemskerk, and Deborah L. Gumucio. Opening the black box: Stem cell–based modeling of human post-implantation development. Journal of Cell Biology, 218 (2):410–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/2/410.

Trotter:2019:SNA

[THM^{+}19] Justin H. Trotter, Junjie Hao, Stephan Maxeiner, Theodoros Tsetsenis, Zhihui Liu, Xiaowei Zhuang, and Thomas C. Südhof. Synaptic neurexin-1 assembles into dynamically regulated active zone nanoclusters. Journal of Cell Biology, 218(8):2677–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/8/2677.

Thomas:2018:TCA

[TJF18] Laura L. Thomas, Aaron M. N. Joiner, and J. Christopher Fromme. The TRAPPIII complex activates the GTPase Ypt1 (Rab1) in the secretory pathway. Journal of Cell Biology, 217 (1):283–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/1/283.

Teixeira:2018:RLD

[TJMM^{+}18] Vitor Teixeira, Lisa Johnsen, Fernando Martínez-Montañés, Alexandra Grippa, Laura Buxó, Fatima-Zahra Idrissi, Christer S. Ejsing, and Pedro Carvalho. Regulation of lipid droplets by metabolically controlled Ldo isoforms. Journal of Cell Biology, 217(1):127–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/1/127.

Tanimoto:2016:SMR

[TKM16] Hirokazu Tanimoto, Akatsuki Kimura, and Nicolas Minc. Shape–motion relationships of centering microtubule asters. Journal of Cell Biology, 212(7):777–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/212/7/777.

Todorov:2017:BRR

[TL17] Vladimir Todorov and Andreas Linkermann. Back to the roots of regulated necrosis. Journal of Cell Biology, 216(2): 303–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/2/303.

Takahashi:2019:VDE

[TLH^{+}19] Yoshinori Takahashi, Xinwen Liang, Tatsuya Hattori, Zhenyuan Tang, Haiyan He, Han Chen, Xiaoming Liu, Thomas Abraham, Yuka Imamura-Kawasawa, Nicholas J. Buchkovich, Megan M. Young, and Hong-Gang Wang. VPS37A directs ESCRT recruitment for phagophore closure. Journal of Cell Biology, 218(10): 3336–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/10/3336.

Treuner-Lange:2015:SGP

[TLMG^{+}15] Anke Treuner-Lange, Eric Macia, Mathilde Guzzo, Edina Hot, Laura M. Faure, Beata Jakobczak, Leon Espinosa, Damien Alcor, Adrien Ducret, Daniela Keilberg, Jean Philippe Castaing, Sandra Lacas Gervais, Michel Franco, Lotte Søgaard-Andersen, and Tâm Mignot. The small G-protein MglA connects to the MreB actin cytoskeleton at bacterial focal adhesions. Journal of Cell Biology, 210(2):243–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/2/243.

Templeman:2018:RRL

[TM18] Nicole M. Templeman and Coleen T. Murphy. Regulation of reproduction and longevity by nutrient-sensing pathways.

Journal of Cell Biology, 217(1):93–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/1/93.

Tonnessen-Murray:2019:CIS

[TMFR^{+}19] Crystal A. Tonnessen-Murray, Wesley D. Frey, Sonia G. Rao, Ashkan Shahbandi, Nathan A. Ungerleider, Joy O. Olayiwola, Lucas B. Murray, Benjamin T. Vinson, Douglas B. Chrisey, Christopher J. Lord, and James G. Jackson. Chemotherapyinduced senescent cancer cells engulf other cells to enhance their survival. Journal of Cell Biology, 218(11):3827–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/11/3827.

Tokarz:2018:CBS

[TMK18] Victoria L. Tokarz, Patrick E. MacDonald, and Amira Klip. The cell biology of systemic insulin function. Journal of Cell Biology, 217(7):2273–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/7/2273.

Tsoumpekos:2018:DBB

[TNK18] Giorgos Tsoumpekos, Linda Nemetschke, and Elisabeth Knust. Drosophila Big bang regulates the apical cytocortex and wing growth through junctional tension. Journal of Cell Biology, 217 (3):1033–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/3/1033.

Tavernier:2015:CPS

[TNP^{+}15] Nicolas Tavernier, Anna Noatynska, Costanza Panbianco, Lisa Martino, Lucie Van Hove, Françoise Schwager, Thibaut Léger, Monica Gotta, and Lionel Pintard. Cdk1 phosphorylates SPAT-1/Bora to trigger PLK-1 activation and drive mitotic entry in C. elegans embryos. Journal of Cell Biology, 208(6):661–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
208/6/661.

Tooze:2018:SAR

[Too18] Sharon A. Tooze. SNAREing an ARP requires a LIR. Journal of Cell Biology, 217(3):803–??, March 2018. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/3/803.

Trainees:2018:GBP

[Tra18] Blobel Laboratory Trainees. Günter Blobel: Pioneer of molecular cell biology (1936–2018). Journal of Cell Biology, 217 (4):1163–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/4/1163.

Timney:2016:SRP

[TRM^{+}16] Benjamin L. Timney, Barak Raveh, Roxana Mironska, Jill M. Trivedi, Seung Joong Kim, Daniel Russel, Susan R. Wente, Andrej Sali, and Michael P. Rout. Simple rules for passive diffusion through the nuclear pore complex. Journal of Cell Biology, 215(1):57–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/1/57.

Thompson:2015:MKD

[TS15a] Barry J. Thompson and Erik Sahai. MST kinases in development and disease. Journal of Cell Biology, 210(6):871– ??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/6/871.

Thumm:2015:MSC

[TS15b] Michael Thumm and Mikael Simons. Myelinophagy: Schwann cells dine in. Journal of Cell Biology, 210(1):9–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/210/1/9.

Torrino:2018:EMC

[TSB^{+}18] Stéphanie Torrino, Wei-Wei Shen, Cédric M. Blouin, Satish Kailasam Mani, Christine Viaris de Lesegno, Pierre Bost, Alexandre Grassart, Darius Köster, Cesar Augusto Valades-Cruz, Valérie Chambon, Ludger Johannes, Paolo Pierobon, Vassili Soumelis, Catherine Coirault, Stéphane Vassilopoulos, and Christophe Lamaze. EHD2 is a mechanotransducer connecting caveolae dynamics with gene transcription. Journal of Cell Biology, 217(12):4092–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/12/4092.

Tsaalbi-Shtylik:2015:ETS

[TSFP^{+}15] Anastasia Tsaalbi-Shtylik, Cristina Ferrás, Bea Pauw, Giel Hendriks, Piya Temviriyanukul, Leone Carlée, Fabienne Calléja, Sandrine van Hees, Jun-Ichi Akagi, Shigenori Iwai, Fumio Hanaoka, Jacob G. Jansen, and Niels de Wind. Excision of translesion synthesis errors orchestrates responses to helixdistorting DNA lesions. Journal of Cell Biology, 209(1):33–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
209/1/33.

Tian:2015:ISH

[TSJ^{+}15] Aiguo Tian, Qing Shi, Alice Jiang, Shuangxi Li, Bing Wang, and Jin Jiang. Injury-stimulated Hedgehog signaling promotes regenerative proliferation of Drosophila intestinal stem cells. Journal of Cell Biology, 208(6):807–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/6/807.

Tsuchiya:2018:IXP

[TSK^{+}18] Yuichi Tsuchiya, Michiko Saito, Hiroshi Kadokura, Jun ichi Miyazaki, Fumi Tashiro, Yusuke Imagawa, Takao Iwawaki, and Kenji Kohno. IRE1–XBP1 pathway regulates oxidative proinsulin folding in pancreatic β cells. Journal of Cell Biology, 217 (4):1287–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress. org/content/217/4/1287. See correction [TSK^{+}19].

Tsuchiya:2019:CIX

[TSK^{+}19] Yuichi Tsuchiya, Michiko Saito, Hiroshi Kadokura, Jun ichi Miyazaki, Fumi Tashiro, Yusuke Imagawa, Takao Iwawaki, and Kenji Kohno. Correction: IRE1–XBP1 pathway regulates oxidative proinsulin folding in pancreatic β cells. Journal of Cell Biology, 218(5):1764–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/5/1764. See [TSK^{+}18].

Taniguchi:2017:AIS

[TST^{+}17] Kenichiro Taniguchi, Yue Shao, Ryan F. Townshend, Chari L. Cortez, Clair E. Harris, Sasha Meshinchi, Sundeep Kalantry, Jianping Fu, K. Sue O’Shea, and Deborah L. Gumucio. An apicosome initiates self-organizing morphogenesis of human pluripotent stem cells. Journal of Cell Biology, 216(12):3981– ??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/12/3981.

Tourriere:2019:RPG

[TT19] Hélène Tourrière and Jamal Tazi. Reply to “Phosphorylation of G3BP1-S149 does not influence stress granule assembly”. Journal of Cell Biology, 218(7):2433–??, July 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/7/2433. See

[PKS^{+}19].

Tatomer:2016:CPM

[TTC^{+}16] Deirdre C. Tatomer, Esteban Terzo, Kaitlin P. Curry, Harmony Salzler, Ivan Sabath, Grzegorz Zapotoczny, Daniel J. McKay, Zbigniew Dominski, William F. Marzluff, and Robert J. Duronio. Concentrating pre-mRNA processing factors in the histone locus body facilitates efficient histone mRNA biogenesis. Journal of Cell Biology, 213(5):557–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/213/5/557.

Tsuyama:2017:MDI

[TTU^{+}17] Taiichi Tsuyama, Asako Tsubouchi, Tadao Usui, Hiromi Imamura, and Tadashi Uemura. Mitochondrial dysfunction induces dendritic loss via eIF2α phosphorylation. Journal of Cell Biology, 216(3):815–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/3/815.

Tiscione:2019:DAM

[TVG^{+}19] Scott A. Tiscione, Oscar Vivas, Kenneth S. Ginsburg, Donald M. Bers, Daniel S. Ory, Luis F. Santana, Rose E. Dixon, and Eamonn J. Dickson. Disease-associated mutations in Niemann– Pick type C1 alter ER calcium signaling and neuronal plasticity. Journal of Cell Biology, 218(12):4141–4156, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/218/12/
4141/132543/Disease-associated-mutations-in-Niemann-
Pick-type.

Tipton:2017:GRS

[TWD^{+}17] Aaron R. Tipton, Jonathan D. Wren, John R. Daum, Joseph C. Siefert, and Gary J. Gorbsky. GTSE1 regulates spindle microtubule dynamics to control Aurora B kinase and Kif4A chromokinesin on chromosome arms. Journal of Cell Biology, 216(10):3117–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/10/3117.

Teo:2016:CRT

[TY16] Jessica L. Teo and Alpha S. Yap. Cycling Rho for tissue contraction. Journal of Cell Biology, 214(5):495–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/214/5/495.

Thomas:2015:NMM

[TYD^{+}15] Dustin G. Thomas, Aishwarya Yenepalli, Celine Marie Denais, Andrew Rape, Jordan R. Beach, Yu li Wang, William P. Schiemann, Harihara Baskaran, Jan Lammerding, and Thomas T. Egelhoff. Non-muscle myosin IIB is critical for nuclear translocation during 3D invasion. Journal of Cell Biology, 210(4): 583–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/4/583.

Takao:2019:TCD

[TYK19] Daisuke Takao, Shohei Yamamoto, and Daiju Kitagawa. A theory of centriole duplication based on self-organized spatial pattern formation. Journal of Cell Biology, 218(11):3537– ??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/11/3537.

Tosoni:2015:NPC

[TZC^{+}15] Daniela Tosoni, Silvia Zecchini, Marco Coazzoli, Ivan Colaluca, Giovanni Mazzarol, Alicia Rubio, Michele Caccia, Emanuele Villa, Olav Zilian, Pier Paolo Di Fiore, and Salvatore Pece. The Numb/p53 circuitry couples replicative self-renewal and tumor suppression in mammary epithelial cells. Journal of Cell Biology, 211(4):845–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/4/845.

Uribe:2015:HDK

[UBBSM15] Rosa A. Uribe, Ailín L. Buzzi, Marianne E. Bronner, and Pablo H. Strobl-Mazzulla. Histone demethylase KDM4B regulates otic vesicle invagination via epigenetic control of Dlx3 expression. Journal of Cell Biology, 211(4):815–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/211/4/
815.

Urbancic:2017:FOS

[UBR^{+}17] Vasja Urbančič, Richard Butler, Benjamin Richier, Manuel Peter, Julia Mason, Frederick J. Livesey, Christine E. Holt, and Jennifer L. Gallop. Filopodyan: an open-source pipeline for the analysis of filopodia. Journal of Cell Biology, 216(10):3405–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/10/3405.

Urban:2016:RHP

[UDH^{+}16] Vaclav Urban, Jana Dobrovolna, Daniela Hühn, Jana Fryzelkova, Jiri Bartek, and Pavel Janscak. RECQ5 helicase promotes resolution of conflicts between replication and transcription in human cells. Journal of Cell Biology, 214(4):401–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/214/4/
401.

Uchimura:2015:FIP

[UFT^{+}15] Seiichi Uchimura, Takashi Fujii, Hiroko Takazaki, Rie Ayukawa, Yosuke Nishikawa, Itsushi Minoura, You Hachikubo, Genji Kurisu, Kazuo Sutoh, Takahide Kon, Keiichi Namba, and Etsuko Muto. A flipped ion pair at the dynein–microtubule interface is critical for dynein motility and ATPase activation. Journal of Cell Biology, 208(2):211–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/2/211.

Urbina:2018:SOE

[UGG18] Fabio L. Urbina, Shawn M. Gomez, and Stephanie L. Gupton. Spatiotemporal organization of exocytosis emerges during neuronal shape change. Journal of Cell Biology, 217(3):1113–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/3/1113.

Ulmschneider:2016:IIP

[UGHB^{+}16] Bryne Ulmschneider, Bree K. Grillo-Hill, Marimar Benitez, Dinara R. Azimova, Diane L. Barber, and Todd G. Nystul. Increased intracellular pH is necessary for adult epithelial and embryonic stem cell differentiation. Journal of Cell Biology, 215 (3):345–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/215/3/345.

Ungricht:2015:DRM

[UKHK15] Rosemarie Ungricht, Michael Klann, Peter Horvath, and Ulrike Kutay. Diffusion and retention are major determinants of protein targeting to the inner nuclear membrane. Journal of Cell Biology, 209(5):687–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/5/687.

Ullal:2015:DFK

[UMC^{+}15] Pranav Ullal, Nathan A. McDonald, Jun-Song Chen, Libera Lo Presti, Rachel H. Roberts-Galbraith, Kathleen L. Gould, and Sophie G. Martin. The DYRK-family kinase Pom1 phosphorylates the F–BAR protein Cdc15 to prevent division at cell poles. Journal of Cell Biology, 211(3):653–??, November 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/211/3/653.

Ullal:2017:RDF

[UMC^{+}17] Pranav Ullal, Nathan A. McDonald, Jun-Song Chen, Libera Lo Presti, Rachel H. Roberts-Galbraith, Kathleen L. Gould, and Sophie G. Martin. Retraction: The DYRK-family kinase Pom1 phosphorylates the F–BAR Cdc15 to prevent division at cell poles. Journal of Cell Biology, 216(3):849–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/3/
849.

Uematsu:2016:ARS

[UOT^{+}16] Keiji Uematsu, Fumihiko Okumura, Syunsuke Tonogai, Akiko Joo-Okumura, Dawit Hailu Alemayehu, Akihiko Nishikimi, Yoshinori Fukui, Kunio Nakatsukasa, and Takumi Kamura.

ASB7 regulates spindle dynamics and genome integrity by targeting DDA3 for proteasomal degradation. Journal of Cell Biology, 215(1):95–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/215/1/95.

Varlakhanova:2018:SFD

[VAB^{+}18] Natalia V. Varlakhanova, Frances J. D. Alvarez, Tyler M. Brady, Bryan A. Tornabene, Christopher J. Hosford, Joshua S. Chappie, Peijun Zhang, and Marijn G. J. Ford. Structures of the fungal dynamin-related protein Vps1 reveal a unique, open helical architecture. Journal of Cell Biology, 217(10):3608–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/10/3608.

Vion:2018:PCS

[VAKB^{+}18] Anne-Clémence Vion, Silvanus Alt, Alexandra Klaus-Bergmann, Anna Szymborska, Tuyu Zheng, Tijana Perovic, Adel Hammoutene, Marta Bastos Oliveira, Eireen Bartels-Klein, Irene Hollfinger, Pierre-Emmanuel Rautou, Miguel O. Bernabeu, and Holger Gerhardt. Primary cilia sensitize endothelial cells to BMP and prevent excessive vascular regression. Journal of Cell Biology, 217(5):1651–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/5/1651.

VanDenBosch:2019:HMA

[Van19] Ludo Van Den Bosch. HDAC6 and Miro1: Another interaction causing trouble in neurons. Journal of Cell Biology, 218 (6):1769–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/6/1769.

Verweij:2018:CQE

[VBJ^{+}18a] Frederik Johannes Verweij, Maarten P. Bebelman, Connie R. Jimenez, Juan J. Garcia-Vallejo, Hans Janssen, Jacques Neefjes, Jaco C. Knol, Richard de Goeij-de Haas, Sander R. Piersma, S. Rubina Baglio, Matthijs Verhage, Jaap M. Middeldorp, Anoek Zomer, Jacco van Rheenen, Marc G. Coppolino, Ilse Hurbain, Graça Raposo, Martine J. Smit, Ruud F. G. Toonen, Guillaume van Niel, and D. Michiel Pegtel. Correction:

Quantifying exosome secretion from single cells reveals a modulatory role for GPCR signaling. Journal of Cell Biology, 217 (3):1157–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/3/1157. See [VBJ^{+}18b].

Verweij:2018:QES

[VBJ^{+}18b] Frederik Johannes Verweij, Maarten P. Bebelman, Connie R. Jimenez, Juan J. Garcia-Vallejo, Hans Janssen, Jacques Neefjes, Jaco C. Knol, Richard de Goeij-de Haas, Sander R. Piersma, S. Rubina Baglio, Matthijs Verhage, Jaap M. Middeldorp, Anoek Zomer, Jacco van Rheenen, Marc G. Coppolino, Ilse Hurbain, Graça Raposo, Martine J. Smit, Ruud F. G. Toonen, Guillaume van Niel, and D. Michiel Pegtel. Quantifying exosome secretion from single cells reveals a modulatory role for GPCR signaling. Journal of Cell Biology, 217(3):1129–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/

217/3/1129. See correction [VBJ^{+}18a].

Villeneuve:2018:USF

[VBL^{+}18] Julien Villeneuve, Laia Bassaganyas, Sebastien Lepreux, Marioara Chiritoiu, Pierre Costet, Jean Ripoche, Vivek Malhotra, and Randy Schekman. Unconventional secretion of FABP4 by endosomes and secretory lysosomes. Journal of Cell Biology, 217(2):649–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/2/649.

vonBudingen:2015:MOG

[vBMG^{+}15] H.-Christian von Büdingen, Feng Mei, Ariele Greenfield, Sarah Jahn, Yun-An A. Shen, Hugh H. Reid, David D. McKemy, and Jonah R. Chan. The myelin oligodendrocyte glycoprotein directly binds nerve growth factor to modulate central axon circuitry. Journal of Cell Biology, 210(6):891–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/210/6/891.

Vandersmissen:2015:EMT

[VCD^{+}15] Ine Vandersmissen, Sander Craps, Maarten Depypere, Giulia Coppiello, Nick van Gastel, Frederik Maes, Geert Carmeliet, Jan Schrooten, Elizabeth A. V. Jones, Lieve Umans, Roland Devlieger, Michel Koole, Olivier Gheysens, An Zwijsen, Xabier L. Aranguren, and Aernout Luttun. Endothelial Msx1 transduces hemodynamic changes into an arteriogenic remodeling response. Journal of Cell Biology, 210(7):1239–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/210/7/
1239.

vanDrogen:2019:MSI

[vDMR^{+}19] Frank van Drogen, Ranjan Mishra, Fabian Rudolf, Michal J. Walczak, Sung Sik Lee, Wolfgang Reiter, Björn Hegemann, Serge Pelet, Ilse Dohnal, Andres Binolfi, Zinaida Yudina, Philipp Selenko, Gerhard Wider, Gustav Ammerer, and Matthias Peter. Mechanical stress impairs pheromone signaling via Pkc1-mediated regulation of the MAPK scaffold Ste5. Journal of Cell Biology, 218(9):3117–??, September 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/9/3117.

vanderVaart:2017:TSE

[vdVFM^{+}17] Babet van der Vaart, Josef Fischböck, Christine Mieck, Peter Pichler, Karl Mechtler, René H. Medema, and Stefan Westermann. TORC1 signaling exerts spatial control over microtubule dynamics by promoting nuclear export of Stu2. Journal of Cell Biology, 216(11):3471–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/11/3471.

Verlhac:2016:MCK

[Ver16] Marie-Hélène Verlhac. Mother centrioles are kicked out so that starfish zygote can grow. Journal of Cell Biology, 212 (7):759–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/212/7/759.

Verlhac:2018:ASD

[Ver18] Marie-Hélène Verlhac. An actin shell delays oocyte chromosome capture by microtubules. Journal of Cell Biology, 217(8): 2601–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/8/2601.

Volkov:2015:CPF

[VGA^{+}15] Vladimir A. Volkov, Paula M. Grissom, Vladimir K. Arzhanik, Anatoly V. Zaytsev, Kutralanathan Renganathan, Tristan McClure-Begley, William M. Old, Natalie Ahn, and J. Richard McIntosh. Centromere protein F includes two sites that couple efficiently to depolymerizing microtubules. Journal of Cell Biology, 209(6):813–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/6/813.

Vos:2017:CPE

[VGB^{+}17] Melissa Vos, Ann Geens, Claudia Böhm, Liesbeth Deaulmerie, Jef Swerts, Matteo Rossi, Katleen Craessaerts, Elvira P. Leites, Philip Seibler, Aleksandar Rakovic, Thora Lohnau, Bart De Strooper, Sarah-Maria Fendt, Vanessa A. Morais, Christine Klein, and Patrik Verstreken. Cardiolipin promotes electron transport between ubiquinone and complex I to rescue PINK1 deficiency. Journal of Cell Biology, 216(3):695–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/216/3/695.

vanGisbergen:2018:ASF

[vGWC^{+}18] Peter A. C. van Gisbergen, Shu-Zon Wu, Mingqin Chang, Kelli A. Pattavina, Madelaine E. Bartlett, and Magdalena Bezanilla. An ancient Sec10–formin fusion provides insights into actin-mediated regulation of exocytosis. Journal of Cell Biology, 217(3):945–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/3/945.

Vasileva:2017:MMF

[VGY^{+}17] Vanya Vasileva, Marek Gierlinski, Zuojun Yue, Nicola O’Reilly, Etsushi Kitamura, and Tomoyuki U. Tanaka. Molecular mechanisms facilitating the initial kinetochore encounter with spindle microtubules. Journal of Cell Biology, 216(6):1609–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/6/
1609.

Vargas-Hurtado:2018:WCC

[VHB18] Diana Vargas-Hurtado and Renata Basto. When E-cadherin is away, centrosomes can play. Journal of Cell Biology, 217 (1):11–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/1/11.

vanHelden:2015:TNC

[vHGD^{+}15] Mary J. van Helden, Steven Goossens, Cécile Daussy, Anne-Laure Mathieu, Fabrice Faure, Antoine Marçais, Niels Vandamme, Natalie Farla, Katia Mayol, Sébastien Viel, Sophie Degouve, Emilie Debien, Eve Seuntjens, Andrea Conidi, Julie Chaix, Philippe Mangeot, Simon de Bernard, Laurent Buffat, Jody Haigh, Danny Huylebroeck, Bart N. Lambrecht, Geert Berx, and Thierry Walzer. Terminal NK cell maturation is controlled by concerted actions of T-bet and Zeb2 and is essential for melanoma rejection. Journal of Cell Biology, 211 (3):??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/3/2113OIA260.

Vogel:2015:CSA

[VKJ^{+}15] Georg F. Vogel, Katharina M. C. Klee, Andreas R. Janecke, Thomas Müller, Michael W. Hess, and Lukas A. Huber. Cargoselective apical exocytosis in epithelial cells is conducted by Myo5B, Slp4a, Vamp7, and Syntaxin 3. Journal of Cell Biology, 211(3):587–??, November 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/211/3/587.

Vogtle:2015:FLP

[VKT^{+}15] F.-Nora Vögtle, Michael Keller, Asli A. Taskin, Susanne E. Horvath, Xue Li Guan, Claudia Prinz, Magdalena Opalińska, Carina Zorzin, Martin van der Laan, Markus R. Wenk, Rolf Schubert, Nils Wiedemann, Martin Holzer, and Chris Meisinger. The fusogenic lipid phosphatidic acid promotes the biogenesis of mitochondrial outer membrane protein Ugo1. Journal of Cell Biology, 210(6):951–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/6/951.

Viero:2015:TDR

[VLP^{+}15] Gabriella Viero, Lorenzo Lunelli, Andrea Passerini, Paolo Bianchini, Robert J. Gilbert, Paola Bernabò, Toma Tebaldi, Alberto Diaspro, Cecilia Pederzolli, and Alessandro Quattrone. Three distinct ribosome assemblies modulated by translation are the building blocks of polysomes. Journal of Cell Biology, 208 (5):581–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/208/5/581.

vanLeeuwen:2018:MSP

[vLvdKR18] Wessel van Leeuwen, Felix van der Krift, and Catherine Rabouille. Modulation of the secretory pathway by aminoacid starvation. Journal of Cell Biology, 217(7):2261–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/7/
2261.

Vichas:2015:SFH

[VLZ15] Athea Vichas, Matthew T. Laurie, and Jennifer A. Zallen. The Ski2-family helicase Obelus regulates Crumbs alternative splicing and cell polarity. Journal of Cell Biology, 211(5): 1011–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/211/5/1011.

Verma:2019:DOB

[VM19] Vikash Verma and Thomas J. Maresca. Direct observation of branching MT nucleation in living animal cells. Journal of Cell Biology, 218(9):2829–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/9/2829.

Valansi:2017:AHG

[VML^{+}17] Clari Valansi, David Moi, Evgenia Leikina, Elena Matveev, Martín Graña, Leonid V. Chernomordik, Héctor Romero, Pablo S. Aguilar, and Benjamin Podbilewicz. Arabidopsis HAP2/GCS1 is a gamete fusion protein homologous to somatic and viral fusogens. Journal of Cell Biology, 216(3):571–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/3/571.

Vlahakis:2016:CCR

[VMP16] Ariadne Vlahakis, Nerea Lopez Muniozguren, and Ted Powers. Calcium channel regulator Mid1 links TORC2-mediated changes in mitochondrial respiration to autophagy. Journal of Cell Biology, 215(6):779–??, December 2016. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/215/6/779.

Venditti:2019:ASA

[VMR^{+}19] Rossella Venditti, Maria Chiara Masone, Laura Rita Rega, Giuseppe Di Tullio, Michele Santoro, Elena Polishchuk, Ivan Castello Serrano, Vesa M. Olkkonen, Akihiro Harada, Diego L. Medina, Raffaele La Montagna, and Maria Antonietta De Matteis. The activity of Sac1 across ER–TGN contact sites requires the four-phosphate-adaptor-protein-1. Journal of Cell Biology, 218(3):783–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/3/783.

Vieillard:2016:TZA

[VPD^{+}16] Jennifer Vieillard, Marie Paschaki, Jean-Luc Duteyrat, Céline Augière, Elisabeth Cortier, Jean-André Lapart, Joëlle Thomas, and Bénédicte Durand. Transition zone assembly and its contribution to axoneme formation in Drosophila male germ cells. Journal of Cell Biology, 214(7):875–??, September 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/214/7/875.

Vizcarra:2017:AFA

[VQ17] Christina L. Vizcarra and Margot E. Quinlan. Actin filament assembly by bacterial factors VopL/F: Which end is up? Journal of Cell Biology, 216(5):1211–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/5/1211.

Verlhac:2018:STM

[VR18] Pauline Verlhac and Fulvio Reggiori. Sorting the trash: Micronucleophagy gets selective. Journal of Cell Biology, 217(8): 2605–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/8/2605.

VanderLugt:2017:TDT

[VRK^{+}17] Bryan Vander Lugt, Jeremy Riddell, Aly A. Khan, Jason A. Hackney, Justin Lesch, Jason DeVoss, Matthew T. Weirauch, Harinder Singh, and Ira Mellman. Transcriptional determinants of tolerogenic and immunogenic states during dendritic cell maturation. Journal of Cell Biology, 216(3):779–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/3/
779.

Venditti:2019:MDG

[VRM^{+}19] Rossella Venditti, Laura Rita Rega, Maria Chiara Masone, Michele Santoro, Elena Polishchuk, Daniela Sarnataro, Simona Paladino, Sabato D’Auria, Antonio Varriale, Vesa M. Olkkonen, Giuseppe Di Tullio, Roman Polishchuk, and Maria Antonietta De Matteis. Molecular determinants of ER–Golgi contacts identified through a new FRET–FLIM system. Journal of Cell Biology, 218(3):1055–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/3/1055.

vanSteensel:2015:SGT

[vS15] Bas van Steensel. A short guide to technology development in cell biology. Journal of Cell Biology, 208(6):655–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/208/6/
655.

Velazquez:2016:LDM

[VTG^{+}16] Ariadna P. Velázquez, Takashi Tatsuta, Ruben Ghillebert, Ingmar Drescher, and Martin Graef. Lipid droplet–mediated ER homeostasis regulates autophagy and cell survival during starvation. Journal of Cell Biology, 212(6):621–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/212/6/621.

vanVugt:2017:SPS

[vV17a] Marcel A. T. M. van Vugt. Shutting down the power supply for DNA repair in cancer cells. Journal of Cell Biology, 216(2): 295–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/2/295.

Vijayan:2017:APC

[VV17b] Vinoy Vijayan and Patrik Verstreken. Autophagy in the presynaptic compartment in health and disease. Journal of Cell Biology, 216(7):1895–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/7/1895.

Viotti:2018:SER

[VWM^{+}18] Manuel Viotti, Catherine Wilson, Mark McCleland, Hartmut Koeppen, Benjamin Haley, Suchit Jhunjhunwala, Christiaan Klijn, Zora Modrusan, David Arnott, Marie Classon, Jean-Philippe Stephan, and Ira Mellman. SUV420H2 is an epigenetic regulator of epithelial/mesenchymal states in pancreatic cancer. Journal of Cell Biology, 217(2):763–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/217/2/763.

Vossel:2015:TRP

[VXF^{+}15] Keith A. Vossel, Jordan C. Xu, Vira Fomenko, Takashi Miyamoto, Elsa Suberbielle, Joseph A. Knox, Kaitlyn Ho, Daniel H. Kim, Gui-Qiu Yu, and Lennart Mucke. Tau reduction prevents Aβ-induced axonal transport deficits by blocking activation of GSK3β. Journal of Cell Biology, 209(3):419–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/209/3/
419.

Venkei:2018:EMA

[VY18] Zsolt G. Venkei and Yukiko M. Yamashita. Emerging mechanisms of asymmetric stem cell division. Journal of Cell Biology, 217(11):3785–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/11/3785.

Valverde:2019:ATL

[VYB^{+}19] Diana P. Valverde, Shenliang Yu, Venkata Boggavarapu, Nikit Kumar, Joshua A. Lees, Thomas Walz, Karin M. Reinisch, and Thomas J. Melia. ATG2 transports lipids to promote autophagosome biogenesis. Journal of Cell Biology, 218(6):1787– ??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/6/1787.

Victoria:2017:SPL

[VZ17] Guiliana Soraya Victoria and Chiara Zurzolo. The spread of prion-like proteins by lysosomes and tunneling nanotubes: Implications for neurodegenerative diseases. Journal of Cell Biology, 216(9):2633–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/216/9/2633.

Vermunt:2019:IGR

[VZB19] Marit W. Vermunt, Di Zhang, and Gerd A. Blobel. The interdependence of gene-regulatory elements and the 3D genome. Journal of Cell Biology, 218(1):12–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/1/12.

Volin:2018:MSP

[VZFG^{+}18] Marina Volin, Maayan Zohar-Fux, Oren Gonen, Lilach Porat-Kuperstein, and Hila Toledano. microRNAs selectively protect hub cells of the germline stem cell niche from apoptosis. Journal of Cell Biology, 217(11):3829–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/11/3829.

Warren:2015:POM

[War15] Graham Warren. In praise of other model organisms. Journal of Cell Biology, 208(4):387–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/208/4/387.

Wu:2018:AMC

[WB18] Shu-Zon Wu and Magdalena Bezanilla. Actin and microtubule cross talk mediates persistent polarized growth. Journal of Cell Biology, 217(10):3531–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/217/10/3531.

Wilmes:2015:RDD

[WBL^{+}15] Stephan Wilmes, Oliver Beutel, Zhi Li, Véronique François-Newton, Christian P. Richter, Dennis Janning, Cindy Kroll, Patrizia Hanhart, Katharina Hötte, Changjiang You, Gilles Uzé, Sandra Pellegrini, and Jacob Piehler. Receptor dimerization dynamics as a regulatory valve for plasticity of type I interferon signaling. Journal of Cell Biology, 209(4):579–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/209/4/579.

Wang:2018:SDD

[WBNH18] Ying Wang, Robyn Branicky, Alycia Noë, and Siegfried Hekimi. Superoxide dismutases: Dual roles in controlling ROS damage and regulating ROS signaling. Journal of Cell Biology, 217 (6):1915–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/6/1915.

Wang:2018:KRC

[WCL^{+}18] Fengrong Wang, Song Chen, Hans B. Liu, Carole A. Parent, and Pierre A. Coulombe. Keratin 6 regulates collective keratinocyte migration by altering cell–cell and cell–matrix adhesion. Journal of Cell Biology, 217(12):4314–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/12/4314.

Wang:2016:CEP

[WCY^{+}16a] Chenran Wang, Song Chen, Syn Yeo, Gizem Karsli-Uzunbas, Eileen White, Noboru Mizushima, Herbert W. Virgin, and Jun-Lin Guan. Correction: Elevated p62/SQSTM1 determines the fate of autophagy-deficient neural stem cells by increasing superoxide. Journal of Cell Biology, 212(7):879–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/212/7/879.

Wang:2016:EPS

[WCY^{+}16b] Chenran Wang, Song Chen, Syn Yeo, Gizem Karsli-Uzunbas, Eileen White, Noboru Mizushima, Herbert W. Virgin, and Jun-Lin Guan. Elevated p62/SQSTM1 determines the fate of autophagy-deficient neural stem cells by increasing superoxide. Journal of Cell Biology, 212(5):545–??, February 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/212/5/545.

Wang:2015:YRR

[WDM^{+}15] Juan Wang, Saralin Davis, Shekar Menon, Jinzhong Zhang, Jingzhen Ding, Serena Cervantes, Elizabeth Miller, Yu Jiang, and Susan Ferro-Novick. Ypt1/Rab1 regulates Hrr25/CK1δ kinase activity in ER–Golgi traffic and macroautophagy. Journal of Cell Biology, 210(2):273–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/2/273.

Webb:2017:GEP

[WDW^{+}17] Bradley A. Webb, Anne M. Dosey, Torsten Wittmann, Justin M. Kollman, and Diane L. Barber. The glycolytic enzyme phosphofructokinase-1 assembles into filaments. Journal of Cell Biology, 216(8):2305–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/8/2305.

Wenz:2015:SCF

[WEQ^{+}15] Lena-Sophie Wenz, Lars Ellenrieder, Jian Qiu, Maria Bohnert, Nicole Zufall, Martin van der Laan, Nikolaus Pfanner, Nils Wiedemann, and Thomas Becker. Sam37 is crucial for formation of the mitochondrial TOM–SAM supercomplex, thereby promoting β-barrel biogenesis. Journal of Cell Biology, 210(7): 1047–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/7/1047.

Wynne:2015:KFC

[WF15] David J. Wynne and Hironori Funabiki. Kinetochore function is controlled by a phospho-dependent coexpansion of inner and outer components. Journal of Cell Biology, 210(6): 899–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/6/899.

Wakatsuki:2015:OSD

[WFOA15] Shuji Wakatsuki, Akiko Furuno, Makiko Ohshima, and Toshiyuki Araki. Oxidative stress-dependent phosphorylation activates ZNRF1 to induce neuronal/axonal degeneration. Journal of Cell Biology, 211(4):881–??, November 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/211/4/881.

Westhorpe:2015:CFC

[WFS15] Frederick G. Westhorpe, Colin J. Fuller, and Aaron F. Straight. A cell-free CENP–A assembly system defines the chromatin requirements for centromere maintenance. Journal of Cell Biology, 209(6):789–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/209/6/789.

Wagner:2016:LRA

[WG16] Elizabeth Wagner and Michael Glotzer. Local RhoA activation induces cytokinetic furrows independent of spindle position and cell cycle stage. Journal of Cell Biology, 213(6):641–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/213/6/
641.

White:2018:CPS

[WGHE^{+}18] Katharine A. White, Bree K. Grillo-Hill, Mario Esquivel, Jobelle Peralta, Vivian N. Bui, Ismahan Chire, and Diane L. Barber. β-catenin is a pH sensor with decreased stability at higher intracellular pH. Journal of Cell Biology, 217(11): 3965–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/11/3965.

Wissel:2018:TRT

[WHB^{+}18] Sebastian Wissel, Heike Harzer, François Bonnay, Thomas R. Burkard, Ralph A. Neumüller, and Juergen A. Knoblich. Timeresolved transcriptomics in neural stem cells identifies a v-ATPase/ Notch regulatory loop. Journal of Cell Biology, 217(9):3285–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/9/3285.

Wang:2019:RRP

[WHC^{+}19] Guang Wang, Huai-Bin Hu, Yan Chang, Yan Huang, Zeng-Qing Song, Shi-Bo Zhou, Liang Chen, Yu-Cheng Zhang, Min Wu, Hai-Qing Tu, Jin-Feng Yuan, Na Wang, Xin Pan, Ai-Ling Li, Tao Zhou, Xue-Min Zhang, Kun He, and Hui-Yan Li. Rab7 regulates primary cilia disassembly through cilia excision. Journal of Cell Biology, 218(12):4030–4041, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

https://rupress.org/jcb/article/218/12/4030/132511/
Rab7-regulates-primary-cilia-disassembly-through.

Watanabe:2019:CMC

[WHiO^{+}19] Reito Watanabe, Masatoshi Hara, Ei ichi Okumura, Solène Hervé, Daniele Fachinetti, Mariko Ariyoshi, and Tatsuo Fukagawa. CDK1-mediated CENP-C phosphorylation modulates CENP-A binding and mitotic kinetochore localization. Journal of Cell Biology, 218(12):4042–4062, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

https://rupress.org/jcb/article/218/12/4042/132534/
CDK1-mediated-CENP-C-phosphorylation-modulates.

Winsor:2017:CCS

[WHL17] James Winsor, David D. Hackney, and Tina H. Lee. The crossover conformational shift of the GTPase atlastin provides the energy driving ER fusion. Journal of Cell Biology, 216 (5):1321–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/5/1321.

Walther:2018:QMH

[WHP^{+}18] Nike Walther, M. Julius Hossain, Antonio Z. Politi, Birgit Koch, Moritz Kueblbeck, Øyvind Ødegård-Fougner, Marko Lampe, and Jan Ellenberg. A quantitative map of human Condensins provides new insights into mitotic chromosome architecture. Journal of Cell Biology, 217(7):2309–??, July 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/7/2309.

Wang:2019:CCA

[WHS^{+}19] Yuxiao Wang, Walter Huynh, Taylor D. Skokan, Wen Lu, Arthur Weiss, and Ronald D. Vale. CRACR2a is a calciumactivated dynein adaptor protein that regulates endocytic traffic. Journal of Cell Biology, 218(5):1619–??, May 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/5/1619.

Wilkinson:2015:BCB

[Wil15] David G. Wilkinson. Balancing cell behavior at boundaries. Journal of Cell Biology, 208(6):659–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/6/659.

Wood:2017:CHR

[WIS^{+}17] Megan N. Wood, Noboru Ishiyama, Indira Singaram, Connie M. Chung, Annette S. Flozak, Alex Yemelyanov, Mitsu Ikura, Wonhwa Cho, and Cara J. Gottardi. α-catenin homodimers are recruited to phosphoinositide-activated membranes to promote adhesion. Journal of Cell Biology, 216(11): 3767–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/11/3767.

Watanabe:2015:TER

[WKM^{+}15] Takashi Watanabe, Mai Kakeno, Toshinori Matsui, Ikuko Sugiyama, Nariko Arimura, Kenji Matsuzawa, Aya Shirahige, Fumiyoshi Ishidate, Tomoki Nishioka, Shinichiro Taya, Mikio Hoshino, and Kozo Kaibuchi. TTBK2 with EB1/3 regulates microtubule dynamics in migrating cells through KIF2A phosphorylation. Journal of Cell Biology, 210(5):737–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/210/5/
737.

Woods:2015:PER

[WKW^{+}15] Benjamin Woods, Chun-Chen Kuo, Chi-Fang Wu, Trevin R. Zyla, and Daniel J. Lew. Polarity establishment requires localized activation of Cdc42. Journal of Cell Biology, 211(1):19–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
211/1/19.

Wood:2017:NTH

[WLC^{+}17] Laura A. Wood, Gabrielle Larocque, Nicholas I. Clarke, Sourav Sarkar, and Stephen J. Royle. New tools for “hot-wiring” clathrin-mediated endocytosis with temporal and spatial precision. Journal of Cell Biology, 216(12):4351–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/
12/4351.

Wu:2016:DTU

[WLJ16] Xi Wu, Lanlan Li, and Hui Jiang. Doa1 targets ubiquitinated substrates for mitochondria-associated degradation. Journal of Cell Biology, 213(1):49–??, April 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/213/1/49.

Wu:2018:MIM

[WLJ18] Xi Wu, Lanlan Li, and Hui Jiang. Mitochondrial innermembrane protease Yme1 degrades outer-membrane proteins Tom22 and Om45. Journal of Cell Biology, 217(1):139–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/1/139.

West:2015:RPT

[WLM^{+}15] Ryan J. H. West, Yubing Lu, Bruno Marie, Fen-Biao Gao, and Sean T. Sweeney. Rab8, POSH, and TAK1 regulate synaptic growth in a Drosophila model of frontotemporal dementia. Journal of Cell Biology, 208(7):931–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/7/931.

Willet:2015:FBC

[WMB^{+}15] Alaina H. Willet, Nathan A. McDonald, K. Adam Bohnert, Michelle A. Baird, John R. Allen, Michael W. Davidson, and Kathleen L. Gould. The F–BAR Cdc15 promotes contractile ring formation through the direct recruitment of the formin Cdc12. Journal of Cell Biology, 208(4):391–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/208/4/391.

Winsor:2018:GHP

[WMH^{+}18] James Winsor, Ursula Machi, Qixiu Han, David D. Hackney, and Tina H. Lee. GTP hydrolysis promotes disassembly of the atlastin crossover dimer during ER fusion. Journal of Cell Biology, 217(12):4184–??, December 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/12/4184.

West:2016:SSR

[WMK^{+}16] Jason A. West, Mari Mito, Satoshi Kurosaka, Toru Takumi, Chiharu Tanegashima, Takeshi Chujo, Kaori Yanaka, Robert E. Kingston, Tetsuro Hirose, Charles Bond, Archa Fox, and Shinichi Nakagawa. Structural, super-resolution microscopy analysis of paraspeckle nuclear body organization. Journal of Cell Biology, 214(7):817–??, September 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/7/817.

Woodruff:2018:PSB

[Woo18] Jeffrey B. Woodruff. Phase separation of BuGZ promotes Aurora A activation and spindle assembly. Journal of Cell Biology, 217(1):09–??, January 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/1/09.

Wordeman:2019:GTL

[Wor19] Linda Wordeman. GTP-tubulin loves microtubule plus ends but marries the minus ends. Journal of Cell Biology, 218(9): 2822–??, September 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/9/2822.

Wang:2018:TCR

[WPA^{+}18] Zhao Wang, Lindsey W. Plasschaert, Shivani Aryal, Nicole A. Renaud, Zinger Yang, Rayman Choo-Wing, Angelica D. Pessotti, Nathaniel D. Kirkpatrick, Nadire R. Cochran, Walter Carbone, Rob Maher, Alicia Lindeman, Carsten Russ, John Reece-Hoyes, Gregory McAllister, Gregory R. Hoffman, Guglielmo Roma, and Aron B. Jaffe. TRRAP is a central regulator of human multiciliated cell formation. Journal of Cell Biology, 217(6):1941–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/6/1941.

Wigerius:2018:PPA

[WQD^{+}18] Michael Wigerius, Dylan Quinn, Antonios Diab, Leanne Clattenburg, Annette Kolar, Jiansong Qi, Stefan R. Krueger, and James P. Fawcett. The polarity protein Angiomotin p130 controls dendritic spine maturation. Journal of Cell Biology, 217 (2):715–??, February 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/2/715.

Wright:2015:UAP

[WRGB^{+}15] Margaret C. Wright, Erin G. Reed-Geaghan, Alexa M. Bolock, Tomoyuki Fujiyama, Mikio Hoshino, and Stephen M. Maricich. Unipotent, Atoh1^{+} progenitors maintain the Merkel cell population in embryonic and adult mice. Journal of Cell Biology, 208(3):367–??, February 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/3/367.

Wang:2016:DRB

[WRH^{+}16] Wei Wang, Asit Rai, Eun-Mi Hur, Zeev Smilansky, Karen T. Chang, and Kyung-Tai Min. DSCR1 is required for both axonal growth cone extension and steering. Journal of Cell Biology, 213(4):451–??, May 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/213/4/451.

Wang:2015:NPE

[WRV15] Shuoshuo Wang, Adriana Reuveny, and Talila Volk. Nesprin provides elastic properties to muscle nuclei by cooperating with spectraplakin and EB1. Journal of Cell Biology, 209(4):529–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
209/4/529.

Watkins:2018:LSI

[WS18] Simon C. Watkins and Claudette M. St. Croix. Light sheet imaging comes of age. Journal of Cell Biology, 217(5):1567–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/5/1567.

Wang:2017:PCM

[WSDY17] Shaohe Wang, Rei Sekiguchi, William P. Daley, and Kenneth M. Yamada. Patterned cell and matrix dynamics in branching morphogenesis. Journal of Cell Biology, 216(3): 559–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/3/559.

Wang:2018:MLC

[WSP^{+}18] Shuoshuo Wang, Elizabeth Stoops, Unnikannan C. P., Barak Markus, Adriana Reuveny, Elly Ordan, and Talila Volk. Mechanotransduction via the LINC complex regulates DNA replication in myonuclei. Journal of Cell Biology, 217(6):2005– ??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/6/2005.

Wang:2019:MYC

[WTB^{+}19] Xin Wang, Wei Tian, Bryan T. Banh, Bethanie-Michelle Statler, Jie Liang, and David E. Stone. Mating yeast cells use an intrinsic polarity site to assemble a pheromone-gradient tracking machine. Journal of Cell Biology, 218(11):3730–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/11/3730.

Wang:2019:AFH

[WTC^{+}19] Ruoxi Wang, Jieqiong Tan, Tingting Chen, Hailong Han, Runyi Tian, Ya Tan, Yiming Wu, Jingyi Cui, Fang Chen, Jie Li, Lu Lv, Xinjie Guan, Shuai Shang, Jiahong Lu, and Zhuohua Zhang. ATP13A2 facilitates HDAC6 recruitment to lysosome to promote autophagosome–lysosome fusion. Journal of Cell Biology, 218(1):267–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/1/267.

Wakatsuki:2017:GMP

[WTSA17] Shuji Wakatsuki, Shinji Tokunaga, Megumi Shibata, and Toshiyuki Araki. GSK3B-mediated phosphorylation of MCL1 regulates axonal autophagy to promote Wallerian degeneration. Journal of Cell Biology, 216(2):477–??, February 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/2/477.

Wu:2017:PWC

[Wu17] Min Wu. Pulses and waves of contractility. Journal of Cell Biology, 216(12):3899–??, December 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/12/3899.

Williamson:2018:SCD

[WV18a] Adam P. Williamson and Ronald D. Vale. Spatial control of Draper receptor signaling initiates apoptotic cell engulfment. Journal of Cell Biology, 217(11):3977–??, November 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/11/3977.

Wynne:2018:CPD

[WV18b] Caitlin L. Wynne and Richard B. Vallee. Cdk1 phosphorylation of the dynein adapter Nde1 controls cargo binding from G2 to anaphase. Journal of Cell Biology, 217(9):3019–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/9/
3019.

Weng:2016:MDR

[WW16] Mo Weng and Eric Wieschaus. Myosin-dependent remodeling of adherens junctions protects junctions from Snail-dependent disassembly. Journal of Cell Biology, 212(2):219–??, January 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/212/2/
219.

Wang:2018:ETS

[WWT18] Qiao-Chu Wang, Xi Wang, and Tie-Shan Tang. EB1 traps STIM1 and regulates local store-operated Ca^{2+} entry. Journal of Cell Biology, 217(6):1899–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/6/1899.

Wheelock:2017:DRC

[WWTF17] Michael S. Wheelock, David J. Wynne, Boo Shan Tseng, and Hironori Funabiki. Dual recognition of chromatin and microtubules by INCENP is important for mitotic progression. Journal of Cell Biology, 216(4):925–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/4/925.

Wang:2018:IMT

[WWW^{+}18] Ke Wang, Lantian Wang, Jianshu Wang, Suli Chen, Min Shi, and Hong Cheng. Intronless mRNAs transit through nuclear speckles to gain export competence. Journal of Cell Biology, 217(11):3912–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/11/3912.

Wang:2018:SMP

[WWY^{+}18] Xiang Wang, Xiaofan Wei, Yang Yuan, Qingrui Sun, Jun Zhan, Jing Zhang, Yan Tang, Feng Li, Lihua Ding, Qinong Ye, and Hongquan Zhang. Src-mediated phosphorylation converts FHL1 from tumor suppressor to tumor promoter. Journal of Cell Biology, 217(4):1335–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/4/1335.

Wei:2017:SII

[WWZ^{+}17] Xiaofan Wei, Xiang Wang, Jun Zhan, Yuhan Chen, Weigang Fang, Lingqiang Zhang, and Hongquan Zhang. Smurf1 inhibits integrin activation by controlling Kindlin-2 ubiquitination and degradation. Journal of Cell Biology, 216(5):1455–??, May 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/5/1455.

Wang:2018:ISL

[WWZ^{+}18] ShiHui Wang, ChenYu Wu, YueBin Zhang, QingLu Zhong, Hao Sun, WenPeng Cao, GaoXiang Ge, GuoHui Li, X. Frank Zhang, and JianFeng Chen. Integrin α4β7 switches its ligand specificity via distinct conformer-specific activation. Journal of Cell Biology, 217(8):2799–??, August 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/8/2799.

Wan:2018:PPD

[WXC^{+}18] Zhengpeng Wan, Chenguang Xu, Xiangjun Chen, Hengyi Xie, Zongyu Li, Jing Wang, Xingyu Ji, Haodong Chen, Qinghua Ji, Samina Shaheen, Yang Xu, Fei Wang, Zhuo Tang, Ji-Shen Zheng, Wei Chen, Jizhong Lou, and Wanli Liu. PI(4,5)P2 determines the threshold of mechanical force–induced B cell activation. Journal of Cell Biology, 217(7):2565–??, July 2018. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/217/7/2565.

Wang:2017:TNE

[WXFS17] Wei Wang, Zhi-Jie Xia, Jean-Claude Farré, and Suresh Subramani. TRIM37, a novel E3 ligase for PEX5-mediated peroxisomal matrix protein import. Journal of Cell Biology, 216(9): 2843–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/9/2843.

Wang:2017:AGF

[WYHG17] Chenran Wang, Syn Yeo, Michael A. Haas, and Jun-Lin Guan. Autophagy gene FIP200 in neural progenitors non-cell autonomously controls differentiation by regulating microglia. Journal of Cell Biology, 216(8):2581–??, August 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/8/2581.

Wang:2017:SMC

[WYoS17] Shuai Wang, Chien-I. Yang, and Shu ou Shan. SecA mediates cotranslational targeting and translocation of an inner membrane protein. Journal of Cell Biology, 216(11):3639– ??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/11/3639.

Wang:2019:FBD

[WYV^{+}19] Liang Wang, Ziyi Yan, Helena Vihinen, Ove Eriksson, Weihuan Wang, Rabah Soliymani, Yao Lu, Yaxin Xue, Eija Jokitalo, Jing Li, and Hongxia Zhao. FAM92A1 is a BAR domain protein required for mitochondrial ultrastructure and function. Journal of Cell Biology, 218(1):97–??, January 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/1/97.

Wan:2015:SFM

[WZC^{+}15] Yihan Wan, Xiaobin Zheng, Haiyang Chen, Yuxuan Guo, Hao Jiang, Xiaonan He, Xueliang Zhu, and Yixian Zheng. Splicing function of mitotic regulators links R-loop–mediated DNA damage to tumor cell killing. Journal of Cell Biology, 209(2):235–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
209/2/235.

Wang:2017:VDK

[WZG^{+}17] Wuyang Wang, Xiaoli Zhang, Qiong Gao, Maria Lawas, Lu Yu, Xiping Cheng, Mingxue Gu, Nirakar Sahoo, Xinran Li, Ping Li, Stephen Ireland, Andrea Meredith, and Haoxing Xu. A voltagedependent K^{+} channel in the lysosome is required for refilling lysosomal Ca^{2+} stores. Journal of Cell Biology, 216(6):1715–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/6/1715.

Wang:2019:HOA

[WZR19] ShiYu Wang, Zechuan Zhao, and Avital A. Rodal. Higher-order assembly of Sorting Nexin 16 controls tubulation and distribution of neuronal endosomes. Journal of Cell Biology, 218(8): 2600–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/8/2600.

Xia:2018:NRS

[XIZ^{+}18] Yuntao Xia, Irena L. Ivanovska, Kuangzheng Zhu, Lucas Smith, Jerome Irianto, Charlotte R. Pfeifer, Cory M. Alvey, Jiazheng Ji, Dazhen Liu, Sangkyun Cho, Rachel R. Bennett, Andrea J. Liu, Roger A. Greenberg, and Dennis E. Discher. Nuclear rupture at sites of high curvature compromises retention of DNA repair factors. Journal of Cell Biology, 217(11):3796–??, November 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/11/3796.

Xiong:2017:OLP

[XJG^{+}17] Lei Xiong, Ji-Ung Jung, Hao-Han Guo, Jin-Xiu Pan, Xiang-Dong Sun, Lin Mei, and Wen-Cheng Xiong. Osteoblastic Lrp4 promotes osteoclastogenesis by regulating ATP release and adenosine–A_{2A} R signaling. Journal of Cell Biology, 216 (3):761–??, March 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/3/761.

Xu:2018:RPL

[XLW^{+}18] Dijin Xu, Yuqi Li, Lizhen Wu, Ying Li, Dongyu Zhao, Jinhai Yu, Tuozhi Huang, Charles Ferguson, Robert G. Parton, Hongyuan Yang, and Peng Li. Rab18 promotes lipid droplet (LD) growth by tethering the ER to LDs through SNARE and NRZ interactions. Journal of Cell Biology, 217(3):975–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/3/975.

Xing:2019:GSS

[XMJ^{+}19] Bowen Xing, Jian Ma, Zongzhe Jiang, Zijie Feng, Sunbin Ling, Katy Szigety, Wen Su, Longmei Zhang, Ruirui Jia, Yanmei Sun, Lin Zhang, Xiangchen Kong, Xiaosong Ma, and Xianxin Hua. GLP-1 signaling suppresses menin’s transcriptional block by phosphorylation in β cells. Journal of Cell Biology, 218 (3):855–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/3/855.

Xia:2019:RDD

[XPZ^{+}19] Yuntao Xia, Charlotte R. Pfeifer, Kuangzheng Zhu, Jerome Irianto, Dazhen Liu, Kalia Pannell, Emily J. Chen, Lawrence J.

Dooling, Michael P. Tobin, Mai Wang, Irena L. Ivanovska, Lucas R. Smith, Roger A. Greenberg, and Dennis E. Discher. Rescue of DNA damage after constricted migration reveals a mechano-regulated threshold for cell cycle. Journal of Cell Biology, 218(8):2545–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/8/2545.

Xiang:2018:CCL

[XRH^{+}18a] Wanqing Xiang, M. Julia Roberti, Jean-Karim Hériché, Sébastien Huet, Stephanie Alexander, and Jan Ellenberg. Correction: Correlative live and super-resolution imaging reveals the dynamic structure of replication domains. Journal of Cell Biology, 217(9):3315–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/9/3315. See [XRH^{+}18b].

Xiang:2018:CLS

[XRH^{+}18b] Wanqing Xiang, M. Julia Roberti, Jean-Karim Hériché, Sébastien Huet, Stephanie Alexander, and Jan Ellenberg. Correlative live and super-resolution imaging reveals the dynamic structure of replication domains. Journal of Cell Biology, 217 (6):1973–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress. org/content/217/6/1973. See correction [XRH^{+}18a].

Xue:2016:BBM

[XS16] Zenghui Xue and Anna Marie Sokac. Back-to-back mechanisms drive actomyosin ring closure during Drosophila embryo cleavage. Journal of Cell Biology, 215(3):335–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/215/3/335.

Xia:2018:CIP

[XSJ18] Wenlong Xia, Libo Su, and Jianwei Jiao. Cold-induced protein RBM3 orchestrates neurogenesis via modulating Yap mRNA stability in cold stress. Journal of Cell Biology, 217(10): 3464–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/10/3464.

Xiong:2015:SFS

[XTS^{+}15] Jianhua Xiong, Dilyana Todorova, Ning-Yuan Su, Jinchul Kim, Pei-Jen Lee, Zhouxin Shen, Steven P. Briggs, and Yang Xu. Stemness factor Sall4 is required for DNA damage response in embryonic stem cells. Journal of Cell Biology, 208(5):513–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
208/5/513.

Xu:2018:KMD

[XTT^{+}18] Fang Xu, Hironori Takahashi, Yosuke Tanaka, Sotaro Ichinose, Shinsuke Niwa, Matthew P. Wicklund, and Nobutaka Hirokawa. KIF1Bβ mutations detected in hereditary neuropathy impair IGF1R transport and axon growth. Journal of Cell Biology, 217(10):3480–??, October 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/10/3480.

Xie:2015:ENS

[XWZ^{+}15] Chuan-Ming Xie, Dongping Wei, Lili Zhao, Sylvie Marchetto, Lin Mei, Jean-Paul Borg, and Yi Sun. Erbin is a novel substrate of the Sag-βTrCP E3 ligase that regulates Kras^{G12D}-induced skin tumorigenesis. Journal of Cell Biology, 209(5):721–??, June 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
209/5/721.

Yang:2015:NMM

[YAHH15] Hui-Ju Yang, Haruhiko Asakawa, Tokuko Haraguchi, and Yasushi Hiraoka. Nup132 modulates meiotic spindle attachment in fission yeast by regulating kinetochore assembly. Journal of Cell Biology, 211(2):295–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/211/2/295.

Yue:2018:ACC

[YBZ^{+}18] Yang Yue, T. Lynne Blasius, Stephanie Zhang, Shashank Jariwala, Benjamin Walker, Barry J. Grant, Jared C. Cochran, and Kristen J. Verhey. Altered chemomechanical coupling causes impaired motility of the kinesin-4 motors KIF27 and KIF7. Journal of Cell Biology, 217(4):1319–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/4/1319.

Yasuda:2017:FID

[YCSJ^{+}17] Kyota Yasuda, Sarah F. Clatterbuck-Soper, Meredith E. Jackrel, James Shorter, and Stavroula Mili. FUS inclusions disrupt RNA localization by sequestering kinesin-1 and inhibiting microtubule detyrosination. Journal of Cell Biology, 216 (4):1015–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/4/1015.

Yap:2018:DDC

[YDM^{+}18] Chan Choo Yap, Laura Digilio, Lloyd P. McMahon, A. Denise R. Garcia, and Bettina Winckler. Degradation of dendritic cargos requires Rab7-dependent transport to somatic lysosomes. Journal of Cell Biology, 217(9):3141–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/9/3141.

Yellen:2018:FTM

[Yel18] Gary Yellen. Fueling thought: Management of glycolysis and oxidative phosphorylation in neuronal metabolism. Journal of Cell Biology, 217(7):2235–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/7/2235.

Yanakieva:2019:CTM

[YEM^{+}19] Iskra Yanakieva, Anna Erzberger, Marija Matejčić, Carl D. Modes, and Caren Norden. Cell and tissue morphology determine actin-dependent nuclear migration mechanisms in neuroepithelia. Journal of Cell Biology, 218(10):3272–??, October 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
218/10/3272.

Yang:2017:DRF

[YGMR^{+}17] Yang Yang, Yanzhe Gao, Liz Mutter-Rottmayer, Anastasia Zlatanou, Michael Durando, Weimin Ding, David Wyatt, Dale Ramsden, Yuki Tanoue, Satoshi Tateishi, and Cyrus Vaziri. DNA repair factor RAD18 and DNA polymerase Polκ confer tolerance of oncogenic DNA replication stress. Journal of Cell Biology, 216(10):3097–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/216/10/3097.

Yao:2017:VEP

[YGW^{+}17] Jiayi Yao, Pierre J. Guihard, Xiuju Wu, Ana M. Blazquez-Medela, Melissa J. Spencer, Medet Jumabay, Peter Tontonoz, Alan M. Fogelman, Kristina I. Boström, and Yucheng Yao. Vascular endothelium plays a key role in directing pulmonary epithelial cell differentiation. Journal of Cell Biology, 216(10): 3369–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/10/3369.

Yamada:2015:RCB

[YH15] Kenneth M. Yamada and Alan Hall. Reproducibility and cell biology. Journal of Cell Biology, 209(2):191–??, April 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/209/2/191.

Yin:2017:GPA

[YHG^{+}17] Jianhua Yin, Yaling Huang, Pengfei Guo, Siqi Hu, Sawako Yoshina, Nan Xuan, Qiwen Gan, Shohei Mitani, Chonglin Yang, and Xiaochen Wang. GOP-1 promotes apoptotic cell degradation by activating the small GTPase Rab2 in C. elegans. Journal of Cell Biology, 216(6):1775–??, June 2017. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/216/6/1775.

Yasunaga:2015:PPI

[YHS^{+}15] Takayuki Yasunaga, Sylvia Hoff, Christoph Schell, Martin Helmstädter, Oliver Kretz, Sebastian Kuechlin, Toma A. Yakulov, Christina Engel, Barbara Müller, Robert Bensch, Olaf Ronneberger, Tobias B. Huber, Soeren S. Lienkamp, and Gerd Walz. The polarity protein Inturned links NPHP4 to Daam1 to control the subapical actin network in multiciliated cells. Journal of Cell Biology, 211(5):963–??, December 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/211/5/963.

Yukawa:2015:MWP

[YIT15] Masashi Yukawa, Chiho Ikebe, and Takashi Toda. The Msd1– Wdr8–Pkl1 complex anchors microtubule minus ends to fission yeast spindle pole bodies. Journal of Cell Biology, 209(4):549– ??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
209/4/549.

Yu-Kemp:2017:CEE

[YKKB17] Hui-Chia Yu-Kemp, James P. Kemp, and William M. Brieher. CRMP-1 enhances EVL-mediated actin elongation to build lamellipodia and the actin cortex. Journal of Cell Biology, 216 (8):2463–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/8/2463.

Yin:2016:PPD

[YKO^{+}16] Xinghua Yin, Grahame J. Kidd, Nobuhiko Ohno, Guy A. Perkins, Mark H. Ellisman, Chinthasagar Bastian, Sylvain Brunet, Selva Baltan, and Bruce D. Trapp. Proteolipid protein– deficient myelin promotes axonal mitochondrial dysfunction via altered metabolic coupling. Journal of Cell Biology, 215(4): 531–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/215/4/531.

Yousefi:2016:MRB

[YLND^{+}16] Maryam Yousefi, Ning Li, Angela Nakauka-Ddamba, Shan Wang, Kimberly Davidow, Jenna Schoenberger, Zhengquan Yu, Shane T. Jensen, Michael G. Kharas, and Christopher J. Lengner. Msi RNA-binding proteins control reserve intestinal stem cell quiescence. Journal of Cell Biology, 215(3): 401–??, November 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/215/3/401.

Yan:2015:DEC

[YLW^{+}15] Kaowen Yan, Li Li, Xiaojian Wang, Ruisha Hong, Ying Zhang, Hua Yang, Ming Lin, Sha Zhang, Qihua He, Duo Zheng, Jun Tang, Yuxin Yin, and Genze Shao. The deubiquitinating enzyme complex BRISC is required for proper mitotic spindle assembly in mammalian cells. Journal of Cell Biology, 210 (2):209–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/2/209.

Ye:2018:BTR

[YNN18] Fan Ye, Andrew R. Nager, and Maxence V. Nachury. BB-Some trains remove activated GPCRs from cilia by enabling passage through the transition zone. Journal of Cell Biology, 217(5):1847–??, May 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/5/1847.

Yoshida:2015:GFS

[YPY^{+}15] Sei Yoshida, Regina Pacitto, Yao Yao, Ken Inoki, and Joel A. Swanson. Growth factor signaling to mTORC1 by amino acid– laden macropinosomes. Journal of Cell Biology, 211(1):159–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
211/1/159.

Yeyati:2017:KCA

[YSM^{+}17] Patricia L. Yeyati, Rachel Schiller, Girish Mali, Ioannis Kasioulis, Akane Kawamura, Ian R. Adams, Christopher Playfoot, Nick Gilbert, Veronica van Heyningen, Jimi Wills, Alex von Kriegsheim, Andrew Finch, Juro Sakai, Christopher J. Schofield, Ian J. Jackson, and Pleasantine Mill. KDM3A coordinates actin dynamics with intraflagellar transport to regulate cilia stability. Journal of Cell Biology, 216(4):999–??, April 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/4/
999.

Yurchenko:2018:SRT

[YSR^{+}18] Maria Yurchenko, Astrid Skjesol, Liv Ryan, Gabriel Mary Richard, Richard Kumaran Kandasamy, Ninghai Wang, Cox Terhorst, Harald Husebye, and Terje Espevik. SLAMF1 is required for TLR4-mediated TRAM–TRIF-dependent signaling in human macrophages. Journal of Cell Biology, 217(4): 1411–??, April 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/4/1411.

Yan:2015:SYD

[YSW^{+}15] Liming Yan, Sha Sun, Wei Wang, Juanming Shi, Xiaoyu Hu, Shiyan Wang, Dan Su, Zihe Rao, Junjie Hu, and Zhiyong Lou. Structures of the yeast dynamin-like GTPase Sey1p provide insight into homotypic ER fusion. Journal of Cell Biology, 210(6): 961–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/6/961.

Yoshida:2016:SDT

[YTGA16] Makoto M. Yoshida, Lily Ting, Steven P. Gygi, and Yoshiaki Azuma. SUMOylation of DNA topoisomerase IIα regulates histone H3 kinase Haspin and H3 phosphorylation in mitosis. Journal of Cell Biology, 213(6):665–??, June 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/213/6/665.

Yang:2015:BPC

[YTL15] Yang Yang, Dai Tsuchiya, and Soni Lacefield. Bub3 promotes Cdc20-dependent activation of the APC/C in S. cerevisiae. Journal of Cell Biology, 209(4):519–??, May 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/209/4/519.

Yamada:2017:MKF

[YTTH^{+}17] Moé Yamada, Yohko Tanaka-Takiguchi, Masahito Hayashi, Momoko Nishina, and Gohta Goshima. Multiple kinesin-14 family members drive microtubule minus end–directed transport in plant cells. Journal of Cell Biology, 216(6):1705–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/6/1705.

Yudushkin:2019:TPN

[Yud19] Ivan Yudushkin. TAPPing into PIPs: a new reporter reveals the origin of plasma membrane PI(3,4)P_{2}. Journal of Cell Biology, 218(3):735–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/3/735.

Yoon:2018:EET

[YVlMS18] Ki-Jun Yoon, Caroline Vissers, Guo li Ming, and Hongjun Song. Epigenetics and epitranscriptomics in temporal patterning of cortical neural progenitor competence. Journal of Cell Biology, 217(6):1901–??, June 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/6/1901.

Ye:2018:NPE

[YVM18] Anna A. Ye, Vikash Verma, and Thomas J. Maresca. NOD is a plus end–directed motor that binds EB1 via a new microtubule tip localization sequence. Journal of Cell Biology, 217(9):3007–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/9/3007.

Yang:2017:EER

[YWdH^{+}17] Chao Yang, Jingchao Wu, Cecilia de Heus, Ilya Grigoriev, Nalan Liv, Yao Yao, Ihor Smal, Erik Meijering, Judith Klumperman, Robert Z. Qi, and Anna Akhmanova. EB1 and EB3 regulate microtubule minus end organization and Golgi morphology. Journal of Cell Biology, 216(10):3179–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/10/3179.

Yau:2017:SRO

[YWW17] Richard G. Yau, Sara Wong, and Lois S. Weisman. Spatial regulation of organelle release from myosin V transport by p21activated kinases. Journal of Cell Biology, 216(6):1557–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/216/6/
1557.

Yaguchi:2018:UCC

[YYM^{+}18] Kan Yaguchi, Takahiro Yamamoto, Ryo Matsui, Yuki Tsukada, Atsuko Shibanuma, Keiko Kamimura, Toshiaki Koda, and Ryota Uehara. Uncoordinated centrosome cycle underlies the instability of non-diploid somatic cells in mammals. Journal of Cell Biology, 217(7):2463–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/217/7/2463.

Yakymovych:2015:CMT

[YYZ^{+}15] Ihor Yakymovych, Mariya Yakymovych, Guangxiang Zang, Yabing Mu, Anders Bergh, Maréne Landström, and Carl-Henrik Heldin. CIN85 modulates TGFβ signaling by promoting the presentation of TGFβ receptors on the cell surface. Journal of Cell Biology, 210(2):319–??, July 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/210/2/319.

Zilberman:2017:CRJ

[ZAAN17] Yuliya Zilberman, Joshua Abrams, Dorian C. Anderson, and Jeremy Nance. Cdc42 regulates junctional actin but not cell polarization in the Caenorhabditis elegans epidermis. Journal of Cell Biology, 216(11):3729–??, November 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/11/3729.

Zhu:2017:MCL

[ZAT^{+}17] Yili Zhu, Xiaojing An, Alexis Tomaszewski, Peter K. Hepler, and Wei-Lih Lee. Microtubule cross-linking activity of She1 ensures spindle stability for spindle positioning. Journal of Cell Biology, 216(9):2759–??, September 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/216/9/2759.

Zhang:2019:RAG

[ZAT^{+}19] Xiao-Feng Zhang, Visar Ajeti, Nicole Tsai, Arash Fereydooni, William Burns, Michael Murrell, Enrique M. De La Cruz, and Paul Forscher. Regulation of axon growth by myosin II– dependent mechanocatalysis of cofilin activity. Journal of Cell Biology, 218(7):2329–??, July 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/7/2329.

Zaidel-Bar:2018:CCP

[ZB18] Ronen Zaidel-Bar. Cell cycle pacemaker keeps adhesion in step with division. Journal of Cell Biology, 217(9):2981–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/217/9/
2981.

Zhang:2019:BBR

[ZB19] Ting Zhang and Siddharth Balachandran. Bayonets over bombs: RIPK3 and MLKL restrict Listeria without triggering necroptosis. Journal of Cell Biology, 218(6):1773–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/6/
1773.

Zulkipli:2018:SRH

[ZCH^{+}18] Ihsan Zulkipli, Joanna Clark, Madeleine Hart, Roshan L. Shrestha, Parveen Gul, David Dang, Tami Kasichiwin, Izabela Kujawiak, Nishanth Sastry, and Viji M. Draviam. Spindle rotation in human cells is reliant on a MARK2-mediated equatorial spindle-centering mechanism. Journal of Cell Biology, 217(9):

3057–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/217/9/3057.

Zou:2015:MPC

[ZCL^{+}15] Zhipeng Zou, Juan Chen, Anling Liu, Xuan Zhou, Qiancheng Song, Chunhong Jia, Zhenguo Chen, Jun Lin, Cuilan Yang, Ming Li, Yu Jiang, and Xiaochun Bai. mTORC2 promotes cell survival through c–Myc-dependent up-regulation of E2F1. Journal of Cell Biology, 211(1):105–??, October 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/211/1/105.

Zellweger:2015:RMR

[ZDM^{+}15] Ralph Zellweger, Damian Dalcher, Karun Mutreja, Matteo Berti, Jonas A. Schmid, Raquel Herrador, Alessandro Vindigni, and Massimo Lopes. Rad51-mediated replication fork reversal is a global response to genotoxic treatments in human cells. Journal of Cell Biology, 208(5):563–??, March 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/208/5/563.

Zobel:2018:NEA

[ZDSM^{+}18] Martina Zobel, Andrea Disanza, Francesca Senic-Matuglia, Michel Franco, Ivan Nicola Colaluca, Stefano Confalonieri, Sara Bisi, Elisa Barbieri, Giusi Caldieri, Sara Sigismund, Salvatore Pece, Philippe Chavrier, Pier Paolo Di Fiore, and Giorgio Scita. A NUMB–EFA6B–ARF6 recycling route controls apically restricted cell protrusions and mesenchymal motility. Journal of Cell Biology, 217(9):3161–??, September 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/217/9/3161.

Zhu:2016:MPC

[ZGDS^{+}16] Lili Zhu, Aurora Gomez-Duran, Gabriele Saretzki, Shibo Jin, Katarzyna Tilgner, Dario Melguizo-Sanchis, Georgios Anyfantis, Jumana Al-Aama, Ludovic Vallier, Patrick Chinnery, Majlinda Lako, and Lyle Armstrong. The mitochondrial protein CHCHD2 primes the differentiation potential of human induced pluripotent stem cells to neuroectodermal lineages. Journal of Cell Biology, 215(2):187–??, October 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/215/2/187.

Zhuo:2015:URK

[ZGZ^{+}15] Xiaolong Zhuo, Xiao Guo, Xiaoyan Zhang, Guihua Jing, Yao Wang, Qiang Chen, Qing Jiang, Junjun Liu, and Chuanmao Zhang. Usp16 regulates kinetochore localization of Plk1 to promote proper chromosome alignment in mitosis. Journal of Cell Biology, 210(5):727–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/210/5/727.

Zhang:2019:KSL

[Zha19] Hongquan Zhang. Kindlin-3 stokes the life span of podosomes. Journal of Cell Biology, 218(10):3166–??, October 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/10/3166.

Zhao:2019:PCA

[ZHP^{+}19] Lei Zhao, Yuqing Hou, Tyler Picariello, Branch Craige, and George B. Witman. Proteome of the central apparatus of a ciliary axoneme. Journal of Cell Biology, 218(6):2051–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/218/6/
2051.

Zhu:2017:WKP

[Zhu17] Michael X. Zhu. A well-known potassium channel plays a critical role in lysosomes. Journal of Cell Biology, 216(6):1513–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/6/1513.

Zhang:2017:BCD

[ZJM^{+}17] Xingmin Zhang, Shan Jiang, Kelly A. Mitok, Lingjun Li, Alan D. Attie, and Thomas F. J. Martin. BAIAP3, a C2 domain–containing Munc13 protein, controls the fate of densecore vesicles in neuroendocrine cells. Journal of Cell Biology, 216(7):2151–??, July 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/7/2151.

Zhao:2015:MNS

[ZLG^{+}15] Xiao-Di Zhao, Yuan-Yuan Lu, Hao Guo, Hua-Hong Xie, Li-Jie He, Gao-Fei Shen, Jin-Feng Zhou, Ting Li, Si-Jun Hu, Lin Zhou, Ya-Nan Han, Shu-Li Liang, Xin Wang, Kai-Chun Wu, Yong-Quan Shi, Yong-Zhan Nie, and Dai-Ming Fan. MicroRNA-7/ NF-κB signaling regulatory feedback circuit regulates gastric carcinogenesis. Journal of Cell Biology, 210(4): 613–??, August 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/4/613.

zurLage:2018:CDM

[zLSSS^{+}18] Petra zur Lage, Panagiota Stefanopoulou, Katarzyna Styczynska-Soczka, Niall Quinn, Girish Mali, Alex von Kriegsheim, Pleasantine Mill, and Andrew P. Jarman. Ciliary dynein motor preassembly is regulated by Wdr92 in association with HSP90 co-chaperone, R2TP. Journal of Cell Biology, 217(7):2583–??, July 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
217/7/2583.

Zhao:2016:MMS

[ZLZD16] Wei Zhao, Jie Liu, Xiaoming Zhang, and Lih-Wen Deng. MLL5 maintains spindle bipolarity by preventing aberrant cytosolic aggregation of PLK1. Journal of Cell Biology, 212(7):829–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
212/7/829.

Zhang:2018:BBL

[ZNR^{+}18] Wanlu Zhang, Annett Neuner, Diana Rüthnick, Timo Sachsenheimer, Christian Lüchtenborg, Britta Brügger, and Elmar Schiebel. Brr6 and Brl1 locate to nuclear pore complex assembly sites to promote their biogenesis. Journal of Cell Biology, 217(3):877–??, March 2018. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/217/3/877.

Zhang:2015:FDR

[ZPT^{+}15] Chenying Zhang, Bhaskar Ponugoti, Chen Tian, Fanxing Xu, Rohinton Tarapore, Angelika Batres, Sarah Alsadun, Jason Lim, Guangyu Dong, and Dana T. Graves. FOXO1 differentially regulates both normal and diabetic wound healing. Journal of Cell Biology, 209(2):289–??, April 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/209/2/289.

Zhang:2019:TLD

[ZQZ19] Yujie Zhang, Linxiang Qi, and Hong Zhang. TGFβ-like DAF-7 acts as a systemic signal for autophagy regulation in C. elegans. Journal of Cell Biology, 218(12):3998–4006, December 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/218/12/3998/132544/TGF-like-DAF-7-acts-as-a-
systemic-signal-for.

Zaoui:2019:ARR

[ZRDP19] Kossay Zaoui, Charles V. Rajadurai, Stéphanie Duhamel, and Morag Park. Arf6 regulates RhoB subcellular localization to control cancer cell invasion. Journal of Cell Biology, 218(11): 3812–??, November 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/11/3812.

Zatulovskiy:2015:MS

[ZS15] Evgeny Zatulovskiy and Jan M. Skotheim. Mitosis is swell. Journal of Cell Biology, 211(4):733–??, November 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/211/4/733.

Zhou:2015:PFS

[ZSdO^{+}15] Xiaolai Zhou, Lirong Sun, Francisco Bastos de Oliveira, Xiaoyang Qi, William J. Brown, Marcus B. Smolka, Ying Sun, and Fenghua Hu. Prosaposin facilitates sortilin-independent lysosomal trafficking of progranulin. Journal of Cell Biology, 210(6): 991–??, September 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/210/6/991.

Zaganjor:2017:SNP

[ZSH17] Elma Zaganjor, Jessica B. Spinelli, and Marcia C. Haigis. Strength in numbers: Phosphofructokinase polymerization prevails in the liver. Journal of Cell Biology, 216(8):2239–??, August 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
216/8/2239.

Zou:2015:ADI

[ZT15] Wei Zou and Steven L. Teitelbaum. Absence of Dap12 and the α vβ3 integrin causes severe osteopetrosis. Journal of Cell Biology, 208(1):125–??, January 2015. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/208/1/125.

Zaritsky:2017:DRG

[ZTR^{+}17] Assaf Zaritsky, Yun-Yu Tseng, M. Angeles Rabadán, Shefali Krishna, Michael Overholtzer, Gaudenz Danuser, and Alan Hall. Diverse roles of guanine nucleotide exchange factors in regulating collective cell migration. Journal of Cell Biology, 216(6):1543–??, June 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/216/6/1543.

Zhang:2019:MDK

[ZWB^{+}19] Junya Zhang, Shan Wu, Susan K. Boehlein, Donald R. Mc-Carty, Gaoyuan Song, Justin W. Walley, Alan Myers, and A. Mark Settles. Maize defective kernel5 is a bacterial TamB homologue required for chloroplast envelope biogenesis. Journal of Cell Biology, 218(8):2638–??, August 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/8/2638.

Zhou:2016:VBF

[ZWS^{+}16] Mo Zhou, Heidi Wiener, Wenjuan Su, Yong Zhou, Caroline Liot, Ian Ahearn, John F. Hancock, and Mark R. Philips. VPS35 binds farnesylated N–Ras in the cytosol to regulate N– Ras trafficking. Journal of Cell Biology, 214(4):445–??, August 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/214/4/
445.

Zhou:2019:LCS

[ZWW^{+}19] Junxiang Zhou, Xin Wang, Min Wang, Yuwei Chang, Fengxia Zhang, Zhaonan Ban, Ruofeng Tang, Qiwen Gan, Shaohuan Wu, Ye Guo, Qian Zhang, Fengyang Wang, Liyuan Zhao, Yudong Jing, Wenfeng Qian, Guodong Wang, Weixiang Guo, and Chonglin Yang. The lysine catabolite saccharopine impairs development by disrupting mitochondrial homeostasis. Journal of Cell Biology, 218(2):580–??, February 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/218/2/580.

Zhou:2019:RDA

[ZWZ^{+}19] Fan Zhou, Zulin Wu, Mengzhu Zhao, Rakhilya Murtazina, Juan Cai, Ao Zhang, Rui Li, Dan Sun, Wenjing Li, Lei Zhao, Qunli Li, Jing Zhu, Xiaoxia Cong, Yiting Zhou, Zhiping Xie, Valeriya Gyurkovska, Liuju Li, Xiaoshuai Huang, Yanhong Xue, Liangyi Chen, Hui Xu, Haiqian Xu, Yongheng Liang, and Nava Segev. Rab5-dependent autophagosome closure by ESCRT. Journal of Cell Biology, 218(6):1908–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/218/6/1908.

Zhang:2016:DMA

[ZY16] Ting Zhang and Yihong Ye. Doa1 is a MAD adaptor for Cdc48. Journal of Cell Biology, 213(1):7–??, April 2016. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/213/1/7.

Zhang:2017:CAM

[ZYA^{+}17] Qing-Hua Zhang, Wai Shan Yuen, Deepak Adhikari, Jennifer A. Flegg, Greg FitzHarris, Marco Conti, Piotr Sicinski, Ibtissem Nabti, Petros Marangos, and John Carroll. Cyclin A2 modulates kinetochore–microtubule attachment in meiosis II. Journal of Cell Biology, 216(10):3133–??, October 2017. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/216/10/3133.

Zhou:2016:FAR

[ZYL^{+}16] Bing Zhou, Panpan Yu, Mei-Yao Lin, Tao Sun, Yanmin Chen, and Zu-Hang Sheng. Facilitation of axon regeneration by enhancing mitochondrial transport and rescuing energy deficits. Journal of Cell Biology, 214(1):103–??, July 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/214/1/103.

Zhao:2016:UCU

[ZZ16] Yan G. Zhao and Hong Zhang. ULK1 cycling: The ups and downs of the autophagy response. Journal of Cell Biology, 215 (6):757–??, December 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/215/6/757.

Zhao:2019:AME

[ZZ19] Yan G. Zhao and Hong Zhang. Autophagosome maturation: an epic journey from the ER to lysosomes. Journal of Cell Biology, 218(3):757–??, March 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/218/3/757.

Zlotek-Zlotkiewicz:2015:OVM

[ZZMC^{+}15] Ewa Zlotek-Zlotkiewicz, Sylvain Monnier, Giovanni Cappello, Mael Le Berre, and Matthieu Piel. Optical volume and mass measurements show that mammalian cells swell during mitosis. Journal of Cell Biology, 211(4):765–??, November 2015. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/211/4/765.

Zheng:2019:SAA

[ZZW^{+}19] Yanrong Zheng, Xiangnan Zhang, Xiaoli Wu, Lei Jiang, Anil Ahsan, Shijia Ma, Ziyu Xiao, Feng Han, Zheng-Hong Qin, Weiwei Hu, and Zhong Chen. Somatic autophagy of axonal mitochondria in ischemic neurons. Journal of Cell Biology, 218 (6):1891–??, June 2019. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/218/6/1891.